A Complete Bibliography of Publications in The Journal of Scientific Computing

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

09 May 2024 Version 2.06

Title word cross-reference

(ν, μ, s, 1 − σ) [2995]. −1 [1754]. 0 [1255]. 1 [812, 955, 1192, 1359, 1558, 1641, 1795, 1858, 2652, 2707, 3276]. 1_{σ} [3800]. 2 [177, 316, 384, 486, 593, 600, 658, 742, 769, 823, 865, 1174, 1260, 1276, 1306, 1310, 1658, 1748, 1779, 2002, 2252, 2581, 2583, 2597, 2644]. 3 [248, 263, 336, 384, 403, 530, 600, 646, 869, 1002, 1006, 1021, 1061, 1078, 1236, 1276, 1310, 1339, 1579, 1737, 1779, 1780, 1838, 1888, 2107, 2529, 2581, 2746, 2755, 3367, 3434, 3479]. 4 [2002]. 5 [1629]. 7 [982]. _{p} [1628]. _{σ} [3481]. A [276, 480, 629, 2767]. A^{−1} [629]. α [2740, 3118, 3567, 3650, 3768]. A\underline{x} = \underline{b} [276]. Ax = b [228]. B [1612, 1776]. β [3022]. C [1612, 1776]. C^{0} [440, 1161, 1273, 1639, 1729, 1974, 2013, 3059, 3256]. C^{1} [3722]. C_{∞} [578]. D [737, 1400, 2024, 2630]. Δ · B [582]. Div(Y) = f [571]. ℓ^{2} [2316]. ℓ^{q} [2316]. ℓ_{0} [2353, 3546]. ℓ_{1} [1190, 1318, 1583, 2039, 2713]. ℓ_{2} [1190, 1628]. ℓ_{p} [2963]. ϵ [104, 247]. EQ_{1}^{rot} [1195, 1739]. f [948]. F_{j}(x) [278, 283]. f ∈ L^{2} [571]. fp [790]. G^{1} [1373]. H [473, 1194, 1213, 1491, 1733, 1951, 1999, 2064, 2253, 2535]. H(curl) [3676].

H(div) [3347]. h/p [2205]. H^{−1} [2928, 3553]. H^{1} [456, 1573, 2267, 3481, 3575]. H^{1/2} [631]. H^{2} [742, 1355, 2882]. H^{∞} [116]. H_{N}^{T} [2980]. hp [301, 307, 344, 447, 500, 547, 608, 757, 894, 896, 971, 1155, 1470, 1988, 3056, 3079, 3156]. j = −1/2 [278]. j = −3/2 [283]. j = 1/2 [278]. K [1147, 2223, 2348, 2724, 3283]. L [3622]. L1 [3797]. L2 [3800]. L2 − 1_{σ} [3118]. L^{1} [716, 1782, 2292]. L^{1−2} [3778]. L^{2} [456, 812, 1915, 2288, 3003]. L^{∞} [571]. L^{∞}(H^{1}) [3118]. L^{p} [1683, 2595]. l_{0} [1121]. L_{1} [1426, 1669, 1685, 1986, 1990, 2077, 2094, 2494, 2902, 3271, 3516]. L_{1/2} [1595]. L_{1}/L_{2} [3786]. L_{2} [31, 736, 1426, 1986, 2072, 3271, 3516]. L_{∞} [736]. L_{k} [3293]. M [110, 2633, 3752]. M_{1} [625]. M_{2} [1793]. R^{2} [2640]. H^{1} [2890]. L^{2}(Γ) [3120]. H [1476, 2302]. L_{∈} [998]. M [1653, 1970, 2171]. O(N log N) [1218]. N [199, 621, 1611, 2557, 3348]. n^{+}-n-n^{+} [1284]. O [2348]. o(1/k) [1818]. O(2) [1048]. p [295, 547, 640, 673, 808, 1049, 1491, 1893, 1895, 1999, 2029, 2147, 2253, 2273, 2304, 2420, 3045, 3079, 3343, 3789, 3799]. p(x) [1490]. P_{1} [1064, 1165, 1371, 1659, 2835]. P_{N} [976]. P_{N}P_{M} [1107, 1329, 2751]. Q [3358]. Q_{1} [2835, 3765]. Q_{4} [181]. qd [2080]. R^{2} [978]. R^{n} [998]. S [3752]. S_{N} [3265]. SO(3) [2024]. t [3244]. τ [1202]. θ [576]. V [1997]. H⃗ [2749]. W^{−1,p′} [3789]. x^{(k)} = b^{(k)} [629]. Z [2497]. Z_{2} [933].

-Adaptive [473, 1194, 1733, 1893, 1951, 2147, 2535, 3079, 3343]. -Adaptivity [307]. -Analysis [344]. -Anisotropic [2273]. -Block [2107]. -body [2557]. -Conforming [2749, 3676, 3722]. -Consistency [1612, 1776]. -Continuous [1974]. -Control [3778]. -Coupled [3348]. -CPR [1329]. -cycle [1997]. -D [336, 600, 1006, 1061, 1078, 1339, 2644, 2746, 3276, 3367, 3434]. -D/ [600]. -dimensional [1611]. -Dimensions [621]. -Discontinuous [301, 896, 3156]. -Dissipative [3768]. -elements [3347]. -Estimation [1202]. -Expansion [104, 247]. -FEM [971, 2064]. -Finite [757, 1412, 1470]. -Fitting [716]. -Laplacian [295, 640, 1049, 1490, 2304, 2420, 3045]. -local [894]. -Matrices [2302]. -Matrix [1476, 2024, 2633]. -Means [2223, 3283]. -Multigrid [790, 3343]. -Nonconforming [1064, 1161]. -Norm [1573, 1583, 1685, 2595, 2963]. -Optimality [812]. -Order [673]. -Partition [2724]. -Rectangular [2630]. -Refinement [1213, 1895]. -Regularized [1318]. -Robust [2740, 3022, 3118, 3567, 3650]. -Scheme [576]. -Shock [3079]. -Spectral [2029]. -Stability [742, 1355, 1612, 1776, 3481]. -Stable [480, 571]. -Stage [982]. -tensor [1970, 3358]. -Tensors [1653, 2171]. -Type [2080, 2497, 3622]. -URV [3244]. -variable [110]. -Version [500, 608, 808, 1155]. -Versions [2253]. -Wave [948, 1629]. -Weak [1273].

/3 [1400].

0rder [2391].

1 [439, 520, 3481]. 1992-2002 [375]. 1D [3039].

2 [1400]. 2D [2106, 2134, 2319, 3041, 3105, 3159, 3247, 3491, 3511, 3740]. 2D/3D [3041]. 2nd [3276]. 2nd-MGFM [3276].

3D [1557, 2230, 2389, 2977, 2995, 3041, 3082, 3148, 3255, 3774].

4-Point [3172]. 40/40 [2348]. 40Sedimentation [2692]. 4D [3134]. 4th [1937, 3426].

5 [182].

6000 [152]. 60th [2231].

90 [182].

A-Posteriori [778]. AAA [3091]. Abarbanel [2488, 2489]. Absolute [1302]. Absorbing [339, 1446, 2066]. Abstract [3269]. Accelerated [286, 1124, 1190, 1369, 1476, 1704, 1838, 2348, 2431, 2923, 2972, 2989, 3009, 3314, 3456, 3472, 3546, 3581, 3632, 3647]. Accelerating [200, 1199, 1538, 2969]. Acceleration [259, 2357, 2741, 2746, 2861, 2985, 3523, 3547, 3599, 3733, 3766]. Acceptance [1025]. Accumulation [2877]. Accuracy [24, 69, 174, 429, 502, 508, 565, 718, 735, 831, 882, 883, 887, 1015, 1017, 1043, 1046, 1060, 1071, 1079, 1103, 1184, 1254, 1332, 1368, 1453, 1456, 1509, 1562, 1581, 1622, 1688, 1753, 1817, 1940, 2080, 2090, 2201, 2288, 2491, 2574, 2604, 2640, 2730, 2997, 3275, 3291, 3428, 3473, 3541, 3590, 3617]. Accuracy-Conserving [735, 1079, 1254, 1453, 1581, 1940, 2491]. Accuracy-Enhancement [3275]. Accuracy-Preserving [3541]. Accurate [209, 238, 419, 487, 519, 544, 598, 662, 681, 744, 817, 890, 1027, 1044, 1082, 1316, 1457, 1492, 1503, 1719, 1750, 1778, 1799, 1801, 1822, 1841, 1843, 1890, 1965, 2173, 2314, 2492, 2504, 2514, 2515, 2532, 2543, 2551, 2601, 2633, 2652, 2742, 2833, 2866, 2884, 2949, 3207, 3298, 3366, 3485, 3630, 3685, 3727, 3772, 3802]. Acoustic [1321, 1784, 1965, 2004, 2054, 2140, 2254, 2319, 2628, 3594, 3660]. Acoustics [508]. Across [1965]. Acting [2952]. Actions [3694]. Active [1349, 1897, 2432, 2705, 2713, 2732, 3340, 3403, 3610, 3736, 3770]. Active-Set [2713, 3610]. Actively [284, 644]. AD [1239]. Adams [738]. Adaptation [610, 1348, 1415, 1650, 2194, 3305, 3459]. Adapted [638, 3187]. Adapting [3506]. Adaptive [190, 206, 221, 282, 329, 411, 464, 473, 505, 510, 530, 536, 564, 609, 613, 615, 675, 763, 792, 793, 837, 913, 976, 1014, 1031, 1088, 1141, 1151, 1159, 1165, 1216, 1230, 1267, 1293, 1317, 1413, 1424, 1481, 1520, 1580, 1599, 1635, 1659, 1730, 1736, 1766, 1876, 1881, 1883, 1890, 1893, 1915, 1933, 1958, 1998, 2011, 2025, 2034, 2044, 2050, 2071, 2082, 2117, 2147, 2236, 2239, 2276, 2296, 2441, 2512, 2531, 2557, 2655, 2657, 2662, 2663, 2677, 2745, 2763, 2764, 2798, 2863, 2894, 2955, 2972, 3023, 3079, 3126, 3166, 3173, 3321, 3330, 3370, 3397, 3403, 3452, 3523, 3538, 3577, 3644, 3657, 3670, 3691, 3719]. Adaptive [105, 134, 277, 309, 407, 470, 823, 827, 1136, 1146, 1186, 1194, 1258, 1357, 1384, 1543, 1705, 1733, 1891, 1951, 2174, 2200, 2300, 2327, 2346, 2404, 2422, 2469, 2500, 2535, 2559, 2586, 2718, 2780, 2837, 2959, 2970, 3047, 3065, 3078, 3175, 3183, 3193, 3259, 3262, 3343, 3386, 3467, 3580, 3641, 3675, 3708]. Adaptivity [307, 933, 1048, 1050, 1308, 2256, 2340, 2990, 3317]. ADCIRC [1062]. Additive [139, 208, 749, 916, 1057, 1187, 1542, 1614, 1760, 1977, 1994, 2426, 2754, 2972]. ADER [316, 348, 471, 946, 1525, 2432, 2796, 2823, 3030, 3467]. ADER-DG [3030]. ADG [1743]. adhesion [150]. ADI [21, 212, 213, 1260, 1387, 1512, 1726, 1863, 2775, 3313]. Adini [1734]. Adjoint [635, 1924, 2104, 2703, 2752, 2921, 3408]. Adjoint-Based [1924, 2104]. Adjustable [3635]. Admissibility [3785]. Admissible [2687, 3809]. ADMM [1564, 1684, 1807, 1818, 1985, 2107, 2117, 2258, 2861, 3409, 3656]. Advanced [146]. Advancement [426]. Advances [122, 540, 1076]. Advantage [957]. Advection [52, 140, 190, 196, 255, 296, 324, 495, 561, 591, 627, 705, 740, 830, 875, 893, 946, 1040, 1147, 1210, 1282, 1283, 1425, 1603, 1604, 1747, 1777, 1858, 1866, 1899, 1942, 2026, 2081, 2108, 2163, 2207, 2232, 2282, 2556, 2759, 2795, 2828, 2983, 3075, 3273, 3295, 3354, 3385, 3463, 3519, 3577, 3736]. advection-diffusion [52, 190]. Advection-Dominant [2795]. Advection-Robust [2232]. Advective [810]. aeroautoelastics [91]. Aerodynamic [42]. Aerodynamics [318, 490]. Aeronautical [574]. Affine [210, 1649, 2365]. Ag [886]. agent [3568]. Aggregation [3738, 3809]. Aggregation-Based [3738]. Aided [3568]. Airfoil [181, 722]. ALE [1403, 2876, 3024, 3729]. Algebraic [112, 181, 532, 1246, 1419, 1488, 2306, 2587, 2633, 2654, 3100, 3238, 3613, 3810]. algebraic- [181]. algebraic-Q4 [112]. Algorithm [201, 241, 283, 313, 337, 374, 553, 684, 722, 785, 807, 853, 884, 892, 1060, 1092, 1121, 1178, 1184, 1199, 1223, 1257, 1277, 1318, 1338, 1341, 1422, 1430, 1454, 1478, 1538, 1580, 1583, 1595, 1598, 1649, 1727, 1731, 1742, 1807, 1872, 1913, 1923, 1964, 1989, 2017, 2027, 2044, 2053, 2084, 2090, 2136, 2170, 2171, 2182, 2223, 2225, 2310, 2333, 2351, 2388, 2391, 2427, 2479, 2548, 2578, 2638, 2690, 2713, 2723, 2779, 2797, 2813, 2825, 2835, 2857, 2858, 2867, 2871, 2878, 2895, 2941, 2976, 3066, 3073, 3116, 3150, 3194, 3244, 3257, 3291, 3332, 3372, 3420, 3500, 3521, 3528, 3553, 3560, 3561, 3574, 3581, 3656, 3707, 3710, 3749, 3753, 3768, 3795, 3808]. Algorithm [20, 52, 114, 119, 136, 148, 150, 1482, 2079, 2091, 3646, 3700]. Algorithmic [1707, 2187, 2302]. Algorithms [46, 97, 126, 190, 320, 359, 460, 512, 638, 640, 699, 826, 914, 1126, 1171, 1200, 1375, 1489, 1496, 1516, 1562, 1577, 1642, 1781, 1820, 1842, 1864, 1868, 1873, 1997, 2116, 2156, 2270, 2281, 2389, 2409, 2618, 2649, 2658, 2734, 2780, 2807, 2856, 2912, 2937, 2943, 3189, 3210, 3366, 3390, 3475, 3488, 3497, 3546, 3568, 3585, 3606, 3655, 3745]. Aliasing [195, 2081, 3140]. Aligned [2021]. Alignment [2776]. Alikhanov [2702, 3118]. All-at-Once [2582, 2840]. All-Speed [3476]. Allen [464, 1892, 2347, 2406, 2426, 2439, 2699, 2743, 2754, 2770, 2802, 2986, 3029, 3054, 3088, 3118, 3124, 3170, 3220, 3248, 3382, 3493, 3552, 3562, 3614, 3800]. Allowing [2855, 3661]. Alloys [1396]. ALM [1564]. Along [398, 1090, 1156, 2779]. ALORA [2365]. Also [666]. Alternate [1421]. Alternating [1022, 1307, 1387, 1510, 1549, 1551, 1568, 1648, 1702, 1875, 2124, 2322, 2333, 2348, 3038, 3368, 3673]. Alternative [138, 1539, 2295, 2307, 2839, 2978, 3364]. Aluminum [840]. Ambiguity [3361]. American [654, 852, 1242, 2096, 2545, 3649]. AMG [3647]. Among [898]. Ampère [929, 1114, 1716, 2175, 2181, 2328, 2329, 2405, 3479, 3629]. Amplitude [3289]. AMR [357, 2335]. Analogy [147]. Analyses [1649, 1728, 3221, 3652]. Analysis [234, 242, 303, 322, 423, 520, 636, 649, 660, 773, 856, 858, 871, 880, 929, 966, 974, 1026, 1074, 1075, 1078, 1107, 1113, 1152, 1251, 1286, 1351, 1401, 1407, 1411, 1421, 1476, 1514, 1521, 1576, 1626, 1631, 1643, 1657, 1664, 1698, 1772, 1802, 1875, 1892, 1928, 2036, 2048, 2178, 2187, 2245, 2247, 2248, 2290, 2313, 2436, 2442, 2443, 2462, 2470, 2494, 2540, 2545, 2574, 2602, 2624, 2685, 2698, 2699, 2809, 2868, 2891, 2921, 2926, 3001, 3007, 3010, 3022, 3039, 3098, 3110, 3123, 3135, 3188, 3205, 3209, 3312, 3494, 3527, 3537, 3603, 3650, 3663, 3667, 3683, 3702, 3709, 3712, 3721, 3742, 3766]. Analysis [247, 261, 266, 306, 309, 344, 434, 455, 467, 500, 503, 568, 593, 608, 637, 852, 891, 915, 1091, 1278, 1303, 1334, 1417, 1422, 1428, 1511, 1540, 1548, 1591, 1598, 1601, 1606, 1630, 1684, 1699, 1706, 1717, 1739, 1743, 1749, 1759, 1879, 1887, 1934, 1945, 1975, 1980, 1994, 2026, 2051, 2075, 2097, 2114, 2160, 2303, 2308, 2326, 2344, 2348, 2401, 2416, 2418, 2420, 2471, 2476, 2477, 2527, 2538, 2563, 2608, 2614, 2625, 2664, 2694, 2700, 2711, 2883, 2884, 2897, 2954, 2960, 2963, 3025, 3042, 3070, 3092, 3117, 3124, 3129, 3143, 3272, 3288, 3335, 3362, 3378, 3394, 3398, 3401, 3404, 3439, 3474, 3482, 3484, 3495, 3520, 3556, 3581, 3630, 3656, 3751]. Analysis [2, 12, 15, 27, 28, 38, 113, 138, 147, 154, 161, 176, 971, 1067, 1070, 1260, 1372, 1489, 1532, 1596, 1681, 1810, 1971, 2079, 2188, 2218, 2419, 2481, 2647, 2665, 2738, 2766, 2781, 2790, 2860, 2947, 3131, 3132, 3179, 3198, 3295, 3503, 3562, 3587, 3745, 3809]. Analytic [3592, 3709]. Analytical [10, 110, 278, 1397, 2580]. Analyzing [2753]. and/or [3057]. Anderson [2861, 3523, 3547, 3733, 3766]. Anelastic [1139]. Angle [806, 3747]. Angular [1361]. Anisotropic [23, 496, 543, 599, 834, 881, 1057, 1145, 1195, 1434, 1507, 1696, 1747, 2044, 2106, 2273, 2412, 2529, 2651, 2835, 2884, 2977, 3023, 3072, 3108, 3288, 3304, 3397, 3459, 3540]. Anisotropy [1331, 2917]. annealing [131]. Annihilation [1589, 3620]. announcement [5, 11, 189]. Annular [1513, 1557]. Anomalous [2594]. ANOVA [1694, 3708]. Ansatz [3500]. Anti [591]. Anti-Dissipative [591]. Antidiffusive [417]. Any [1192, 1240, 1822, 3443, 3603]. Any-Order [3443]. AO [3615]. Apertures [2667]. Appearance [1231]. Application [26, 28, 77, 118, 208, 229, 274, 309, 318, 345, 349, 378, 408, 416, 431, 591, 624, 672, 677, 688, 692, 705, 719, 758, 781, 999, 1028, 1041, 1160, 1254, 1361, 1470, 1499, 1516, 1524, 1552, 1626, 1693, 1700, 1722, 1793, 1904, 1934, 1974, 2002, 2016, 2081, 2127, 2173, 2269, 2313, 2392, 2482, 2538, 2589, 2622, 2686, 2742, 2744, 2798, 2857, 2974, 3011, 3052, 3102, 3159, 3289, 3317, 3380, 3461, 3689, 3716, 3725, 3747, 3766, 3795]. Applications [51, 122, 141, 150, 214, 233, 265, 298, 418, 484, 485, 568, 632, 737, 833, 835, 879, 1083, 1149, 1153, 1433, 1464, 1644, 1655, 1807, 1810, 1842, 1956, 1962, 1987, 2008, 2020, 2024, 2029, 2035, 2064, 2184, 2233, 2238, 2361, 2422, 2547, 2618, 2747, 2751, 2780, 2875, 2876, 2923, 2945, 2982, 2984, 2987, 3001, 3008, 3031, 3103, 3126, 3133, 3160, 3184, 3202, 3210, 3244, 3257, 3269, 3299, 3350, 3421, 3422, 3429, 3615, 3620, 3633, 3671, 3807]. Applied [143, 174, 442, 445, 545, 660, 1055, 1536, 1590, 1613, 1802, 1852, 2176, 2475, 2861, 3374, 3470]. Applying [128]. Approach [19, 36, 107, 135, 138, 141, 271, 316, 348, 377, 380, 384, 478, 543, 558, 602, 715, 725, 792, 829, 843, 932, 1061, 1153, 1224, 1237, 1327, 1371, 1379, 1393, 1423, 1432, 1539, 1547, 1721, 1723, 1859, 1896, 2004, 2053, 2122, 2158, 2172, 2278, 2314, 2323, 2339, 2349, 2407, 2438, 2490, 2496, 2517, 2520, 2562, 2572, 2637, 2692, 2703, 2746, 2783, 2831, 2939, 2964, 2969, 2983, 3003, 3082, 3106, 3124, 3134, 3142, 3157, 3188, 3190, 3261, 3382, 3412, 3444, 3543, 3576, 3645, 3784]. Approaches [335, 2454, 3133, 3200, 3392]. appropriate [96]. Approximants [1953]. Approximate [247, 804, 1216, 1249, 1799, 1857, 2215, 2720, 2854, 2867, 2941, 3015, 3130, 3244]. Approximating [382, 1842, 2961, 2975]. Approximation [254, 265, 278, 355, 448, 449, 506, 571, 586, 594, 631, 672, 680, 720, 747, 793, 800, 803, 804, 812, 831, 887, 891, 931, 988, 1011, 1068, 1072, 1142, 1149, 1195, 1227, 1285, 1293, 1299, 1301, 1316, 1347, 1371, 1470, 1567, 1573, 1642, 1698, 1793, 1805, 1836, 1895, 1902, 1917, 1961, 2033, 2036, 2038, 2090, 2099, 2112, 2114, 2146, 2187, 2191, 2195, 2222, 2274, 2294, 2414, 2436, 2504, 2549, 2609, 2627, 2634, 2712, 2772, 2786, 2900, 2957, 2960, 2963, 2980, 2993, 3053, 3099, 3116, 3169, 3208, 3231, 3256, 3293, 3304, 3306, 3319, 3323, 3339, 3398, 3446, 3474, 3475, 3480, 3551, 3559, 3631]. Approximation [23, 51, 67, 93, 95, 138, 151, 171, 190, 396, 547, 759, 801, 824, 1160, 2395, 2419, 2558, 2573, 2759, 2787, 2846, 2854, 3015, 3120, 3445, 3675]. Approximation-Based [2099]. Approximations [22, 173, 209, 238, 259, 264, 284, 290, 332, 333, 366, 367, 369, 450, 456, 608, 635, 644, 665, 775, 778, 784, 788, 839, 883, 987, 1038, 1042, 1069, 1084, 1114, 1137, 1153, 1206, 1302, 1451, 1538, 1631, 1735, 1760, 1771, 1811, 1822, 1846, 1856, 1876, 1880, 1924, 1939, 1959, 1972, 1977, 2018, 2024, 2045, 2128, 2188, 2196, 2291, 2306, 2313, 2365, 2375, 2450, 2478, 2532, 2579, 2621, 2641, 2643, 2822, 2835, 2859, 2948, 2984, 3001, 3034, 3065, 3109, 3128, 3157, 3270, 3296, 3442, 3513, 3584, 3634]. Aquifer [2522]. Arbitrarily [3054, 3300, 3381, 3682]. Arbitrary [348, 529, 546, 607, 898, 1012, 1238, 1246, 1336, 1525, 1838, 1935, 2122, 2140, 2303, 2397, 2493, 2596, 2613, 2639, 2681, 2843, 2903, 2999, 3030, 3063, 3097, 3566, 3578, 3707, 3722]. Arbitrary-Lagrangian [1525, 2843]. Arbitrary-Order [2397, 2613, 3722]. Arc [2871]. Arc-Search [2871]. Arch [1045]. Arches [1264]. Architecture [3473]. Architectures [1306, 3189]. Area [3524]. Area-Conserved [3524]. Areas [3184]. Arguments [1975, 3069]. Arisen [1697]. Arising [34, 377, 1433, 1507, 1826, 1923, 2132, 2587, 2715, 2841, 3208, 3474, 3497, 3622]. Arnoldi [1199, 3037, 3111]. Arnoldi-Tikhonov [3037]. Arnoldi-Type [1199]. Arrays [886]. Arrow [2916, 3165, 3565]. Arrow-Hurwicz [2916, 3165]. Arterial [558]. Artery [3387]. Article [3218]. Artificial [227, 289, 419, 433, 590, 764, 1531, 2099, 2160, 2300, 2571, 2622, 2841, 3366, 3399, 3672]. ased [679]. Asian [3731]. Askey [546]. Aspects [292, 401, 413, 481, 715, 840, 2187]. Assessment [24, 963, 3006]. Asset [654, 755, 824]. Assets [759]. Assimilation [19, 1056, 1327, 2240, 2351, 3134, 3360, 3374, 3470, 3654]. Assisted [844, 854, 3254]. Assumptions [290, 2168]. Astrophysical [459, 670, 1123]. asymmetric [13, 48]. Asymptotic [12, 578, 900, 945, 1305, 1361, 1368, 1492, 1630, 1829, 2284, 2412, 2471, 2522, 2543, 2624, 2756, 2861, 2985, 3181, 3265, 3406, 3510, 3607, 3680, 3745]. Asymptotic-Preserving [1368, 3181, 3406]. Asymptotically [798, 2112, 2176, 3800]. Asymptotics [64, 402, 697, 1609, 2346]. Asymptotics-Based [2346]. Asynchronous [128, 2116, 2912]. asynoptic [19]. Atlas [1005]. Atmosphere [1965, 2230, 2351, 2665]. Atmospheric [279, 325, 1333]. Atom [401, 413, 423]. Atomic [674, 745]. attached [181]. Attention [3416]. Attenuating [227]. Attraction [1642]. Augmentation [2836]. Augmented [986, 990, 1032, 1102, 1508, 1514, 1520, 1728, 1731, 1815, 1857, 1937, 2750, 2850, 2854, 3015, 3136, 3233, 3383, 3389, 3558, 3641, 3807]. Auto [309, 2277]. Auto-Adaptive [309]. Autoencoders [3423]. Automatic [307, 2507]. Automatically [3465]. Auxiliary [492, 1379, 1848, 2572, 2831, 2866, 2905, 3256, 3278, 3301, 3664]. Average [2275]. Averaged [3797]. Averaging [562, 2144]. AVF [1521]. Avoiding [1917]. Aware [2351, 3365]. Axially [2471]. Axis [3258]. Axisymmetric [345, 520, 553, 934, 1019, 1028, 1803, 3083, 3621].

B [328, 1699]. B-Spline [1699]. Babich [1609]. Babuska [1176]. Back [1337, 3291]. Background [2937]. Backscattering [326]. backtracking [2430]. Backward [1375, 1698, 1708, 2132, 2223, 2370, 2375, 2540, 2590, 2642, 2806, 2819, 2875, 2925, 3221, 3288, 3402, 3422, 3550]. Bad [2260]. Baer [884]. Bakhvalov [2688, 2824]. Bakhvalov-Type [2688, 2824]. Balance [939, 952, 1305, 1621, 2530, 2654, 2684, 3140]. Balanced [548, 936, 948, 952, 957, 1033, 1102, 1133, 1179, 1197, 1409, 1558, 1592, 1597, 1621, 1830, 1832, 2460, 2530, 2605, 2710, 2812, 2824, 2852, 2876, 2903, 2937, 3030, 3221, 3223, 3237, 3262, 3517, 3536, 3654, 3718]. Balanced-Norm [3221]. Balancing [201, 2350, 2987, 3104]. Ball [3477, 3572, 3606]. Banach [2723, 3254, 3537]. Band [2592, 2773, 3017, 3774]. Banded [2757]. Bandwidth [1822]. Barotropic [874, 3680]. Barrier [2046]. Barriers [314]. Barycentric [2101, 2277, 2637, 3056]. Barycentric-Based [3056]. Barzilai [1318, 1595, 2895]. Base [1559]. Based [238, 257, 276, 356, 376, 414, 564, 618, 668, 721, 943, 967, 1120, 1121, 1138, 1224, 1237, 1248, 1294, 1302, 1307, 1348, 1364, 1523, 1533, 1552, 1575, 1641, 1691, 1700, 1713, 1746, 1758, 1836, 1861, 1883, 1902, 1913, 1924, 1964, 1972, 1992, 1998, 2028, 2060, 2088, 2090, 2098, 2099, 2127, 2173, 2260, 2335, 2350, 2380, 2389, 2438, 2451, 2464, 2480, 2520, 2564, 2571, 2572, 2580, 2662, 2674, 2682, 2686, 2697, 2769, 2783, 2822, 2835, 2836, 2943, 3023, 3027, 3159, 3166, 3190, 3203, 3218, 3223, 3228, 3237, 3254, 3319, 3333, 3341, 3372, 3440, 3442, 3459, 3500, 3517, 3543, 3561, 3613, 3628, 3683, 3738, 3773]. Based [225, 231, 236, 241, 295, 407, 418, 603, 634, 823, 892, 906, 911, 912, 970, 986, 1007, 1008, 1036, 1071, 1130, 1178, 1208, 1252, 1289, 1305, 1310, 1426, 1543, 1570, 1571, 1587, 1669, 1685, 1699, 1742, 1810, 1834, 1851, 1872, 1956, 1984, 2006, 2077, 2104, 2114, 2138, 2153, 2214, 2274, 2280, 2283, 2300, 2346, 2363, 2401, 2459, 2461, 2500, 2542, 2546, 2622, 2684, 2703, 2704, 2722, 2788, 2825, 2850, 2881, 2888, 2931, 2936, 2954, 2956, 3004, 3056, 3078, 3101, 3116, 3124, 3127, 3175, 3216, 3227, 3238, 3245, 3326, 3328, 3353, 3354, 3391, 3412, 3420, 3444, 3492, 3520, 3539, 3558, 3581, 3606, 3624, 3662, 3704, 3716, 3741, 3752, 3774, 3808]. Based [132, 218, 671, 732, 951, 999, 1160, 1185, 1482, 1678, 1754, 2222, 2340, 2728, 2860, 2883, 2940, 2994, 3233, 3271, 3404, 3448, 3482, 3797]. Bases [998, 3158, 3465, 3530]. Basic [2, 2701, 3502]. Basins [1642]. Basis [125, 153, 244, 333–335, 631, 863, 883, 988, 992, 1010, 1039, 1164, 1263, 1358, 1412, 1416, 1433, 1513, 1518, 1617, 1626, 1694, 1701, 1758, 1767, 1883, 1904, 1933, 1961, 1993, 2016, 2041, 2071, 2092, 2194, 2347, 2452, 2454, 2493, 2512, 2531, 2630, 2758, 2760, 2967, 3080, 3116, 3353, 3413, 3714]. Batch [1306]. Batteries [2746]. Battle [241]. Bayesian [162, 1250, 1289, 1393, 2410, 3104, 3453, 3522, 3593, 3660]. BBM [2052, 2476]. BDDC [3217]. BDF [638, 971, 1535, 1762, 2886, 3122, 3388, 3510, 3575]. BDF2 [1769, 2563, 2588, 3206, 3219, 3317, 3460, 3495, 3702, 3751]. BDF3 [2958, 2968]. BDF3/EP3 [2958, 2968]. Be [315]. Beam [1548, 3122, 3480]. Beams [526, 850]. Beavers [2994, 3046]. Bed [955, 2652]. Behavior [129, 160, 264, 331, 745, 1784, 2336]. Behaviour [1811, 2756]. Bell [578]. Bellman [1159, 1561, 1852, 1959, 2666, 2715, 2841, 3208, 3256]. Bellmann [591]. Beltrami [121, 1004, 1020, 1297, 1405, 1496, 2222, 2248, 2758, 2760, 2934, 3240, 3619, 3632]. BEM [1638, 2284, 3247, 3464]. Bénard [2240, 2271]. Benchmarks [973]. Bending [1548, 2126, 2129, 2883]. Benjamin [1868, 2437, 2575, 2785, 3548]. Bernardo [2231]. Bernoulli [1286, 1287, 2597, 2864]. Bernstein [2049, 3212, 3676]. Bessel [110, 122, 2035, 3378]. Better [315]. Between [183, 1115, 1256, 1263, 1337, 1408, 1624, 1651, 2574, 2655, 2950, 3528]. Beyond [546, 702]. Bézier [3212, 3676]. BGK [129, 130, 204, 299, 628, 900, 984, 1177, 3502]. BGK-Collision [1177]. BGK-type [3502]. Bi [72, 511, 1023, 1641, 2266]. Bi-Harmonic [1641]. Bi-Hyperbolic [511]. Bi-level [72]. Bi-objective [2266]. Bi-Orthogonal [1023]. Biased [2308, 2583, 3086, 3147, 3463]. BiCGStab [1657]. Bicharacteristics [3770]. Bidiagonal [1421]. Bidomain [406, 3721]. Bifurcating [844, 854]. Bifurcation [933, 1048, 1090, 1182, 1449, 1904, 2454, 2947, 3188]. Big [1807]. Biharmonic [448, 608, 717, 771, 908, 1076, 1273, 1442, 1734, 1775, 1977, 2013, 2061, 2087, 2166, 2169, 2440, 2526, 3059, 3178, 3270, 3328, 3447, 3567, 3726]. Bilinear [1424, 3500]. Bimaterial [1865]. Binary [252, 1396, 2018, 2262, 2439, 2621, 2650, 2900, 2973, 3365, 3451, 3452, 3571]. Bingham [2084, 2236]. Bio [2786]. Biofilm [1998]. Biofluids [1041]. Biological [1679]. Biology [781]. Bioluminescence [1002, 1021]. Bioreactor [829]. Biorthogonal [908, 1571]. Biot [1706, 2249, 3339, 3655, 3667, 3737]. Biot/Elasticity [3737]. Birkhoff [982]. Birkhoffian [1468]. Birthday [2231]. Bisection [785]. Bisection-Like [785]. Bisingular [2332]. Bivariate [695, 2063, 2544]. Black [759, 1063, 1183, 2460, 2545]. Blast [445]. Blind [586, 1041, 2976, 3368]. Bloch [1771]. Block [432, 451, 632, 1481, 1564, 1580, 1637, 1684, 1781, 1818, 1872, 1878, 1900, 1901, 1926, 2000, 2027, 2107, 2216, 2280, 2285, 2286, 2335, 2424, 2726, 2746, 2814, 3418, 3738, 3753]. Block-Based [2335]. Block-Centered [1878, 2000, 2814]. Block-Oriented [1481]. Blocks [529]. Blood [2507]. Blow [978, 1680]. Blow-Up [978, 1680]. Blurred [1003]. Bodies [355]. Body [289, 647, 1838, 2557]. Bogner [3266]. Boiling [379]. Boltzmann [129, 183, 255, 286, 593, 600, 604, 663, 900, 912, 931, 970, 984, 1011, 1043, 1050, 1071, 1100, 1106, 1172, 1177, 1259, 1315, 1326, 1330, 1337, 1368, 1696, 2539, 2615, 2647, 2703, 2793, 3188, 3500]. Bona [856, 2437, 2575, 2785]. Bond [1984, 2138]. Bond-Based [1984]. Boolean [3200]. boosts [90]. Bootstrap [1137]. Borwein [1318, 1595, 2895]. Borwein-Like [1595]. Bose [734, 1420, 1910, 2485, 3386]. Both [1293, 3281, 3672]. Bottom [955, 2812]. Bounce [1337]. Bounce-Back [1337]. Bound [1249, 1357, 1939, 1981, 2802, 2933, 3051, 3054, 3096, 3118, 3174, 3220, 3248, 3392, 3488, 3512, 3769]. Bound-Preserving [1939, 1981, 2933, 3096, 3174, 3512, 3769]. Boundaries [105, 674, 1635, 2428, 2728, 3337]. Boundary [197, 224, 262, 272, 280, 289, 297, 317, 339, 363, 412, 440, 563, 568, 597, 601, 616, 624, 688, 757, 762, 787, 813, 831, 843, 865, 994, 1097, 1144, 1182, 1189, 1193, 1235, 1312, 1337, 1356, 1431, 1446, 1457, 1480, 1492, 1513, 1538, 1546, 1553, 1559, 1562, 1567, 1610, 1643, 1656, 1659, 1689, 1749, 1759, 1790, 1881, 1889, 1913, 1973, 2066, 2083, 2099, 2108, 2119, 2125, 2137, 2166, 2458, 2462, 2468, 2476, 2483, 2549, 2576, 2664, 2670, 2714, 2740, 2748, 2755, 2864, 2898, 2931, 2955, 2992, 3092, 3149, 3156, 3158, 3210, 3232, 3242, 3279, 3429, 3446, 3465, 3482, 3555, 3627, 3629, 3690, 3726, 3729, 3796]. Boundary [4, 7, 15, 25, 28, 40, 56, 61, 81, 99, 112, 144, 185, 227, 291, 435, 543, 796, 799, 819, 888, 890, 895, 904, 1093, 1184, 1229, 1346, 1373, 1484, 1494, 1544, 1619, 1623, 1640, 1697, 1743, 1753, 1784, 1808, 1827, 1838, 1844, 1852, 1909, 1955, 1996, 2111, 2115, 2320, 2321, 2352, 2428, 2440, 2558, 2669, 2794, 2842, 2846, 2949, 2981, 3120, 3178, 3179, 3379, 3647, 3705, 3803]. Boundary-Conforming [813]. boundary-divergence [61]. Boundary-Finite [1784]. Boundary-Unfitted [3410]. Boundary-Value [1749, 1753, 2670]. Bounded [16, 256, 504, 1148, 1671, 1672, 2279, 2355, 2759, 3242]. Boundedness [991, 1856, 2376]. Bounds [264, 657, 1217, 1325, 1367, 1620, 2154, 2196, 2259, 2296, 2851, 3281, 3466, 3665]. Boussinesq [308, 368, 856, 1537, 1692, 2052, 2058, 2100, 2290, 2399, 2609, 2647, 2676, 2696, 3348, 3496]. Boussinesq-Type [308]. Box [566, 972, 2460, 2713, 2825]. Box-Scheme [972]. Boxplot [1928]. BR1 [2185, 2186]. Branch [785]. Branches [1090, 1725]. Bratu [10]. Break [1497]. Breakdowns [2285, 2286]. Bregman [727, 879, 892, 1118, 1124, 1135, 1190, 3537]. Bregmanized [1129]. Bremmer [3289]. Brezzi [1176]. Brillouin [1157]. Brinkman [1630, 1803, 2250, 2259, 2697, 2798, 3066, 3312, 3339, 3501, 3668, 3793, 3812]. Broadband [414]. Broken [3659, 3681]. Brownian [312, 1686, 3528].

Brusselator [2924]. Bubble [36, 226]. Buckley [1455]. budget [65]. Budgets [3351]. Building [529]. Bulk [2464, 3295]. Bulk-Surface [2464]. Bulkley [1102]. Burgers’ [158, 497, 843, 1078, 2151, 2476, 2487, 2664, 2785, 2908, 3466, 3525, 3806]. Burnett [3500]. BV [569]. BV-duality [569]. BVP [3011]. BVPs [2005, 2893]. Bypass [558, 3387].

C [182, 954]. C-90 [182]. C-property [954]. C.R.F. [152]. C0 [2872, 3159]. C0-IPM [3159]. Cable [1636]. Cadzow [2858]. Cahn [464, 803, 851, 1054, 1241, 1418, 1560, 1601, 1724, 1748, 1892, 1945, 2018, 2178, 2224, 2226–2228, 2347, 2362, 2406, 2426, 2439, 2569, 2614, 2656, 2696, 2699, 2743, 2754, 2755, 2766, 2770, 2802, 2931, 2986, 3029, 3040, 3054, 3088, 3118, 3124, 3170, 3185, 3195, 3206, 3220, 3248, 3382, 3439, 3460, 3493, 3552, 3562, 3588, 3614, 3664, 3671, 3702, 3719, 3800]. Calculating [2877]. Calculation [223, 498, 2457, 2773, 3340]. Calculations [615, 1006, 1054, 1153, 1565, 1932]. Calculus [2110, 2184, 2195, 2922]. Calibrationless [1721]. Camassa [2572]. Can [1479, 3281]. Canal [1670]. Cancellation [2414]. Canonical [1606, 1671, 1672]. Cantor [100]. Capability [232, 2885]. Capacity [1734]. Capture [2965]. Capturing [201, 272, 296, 589, 616, 764, 943, 1783, 2338, 2458, 2475, 2506, 2721, 2966, 3079]. Caputo [1358, 1855, 2111, 2119, 2125, 2141, 2391, 2589, 2700, 3250]. Carbuncle [2229, 3717]. Cardiac [406, 1288, 3721]. Carleman [1087]. Carlo [149, 298, 714, 920, 1538, 2433, 2564, 2878, 2912, 3046, 3173, 3324, 3351, 3748, 3795]. Cartesian [410, 613, 787, 1009, 1136, 1332, 1492, 1546, 1664, 1689, 2499, 2644, 3403]. Cascadic [1780]. Case [82, 158, 175, 226, 358, 635, 804, 910, 936, 1031, 1306, 1413, 1649, 2121, 2544]. Cases [1777, 1863, 2372]. Cat [1154]. Catalytic [250]. Categories [1671, 1672]. Cattaneo [1376]. Cauchy [1087, 1350, 1427, 1808, 1985, 2332, 2577, 3589]. Caused [1469]. Cavities [3699, 3745]. Cavity [403, 524, 537, 1092]. CBS [2219]. Cell [168, 424, 688, 1182, 1232, 1408, 1505, 1647, 2338, 2673, 3121, 3283, 3467, 3647]. Cell-Centered [1647]. Cell/Level [688]. Cells [153, 1409, 1783, 1827, 3742]. Cellular [1154]. CENO [3397]. Center [1157, 1951, 2189, 2364]. Centered [188, 1027, 1647, 1809, 1878, 1901, 2000, 2027, 2673, 2814]. Centers [1039]. Central [360, 433, 525, 536, 552, 885, 1150, 1156, 1367, 1409, 1596, 1693, 1759, 1830, 2068, 2275, 2368, 2469, 2498, 2728, 2943, 3073, 3223, 3517, 3569, 3635, 3794]. Central-Upwind [536, 1693, 2469, 2943, 3223, 3517, 3569]. Central-WENO [1150]. Centroidal [843, 1390]. Century [2489]. Certain [245, 292, 2801]. Certified [1010, 1961, 2041, 2053, 2493]. CFD [38, 574]. CFL [581, 1891, 2457, 2685]. CFL-Dependent [2457]. CFL-Violating [581]. CG [1016, 1579, 2938]. CG/DG [2938]. CG/DG-SEM [2938]. Chain [2323]. chained [141]. chained-matrices [141]. Chains [873, 3280]. Challenge [574, 2252]. Chambolle [3332]. Chandrashekar [3559]. Change [2393, 3501, 3748]. Changes [1897]. Changing [2501]. Channel [25, 568, 1084, 1139, 1909]. Channels [947, 1156, 2615, 2710, 3063]. Chaos [324, 1024]. Chaotic [1034]. Characteristic [54, 128, 721, 728, 845, 846, 911, 924, 1681, 1806, 1862, 1899, 2167, 2571, 2852, 2873, 2877, 2994, 3279, 3356]. Characteristic-Based [721]. Characteristic-Curve [728]. Characteristic-Featured [2571]. Characteristic-Wise [2852]. Characteristics [1138, 1226, 1501, 2189, 3212]. Characterization [184]. Characterizing [749]. Charge [2860, 2878, 3520]. Charge-Conservative [2860]. Charlier [2894]. Chebshev [16]. Chebyshev [31, 35, 50, 67, 74, 79, 81, 124, 140, 218, 339, 365, 367, 370, 537, 664, 669, 756, 1067, 1202, 1256, 1572, 1725, 2525, 2544, 2717, 3584, 3628, 3753]. Chemical [32, 243, 1250]. Chemically [3512, 3769]. Chemistry [489]. Chemotaxis [773, 1105, 1408, 1597, 1936, 2311, 2389, 2790, 3243, 3424]. Chern [3763]. Chi [1916]. Chi-Wang [1916]. Choice [2935]. Choices [857, 3431]. Choosing [559]. Chorin [3326]. Chosen [1389]. CIMA [285]. Circle [1347, 1545, 1701, 2948]. Circuits [2624]. Circulant [217]. Circular [401, 544, 903, 2703]. Circularly [423]. Circulation [172, 279, 408, 416, 1268, 1699, 2240]. Class [46, 199, 273, 474, 548, 626, 721, 741, 744, 802, 803, 896, 917, 1018, 1024, 1067, 1220, 1302, 1331, 1376, 1514, 1690, 1709, 1728, 1751, 1835, 1877, 1905, 1931, 2101, 2102, 2107, 2121, 2164, 2188, 2272, 2468, 2479, 2649, 2721, 2767, 2821, 2838, 2906, 2937, 3018, 3070, 3128, 3172, 3196, 3215, 3359, 3497, 3568, 3690, 3693, 3784]. Classes [3527]. Classical [401, 413, 423, 745, 1053, 1456, 2573, 3758]. Classification [2157, 3444]. Climate [491]. Cloaks [2742]. Close [366]. Closed [941, 1284, 2222, 3810]. Closest [686]. Closure [1793, 2076, 2948, 3356, 3442]. Closures [157, 191, 697]. Cloud [2202]. Clouds [1130, 2762, 2992, 3017, 3210, 3602]. Cluster [152, 864, 3010, 3412]. Clustering [1960, 2223, 3269, 3283, 3753]. CM [182]. CM-5 [182]. CNN [3454]. Coarse [713, 3144, 3407]. Coarse-Graining [713]. Coarse-Level [3407]. Coarsening [7, 3407]. Coastal [172, 408, 416]. Coaxial [1737]. Cockburn [2231]. Code [168, 478, 847, 995]. codes [76, 109]. Coefficient [361, 987, 1239, 1544, 2141, 2250, 2283, 2848, 2993, 3149, 3426, 3465, 3531]. Coefficients [162, 519, 560, 578, 590, 732, 791, 909, 1038, 1042, 1319, 1350, 1405, 1434, 1523, 1585, 1762, 1937, 1957, 2075, 2145, 2308, 2349, 2376, 2433, 2525, 2583, 2620, 2634, 2772, 2907, 2942, 3101, 3151, 3256, 3306, 3417, 3712]. Coercivity [561]. COFRB [2493]. coherence [49]. Coherent [177, 1426, 1956]. Cohesive [751]. Cold [797]. Collective [2991]. Collinear [3685]. Collision [472, 1177, 1967, 2533, 3500]. Collisional [2211]. collisions [109]. Collocated [1495]. Collocation [31, 33, 81, 117, 140, 158, 159, 175, 340, 365, 370, 420, 641, 669, 747, 794, 819, 857, 1051, 1078, 1104, 1164, 1202, 1219, 1246, 1263, 1312, 1356, 1369, 1433, 1512, 1582, 1618, 1646, 1709, 1767, 1790, 1870, 1943, 2069, 2111, 2119, 2199, 2354, 2434, 2464, 2636, 2683, 2701, 2870, 3423, 3449, 3479, 3577, 3602, 3636]. Colloidal [2900]. Color [52, 1896, 3240, 3271, 3350, 3461]. Colorization [1660, 2214]. Combined [232, 355, 1100, 1365, 1523, 1613, 1780, 2638, 2736, 2819, 2879, 2896, 3526].

Combining [2835]. Combustion [216, 231, 236, 962]. Common [2723]. Commutative [306, 1096]. Commutator [1185]. Commuting [1028]. Compact [44, 60, 281, 343, 356, 429–433, 486, 501, 605, 702, 724, 816, 868, 944, 1029, 1076, 1080, 1095, 1110, 1112, 1169, 1174, 1180, 1204, 1260, 1262, 1320, 1331, 1365, 1376, 1421, 1457, 1460, 1464, 1543, 1544, 1587, 1615, 1652, 1780, 1863, 1928, 1981, 2040, 2124, 2133, 2141, 2153, 2163, 2492, 2499, 2513, 2593, 2622, 2673, 2730, 2775, 2785, 2814, 2893, 2925, 2970, 3026, 3183, 3307, 3426, 3526, 3590, 3797, 3806]. Compact-WENO [1981, 3183]. Compactly [1416, 2338]. Companion [3081]. Companion-Like [3081]. Comparative [426, 1016, 1579]. Comparing [2507]. Comparison [182, 287, 350, 603, 760, 777, 876, 1242, 1263, 1337, 1408, 1507, 1531, 1578, 1833, 2113, 2415, 2574, 2588, 2864, 3357]. Comparisons [540]. Compartmental [3374, 3470]. Compatibility [36, 74, 3216]. Compatible [1462, 2823, 3318, 3616, 3800]. Compensation [3291]. compensator [139]. Competition [3528]. Complement [1485]. Complementarity [927, 2377, 2660, 3687, 3799]. Complementary [2503, 3420]. Complete [252, 893, 1099, 1420, 1858, 2996]. Completion [1186, 2031, 2480, 2520, 2695, 2895, 2945, 2976, 3093, 3162, 3234, 3309, 3323, 3604]. Complex [202, 305, 504, 644, 1071, 1122, 1140, 1315, 1546, 1598, 1689, 1888, 2352, 2482, 2592, 2612, 2753, 2881, 3094, 3101, 3477, 3511, 3632, 3791, 3795]. Complexes [3413, 3804]. Complexity [1354, 1684, 1727, 1938, 2117, 2750]. Complicated [1295, 1321]. Component [734, 2689, 2804, 2900, 2963, 2978, 3024, 3180, 3512, 3769]. Components [2596]. Composite [862, 1272, 1289, 1402, 1859, 2118, 2134, 2138, 2378, 2471, 2632, 3314, 3414, 3462, 3641, 3745]. Composition [2411]. Comprehensive [2783]. Compressed [3412]. Compressibility [2692, 3366]. Compressible [15, 65, 68, 103, 168, 186, 230, 232, 261, 296, 309, 597, 610, 634, 667, 698, 787, 809, 842, 923, 1083, 1106, 1208, 1320, 1348, 1451, 1452, 1478, 1494, 1616, 1632, 1645, 1869, 1901, 1948, 1979, 1981, 2000, 2011, 2027, 2042, 2185, 2186, 2500, 2510, 2513, 2596, 2623, 2644, 2678, 2755, 2926, 3064, 3146–3148, 3183, 3276, 3343, 3345, 3367, 3393, 3396, 3399, 3434, 3440, 3445, 3467, 3590, 3597, 3616, 3637]. Compression [586, 833, 1288, 1456, 1890, 2160, 2284, 2337]. Compressive [705, 1318, 3451, 3571]. Computable [2196, 2485, 2765]. Computation [39, 45, 53, 66, 90, 116, 141, 192, 227, 241, 246, 253, 283, 326, 360, 385, 422, 478, 514, 602, 631, 677, 681, 683, 698, 763, 902, 969, 1061, 1154, 1383, 1443, 1563, 1649, 1755, 1841, 1898, 1956, 2035, 2266, 2580, 2637, 2894, 2907, 2917, 3031, 3268, 3435, 3458, 3490, 3592, 3655, 3685]. Computational [37, 161, 331, 444, 490, 619, 1122, 1354, 1396, 1590, 1746, 1804, 1938, 2095, 2240, 2378, 2574, 3351, 3355, 3576]. Computationally [1839]. Computations [96, 195, 335, 379, 410, 461, 549, 565, 658, 825, 995, 1227, 1374, 2093, 2471, 2843, 3082, 3140, 3408, 3555, 3772]. Compute [2405, 3281, 3500]. Computed [2495]. Computer [3, 844, 854]. computers [21, 114]. Computing [13, 78, 82, 205, 657, 706, 724, 820, 1049, 1062, 1083, 1157, 1297, 1390, 1414, 1426, 1469, 1545, 1654, 1671, 1672, 1771, 1817, 1910, 2029, 2030, 2080, 2085, 2109, 2184, 2279, 2323, 2325, 2399, 2512, 2592, 2701, 2730, 2782, 2827, 2867, 3094, 3105, 3142, 3152, 3244, 3303, 3352, 3366, 3384, 3628, 3648, 3706]. Comsol [2002]. Concatenated [3458]. Concave [1516, 2297]. Concentrated [393]. Concentration [693, 1037, 2687]. Concepts [378, 2404]. Concurrent [2149, 3658]. Condensates [734, 1420, 1910, 2485, 3386]. Condition [88, 159, 272, 597, 616, 617, 831, 904, 1176, 1312, 1353, 1640, 1656, 1697, 1844, 1854, 1881, 1891, 2066, 2083, 2132, 2137, 2458, 2598, 2714, 2994, 3046, 3178, 3295, 3332, 3429]. Conditional [170, 3435]. Conditionally [810]. Conditioned [998, 1993]. Conditioning [245, 3225, 3487]. Conditions [4, 36, 74, 81, 99, 171, 196, 224, 262, 280, 289, 339, 412, 543, 595, 601, 737, 757, 788, 796, 819, 994, 1184, 1193, 1229, 1235, 1337, 1356, 1367, 1373, 1446, 1492, 1494, 1544, 1559, 1562, 1567, 1619, 1643, 1659, 1687, 1689, 1909, 1996, 2099, 2108, 2115, 2462, 2483, 2576, 2639, 2665, 2685, 2748, 2755, 2767, 2794, 2842, 3216, 3232, 3279, 3465, 3514, 3704]. Conducting [2878]. Conduction [1226, 2083, 2590]. Conductive [2549]. Conductivity [1393, 3046, 3465]. Conductors [2888]. Cone [2377, 3430]. Cone-Preinvexity [3430]. Cones [1649]. Configuration [312]. Configurations [248, 2378]. confined [133]. Confluence [2496]. Conformal [185, 863, 1255, 1341, 1506, 1566, 1611, 1671, 1672, 1831, 1906, 2758, 2992, 3194, 3713]. Conforming [124, 155, 294, 813, 909, 919, 1240, 1434, 1666, 2053, 2057, 2205, 2208, 2253, 2331, 2369, 2676, 2749, 2908, 2935, 2940, 3163, 3294, 3484, 3491, 3513, 3514, 3539, 3676, 3722, 3776]. Conjugate [228, 276, 509, 2269, 2424, 2779, 3014, 3456, 3599]. Connected [1148, 1297, 1506, 1671, 1672, 2279, 3162, 3311, 3700]. Connection [101, 1115, 3169, 3310, 3332]. Connections [1624]. consequences [161]. Conservation [206, 267, 269, 274, 362, 374, 404, 417, 427, 525, 538, 541, 542, 555, 609, 617, 655, 659, 738, 904, 907, 922, 981, 1024, 1058, 1085, 1094, 1099, 1128, 1181, 1194, 1228, 1244, 1291, 1329, 1404, 1407, 1438, 1444, 1445, 1448, 1458, 1464, 1533, 1539, 1664, 1693, 1715, 1745, 1787, 1795, 1799, 1855, 1886, 1928, 1952, 1974, 2016, 2032, 2040, 2105, 2153, 2200, 2221, 2237, 2249, 2331, 2356, 2368, 2394, 2444, 2494, 2536, 2559, 2560, 2571, 2822, 2970, 3000, 3018, 3062, 3074, 3085, 3096, 3107, 3126, 3214, 3215, 3237, 3298, 3334, 3346, 3363, 3413, 3441, 3478, 3488, 3520, 3615, 3677, 3701, 3759, 3762, 3767, 3794]. conservation [174, 3071]. Conservations [3694]. Conservative [386, 555, 590, 597, 621, 666, 667, 698, 809, 846, 940, 952, 953, 1497, 1894, 1899, 1919, 1939, 1948, 2007, 2014, 2043, 2057, 2105, 2113, 2141, 2148, 2209, 2210, 2243, 2251, 2293, 2421, 2678, 2679, 2784, 2785, 2802, 2804, 2811, 2823, 2849, 2860, 2862, 2913, 2943, 2986, 3006, 3024, 3029, 3058, 3075, 3088, 3223, 3226, 3246, 3290, 3345, 3348, 3359, 3375, 3419, 3440, 3548, 3586, 3611, 3637, 3754]. Conserve [666]. Conserved [985, 1710, 3055, 3524]. Conserving [735, 887, 1046, 1079, 1254, 1453, 1487, 1581, 1940, 2201, 2336, 2491, 2575, 2606, 2766]. Considering [1777, 2959]. Consistency [956, 1612, 1776]. Consistent [1340, 1406, 1458, 1787, 2067, 2360, 2730, 2856]. Consolidation [1706, 2249, 3667]. Constant [211, 361, 828, 873, 1409, 2144, 2525, 3249, 3570]. Constants [510, 2219, 2765, 2957]. Constrained [752, 793, 1096, 1132, 1190, 1243, 1388, 1485, 1588, 1625, 1628, 1987, 2199, 2265, 2578, 2704, 2825, 3073, 3137, 3150, 3314, 3558, 3646, 3705, 3778]. Constraint [478, 1258, 1307, 1391, 1573, 2331, 2644, 2802, 3176, 3521, 3597]. Constraint-Defined [478]. Constraint-Free [1258]. Constraints [376, 668, 709, 899, 992, 1005, 1212, 1293, 1857, 2002, 2318, 2350, 2406, 2425, 2593, 2713, 2750, 3131, 3136, 3137, 3170, 3358, 3771, 3783]. Constructing [1820, 2077, 2637]. Construction [377, 479, 666, 1395, 1465, 1614, 1893, 1933, 1956, 2306, 2392, 2546, 2566, 3075, 3115, 3139, 3426]. Constructive [3053, 3194]. Contact [296, 647, 1402, 1454, 1888, 2655, 2982, 3174, 3509, 3521, 3538, 3595, 3600, 3747, 3783]. Container [3783]. Containing [3275]. Contamination [1862]. Context [585]. Continuation [712, 1090, 1420, 1448, 1549, 1990, 2904, 3771]. continued [141]. Continuous [94, 500, 698, 777, 839, 906, 1210, 1245, 1491, 1893, 1917, 1974, 1984, 1999, 2007, 2167, 2188, 2269, 2639, 2892, 2951, 2954, 2960, 3069, 3192, 3199, 3346, 3380, 3398, 3532, 3573, 3674]. Continuous/Discontinuous [3199]. Continuum [2981, 3318]. Contour [1517, 2459, 2918, 3652]. Contour-Integral [2459]. Contours [1006]. Contractile [1610]. Contraction [1413]. Contractive [2672]. Contrast [1911, 2250, 2507, 2853, 3175]. Contrast-Enhanced [2507]. Control [27, 42, 407, 477, 495, 652, 740, 754, 767, 793, 795, 822, 843, 899, 933, 992, 999, 1048, 1096, 1167, 1182, 1288, 1293, 1461, 1485, 1522, 1553– 1555, 1561, 1573, 1588, 1625, 1736, 1764, 1876, 1898, 1905, 1915, 1923, 1991, 2041, 2159, 2199, 2259, 2266, 2318, 2367, 2419, 2451, 2457, 2468, 2476, 2597, 2616, 2664, 2690, 2798, 2837, 2846, 2952, 2955, 2990, 3092, 3120, 3131, 3139, 3151, 3242, 3292, 3317, 3411, 3422, 3459, 3543, 3658, 3670, 3705, 3775, 3778]. Controllability [1768]. Controllable [204]. Controllers [277]. Controls [2041, 2952, 3139]. Convection [58, 145, 211, 248, 288, 393, 403, 437, 466, 568, 636, 761, 786, 789, 795, 800, 827, 836, 841, 844, 846, 848, 854, 855, 924, 932, 971, 1013, 1026, 1134, 1172, 1207, 1226, 1230, 1275, 1317, 1330, 1362, 1433, 1466, 1529, 1536, 1602, 1605, 1613, 1658, 1696, 1701, 1711, 1757, 1825, 1882, 1938, 1982, 2010, 2106, 2159, 2176, 2199, 2218, 2239, 2393, 2468, 2679, 2697, 2707, 2761, 2826, 2862, 2901, 2918, 2927, 2999, 3050, 3051, 3191, 3212, 3213, 3235, 3253, 3301, 3501, 3550, 3643, 3743, 3773]. Convection-Diffusion [437, 1026, 1658, 2901, 3235]. Convection-Dominated [288, 827, 924, 971, 1013, 1825, 3643]. Convective [370, 1578, 1963, 3582, 3614, 3668, 3812]. Convergence [196, 259, 264, 267, 333, 362, 458, 496, 595, 623, 716, 775, 952, 1078, 1087, 1091, 1141, 1166, 1187, 1267, 1278, 1344, 1372, 1411, 1428, 1511, 1516, 1551, 1596, 1598, 1661, 1731, 1762, 1798, 1802, 1805, 1818, 1858, 1915, 1980, 1985, 1989, 1994, 1997, 2003, 2055, 2075, 2097, 2107, 2114, 2150, 2151, 2178, 2258, 2326, 2388, 2418, 2442, 2444, 2526, 2529, 2602, 2623, 2649, 2718, 2729, 2754, 2780, 2794, 2809, 2830, 2861, 2890, 2910, 2954, 2964, 2985, 2988, 2995, 3071, 3076, 3098, 3107, 3124, 3135, 3143, 3250, 3301, 3335, 3362, 3396, 3409, 3419, 3422, 3510, 3520, 3537, 3545, 3547, 3559, 3575, 3630, 3656, 3673, 3713, 3721, 3751, 3766]. Convergence [70, 146, 925, 1949, 2079, 2188, 2688, 2785, 2824, 3132]. Convergent [16, 613, 766, 1465, 1703, 1747, 1853, 1872, 1927, 2131, 2171, 2408, 2510, 2578, 2707, 3013, 3072, 3077, 3130, 3299, 3414, 3560, 3589, 3655, 3696, 3768]. Conversion [2292]. Converter [250]. Convex [620, 622, 652, 822, 1025, 1516, 1564, 1648, 1649, 1669, 1684, 1731, 1781, 1923, 2030, 2107, 2118, 2180, 2421, 2708, 2797, 2817, 2857, 2972, 3073, 3178, 3314, 3546, 3604, 3702, 3811]. Convex-Splitting [3702]. Convexification [3149]. Convexified [1130]. Convexity [148, 1416, 1684]. convexity-preserving [148]. Convolution [2277, 2734, 2774, 2922]. Convolutional [1338, 3017, 3115, 3423]. Cooling [459]. Coordinate [400, 602, 1872, 3599]. Coordinates [868, 1174, 1334, 1494, 1717, 2537, 2637, 2899]. Coordination [3261]. Copolymer [2856]. Cordes [3256]. Core [491, 995, 1090, 1557, 2302]. Core-Annular [1557]. Corner [50, 270, 1839, 2372]. Cornered [3749]. Corners [737, 3627]. Coronary [3387]. Corrected [97, 3507, 3622]. Correcting [3268]. Correction [317, 342, 1027, 1115, 1166, 1336, 1500, 1646, 1705, 1815, 1867, 1876, 1978, 2058, 2104, 2186, 2210, 2218, 2286, 2321, 2329, 2370, 2399, 2447, 2611, 2774, 2888, 2968, 3015, 3058, 3085, 3088, 3101, 3145, 3229, 3236, 3291, 3293, 3304, 3470, 3505, 3506, 3508, 3561, 3571, 3639, 3653, 3681, 3729, 3755, 3767, 3769]. Corrector [2722, 3433]. Correlated [1041]. Correlation [1476, 3287]. Correlations [177, 3758]. Corrigenda [2859]. Corrupted [1186]. Cosine [1100, 3416]. Cost [19, 586, 2405, 2568, 3778]. cost-function [19]. Costly [1394]. Coulomb [2414, 3626]. Coupled [113, 178, 455, 509, 646, 770, 805, 1061, 1089, 1125, 1209, 1262, 1268, 1308, 1326, 1369, 1500, 1706, 1965, 1983, 2052, 2097, 2218, 2262, 2351, 2464, 2522, 2550, 2778, 2860, 2982, 3021, 3068, 3165, 3209, 3227, 3295, 3312, 3321, 3348, 3419, 3421, 3428, 3437, 3448, 3489, 3668, 3678, 3763]. Coupling [294, 1097, 1317, 1638, 1687, 1961, 1968, 2150, 2243, 2269, 2496, 2555, 2665, 3156, 3242, 3247, 3324, 3513, 3631]. Couplings [1992]. Covariance [9, 1022, 3487]. Cowell [2005]. CP [3052, 3093]. CPMG [114]. CPR [1329]. Crank [1303, 1356, 1376, 1512, 1537, 1682, 1825, 1879, 1887, 1892, 2044, 2344, 2602, 2774, 2890, 3252, 3399, 3678, 3734]. Cray [182]. Criteria [332, 881, 2460]. Criterion [1244, 1606, 3478]. Critical [715, 1106, 3183]. Critical-adaptive [3183]. Cross [2450, 2544, 2924, 3261, 3352, 3573]. Cross-Diffusion [2924]. Cross-Dimensional [3261]. Cross-Product [3352]. Crouzeix [1357, 1399, 1437, 1441, 2015]. Cryogenic [2704]. Crystal [27, 563, 1379, 1508, 1678, 1995, 2727, 2866, 3031, 3219]. Crystalline [3658]. Crystals [1157, 1771, 2592, 2900, 3230, 3774]. CT [1120]. CU6 [2469]. Cubature [695]. Cubed [1204, 1328]. Cubed-Sphere [1204, 1328]. Cubic [524, 1785, 1853, 2792, 3135, 3506, 3780]. Cubical [537]. Cuboid [3804]. cuboidal [155]. CUDA [1183]. Curl [1673, 1918, 3007, 3331, 3373]. Curl-Free [3373]. Curlcurl [3484]. Curlcurl-Conforming [3484]. Current [244, 293, 1737, 3083, 3792]. Curse [1923]. Curvature [389, 392, 464, 1132, 1732, 1761, 2248, 2595, 3078, 3170, 3268, 3420, 3524]. Curvatures [681]. Curve [728]. Curved [244, 547, 865, 1085, 1094, 1275, 1279, 1496, 1735, 2330, 2369, 2374, 2428, 3028, 3298, 3337, 3383, 3777]. Curved-Interface [1735]. Curves [680, 3783, 3810]. Curvilinear [513, 1319, 1494, 1624, 1884, 2194, 2296, 2508, 2537, 2559, 2899, 3169, 3586, 3746]. Customized [3693]. Cut [1827, 2931, 3054, 3062, 3348]. Cut-Off [3054, 3348].

CutFEM [2864]. CVEM [3247]. CVEM-BEM [3247]. CWENO [1622, 2966]. Cycle [1182, 1997, 2281]. Cylinder [544, 605, 903]. Cylindrical [1253, 3276]. Czochralski [27].

D [177, 248, 263, 316, 336, 384, 403, 486, 530, 593, 600, 646, 658, 742, 769, 812, 823, 865, 869, 955, 1002, 1006, 1021, 1061, 1078, 1174, 1192, 1236, 1260, 1276, 1306, 1310, 1339, 1359, 1558, 1579, 1641, 1658, 1737, 1748, 1779, 1780, 1795, 1838, 1858, 1888, 2002, 2252, 2529, 2581, 2583, 2597, 2644, 2652, 2707, 2746, 2755, 2808, 3276, 3367, 3434, 3479, 3803]. D/ [600, 1310, 1779, 2581]. DA-PredGAN [3374, 3470]. Dafermos [3478]. Dai [3014]. Dam [1497]. Damp [965]. Damped [1566, 2788, 3315]. Damped/Driven [2788]. Damping [433]. Darcy [441, 455, 830, 1066, 1089, 1203, 1285, 1853, 1901, 1969, 1983, 2027, 2150, 2227, 2243, 2535, 2550, 2696, 2738, 2778, 3021, 3046, 3061, 3066, 3209, 3227, 3362, 3366, 3413, 3631]. Darwin [1829]. Data [19, 78, 131, 475, 512, 562, 564, 606, 693, 817, 889, 1002, 1003, 1005, 1015, 1021, 1041, 1056, 1173, 1266, 1327, 1338, 1395, 1483, 1589, 1700, 1751, 1782, 1807, 1819, 1934, 1944, 1987, 1990, 2016, 2038, 2109, 2260, 2351, 2408, 2410, 2418, 2527, 2629, 2642, 2687, 2690, 2777, 2795, 2836, 2890, 2895, 3065, 3086, 3089, 3091, 3107, 3134, 3154, 3166, 3177, 3179, 3258, 3270, 3316, 3349, 3374, 3385, 3387, 3425, 3444, 3470, 3575, 3654, 3656, 3699, 3735]. Data-Driven [1395, 1700, 2836, 3091, 3166, 3316, 3387, 3425]. Data-Fidelity [1266, 1782]. Datasets [910]. Davey [3764]. Davidson [2079, 2717, 3303, 3352, 3753]. DC [1676, 3150]. DCT [3416]. DCT-Former [3416]. DDG [1949]. DDGIC [1949]. De-noising [1008]. Deblurring [1499, 1570, 1786, 1850, 1864, 3799]. Decay [1352, 1488, 1811, 3552]. Decaying [65, 103, 2112]. Decision [3157]. Decision-Making [3157]. Decomposition [14, 33, 62, 102, 108, 117, 126, 135, 153, 154, 156, 190, 196, 201, 205, 251, 336, 371, 509, 553, 641, 737, 760, 832, 896, 1001, 1007, 1093, 1129, 1224, 1236, 1290, 1295, 1328, 1364, 1391, 1446, 1454, 1462, 1564, 1608, 1637, 1703, 1764, 1873, 1918, 1959, 1990, 2112, 2131, 2207, 2283, 2292, 2350, 2453, 2461, 2603, 2662, 2778, 2867, 2873, 2895, 2921, 2934, 2954, 2994, 3052, 3134, 3162, 3166, 3231, 3282, 3303, 3309, 3311, 3326, 3335, 3352, 3466, 3477, 3489, 3492, 3541, 3674, 3695, 3700, 3707]. Decompositions [245, 1817, 2735, 3330, 3355, 3776]. Decontamination [1679]. Deconvolution [1032, 2078, 2091, 2976]. Decoupled [1203, 1370, 1500, 1560, 1788, 1791, 2227, 2614, 2696, 2811, 2900, 2982, 3007, 3066, 3247, 3399, 3437, 3591]. Decoupling [3132, 3655]. Deep [2157, 2783, 2829, 2887, 3053, 3152, 3177, 3189, 3245, 3267, 3282, 3292, 3319, 3342, 3357, 3394, 3454, 3623, 3643, 3675, 3728, 3732]. Deep-HyROMnet [3319]. DeepSets [3138]. Defect [1705, 3236, 3304]. Defect-Correction [3236]. Defective [895]. Defects [3658]. Deferred [317, 342, 1166, 1646, 2058, 2218, 2370, 2611, 3229]. Defibrillation [1288]. Defined [478, 724]. Definite [2327, 2377, 2556, 2877]. Definition [364]. Deflagration [231]. Deflated [2424]. Deflation [760, 2726, 3455]. Deformable [3379]. Deformation [1309, 1611, 1946]. Deformations [329, 674, 2934]. Deformed [746, 988]. Deforming [3094]. Degasperis [1342]. Degenerate [371, 708, 1163, 1196, 1408, 1998, 2308, 2719, 2850, 3000, 3322, 3596, 3671]. Degraded [1896, 2315]. Degree [340, 1725, 2057, 2256, 2689, 3290]. Degree-Increasing [1725]. Degrees [17, 812]. DEIM [1962]. Delay [1104, 1618, 1993, 1999, 2005, 2101, 2635, 2775, 3249, 3510, 3570, 3622]. Delayed [2877]. Delays [2116]. Delta [1886, 2201]. Demi [2672]. Dendritic [380, 763]. Denoising [569, 826, 1032, 1252, 1307, 1364, 1470, 1851, 1864, 2280, 2595, 3035, 3047, 3240, 3656]. Dense [1828, 2284]. Densities [1895]. Density [236, 642, 643, 1127, 1370, 1929, 2427, 2747, 2878, 3030, 3077, 3379, 3495, 3591, 3675]. Density-Equalizing [2747]. Dependence [806]. Dependent [49, 77, 126, 187, 232, 312, 352, 532, 557, 762, 770, 831, 966, 976, 1014, 1134, 1226, 1355, 1371, 1401, 1446, 1453, 1465, 1561, 1613, 1665, 1784, 1923, 1995, 1996, 2044, 2064, 2075, 2197, 2235, 2290, 2313, 2347, 2403, 2423, 2443, 2446, 2447, 2457, 2616, 2620, 2640, 2662, 2694, 2712, 2773, 2783, 2859, 2868, 2911, 2942, 2964, 2994, 3005, 3019, 3050, 3075, 3218, 3333, 3341, 3354, 3369, 3406, 3417, 3459, 3492, 3583, 3595, 3663, 3788]. Depth [806, 2100]. Derivation [104, 987]. Derivative [632, 1393, 1509, 1662, 1663, 1715, 1976, 1996, 2067, 2111, 2190, 2465, 2589, 2740, 2961, 3020, 3138, 3453, 3559, 3704]. Derivative-Informed [3453]. Derivatives [302, 706, 1038, 1042, 1388, 1636, 1820, 1842, 1855, 1874, 1976, 2035, 2288, 2391, 2484, 2864, 3275, 3307, 3526, 3529, 3592, 3650]. Derived [312, 760]. Descent [640, 1034, 2053, 2223, 2507, 2658, 2915, 2923, 3014, 3149, 3241, 3456, 3599, 3783]. Describing [743, 2594, 3528]. Description [3729]. Descriptors [2934]. Design [42, 91, 558, 1037, 1314, 1373, 1421, 1907, 2106, 2304, 2580, 3453]. Designing [2407]. Destruction [2361]. Detached [1578]. Detached-Eddy [1578]. Detailed [489]. Detecting [1819, 2410, 3594, 3643]. Detection [559, 693, 784, 913, 1003, 1031, 1036, 1037, 1338, 1642, 2071, 2236, 2687, 3420, 3589, 3593, 3644]. Detector [2622]. Detectors [564, 3115]. Determinantal [2787]. Determination [1389]. Determining [3805]. Deterministic [2102, 3532]. Detonation [381, 1077, 1150, 3638]. Developing [1644, 2753, 2885]. Development [555, 847, 1854, 2914]. Developmental [781]. Developments [877, 2235]. Deviation [2325]. Devices [2403, 2684]. DG [772, 774, 776, 795, 887, 1221, 1233, 1367, 1550, 1638, 1834, 1921, 2243, 2331, 2368, 2380, 2381, 2468, 2474, 2676, 2812, 2908, 2920, 2983, 3030, 3061, 3095, 3198, 3300, 3463, 3467, 3625, 3729]. DG-IMEX [1921]. DG-Methods [1221]. DG-Mixed [2676]. DG-SEM [2938]. DG/Mixed [2243]. DGP [3623]. DGSEM [1436]. Diagnosing [2435]. Diagnostics [187, 575]. Diagonal [1707, 2366, 2743]. Diagonally [696, 3253]. Diagrams [1420]. Diamond [3304]. Diblock [2856]. Dictionary [1426]. Dielectric [260, 498]. Dielectrics [859, 1084]. Diffeomorphism [1405]. Difference [239, 264, 270, 281, 282, 332, 343, 360, 366, 386, 486, 542, 552, 573, 590, 593, 613, 762, 882, 928, 987, 994, 1038, 1042, 1080, 1095, 1105, 1140, 1143, 1150, 1350, 1404, 1426, 1439, 1448, 1457, 1460, 1481, 1495, 1539, 1546, 1568, 1635, 1689, 1702, 1716, 1758, 1759, 1860, 1862, 1878, 1882, 1901, 1902, 1913, 1928, 1965, 1979, 1981, 1984, 2000, 2027, 2066, 2067, 2075, 2082, 2133, 2175, 2324, 2389, 2446, 2447, 2466, 2490, 2505, 2513, 2516, 2542, 2576, 2617, 2623, 2673, 2715, 2719, 2728, 2730, 2806, 2815, 2833, 2857, 2933, 2961, 2978, 3013, 3185, 3199, 3255, 3279, 3322, 3447, 3466, 3472, 3567, 3763, 3790, 3806]. Difference [39, 113, 160, 188, 246, 315, 364, 391, 503, 548, 630, 718, 764, 851, 1040, 1044, 1110, 1169, 1244, 1262, 1365, 1412, 1477, 1494, 1544, 1596, 1616, 1632, 1666, 1798, 1996, 2003, 2141, 2203, 2208, 2274, 2382, 2497, 2508, 2585, 2622, 2738, 2814, 2959, 3002, 3045, 3082, 3127, 3307, 3310, 3334, 3426, 3550, 3586, 3696, 3797]. Difference/Compact [3806]. Difference/Finite [1860]. Differencefor [2153]. Differences [405, 1587, 1719, 2173]. Differencing [1397, 1541, 1615, 2456, 2498, 2541, 2739, 2770, 3284, 3614, 3760, 3780]. Different [603, 1033, 1416, 1488, 2363, 2378, 2615, 3327, 3741]. differentiable [2479]. Differential [202, 212, 213, 264, 265, 333, 340, 479, 731, 741, 744, 747, 756, 815, 832, 1068, 1086, 1168, 1173, 1246, 1286, 1287, 1313, 1366, 1461, 1518, 1530, 1635, 1665, 1680, 1694, 1708, 1725, 1749, 1750, 1821, 1874, 1890, 1898, 1923, 1926, 1943, 1962, 1999, 2008, 2009, 2019, 2039, 2043, 2101, 2110, 2124, 2128, 2173, 2216, 2235, 2324, 2337, 2344, 2349, 2354, 2375, 2391, 2404, 2464, 2493, 2512, 2517, 2611, 2654, 2662, 2700, 2793, 2801, 2919, 2955, 3110, 3157, 3199, 3203, 3218, 3229, 3250, 3288, 3320, 3333, 3342, 3402, 3458, 3473, 3492, 3510, 3529, 3535, 3549, 3563, 3576, 3602, 3639, 3645, 3669, 3698, 3782, 3788, 3802]. Differential [63, 90, 136, 142, 164, 302, 398, 894, 1051, 1093, 1104, 1206, 1322, 1352, 1491, 1618, 1751, 1870, 1993, 2005, 2086, 2188, 2367, 2370, 2744, 2818, 2819, 2947, 3034, 3282, 3526, 3636, 3708, 3756]. Differential-Algebraic [1246]. Differentially [403]. Differentiation [132, 474, 624, 1358, 1698, 2507, 3288]. Diffraction [220, 2174]. Diffuse [2070, 3041]. Diffusion [196, 255, 442, 452, 495, 561, 568, 599, 636, 697, 740, 761, 800, 841, 848, 946, 979, 1009, 1026, 1063, 1191, 1210, 1230, 1275, 1279, 1330, 1362, 1466, 1503, 1507, 1512, 1562, 1568, 1585, 1602, 1613, 1696, 1746, 1759, 1777, 1858, 1860, 1866, 1882, 1899, 1975, 1998, 2089, 2096, 2106, 2108, 2143, 2159, 2176, 2180, 2187, 2199, 2207, 2217, 2239, 2313, 2403, 2408, 2421, 2443, 2468, 2521, 2527, 2541, 2570, 2620, 2675, 2679, 2772, 2826, 2828, 2829, 2918, 2920, 2924, 2927, 2953, 2991, 3019, 3022, 3050, 3051, 3129, 3191, 3212, 3254, 3304, 3312, 3536, 3557, 3577, 3603, 3612, 3643, 3668, 3671, 3723, 3773, 3787]. Diffusion [97, 142, 247, 344, 392, 437, 755, 780, 781, 786, 795, 830, 836, 846, 855, 893, 909, 924, 971, 1043, 1070, 1134, 1169, 1184, 1282, 1283, 1317, 1412, 1433, 1434, 1477, 1517, 1529, 1536, 1647, 1658, 1711, 1732, 1747, 1752, 1753, 1757, 1877, 1880, 1938, 1942, 1955, 1996, 2026, 2075, 2162, 2163, 2218, 2282, 2283, 2298, 2326, 2330, 2425, 2429, 2446, 2447, 2486, 2518, 2582, 2627, 2635, 2671, 2707, 2757, 2759, 2761, 2824, 2884, 2930, 2949, 2977, 2983, 2997, 2999, 3025, 3043, 3089, 3109, 3131, 3143, 3172, 3179, 3221, 3224, 3249, 3253, 3295, 3296, 3306, 3354, 3385, 3411, 3550, 3705, 3734, 3740, 3743, 3751, 3797, 3801, 3805, 3809]. diffusion [52, 60, 134, 190, 1460, 1544, 1902, 2141, 2901, 3069, 3235]. Diffusion- [3723]. Diffusion-Dispersion [830]. Diffusion-Reaction [836, 855, 893, 924, 946, 1210, 1602, 1613, 2282]. Diffusion-Wave [1184, 1477, 1512, 1880, 2527, 3296]. Diffusions [2355, 2373, 2594, 3528].

Diffusive [1921, 2697, 2821, 3306, 3454, 3474, 3531]. Diffusivity [1747]. Digital [250, 574]. Dike [999]. dilation [72]. Dimension [439, 555, 634, 1240, 1262, 1286, 1287, 1533, 1711, 1823, 1938, 2060, 2192, 2287, 2509, 2589, 2704, 3086, 3369, 3451, 3571]. Dimensional [212, 213, 219, 243, 249, 280, 293, 367, 369, 446, 478, 494, 496, 552, 594, 627, 649, 667, 672, 720, 730, 748, 800, 944, 988, 1029, 1031, 1066, 1076, 1077, 1156, 1157, 1161, 1165, 1228, 1232, 1261, 1312, 1362, 1411, 1413, 1414, 1512, 1521, 1528, 1568, 1608, 1610, 1652, 1659, 1664, 1702, 1710, 1768, 1806, 1812, 1873, 1875, 1897, 1919, 1920, 1978, 1997, 2099, 2121, 2180, 2181, 2214, 2288, 2308, 2319, 2328, 2329, 2364, 2368, 2375, 2443, 2496, 2538, 2553, 2591, 2592, 2651, 2664, 2761, 2770, 2843, 2863, 2917, 2952, 2981, 3055, 3066, 3076, 3268, 3297, 3311, 3397, 3420, 3478, 3638, 3669, 3722]. Dimensional [3, 7, 10, 15, 28, 52, 54, 56, 101, 158, 179, 191, 207, 260, 349, 404, 465, 617, 632, 635, 729, 746, 820, 827, 830, 1083, 1110, 1271, 1315, 1427, 1468, 1574, 1611, 1636, 1693, 1950, 1951, 1996, 2033, 2056, 2068, 2162, 2189, 2322, 2352, 2421, 2492, 2612, 2623, 2729, 2800, 2881, 2897, 3002, 3042, 3261, 3310, 3623, 3637, 3708, 3740]. Dimensional-Splitting [3740]. Dimensional/Two [2496]. Dimensionality [1923, 3446]. Dimensions [120, 199, 499, 507, 557, 571, 621, 625, 671, 714, 791, 1142, 1149, 1206, 1486, 1623, 1899, 2156, 2241, 2391, 2429, 2552, 2570, 2873, 2946, 3013, 3028, 3274, 3313, 3357, 3608, 3632]. Diode [779, 1284]. Dirac [278, 283, 677, 1227, 1833, 1886, 2201, 2803, 3415, 3763]. Dirac-Delta [1886]. Direct [23, 32, 136, 169, 170, 186, 236, 298, 502, 531, 537, 610, 667, 741, 1070, 1218, 1320, 1565, 1765, 1775, 1942, 2040, 2042, 2224, 2554, 3090, 3296, 3337, 3454, 3557]. Direction [1022, 1307, 1387, 1421, 1510, 1549, 1551, 1568, 1702, 1861, 2124, 2348, 2402, 3038, 3368, 3433, 3502, 3673]. Directional [1650]. Directly [1198]. Dirichlet [796, 1516, 1527, 1553, 1562, 2061, 2468, 2724, 2846, 2955, 3120, 3179, 3705, 3713]. DIRK [2046]. Disappearance [1231]. Discontinous [2295]. Discontinuities [24, 256, 366, 374, 696, 1196, 1839, 2090, 2661]. Discontinuity [296, 2599, 3060, 3115, 3644, 3706]. Discontinuous [288, 437–439, 441– 443, 445, 446, 516, 526, 559, 614, 636, 649, 653, 660, 723, 729, 769– 771, 813, 839, 850, 896, 932, 975, 1045, 1047, 1072, 1155, 1162, 1175, 1210, 1234, 1286, 1287, 1295, 1362, 1392, 1401, 1419, 1435, 1528, 1533, 1552, 1567, 1577, 1631, 1664, 1745, 1760, 1832, 1849, 1874, 1875, 1919, 1922, 1968, 2011, 2014, 2052, 2074, 2176, 2196, 2231, 2233, 2235, 2237, 2281, 2287, 2336, 2364, 2415, 2467, 2470, 2475, 2509, 2644, 2678, 2771, 2796, 2823, 2940, 3001, 3123, 3199, 3205, 3211, 3318, 3376, 3557, 3667, 3679, 3712]. Discontinuous [301, 308, 322, 414, 447–449, 452, 513, 519, 529, 531, 617, 671, 717, 773, 775, 777, 797, 802, 808, 857, 883, 926, 974, 1009, 1079, 1097, 1180, 1241, 1245, 1261, 1300, 1306, 1348, 1418, 1472, 1487, 1502, 1569, 1576, 1585, 1667, 1713–1715, 1729, 1765, 1814, 1856, 1865, 1935, 1977, 2022, 2053, 2073, 2120, 2137, 2147, 2164, 2172, 2205, 2239, 2241, 2242, 2244, 2259, 2261, 2268, 2288, 2311, 2338, 2369, 2394, 2433, 2477, 2505, 2554, 2600, 2677, 2709, 2848, 2909, 3023, 3026, 3062, 3101, 3156, 3237, 3297, 3442, 3478, 3512, 3525, 3531, 3564, 3631, 3643, 3663, 3746, 3754, 3769, 3773].

Discontinuous [266, 294, 344, 394, 450, 453– 456, 538, 603, 607, 608, 634, 648, 735, 778, 779, 788, 861, 885, 887, 894, 909, 1046, 1096, 1222, 1254, 1268, 1278, 1281, 1378, 1473, 1498, 1531, 1540, 1592, 1624, 1633, 1658, 1677, 1747, 1757, 1762, 1771, 1783, 1792, 1830, 1936, 1937, 1941, 1943, 1945, 1959, 1983, 2081, 2122, 2138, 2140, 2161, 2168, 2197, 2251, 2254, 2255, 2257, 2269, 2273, 2308, 2376, 2397, 2401, 2491, 2495, 2527, 2535, 2539, 2560, 2575, 2583, 2634, 2639, 2700, 2761, 2804, 2842, 2897, 2898, 2914, 2946, 3024, 3077, 3079, 3124, 3163, 3238, 3260, 3275, 3396, 3432, 3509, 3540, 3550, 3579, 3704, 3787, 3798]. Discontinuous [53, 89, 162, 287, 302, 457, 687, 732, 733, 758, 766, 780, 781, 786, 820, 897, 904, 981, 1013, 1070, 1108, 1134, 1158, 1181, 1197, 1247, 1319, 1342, 1388, 1410, 1434, 1655, 1810, 1947, 1948, 1950, 2198, 2218, 2322, 2419, 2519, 2529, 2605, 2625, 2731, 2799, 2879, 2903, 2962, 2999, 3086, 3250, 3276, 3308, 3428, 3471, 3570, 3597, 3711]. Discovery [3177]. Discrepancy [1307, 1389]. Discrete [520, 525, 592, 599, 663, 773, 774, 804, 806, 1013, 1029, 1078, 1138, 1142, 1158, 1188, 1244, 1276, 1355, 1359, 1389, 1462, 1678, 1690, 1772, 1895, 1957, 2047, 2048, 2066, 2109, 2166, 2313, 2316, 2417, 2452, 2558, 2569, 2608, 2635, 2788– 2790, 2859, 2868, 2889, 2954, 2957, 2959, 2980, 3018, 3084, 3102, 3124, 3132, 3200, 3300, 3304, 3362, 3408, 3413, 3416, 3428, 3437, 3439, 3448, 3496, 3502, 3520, 3521, 3552, 3573, 3627, 3629, 3672, 3739]. Discrete-Velocity-Direction [3502]. Discretely [666, 1388, 3731]. Discretisation [1571, 2312, 3694]. Discretisations [648, 861, 2259]. Discretised [3118]. Discretization [268, 314, 361, 425, 488, 520, 601, 708, 834, 921, 934, 982, 1019, 1026, 1098, 1241, 1245, 1310, 1328, 1348, 1386, 1419, 1429, 1578, 1613, 1667, 1673, 1687, 1688, 1695, 1784, 1912, 1988, 1994, 2203, 2224, 2236, 2263, 2414, 2536, 2661, 2754, 2755, 2920, 2997, 3113, 3131, 3179, 3203, 3272, 3308, 3357, 3431, 3489, 3616, 3666]. Discretizations [136, 304, 346, 412, 439, 451, 483, 534, 739, 766, 909, 926, 1167, 1300, 1417, 1434, 1552, 1716, 1791, 1792, 1943, 1963, 1992, 2043, 2053, 2073, 2086, 2088, 2528, 2537, 2559, 2567, 2608, 2711, 2752, 2803, 2828, 3020, 3096, 3199, 3214, 3217, 3227, 3253, 3298, 3358, 3519, 3612, 3704, 3710, 3721, 3812]. Discretized [83, 1388, 1718, 1785, 1844, 1892, 2138, 3230]. Discretizing [2172]. Discussion [716]. DISH [1725]. Disk [1506, 3103, 3194]. Disk-Like [3103]. Dislocation [2981]. Disordered [2951]. Dispersion [161, 266, 648, 660, 830, 1107, 1158, 1365, 1382, 1385, 1421, 2193, 2247, 2317, 2730, 2959, 3183, 3527]. Dispersionless [1239]. Dispersive [246, 447, 501, 516, 915, 942, 2234, 2336, 2339, 2609, 2731, 2796, 2975, 3226]. Displacement [1683, 1814, 2257, 2416, 2602, 2738, 2879]. Displacements [1948]. Disposal [1862]. dissection [136]. Dissipating [3809]. Dissipation [56, 227, 237, 407, 419, 590, 632, 636, 648, 764, 1107, 1158, 1170, 1382, 1385, 2179, 2200, 2382, 2457, 2832, 2959, 2984, 3033, 3183, 3220, 3315, 3527, 3569, 3586, 3635, 3800]. Dissipation-Preserving [2382, 3315]. Dissipative [516, 582, 591, 705, 2272, 2311, 2313, 2449, 2466, 2575, 3003, 3768]. Distance [187, 2037, 2093, 2444, 3010]. distance-dependent [187]. Distinguish [3758]. Distorted [1068, 1361, 2143, 3043]. Distortions [2607]. Distributed [108, 205, 652, 1485, 1568, 1702, 1763, 1860, 1902, 1988, 2041, 2159, 2162, 2429, 2675, 2705, 2775, 3022, 3753]. Distributed-Order [1568, 1702, 1860, 1902, 2429]. Distribution [88, 596, 783, 867, 954, 1377, 1403, 1893, 2084, 2163, 2356, 3152]. Distributional [2354]. Distributionally [3361]. Distributions [73, 1227, 3204]. Distributive [137, 1185]. Disturbance [3139, 3758]. disturbances [139]. Div [517, 1556, 1673, 2064, 2563, 2749, 2859, 3259, 3347]. div-curl [1673]. Divergence [22, 61, 303, 453, 614, 1052, 1289, 1559, 2022, 2063, 2064, 2161, 2244, 2289, 2331, 2381, 2676, 2807, 3048, 3169, 3217, 3421, 3513, 3598, 3653]. Divergence-Conforming [2676, 3513]. Divergence-Free [22, 303, 453, 614, 1052, 1559, 2064, 2381, 3048, 3421]. Divergence/Connection [3169]. Divergency [912]. Diverse [2378]. DNS [508]. Does [965]. Domain [33, 35, 62, 102, 108, 117, 126, 135, 156, 190, 196, 201, 205, 245, 246, 251, 263, 291, 327, 336, 355, 365, 371, 509, 519, 520, 641, 703, 707, 725, 737, 760, 832, 863, 872, 896, 915, 1012, 1061, 1093, 1120, 1129, 1236, 1290, 1295, 1328, 1343, 1364, 1371, 1372, 1381, 1446–1448, 1454, 1521, 1608, 1644, 1680, 1703, 1763, 1764, 1844, 1873, 1959, 1988, 1990, 2004, 2180, 2207, 2292, 2350, 2461, 2603, 2681, 2735, 2742, 2778, 2921, 2954, 2994, 3134, 3169, 3186, 3242, 3282, 3299, 3395, 3489, 3541, 3548, 3627, 3674, 3695, 3745, 3763, 3776]. Domains [81, 98, 124, 155, 323, 432, 440, 469, 504, 553, 613, 746, 757, 762, 796, 805, 856, 978, 988, 1068, 1088, 1093, 1113, 1275, 1295, 1297, 1299, 1511, 1513, 1518, 1546, 1590, 1689, 1738, 1871, 1888, 1895, 1918, 2047, 2066, 2134, 2213, 2279, 2330, 2374, 2421, 2428, 2440, 2458, 2499, 2653, 2667, 2675, 2893, 2935, 2983, 3103, 3178, 3235, 3298, 3315, 3337, 3369, 3395, 3472, 3474, 3551, 3578, 3659, 3681, 3749, 3801]. Dominance [2743]. Dominant [2795]. Dominated [145, 288, 466, 789, 827, 924, 932, 971, 1013, 1207, 1605, 1825, 2010, 2106, 3643]. Dose [1793]. Double [637, 847, 1194, 2213, 2307, 2697, 2954, 3049, 3312, 3605, 3668, 3739]. Double-Diffusion [3312, 3668]. Double-Scale [847]. Doubling [340]. Doubly [888]. Douglas [3693]. downstream [36]. Downwind [534]. DP [350]. dPDEs [1257]. DPG [1866]. Dramatic [2965]. Drift [2403, 2884, 3787]. Drinfel’d [2833]. Driven [211, 524, 537, 1187, 1395, 1699, 1700, 1732, 1994, 2081, 2493, 2690, 2788, 2836, 2984, 3091, 3166, 3316, 3385, 3387, 3425, 3501, 3528]. Droplet [200, 692]. Dry [1179, 2710, 2737, 2987, 3184, 3262, 3601]. Drying [439, 985]. DtN [2174]. Dual [408, 416, 661, 892, 1278, 1364, 1637, 1731, 1914, 2045, 2067, 2083, 2114, 2170, 2292, 2343, 2484, 2500, 2632, 2741, 2786, 2797, 2994, 3332, 3489, 3498, 3542, 3558, 3646, 3720, 3779]. Dual-Level [408, 416]. Dual-Petrov [661]. Dual-Phase-Lag [2786]. Dual-Porosity [3489]. Dual-porosity-Navier [2994]. Dual-Porosity-Stokes [2343]. Dual-Time [2500]. Dual-Weighted [3779]. Dual-Wind [1278]. Duality [147, 569, 951, 3414, 3430, 3610]. Duality-Gap [3610]. Due [955]. DWR [3051]. DWR-Type [3051]. Dyadic [1841]. Dynamic [40, 235, 298, 512, 751, 989, 1140, 1415, 1567, 1721, 2002, 2372, 2378, 2400, 2561, 2912, 3098, 3166, 3492, 3521]. Dynamical [17, 120, 129, 491, 497, 682, 1117, 1339, 1580, 1642, 2095, 2745, 3091, 3104, 3113, 3152, 3200, 3372, 3425, 3561, 3611, 3613, 3626]. Dynamically [615]. Dynamics [8, 27, 46, 68, 103, 296, 298, 347, 619, 667, 685, 743, 825, 895, 973, 1127, 1133, 1270, 1271, 1374, 1394, 1431, 1600, 1869, 1904, 1973, 2064, 2121, 2212, 3137, 3208, 3477, 3665, 3796]. dynamo [121].

Early [859]. Earth [1965, 2037]. Earthquake [1140]. Eddy [55, 112, 151, 179, 232, 320, 431, 524, 539, 583, 847, 880, 959, 963, 965, 1578, 1737, 2693, 3083, 3792]. EDG [3415]. Edge [559, 564, 784, 913, 1003, 1031, 1036, 1338, 1351, 2051, 2303, 2622, 2687, 3230, 3331, 3593]. Edges [693, 1483, 1819, 1990, 2071, 2410, 2855, 3028, 3061, 3661, 3777]. Editorial [85]. Effect [434, 623, 960, 1787, 2552, 3547]. Effective [1960, 2382, 2680, 2901, 3158, 3692]. Effectively [3368]. Effects [177, 195, 230, 243, 248, 310, 423, 472, 572, 1688, 1796, 2260]. Efficiency [152, 3112, 3324]. Efficient [9, 61, 109, 512, 520, 629, 642, 643, 664, 689, 722, 752, 817, 832, 869, 914, 984, 1006, 1027, 1046, 1121, 1126, 1130, 1145, 1171, 1216, 1222, 1241, 1257, 1330, 1360, 1384, 1389, 1398, 1525, 1538, 1643, 1652, 1667, 1770, 1783, 1831, 1839, 1841, 1846, 1973, 2036, 2042, 2128, 2179, 2182, 2204, 2226, 2314, 2339, 2371, 2377, 2389, 2396, 2409, 2460, 2472, 2486, 2514, 2525, 2586, 2587, 2690, 2762, 2808, 2825, 2847, 2867, 2900, 2941, 2948, 3031, 3033, 3041, 3050, 3051, 3072, 3105, 3106, 3116, 3180, 3184, 3244, 3306, 3326, 3378, 3397, 3416, 3455, 3486, 3519, 3528, 3577, 3651, 3657, 3658, 3811]. Efficient [13, 194, 582, 902, 1871, 1906, 1914, 2109, 2429, 2469, 2683, 2944, 3271, 3359, 3386, 3392, 3511, 3772, 3801, 3808]. Efficiently [3500]. Eigen [794]. Eigen-Problems [794]. Eigenfunctions [1811, 2065, 3016]. Eigenmodes [1139]. Eigenpair [1049, 3153]. Eigenpairs [3648]. Eigenproblem [48, 769, 2523]. Eigensolution [1423]. Eigensolver [3230]. Eigensolvers [768]. Eigenspace [1563]. Eigenvalue [517, 870, 871, 1029, 1126, 1153, 1224, 1225, 1298, 1422, 1639, 1730, 1741, 1755, 1775, 1867, 1908, 2015, 2036, 2079, 2087, 2191, 2204, 2220, 2270, 2304, 2305, 2357, 2385, 2459, 2511, 2562, 2604, 2717, 2726, 2855, 2905, 2973, 2996, 3059, 3095, 3178, 3236, 3281, 3325, 3327, 3328, 3344, 3383, 3386, 3391, 3482, 3534, 3572, 3579, 3619, 3695, 3724, 3732, 3745, 3774]. Eigenvalues [13, 612, 624, 724, 1060, 1157, 1249, 1325, 1357, 1479, 1598, 1620, 1654, 1720, 1811, 1908, 2080, 2549, 2551, 3350, 3540, 3551]. eigenvectors [13, 93]. Eikonal [395, 1057, 1354, 1607, 2651, 3108]. Einfeldt [2899]. Einstein [734, 1420, 1910, 2485, 3386]. Elastic [510, 518, 658, 698, 823, 1044, 1187, 1610, 1714, 1784, 1816, 2122, 2882, 2909, 3327, 3448, 3749, 3774, 3783]. Elastica [1215, 3047]. Elasticity [499, 518, 545, 858, 919, 983, 1005, 1180, 1240, 1350, 1427, 1504, 1540, 2529, 2749, 3163, 3175, 3222, 3272, 3410, 3427, 3443, 3503, 3515, 3661, 3737, 3804]. Elastodynamics [1631, 1885, 2252]. Elastoplasticity [330]. Elastostatic [1838]. Electric [1446, 1841, 2856, 2888]. Electrical [2663, 3036, 3733]. Electro [1326]. Electro-osmotic [1326]. Electrochemical [3421]. Electrochemistry [253, 2746]. Electrodynamics [1829].

Electrohydrodynamic [3591]. Electrolytes [1084]. Electrolytes/Dielectrics [1084]. Electromagnetic [244, 246, 260, 366, 401, 413, 414, 423, 560, 878, 888, 1092, 1771, 2742, 2792, 3012, 3121, 3659, 3676, 3681]. Electromagnetics [331, 444, 557, 1222]. Electromigration [720]. Electron [220, 825, 1359, 2441, 2704]. Electronic [335, 1153, 1565]. Electrons [424]. Electrostatic [1084]. Element [242, 312, 345, 358, 437, 439, 495, 497, 499, 516, 521–524, 526, 601, 653, 659, 679, 709, 720, 763, 773, 811, 822, 839, 896, 931, 979, 1063, 1144, 1151, 1155, 1165, 1225, 1237, 1274, 1293, 1386, 1402, 1424, 1480, 1641, 1659, 1692, 1723, 1805, 1836, 1853, 1860, 1912, 1954, 1972, 2183, 2191, 2196, 2219, 2228, 2233, 2234, 2270, 2289, 2290, 2393, 2445, 2510, 2526, 2619, 2648, 2698, 2706, 2716, 2786, 2889, 2924, 2955, 3007, 3011, 3021, 3028, 3076, 3098, 3151, 3171, 3197, 3212, 3217, 3228, 3243, 3274, 3312, 3421, 3521, 3608, 3661, 3668, 3710, 3719, 3722, 3735, 3812]. Element [290, 301, 305, 308, 322, 329, 352, 355, 447, 464, 510, 513, 530, 531, 717, 793, 802, 853, 857, 899, 976, 1054, 1097, 1141, 1153, 1173, 1214, 1227, 1240, 1249, 1306, 1325, 1340, 1437, 1515, 1524, 1553– 1555, 1656, 1698, 1730, 1734, 1801, 1856, 1876, 1881, 1885, 1894, 1911, 1934, 1969, 1992, 2013, 2015, 2053, 2180, 2193, 2199, 2209, 2210, 2245, 2248, 2249, 2299, 2328, 2329, 2349, 2390, 2521, 2538, 2569, 2574, 2674, 2753, 2784, 2791, 2816, 2820, 2835, 2868, 2901, 2932, 2995, 3012, 3022, 3062, 3067, 3069, 3117, 3129, 3141, 3156, 3158, 3165, 3246, 3254, 3256, 3489, 3494, 3538, 3542, 3591, 3621, 3626, 3642, 3721, 3804]. Element [200, 294, 321, 334, 354, 406, 455, 460, 469, 518, 540, 541, 608, 633, 637, 654, 725, 728, 777, 845, 846, 891, 908, 1012, 1066, 1114, 1187, 1207, 1222, 1273, 1324, 1417, 1441, 1475, 1498, 1501, 1502, 1537, 1548, 1571, 1699, 1749, 1756, 1784, 1862, 1880, 1888, 1915, 1958, 1966, 1983, 1994, 2120, 2253, 2255, 2273, 2282, 2301, 2342, 2344, 2345, 2397, 2426, 2476, 2506, 2645, 2653, 2657, 2664, 2676, 2681, 2693, 2712, 2714, 2740, 2768, 2848, 2855, 2914, 2950, 2993, 3092, 3113, 3128, 3167, 3173, 3227, 3252, 3255, 3281, 3290, 3301, 3396, 3405, 3437, 3467, 3495, 3524, 3556, 3612, 3619, 3634, 3670, 3678, 3737, 3751, 3798]. Element [27, 30, 47, 76, 86, 136, 143, 190, 266, 279, 315, 325, 349, 351, 394, 465, 491, 547, 549, 687, 700, 733, 755, 759, 805, 824, 827, 918, 1069, 1146, 1154, 1195, 1211, 1226, 1233, 1310, 1346, 1357, 1398, 1442, 1482, 1491, 1511, 1601, 1630, 1644, 1673, 1739, 1742, 1751–1753, 1826, 1838, 1847, 1988, 2012, 2087, 2114, 2126, 2152, 2174, 2346, 2419, 2463, 2475, 2547, 2598, 2604, 2610, 2635, 2643, 2646, 2688, 2702, 2707, 2718, 2738, 2773, 2824, 2837, 2860, 2872, 2879, 2880, 3048, 3120, 3163, 3175, 3201, 3222, 3266, 3278, 3294, 3327, 3331, 3337, 3404, 3424, 3491, 3539, 3595, 3705, 3730, 3740, 3750, 3765, 3777]. element [26, 34, 52, 159]. Element-by-Element [526]. Element-Characteristic [1862]. Element-Free [3556]. element/Newton [164]. Element/Operator [2328, 2329]. Element/Operator-Splitting [2328, 2329]. Elements [14, 330, 350, 438, 466, 547, 751, 757, 821, 847, 867, 869, 919, 983, 998, 1018, 1058, 1161, 1213, 1245, 1334, 1351, 1353, 1399, 1467, 1470, 1486, 1556, 1603, 1604, 1717, 1738, 1971, 2026, 2044, 2051, 2252, 2296, 2303, 2334, 2369, 2482, 2570, 2613, 2765, 2882, 2931, 3056, 3082, 3192, 3230, 3310, 3347, 3427, 3484, 3533, 3676]. Ellipse [2049]. Elliptic [333, 421, 440, 496, 724, 741, 754, 778, 975, 1047, 1060, 1064, 1086, 1122, 1126, 1145, 1155, 1161, 1162, 1167, 1192, 1193, 1201, 1209, 1249, 1261, 1263, 1274, 1293– 1295, 1299, 1306, 1325, 1369, 1424, 1513, 1554, 1555, 1573, 1577, 1620, 1643, 1716, 1741, 1760, 1765, 1801, 1802, 1876, 1900, 1954, 2015, 2022, 2041, 2045, 2047, 2063, 2066, 2082, 2168, 2181, 2215, 2241, 2306, 2312, 2328, 2329, 2340, 2345, 2390, 2452, 2458, 2483, 2538, 2542, 2584, 2626, 2634, 2706, 2784, 2816, 2865, 2898, 2905, 2955, 3011, 3013, 3076, 3087, 3090, 3141, 3151, 3155, 3173, 3274, 3383, 3429, 3465, 3533, 3634, 3690, 3710, 3712, 3728, 3810]. Elliptic [88, 105, 171, 182, 700, 897, 1108, 1109, 1195, 1308, 1319, 1372, 1516, 1582, 1588, 1635, 1704, 1740, 1785, 1797, 2012, 2115, 2318, 2367, 2428, 2455, 2639, 2667, 2718, 2814, 2818, 2837, 2893, 3070, 3266, 3337, 3511, 3539, 3598, 3653]. Elliptical [413, 2735, 3732]. ELPIC [168]. Embedded [246, 762, 787, 1827, 2468, 2629, 2888, 2931, 3205, 3647, 3663, 3667]. Embedded-Hybridizable [3667]. Embedded-Hybridized [3663]. Embedding [2319, 3346]. EMI [3742]. Emission [1008]. Empirical [3245]. Employing [609]. Enclosures [2782]. Endpoint [3584]. Energy [68, 103, 450, 552, 590, 666, 667, 917, 969, 1018, 1065, 1255, 1277, 1282, 1283, 1370, 1431, 1441, 1487, 1495, 1516, 1678, 1710, 1718, 1724, 1772, 1788, 1796, 1834, 1868, 1906, 1945, 1975, 2048, 2070, 2172, 2177, 2178, 2192, 2310, 2311, 2336, 2406, 2438, 2439, 2456, 2466, 2474, 2575, 2606, 2656, 2658, 2696, 2766, 2806, 2808, 2831, 2832, 2886, 2900, 2913, 2928, 2946, 2984, 3003, 3029, 3041, 3055, 3058, 3122, 3124, 3127, 3174, 3180, 3185, 3206, 3220, 3221, 3248, 3379, 3381, 3404, 3424, 3460, 3488, 3517, 3532, 3552, 3588, 3591, 3625, 3680, 3682, 3694, 3713, 3714, 3780, 3800, 3809]. Energy-Based [3127, 3404]. Energy-Conserved [3055]. Energy-Conserving [2336, 2606]. Energy-Dissipating [3809]. Energy-Norm [3221]. Energy-Preserving [1868, 2438, 2831, 3180, 3381]. Energy-Stable [552, 1678, 1724, 3424, 3714]. Enforced [2246, 2483]. Enforcement [994]. Enforcing [1003, 3356]. Engine [381]. Engineering [3697]. Enhanced [561, 735, 1478, 1532, 1732, 1928, 1933, 1990, 2507, 3035, 3389, 3487, 3614]. Enhancement [1118, 2157, 2817, 3275, 3720]. Enhancing [3291, 3499, 3508, 3733]. ENO [69, 70, 186, 269, 559, 580, 620, 999, 1758, 2553, 3393, 3498]. ENO-Based [999]. ENO-Like [3498]. Enriched [1399, 2010, 2238, 2639, 2641, 3375, 3793]. Enrichment [1865]. Ensemble [1430, 1565, 2399, 3366, 3390]. Ensembles [1991, 2400, 3366]. Enstrophy [65]. Ensures [3413]. Entangled [3099]. enthalpy [71]. Entries [1186]. Entropic [604]. Entropy [26, 311, 319, 417, 617, 666, 667, 804, 922, 961, 1059, 1229, 1378, 1406, 1458, 1558, 1593, 1677, 1823, 2076, 2113, 2123, 2205, 2293, 2386, 2394, 2401, 2417, 2467, 2498, 2528, 2559, 2560, 2605, 2752, 2789, 2940, 2956, 3057, 3085, 3168, 3237, 3285, 3358, 3367, 3434, 3478, 3502, 3616, 3672, 3679, 3684, 3727, 3746, 3752, 3755]. Entropy-Preserving [3285]. Entropy-Stable [1229, 2528, 2559, 2752, 2940, 3057, 3285]. Entropy-TVD [922, 1823].

Envelope [29, 79, 3136]. Environment [1679]. Environments [229, 1123]. EP3 [2958, 2968]. Epidemic [1903]. Epidemiology [3374, 3470]. EPIRK [2263]. EPIRK-K [2263]. EPIRK-W [2263]. Epitaxial [1840, 2314, 2456, 3122, 3284]. Equal [772, 1502, 3123]. Equal-Order [772]. Equality [2265, 2750, 3137, 3771]. Equalizing [2747]. Equation [196, 338, 340, 358, 422, 448, 464, 482, 495, 497, 504, 516, 563, 601, 624, 665, 722, 868, 929, 931, 966, 980, 984, 1026, 1050, 1144, 1151, 1286, 1287, 1374, 1446, 1512, 1528, 1576, 1590, 1682, 1712, 1716, 1724, 1822, 1833, 1858, 1866, 1873, 1921, 1922, 1997, 2007, 2044, 2099, 2133, 2136, 2143, 2169, 2175, 2178, 2180, 2181, 2228, 2239, 2287, 2336, 2341, 2359, 2380, 2391, 2509, 2514, 2526, 2569, 2576, 2678, 2682, 2685, 2728, 2754, 2806, 2809, 2844, 2851, 2891, 2902, 2920, 3006, 3028, 3087, 3098, 3156, 3207, 3209, 3270, 3277, 3279, 3370, 3504, 3514, 3588, 3629, 3719, 3722, 3768, 3780]. Equation [255, 324, 369, 399, 600, 642, 643, 683, 705, 717, 743, 774, 799, 802, 803, 830, 900, 974, 977, 1054, 1076, 1078, 1080, 1095, 1100, 1245, 1330, 1335, 1347, 1368, 1418, 1460, 1487, 1510, 1878, 1882, 1958, 1977, 1979, 2017, 2070, 2083, 2100, 2132, 2166, 2179, 2224, 2248, 2261, 2324, 2328, 2392, 2437, 2438, 2451, 2461, 2474, 2504, 2572, 2575, 2627, 2651, 2664, 2670, 2700, 2729, 2743, 2748, 2772, 2802, 2829, 2831, 2833, 2839, 2901, 2908, 2954, 2964, 2993, 2997, 3019, 3022, 3029, 3044, 3103, 3108, 3129, 3170, 3181, 3248, 3338, 3356, 3395, 3447, 3460, 3466, 3479, 3496, 3549, 3614, 3652, 3665, 3671, 3678, 3683, 3692, 3763, 3773, 3805]. Equation [239, 260, 394, 608, 700, 779, 820, 912, 953, 970, 1044, 1082, 1098, 1114, 1143, 1184, 1187, 1273, 1303, 1342, 1352, 1366, 1372, 1429, 1484, 1529, 1596, 1601, 1603, 1604, 1633, 1647, 1666, 1739, 1747, 1780, 1798, 1808, 1844, 1868, 1870, 1880, 1887, 1891, 1945, 2003, 2026, 2039, 2061, 2092, 2140, 2151, 2154, 2162, 2189, 2208, 2226, 2274, 2320, 2321, 2346, 2347, 2352, 2362, 2384, 2426, 2440, 2486, 2487, 2492, 2517, 2539, 2625, 2635, 2642, 2744, 2774, 2801, 2811, 2832, 2837, 2866, 2886, 2957, 2977, 3002, 3080, 3084, 3086, 3109, 3124, 3127, 3186, 3201, 3288, 3463, 3481, 3507, 3548, 3550, 3664, 3727, 3740, 3744, 3750, 3803]. Equation [10, 52, 60, 87, 158, 165, 167, 183, 240, 733, 1013, 1067, 1071, 1110, 1260, 1315, 1376, 1387, 1477, 1636, 1655, 1697, 1751, 2058, 2087, 2203, 2237, 2298, 2322, 2329, 2481, 2573, 2582, 2683, 2727, 2759, 2766, 2781, 2785, 2846, 2913, 2951, 3058, 3131, 3179, 3308, 3426, 3534, 3552, 3562, 3801]. equation- [158]. Equations [264, 268, 331, 343, 374, 412, 421, 439, 520, 552, 554, 557, 558, 591, 697, 816, 832, 839, 936, 967, 1014, 1025, 1068, 1138, 1145, 1152, 1159, 1163, 1165, 1179, 1192, 1237, 1311, 1350, 1492, 1519, 1521, 1526, 1552, 1613, 1645, 1659, 1665, 1670, 1680, 1696, 1709, 1719, 1725, 1750, 1803, 1805, 1860, 1874, 1890, 2069, 2089, 2101, 2192, 2212, 2216, 2234, 2235, 2294, 2337, 2364, 2434, 2436, 2499, 2543, 2568, 2591, 2648, 2697, 2725, 2755, 3039, 3052, 3076, 3182, 3208, 3217, 3312, 3322, 3390, 3476, 3493, 3535, 3536, 3596, 3603, 3622, 3668, 3679, 3691, 3709, 3735, 3742, 3802]. Equations [202, 212, 213, 263, 274, 308, 327, 409, 446, 528, 613, 614, 664, 712, 731, 753, 754, 761, 772, 874, 979, 988, 1039, 1056, 1064, 1168, 1175, 1189, 1219, 1261, 1279, 1313, 1369, 1375, 1411, 1466, 1507, 1518, 1634, 1714, 1723, 1746, 1774, 1823, 1824, 1832, 1843, 1849, 1923, 1926, 1927, 1972, 2008, 2048, 2106, 2164, 2173, 2206, 2241, 2272, 2312, 2375, 2396, 2407, 2408, 2421, 2427, 2443, 2464, 2470, 2540, 2561, 2603, 2662, 2691, 2699, 2706, 2770, 2916, 2919, 2955, 2987, 3110, 3157, 3171, 3203, 3218, 3228, 3232, 3250, 3333, 3341, 3402, 3421, 3517, 3557, 3563, 3575, 3627, 3645, 3703, 3746, 3782, 3790]. Equations [215, 249, 262, 265, 290, 420, 479, 532, 597, 671, 701, 740, 744, 797, 853, 926, 946, 976, 1028, 1173, 1191, 1235, 1241, 1246, 1251, 1269, 1304, 1355, 1530, 1535, 1542, 1559, 1568, 1694, 1695, 1702, 1708, 1718, 1734, 1736, 1765, 1768, 1787, 1892, 1894, 1899, 1902, 1931, 1962, 2002, 2009, 2010, 2019, 2066, 2082, 2098, 2100, 2110, 2185, 2186, 2199, 2209, 2210, 2215, 2244, 2245, 2332, 2349, 2361, 2412, 2423, 2433, 2512, 2600, 2675, 2677, 2753, 2813, 2820, 2862, 2868, 2906, 2911, 2918, 2932, 2995, 3013, 3027, 3069, 3195, 3229, 3237, 3242, 3256, 3313, 3419, 3483, 3501, 3544, 3576, 3626, 3643, 3650, 3663, 3682, 3810]. Equations [368, 453, 519, 536, 568, 589, 593, 620, 694, 747, 783, 787, 800, 843, 848, 885, 915, 1033, 1059, 1086, 1198, 1312, 1317, 1344, 1363, 1451, 1455, 1465, 1474, 1562, 1572, 1592, 1677, 1710, 1726, 1748, 1756, 1759, 1789, 1821, 1830, 1845, 1935, 1957, 1981, 2043, 2063, 2075, 2135, 2172, 2250, 2251, 2277, 2291, 2295, 2308, 2388, 2399, 2406, 2476, 2477, 2493, 2527, 2542, 2583, 2606, 2633, 2636, 2654, 2657, 2668, 2670, 2671, 2684, 2711, 2719, 2761, 2763, 2777, 2798, 2803, 2810, 2850, 2852, 2888, 2890, 2927, 2956, 3025, 3060, 3128, 3147, 3220, 3260, 3267, 3301, 3306, 3326, 3328, 3342, 3348, 3364, 3401, 3406, 3411, 3415, 3474, 3492, 3533, 3624, 3669, 3670, 3688, 3723, 3793]. Equations [219, 333, 341, 426, 572, 580, 622, 628, 648, 686, 702, 728, 741, 756, 788, 810, 861, 891, 894, 928, 954, 1012, 1057, 1324, 1339, 1380, 1384, 1406, 1412, 1414, 1417, 1433, 1440, 1475, 1501, 1517, 1537, 1591, 1609, 1632, 1635, 1711, 1749, 1790, 1796, 1852, 1898, 1943, 1999, 2020, 2076, 2081, 2113, 2124, 2130, 2161, 2255, 2344, 2354, 2404, 2420, 2454, 2490, 2516, 2550, 2563, 2594, 2611, 2617, 2645, 2666, 2715, 2793, 2801, 2838, 2884, 2930, 2942, 2946, 2965, 2975, 3008, 3055, 3143, 3193, 3252, 3255, 3262, 3265, 3289, 3296, 3320, 3336, 3357, 3449, 3458, 3473, 3529, 3640, 3694, 3698, 3751, 3788, 3806]. Equations [261, 302, 395, 398, 579, 607, 630, 651, 661, 707, 742, 795, 815, 849, 851, 890, 957, 1001, 1051, 1083, 1091, 1104, 1112, 1133, 1134, 1146, 1185, 1197, 1206, 1229, 1262, 1349, 1354, 1452, 1461, 1484, 1491, 1544, 1561, 1593, 1607, 1616, 1658, 1753, 1762, 1779, 1909, 1959, 1993, 2051, 2086, 2115, 2141, 2198, 2303, 2326, 2360, 2363, 2367, 2370, 2382, 2472, 2498, 2518, 2565, 2589, 2605, 2647, 2742, 2757, 2818, 2819, 2842, 2897, 2949, 2998, 3014, 3043, 3054, 3071, 3172, 3213, 3249, 3282, 3294, 3307, 3335, 3343, 3354, 3359, 3362, 3367, 3400, 3434, 3491, 3510, 3511, 3526, 3565, 3598, 3607, 3636, 3653, 3666, 3675, 3684, 3708, 3752, 3765, 3781, 3809]. Equations [9, 18, 23, 41, 44, 56, 63, 67, 90, 101, 105, 113, 127, 136, 138, 142, 163, 164, 180, 190, 192, 193, 218, 805, 893, 925, 999, 1043, 1069, 1070, 1093, 1169, 1174, 1271, 1322, 1378, 1442, 1536, 1615, 1618, 1619, 1686, 1740, 1752, 1785, 1863, 1950, 1951, 1955, 1971, 1996, 2005, 2145, 2188, 2218, 2429, 2558, 2585, 2612, 2702, 2731, 2812, 2840, 2859, 2903, 2944, 2947, 2979, 2986, 2999, 3005, 3034, 3048, 3088, 3120, 3145, 3198, 3224, 3253, 3381, 3417, 3471, 3734, 3756]. Equiareal [2310]. Equilibrated [1419, 2053, 2570, 2613]. Equilibration [1602]. Equilibria [957, 2266, 2407, 3471, 3566]. Equilibrium [712, 1695, 1712, 2361, 2659, 2948, 3436, 3505, 3555]. Equivalent [595, 1399, 2666]. ERBA [3116]. Ergodic [873, 2323]. erosion [72]. Erratum [416, 854, 855, 1019– 1021, 1042, 1094, 1095, 1283, 1445, 1555, 1604, 1663, 1672, 1675]. Error [95, 264, 290, 415, 442, 443, 464, 504, 649, 695, 726, 729, 740, 767, 774, 802, 803, 822, 834, 933, 974, 992, 1035, 1048, 1055, 1162, 1173, 1251, 1274, 1286, 1287, 1294, 1362, 1374, 1419, 1466, 1522, 1533, 1535, 1553, 1576, 1577, 1602, 1664, 1682, 1683, 1698, 1728, 1729, 1835, 1911, 1926, 1954, 1995, 2036, 2074, 2176, 2196, 2219, 2245, 2261, 2290, 2305, 2358, 2414, 2443, 2462, 2545, 2570, 2587, 2613, 2619, 2648, 2716, 2765, 2769, 2786, 2851, 2889, 2891, 2905, 2925, 3007, 3023, 3051, 3118, 3169, 3219, 3270, 3291, 3312, 3330, 3341, 3389, 3421, 3459, 3527, 3551, 3567, 3600, 3709, 3712, 3719, 3735, 3775]. Error [273, 295, 418, 450, 456, 467, 477, 500, 608, 647, 648, 652, 778, 815, 909, 963, 1009, 1202, 1211, 1262, 1303, 1407, 1436, 1441, 1461, 1498, 1501, 1520, 1588, 1601, 1626, 1630, 1699, 1706, 1756, 1856, 1879, 1975, 2132, 2135, 2154, 2168, 2259, 2273, 2296, 2344, 2347, 2363, 2376, 2413, 2426, 2435, 2441, 2476, 2485, 2583, 2614, 2623, 2635, 2636, 2642, 2645, 2646, 2653, 2694, 2803, 2832, 2883, 2892, 2952, 2957, 2993, 3025, 3070, 3086, 3089, 3092, 3095, 3102, 3117, 3146, 3221, 3268, 3284, 3288, 3317, 3348, 3378, 3396, 3401, 3415, 3439, 3447, 3463, 3488, 3491, 3495, 3550, 3584, 3614, 3634, 3650, 3658, 3665, 3678, 3728]. Error [38, 79, 582, 971, 1067, 1070, 1109, 1135, 1260, 1434, 1536, 1971, 2218, 2298, 2419, 2481, 2573, 2643, 2727, 2781, 2811, 2979, 3005, 3034, 3120, 3308, 3503, 3562, 3587]. Error-Correcting [3268]. Error-Estimator-Based [418]. Error-Landscape [963]. Errors [61, 195, 586, 887, 1079, 1107, 1158, 1915]. Escape [1081]. ESDIRK [2588]. Essential [2970]. Essentially [675, 928, 1030, 1077, 1448, 1584, 1594, 1825, 1884, 1979, 2098, 2478, 2720, 3377]. Estimate [442, 443, 1109, 1262, 1588, 2413, 2441, 2573, 2598, 2645, 2727, 2889, 2993, 3219, 3284, 3614, 3728]. Estimates [31, 262, 290, 464, 590, 652, 695, 726, 740, 774, 778, 802, 815, 822, 909, 1009, 1035, 1055, 1128, 1162, 1173, 1211, 1274, 1294, 1434, 1454, 1461, 1498, 1501, 1535, 1536, 1553, 1577, 1630, 1682, 1683, 1729, 1756, 1848, 1911, 1954, 1995, 2074, 2135, 2168, 2176, 2215, 2218, 2261, 2298, 2305, 2363, 2426, 2485, 2537, 2583, 2587, 2619, 2623, 2635, 2642, 2643, 2646, 2648, 2653, 2716, 2786, 2803, 2811, 2832, 2865, 2892, 2905, 2952, 2979, 3005, 3034, 3086, 3089, 3095, 3102, 3120, 3146, 3270, 3278, 3308, 3330, 3341, 3389, 3396, 3415, 3421, 3447, 3463, 3491, 3536, 3550, 3567, 3584, 3634, 3678, 3735, 3775]. Estimating [2507]. Estimation [162, 295, 367, 384, 450, 456, 477, 729, 736, 803, 834, 992, 1127, 1202, 1265, 1287, 1362, 1419, 1436, 1520, 1602, 1626, 2009, 2219, 2273, 2347, 2549, 2636, 3210, 3534, 3551, 3554, 3600, 3719, 3720, 3788]. Estimator [418, 647, 1466, 1533, 1835, 2570, 2613, 3023]. Estimators [273, 295, 1441, 2358, 3779]. ETDRK32 [3760]. Euler [18, 36, 44, 56, 261, 274, 630, 677, 705, 742, 865, 925, 945, 953, 1011, 1083, 1138, 1215, 1229, 1286, 1287, 1451, 1452, 1552, 1592, 1598, 1612, 1616, 1632, 1645, 1695, 1789, 1796, 1824, 1832, 1843, 1951, 1981, 2113, 2189, 2206, 2212, 2223, 2364, 2423, 2472, 2498, 2540, 2543, 2568, 2600, 2694, 2852, 2975, 2979, 3025, 3047, 3146, 3147, 3182, 3221, 3364, 3476, 3680, 3781]. Euler-Type [1612]. Euler/FEM [3221]. Euler/Navier [36]. Eulerian [375, 395, 398, 491, 588, 1230, 1525, 1609, 1935, 1956, 2122, 2158, 2221, 2546, 2596, 2681, 2721, 2843, 2903, 2999, 3018, 3063, 3754]. European [755, 759, 824, 1063]. Eutrophicated [829]. Evaluating [3518, 3795]. Evaluation [214, 502, 1216, 1276, 1447, 1456, 1932, 2371, 3056]. Evaluations [3111, 3226]. Evolution [9, 237, 584, 708, 783, 949, 1001, 1007, 1417, 1510, 1777, 2711, 2846, 2942, 2993, 3193, 3640]. Evolutionary [1536, 1556, 2399, 2801, 3529]. Evolutions [632, 2965]. Evolving [1379, 1727, 2217, 2224, 3452]. Exact [757, 819, 1186, 1388, 1844, 2102, 2176, 2373, 2725, 3153, 3345, 3413]. Exactly [3373]. Example [867, 2487]. Exceeding [1845]. Exceptions [1256]. Exchange [959]. Excitable [837, 1467, 3742]. Excitation [1469]. Excitations [3083]. Existence [1309, 1707, 3392]. Existing [2278]. Exp [2050]. Expanded [1722]. Expansion [104, 247, 278, 634, 671, 1609, 1713, 1773, 2565, 2850, 2974]. Expansions [546, 706, 1694, 3518]. Expectations [3435]. Experimental [3453]. Experiments [69, 3431]. Explicit [117, 261, 263, 405, 484, 628, 738, 782, 875, 904, 1014, 1322, 1363, 1403, 1407, 1455, 1459, 1612, 1632, 1662, 1663, 1776, 1778, 1827, 1855, 1925, 1931, 1945, 2072, 2073, 2123, 2140, 2193, 2220, 2257, 2370, 2381, 2608, 2754, 2986, 3074, 3088, 3096, 3388, 3476, 3535, 3550, 3716, 3723]. Explicit-SAV [2979]. explicit/implicit [117]. Exploiting [1306]. Exploration [101, 1707]. Explosive [379]. Exponent [2158]. Exponential [63, 233, 789, 1015, 1214, 1316, 1368, 1429, 1509, 1615, 1691, 1994, 2098, 2456, 2504, 2541, 2612, 2739, 2770, 2777, 2788, 2826, 3044, 3071, 3111, 3180, 3220, 3250, 3284, 3320, 3382, 3496, 3563, 3576, 3614, 3665, 3716, 3760, 3780]. Exponential-Type [3576]. Exponentially [16, 209, 238, 1801, 2408, 3130]. Exponentials [3302]. Exponents [3789]. Expression [1005]. Extended [168, 1569, 1842, 1865, 2299, 2463, 2573, 2626, 2983, 3111, 3410, 3530]. Extended-Rational [3111]. Extending [872]. Extensible [689]. Extension [578, 602, 627, 630, 701, 940, 1099, 1624, 1863, 3337]. Extensions [748, 813, 1275, 1769, 1858, 2666]. Exterior [551, 704, 862, 1097, 1272, 2066, 2110, 2483, 3247, 3395]. External [543, 1695]. Extracellular [406]. Extracting [674]. Extraction [1956, 3412]. Extragradient [2431, 2910, 2936, 2988, 3049, 3537]. Extrapolated [1168, 1356, 2890, 3103, 3300]. Extrapolation [133, 249, 269, 891, 1029, 1196, 1292, 1546, 1689, 1780, 2353, 2601, 2836, 3696]. Extremal [218, 1157, 1297, 2748]. Extreme [1654, 2947, 3325, 3458, 3781]. Extrinsic [751]. Eye [1154].

Faber [3716]. Face [3461]. Faces [3225]. Facet [2142, 3513]. Facing [349, 3741]. Factor [77, 1363, 1711, 1938, 2502]. Factored [1057]. Factoring [2372]. Factorization [1485, 2031, 2480, 2691, 2807, 3125, 3289, 3461]. Factorizations [2618]. Factors [1037, 1057, 2687, 3052, 3162, 3407]. Families [888]. Family [89, 438, 967, 1110, 1238, 1495, 1684, 1837, 2880, 3318, 3328, 3393]. Far [1229, 1447]. Far-Field [1229, 1447]. Farin [3469]. Fast [72, 150, 251, 409, 421, 534, 579, 622, 714, 741, 746, 753, 794, 826, 875, 890, 930, 972, 1057, 1092, 1142, 1189, 1218, 1219, 1276, 1354, 1363, 1390, 1444, 1445, 1447, 1455, 1484, 1506, 1517, 1607, 1615, 1625, 1671, 1672, 1690, 1838, 1864, 1882, 1955, 1984, 1986, 1996, 2071, 2084, 2091, 2096, 2138, 2165, 2166, 2190, 2216, 2233, 2324, 2333, 2352, 2388, 2546, 2617, 2640, 2651, 2720, 2800, 2838, 2858, 2906, 2911, 2920, 2927, 3000, 3016, 3038, 3056, 3090, 3129, 3172, 3240, 3265, 3366, 3393, 3472, 3493, 3541, 3554, 3613, 3627, 3674, 3685, 3743, 3774, 3795, 3797, 3806]. Faster [1034]. Fat [291]. Fault [2213]. Faulty [252]. FC [1239]. FD [1412, 2893, 3363, 3431, 3620]. FD/ [3620]. FDEM [315]. FDEs [2525]. FDM [1899]. FDTD [750, 1710, 2192, 2888, 3101, 3313]. Feasibility [1583]. FEAST [3628]. Feature [2157, 2836, 3420, 3697]. Featured [2571]. FEEC [3659, 3681]. Feedback [139, 843, 1679, 1736, 2476, 2664, 2690, 3139]. FEM [353, 500, 795, 971, 1053, 1061, 1089, 1599, 1638, 1846, 1887, 1907, 1975, 2007, 2064, 2243, 2330, 2380, 2483, 2555, 2620, 2743, 2803, 2960, 3025, 3118, 3221, 3272, 3370, 3398, 3549, 3664]. FEM/DG [795]. FEM/SBFEM [1061]. FEMs [1251, 1258, 1303, 1535, 1879, 2402, 2416, 2602, 2811, 2994, 3443]. Fermi [278, 283, 677]. FETI [350]. Few [162, 2174]. Feynman [83, 1375, 1686, 1997, 2136, 2642, 2902]. FFT [241, 2728]. FFT-Based [2728]. FFTs [1006]. FGMRES [223]. Fiber [743]. Fibers [563]. Fictitious [355, 725, 1343, 1763, 1988, 3395]. Fidelity [1266, 1394, 1685, 1782, 2230, 2292]. Field [187, 247, 312, 715, 720, 801, 943, 1229, 1370, 1420, 1446, 1772, 1788, 1841, 2018, 2021, 2161, 2226, 2395, 2462, 2727, 2756, 2841, 2866, 2931, 3161, 3174, 3219, 3246, 3300, 3382, 3399, 3452, 3711, 3722, 3760, 3763]. Field-Aligned [2021]. Fields [735, 887, 1133, 1592, 2852, 2856]. Fifth [533, 580, 890, 921, 1438, 1457, 1584, 2478, 2497, 2800, 2897, 3122, 3550]. Fifth-Order [580, 921, 1584, 2478, 3122]. Film [1065, 1277, 1840, 2177, 2314, 2442, 2456]. Films [711, 849]. Filter [498, 1248, 2409, 2475]. Filtered [2291, 2359, 2540]. Filtering [550, 735, 887, 1007, 1079, 1453, 1581, 1940, 2491, 2515, 3287]. Filters [89, 432, 1046, 1254, 1531, 2300]. Final [2846]. Final-Value [723]. Finance [672, 3497]. Financial [1433]. Finding [48, 1193, 1372, 1452, 1516, 1815, 2264, 2265, 2668, 3027, 3153, 3155]. Fine [2636, 2885]. Finite [264, 343, 405, 433, 437–439, 499, 516, 587, 613, 653, 659, 679, 720, 757, 811, 821, 822, 834, 839, 869, 873, 896, 919, 931, 936, 979, 983, 1058, 1102, 1151, 1165, 1192, 1237, 1274, 1311, 1332, 1350, 1386, 1402, 1424, 1467, 1525, 1546, 1597, 1641, 1645, 1659, 1674, 1675, 1689, 1716, 1719, 1805, 1860, 1912, 1920, 1954, 1978, 2032, 2133, 2143, 2146, 2175, 2183, 2191, 2196, 2228, 2233, 2234, 2252, 2270, 2290, 2323, 2393, 2445, 2530, 2543, 2570, 2648, 2679, 2698, 2823, 3007, 3011, 3076, 3151, 3171, 3185, 3197, 3243, 3312, 3322, 3521, 3608, 3668, 3676, 3718, 3735, 3812]. Finite [27, 282, 330, 332, 464, 488, 495, 497, 510, 527, 703, 717, 762, 763, 773, 793, 802, 847, 899, 976, 1054, 1064, 1150, 1153, 1155, 1173, 1227, 1293, 1464, 1505, 1524, 1553–1555, 1692, 1695, 1698, 1723, 1758, 1843, 1861, 1876, 1885, 1889, 1894, 1900, 1911, 1928, 1972, 1991, 2013, 2044, 2053, 2067, 2082, 2173, 2180, 2193, 2199, 2209, 2210, 2213, 2219, 2245, 2248, 2324, 2328, 2329, 2334, 2335, 2510, 2526, 2569, 2576, 2615, 2641, 2706, 2751, 2786, 2791, 2868, 2889, 2901, 2955, 2961, 2978, 3013, 3021, 3062, 3158, 3212, 3228, 3254, 3256, 3279, 3373, 3397, 3559, 3566, 3621, 3626, 3642, 3710, 3719, 3790]. Finite [270, 281, 290, 301, 355, 593, 637, 654, 666, 777, 791, 800, 853, 908, 949, 977, 1140, 1240, 1249, 1306, 1325, 1340, 1404, 1448, 1457, 1460, 1498, 1502, 1539, 1548, 1571, 1730, 1738, 1756, 1811, 1878, 1881, 1908, 1913, 1915, 1969, 1979, 1981, 1983, 1984, 1992, 2000, 2027, 2028, 2075, 2249, 2259, 2299, 2342, 2345, 2349, 2400, 2476, 2513, 2521, 2542, 2600, 2657, 2664, 2671, 2673, 2681, 2719, 2740, 2753, 2765, 2784, 2815, 2820, 2833, 2835, 2850, 2852, 2862, 2931–2933, 2995, 3012, 3022, 3069, 3117, 3128, 3129, 3141, 3165, 3167, 3215, 3227, 3281, 3301, 3427, 3472, 3489, 3494, 3495, 3498, 3533, 3538, 3542, 3591, 3612, 3634, 3678, 3715, 3763, 3804]. Finite [239, 294, 315, 354, 366, 386, 391, 406, 455, 466, 469, 511, 542, 590, 608, 633, 728, 827, 836, 845, 846, 855, 891, 924, 928, 1012, 1038, 1040, 1042, 1044, 1114, 1143, 1146, 1154, 1187, 1195, 1226, 1273, 1310, 1398, 1412, 1417, 1475, 1495, 1501, 1537, 1632, 1635, 1647, 1651, 1699, 1739, 1749, 1783, 1784, 1796, 1880, 1888, 1901, 1958, 1965, 1994, 2104, 2255, 2296, 2344, 2346, 2426, 2446, 2447, 2490, 2516, 2617, 2623, 2645, 2653, 2676, 2712, 2715, 2814, 2884, 2914, 2950, 2993, 3092, 3113, 3173, 3192, 3201, 3230, 3252, 3255, 3290, 3396, 3437, 3447, 3467, 3524, 3586, 3619, 3740, 3751, 3798, 3806]. Finite [34, 246, 287, 364, 404, 465, 547, 548, 687, 700, 733, 764, 851, 893, 907, 1069, 1169, 1211, 1238, 1259, 1353, 1442, 1477, 1482, 1491, 1494, 1511, 1556, 1593, 1596, 1601, 1605, 1616, 1630, 1644, 1666, 1673, 1688, 1740, 1742, 1747, 1751– 1753, 1798, 1826, 1946, 1952, 1971, 1996, 2003, 2012, 2068, 2114, 2152, 2158, 2174, 2203, 2208, 2419, 2429, 2463, 2497, 2508, 2547, 2585, 2598, 2604, 2622, 2635, 2643, 2688, 2702, 2707, 2718, 2738, 2773, 2790, 2824, 2860, 2872, 2879, 2880, 2959, 2966, 2977, 3002, 3043, 3045, 3048, 3120, 3127, 3148, 3163, 3175, 3266, 3327, 3334, 3337, 3404, 3424, 3426, 3590, 3701, 3750, 3765, 3767, 3797, 3809]. finite [14, 26, 39, 52, 60, 76, 96, 113, 136, 159, 160, 164, 177, 190, 236, 1470]. Finite-Difference [39, 246, 282, 364, 3426]. Finite-Element [27, 164, 439, 1069, 1324, 1699, 1972]. finite-element/Newton [164]. Finite-Size [177, 931]. Finite-Volume [1796, 2335, 2615, 2966, 3215, 3397, 3467, 3498, 3559]. Finite-Volume-Particle [527]. First [468, 505, 589, 967, 977, 1049, 1051, 1159, 1355, 1562, 2094, 2332, 2373, 2721, 2796, 2853, 2917, 2953, 2961, 3187, 3328, 3367, 3373, 3496]. First-Order [468, 589, 967, 2094, 2721, 2853, 3367, 3496]. First-Passage [2373]. Fisher [2093, 3689]. FISTA [3414]. Fitted [1459, 2125]. Fitting [716, 1691, 2835]. Five [2678, 2839]. Five-Equation [2839]. Fixed [955, 1781, 2116, 2334, 2479, 2632, 2660, 2672, 2723, 3336, 3537, 3639]. Fixed-Point [1781, 2479, 3336]. Flame [329]. flames [28, 134]. Flat [112, 337, 2812]. Flexible [252, 689, 1657, 1807, 3063]. Flexoelectricity [2872, 3159]. Flexoelectricity-Based [3159]. Flight [574]. Flooding [439]. Flory [3185]. Flow [200, 248, 250, 270, 272, 286, 289, 293, 313, 328, 345, 349, 371, 410, 411, 441, 451, 458, 464, 465, 488, 489, 493, 523, 524, 544, 576, 583, 605, 634, 667, 671, 770, 780, 809, 841, 876, 903, 920, 933, 1004, 1027, 1048, 1163, 1297, 1365, 1386, 1392, 1511, 1524, 1550, 1557, 1600, 1712, 1742, 1761, 1764, 1791, 1895, 1901, 1932, 1956, 1961, 1968, 2007, 2109, 2147, 2152, 2213, 2240, 2243, 2246, 2293, 2299, 2358, 2374, 2487, 2501, 2507, 2546, 2630, 2641, 2733, 2771, 2926, 2989, 3068, 3077, 3123, 3165, 3216, 3321, 3358, 3366, 3413, 3485, 3489, 3524, 3608, 3621, 3679, 3747]. Flow [3, 8, 36, 48, 59, 77, 108, 112, 123, 133, 145, 156, 240, 1020, 1788, 2395, 2766]. Flow-Porous [3068]. Flows [232, 277, 281, 303, 305, 309, 312, 320, 352, 430, 509, 539, 602, 615, 618, 633, 662, 676, 679, 792, 842, 941, 947, 950, 955, 964, 1125, 1154, 1156, 1170, 1207, 1231, 1236, 1309, 1310, 1320, 1336, 1337, 1348, 1370, 1430, 1464, 1469, 1478, 1514, 1547, 1578, 1652, 1718, 1728, 1772, 1825, 1929, 1995, 2011, 2042, 2064, 2065, 2160, 2236, 2256, 2335, 2342, 2361, 2495, 2500, 2506, 2513, 2580, 2609, 2673, 2678, 2710, 2737, 2778, 2804, 2823, 2978, 2984, 3003, 3024, 3033, 3063, 3064, 3183, 3318, 3373, 3388, 3393, 3397, 3399, 3440, 3467, 3495, 3512, 3601, 3633, 3637, 3755, 3769, 3796]. Flows [4, 25, 32, 39, 50, 54, 86, 121, 143, 168, 170, 180, 181, 203, 351, 492, 549, 582, 692, 1083, 1208, 1326, 1482, 1494, 1574, 1678, 1732, 1810, 2482, 2588, 2596, 2614, 2621, 2966, 3061, 3300, 3345, 3445, 3590, 3597]. Fluctuation [373]. Fluid [9, 77, 277, 328, 355, 422, 485, 522, 581, 619, 633, 694, 796, 895, 898, 933, 1048, 1059, 1125, 1311, 1430, 1610, 1718, 1788, 1859, 1869, 1904, 2014, 2018, 2121, 2122, 2131, 2342, 2395, 2506, 2596, 2621, 2926, 3001, 3063, 3276, 3353, 3424, 3445, 3452, 3616, 3633, 3657, 3679]. Fluid-dynamics [895]. Fluid-Structure [522, 694, 3001, 3063, 3353]. Fluid-Surfactant [3452]. Fluids [120, 224, 352, 383, 531, 989, 1281, 1428, 2122, 2937, 3513]. Fluorescence [1826]. Flux [97, 360, 452, 533, 633, 790, 893, 905, 917, 997, 1018, 1099, 1228, 1282, 1283, 1334, 1336, 1382, 1440, 1458, 1578, 1586, 1603, 1604, 1616, 1624, 1717, 1858, 1875, 1942, 2026, 2040, 2105, 2432, 2457, 2536, 2552, 2554, 2576, 2588, 2657, 2732, 2768, 2804, 2987, 3024, 3085, 3096, 3223, 3238, 3253, 3304, 3310, 3343, 3403, 3441, 3517, 3530, 3566, 3638, 3712, 3736, 3755, 3770, 3785]. Flux-Computation [360]. Flux-Conservative [2105]. flux-corrected [97]. Flux-free [633]. Flux-Limiters [2457]. Fluxes [471, 603, 1419, 2053, 2113, 2308, 2322, 2583, 2909, 3086, 3525, 3677]. FMM [3632]. FMM-Accelerated [3632]. Focalization [472]. focusing [75]. Fokker [743, 1339, 1378, 1677, 2020, 2594, 2740, 2841, 3109, 3675, 3688]. Folding [1255]. Folding-Free [1255]. Following [3073]. FOM [2969]. FOM/ROM [2969]. Force [1960]. Forced [68, 2627]. Forces [1695, 1838, 3031]. Forcheiner [2535]. Forchheimer [1066, 1901, 1969, 2027, 2697, 2738, 3209, 3312, 3631, 3668, 3812]. Forcing [586, 2721, 3812]. Forest [3444]. Foreword [300, 372, 428, 436, 476, 515, 765, 866, 2231]. Forgetting [1135]. Form [386, 799, 857, 1265, 1309, 1909, 2022, 2062, 2063, 2141, 2154, 2222, 2677, 2938, 3057, 3127, 3279, 3598, 3653, 3714]. Formalism [2017, 2100]. Formally [1027]. Format [1755, 2443]. Formation [177, 711, 2924, 3243]. Former [3416]. Forms [492]. Formula [106, 677, 1686, 1698, 2222, 2427, 3288, 3491, 3797]. Formulae [1691, 2806]. Formulas [1112, 1292, 2267, 3550]. Formulation [30, 55, 313, 321, 370, 398, 433, 449, 491, 667, 702, 730, 816, 877, 905, 934, 1019, 1044, 1052, 1132, 1174, 1398, 1403, 1480, 1525, 1537, 1548, 1586, 1652, 1713, 1761, 1803, 1854, 2025, 2026, 2061, 2283, 2295, 2436, 2462, 2490, 2499, 2525, 2542, 2561, 2588, 2609, 2616, 2697, 2771, 2849, 3198, 3310, 3482, 3582, 3726]. Formulations [382, 590, 790, 1646, 2467, 2682]. Fort [1075, 1344]. Forth [3291]. Forward [349, 1257, 1375, 1708, 2370, 2435, 2546, 2819, 2875, 3402, 3422]. Forward-Backward [2875]. Forward-Facing [349]. Forward-Reflected-Backward [3422]. FOSLS [3347]. Four [52, 2462, 3006]. four-color [52]. Four-Equation [3006]. Four-Field [2462]. Fourier [16, 35, 67, 78, 125, 153, 158, 162, 170, 173, 175, 238, 256, 280, 336, 341, 494, 506, 578, 704, 756, 1003, 1015, 1036, 1070, 1078, 1189, 1338, 1342, 1448, 1483, 1509, 1589, 1725, 1748, 1811, 1819, 1840, 1949, 1990, 2146, 2165, 2166, 2410, 2470, 2544, 2687, 2772, 2832, 3154, 3202, 3482, 3490]. Fourier-collocation [175]. Fourier-Continuation [1448]. Fourth [239, 281, 368, 440, 485, 605, 862, 868, 972, 1044, 1076, 1110, 1113, 1161, 1286, 1287, 1296, 1346, 1373, 1522, 1534, 1562, 1619, 1623, 1671, 1672, 1690, 1780, 1794, 1801, 1826, 2115, 2197, 2312, 2320, 2321, 2382, 2683, 2785, 2793, 2816, 3067, 3266, 3722, 3723, 3725]. Fourth-Order [281, 485, 605, 862, 1076, 1110, 1286, 1287, 1373, 1534, 1562, 1619, 1623, 1780, 1794, 2320, 2321, 2382, 2683, 2785, 3067, 3266, 3722, 3725]. Fox [3266]. Fractal [3562]. Fractional [532, 1344, 1375, 1460, 1470, 1512, 1562, 1568, 1625, 1749, 1750, 1756, 1821, 1860, 1862, 1866, 1878, 1882, 1890, 1892, 1902, 1913, 1962, 1975, 1997, 2008, 2019, 2069, 2075, 2083, 2096, 2128, 2135, 2136, 2180, 2184, 2187, 2191, 2216, 2225, 2237, 2324, 2337, 2349, 2392, 2402, 2408, 2418, 2421, 2434, 2437, 2443, 2514, 2521, 2527, 2579, 2590, 2597, 2620, 2627, 2671, 2675, 2699, 2700, 2711, 2740, 2743, 2770, 2772, 2851, 2901, 2902, 2922, 2927, 2942, 2993, 2997, 3003, 3019, 3022, 3076, 3089, 3100, 3102, 3110, 3118, 3129, 3130, 3158, 3195, 3229, 3250, 3270, 3296, 3306, 3315, 3411, 3419, 3493, 3531, 3549, 3567, 3576, 3592, 3612, 3622, 3624, 3645, 3665, 3800, 3802, 3805]. Fractional [1169, 1184, 1260, 1376, 1428, 1477, 1517, 1544, 1636, 1658, 1751– 1753, 1842, 1877, 1880, 1955, 1993, 1996, 2111, 2141, 2151, 2162, 2179, 2188, 2190, 2274, 2283, 2298, 2382, 2419, 2429, 2481, 2486, 2518, 2548, 2582, 2589, 2612, 2642, 2702, 2738, 2739, 2744, 2757, 2759, 2832, 2846, 2847, 2930, 2949, 3002, 3109, 3131, 3143, 3172, 3179, 3249, 3382, 3449, 3507, 3528, 3552, 3584, 3636, 3652, 3675, 3683, 3740, 3743, 3772, 3797, 3801, 3806]. Fractional-in-Space [1892, 2382, 2675]. Fractional-Order [2548, 3296]. Fractional-Step [532]. Fractional/Normal [2846, 3179]. fractions [141]. Fracture [751, 3066, 3541]. Fractured [2152, 2522, 2733, 3061]. Fractures [2358]. Frame [1072, 1120, 1121, 1212, 1395, 1523, 1700, 2146]. Frames [1085, 1094, 1279, 1819]. Framework [589, 681, 892, 1144, 1164, 1214, 1329, 1337, 1369, 1448, 1452, 1530, 1646, 1746, 2601, 2893, 2912, 3024, 3091, 3106, 3121, 3238, 3321, 3557, 3589, 3646, 3659, 3681, 3689]. Frankel [1075, 1344]. Fréchet [84]. Fredholm [41, 249, 1634]. Free [22, 48, 59, 303, 453, 457, 465, 487, 523, 614, 633, 692, 784, 903, 1045, 1052, 1182, 1255, 1258, 1270, 1309, 1525, 1552, 1559, 1635, 1668, 1836, 1912, 1955, 2064, 2161, 2229, 2244, 2289, 2320, 2321, 2369, 2381, 2440, 2466, 2490, 2493, 2497, 2508, 2658, 2709, 2749, 2864, 2874, 2975, 2992, 3048, 3068, 3074, 3147, 3217, 3260, 3325, 3352, 3373, 3375, 3421, 3443, 3497, 3504, 3512, 3517, 3555, 3556, 3586, 3729, 3762, 3769, 3784]. Free-Boundary [1955, 2992, 3555, 3729]. Free-Form [1309]. Free-Stream [2369, 2508, 3586]. Free-Stream-Preserving [3147]. Free-Surface [457, 465, 523, 692, 2709, 2975]. freedom [17]. Freeform [2405]. freely [28]. Freestream [1884]. Frequency [46, 91, 514, 519, 685, 1371, 1714, 1813, 2004, 2284, 2412, 2573, 2773, 3140, 3336, 3699]. Frequency-Domain [1371]. Friction [938, 945, 2137, 2462, 2619]. Frictional [936, 994, 3600]. Friedrichs [261, 1361, 1444, 1445, 3376]. Frog [915]. Front [94, 555, 621, 753, 817, 1154]. Front-Tracking [555]. frontal [154]. Fronts [792, 1179, 1414, 1727, 2710, 2987, 3262]. frontwidth [20]. Frozen [2099]. Full [1265, 1557, 2090, 2294, 2568, 2754, 3179, 3694]. Full-Discretization [2754]. Fully [255, 307, 609, 731, 773, 774, 909, 929, 985, 1013, 1078, 1158, 1261, 1328, 1355, 1425, 1647, 1718, 1772, 1892, 1932, 2048, 2082, 2166, 2196, 2215, 2241, 2313, 2417, 2569, 2608, 2635, 2697, 2828, 2859, 2868, 2889, 2959, 3013, 3048, 3084, 3124, 3132, 3162, 3300, 3311, 3362, 3399, 3428, 3437, 3439, 3520, 3521, 3627, 3700]. Fully-Connected [3162, 3311, 3700]. Fully-Decoupled [3399]. Fully-Discrete [1013, 2608, 3132, 3521]. Function [19, 162, 253, 256, 283, 863, 883, 1039, 1302, 1307, 1323, 1379, 1412, 1433, 1513, 1518, 1594, 1617, 1701, 1758, 1803, 1841, 1883, 1909, 1976, 2071, 2090, 2092, 2184, 2194, 2325, 2405, 2490, 2531, 2532, 2630, 2703, 2705, 2760, 2769, 2782, 2888, 3101, 3331, 3662, 3698, 3714]. Functional [2017, 2104, 2537, 2928, 3020, 3298, 3485]. Functionalized [1418, 2178, 2766, 3588]. Functionally [1459]. Functionals [359, 635, 1924]. Functions [22, 51, 110, 122, 209, 238, 244, 271, 506, 606, 704, 706, 860, 1115, 1358, 1416, 1509, 1559, 1685, 1754, 1767, 1800, 1812, 1820, 1866, 1886, 1988, 2008, 2035, 2071, 2170, 2347, 2412, 2639, 2789, 2834, 3251, 3420, 3430, 3584, 3592, 3795]. Functions-I [2071]. Fundamental [292, 592, 1427]. Further [555, 1450]. Fusion [3103]. Fuzzy [1685]. FV [2652, 3620].

Galerkin [288, 323, 438, 439, 441– 443, 445, 516, 526, 559, 614, 635, 636, 649, 653, 660, 698, 729, 769– 771, 850, 896, 975, 1045, 1047, 1162, 1181, 1189, 1210, 1234, 1274, 1286, 1287, 1295, 1362, 1392, 1401, 1413, 1435, 1533, 1535, 1552, 1567, 1573, 1577, 1631, 1664, 1682, 1745, 1760, 1832, 1836, 1849, 1874, 1875, 1919, 1921, 1922, 1968, 1995, 2004, 2014, 2036, 2052, 2074, 2176, 2195, 2196, 2233, 2235, 2237, 2281, 2287, 2336, 2364, 2415, 2467, 2470, 2475, 2509, 2602, 2644, 2678, 2796, 2823, 2940, 3001, 3123, 3199, 3205, 3250, 3318, 3464, 3486, 3667, 3679]. Galerkin [127, 175, 301, 409, 446–449, 529, 535, 600, 617, 664, 671, 717, 775, 797, 802, 808, 813, 926, 929, 932, 949, 974, 1097, 1155, 1175, 1180, 1241, 1245, 1251, 1261, 1348, 1418, 1419, 1472, 1487, 1524, 1528, 1569, 1576, 1643, 1667, 1713, 1714, 1729, 1765, 1814, 1865, 1924, 2007, 2011, 2022, 2045, 2053, 2147, 2164, 2205, 2219, 2239, 2241, 2242, 2244, 2261, 2268, 2288, 2311, 2338, 2349, 2394, 2396, 2433, 2505, 2526, 2554, 2675, 2677, 2771, 2909, 3023, 3069, 3087, 3110, 3156, 3191, 3211, 3212, 3237, 3315, 3376, 3419, 3512, 3525, 3557, 3564, 3627, 3631, 3643, 3663, 3746, 3754, 3769, 3773]. Galerkin [194, 414, 452–454, 538, 608, 634, 778, 857, 885, 887, 1009, 1046, 1079, 1096, 1222, 1254, 1268, 1273, 1278, 1300, 1306, 1324, 1378, 1475, 1498, 1510, 1531, 1540, 1572, 1592, 1621, 1624, 1677, 1715, 1757, 1771, 1792, 1816, 1830, 1856, 1879, 1935, 1936, 1957, 1974, 1977, 1999, 2034, 2062, 2073, 2081, 2120, 2135, 2137, 2138, 2166, 2168, 2172, 2197, 2251, 2295, 2308, 2369, 2397, 2402, 2418, 2477, 2491, 2527, 2575, 2583, 2600, 2634, 2700, 2709, 2711, 2761, 2804, 2848, 2898, 2914, 2946, 2957, 3024, 3026, 3077, 3079, 3097, 3238, 3260, 3267, 3281, 3297, 3375, 3396, 3411, 3432, 3442, 3478, 3482, 3549, 3556, 3579, 3634, 3704, 3787, 3793]. Galerkin [294, 302, 341, 344, 450, 456, 500, 603, 607, 648, 661, 733, 779, 780, 786, 788, 861, 894, 897, 904, 906, 930, 981, 1013, 1029, 1070, 1091, 1108, 1134, 1158, 1197, 1281, 1303, 1305, 1319, 1388, 1410, 1442, 1473, 1491, 1588, 1633, 1658, 1673, 1744, 1783, 1797, 1810, 1826, 1887, 1943, 1945, 1947, 1948, 1959, 2012, 2122, 2140, 2152, 2161, 2198, 2254, 2255, 2257, 2269, 2273, 2298, 2318, 2357, 2376, 2401, 2413, 2416, 2419, 2426, 2518, 2529, 2535, 2539, 2560, 2605, 2616, 2625, 2639, 2718, 2811, 2842, 2879, 2892, 2897, 2962, 3086, 3124, 3163, 3201, 3275, 3441, 3443, 3509, 3511, 3540, 3550, 3570, 3572, 3708, 3711, 3740]. Galerkin [287, 457, 732, 758, 766, 781, 1023, 1247, 1751, 1950, 2145, 2218, 2322, 2519, 2702, 2707, 2731, 2759, 2799, 2903, 2999, 3168, 3276, 3308, 3380, 3417, 3428, 3471, 3597]. Galerkin-Characteristics [3212]. Galerkin-Finite [3396]. Galerkin/Arbitrary [2122]. Galerkin/Finite [1324]. Galerkin/Finite-Element [1324]. Game [3161]. Games [1923, 2841]. GAN [3374, 3470]. Gap [351, 2220, 3610]. Gas [120, 296, 298, 347, 667, 1133, 2212, 2843, 3445, 3607]. Gas-Kinetic [2843]. Gasdynamics [2600, 2751]. Gauge [1226, 2423, 3626, 3763]. Gauge-Invariant [2423]. Gauss [258, 369, 528, 1051, 1142, 1185, 1280, 1691, 1834, 2081, 2337, 3463, 3585, 3639, 3651, 3808]. Gauss-Type [1280]. Gaussian [1575, 2076, 2099, 2121, 2325, 2627, 3487]. Gegenbauer [415]. Gene [1005]. General [279, 299, 479, 588, 1166, 1192, 1214, 1221, 1322, 1471, 1659, 1696, 1778, 2108, 2119, 2139, 2144, 2284, 2355, 2465, 2519, 2524, 2609, 2751, 2837, 2917, 3264, 3267, 3290, 3298, 3314, 3378, 3481, 3517, 3533, 3687, 3704, 3727]. Generalised [1024, 3159, 3360]. Generalization [356, 677, 1276, 2056, 2214]. Generalizations [3430]. Generalized [48, 59, 87, 99, 100, 110, 122, 212, 354, 507, 535, 551, 704, 722, 835, 954, 990, 1147, 1149, 1205, 1206, 1303, 1389, 1488, 1551, 1713, 1754, 1816, 1851, 1857, 1875, 1989, 2005, 2078, 2322, 2387, 2427, 2453, 2459, 2497, 2513, 2532, 2537, 2551, 2562, 2597, 2635, 2637, 2717, 2739, 2832, 2904, 2908, 3011, 3020, 3038, 3097, 3100, 3132, 3175, 3286, 3298, 3303, 3309, 3339, 3349, 3352, 3378, 3384, 3436, 3438, 3463, 3505, 3524, 3525, 3544, 3548, 3557, 3604, 3613, 3651, 3803, 3806]. Generate [1946]. Generated [1323]. Generating [543, 1820, 2065]. Generation [105, 225, 469, 2701, 3689]. Generative [3451, 3571, 3623]. Generic [932, 1256, 1257, 1619]. Genetic [359]. Genuinely [1138, 2899]. Genus [1255]. Genus- [1255]. Genvar [1606]. Geodesic [1761]. Geodesics [3689]. Geomechanics [1968]. Geometric [210, 358, 481, 879, 1127, 1373, 1457, 1496, 1810, 2264, 2265, 2529, 2765, 3103, 3169, 3176]. Geometrical [375, 955, 1655, 2913, 3058, 3387]. Geometrically [2369, 2653]. Geometries [202, 305, 588, 702, 1122, 1140, 1321, 1345, 2036, 2284, 2615, 2881, 3034, 3511].

Geometry [646, 2352]. Geophysics [3474]. Geostrophic [980]. GePUP [1619]. Get [2417, 3672]. Ghost [688, 694, 796, 3276, 3657]. Ghost-Fluid [3657]. Gibb [429]. Gibbs [567, 570]. Ginzburg [1071, 1315, 1414, 2423, 2612, 3341, 3553, 3626]. Given [2155]. Glerkin [1369]. Global [362, 910, 941, 1255, 1551, 2095, 2107, 2149, 2258, 2275, 2331, 2425, 2476, 2532, 2664, 2778, 2830, 3135, 3481, 3701, 3714, 3767]. Global-in-Time [2778, 3481]. Globalization [2987, 3223, 3517, 3523, 3566]. Globally [666, 1284, 1872, 1917, 2171, 2244, 2578]. Globally-Hyperbolically-Closed [1284]. GMRES [2285, 2286, 2726, 3037]. GMRES-Type [3037]. Goal [1050, 1626, 1835, 2453, 2587, 3394]. Goal-Oriented [1050, 1626, 1835, 2453, 2587, 3394]. Godunov [348, 662, 1205, 3146, 3677]. Godunov-Like [3677]. Golden [2797]. Golub [1493]. Good [1941, 2058, 3496]. Gordon [1100, 2438, 2811, 2944, 3207, 3419, 3665, 3682]. Gordon-Schrödinger [2944]. Governed [740, 754, 795, 1461, 1553, 1588, 1898, 2199, 2318, 2367, 2837, 3411, 3670, 3705]. Governing [382, 1718]. GPU [1183, 1369, 1384, 3184]. GPU-Accelerated [1369]. GPU-Implementation [1384]. GPUs [944, 1276]. Grad [517, 1556, 2563, 2859, 3259, 3804]. Grad-div [1556, 2859]. Graded [613, 1136, 2248, 2711, 2840, 2902, 3536, 3562]. Gradient [111, 211, 228, 236, 276, 295, 509, 1034, 1200, 1276, 1318, 1391, 1437, 1499, 1520, 1625, 1648, 1800, 1861, 1889, 1980, 1992, 2055, 2223, 2299, 2424, 2507, 2569, 2736, 2749, 2766, 2779, 2858, 2882, 2895, 2896, 2915, 2923, 2984, 2985, 2989, 3003, 3009, 3014, 3108, 3149, 3176, 3196, 3245, 3264, 3286, 3300, 3358, 3372, 3388, 3418, 3427, 3456, 3485, 3487, 3515, 3554, 3556, 3561, 3610, 3686]. Gradient-Augmented [1520]. Gradient-Enhanced [3487]. Gradient-Preserving [1800]. Gradient-Robust [2749]. Gradient-Type [2569]. Gradient/Semismooth [3610]. Gradients [1929, 2708, 2788, 3440, 3599]. Grafts [3387]. Grain [40, 674, 1394, 2981]. Graining [713]. Gramicidin [293]. Grand [574]. Granular [120, 2737, 3601]. Graph [3010, 3264]. Graph-Laplacians [3010]. Graphene [2984]. graphical [38]. Graphs [780, 1009]. Grating [2174, 2195]. Gratings [888]. Gravitation [3790]. Gravitational [1133, 1592, 2852]. Gravity [457, 1832]. Greedy [1223]. Green [458, 942, 1841, 2667, 3280]. Grey [3261]. Grid [105, 201, 225, 298, 340, 408, 410, 414, 416, 469, 512, 583, 606, 615, 906, 927, 930, 975, 979, 1069, 1098, 1155, 1159, 1178, 1204, 1308, 1310, 1317, 1348, 1577, 1582, 1666, 1683, 1741, 1824, 1878, 1954, 1972, 2027, 2088, 2199, 2208, 2662, 2800, 2838, 2879, 2880, 2916, 2973, 3066, 3175, 3227, 3229, 3403, 3567, 3738]. Grid-Based [414]. Grid-Insensitive [1824]. Gridding [1338]. Grids [262, 404, 613, 630, 730, 750, 956, 1027, 1136, 1240, 1244, 1332, 1336, 1419, 1424, 1481, 1492, 1495, 1558, 1574, 1664, 1674, 1675, 1705, 1711, 1736, 1745, 1809, 1824, 1858, 1884, 1900, 1942, 2016, 2028, 2048, 2550, 2553, 2559, 2644, 2679, 2711, 2862, 3068, 3082, 3119, 3448, 3541, 3586, 3725]. Grönwall [3102]. Grönwall-Type [3102]. Gross [1262, 3027]. Grossly [2695]. Ground [1420, 1910, 2485, 2668, 3027]. Group [2, 37, 55, 64, 104, 231, 236, 247, 2457, 2991, 3558]. Group-Based [231, 236].

Growing [469]. Growth [27, 380, 688, 711, 763, 1182, 1232, 1840, 2314, 2984, 3284]. GRP [3445]. GRPDA [3332]. GSVD [3384]. Guaranteed [909, 1441, 1620, 2074]. Guarantees [1985, 3210, 3435]. Guderley [1450]. guide [29]. Guidelines [796, 3431]. Guiding [1951, 2189, 2364]. Gyrokinetic [2021, 3103].

H2N2 [2589]. Haar [3757]. Haar-Like [3757]. Hadamard [2700]. Half [265, 1595, 2489, 2834, 3574, 3733]. Half-Quadratic [3733]. Hamilton [268, 387, 536, 579, 580, 591, 620, 622, 753, 885, 890, 926, 1025, 1132, 1159, 1465, 1561, 1852, 1923, 1935, 1959, 2098, 2295, 2427, 2585, 2617, 2666, 2715, 2838, 2841, 2844, 3060, 3208, 3256]. Hamiltonian [687, 1341, 1566, 1655, 1912, 2148, 2449, 2631, 2870, 3180, 3359, 3488]. Hamiltonians [536, 2414]. Hammerstein [1634]. Hand [2285, 2286]. Handling [358, 546]. Hankel [1654, 2412]. Hard [2353, 3546]. Harmonic [185, 292, 551, 861, 891, 1015, 1255, 1590, 1641, 1659, 1837, 2051, 2253, 2433, 2842, 3140, 3303, 3676]. Harmonic-Generalized [551]. Harmonics [147, 2035]. Harten [2899]. HDG [876, 1016, 1035, 1052, 1152, 1264, 1306, 1514, 1529, 1569, 1579, 1591, 1728, 1735, 1774, 1927, 2054, 2142, 2159, 2246, 2247, 2250, 2256, 2374, 2383, 2403, 2420, 2626, 2682, 2749, 2763, 2781, 2798, 3070, 3401, 3410, 3415, 3429, 3691]. HDG-NEFEM [2256]. Heat [200, 205, 230, 236, 613, 844, 854, 1347, 1387, 1768, 1844, 1865, 2083, 2269, 2352, 2451, 2590, 2615, 2786, 2794, 2957, 3068, 3209]. Heated [403]. Heating [1251]. Heaviside [1865]. Heavy [109, 3477, 3606]. Heavy-ball [3477]. Heavy-Ball-Based [3606]. heavy-ion [109]. held [2454]. Hele [851, 1560]. Hellan [2129, 3328]. Helmholtz [341, 868, 974, 1080, 1095, 1144, 1221, 1298, 1462, 1484, 1576, 1609, 1714, 2320, 2321, 2380, 2412, 2461, 2534, 2603, 2604, 2781, 2954, 3156, 3336, 3370, 3395, 3504, 3518, 3534, 3624, 3744, 3750]. Hemivariational [2714, 2932, 3067, 3167, 3509, 3595]. Hemodynamics [1056, 3387]. Hermite [284, 644, 758, 982, 1206, 1404, 1488, 1539, 2105, 2398, 2585, 2675, 2822, 3060, 3157, 3265, 3474, 3506, 3701, 3767, 3794]. Hermite-Based [2822]. Hermitian [595, 972, 1204]. Herrmann [2129, 3328]. Herrmann-Johnson [2129]. Herschel [1102]. Hessian [2312, 2825, 3241]. Heteroepitaxy [714]. Heterogeneities [2935]. Heterogeneous [1321, 1965, 2250, 2786, 2917, 3213, 3304, 3338, 3360, 3691]. Heuristic [2787]. Hexagonal [1940]. Hexahedral [2940, 3347, 3765]. HHO [3182]. Hidden [3129, 3458]. Hidden-Layer [3458]. Hidden-Memory [3129]. Hierarchial [1109]. Hierarchical [111, 336, 497, 656, 825, 911, 998, 1308, 1846, 2016, 2506, 2691, 2912, 3090, 3125, 3385, 3522, 3593, 3745, 3757]. Hierarchically [1218]. Hierarchy [665]. High [202, 244, 268, 282, 326, 330, 373, 378, 422, 444, 480, 483, 525, 529, 552, 565, 632, 739, 748, 762, 816, 865, 867, 869, 923, 995, 1018, 1054, 1077, 1122, 1170, 1191, 1205, 1236, 1323, 1332, 1348, 1382, 1446, 1496, 1519, 1525, 1546, 1576, 1645, 1665, 1687, 1689, 1695, 1707, 1719, 1750, 1760, 1820, 1843, 1890, 1899, 1917, 1919, 1925, 2023, 2133, 2136, 2173, 2192, 2232, 2241, 2288, 2335, 2364, 2375, 2391, 2443, 2475, 2530, 2584, 2593, 2640, 2652, 2728, 2771, 2796, 2809, 2863, 2978, 3030, 3033, 3057, 3062, 3199, 3225, 3228, 3237, 3313, 3397, 3527, 3566, 3567, 3617, 3627, 3682, 3790]. High [188, 193, 251, 275, 311, 320, 357, 366, 459, 486, 501, 502, 512, 519, 537, 583, 670, 685, 730, 883, 987, 1012, 1060, 1080, 1095, 1098, 1140, 1143, 1183, 1222, 1228, 1276, 1284, 1336, 1368, 1458, 1460, 1464, 1465, 1531, 1558, 1562, 1587, 1609, 1646, 1667, 1711, 1735, 1759, 1806, 1813, 1817, 1821, 1911, 1938, 1965, 1974, 1981, 2028, 2080, 2112, 2134, 2156, 2248, 2250, 2251, 2284, 2330, 2356, 2368, 2412, 2437, 2446, 2447, 2513, 2528, 2542, 2606, 2608, 2693, 2719, 2777, 2831, 2833, 2839, 2843, 2850, 2862, 2893, 2929, 2933, 2997, 3029, 3044, 3096, 3101, 3215, 3298, 3336, 3398, 3472, 3498, 3635, 3638, 3669, 3690]. High [44, 58, 75, 142, 146, 160, 232, 240, 306, 316, 347, 348, 388, 407, 427, 429, 432, 471, 508, 544, 547, 548, 550, 579, 582, 607, 648, 656, 681, 736, 764, 820, 917, 981, 1040, 1093, 1133, 1166, 1184, 1208, 1233, 1291, 1494, 1574, 1605, 1615, 1616, 1632, 1666, 1732, 1799, 1884, 1946, 1976, 1982, 2051, 2064, 2121, 2140, 2141, 2189, 2208, 2230, 2352, 2374, 2440, 2492, 2567, 2573, 2588, 2604, 2799, 2800, 2812, 2814, 2819, 2873, 3032, 3054, 3137, 3175, 3265, 3275, 3293, 3300, 3308, 3381, 3384, 3434, 3526, 3529, 3548, 3552, 3587, 3590, 3623, 3708, 3727, 3749, 3781, 3794, 3797]. High-0rder [2391]. High-Accuracy [2997]. High-Contrast [1911, 3175]. High-Dimensional [2375, 2863, 3623, 3669, 3708]. High-fidelity [2230]. High-Frequency [2284]. High-Index [3137]. High-Order [244, 282, 316, 320, 326, 330, 471, 519, 525, 552, 565, 583, 607, 632, 648, 730, 762, 764, 869, 917, 923, 1012, 1018, 1140, 1166, 1191, 1205, 1222, 1284, 1291, 1323, 1332, 1336, 1348, 1368, 1382, 1446, 1460, 1464, 1496, 1574, 1587, 1609, 1632, 1687, 1707, 1750, 1760, 1820, 1821, 1843, 1890, 1982, 2023, 2051, 2121, 2134, 2140, 2141, 2156, 2232, 2335, 2352, 2437, 2446, 2447, 2475, 2528, 2530, 2584, 2606, 2608, 2640, 2693, 2771, 2809, 2831, 2843, 2873, 2929, 2933, 3029, 3032, 3033, 3044, 3096, 3101, 3199, 3225, 3228, 3237, 3293, 3298, 3313, 3397, 3472, 3498, 3529, 3567, 3627, 3690, 3781, 3794]. High-Order [44, 160, 188, 193, 432, 444, 501, 502, 548, 681, 736, 1208, 1233, 1615, 1732, 1976, 2440, 2588, 2777, 2799, 2812, 2819, 3054, 3057, 3215, 3308, 3381, 3384, 3434, 3526, 3552, 3587, 3590, 3727, 3797]. High-Precision [1183]. High-Resolution [995, 1946, 2593, 3635]. High-Speed [232, 1170]. Higher [302, 315, 343, 528, 638, 663, 679, 821, 932, 1043, 1071, 1079, 1118, 1142, 1430, 1613, 1660, 1791, 1874, 1908, 2068, 2082, 2092, 2122, 3123, 3166, 3457, 3529, 3620, 3666, 3750]. Higher-Order [528, 638, 821, 1043, 1071, 1079, 1118, 1660, 2082, 2122, 3457, 3620, 3750]. Highest [1515, 1743]. Highly [496, 744, 828, 1068, 1082, 1117, 1221, 1383, 1426, 1667, 1722, 1822, 1998, 2314, 2532, 2851, 3041, 3071, 3444, 3486, 3488, 3801]. Hilbert [2672, 2910, 2941, 3392]. Hilliard [803, 851, 1054, 1241, 1418, 1560, 1601, 1724, 1748, 1945, 2018, 2178, 2224, 2226– 2228, 2362, 2569, 2614, 2656, 2696, 2755, 2766, 2931, 3040, 3185, 3195, 3206, 3382, 3439, 3460, 3588, 3664, 3671, 3702, 3719]. Hilliard-Darcy [2227]. HiMod [2108]. Hirota [2203]. History [66, 422, 3595]. History-Dependent [3595]. HJB [1736, 3543]. HJB-RBF [3543]. HLLC [625, 3717]. HOC [903]. Hodge [1462, 1918, 2131]. Hodgkin [2638]. Hohenberg [2474, 2886, 3308, 3780]. Hole [1183]. Holes [291]. Hollow [553]. Holm [2572]. Holomorphic [1004, 1020, 1297]. Homogeneous [82, 166, 230, 235, 260, 701, 873, 1548, 1724]. Homogenization [353, 354, 560, 1835, 2215]. Homology [3609]. Homotopy [1725, 1785, 2009, 2512, 2813, 2896, 2996, 3436, 3505]. Honor [1916, 2231]. Hood [355, 1486, 3278, 3427]. Hopf [2427]. Horizontal [248, 959, 1557, 1929]. hp [1846, 1870, 2034, 2559, 2913, 3058, 3250, 3459, 3657]. hp-Adaptive [2034, 2559, 3657]. hp}-discontinuous [3250]. hp-FEM [1846]. hp}-mesh [3459]. hp}-method [2913, 3058]. hp-version [1870]. HPX [2415]. Huggins [3185]. Hull [2030]. Humidity [2591]. Hunter [3768]. Hurricane [906, 1062]. Hurwicz [2916, 3165, 3565]. Huxley [2638, 2908]. Huygens [2906, 2911]. HWENO [1645, 2032, 2153, 3596]. Hybird [3183]. Hybrid [86, 263, 271, 460, 699, 756, 821, 884, 939, 950, 1014, 1033, 1150, 1439, 1448, 1478, 1681, 1928, 1929, 1979, 1981, 2023, 2066, 2167, 2198, 2200, 2213, 2232, 2236, 2243, 2445, 2472, 2519, 2584, 2599, 2617, 2622, 2651, 2809, 2849, 2943, 2956, 2969, 2980, 3042, 3119, 3164, 3225, 3587, 3596, 3690]. Hybridizable [771, 850, 974, 1245, 1550, 1569, 1865, 2140, 2161, 2196, 2268, 2269, 2709, 2771, 3124, 3432, 3667, 3787]. Hybridization [1729, 2164]. Hybridized [1306, 1473, 1633, 1658, 2168, 2242, 2255, 2477, 2815, 3205, 3663]. Hydraulic [3046]. Hydrodynamic [434, 936, 1678, 1718, 2621, 3517]. Hydrodynamically [3428]. Hydrodynamics [642, 643, 646, 1259, 2694, 2712, 3439, 3727]. Hydrogen [401, 413, 423, 745]. hydromagnetic [113]. Hydrostatic [1304, 1695, 2339, 2417, 2591, 2609, 2737, 3601, 3755]. Hyper [3319]. Hyper-Reduction [3319]. Hyperbolic [215, 221, 259, 274, 316, 374, 382, 427, 454, 484, 511, 528, 538, 565, 592, 617, 649, 653, 655, 659, 729, 748, 783, 839, 907, 940, 944, 948, 1039, 1128, 1194, 1228, 1234, 1244, 1291, 1329, 1388, 1404, 1408, 1415, 1438, 1443–1445, 1448, 1449, 1458, 1530, 1543, 1597, 1621, 1664, 1693, 1705, 1856, 1886, 1917, 1928, 1957, 2068, 2153, 2172, 2200, 2251, 2308, 2331, 2356, 2359, 2368, 2394, 2505, 2536, 2559, 2560, 2567, 2583, 2609, 2648, 2721, 2796, 2810, 2938, 2960, 2970, 3018, 3039, 3062, 3074, 3085, 3096, 3184, 3215, 3289, 3298, 3346, 3373, 3376, 3398, 3662, 3754, 3770, 3782, 3794]. Hyperbolic [18, 24, 128, 269, 548, 687, 1023, 1799, 1863, 1952, 2544, 2625, 2800, 3086, 3168, 3334, 3441, 3762]. Hyperbolic/Parabolic [1597]. Hyperbolically [1284]. Hyperbolicity [937, 3356]. Hyperconcentrated [955]. Hypercube [78, 82, 169, 252]. Hypercubes [217]. Hypergraph [3269, 3761]. Hypergraphs [2029]. Hyperparameters [3394]. Hypersingular [1112, 1292]. Hypersurfaces [724, 2666]. Hyperviscous [142]. Hypothesis [1036]. HyROMnet [3319].

I/O [475]. IBM [152, 169]. iDCA [3778]. Idea [1941, 3502]. Ideal [550, 1869, 1979, 2244, 2508, 2644, 3148, 3228, 3718, 3752]. Idealized [2351]. Identification [510, 521, 1394, 1636, 1962]. Identities [513]. IgA [2920]. IgA-DG [2920]. II [47, 59, 153, 236, 374, 582, 630, 671, 729, 759, 933, 1031, 1287, 1539, 1550, 2265, 3024, 3226, 3685, 3764]. IIF [2927]. IIF-WENO [2927]. III [404, 693, 758, 1048, 3356]. IIPG [812]. Ill [828, 1188, 1389, 1432, 2316, 2864, 3037, 3134, 3371, 3487, 3739]. Ill-Conditioning [3487]. Ill-Posed [828, 1188, 1432, 2316, 2864, 3037, 3371, 3739]. ILU [178]. Image [26, 195, 310, 376, 397, 650, 727, 801, 814, 826, 833, 881, 986, 1032, 1084, 1120, 1121, 1129, 1130, 1178, 1212, 1215, 1252, 1266, 1290, 1360, 1364, 1391, 1395, 1470, 1483, 1499, 1523, 1570, 1589, 1628, 1660, 1685, 1700, 1703, 1721, 1782, 1786, 1850, 1864, 2006, 2144, 2214, 2280, 2297, 2309, 2362, 2548, 2578, 2595, 2776, 2817, 2853, 2976, 3009, 3035, 3047, 3154, 3187, 3233, 3240, 3257, 3271, 3311, 3422, 3522, 3553, 3560, 3662, 3720, 3741, 3808, 3811]. Images [72, 131, 146, 384, 494, 674, 1394, 1851, 1896, 1897, 2315, 3799]. Imaginary [3258]. Imaging [1000, 2507, 2857]. IMEX [1503, 1921, 2543, 3032, 3075, 3181, 3285, 3317, 3718]. Immediate [3026]. Immersed [383, 1337, 1340, 1431, 1498, 1610, 1920, 1978, 2090, 2334, 2345, 2428, 2598, 2648, 2784, 2880, 3141, 3274, 3379, 3710]. Immiscible [849, 2596]. Immobile [3306, 3562, 3797]. Impact [692]. Impedance [2663, 3036, 3733]. Imperfect [2143]. Implementable [1583]. Implementation [21, 164, 354, 409, 479, 585, 645, 752, 887, 1046, 1053, 1384, 1746, 1854, 2002, 2128, 2355, 2586, 2998, 3110, 3145, 3184, 3321, 3651, 3692, 3779]. Implications [2296]. Implicit [71, 117, 118, 126, 215, 261, 279, 325, 392, 484, 491, 533, 628, 680, 690, 742, 968, 1014, 1130, 1175, 1322, 1328, 1381, 1387, 1510, 1568, 1607, 1612, 1614, 1652, 1665, 1667, 1668, 1702, 1715, 1748, 1776, 1778, 1827, 1855, 1882, 1925, 1931, 1932, 1943, 1945, 1995, 2042, 2124, 2127, 2179, 2203, 2211, 2212, 2218, 2221, 2223, 2226, 2257, 2282, 2386, 2422, 2423, 2438, 2439, 2490, 2500, 2518, 2539, 2568, 2572, 2600, 2606, 2675, 2686, 2693, 2694, 2766, 2979, 2980, 3044, 3214, 3219, 3253, 3343, 3353, 3373, 3388, 3440, 3527, 3529, 3550, 3640, 3649, 3718, 3736, 3751]. Implicit-Explicit [484, 628, 1014]. Implicit/Explicit [2979]. Implicit/Explicit-SAV [2979]. Implicitly [3103]. Important [5, 11, 189, 481, 1256]. Imposing [224]. Imposition [74, 819]. Improve [2039]. Improved [360, 676, 737, 1079, 1332, 1794, 1863, 1895, 1913, 2103, 2158, 2208, 2452, 2506, 2566, 2851, 3112, 3155, 3189, 3282, 3324, 3565, 3665]. Improvement [588, 925, 2417]. Improvements [1745]. Improving [215]. Impulse [1575]. Incident [246]. Including [3573]. Inclusion [2780, 3133, 3605]. Inclusions [657, 2875, 3422]. Incomplete [136, 1302, 3184, 3202]. Incompressibile [895]. Incompressible [4, 77, 86, 87, 127, 143, 156, 170, 224, 231, 270, 272, 281, 320, 327, 343, 345, 355, 368, 371, 430, 451, 458, 492, 493, 532, 549, 576, 615, 728, 772, 776, 792, 841, 903, 933, 988, 989, 1048, 1056, 1125, 1174, 1180, 1236, 1310, 1349, 1370, 1400, 1428, 1464, 1482, 1500, 1511, 1519, 1537, 1619, 1742, 1772, 1779, 1788, 1814, 1843, 1904, 1951, 2023, 2048, 2064, 2122, 2189, 2257, 2363, 2364, 2381, 2506, 2580, 2581, 2614, 2763, 2879, 2889, 2916, 2933, 3077, 3165, 3182, 3232, 3255, 3290, 3399, 3401, 3427, 3437, 3494, 3495, 3513, 3515, 3663, 3680, 3796]. Inconsistent [2607]. Incorporating [2981]. Increasing [735, 887, 1046, 1079, 1254, 1453, 1581, 1725, 1940, 2201, 2491]. Incurred [2925].

Indefinite [595, 1108, 1221, 2921, 3409]. Indentation [1930]. Independent [17, 1911, 2413, 2787, 2852, 3302]. Index [1246, 1710, 2549, 3137, 3454]. Indication [3283]. Indicator [206, 623, 2385, 2566, 2571, 2599, 3615]. Indicators [147, 1033, 1839]. Indices [2371]. Indirect [1480, 1808]. Induced [801, 2217, 3234, 3758]. Inductionless [2860, 3048, 3520]. Inequalities [273, 726, 2390, 2417, 2672, 2790, 2941, 3049, 3509, 3616, 3672]. Inequality [1087, 2431, 2714, 2882, 2910, 2932, 2936, 2988, 3067, 3102, 3167, 3537, 3595, 3646, 3715]. Inertial [2723, 2780, 2910, 2941, 2988, 3049, 3286, 3605, 3693]. Inertial-Type [2723]. Inexact [2079, 2285, 2286, 2655, 3150, 3153, 3397, 3641, 3808]. Inextensible [1431]. Inf [1556, 1971]. Inf-sup [1556, 1971]. Infeasible [2327, 2722, 2871, 3784]. Inference [2869, 3104, 3623]. inferior [35]. Infiltration [234]. Infinite [16, 35, 365, 672, 1139, 1468, 1649, 1909, 2008, 3378]. Infinite-Dimensional [1468]. infinities [1488]. Infinity [3724]. Inflow [904, 1861, 2669]. Inflow-Based [1861]. Influence [196, 1072, 2960, 3398]. Information [2093, 2117, 2989]. Informed [3239, 3241, 3251, 3453, 3642, 3706]. Inherent [1778]. Inhibiting [1926]. Inhibition [75]. Inhomogeneous [168, 187, 454, 914, 1093, 1346, 1376, 1623, 1720, 1749, 2648, 2924, 3698]. Initial [28, 95, 106, 198, 222, 317, 500, 503, 723, 762, 1051, 1173, 1934, 2034, 2125, 2408, 2531, 2642, 2740, 2767, 2777, 2890, 2965, 3086, 3380, 3575, 3735]. Initial-Boundary [28, 2740]. Initial-Value [222]. initial-valued [95]. Initial/Final [723]. Initial/Final-Value [723]. Initializing [752]. initio [335]. inlet [123]. Inner [462, 1668]. Inner-Iteration [1668]. Innovation [3208]. Inpainting [1120, 3553]. Input [1582]. Inputs [1621, 1921, 2396, 2455, 3168, 3577, 3708]. Insensitive [1824]. Insight [3807]. Insights [2081]. Inspired [2432]. Instabilities [936, 964, 2081]. Instability [101, 1845, 3717]. Integer [2038, 2451]. Integral [7, 41, 249, 260, 409, 563, 794, 948, 1029, 1091, 1112, 1189, 1219, 1293, 1484, 1517, 1545, 1572, 1634, 1709, 1790, 1808, 1854, 1956, 1973, 2017, 2058, 2061, 2069, 2085, 2100, 2166, 2190, 2277, 2320, 2321, 2332, 2352, 2384, 2434, 2440, 2459, 2483, 2565, 2636, 2877, 2918, 3242, 3429, 3449, 3482, 3627, 3652, 3726, 3803]. Integrals [278, 677, 1101, 1112, 1292, 2109, 2142, 2325]. Integrated [1102, 1674, 1675]. Integrating [146, 1646, 2502]. Integration [77, 83, 306, 481, 487, 528, 626, 735, 738, 1175, 1222, 1257, 1363, 1469, 1711, 1812, 1838, 1938, 2127, 2130, 2211, 2233, 2422, 2565, 2767, 2805, 3071, 3193, 3343, 3476, 3556, 3640, 3748]. Integrator [1429, 2777, 3084, 3496, 3665, 3716, 3764]. Integrators [789, 968, 1291, 1316, 2504, 2788, 2826, 2951, 3044, 3180, 3207, 3320, 3563]. Integro [894, 1104, 1173, 1461, 1491, 1618, 1680, 1870, 2005, 2101, 2124, 2344, 3288, 3549]. Integro-Differential [1104, 1173, 1491, 1618, 1680, 1870, 2005]. Intel [169]. Intense [1446]. Interacting [401, 413, 423, 1414, 1431, 2658, 2856]. Interaction [23, 207, 522, 694, 878, 898, 994, 1061, 1859, 2014, 2122, 2131, 3001, 3063, 3353]. Interactions [388, 2650, 3213, 3729]. Interconnected [3090]. Interface [196, 383, 398, 676, 690, 737, 788, 872, 956, 1209, 1340, 1349, 1431, 1582, 1599, 1698, 1735, 1911, 1920, 1978, 2011, 2070, 2090, 2143, 2168, 2195, 2334, 2345, 2458, 2463, 2528, 2626, 2643, 2665, 2784, 2835, 2880, 2966, 2994, 3041, 3062, 3101, 3141, 3163, 3274, 3291, 3295, 3410, 3514, 3539, 3545, 3634, 3637, 3658, 3710, 3737]. Interface/Boundary [3410]. Interface/Boundary-Unfitted [3410]. Interfaces [171, 391, 616, 994, 1086, 1098, 1380, 1387, 1666, 1763, 2208, 2893, 3043, 3094, 3101, 3279, 3313, 3337, 3383, 3703, 3776]. Interfacial [1600]. Interior [440, 608, 732, 767, 774, 797, 808, 839, 916, 1113, 1211, 1237, 1267, 1576, 1867, 2022, 2086, 2167, 2197, 2327, 2433, 2677, 2722, 2871, 2872, 2945, 3073, 3196, 3211, 3433, 3579]. Interior-Point [2327, 2871, 3433]. Intermediate [64, 943]. intermittency [49, 166]. Intermittent [2991]. Internal [318, 2594]. Internet [2895]. Interplay [3431]. Interpolants [567]. Interpolated [1337, 3201]. Interpolation [79, 119, 148, 198, 225, 253, 356, 391, 863, 901, 1204, 1280, 1581, 1902, 1944, 1953, 1956, 1964, 2016, 2021, 2109, 2173, 2194, 2274, 2296, 2546, 2554, 2589, 2629, 2744, 2765, 2792, 2977, 3204, 3299, 3377, 3405, 3459]. Interpolative [2691]. Interpolatory [1101, 2383]. Interval [35, 176, 365, 2759]. Intervals [16, 2008]. Intrinsic [3451, 3571]. Intrinsically [811]. Introduced [887]. Introduction [375, 556]. Introductory [1]. Intrusive [993, 1530, 2278, 2792, 3487]. Invariant [1026, 1427, 2217, 2423, 2438, 3124, 3152, 3364]. Invariants [884, 1496, 1974, 2237]. Inverse [494, 1049, 1393, 1490, 1608, 1671, 1672, 1759, 2059, 2146, 2334, 2532, 2590, 2663, 2810, 2939, 3095, 3134, 3149, 3161, 3392, 3446, 3465, 3534, 3568, 3579, 3606, 3623, 3792]. Inverses [2961]. Inversion [199, 1265, 1737, 1987]. Inverting [3202]. Investigation [330, 1096, 1505, 1905, 2487, 3233]. Investigations [315, 1381, 3619]. Inviscid [458, 634, 1961, 2513, 2591]. Invisibility [2742]. Involved [1777, 2325]. Involving [260, 1266, 1509, 1516, 2748, 2893, 3063, 3315]. Ion [109, 1084, 2746]. Ion-Channel [1084]. Ionic [801]. Ionization [859, 1446]. IP [1639]. IPDG [1762]. IPG [3059]. IPM [3159]. Irreducible [3153, 3500]. Irregular [613, 762, 796, 863, 1333, 1518, 2180, 2440, 2458, 2499, 2893, 3369, 3472]. Irregularly [920]. Irreversible [3688]. Isentropic [1451, 1789, 2543]. Isogeometric [1810, 2574, 3683]. Isospectral [2047]. Isothermal [597, 953, 2621, 2755]. Isotropic [179, 235, 658, 1276, 1279, 1341, 1427, 2252, 2651, 3803]. Isotropy [663]. Isoviscous [1992]. Issue [1000, 1916, 2231, 2448, 2488]. Issued [2685]. Issues [716, 962, 2340, 3057]. Iterated [1029, 1219, 1242, 3299]. Iteration [411, 727, 892, 1049, 1118, 1135, 1242, 1277, 1668, 1684, 1724, 1742, 2117, 2490, 2581, 2871, 3172, 3336, 3391, 3450, 3521]. Iteration-Free [2490]. Iterations [2661, 3081]. Iterative [106, 183, 218, 257, 554, 712, 766, 853, 1007, 1037, 1092, 1252, 1400, 1489, 1595, 1779, 1882, 2017, 2071, 2097, 2115, 2353, 2363, 2443, 2460, 2534, 2659, 2660, 2668, 2706, 2762, 2860, 2941, 2995, 3050, 3072, 3153, 3165, 3207, 3214, 3560, 3742, 3792]. IV [371].

Jackson [3628]. Jacobi [268, 387, 467, 535, 536, 579, 580, 591, 620, 622, 753, 835, 885, 890, 926, 1025, 1132, 1149, 1159, 1256, 1280, 1465, 1561, 1709, 1852, 1923, 1935, 1959, 2049, 2079, 2098, 2295, 2337, 2384, 2427, 2585, 2617, 2666, 2715, 2834, 2838, 2841, 2844, 3060, 3208, 3256, 3303, 3352]. Jacobi-Bellman [2666, 2715]. Jacobian [487, 1564, 1668, 3074, 3277]. Jacobian-Free [487, 1668, 3074]. JCP [3499, 3508]. Jet [3105]. Jets [431, 459, 670, 1123, 1722]. Johnson [2129, 3328]. Joint [2260, 2315, 2841, 3522, 3593]. Joseph [2994, 3046]. Joule [1251]. JSC [3499, 3508]. Jump [755, 1037, 1063, 1503, 2083, 2102, 2323, 2634, 2639]. Jump-Diffusion [1503]. Jumps [1708, 2610, 2848, 3731]. Junctions [1270]. Just [3451, 3571]. justification [47].

Kac [1375, 1686, 1997, 2136, 2642, 2902]. Kahan [1493]. Kahan-Type [1493]. Kansa [1513, 1831]. Kansa-Radial [1513]. Kantorovich [2094, 2561]. Kármán [2454]. Kawahara [661]. Kawahara-Type [661]. Kawasaki [2439]. KdV [338, 482, 1374, 2198, 2731]. KdV-Type [2731]. Keller [773, 1105, 1936, 2311, 2445, 2686, 3625]. Kelvin [660]. Kemeny [873]. Kernel [1007, 1191, 1359, 1383, 1709, 1890, 2320, 2321, 2337, 2440, 2464, 3116, 3302, 3354, 3556]. Kernel-Based [2464, 3116, 3354]. Kernel-Free [2320, 2321, 2440]. Kernel-Independent [3302]. Kernels [695, 1837, 1870, 2069, 2201, 2340, 2434, 3486]. Kerr [1183]. Khokhlov [1922]. Kind [409, 1219, 2069, 2283, 2332, 2384, 2390, 2434, 2565, 3449, 3486]. Kinematic [121, 936]. Kinetic [163, 214, 216, 424, 572, 628, 667, 697, 714, 941, 1138, 1198, 1208, 1492, 1547, 1796, 1834, 1903, 1931, 2130, 2211, 2272, 2291, 2558, 2725, 2789, 2804, 2843, 2956, 3024, 3442, 3607, 3759]. kinetic-type [163]. Kinetics [39, 1250]. Kirchhoff [1729, 2129, 2268, 3067, 3549, 3725]. Klein [2811, 2944, 3207, 3419, 3665, 3682]. Koebe [1671, 1672]. Kohn [1151, 2346]. Kolmogorov [2841, 2887, 3357]. Koornwinder [3097]. Korn [2882]. Korteweg [1633, 1927, 1974, 3186, 3359]. Kou [1503]. KPP [1154]. Krylov [196, 487, 645, 752, 927, 1711, 1938, 2095, 2326, 2925, 3038, 3425, 3651, 3766]. Kullback [1289, 2807]. Kuramoto [101, 970]. Kutta [222, 287, 288, 361, 462, 483– 485, 648, 696, 749, 758, 789, 860, 904, 1181, 1220, 1300, 1377, 1392, 1403, 1407, 1522, 1614, 1615, 1778, 1828, 1925, 1931, 1932, 2011, 2072, 2123, 2127, 2139, 2155, 2220, 2386, 2411, 2502, 2608, 2612, 2625, 2631, 2826, 2929, 2962, 3026, 3074, 3086, 3096, 3160, 3248, 3253, 3284, 3408, 3527, 3535, 3611, 3754, 3760, 3782]. Kuznetzov [1922].

L1 [3249, 3507, 3549]. L1-Galerkin [3549]. L2 [3481]. L2-1 [3481]. Labeling [3560, 3674]. Ladyzenskaja [1176]. Lag [2786]. Lagging [2083]. Lagrange [355, 633, 725, 867, 1763, 1988, 2413, 3029, 3191]. Lagrangian [168, 255, 346, 492, 743, 761, 789, 798, 920, 971, 986, 990, 1032, 1102, 1159, 1230, 1339, 1425, 1508, 1519, 1525, 1852, 1857, 1919, 1935, 1956, 2021, 2089, 2109, 2122, 2221, 2364, 2679, 2681, 2721, 2750, 2795, 2828, 2843, 2854, 2903, 2999, 3015, 3018, 3063, 3075, 3136, 3148, 3233, 3273, 3558, 3614, 3641, 3754, 3807]. Laguerre [551, 704, 805, 862, 1691, 2008]. Laminar [28, 134, 1337, 2002, 2580]. Laminates [385]. Lanczos [1341, 1580]. Lanczos-Type [1580]. Landau [1071, 1315, 1414, 1682, 2423, 2612, 3338, 3341, 3553, 3626]. Landau [?]Bai:2023:GLH. Landmark [1005, 1611]. Landmark-Matching [1611]. Landscape [963, 3137]. Lane [2501]. Laplace [358, 912, 1496, 1686, 1808, 1955, 2097, 2222, 2248, 2758, 2760, 3578, 3619, 3632]. Laplacian [199, 295, 640, 1049, 1470, 1490, 1944, 2056, 2183, 2304, 2357, 2420, 2514, 2743, 2847, 3002, 3045, 3269, 3624, 3724, 3772, 3803]. Laplacians [3010, 3315]. Large [13, 55, 63, 93, 101, 179, 218, 232, 320, 402, 431, 524, 539, 583, 847, 945, 949, 959, 963, 1092, 1188, 1333, 1563, 1580, 1611, 1654, 1714, 1746, 2095, 2316, 2693, 2781, 3031, 3111, 3134, 3175, 3243, 3280, 3303, 3352, 3357, 3370, 3425, 3453, 3457, 3475, 3558, 3628, 3700, 3754, 3759, 3789]. Large-Eddy [55, 232, 320, 524, 539, 847, 963]. Large-Scale [101, 1188, 1333, 1563, 1746, 3425, 3453, 3558, 3700]. Large-Strain [3243]. Large-Time [402]. Largest [3153]. Laser [284, 644, 646, 859]. Lasso [3558]. Last [574, 2878]. Last-passage [2878]. Latent [3360]. Lattice [129, 130, 204, 216, 255, 286, 299, 604, 663, 912, 970, 1043, 1071, 1100, 1106, 1172, 1177, 1259, 1315, 1326, 1330, 1337, 1696, 2615, 2647, 2703, 2793, 3188, 3803]. Lattice-Boltzmann [1177]. Lattices [3031]. Lavrentiev [1087]. Law [904, 1723, 3220, 3701, 3767, 3800]. Laws [174, 206, 267, 269, 274, 374, 404, 417, 427, 525, 538, 542, 555, 609, 617, 655, 659, 738, 907, 922, 939, 952, 981, 1024, 1085, 1094, 1099, 1128, 1194, 1228, 1244, 1291, 1305, 1329, 1404, 1407, 1438, 1444, 1445, 1448, 1458, 1464, 1533, 1539, 1621, 1664, 1693, 1715, 1745, 1795, 1799, 1855, 1886, 1928, 1952, 2016, 2032, 2040, 2105, 2153, 2200, 2221, 2331, 2356, 2368, 2394, 2444, 2476, 2494, 2530, 2536, 2559, 2560, 2571, 2970, 3000, 3018, 3062, 3074, 3085, 3096, 3107, 3126, 3214, 3215, 3298, 3334, 3346, 3363, 3441, 3478, 3615, 3677, 3759, 3762, 3794]. Lax [261, 603, 926, 928, 1444, 1445, 1472, 1759, 1792, 1799, 2536, 2720, 2899, 3442, 3785]. Layer [25, 99, 112, 331, 692, 806, 937, 943, 1103, 1156, 1227, 2213, 2339, 2686, 3012, 3458, 3518, 3746]. Layered [246, 1084, 1656, 1841, 2534, 2939]. Layers [15, 142, 227, 435, 568, 923, 1681, 2167, 2911]. LBGK [285]. LDA [1824]. LDG [451, 905, 1017, 1055, 1590, 1638, 3143]. LDG2 [905]. leading [13, 48]. Leak [521]. Leap [915]. Leap-Frog [915]. Leapfrog [1381, 2398, 3399]. Learn [3454]. Learnable [3311]. Learned [3733]. Learning [1700, 1960, 2006, 2379, 2410, 2479, 2783, 2869, 2887, 2915, 2947, 3036, 3152, 3157, 3177, 3188, 3190, 3218, 3239, 3245, 3264, 3267, 3292, 3319, 3333, 3340, 3356, 3357, 3360, 3385, 3394, 3446, 3458, 3477, 3497, 3522, 3568, 3593, 3642, 3689, 3807]. Learning-Based [2783, 3319]. Least [301, 321, 322, 420, 468, 531, 541, 725, 1086, 1185, 1552, 1668, 1836, 2181, 2283, 2342, 2452, 2657, 2825, 2977, 3125, 3150, 3176, 3277, 3412, 3451, 3533, 3571, 3730]. Least-Squared [2283]. Least-Squares [301, 321, 420, 468, 531, 541, 1086, 2181, 2342, 2452, 3277, 3533, 3730]. Least-Squares/Reduced-Basis [2452]. Least-Squares/Relaxation [2181]. Leer [2899]. Leffler [2184, 2532, 2782]. LeFloch [1713]. Legacy [478].

Legendre [150, 200, 258, 367, 567, 641, 707, 805, 815, 862, 1051, 1067, 1104, 1256, 1413, 1572, 1618, 1634, 1775, 1834, 2518, 3055, 3087, 3639]. Leibler [1289, 2807]. Lemarie [241]. Length [680]. Lens [2405]. Level [72, 175, 374, 377, 380, 384, 392, 394, 408, 416, 448, 497, 512, 589, 615, 618, 679– 684, 686, 688–690, 724, 760, 763, 828, 902, 916, 995, 1123, 1125, 1201, 1345, 1379, 1400, 1478, 1500, 1520, 1557, 1608, 1732, 1761, 1764, 1779, 1800, 1977, 2030, 2087, 2147, 2262, 2388, 2517, 2581, 2601, 2634, 2649, 2724, 2799, 2973, 3078, 3268, 3407, 3420, 3432, 3493, 3657]. Level-Set [690, 724, 902, 1345, 1732, 3268, 3420, 3657]. Levels [999]. Levenberg [2001, 2276]. Leverett [1455]. Lévy [1214]. Li [2746]. Li-Ion [2746]. Lid [524, 537]. Lid-Driven [524, 537]. Lie [2845]. Life [2231]. Lifshitz [402, 1682, 3338]. Lifting [1896]. Light [1987, 3280, 3720]. Lighthill [721]. Like [785, 1595, 3081, 3103, 3405, 3427, 3498, 3677, 3757]. Limit [683, 798, 820, 1011, 1550, 1829, 1833, 1931, 3680]. Limited [985, 2216, 3263, 3654]. Limited-Memory [3263]. Limiter [564, 1458, 1587, 2556, 2600, 2751, 3590, 3638, 3752, 3785]. Limiters [378, 758, 1160, 1228, 1531, 1616, 2291, 2457, 3620]. Limiting [550, 1651, 1693, 3026, 3096, 3253, 3467]. Line [265, 807, 911, 1888, 1956, 2085, 2109, 2401, 2409, 2834, 2910, 3186, 3372, 3561]. Line-Based [2401]. Line-Search [2910]. Line-SIAC [2409]. Linear [210, 257, 296, 361, 442, 451, 479, 528, 557, 600, 627, 635, 705, 969, 975, 1065, 1141, 1147, 1180, 1205, 1230, 1240, 1277, 1351, 1361, 1471, 1540, 1551, 1567, 1613, 1641, 1664, 1674, 1675, 1724, 1778, 1794, 1849, 1873, 1921, 1924, 1925, 1984, 2013, 2022, 2025, 2070, 2101, 2106, 2138, 2139, 2177, 2249, 2251, 2252, 2277, 2288, 2291, 2308, 2312, 2379, 2380, 2406, 2470, 2509, 2538, 2564, 2583, 2587, 2657, 2663, 2698, 2706, 2709, 2729, 2749, 2791, 2802, 2831, 2894, 2938, 3025, 3039, 3062, 3091, 3100, 3149, 3246, 3272, 3304, 3376, 3390, 3406, 3410, 3519, 3555, 3599, 3685, 3709, 3736, 3738, 3739, 3754, 3761, 3775, 3799]. Linear [15, 91, 95, 149, 309, 316, 454, 457, 506, 518, 550, 687, 736, 754, 820, 904, 921, 997, 1024, 1082, 1108, 1162, 1305, 1320, 1322, 1354, 1427, 1515, 1530, 1584, 1603, 1604, 1653, 1731, 1790, 1863, 1898, 1943, 1950, 1989, 2026, 2086, 2130, 2157, 2170, 2203, 2218, 2285, 2286, 2298, 2322, 2354, 2377, 2384, 2422, 2450, 2454, 2460, 2465, 2524, 2529, 2614, 2616, 2625, 2661, 2685, 2726, 2861, 2880, 2897, 2957, 2971, 2999, 3002, 3070, 3086, 3104, 3110, 3136, 3163, 3175, 3213, 3222, 3249, 3322, 3357, 3392, 3423, 3443, 3463, 3503, 3598, 3606, 3639, 3653, 3698, 3704, 3751, 3774]. Linear-in-the-Parameters [736]. Linearisations [2422]. Linearity [1809]. Linearity-Preserving [1809]. Linearization [2925]. Linearized [776, 1124, 1307, 1682, 1762, 1894, 1967, 2135, 2151, 2209, 2210, 2388, 2402, 2576, 2686, 2702, 2775, 2785, 2811, 3368, 3549]. Linearly [698, 1190, 1196, 1522, 2438, 2572, 2606, 2788, 2805, 2825, 2890, 3044, 3073, 3414, 3558]. Linearly-Convergent [3414]. Lines [54, 654, 1770, 3174, 3519]. Liouville [1655, 1905, 2913, 3058]. Lipid [3428]. Lipschitz [238, 2578, 2923, 3196, 3560]. Liquid [692, 849, 1508, 1678, 1995]. Ljusternik [1516]. Load [201]. Lobatto [258, 528, 2081, 3639]. Local [125, 153, 201, 269, 289, 294, 302, 344, 371, 376, 434, 442, 443, 543, 559, 636, 668, 690, 695, 733, 779, 780, 802, 841, 926, 1007, 1009, 1013, 1017, 1047, 1058, 1087, 1097, 1146, 1212, 1233, 1242, 1261, 1281, 1286, 1287, 1312, 1362, 1391, 1401, 1418, 1487, 1667, 1695, 1700, 1713, 1731, 1742, 1750, 1814, 1874, 1875, 1903, 1922, 1935, 1936, 1945, 1948, 1953, 1976, 1982, 2052, 2147, 2149, 2176, 2209, 2210, 2237, 2240, 2241, 2257, 2264, 2265, 2270, 2280, 2311, 2438, 2470, 2511, 2535, 2575, 2680, 2700, 2760, 2761, 2835, 2897, 2971, 2984, 3021, 3078, 3086, 3092, 3135, 3155, 3227, 3236, 3237, 3412, 3413, 3550, 3729]. Local [894, 1395, 1482, 1536, 1643, 1755, 1950, 2218, 2322, 2643, 2731, 2870, 2999, 3308]. Local-Min-Orthogonal [3155]. Local-Structure-Preserving [926, 1017]. Localization [3016]. Localizations [3779]. Localized [2156, 2541, 3335]. Locally [118, 453, 609, 698, 985, 1175, 1800, 2007, 2138, 2784, 2923, 3375]. Locally-Conservative [3375]. Locating [256]. loci [184]. Lock [959]. Locked [284, 644]. Locking [1045, 2749, 3375]. Locking-Free [1045, 2749, 3375]. Loewner [3106]. Log [3329]. Logarithmic [2035, 3460, 3588]. Long [192, 331, 742, 782, 1313, 1845, 1958, 2017, 2154, 2650, 3031, 3071, 3488, 3665]. Long-Range [3031]. Long-Term [192, 1313]. Long-Time [742, 782, 1845, 3071, 3488, 3665]. Longitudinal [2934]. Loop [1373, 2652]. Lorentz [1111, 1644]. Loss [937, 2736, 3496]. Lossy [498, 1288, 2939]. Love [3725]. Low [18, 86, 133, 151, 203, 311, 478, 489, 582, 897, 1170, 1186, 1486, 1549, 1614, 1721, 1727, 1755, 1786, 1797, 1846, 1869, 2006, 2046, 2056, 2160, 2168, 2195, 2200, 2214, 2272, 2278, 2319, 2365, 2480, 2543, 2568, 2643, 2680, 2689, 2776, 2854, 2899, 2945, 3015, 3033, 3035, 3084, 3089, 3093, 3231, 3451, 3475, 3496, 3527, 3554, 3569, 3571, 3656, 3720, 3764, 3788]. Low-CP-Rank [3093]. Low-Degree [2689]. Low-Dimensional [478, 2319]. Low-Dissipation [1170, 2200, 3033, 3569]. Low-light [3720]. Low-Mach [1869, 2568]. Low-Memory [2278]. Low-Rank [1186, 1755, 1846, 2006, 2195, 2365, 2480, 2854, 2945, 3015, 3035, 3231, 3475, 3554, 3656, 3788]. Low-Regularity [897, 3084, 3764]. Low-Resolution [2680]. Low-Reynolds [151]. Low-Speed [2899]. Low-Storage [1614, 2046]. Lower [1249, 1325, 1357, 1620, 2952, 3281]. Lower/Upper [1249, 1357]. Lowest [3290, 3331]. Lowest-Degree [3290]. LPS [1613]. LSQ [111]. LSQR [3739]. LST [3467]. Lucy [1987]. Lumped [3192]. Lyapunov [2158, 2790].

MAC [1411, 1474, 2150, 2550, 3068, 3545]. MacCormack [691]. Mach [18, 75, 133, 203, 311, 388, 459, 489, 670, 1450, 1869, 2212, 2307, 2543, 2568, 3718, 3790]. Machine [101, 2379, 2479, 3157, 3188, 3218, 3239, 3333, 3356, 3360, 3385, 3446, 3458, 3497, 3689, 3807]. machines [2947]. Maclaurin [677]. Macroscopic [674]. Magma [973]. Magnetic [1420, 1659, 1721, 2160, 2856]. Magnetized [2211]. Magneto [383, 490, 2694, 2712, 3439]. Magneto-Aerodynamics [490]. Magneto-Hydrodynamics [2694, 2712]. Magneto-Rheological [383]. Magnetohydrodynamic [607, 2160, 2244, 2860, 2889, 3048, 3437]. Magnetohydrodynamics [742, 776, 1400, 1500, 1593, 1779, 1979, 2335, 2363, 2381, 2508, 2510, 2751, 3041, 3148, 3165, 3390, 3494, 3520, 3718]. Magnetohydronamics [2581]. Magnetostatic [2613]. Magnus [2974].

Mahony [2437, 2575, 2785]. Making [3157]. Management [829, 3483]. Manakov [2845]. Manifold [752, 2006, 2056, 2214, 2425, 3423, 3446, 3462, 3580, 3784]. Manifolds [478, 1178, 2110, 2952, 3193, 3346, 3354, 3640]. Manner [941]. Many [713, 1479, 2302, 3782]. Many-Particle [713]. Many-Stage [3782]. Map [1117, 1297, 1956, 2747]. Mapped [467, 1030, 1150, 1594, 3334, 3659, 3681]. Mapping [185, 1255, 1594, 1671, 1672, 1831, 1895, 2607, 2885]. Mappings [1072, 2202, 2672]. Maps [1148, 1611, 2109, 2533, 2546]. Marangoni [403, 434]. Marcher [3105]. Marching [117, 144, 197, 219, 753, 2257]. Markov [873, 2102, 2323, 3190, 3280]. Markovian [9, 2869]. Marquardt [2001, 2276]. Martensitic [385]. Martian [910]. Mass [541, 846, 1058, 1787, 1899, 1917, 2014, 2182, 2249, 2262, 2445, 2766, 2896, 3025, 3191, 3192, 3246, 3413, 3633, 3682]. Mass- [541, 3682]. Mass-Conservative [846]. Mass-conserving [2766]. Mass-Lumped [3192]. Mass-Preserving [3191]. Masses [1227]. Massively [6, 172, 964]. Master [665]. Matched [331, 1103, 1387, 2911]. Matching [395, 647, 1488, 1611, 1900, 2280, 3119]. Material [366, 1659, 2350, 2506, 3094, 3313]. Materials [560, 698, 710, 2773, 2827, 3445, 3555]. Mathematical [234, 558, 672, 715, 840, 1000, 2700]. Mathematics [2654]. Matrices [13, 51, 93, 141, 245, 474, 595, 825, 1218, 1817, 1846, 1917, 2284, 2302, 2551, 2848, 3628]. Matrix [132, 233, 553, 930, 943, 1039, 1186, 1217, 1476, 1493, 1552, 1787, 2024, 2031, 2038, 2080, 2132, 2182, 2184, 2480, 2562, 2633, 2743, 2782, 2801, 2807, 2830, 2854, 2945, 3015, 3111, 3303, 3352, 3384, 3461, 3487, 3504, 3554, 3685, 3795]. Matrix-Free [1552, 3504]. Matrix-Valued [2182]. Matrix-Vector [930]. Matter [878]. Max [1486]. Max-Norm [1486]. Maximal [1314]. Maximization [2715, 3384]. Maximum [26, 981, 1228, 1464, 1605, 1787, 1980, 2055, 2151, 2368, 2650, 2743, 2785, 2799, 2802, 2956, 3054, 3220, 3248, 3253, 3292, 3304, 3552]. Maximum-Principle [1787, 3552]. Maximum-Principle-Satisfying [981, 2368, 2799]. Maxwell [263, 449, 519, 648, 707, 768, 769, 797, 861, 891, 915, 1012, 1028, 1165, 1175, 1475, 1521, 1590, 1599, 1659, 1710, 1774, 1829, 1947, 2036, 2051, 2192, 2299, 2303, 2305, 2396, 2433, 2523, 2742, 2753, 2888, 2946, 3055, 3101, 3313, 3544, 3691]. May [315]. MBE [2958, 2968]. MBO [1973]. MDCD [3042]. MDCD-TENO [3042]. Mean [389, 464, 2077, 2595, 2841, 3161, 3170]. Means [995, 1575, 2223, 2252, 2887, 3283]. Measure [2896, 3748]. Measurement [3758]. Measurement-Induced [3758]. Measurements [657, 1737, 2240, 2680]. Measures [2317, 2803, 2894, 3415]. Mechanical [2378, 2982]. Mechanics [485, 1984, 3318, 3509]. Media [108, 246, 454, 543, 823, 837, 915, 1163, 1253, 1259, 1321, 1402, 1467, 1656, 1814, 1841, 1901, 1948, 2027, 2152, 2192, 2257, 2358, 2393, 2412, 2416, 2522, 2534, 2602, 2630, 2648, 2651, 2733, 2753, 2771, 2879, 2917, 2939, 3061, 3068, 3304, 3691]. Medical [1000]. Medium [733, 1083, 1376, 2011, 2678, 2786, 3012, 3276, 3540]. Membership [1685]. Membrane [3379]. Membranes [2655]. Memory [108, 902, 1119, 1352, 1510, 2216, 2278, 2488, 3106, 3129, 3263].

Memory-Saving [3106]. Merton [1063, 1503]. Mesh [96, 297, 418, 468, 473, 495, 523, 538, 659, 763, 792, 1062, 1119, 1151, 1194, 1208, 1384, 1543, 1861, 1958, 2071, 2104, 2138, 2180, 2194, 2296, 2301, 2334, 2437, 2441, 2552, 2707, 2745, 2764, 2812, 3206, 3266, 3386, 3397, 3403, 3459, 3497, 3642, 3752, 3809]. Mesh-Free [3497]. Mesh-Informed [3642]. Mesh-Robustness [3206]. Meshes [208, 268, 373, 470, 511, 513, 530, 540, 596, 598, 647, 662, 758, 793, 821, 834, 981, 1009, 1089, 1197, 1208, 1233, 1234, 1238, 1254, 1258, 1357, 1361, 1434, 1453, 1457, 1462, 1474, 1525, 1546, 1578, 1622, 1624, 1681, 1689, 1740, 1827, 1843, 1891, 2057, 2089, 2106, 2143, 2167, 2193, 2233, 2248, 2281, 2289, 2368, 2470, 2491, 2508, 2519, 2529, 2560, 2636, 2645, 2688, 2791, 2824, 2835, 2844, 2884, 2902, 2927, 2940, 2966, 3026, 3043, 3141, 3215, 3225, 3249, 3398, 3443, 3463, 3481, 3484, 3498, 3536, 3539, 3562, 3703, 3730, 3746, 3765, 3781, 3804]. Meshfree [1144, 3273, 3598, 3602, 3653, 3772]. Meshless [1381, 2148, 2464, 2870, 2874]. Metamaterial [1644]. Metamaterials [637, 1111, 1710, 2628, 3159]. Method [210, 323, 343, 442, 499, 517, 521–524, 559, 585, 641, 649, 709, 724, 729, 767, 769, 811, 916, 966, 975, 984, 1025, 1047, 1055, 1087, 1102, 1117, 1162, 1165, 1225, 1234, 1237, 1286, 1287, 1356, 1381, 1392, 1413, 1446, 1490, 1533, 1577, 1659, 1664, 1682, 1709, 1769, 1919, 1995, 2004, 2069, 2090, 2095, 2101, 2169, 2176, 2181, 2195, 2228, 2232, 2290, 2316, 2323, 2336, 2355, 2364, 2415, 2564, 2630, 2648, 2679, 2760, 2945, 3011, 3038, 3098, 3111, 3123, 3277, 3280, 3403, 3413, 3429, 3486, 3493, 3596, 3600, 3637, 3651, 3661, 3691, 3722, 3735, 3738, 3802, 3807]. Method [312, 338, 358, 443, 446, 563, 636, 650, 698, 719, 753, 763, 771, 772, 794, 875, 979, 992, 1032, 1063, 1074, 1085, 1175, 1236, 1279, 1296, 1307, 1337, 1401, 1456, 1480, 1505, 1518, 1550, 1551, 1641, 1683, 1724, 1766, 1770, 1774, 1824, 1832, 1836, 1849, 1875, 1968, 2014, 2060, 2074, 2106, 2129, 2146, 2237, 2287, 2289, 2331, 2359, 2383, 2403, 2443, 2445, 2458, 2514, 2526, 2540, 2555, 2619, 2644, 2660, 2662, 2672, 2678, 2724, 2841, 2844, 2924, 2967, 2994, 3007, 3028, 3144, 3155, 3196, 3197, 3203, 3218, 3228, 3250, 3312, 3333, 3383, 3389, 3519, 3575, 3583, 3622, 3667, 3719, 3760, 3766, 3770, 3792]. Method [21, 202, 224, 276, 282, 322, 326, 327, 340, 415, 464, 530, 531, 600, 664, 693, 723, 731, 761, 773, 802, 808, 812, 879, 1010, 1105, 1191, 1304, 1400, 1512, 1514, 1522, 1524, 1528, 1553– 1555, 1625, 1627, 1714, 1728, 1866, 1885, 1900, 1927, 1928, 2023, 2067, 2071, 2164, 2180, 2199, 2205, 2209, 2210, 2239, 2246, 2276, 2312, 2324, 2328, 2329, 2335, 2378, 2381, 2423, 2433, 2473, 2510, 2523, 2525, 2569, 2603, 2706, 2791, 2830, 2851, 2889, 2901, 2909, 2911, 2955, 2996, 3021, 3050, 3067, 3158, 3181, 3212, 3240, 3254, 3279, 3291, 3302, 3305, 3325, 3347, 3397, 3459, 3512, 3567, 3626, 3627, 3746, 3747, 3769]. Method [198, 228, 383, 496, 533, 588, 617, 703, 717, 774, 776, 791, 797, 835, 862, 864, 882, 974, 977, 990, 1078, 1097, 1113, 1124, 1141, 1180, 1201, 1230, 1235, 1245, 1269, 1272, 1431, 1472, 1488, 1493, 1508, 1559, 1643, 1656, 1729, 1730, 1737, 1765, 1814, 1861, 1876, 1882, 1911, 1929, 1969, 2009, 2010, 2013, 2015, 2147, 2149, 2155, 2159, 2248, 2268, 2271, 2319, 2334, 2349, 2432, 2452, 2503, 2512, 2521, 2553, 2554, 2580, 2674, 2675, 2677, 2705, 2778, 2816, 2820, 2827, 2928, 2932, 2935, 3022, 3108, 3135, 3229, 3235, 3246, 3263, 3286, 3315, 3351, 3352, 3395, 3410, 3452, 3454, 3485, 3489, 3537, 3542, 3663, 3702, 3754, 3773, 3810]. Method [223, 272, 410, 414, 418, 507, 513, 519, 537, 538, 593, 654, 683, 743, 787, 848, 885, 905, 908, 948, 989, 1009, 1017, 1027, 1086, 1100, 1125, 1190, 1194, 1198, 1216, 1243, 1278, 1298, 1340, 1448, 1455, 1478, 1520, 1540, 1549, 1557, 1587, 1610, 1621, 1626, 1801, 1815, 1830, 1840, 1862, 1878, 1881, 1915, 1933– 1937, 1957, 1973, 2000, 2043, 2063, 2083, 2096, 2166, 2224, 2249, 2392, 2453, 2477, 2538, 2628, 2673, 2681, 2703, 2709, 2719, 2722, 2729, 2740, 2784, 2810, 2855, 2888, 2927, 2954, 3012, 3097, 3101, 3117, 3129, 3141, 3149, 3165, 3268, 3292, 3375, 3433, 3436, 3505, 3538, 3665, 3673, 3723, 3732, 3787, 3793]. Method [208, 225, 291, 334, 454, 460, 469, 621, 686, 688, 694, 750, 1080, 1095, 1178, 1207, 1222, 1317, 1324, 1343, 1378, 1405, 1513, 1648, 1677, 1711, 1790, 1831, 1888, 1938, 1983, 1984, 1999, 2037, 2135, 2138, 2168, 2179, 2197, 2253, 2262, 2264, 2265, 2277, 2301, 2345, 2348, 2385, 2397, 2418, 2437, 2446, 2447, 2461, 2522, 2527, 2536, 2583, 2668, 2671, 2700, 2714, 2717, 2761, 2804, 2833, 2890, 2956, 3024, 3040, 3060, 3077, 3079, 3089, 3092, 3094, 3095, 3114, 3119, 3146, 3167, 3255, 3260, 3267, 3281, 3296, 3301, 3328, 3336, 3368, 3396, 3409, 3411, 3437, 3449, 3455, 3472, 3482, 3520, 3554, 3579, 3581, 3614, 3630, 3639, 3652, 3657, 3670, 3678, 3690, 3728, 3771, 3805]. Method [200, 205, 292, 315, 341, 344, 391, 406, 455, 511, 539, 541, 573, 603, 622, 633, 690, 691, 725, 756, 796, 815, 836, 855, 888, 894, 924, 953, 1029, 1066, 1187, 1202, 1226, 1267, 1273, 1305, 1310, 1339, 1349, 1412, 1433, 1473, 1475, 1504, 1517, 1527, 1529, 1586, 1591, 1632, 1633, 1661, 1699, 1704, 1735, 1780, 1901, 1980, 1982, 1990, 2062, 2094, 2142, 2160, 2161, 2255, 2273, 2318, 2343, 2344, 2346, 2354, 2374, 2401, 2420, 2428, 2486, 2498, 2506, 2563, 2617, 2645, 2667, 2676, 2842, 2950, 2981, 3055, 3070, 3080, 3143, 3164, 3252, 3265, 3310, 3353, 3423, 3467, 3473, 3481, 3524, 3647, 3693, 3740]. Method [349, 351, 394, 395, 404, 551, 584, 604, 630, 676, 718, 779, 819, 820, 895, 912, 986, 1003, 1041, 1051, 1136, 1146, 1206, 1211, 1281, 1354, 1388, 1396, 1427, 1588, 1594, 1607, 1623, 1630, 1658, 1743, 1744, 1747, 1763, 1779, 1797, 1808, 1826, 1838, 1870, 1910, 1945, 2055, 2059, 2111, 2118, 2144, 2208, 2257, 2285, 2286, 2377, 2384, 2424, 2455, 2463, 2472, 2535, 2539, 2598, 2610, 2635, 2646, 2726, 2762, 2832, 2837, 2866, 2872, 2897, 2998, 3009, 3043, 3059, 3078, 3124, 3127, 3145, 3163, 3176, 3186, 3201, 3233, 3249, 3314, 3317, 3335, 3346, 3379, 3404, 3462, 3511, 3539, 3540, 3558, 3565, 3605, 3623, 3641, 3658, 3666, 3705, 3708, 3730]. Method [218, 260, 269, 346, 492, 555, 661, 669, 687, 692, 700, 733, 780, 823, 898, 904, 918, 920, 927, 930, 970, 1002, 1013, 1021, 1022, 1067, 1070, 1093, 1115, 1134, 1137, 1319, 1326, 1346, 1372, 1387, 1491, 1534, 1582, 1596, 1605, 1644, 1732, 1751, 1785, 1867, 1909, 1948, 1988, 1996, 2087, 2131, 2190, 2198, 2320, 2321, 2327, 2353, 2429, 2440, 2459, 2481, 2511, 2519, 2547, 2612, 2647, 2683, 2688, 2727, 2799, 2812, 2860, 2877, 2879, 2892, 2896, 2910, 2936, 2973, 3046, 3049, 3061, 3136, 3148, 3175, 3198, 3222, 3236, 3276, 3287, 3327, 3337, 3380, 3386, 3391, 3418, 3471, 3503, 3507, 3580, 3587, 3595, 3597, 3604, 3695, 3701, 3752, 3762, 3767, 3791, 3803]. Method [7, 13, 26, 33, 43, 50, 52, 63, 77, 93, 117, 118, 127, 152, 156, 163, 164, 170, 174, 175, 178, 183, 185, 192, 194, 457, 755, 758, 805, 902, 951, 971, 999, 1094, 1172, 1247, 1536, 1640, 1733, 1857, 1914, 1950, 1951, 2174, 2218, 2632, 2707, 2759, 2781, 2854, 2859, 2913, 2999, 3015, 3045, 3058, 3308, 3369, 3371, 3777, 3801]. Method-of-Lines [654, 3519]. Methodology [963]. Methods [440, 441, 444, 480, 516, 526, 653, 675, 679, 770, 822, 837, 850, 870, 896, 967, 979, 1039, 1045, 1118, 1145, 1151, 1152, 1164, 1189, 1210, 1219, 1263, 1264, 1274, 1290, 1295, 1332, 1369, 1402, 1420, 1424, 1444, 1445, 1509, 1519, 1521, 1526, 1546, 1590, 1613, 1670, 1689, 1725, 1778, 1803, 1812, 1833, 1860, 1874, 1883, 1918, 1920, 1954, 1967, 1978, 2052, 2127, 2129, 2139, 2183, 2192, 2220, 2233–2236, 2238, 2281, 2393, 2475, 2530, 2640, 2686, 2698, 2720, 2725, 2809, 2823, 2894, 2940, 3039, 3076, 3151, 3182, 3205, 3214, 3243, 3274, 3318, 3408, 3476, 3515, 3603, 3611, 3668, 3736]. Methods [288, 305, 339, 352, 361, 409, 445, 488, 495, 510, 527, 655, 701, 813, 860, 876, 929, 932, 1000, 1035, 1052, 1064, 1075, 1122, 1144, 1155, 1168, 1173, 1261, 1313, 1419, 1576, 1608, 1634, 1639, 1692, 1703, 1894, 1922, 1925, 2011, 2022, 2054, 2085, 2110, 2193, 2206, 2219, 2241, 2242, 2244, 2245, 2247, 2288, 2332, 2338, 2341, 2387, 2390, 2396, 2464, 2468, 2470, 2509, 2545, 2574, 2584, 2624, 2655, 2682, 2699, 2716, 2735, 2749, 2756, 2863, 2868, 2875, 2916, 3023, 3062, 3072, 3074, 3087, 3156, 3160, 3171, 3211, 3248, 3376, 3421, 3501, 3527, 3557, 3608, 3621, 3645, 3683, 3714, 3770, 3782]. Methods [221, 255, 301, 317, 355, 399, 425, 452, 462, 479, 486, 487, 509, 514, 535, 713, 749, 775, 842, 857, 858, 926, 949, 991, 1088, 1123, 1140, 1246, 1249, 1300, 1325, 1335, 1363, 1364, 1368, 1390, 1397, 1418, 1439, 1471, 1487, 1503, 1510, 1541, 1614, 1617, 1667, 1694, 1741, 1758, 1828, 1908, 1969, 1977, 2019, 2025, 2041, 2046, 2066, 2073, 2082, 2105, 2120, 2137, 2172, 2261, 2311, 2502, 2505, 2531, 2541, 2592, 2626, 2651, 2815, 2821, 2834, 2906, 2908, 2918, 2948, 2985, 3013, 3026, 3027, 3069, 3122, 3285, 3303, 3419, 3422, 3457, 3478, 3483, 3494, 3525, 3541, 3564, 3669, 3715, 3763]. Methods [226, 294, 328, 453, 473, 518, 540, 638, 645, 689, 696, 845, 915, 982, 1036, 1057, 1096, 1106, 1143, 1221, 1232, 1406, 1459, 1463, 1481, 1498, 1499, 1502, 1572, 1592, 1606, 1635, 1637, 1669, 1701, 1706, 1748, 1756, 1757, 1792, 1821, 1904, 2034, 2077, 2081, 2097, 2119, 2134, 2204, 2250, 2251, 2254, 2263, 2269, 2282, 2292, 2308, 2340, 2342, 2368, 2376, 2426, 2465, 2493, 2513, 2524, 2575, 2608, 2611, 2634, 2636, 2657, 2715, 2741, 2763, 2776, 2798, 2850, 2898, 2904, 2914, 2923, 2929, 2933, 2946, 2972, 2988, 3008, 3121, 3153, 3173, 3227, 3238, 3273, 3338, 3401, 3415, 3432, 3447, 3465, 3479, 3533, 3573, 3619, 3704, 3725, 3737]. Methods [222, 266, 279, 302, 314, 350, 354, 392, 407, 426, 467, 468, 503, 545, 547, 549, 550, 577, 738, 760, 809, 827, 897, 1040, 1091, 1104, 1108, 1158, 1166, 1185, 1197, 1328, 1342, 1367, 1410, 1468, 1574, 1601, 1646, 1650, 1651, 1666, 1673, 1753, 1764, 1783, 1798, 1947, 1970, 1976, 2003, 2012, 2092, 2104, 2114, 2121, 2140, 2141, 2152, 2163, 2326, 2352, 2363, 2367, 2370, 2398, 2431, 2529, 2581, 2605, 2625, 2631, 2639, 2659, 2718, 2847, 2880, 2930, 2962, 3014, 3037, 3086, 3172, 3245, 3264, 3266, 3275, 3294, 3354, 3363, 3400, 3450, 3463, 3491, 3509, 3550, 3572, 3598, 3610, 3636, 3653, 3686, 3698, 3711, 3750, 3772, 3789]. Methods [12, 27, 34, 39, 41, 53, 59–61, 74, 80, 81, 86, 98, 111, 124, 128, 134, 142, 146, 155, 158, 160, 188, 207, 287, 579, 704, 732, 766, 781, 786, 1023, 1034, 1160, 1271, 1322, 1500, 1615, 1618, 1752, 1755, 1871, 1949, 1955, 1971, 2005, 2058, 2145, 2167, 2298, 2322, 2357, 2702, 2731, 2824, 2870, 2903, 3168, 3355, 3359, 3417, 3674, 3797]. Metric [513, 2928, 3689]. Metrics [2764]. Meyer [951]. MFS [553, 914, 1171]. MGFM [3276]. MHD [163, 453, 470, 550, 582, 1629, 1742, 1791, 2540, 2644, 2995, 3033, 3228, 3752, 3790]. Microflows [1198, 1967]. Microgravity [379]. Microlithographic [195]. Micromagnetics [2422]. Microscope [2441]. Microscopy [2704]. Microstructure [385, 1428, 2746, 3094]. Microwave [575]. Midpoint [2422]. migration [40]. Milstein [1776, 2355]. Milstein-Type [1776]. Mimetic [1167, 2446, 2447]. MIMO [2095]. Min [3155]. Mindlin [438, 967, 1435, 2062]. Minimal [290, 636, 653, 1382, 3205, 3503, 3513, 3527, 3789]. Minimally [2278]. Minimax [359, 1372, 1516, 2264, 2265, 2297]. Minimization [397, 531, 582, 650, 1121, 1190, 1255, 1318, 1549, 1563– 1565, 1648, 1684, 1906, 1986, 2107, 2114, 2292, 2310, 2333, 2353, 2494, 2577, 2637, 3009, 3365, 3516, 3542, 3713, 3786]. Minimized [2959, 3183]. Minimizing [1720, 2170, 2317, 2658, 3140]. Minimum [969, 1795, 3502]. Mining [2038]. miniNode [3]. Minkowski [377]. minmod [564]. Minus [2366]. Mirror [2658, 3241, 3418]. Miscible [1683, 1814, 1948, 2257, 2416, 2602, 2738, 2879]. MISTRAL [847]. Mittag [2184, 2532, 2782]. Mittag-Leffler [2184, 2532, 2782]. Mixed [185, 244, 346, 367, 441, 449, 450, 499, 601, 704, 717, 736, 776, 795, 819, 822, 858, 891, 899, 908, 919, 941, 967, 979, 983, 1066, 1069, 1089, 1093, 1173, 1203, 1237, 1240, 1328, 1346, 1537, 1554, 1555, 1571, 1597, 1601, 1623, 1630, 1706, 1752, 1756, 1803, 1816, 1862, 1885, 1900, 1969, 1971, 1972, 2025, 2087, 2129, 2175, 2183, 2197, 2209, 2210, 2228, 2243, 2245, 2283, 2290, 2330, 2393, 2446, 2447, 2451, 2462, 2523, 2529, 2547, 2569, 2570, 2620, 2628, 2643, 2681, 2697, 2738, 2820, 2830, 2868, 2879, 2932, 3113, 3141, 3160, 3167, 3198, 3254, 3272, 3312, 3315, 3328, 3396, 3494, 3501, 3515, 3540, 3634, 3668, 3728, 3808]. mixed [28, 171, 2676]. Mixed-FEM [2330]. Mixed-Integer [2451]. Mixed-Primal [3254, 3501]. Mixed-Subordinate [2830]. Mixed-Type [1816]. Mixing [75, 923, 959]. Mixture [1575, 2031, 2692]. Mobile [3306, 3562, 3797]. Mobile/Immobile [3562]. Modal [1849, 2300, 2467]. Mode [284, 644, 874, 1001, 2174, 3166, 3492]. Mode-Locked [284, 644]. Model [204, 231, 235, 285, 308, 581, 625, 668, 720, 763, 773, 818, 884, 900, 938, 942, 967, 990, 1011, 1032, 1045, 1062, 1065, 1089, 1090, 1102, 1105, 1111, 1277, 1330, 1359, 1392, 1524, 1580, 1597, 1628, 1692, 1696, 1706, 1712, 1816, 1864, 1936, 1969, 1998, 2007, 2070, 2095, 2177, 2187, 2213, 2214, 2249, 2299, 2311, 2351, 2358, 2364, 2392, 2403, 2442, 2448, 2449, 2456, 2545, 2678, 2737, 2786, 2839, 2882, 2924, 2926, 2931, 2939, 2963, 2990, 2994, 3010, 3021, 3030, 3066, 3083, 3122, 3166, 3188, 3209, 3218, 3219, 3227, 3316, 3326, 3373, 3425, 3452, 3480, 3553, 3591, 3601, 3625, 3655, 3667, 3702, 3721, 3812]. Model [112, 150, 172, 181, 325, 406, 434, 656, 759, 906, 911, 912, 951, 970, 1043, 1066, 1071, 1084, 1203, 1268, 1308, 1315, 1326, 1345, 1370, 1415, 1570, 1630, 1644, 1678, 1788, 1946, 1951, 1965, 1974, 1984, 2018, 2027, 2050, 2056, 2138, 2297, 2343, 2395, 2439, 2501, 2522, 2535, 2548, 2578, 2596, 2614, 2616, 2621, 2710, 2793, 2794, 2853, 2881, 2975, 2976, 3041, 3046, 3063, 3104, 3174, 3202, 3206, 3216, 3264, 3284, 3333, 3345, 3374, 3384, 3385, 3399, 3428, 3470, 3492, 3528, 3560, 3616, 3684, 3776, 3800]. Model-Reduction [2449]. Modeled [2938, 3213]. Modeling [131, 229, 231, 236, 240, 397, 403, 408, 416, 489, 710, 753, 859, 886, 1163, 1182, 1289, 1699, 1777, 1805, 1810, 1903, 1930, 1946, 1962, 1965, 2078, 2121, 2496, 2669, 2705, 2746, 2783, 2792, 3091, 3202, 3489, 3716, 3742, 3779]. Modelling [447, 1679, 2193, 2630, 3387, 3446]. Models [104, 130, 207, 279, 469, 470, 493, 527, 663, 721, 736, 777, 806, 809, 880, 960, 1022, 1063, 1266, 1284, 1361, 1366, 1408, 1503, 1660, 1731, 1772, 1782, 1840, 1851, 1960, 1961, 2096, 2132, 2314, 2315, 2450, 2479, 2670, 2680, 2686, 2756, 2856, 2869, 2900, 3132, 3135, 3223, 3243, 3295, 3317, 3356, 3360, 3382, 3423, 3442, 3466, 3502, 3541, 3692, 3711, 3731, 3760]. Moderately [229]. Modes [48, 1771, 2433, 2827, 3142]. Modica [3485]. Modification [201, 2899]. modifications [56]. Modified [69, 183, 222, 228, 276, 336, 341, 576, 694, 719, 931, 951, 1226, 1344, 1441, 1455, 1484, 1584, 1691, 1836, 2035, 2169, 2320, 2321, 2478, 2593, 2820, 2910, 2994, 3112, 3170, 3276, 3491, 3545, 3582, 3605, 3635, 3638, 3765]. Modular [2563, 3252]. Molecular [8, 120, 298, 3122]. Molecules [1600]. Moment [663, 697, 731, 1125, 1284, 1438, 1476, 1548, 1889, 1967, 2076, 2359, 2506, 2516, 2956, 3356, 3361, 3438, 3442]. Moment-Closure [3442]. Moment-of-Fluid [2506]. Moments [2084]. Momentum [541, 1497, 3456, 3694]. Monge [929, 1114, 1716, 2094, 2175, 2181, 2328, 2329, 2405, 2561, 3479, 3629]. Monitor [418]. Monitored [3731]. Monolithic [1859]. Monotone [417, 753, 807, 1025, 1561, 2115, 2143, 2175, 2444, 2473, 3014, 3107, 3422, 3598, 3605, 3653]. Monotonically [1727]. Monotonicity [142, 462, 991, 1170, 2155, 2593]. Monotonicity-Preserving [1170]. Monte [149, 298, 714, 920, 1538, 2433, 2564, 2878, 2912, 3046, 3173, 3324, 3351, 3748, 3795]. MOOD [2652, 2849]. Moreau [3136]. Morley [2126, 2816]. morphological [34]. Morphology [715]. Mortar [350, 351, 358, 406, 530, 585, 601, 708, 769, 1353, 1900, 2940, 3156]. Mortar-based [2940]. Mortar-Discontinuous [769]. Mortola [3485]. Moser [1452]. Motion [389, 392, 422, 690, 1127, 1132, 1732, 2981, 2991, 3170]. Motions [1686]. Motivated [2909]. Mouse [1005]. Movements [1408]. Mover [2037]. Moving [398, 418, 466, 469, 470, 495, 538, 662, 688, 957, 1085, 1094, 1208, 1233, 1279, 1763, 1836, 1888, 1958, 1965, 2089, 2407, 2560, 2681, 2812, 3174, 3471, 3566, 3703, 3752]. Moving-Water [957, 2407]. Müller [951]. Mullins [3747]. Multi [154, 176, 327, 432, 473, 483, 497, 507, 589, 617, 632, 654, 671, 685, 710, 721, 763, 793, 872, 898, 1228, 1313, 1319, 1395, 1413, 1414, 1438, 1653, 1684, 1693, 1781, 1806, 1818, 1889, 1900, 1902, 1928, 2033, 2068, 2071, 2087, 2180, 2189, 2274, 2276, 2282, 2285, 2286, 2350, 2433, 2457, 2475, 2506, 2521, 2525, 2552, 2596, 2623, 2649, 2719, 2776, 2804, 2808, 2819, 2870, 2873, 2877, 2946, 2956, 2978, 3024, 3066, 3094, 3110, 3180, 3186, 3229, 3276, 3305, 3311, 3323, 3364, 3393, 3512, 3521, 3560, 3568, 3622, 3674, 3699, 3761, 3769, 3780, 3794]. Multi-agent [3568]. Multi-Asset [654]. Multi-Block [432, 632, 1684, 1781, 1818, 1900]. Multi-Class [721]. Multi-Component [2804, 2978, 3024, 3180, 3512, 3769]. Multi-contact [3521]. Multi-D [2808]. Multi-Delayed [2877]. Multi-Dimensional [617, 1228, 1413, 1693, 1806, 2033, 2068, 2189, 2623, 3066, 3311]. Multi-Dimensions [507, 2552, 2873, 2946]. Multi-Domain [327, 872, 3186]. Multi-element [2475]. Multi-Fluid [2596]. Multi-frequency [3699].

Multi-fronts [1414]. Multi-implicit [2282]. Multi-Level [497, 2087, 2649]. Multi-Linear [1653, 3761]. Multi-Material [2350, 2506, 3094]. Multi-medium [3276]. Multi-Mesh [473, 763]. Multi-meshes [793]. Multi-modes [2433]. Multi-Moment [1438, 1889]. Multi-Order [2525, 3110, 3229]. Multi-particle [898]. Multi-phase [2506, 3560, 3674]. Multi-phase/Multi-material [2506]. Multi-quadric [2071]. Multi-Rank [2776, 3323]. Multi-resolution [154, 176, 1928, 2719, 3305, 3364, 3393, 3794]. Multi-scale [710, 1319, 1395, 2956]. Multi-Shifted [2285, 2286]. Multi-singularity [3622]. Multi-stage [2457]. Multi-step [2276, 2819, 3780]. Multi-symplectic [2870]. Multi-term [1902, 2180, 2274, 2521]. Multi-Value [1313]. Multi-Valued [589, 685]. Multiblock [408, 416]. Multiclass [3444]. Multicomponent [662, 2966, 3467]. Multiderivative [1291, 1943, 3032, 3074]. Multidimensional [226, 810, 1138, 1705, 1858, 2179, 2384, 2559, 2675, 2770, 2899, 2929, 3018, 3146, 3184, 3307, 3544, 3615]. Multidimensions [360]. Multidirectional [1773]. Multidomain [107, 125, 140, 153, 171, 309, 391, 3055]. Multifidelity [3351]. Multifluid [1547, 2943]. Multigrid [30, 43, 47, 108, 114, 180, 337, 460, 650, 699, 760, 790, 851, 1136, 1185, 1269, 1298, 1444, 1445, 1780, 1847, 1852, 1969, 1997, 2129, 2183, 2281, 2511, 2567, 2835, 2841, 3046, 3103, 3343, 3391, 3407, 3519, 3613, 3710]. Multigrid/Schwarz [460]. Multilabel [1394]. Multilayer [498, 1304, 2692, 3030, 3601]. Multilevel [134, 208, 235, 488, 1050, 1200, 1585, 1690, 1867, 1876, 2306, 2375, 2461, 2564, 3046, 3236, 3324, 3738]. Multilinear [2171]. Multimodal [3261]. multimode [29]. Multiparameter [1773]. Multipartite [3099]. Multipatch [3659, 3681]. Multiphase [272, 616, 1685, 3006]. multiphoton [122]. Multiphysics [2002, 3197]. Multiple [46, 828, 1149, 1168, 1172, 1193, 1206, 1372, 1516, 1564, 1606, 1696, 1725, 2009, 2130, 2285, 2286, 2512, 2555, 2572, 2726, 3140, 3155, 3430, 3455, 3526, 3611, 3639, 3732]. Multiple-Block [1564]. Multiple-Relaxation [3611]. Multiple-Relaxation-Time [1172, 1696]. Multiple-Sets [1606]. Multiple-Subdomain [2555]. Multiplication [930]. Multiplicative [1057, 1307, 1676, 1994, 2050, 2569, 2618, 2868, 3047, 3128, 3666]. Multiplier [355, 1763, 1988, 3029]. Multiplier/Fictitious [355]. Multipliers [633, 725, 1551, 2348, 3038, 3368, 3673]. Multiply [1148, 1297, 1671, 1672, 2279]. Multiply-Connected [1297]. Multipoint [3304]. Multipoints [225]. Multipole [1838]. Multiquadric [198, 1039]. Multirate [655, 738, 1014, 1168, 1522, 3285, 3321]. Multiresolution [407, 826, 833, 837, 1520, 1530, 2677]. Multiscale [582, 609, 847, 993, 1117, 1415, 1432, 1528, 2149, 2378, 2547, 2673, 2795, 3075, 3114, 3119, 3151, 3175, 3316, 3335, 3338, 3371, 3406, 3568, 3627, 3658]. multischeme [45, 66]. Multistage [1662, 1663, 1715]. Multistep [106, 991, 1708, 2661, 2962, 3402]. Multisymplectic [482]. Multitude [3431]. Multivalued [395]. Multivariate [2148, 2544]. Multiwavelet [1348]. Multiwavelet-Based [1348]. Mumford [1660]. Müntz [3572]. MUSCL [267, 2200, 2849]. Myocardial [3243]. Mysteries [2544].

Naghdi [942, 1045, 1264]. Naghdi-Type [1045]. Nagumo [3723]. Nagumo-Type [3723]. Nano [886]. Nano-Optics [886]. Nanofluid [2615]. Nanomagnetic [2400]. Nanomagnets [1991]. Narrow [1081, 2067, 3017, 3020]. Narrow-Band [3017]. Narrow-Stencil [3020]. Nash [1452, 2266, 3436, 3505]. Natural [176, 1097, 1172, 1454, 1679, 1982, 2393, 2736, 3245, 3301, 3501]. Nature [577, 1034]. Navier [3, 36, 87, 101, 127, 138, 192, 227, 327, 343, 420, 426, 430, 520, 532, 597, 614, 702, 728, 772, 787, 788, 790, 805, 816, 853, 877, 988, 1056, 1146, 1174, 1235, 1304, 1312, 1349, 1350, 1355, 1386, 1406, 1417, 1519, 1526, 1542, 1559, 1619, 1640, 1756, 1768, 1805, 1843, 1854, 1909, 1931, 1972, 2002, 2023, 2048, 2088, 2137, 2161, 2185, 2186, 2360, 2477, 2482, 2490, 2499, 2558, 2563, 2623, 2696, 2755, 2763, 2803, 2820, 2890, 2916, 2932, 2979, 2994, 3025, 3178, 3182, 3185, 3227, 3232, 3252, 3255, 3343, 3367, 3401, 3421, 3434, 3489, 3501, 3565, 3575, 3582, 3663, 3703, 3715, 3735]. Navier [1069, 1971, 2859]. Naviers [694]. Naydenov [1777]. NCAR [491]. Near [1157, 3504, 3594]. Nearly [297, 458, 1180, 2240, 3427, 3515]. Necessary [1480, 3739]. Necrotic [1090]. Nédélec [2613]. NEFEM [2256]. Negative [637, 1017, 1128, 1269, 1516, 1710, 1746, 3461]. Negative-Order [1017, 1128]. Neighbor [3108]. Neighbor-Gradient [3108]. Neighborhood [2327, 2722]. Neighbors [3026]. Neither [705]. Nematic [1508, 1678, 1995]. Nernst [293, 1326, 1894, 2209, 2210, 2466, 2808, 2820, 3005, 3421, 3630]. Nested [136, 223, 2281, 3207]. Nesterov [3314]. net [3017]. Network [2157, 2622, 2686, 3053, 3162, 3203, 3311, 3365, 3453, 3464, 3473, 3700, 3711]. Networks [1392, 1670, 2638, 2652, 2829, 2919, 3017, 3115, 3138, 3189, 3216, 3239, 3251, 3268, 3282, 3342, 3406, 3446, 3532, 3620, 3642, 3643, 3677, 3706, 3733, 3756, 3795]. Neumann [1184, 1454, 1544, 1697, 2664, 2685]. Neural [2157, 2622, 2638, 2652, 2686, 2829, 2919, 3017, 3053, 3115, 3203, 3239, 3251, 3268, 3282, 3342, 3406, 3446, 3453, 3464, 3473, 3477, 3532, 3620, 3642, 3643, 3706, 3756]. Neuron [2571]. Neuroscience [3633]. Neutral [798, 3069]. Newton [164, 487, 752, 853, 927, 1201, 1400, 1779, 1815, 1910, 1972, 2059, 2118, 2422, 2655, 2713, 2904, 3027, 3114, 3135, 3203, 3233, 3391, 3397, 3400, 3462, 3558, 3585, 3610, 3639, 3651]. Newton-Based [3027]. Newton-Correction [1815, 3639]. Newton-Type [1972]. Newtonian [1514, 1610, 1728, 3132, 3631]. Next [2701, 3239]. Next-Generation [2701]. Nicolson [1303, 1356, 1376, 1512, 1537, 1682, 1825, 1879, 1887, 1892, 2044, 2344, 2602, 2774, 2890, 3252, 3399, 3678, 3734]. Nicolson-Leapfrog [3399]. Nitsche [895, 1343, 2463, 2503, 2827, 3119, 3491, 3600]. NMR [1041]. No [224, 845, 1619, 2456, 2533, 2958, 2968]. No-Collision [2533]. No-Slip [1619]. No-Slope-Selection [2456]. Nodal [821, 869, 996, 1832, 2120, 2946, 3204]. Node [1025, 1027, 3468]. Node-Centered [1027]. Nodes [1166]. Noise [310, 1187, 1248, 1542, 1575, 1676, 1985, 1994, 2050, 2059, 2091, 2309, 2426, 2493, 2569, 2577, 2627, 2754, 2868, 3004, 3076, 3128, 3411, 3528, 3589, 3666, 3808]. Noises [3287]. noising [1008]. Noisy [706, 1003, 3154, 3177, 3258, 3385]. Non [323, 403, 412, 448, 449, 528, 557, 595, 598, 606, 613, 675, 701, 716, 757, 775, 807, 870, 940, 941, 993, 997, 1015, 1030, 1077, 1156, 1266, 1320, 1364, 1383, 1448, 1449, 1474, 1530, 1548, 1559, 1584, 1643, 1700, 1703, 1712, 1746, 1819, 1825, 1889, 1900, 1923, 1934, 1957, 1979, 1983, 2022, 2028, 2057, 2063, 2098, 2205, 2281, 2339, 2361, 2369, 2385, 2402, 2408, 2410, 2430, 2450, 2479, 2483, 2609, 2685, 2720, 2737, 2791, 2792, 2817, 2849, 2862, 2921, 2927, 2935, 2940, 2954, 3006, 3019, 3068, 3087, 3110, 3196, 3277, 3279, 3304, 3322, 3377, 3461, 3514, 3549, 3555, 3601, 3631, 3635, 3811]. Non [9, 306, 543, 550, 647, 732, 867, 889, 928, 1108, 1136, 1338, 1345, 1395, 1453, 1511, 1537, 1558, 1561, 1594, 1666, 1669, 1727, 1767, 1884, 1971, 2163, 2208, 2253, 2405, 2422, 2454, 2478, 2550, 2578, 2621, 2642, 2687, 2804, 2812, 2869, 3024, 3064, 3070, 3119, 3142, 3178, 3357, 3379, 3423, 3467, 3487, 3560, 3573, 3598, 3604, 3653, 3674, 3776]. Non-backtracking [2430]. Non-Conforming [1666, 2057, 2205, 2208, 2253, 2369, 2935, 2940, 3514, 3776]. Non-Conservative [940, 2849, 2862, 3006]. Non-Convex [1669, 1923, 2817, 3178, 3604, 3811]. Non-differentiable [2479]. Non-divergence [2022, 2063, 3598, 3653]. Non-equilibrium [1712, 2361]. Non-Flat [2812]. Non-Graded [613, 1136]. Non-harmonic [1015]. Non-Hermitian [595]. Non-Homogeneous [701, 1548]. Non-hydrostatic [2339, 2609, 2737, 3601]. Non-Ideal [550]. Non-Intrusive [993, 2792, 3487]. Non-isothermal [2621]. non-Lagrange [867]. Non-Linear [528, 550, 557, 997, 1320, 1530, 1584, 2422, 2450, 2454, 2685, 3070, 3110, 3304, 3322, 3357, 3423, 3555]. Non-Lipschitz [3560]. Non-Local [543, 1395, 1643, 1700]. Non-Markovian [9, 2869]. Non-Matching [647, 1900, 3119]. Non-monotone [807, 1561]. Non-Monotonically [1727]. Non-Negative [1746, 3461]. Non-nested [2281]. Non-obtuse [2791]. Non-oscillation [1077]. Non-Oscillatory [675, 928, 1030, 1156, 1383, 1448, 1584, 1594, 1825, 1884, 1889, 1979, 2098, 2478, 2720, 2804, 3024, 3377]. Non-Overlapping [448, 1364, 1703, 2954, 3573, 3674]. Non-perturbative [3142]. Non-Polynomial [598, 1767]. Non-quadratic [2405]. Non-rectangular [1511]. Non-Reflecting [757]. Non-self-adjoint [2921]. Non-selfadjoint [1108, 2385]. Non-separable [3087]. Non-slip [1559]. Non-Smooth [716, 732, 870, 1266, 1345, 1934, 1957, 2402, 2408, 2642, 3019, 3277, 3549, 3811]. Non-Stabilized [449]. Non-staggered [3635]. Non-stationary [1537, 1983, 3467]. Non-Stiff [3064, 3279]. Non-strictly [1449]. Non-symmetric [775]. Non-Trivial [323]. Non-unified [2163]. Non-Uniform [403, 606, 889, 1338, 1453, 1474, 1558, 1819, 2028, 2410, 2550, 2687, 2927, 3068, 3379]. Nonclassical [1067, 2789]. Nonconforming [530, 768, 891, 1064, 1086, 1089, 1098, 1141, 1161, 1195, 1249, 1258, 1325, 1353, 1739, 1853, 1966, 2172, 2345, 2463, 2604, 2714, 2768, 2882, 2908, 3067, 3222, 3348, 3533, 3541, 3608, 3751, 3777]. Nonconservative [748, 944, 1083, 3184]. Nonconvergence [1213]. Nonconvex [536, 1360, 1549, 1857, 1872, 1985, 2032, 2077, 2258, 2297, 2695, 2750, 2791, 2963, 3009, 3136, 3641, 3662]. Nonergodic [2348]. Nonet [410]. Nonisometric [2758]. noniterative [135]. Nonlinear [237, 249, 282, 326, 383, 442, 586, 650, 682, 731, 818, 832, 848, 869, 871, 880, 922, 929, 942, 946, 979, 994, 1100, 1128, 1189, 1201, 1220, 1251, 1261, 1374, 1487, 1533, 1535, 1552, 1577, 1667, 1680, 1696, 1709, 1719, 1725, 1777, 1802, 1912, 1954, 2009, 2034, 2082, 2089, 2099, 2105, 2137, 2164, 2176, 2192, 2215, 2221, 2237, 2241, 2245, 2375, 2387, 2434, 2473, 2476, 2597, 2649, 2664, 2675, 2719, 2732, 2778, 2783, 2796, 2805, 2813, 2830, 2845, 2851, 2863, 2891, 2927, 2982, 3013, 3036, 3044, 3052, 3098, 3134, 3155, 3181, 3250, 3279, 3315, 3319, 3419, 3430, 3456, 3483, 3557, 3567, 3596, 3709, 3773, 3780]. Nonlinear [39, 63, 111, 127, 164, 174, 175, 359, 500, 501, 589, 593, 844, 851, 854, 918, 927, 1005, 1082, 1193, 1303, 1423, 1429, 1489, 1491, 1531, 1633, 1635, 1647, 1668, 1677, 1704, 1739, 1757, 1762, 1794, 1809, 1830, 1845, 1863, 1870, 1878, 1879, 1887, 1999, 2005, 2020, 2027, 2065, 2111, 2115, 2135, 2148, 2179, 2382, 2402, 2429, 2460, 2481, 2511, 2518, 2565, 2582, 2605, 2606, 2612, 2668, 2683, 2702, 2750, 2761, 2777, 2793, 2831, 2832, 2850, 2884, 2923, 2946, 2947, 2951, 2965, 3002, 3008, 3014, 3084, 3128, 3139, 3147, 3193, 3216, 3234, 3287, 3343, 3348, 3363, 3380, 3386, 3391, 3400, 3447, 3465, 3640, 3678, 3690, 3694, 3750, 3771]. Nonlinearities [1516, 1785]. Nonlinearity [3665]. Nonlinearly [1209]. Nonlocal [814, 1290, 1356, 1523, 1816, 1944, 2056, 2115, 2313, 2406, 2640, 2802, 2821, 2906, 3029, 3098, 3213, 3510, 3598, 3624, 3653]. Nonmatching [1089, 1419, 1424]. Nonmonotone [1162, 1318, 3523]. Nonnegative [2618, 2807, 2854, 3015, 3153, 3516]. Nonnested [3383]. Nonorthogonal [750]. Nonoverlapping [641, 1959, 2541]. Nonperiodic [144, 1301]. Nonpolar [1600]. Nonpositivity [1817]. Nonprimitive [1854]. Nonreflecting [197]. Nonrelativistic [1833, 3207]. Nonsingular [3734]. Nonsmooth [1173, 1318, 1360, 1751, 1857, 1913, 2258, 2418, 2518, 2527, 2739, 2777, 3009, 3270, 3314, 3449, 3546, 3662, 3735, 3784]. Nonsquare [2562]. Nonstandard [2241, 3407]. Nonstationary [2694, 3326]. nonsteady [3]. Nonsymmetric [626, 1134, 2460, 3236]. Nonuniform [956, 1622, 2437, 2491, 2645, 2702, 2942, 3249, 3296, 3481, 3650, 3800]. Nonuniformly [1847]. Nor [705]. Norm [450, 456, 1017, 1128, 1217, 1486, 1573, 1583, 1685, 1707, 1980, 2031, 2055, 2151, 2327, 2520, 2577, 2595, 2824, 2963, 3221, 3234]. Normal [1861, 2846, 3071, 3142, 3179, 3652]. Normal-Subdiffusion [3652]. Norms [462, 1028, 2785, 2830, 3536]. Note [81, 174, 239, 614, 747, 1176, 2112, 3333, 3469]. Notes [1670]. Notion [1440]. Novel [626, 681, 1002, 1021, 1109, 1341, 1422, 1530, 1655, 1713, 2310, 2681, 2703, 2730, 2769, 2793, 2965, 3094, 3102, 3551, 3695, 3726]. NPTC [3017]. NPTC-net [3017]. NRBC [2742]. NRxx [984]. Nuclear [223, 1862, 2031, 2577, 3234]. Nuclei [715]. NUFFT [1973]. Null [3644]. Number [58, 75, 133, 151, 159, 203, 204, 240, 275, 459, 489, 537, 670, 1236, 1576, 1869, 2064, 2212, 2543, 2596, 2598, 2701, 2781, 3707]. Numbers [18, 88, 311, 388, 2132, 2160, 3718, 3790]. Numerical [23, 25, 58, 92, 163, 202, 237, 268, 330, 402, 403, 472, 480, 520, 581, 680, 697, 723, 737, 829, 856, 870, 871, 923, 937, 938, 942, 980, 1011, 1014, 1145, 1151, 1163, 1170, 1209, 1299, 1309, 1375, 1381, 1446, 1505, 1507, 1521, 1613, 1680, 1687, 1723, 1802, 1805, 1812, 1820, 1833, 1869, 1875, 1967, 2048, 2178, 2181, 2195, 2206, 2233, 2314, 2341, 2375, 2389, 2456, 2514, 2561, 2587, 2591, 2609, 2640, 2656, 2680, 2690, 2696, 2699, 2733, 2909, 2926, 2947, 2969, 3083, 3208, 3209, 3226, 3304, 3440, 3476, 3493, 3531, 3535, 3588, 3637, 3724, 3747, 3768, 3802]. Numerical [101, 253, 278, 283, 289, 311, 326, 355, 380, 381, 452, 459, 481, 496, 502, 537, 568, 642, 643, 667, 670, 694, 696, 713, 818, 830, 840, 848, 878, 887, 952, 1059, 1100, 1111, 1183, 1198, 1232, 1320, 1347, 1358, 1408, 1457, 1470, 1517, 1560, 1562, 1572, 1606, 1768, 1772, 1777, 1816, 1821, 1845, 1892, 1939, 1998, 2042, 2083, 2110, 2134, 2149, 2215, 2217, 2227, 2262, 2279, 2328, 2329, 2332, 2342, 2347, 2494, 2522, 2594, 2627, 2633, 2710, 2711, 2821, 2829, 2900, 3030, 3129, 3188, 3235, 3357, 3388, 3431, 3458, 3479, 3480, 3483, 3525, 3618, 3619, 3630, 3635, 3694, 3702, 3725, 3737, 3775, 3805, 3810]. Numerical [10, 30, 32, 57, 69, 70, 97, 110, 113, 116, 121, 123, 133, 160, 161, 169, 170, 185, 186, 203, 216, 218, 222, 287, 306, 319, 390, 396, 407, 458, 576, 577, 582, 603, 610, 626, 677, 692, 715, 733, 734, 809, 820, 827, 852, 898, 964, 970, 978, 1071, 1112, 1184, 1253, 1271, 1292, 1331, 1352, 1380, 1417, 1468, 1479, 1511, 1600, 1636, 1646, 1686, 1705, 1884, 1905, 1980, 2018, 2039, 2113, 2160, 2322, 2377, 2395, 2457, 2469, 2481, 2563, 2565, 2621, 2647, 2738, 2866, 2886, 2951, 2975, 3086, 3109, 3139, 3213, 3226, 3289, 3345, 3428, 3450, 3528, 3534, 3677, 3779, 3801]. Numerically [382, 705, 1423, 2179, 2810]. Numerics [2654]. Numerov [2054, 2998, 3145]. Nunziato [884]. Nyström [249].

O [475]. Oberbeck [1692]. Object [242]. Object-Oriented [242]. objective [2266]. Objectives [2923, 3430]. Objects [246, 260, 2264, 2265]. Observability [624]. Observables [820]. Observation [2846]. Observations [2695]. Observed [2869]. Obstacle [418, 1109, 1914, 2517, 2584, 2716]. Obstacles [1272, 1784, 2319, 3660, 3699]. Obtain [1027]. Obtained [3513]. obtuse [2791]. Ocean [172, 229, 408, 416, 435, 1699, 2351, 2665]. Octree [1136]. Odd [775, 812]. ODE [3564]. ODEs [477, 1598, 1646, 2127, 3477]. Oedema [3243]. Off [415, 3054, 3348]. Offline [1933, 2493, 2638]. Offline-Enhanced [1933]. Offline-Free [2493]. Ohta [2439]. Oil [618, 1777]. Oldroyd [328]. Oleg [1777]. Once [2582, 2840]. One [28, 71, 119, 280, 369, 404, 439, 555, 559, 594, 634, 635, 667, 735, 792, 800, 802, 820, 824, 1076, 1083, 1156, 1261, 1262, 1271, 1286, 1287, 1414, 1453, 1528, 1533, 1773, 1823, 1875, 1920, 1926, 1950, 1978, 2099, 2121, 2192, 2287, 2288, 2308, 2327, 2333, 2336, 2397, 2496, 2509, 2589, 2729, 2761, 2771, 2897, 3069, 3076, 3086, 3263, 3289, 3478, 3638]. One-Dimension [3086]. One-Dimensional [28, 280, 369, 594, 635, 667, 800, 820, 1076, 1083, 1156, 1261, 1271, 1528, 1875, 1920, 1950, 1978, 2099, 2121, 2496, 2729, 3076, 3478, 3638]. One-Dimensional/Two-Dimensional [2496]. One-Norm [2327]. One-Parameter [1773, 3069]. one-pass [119]. One-Phase [71, 2771]. One-Sided [559, 735, 1453]. One-step [1926]. One-Way [2336, 3289]. Online [2547, 2638, 3585]. Only [2657]. Ono [3548]. onto [1671, 1672, 3610]. Open [933, 1048, 1506, 1777, 3779]. Open-Source [3779]. OpenCL [1964].

Operator [77, 137, 143, 233, 449, 624, 875, 1029, 1177, 1367, 1455, 1487, 1515, 1811, 1840, 1967, 1986, 2058, 2170, 2222, 2280, 2364, 2435, 2545, 2727, 2760, 2831, 2869, 3040, 3189, 3240, 3257, 3289, 3319, 3329, 3454, 3500, 3642]. Operator-integration-factor [77]. Operator-Splitting [2328, 2329, 3040, 3240]. Operators [56, 88, 159, 199, 361, 363, 364, 400, 461, 472, 632, 724, 794, 814, 872, 994, 1038, 1042, 1060, 1249, 1276, 1325, 1462, 1496, 1620, 1707, 2033, 2047, 2057, 2067, 2128, 2190, 2366, 2579, 2701, 2821, 2961, 3130, 3199, 3344, 3392, 3407, 3535, 3602, 3639]. Optical [806, 1359, 2192, 3454]. Optics [375, 886, 1655, 2913, 3058]. Optima [2009]. Optimal [495, 534, 535, 586, 612, 652, 672, 719, 740, 754, 774, 793, 815, 822, 829, 861, 899, 992, 1009, 1045, 1096, 1173, 1242, 1251, 1267, 1288, 1293, 1300, 1461, 1485, 1535, 1554, 1555, 1573, 1625, 1682, 1738, 1866, 1876, 1879, 1898, 1907, 1915, 1923, 1953, 1975, 1991, 1995, 2041, 2093, 2135, 2155, 2182, 2199, 2259, 2261, 2288, 2304, 2318, 2348, 2366, 2387, 2426, 2435, 2442, 2444, 2451, 2491, 2575, 2583, 2597, 2608, 2786, 2798, 2826, 2835, 2894, 2929, 2952, 2964, 3000, 3025, 3076, 3089, 3113, 3114, 3151, 3183, 3226, 3242, 3292, 3301, 3411, 3421, 3422, 3504, 3575, 3613, 3618, 3633, 3634, 3670, 3678, 3685, 3735, 3775, 3782]. Optimal [116, 775, 795, 1132, 1262, 1588, 1754, 1989, 2267, 2363, 2367, 2419, 2544, 2573, 2616, 2643, 2688, 2759, 2811, 2837, 2896, 2959, 2979, 3071, 3131, 3308, 3362, 3453, 3463, 3584, 3606]. Optimal-Order [3575]. Optimality [812, 1141, 1413, 1915]. Optimally [1927, 3655]. Optimisation [468]. Optimization [76, 208, 234, 404, 415, 604, 807, 1004, 1020, 1126, 1248, 1285, 1309, 1360, 1388, 1405, 1746, 1758, 1857, 1872, 2053, 2187, 2258, 2425, 2479, 2507, 2562, 2607, 2632, 2669, 2695, 2708, 2713, 2722, 2750, 2797, 2871, 2972, 2973, 3042, 3133, 3136, 3196, 3200, 3314, 3361, 3414, 3418, 3433, 3462, 3532, 3641, 3686, 3693, 3771, 3778, 3784, 3811]. Optimization-Based [1746]. Optimizations [1794, 3261]. Optimized [218, 514, 632, 1170, 1526, 1552, 1877, 2200, 2367, 2593, 2720, 2735, 2935]. Optimizer [3261]. Option [1214, 1503, 2060, 2097, 2610]. Options [654, 755, 759, 824, 852, 1063, 1242, 2096, 2545, 3317, 3649, 3731]. Orbits [401, 413]. Order [244, 268, 343, 373, 422, 440, 480, 485, 516, 525, 552, 598, 624, 662, 663, 679, 748, 782, 816, 867–869, 923, 967, 972, 1018, 1118, 1147, 1159, 1161, 1192, 1223, 1274, 1277, 1286, 1287, 1299, 1314, 1323, 1332, 1382, 1424, 1446, 1496, 1519, 1525, 1546, 1566, 1613, 1645, 1665, 1687, 1689, 1707, 1724, 1750, 1760, 1860, 1874, 1890, 1917, 1919, 2069, 2133, 2136, 2187, 2192, 2212, 2232, 2364, 2442, 2456, 2458, 2475, 2530, 2540, 2584, 2640, 2648, 2721, 2796, 2809, 2828, 2939, 3052, 3123, 3126, 3166, 3199, 3218, 3225, 3333, 3373, 3390, 3566, 3575, 3603, 3608, 3722, 3790]. Order [202, 251, 254, 282, 318, 326, 330, 339, 483, 486, 505, 528, 529, 533, 565, 586, 632, 664, 731, 739, 762, 772, 862, 865, 932, 939, 944, 1054, 1076, 1077, 1113, 1155, 1191, 1205, 1261, 1296, 1348, 1355, 1464, 1486, 1522, 1568, 1576, 1622, 1649, 1667, 1695, 1702, 1772, 1791, 1820, 1843, 1858, 1882, 1899, 1902, 1925, 1927, 1962, 2022, 2023, 2045, 2082, 2098, 2227, 2241, 2261, 2288, 2312, 2335, 2366, 2528, 2566, 2613, 2626, 2675, 2728, 2771, 2783, 2839, 2891, 2911, 2978, 3010, 3030, 3033, 3062, 3228, 3229, 3237, 3279, 3313, 3397, 3425, 3514, 3527, 3531, 3545, 3567, 3576, 3627, 3682, 3780, 3799]. Order [275, 278, 281, 283, 357, 368, 519, 589, 730, 778, 843, 883, 977, 987, 1017, 1027, 1079, 1128, 1140, 1228, 1269, 1276, 1284, 1368, 1430, 1438, 1457, 1460, 1465, 1502, 1503, 1562, 1584, 1587, 1660, 1748, 1759, 1765, 1794, 1801, 1821, 1840, 1937, 1981, 2015, 2016, 2046, 2063, 2134, 2177, 2197, 2251, 2356, 2361, 2368, 2397, 2406, 2437, 2447, 2471, 2513, 2542, 2606, 2671, 2677, 2719, 2722, 2740, 2792, 2816, 2826, 2831, 2833, 2850, 2862, 2890, 2900, 2931, 2933, 2942, 3003, 3022, 3029, 3044, 3096, 3101, 3122, 3129, 3306, 3316, 3326, 3328, 3388, 3430, 3446, 3457, 3466, 3472, 3492, 3496, 3498, 3533, 3612, 3614, 3620, 3630, 3638, 3650, 3690, 3723, 3725]. Order [239, 277, 306, 311, 315, 316, 320, 366, 407, 544, 580, 583, 605, 607, 648, 656, 673, 728, 764, 799, 815, 921, 950, 1012, 1040, 1044, 1080, 1095, 1098, 1143, 1166, 1195, 1222, 1291, 1336, 1380, 1494, 1531, 1543, 1558, 1591, 1609, 1632, 1646, 1651, 1735, 1762, 1780, 1796, 1806, 1842, 1945, 1965, 1974, 1982, 2051, 2092, 2094, 2115, 2121, 2140, 2154, 2156, 2274, 2330, 2374, 2446, 2492, 2501, 2567, 2586, 2589, 2608, 2616, 2639, 2693, 2793, 2814, 2843, 2873, 2886, 2893, 2897, 2929, 2970, 3032, 3265, 3275, 3293, 3298, 3300, 3307, 3320, 3387, 3398, 3423, 3529, 3698, 3781, 3794, 3806]. Order [302, 348, 432, 468, 471, 547, 548, 550, 579, 582, 736, 820, 890, 897, 904, 917, 981, 1043, 1051, 1071, 1093, 1110, 1133, 1208, 1233, 1238, 1328, 1346, 1354, 1373, 1442, 1477, 1534, 1574, 1605, 1615, 1616, 1619, 1623, 1666, 1690, 1732, 1754, 1826, 1871, 1884, 1976, 2005, 2012, 2068, 2086, 2103, 2122, 2141, 2151, 2162, 2189, 2208, 2226, 2303, 2320, 2321, 2352, 2370, 2377, 2382, 2429, 2440, 2478, 2497, 2548, 2588, 2601, 2683, 2688, 2718, 2738, 2785, 2799, 2812, 2866, 2986, 3034, 3054, 3088, 3127, 3131, 3187, 3266, 3331, 3367, 3381, 3384, 3434, 3443, 3526, 3548, 3550, 3552, 3562, 3587, 3590, 3607, 3666, 3727, 3745, 3750, 3772, 3776, 3797]. Order [44, 86, 88, 142, 160, 188, 193, 444, 501, 502, 638, 681, 732, 801, 821, 971, 999, 1799, 2525, 2777, 2800, 2819, 2853, 2953, 3057, 3067, 3110, 3215, 3276, 3296, 3308, 3426, 3428, 3485]. Order- [3799]. Ordered [1025, 1661, 2085]. Orders [882, 3247]. Ordinary [479, 744, 1051, 1313, 1322, 1366, 1725, 1898, 1926, 2611, 3473, 3698]. Ordinates [806]. Oriented [242, 1050, 1481, 1626, 1835, 2371, 2453, 2587, 3394]. Origins [2460]. Orthogonal [706, 1023, 1356, 1512, 2112, 2142, 2662, 2834, 2907, 3081, 3125, 3155, 3326, 3335, 3355, 3466, 3477]. Orthogonality [105, 1243, 1857, 2425]. Orthogonalization [3325]. Orthogonalization-Free [3325]. Orthomin [178]. Orthonormal [998]. OSC [2162]. Oscillating [222, 397]. Oscillation [1077, 3260, 3512, 3762, 3769]. Oscillation-Free [3260, 3512, 3762, 3769]. Oscillations [2684, 3643]. Oscillator [1837, 3378]. Oscillatory [675, 928, 1030, 1101, 1117, 1156, 1383, 1448, 1584, 1594, 1825, 1884, 1889, 1979, 2098, 2478, 2720, 2804, 2851, 3024, 3071, 3377, 3378, 3486, 3488]. Oseen [1152, 1501, 1556, 1742, 2232, 2413, 2436, 2581, 3171]. Oseen-Type [1742]. Osher [386, 940]. osmotic [1326]. Other [98, 2092]. Out-of-Core [995]. Outflow [280]. Outlier [1186]. Outliers [3589]. Output [1828, 3091]. Outputs [2705]. Over-Collocation [3423]. Over-Penalized [767, 916, 1237, 1267, 1435]. Overapproximating [387]. Overcoming [1923]. Overintegration [2120]. Overlap [2624]. Overlapping [448, 518, 750, 1364, 1409, 1703, 1877, 2368, 2470, 2954, 3573, 3674, 3725]. Oversampling [2547]. Overset [583, 3297, 3404]. p [2520]. Padé [567, 594, 2532, 2579]. Padé-type [2579]. PageRank [1199, 2430, 3457, 3761]. Pair [3303, 3352]. Pairs [1493]. Pairwise [3324]. Pantograph [1104]. Papers [3499, 3508]. Parabolic [117, 126, 212, 213, 332, 503, 568, 584, 726, 832, 894, 943, 978, 1067, 1122, 1173, 1261, 1356, 1363, 1408, 1461, 1553, 1567, 1597, 1615, 1635, 1680, 1690, 1698, 1719, 1739, 1762, 1763, 1809, 1878, 1879, 1934, 2025, 2043, 2124, 2134, 2241, 2245, 2344, 2375, 2383, 2402, 2470, 2639, 2681, 2691, 2702, 2719, 2775, 2777, 2814, 2850, 2874, 2880, 2952, 3000, 3092, 3164, 3221, 3288, 3322, 3380, 3472, 3491, 3570, 3596, 3779]. Paradox [1450]. Parallel [6, 21, 45, 52, 62, 66, 82, 108, 114, 128, 135, 136, 141, 152, 164, 167, 169, 172, 201, 205, 217, 223, 229, 237, 279, 298, 530, 583, 751, 811, 832, 853, 927, 964, 968, 1061, 1236, 1310, 1328, 1482, 1500, 1721, 1742, 1746, 1818, 1888, 1964, 2037, 2088, 2094, 2116, 2270, 2511, 2641, 2706, 2730, 2734, 2912, 2920, 2982, 2994, 2998, 3017, 3021, 3032, 3144, 3145, 3227, 3230, 3450, 3581, 3753, 3759]. Parallel-in-Time [2998, 3032, 3145]. Parallelism [408, 416]. Parallelization [1783, 2415]. parallelized [186]. Parallelizing [126]. Parameter [204, 333, 367, 415, 665, 1366, 1389, 1773, 1822, 1962, 2001, 2009, 2106, 2497, 2549, 2852, 2996, 3050, 3069, 3259, 3431, 3788]. Parameter-Dependent [3050]. Parameter-Independent [2852]. Parameterisation [496]. Parameterization [1506, 2992]. Parameterizations [1341, 1906, 2310, 3194, 3713]. Parameterized [1817, 2452, 2792, 3073, 3218, 3333, 3366, 3492]. Parameters [736, 861, 1758, 2050, 2507, 2694, 2712, 2826, 2935, 3283, 3387, 3618]. Parametric [599, 673, 1386, 1694, 1897, 2813, 2829, 3173, 3342, 3385, 3524]. Parametrization [3446]. Parametrized [988, 992, 1228, 1616, 1961, 2041, 2454, 2616, 2783, 2931, 3319]. Parareal [1598, 2324, 2671]. Parareal-Euler [1598]. Paraxial [646]. Pareto [2266]. Parsimonious [3188]. Part [439, 520, 649, 729, 1257, 1286, 1287, 1372, 1516, 1550, 2264, 2265, 2544, 2698, 3226]. Partial [90, 136, 164, 173, 202, 238, 256, 264, 302, 333, 398, 562, 731, 741, 756, 832, 1068, 1086, 1206, 1218, 1286, 1287, 1352, 1518, 1530, 1635, 1665, 1694, 1815, 1821, 1874, 1943, 1962, 2039, 2043, 2094, 2165, 2216, 2235, 2324, 2354, 2367, 2375, 2391, 2404, 2464, 2517, 2654, 2662, 2700, 2793, 2801, 2919, 2947, 2955, 3034, 3157, 3218, 3282, 3303, 3320, 3333, 3342, 3352, 3458, 3492, 3510, 3529, 3628, 3639, 3648, 3708, 3756, 3782, 3788]. Partially [746, 2869, 2880, 3457, 3693]. Particle [7, 168, 338, 422, 424, 527, 713, 719, 898, 1144, 1178, 1431, 1463, 1478, 2354, 2507, 2704, 2787, 3078, 3121, 3444]. Particle-in-Cell [168, 424, 3121]. Particle-Method [719]. Particle-without-Particle [2354]. Particles [2400]. Particular [1304]. Partition [404, 612, 1296, 1433, 1482, 1617, 2088, 2340, 2630, 2724, 3779]. Partition-of-Unity [3779]. Partitioned [1322, 1407, 2122, 2343, 3353]. Partitioning [191, 3761]. Partitions [587, 1673]. Parts [363, 400, 461, 632, 872, 994, 1038, 1042, 1707, 1798, 2003, 2033, 2086, 2205, 2394, 2528, 2537, 2559, 2767, 2789, 2815, 3020, 3127, 3238, 3279, 3298, 3344, 3714]. Pass [119, 2917, 3108]. Passage [2373, 2878]. Passes [3707]. Passive [166, 247, 393, 875]. Past [8, 544, 605, 903, 2342]. Patch [1575, 3004].

Patch-Based [1575, 3004]. Patches [2630]. Path [83, 948, 952, 2823, 2943, 3073, 3223]. Path-Conservative [952, 2823, 2943, 3223]. Path-Integral [948]. Pathfinding [2002]. Paths [969]. Patient [3387]. Patient-Specific [3387]. Patlak [1105]. Pattern [469, 711, 2924]. Patterning [3711]. Patterns [146, 285, 397, 2302, 3468]. PCA [3585]. PCBC [2202]. PDAEs [2654]. PDE [95, 117, 586, 592, 897, 987, 1096, 1193, 1388, 1485, 1588, 1704, 2318, 2460, 2887, 3324, 3778]. PDE-Constrained [1388, 1485, 3778]. PDE’s [128, 221, 396, 505, 588, 696, 732, 881, 918, 978, 1155, 1201, 1239, 1373, 1439, 1553, 1643, 1667, 1705, 1736, 1937, 1974, 2022, 2092, 2097, 2148, 2149, 2159, 2215, 2288, 2340, 2379, 2383, 2428, 2452, 2455, 2468, 2667, 2783, 2800, 2805, 2847, 2863, 2870, 2874, 2923, 2960, 3051, 3075, 3138, 3149, 3155, 3173, 3188, 3273, 3275, 3299, 3319, 3369, 3398, 3543, 3583, 3618, 3725, 3728]. PDFs [546]. PDHG [3409]. Peaceman [212, 2473]. Peano [695]. Peer [1459, 2073]. Penalization [3064]. Penalized [767, 916, 1237, 1267, 1410, 1435, 2015, 2428, 2716, 2880]. Penalties [732, 872, 2086, 2528]. Penalty [117, 440, 608, 654, 658, 707, 767, 774, 797, 808, 839, 861, 916, 989, 1074, 1113, 1237, 1267, 1296, 1349, 1417, 1431, 1563, 1576, 1640, 1723, 1779, 2022, 2096, 2167, 2197, 2297, 2363, 2433, 2581, 2677, 2872, 3211, 3329, 3443, 3735, 3784]. Penalty-Free [3443]. Pencils [2562, 3081]. Perfect [2888]. Perfectly [331, 475, 1103, 2911]. Performance [116, 350, 777, 782, 1062, 1590, 1745, 1746, 1794, 2415, 2764]. Performances [215]. Peridynamic [2138, 2794]. Peridynamics [1984]. Periodic [41, 209, 238, 888, 1253, 1452, 1656, 2080, 2378, 2414, 3299, 3691, 3709]. Periodically [560]. Periodicity [3159]. Periodizing [626]. Perspective [2117, 3187]. Perturbation [107, 194, 423, 703, 791, 888, 2156, 2816]. Perturbations [2155]. perturbative [3142]. Perturbed [757, 996, 1317, 1347, 1401, 1858, 1875, 1908, 2127, 2167, 2526, 2653, 2953, 3051, 3152, 3160, 3221, 3235, 3266, 3536]. Peter [951]. Petrov [661, 664, 1491, 1999, 2418, 2759, 3411]. Petviashvilli [1527]. PFC [2900]. Phase [71, 177, 371, 423, 615, 618, 676, 683, 715, 720, 792, 809, 849, 1053, 1125, 1231, 1281, 1370, 1386, 1420, 1524, 1557, 1712, 1718, 1772, 1788, 1888, 1895, 2007, 2018, 2083, 2152, 2226, 2333, 2393, 2395, 2614, 2727, 2756, 2771, 2786, 2866, 2926, 2931, 2950, 3006, 3041, 3174, 3219, 3300, 3345, 3373, 3382, 3399, 3452, 3501, 3527, 3560, 3597, 3621, 3637, 3674, 3698, 3711, 3722, 3760]. Phase-Change [2393]. Phase-Field [715, 1370, 2395, 2931, 3174, 3300, 3382, 3711, 3722]. Phase-Lagging [2083]. Phase-Space [1895]. phase/Multi [2506]. Phases [2950]. Phenomena [201, 381, 570, 1777]. Phenomenological [73]. Phenomenon [429, 901, 1474]. Phonon [1359]. Phononic [3774]. photolithographic [92]. Photonic [563, 1157, 1771, 2592, 2773, 3230]. Physical [820, 964, 1072, 3356, 3358, 3387, 3597]. Physical-Constraint-Preserving [3597]. Physically [2909]. Physics [250, 960, 1507, 1804, 2350, 3239, 3251, 3706]. Physics-Based [2350]. Physics-Informed [3239, 3706]. Picard [2375, 3207, 3450]. PIDEs [3317].

Piece [209]. Piece-Wise [209]. Piecewise [271, 385, 564, 606, 828, 1072, 1509, 1782, 2071, 2102, 2144, 2188, 2744, 2791, 2885, 3069]. Piecewise-Smooth [1782]. Piezoelectric [1784, 2471]. Pipe [112, 123, 240, 521, 1557, 2236]. Pipeline [3446]. Pipes [618, 941]. Pitaevskii [1262, 3027]. Planar [1918]. Planck [293, 743, 1326, 1339, 1378, 1677, 1894, 2020, 2209, 2210, 2466, 2594, 2740, 2808, 2820, 2841, 3005, 3109, 3421, 3630, 3675, 3688]. Planck/Navier [2820, 3421]. Plane [856, 2414, 2641]. Plane-Wave [2414]. Planets [3142]. Planning [2991]. Plasma [575, 797, 1059, 1111, 1311, 1507, 3103, 3555, 3679]. Plasmas [2211]. Plasmon [2156]. Plate [8, 112, 337, 438, 967, 1398, 1571, 2062, 2126, 2129, 2836, 2883, 3067, 3098]. Plates [547, 1435, 1720, 1729, 2268, 3725]. platform [152]. Play [3589]. Plug [3589]. Plug-and-Play [3589]. Plus [2473, 3248]. PML [557, 2174, 3012]. PMLs [782, 2954]. Pock [3332]. POD [277, 1962, 2451, 2455, 2616]. POD-Based [2451]. POD/DEIM [1962]. Poincaré [660, 1117]. Point [606, 686, 703, 785, 791, 833, 836, 855, 924, 977, 1130, 1200, 1262, 1334, 1450, 1457, 1609, 1669, 1717, 1743, 1747, 1781, 1908, 1989, 2116, 2119, 2202, 2327, 2473, 2479, 2616, 2632, 2660, 2722, 2723, 2762, 2787, 2871, 2945, 2992, 3011, 3017, 3043, 3073, 3135, 3172, 3210, 3336, 3340, 3433, 3532, 3537, 3602]. Point-Source [1609]. Point-Value [833]. Point-Wise [1262]. Points [258, 275, 1005, 1834, 2672, 2877, 2928, 3573, 3639]. Pointwise [899, 1293, 1577, 2161, 2791, 3102, 3117, 3131, 3775]. Pointwise-in-Time [3102]. Poisson [124, 155, 167, 183, 251, 280, 291, 293, 336, 341, 593, 601, 613, 796, 912, 931, 972, 1002, 1011, 1021, 1050, 1110, 1181, 1248, 1596, 1697, 1733, 1780, 1806, 1831, 1894, 2091, 2209, 2210, 2466, 2728, 2729, 2748, 2788, 2808, 2820, 2822, 2894, 3005, 3103, 3178, 3201, 3247, 3307, 3421, 3630]. Polar [868, 1174]. Polarized [423]. Pole [3258]. Poles [1137]. Policy [1242]. Pollutant [527, 701]. Poly [2818]. Poly-Sinc [2818]. Polyak [2915]. Polycrystal [1394]. Polydisperse [2692]. Polygonal [440, 1113, 1272, 1738, 1809, 3730]. Polygons [2254]. Polyharmonic [914]. Polyhedra [2791]. Polyhedral [1372, 1462, 1590, 1673, 3436, 3505]. Polyhedron [1861, 3610]. Polynomial [51, 117, 144, 197, 219, 324, 494, 598, 775, 812, 901, 944, 1024, 1072, 1301, 1316, 1552, 1589, 1767, 1813, 2079, 2132, 2146, 2188, 2201, 2380, 2871, 3200, 3204, 3468, 3530, 3617, 3620]. Polynomial-Based [1552]. Polynomial-Iteration [2871]. Polynomials [184, 218, 535, 1023, 1256, 1442, 1775, 2049, 2098, 2597, 2689, 2907, 2929, 3097, 3436, 3505, 3572, 3617, 3716]. Polytopal [2289]. polytopes [184]. Polytopic [2233, 2281, 3001]. Pontryagin [1795]. Population [656]. Populations [2701]. Poroelastic [3448]. Poroelasticity [3197, 3243, 3246, 3375]. Poroelasticity/Chemotaxis [3243]. Porosity [2343, 2994, 3489]. Porous [108, 733, 1163, 1259, 1814, 1901, 1948, 2027, 2152, 2257, 2358, 2393, 2416, 2602, 2630, 2733, 2771, 2879, 3061, 3068]. Portfolios [2077]. Portioned [920]. Posed [828, 1188, 1389, 1432, 1804, 2316, 2864, 3037, 3134, 3235, 3371, 3739]. Posedness [87, 99, 2712, 2772, 3631]. Position [1453]. Position-Dependent [1453]. Positive [670, 1026, 1510, 2046, 2377, 2556, 2808, 3018, 3027, 3607, 3625, 3723]. Positive-Type [1510]. Positivity [261, 981, 1197, 1616, 1632, 1645, 1647, 1722, 1792, 1830, 1963, 1981, 2251, 2291, 2361, 2368, 2466, 2516, 2668, 2710, 2812, 2903, 2977, 3185, 3237, 3262, 3367, 3434, 3442, 3460, 3563, 3588, 3617, 3630, 3743, 3781]. Positivity- [1981]. Positivity-Preserving [261, 981, 1197, 1632, 1645, 1722, 1792, 1830, 2368, 2516, 2710, 2812, 2903, 2977, 3262, 3367, 3434, 3442, 3460, 3588, 3743, 3781]. Positron [1008]. Post [526, 559, 1060, 2865, 3077]. Post-processed [3077]. Post-Processing [526, 559, 1060]. Post-processors [2865]. Posteriori [273, 295, 442, 443, 450, 456, 464, 647, 652, 726, 729, 767, 778, 834, 909, 933, 1026, 1035, 1048, 1162, 1274, 1287, 1294, 1362, 1434, 1454, 1461, 1466, 1489, 1602, 1664, 1698, 1728, 1848, 1954, 2074, 2176, 2196, 2290, 2305, 2344, 2358, 2441, 2600, 2613, 2619, 2635, 2646, 2652, 2716, 2751, 2883, 2905, 2952, 3005, 3034, 3051, 3092, 3095, 3117, 3272, 3278, 3288, 3312, 3317, 3341, 3401, 3467, 3503, 3587, 3600, 3638, 3658, 3712, 3719]. Postprocessing [784, 1146, 1601, 1729, 3054, 3380, 3636]. Potential [214, 406, 2085, 2341, 2889, 3460, 3588]. Potentials [364, 573, 820, 1080, 1084, 1095, 1105, 1476, 2252, 2389, 2667, 3016, 3335, 3518, 3709]. Power [46, 580, 1490, 1723, 2323, 2830, 3457]. Power-Law [1723]. Powers [199, 2579, 3130, 3158]. PPE [1619]. PPR [2303]. Practical [24, 645, 1485, 1657, 2354, 3093, 3475]. Prandtl [204]. Preasymptotic [1576, 2781]. Precision [283, 340, 1183, 2112, 3160, 3739, 3749]. Precomputed [1025]. Preconditioned [313, 640, 1092, 1250, 2058, 2096, 2294, 2324, 2443, 2857, 2923, 3087, 3257, 3409]. Preconditioner [1247, 1668, 1760, 2421, 2461, 2691, 2921, 3175, 3343, 3504, 3687]. Preconditioners [33, 143, 448, 451, 708, 760, 896, 916, 993, 1295, 1532, 1585, 1850, 1959, 2183, 2216, 2366, 2534, 2746, 2757, 2982, 3195, 3217, 3256]. Preconditioning [88, 159, 178, 233, 474, 493, 592, 1050, 1203, 1552, 1617, 1846, 1852, 2255, 2582, 2733, 2741, 2840, 3624]. PredGAN [3374, 3470]. Prediction [34, 562, 719]. Predictive [2990, 3374, 3470]. Predictor [3433, 3467]. Preface [463, 611, 678, 710, 838, 935, 958, 1073, 1116, 2448, 2488]. Preimage [2279]. Preinvexity [3430]. Preliminary [602]. Premixed [28, 231, 236, 329]. Prescribed [3277]. Presence [1839, 2090, 3385, 3506, 3555]. Preservation [1884, 2217, 2369, 2853, 3169, 3237, 3566]. Preserving [314, 361, 389, 425, 483, 534, 667, 739, 749, 900, 926, 945, 982, 1017, 1170, 1228, 1300, 1305, 1359, 1361, 1368, 1421, 1471, 1492, 1627, 1645, 1662, 1663, 1769, 1792, 1796, 1800, 1813, 1828–1830, 1834, 1850, 1925, 1931, 1939, 1981, 2020, 2057, 2073, 2139, 2155, 2179, 2251, 2299, 2361, 2368, 2380, 2407, 2411, 2438, 2449, 2465, 2466, 2502, 2516, 2543, 2572, 2592, 2593, 2644, 2668, 2710, 2743, 2802, 2831, 2894, 2933, 2984, 3010, 3044, 3121, 3147, 3180, 3181, 3185, 3191, 3220, 3248, 3262, 3265, 3285, 3315, 3402, 3406, 3442, 3460, 3512, 3524, 3541, 3563, 3588, 3630, 3671, 3680, 3682, 3688, 3704, 3769, 3785, 3790]. Preserving [119, 148, 261, 981, 1197, 1365, 1468, 1605, 1616, 1632, 1647, 1722, 1809, 1868, 2382, 2508, 2621, 2650, 2812, 2896, 2903, 2944, 2962, 2977, 2986, 3054, 3088, 3096, 3174, 3253, 3358, 3367, 3381, 3434, 3471, 3488, 3526, 3552, 3586, 3590, 3597, 3607, 3743, 3781, 3791]. Pressure [143, 224, 313, 934, 1019, 1312, 1410, 1501, 2007, 2399, 2413, 2436, 2477, 2914, 3064, 3167, 3205, 3401, 3601, 3793]. Pressure-Correction [2399]. Pressure-Robust [2477, 2914, 3205, 3401, 3793]. Pressure-Stabilized [1501, 2413]. Pressure-Tight [3064]. Prestige [1777]. Price [755, 759, 824]. Pricing [654, 1063, 1214, 1503, 2060, 2096, 2097, 2610, 3317, 3731]. Primal [770, 892, 1637, 1731, 1914, 2045, 2114, 2170, 2290, 2632, 2741, 2797, 3164, 3254, 3332, 3501, 3542, 3646]. Primal-Dual [1637]. Primary [2062]. Primitive [370, 874, 2591]. Principal [2689, 2963]. Principle [139, 356, 981, 1228, 1389, 1464, 1605, 1787, 1795, 1816, 2368, 2650, 2743, 2799, 2802, 2956, 3220, 3253, 3292, 3304, 3552]. Principle-Satisfying [1464]. Prior [1289, 1523, 1786, 3004, 3035, 3623, 3662]. Priori [333, 500, 608, 740, 1436, 1498, 1553, 1588, 1728, 2419, 2643, 2652, 2716, 2786, 3095, 3503, 3634, 3709, 3728]. Priors [682]. Prism [1161]. Probabilistic [832, 1217, 2564, 3176]. Probabilities [214]. Probability [73, 783]. Probe [1930]. Problem [297, 345, 370, 377, 383, 418, 495, 496, 517, 520, 578, 612, 703, 719, 791, 793, 908, 934, 972, 1052, 1053, 1087, 1141, 1147, 1161, 1182, 1293, 1356, 1401, 1454, 1490, 1502, 1508, 1548, 1583, 1599, 1631, 1639, 1683, 1713, 1730, 1802, 1814, 1853, 1875, 1918, 1934, 1983, 2004, 2023, 2025, 2088, 2125, 2137, 2191, 2195, 2199, 2232, 2262, 2270, 2271, 2289, 2290, 2293, 2305, 2462, 2521, 2584, 2590, 2619, 2655, 2674, 2681, 2716, 2724, 2740, 2810, 2816, 2855, 2864, 3001, 3067, 3117, 3118, 3149, 3151, 3161, 3205, 3272, 3292, 3521, 3538, 3544, 3546, 3629, 3632, 3661, 3712, 3724, 3726, 3737, 3792, 3799]. Problem [22, 95, 121, 205, 291, 321, 450, 541, 708, 1019, 1029, 1081, 1106, 1199, 1247, 1343, 1393, 1405, 1410, 1427, 1489, 1504, 1516, 1527, 1556, 1588, 1595, 1623, 1686, 1743, 1744, 1808, 1826, 1867, 1907, 1914, 2059, 2061, 2087, 2094, 2150, 2174, 2257, 2269, 2304, 2307, 2318, 2357, 2377, 2385, 2405, 2413, 2419, 2517, 2562, 2604, 2646, 2669, 2676, 2694, 2707, 2712, 2799, 2837, 2879, 2892, 2914, 3059, 3094, 3095, 3177, 3178, 3216, 3221, 3222, 3276, 3278, 3327, 3328, 3345, 3379, 3411, 3437, 3439, 3482, 3487, 3503, 3534, 3539, 3579, 3587, 3619, 3670, 3674, 3705, 3774, 3797]. Problems [288, 437, 440, 442, 445, 480, 561, 636, 641, 649, 653, 729, 754, 771, 789, 794, 817, 822, 826, 831, 870, 871, 932, 933, 969, 975, 992, 996, 1047, 1048, 1055, 1162, 1167, 1200, 1209, 1210, 1225, 1234, 1263, 1274, 1275, 1294–1296, 1299, 1362, 1389, 1424, 1444, 1445, 1567, 1573, 1577, 1641, 1760, 1775, 1835, 1917, 1920, 1923, 1954, 1978, 2014, 2036, 2045, 2108, 2129, 2176, 2207, 2220, 2238, 2266, 2316, 2350, 2458, 2483, 2570, 2640, 2660, 2672, 2679, 2690, 2698, 2708, 2721, 2732, 2921, 2945, 3001, 3007, 3134, 3196, 3274, 3304, 3383, 3429, 3568, 3622, 3676, 3687, 3710]. Problems [198, 254, 257, 270, 282, 313, 330, 411, 451, 452, 488, 566, 684, 723, 740, 762, 841, 862, 883, 899, 914, 1097, 1113, 1122, 1126, 1171, 1220, 1224, 1230, 1272, 1298, 1306, 1340, 1397, 1443, 1457, 1553, 1554, 1564, 1569, 1602, 1608, 1698, 1741, 1801, 1865, 1876, 1900, 1908, 1911, 1913, 1975, 2002, 2013, 2015, 2041, 2306, 2313, 2334, 2378, 2473, 2531, 2541, 2556, 2597, 2613, 2620, 2626, 2649, 2748, 2784, 2835, 2865, 2905, 2948, 2953, 2996, 3012, 3050, 3125, 3141, 3158, 3191, 3212, 3235, 3247, 3254, 3325, 3339, 3410, 3438, 3446, 3457, 3459, 3537, 3545, 3567, 3577, 3613, 3645, 3659, 3681, 3729, 3732, 3747, 3775]. Problems [197, 317, 344, 353, 354, 365, 455, 466, 652, 672, 688, 778, 844, 846, 948, 1096, 1188, 1195, 1221, 1243, 1308, 1422, 1513, 1538, 1585, 1701, 1735, 1749, 1781, 1804, 1831, 1856, 1888, 1898, 1904, 1905, 1915, 1938, 1942, 1982, 2034, 2107, 2119, 2134, 2168, 2197, 2204, 2225, 2259, 2282, 2330, 2345, 2356, 2402, 2418, 2422, 2446, 2447, 2454, 2471, 2534, 2557, 2567, 2631, 2634, 2639, 2670, 2713, 2717, 2775, 2792, 2814, 2826, 2883, 2898, 2950, 2952, 2988, 3037, 3063, 3070, 3089, 3092, 3101, 3133, 3139, 3149, 3290, 3297, 3353, 3385, 3427, 3436, 3465, 3472, 3497, 3505, 3556, 3612, 3634, 3654, 3690, 3693, 3739, 3771, 3784, 3787, 3798]. Problems [222, 260, 292, 456, 465, 500, 503, 543, 545, 551, 553, 584, 647, 704, 786, 827, 836, 895, 909, 924, 927, 1023, 1051, 1093, 1108, 1211, 1226, 1282, 1283, 1319, 1328, 1346, 1365, 1388, 1423, 1432, 1434, 1534, 1582, 1605, 1668, 1681, 1690, 1753, 1763, 1764, 1797, 1809, 1838, 1871, 1877, 2012, 2068, 2079, 2111, 2122, 2126, 2163, 2167, 2283, 2352, 2431, 2459, 2463, 2487, 2511, 2616, 2643, 2659, 2718, 2723, 2780, 2795, 2824, 2880, 2896, 2910, 2936, 2983, 3163, 3164, 3175, 3236, 3266, 3287, 3337, 3371, 3380, 3386, 3391, 3448, 3558, 3570, 3572, 3595, 3605, 3606, 3623, 3641, 3686, 3695, 3730, 3743, 3745, 3749, 3778, 3779, 3781, 3803]. Problems [24, 28, 29, 34, 35, 43, 59, 71, 77, 89, 102, 106, 107, 111, 124, 126, 140, 144, 155, 171, 185, 194, 828, 854, 855, 971, 1109, 1555, 1640, 1955, 3646]. Procedure [186, 1056, 1115, 1799, 1825, 2078, 2459]. Procedures [363, 2515, 2741]. Procesi [1342]. Process [1693, 2121, 2910]. processed [3077]. Processes [122, 234, 1214, 2102, 2787, 3487, 3688]. Processing [6, 92, 172, 397, 526, 559, 881, 1060, 1306, 2144, 3009, 3350]. Processors [108, 237, 2302, 2865]. Product [462, 695, 746, 1216, 2306, 2537, 3037, 3298, 3309, 3352]. Production [319, 840, 2123, 2361, 2386]. productivity [90]. Products [2035, 2080]. Professor [1916, 2231, 2488]. Profile [642, 643]. Program [1649]. Programming [1676, 2327, 2387, 2400, 2945, 3073, 3150, 3430]. Programs [210, 2118]. Projected [1188, 2424, 2896, 3286, 3453]. Projection [176, 192, 345, 841, 989, 993, 1619, 1625, 1634, 1650, 1971, 1976, 1982, 2043, 2118, 2674, 2689, 2859, 2928, 3132, 3167, 3326, 3353, 3520, 3521, 3610]. Projection-Decoupling [3132]. Projections [387, 1587, 2142, 3463]. Projective [2130]. Projector [2323]. Projectors [1028]. Prolate [918, 1213, 1754, 1822]. Prolate-Element [918]. Promotes [2853]. Proof [882, 1474]. Proofs [844, 854]. Propagating [28, 3242]. Propagation [63, 102, 161, 299, 339, 366, 486, 637, 646, 817, 823, 1268, 1281, 1321, 1333, 1447, 1770, 1813, 1861, 1965, 2000, 2193, 2230, 2242, 2254, 2535, 2628, 2848, 2933, 2938, 3213, 3289]. Propagation/Circulation [1268]. Propagations [753]. Propagator [1082]. Proper [2453, 2662, 3326, 3466, 3477]. Properties [70, 361, 502, 516, 610, 684, 1072, 1096, 1352, 1413, 1659, 2193, 2411, 2574, 2693, 2794, 2822, 2959, 3214, 3355, 3405, 3426]. Property [218, 952, 954, 1249, 1357, 2205, 2394, 3237, 3238, 3344]. ProteinRank [1199]. Provable [3700]. Proximal [1564, 1648, 1986, 1989, 2353, 2713, 2750, 3150, 3176, 3264, 3372, 3409, 3462, 3561, 3610, 3641, 3673, 3686]. Proximal-Newton [2713]. Proximity [1781, 2479, 3329]. Pseudo [340, 545, 1067, 1572, 1650, 1724, 3404, 3761]. Pseudo-PageRank [3761]. Pseudo-Spectral [340, 1067, 1572, 1724]. Pseudo-static [3404]. Pseudodifferential [2366]. Pseudomontone [2659]. Pseudospectral [50, 62, 135, 142, 245, 327, 339, 369, 399, 658, 704, 707, 784, 886, 1822, 1993, 2354, 2772, 2918, 3639]. Pseudostress [3171, 3254]. Pseudostress-Based [3254]. Pseudostress-Velocity [3171]. Ptychography [3368]. PU [1953]. Pucci [2215]. Puiseux [2565]. Pulse [381]. Pulses [859]. PUM [1964]. Pumping [1777]. Pure [561, 816, 877, 2603]. Pure-Streamfunction [816]. Pursuit [1186, 2758]. Put [3649]. Pyramidal [821].

Q4 [112]. QLP [3231]. QSC [3306]. QSVD [3791]. Quad [1918, 3007]. Quad-Curl [1918, 3007]. Quadratic [119, 860, 1114, 1649, 1740, 1974, 2405, 2616, 2857, 2996, 3091, 3342, 3538, 3733, 3775, 3780, 3798]. Quadratically [2171]. Quadratization [2438, 3124]. Quadrature [16, 41, 74, 360, 845, 857, 1101, 1112, 1166, 1292, 1525, 2267, 2277, 2774, 2922, 3168, 3378, 3442, 3518]. Quadrature-Based [3442]. Quadrature-Free [1525]. Quadratures [1223, 3251]. quadric [2071]. Quadrilateral [511, 596, 1064, 1195, 1258, 1603, 1604, 1738, 1966, 2026, 2496, 2940, 3347, 3443, 3765]. Quadrilaterals [1534, 3530]. Quadtree [1136]. Quality [962, 2552]. Quantification [887, 1530, 1650, 2319, 2764, 2912, 3788]. Quantifying [97, 3609, 3758]. Quantitative [220, 1737]. Quantized [3365]. Quantum [1368, 3758]. Quasi [798, 908, 969, 975, 980, 1162, 1169, 1267, 1436, 1514, 1581, 1611, 1728, 1915, 2085, 2206, 2366, 2678, 2680, 2804, 2810, 2930, 2941, 2948, 3024, 3149, 3173, 3197, 3299, 3307, 3358, 3461, 3462, 3691]. Quasi-A [1436]. Quasi-Biorthogonal [908]. Quasi-Compact [1169, 3307]. Quasi-conformal [1611]. Quasi-Conservative [2678, 2804, 3024]. Quasi-diagonal [2366]. Quasi-Entropy [3358]. Quasi-Equilibrium [2948]. Quasi-Geostrophic [980]. Quasi-Interpolation [1581, 3299]. Quasi-Linear [969, 975, 1162, 3149]. Quasi-Local [2680]. Quasi-Monte [3173]. Quasi-neutral [798]. Quasi-Newton [3462]. Quasi-Newtonian [1514, 1728]. Quasi-Optimal [1267]. Quasi-Optimality [1915]. Quasi-periodic [3691]. Quasi-reversibility [2810]. Quasi-Static [3197]. Quasi-Toeplitz [2930]. Quasiconformality [2202]. Quasicontinuum [712]. quasigeostrophic [180]. Quasilinear [911, 1155, 1356, 1408, 3011, 3164, 3549]. Quasimonotone [2988]. Quasineutral [1011]. Quasiperiodic [2534]. Quasipotential [3105]. Quaternion [3231, 3271, 3350, 3461, 3648]. Quaternion-based [3271]. Quintic [1785, 2028]. Quotients [3466].

Rachford [212, 2473, 3693]. Radau [369, 1051, 3463]. Radial [863, 883, 1039, 1148, 1358, 1412, 1416, 1433, 1497, 1513, 1518, 1617, 1701, 1758, 1883, 2066, 2071, 2092, 2110, 2194, 2347, 2531, 2630, 2760, 3714]. Radiance [384]. Radiating [59]. Radiation [401, 413, 423]. Radiative [459, 625, 1371, 2145, 2251, 2842, 2980, 3072, 3080, 3181, 3356, 3417, 3607]. Radii [2029]. Radiotherapy [1793]. Radon [1120]. Randles [253]. Random [84, 100, 247, 1542, 1582, 1621, 1921, 1957, 2195, 2199, 2318, 2396, 2433, 2455, 2667, 3046, 3050, 3168, 3444, 3699, 3708]. Randomized [2867, 3203, 3231, 3707]. Randomly [3152]. Range [1366, 2650, 3031, 3530]. Rank [1186, 1549, 1721, 1755, 1786, 1846, 2006, 2195, 2333, 2365, 2480, 2776, 2854, 2863, 2945, 3015, 3035, 3093, 3099, 3150, 3231, 3263, 3323, 3355, 3475, 3554, 3656, 3788]. Rank-1 [3099]. Rank-Adaptive [2863]. Rank-One [2333, 3263]. Rannacher [2768]. RANS [232, 2294, 2670]. Rao [3689]. Rapid [3747]. Rate [264, 1267, 1661, 1989, 2444, 2915, 2984, 3478]. Rates [146, 2754, 3107, 3673]. Ratio [2797]. Rational [173, 253, 265, 506, 554, 835, 1302, 1580, 1594, 2101, 2128, 2277, 2834, 2885, 2925, 3111, 3315]. Ratios [2035, 3384]. Raviart [1357, 1399, 1437, 1441, 1713, 2015, 3082, 3141, 3310]. Ray [1120, 2284]. Rayleigh [58, 403, 2240, 2271, 2459]. Rayleigh-Marangoni [403]. RBF [1412, 1831, 1953, 1964, 2173, 2222, 2340, 2553, 2870, 3126, 3322, 3363, 3371, 3431, 3543]. RBF-Based [2340]. RC [2624]. Reachable [387]. Reacting [39, 274, 489, 964, 3512, 3769]. Reaction [32, 230, 243, 781, 801, 836, 855, 893, 924, 946, 979, 1043, 1191, 1210, 1412, 1536, 1585, 1602, 1613, 1757, 1777, 1998, 2108, 2135, 2162, 2217, 2282, 2298, 2326, 2518, 2645, 2675, 2683, 2759, 2824, 2828, 2927, 2953, 3069, 3221, 3385, 3536, 3723, 3751, 3773]. Reaction-Diffusion [1536, 3069]. reactions [109]. Reactive [203, 250, 456, 3467]. Reactor [223]. Real [78, 1414, 2941, 3099, 3445, 3810]. Realistic [1965, 3351]. realizability [157]. Realization [323, 3692]. Really [1941]. Rearrangement [1126, 2729, 2748]. Rebuilding [912]. Rebuilding-Divergency [912]. Receding [2293]. Recipes [910]. Recognition [3461]. Recombination [1446]. Reconstructing [1990, 2762, 3660]. Reconstruction [26, 162, 187, 271, 310, 415, 494, 521, 567, 587, 598, 606, 828, 879, 889, 917, 997, 1002, 1018, 1021, 1115, 1120, 1130, 1131, 1179, 1208, 1282, 1283, 1334, 1382, 1483, 1543, 1586, 1589, 1603, 1604, 1624, 1627, 1674, 1675, 1717, 1721, 1737, 1883, 1941, 1942, 2026, 2028, 2040, 2146, 2200, 2315, 2417, 2536, 2552, 2554, 2588, 2680, 2704, 2779, 2795, 2901, 3019, 3187, 3238, 3262, 3297, 3310, 3343, 3441, 3530, 3531, 3638, 3644, 3656, 3699, 3712, 3733, 3741, 3755, 3785]. Reconstructions [944, 1205, 1622, 2586, 2663, 2720, 2873, 3126, 3615]. Recovering [565, 1015, 1509]. Recovery [295, 635, 652, 814, 864, 883, 1153, 1207, 1395, 1410, 1437, 1533, 1641, 1700, 1813, 1976, 2260, 2303, 2380, 2768, 2769, 2883, 3023, 3154, 3258, 3311, 3331, 3456, 3516, 3522, 3719, 3786]. Recovery-Based [1533, 2883, 3023]. Rectangular [81, 124, 598, 700, 702, 919, 983, 1177, 1238, 1240, 1474, 1511, 1844, 2051, 2630, 3484]. Rectified [3342]. Recurrence [1181, 2907]. Recursive [1041, 2035]. Red [3280]. Redistancing [1394]. Redistribution [659, 1151, 3386]. Reduced [277, 333–335, 843, 988, 992, 1010, 1164, 1223, 1263, 1442, 1473, 1518, 1591, 1626, 1694, 1904, 1933, 1961, 1962, 2041, 2213, 2454, 2493, 2783, 2792, 2869, 2931, 2939, 2967, 3080, 3116, 3166, 3316, 3326, 3353, 3360, 3387, 3413, 3423, 3446, 3466, 3492, 3541, 3574, 3773, 3776]. Reduced-Basis [333, 334, 2452]. Reduced-Order [277, 843, 1591, 1962, 2792, 3326, 3423, 3492]. Reducing [310, 2260].

Reduction [20, 415, 1308, 1366, 1493, 1580, 2060, 2080, 2095, 2108, 2448, 2449, 2616, 2881, 2971, 3010, 3104, 3216, 3218, 3296, 3319, 3333, 3372, 3385, 3425, 3446, 3519, 3561]. Reduction-in-Time [3519]. Reference [1072, 2747]. Reference-to-Physical [1072]. Refined [2138]. Refinement [371, 693, 792, 1213, 1384, 1543, 1895, 2296, 2441, 2745, 2764, 3397, 3403]. Refinements [1755]. Reflectance [806, 3720]. Reflected [3286, 3422]. Reflecting [757, 1379, 1686, 2483]. Reflection [1450, 2307]. Reflexive [3537]. Reformulated [3586]. Reformulation [3724]. Reformulations [2796]. Regime [852, 1242, 1833, 1869, 2096, 2097, 2099, 2573, 2610, 2715, 3071, 3207, 3731]. Regime-Switching [2610, 3731]. Region [123, 1960, 2496, 2904, 3263, 3455, 3580]. Regions [133, 1148, 1671, 1672, 2217, 2248]. Registration [1005, 2607, 2758, 2881]. Registration-Based [2881]. Registrations [1004, 1020]. Regression [1250, 2825, 3176, 3546]. Regular [2100, 2524, 2551]. Regularisation [1998]. Regularity [290, 897, 1797, 2168, 2351, 2643, 3084, 3089, 3205, 3469, 3480, 3496, 3503, 3697, 3764]. Regularity-Aware [2351]. Regularization [396, 727, 901, 960, 1120, 1188, 1212, 1215, 1389, 1712, 1773, 1786, 1886, 1990, 2006, 2039, 2050, 2093, 2201, 2280, 2316, 2360, 2590, 2769, 2857, 2963, 3047, 3093, 3187, 3202, 3252, 3271, 3323, 3523, 3771, 3792]. Regularized [390, 1318, 1583, 1595, 1837, 1910, 1958, 2017, 2118, 2317, 2713, 2755, 3134, 3135, 3162, 3400, 3546, 3633]. Regularizers [1660]. regulating [39]. Regulation [2885]. Reinforcement [3190, 3568]. Reinitialization [681, 1800, 2601]. Reissner [438, 967, 1435, 2062]. Reject [115]. rejection [115]. Related [435, 716, 826, 1679, 3564]. Relation [176, 1365, 1421]. Relations [1651, 3226]. Relationship [1564]. Relationships [1256]. Relative [806, 1817, 2080, 3617]. Relatively [2551]. Relativistic [1947, 2751, 2937, 3679, 3727]. Relaxation [274, 399, 484, 804, 1172, 1361, 1415, 1629, 1696, 1877, 2130, 2181, 2568, 2624, 2631, 3285, 3408, 3581, 3598, 3611, 3653]. Relaxations [137, 1185]. Relaxed [267, 347, 427, 1797, 2875, 2945, 2988, 3332]. Relaxing [267, 1891]. Release [236]. Reliability [2371]. ReLU [3053, 3464]. Remarks [956, 983]. Removal [61, 1676, 1985, 2050, 2091, 2577, 3004, 3589]. Removing [1575]. Renormalization [2, 37, 55, 64, 104, 231, 236, 247, 562]. Representation [96, 1345, 1426, 2450, 3500]. Representations [865, 2551]. Represented [1178]. Reproducibility [3499, 3508]. Reproducing [3556]. Reproduction [2701]. Reprojection [1015]. Requirements [57, 1614]. Requiring [845]. Research [3499, 3508]. Reseeding [1478]. Reservoir [627]. Residual [97, 295, 596, 867, 954, 1216, 1377, 1403, 1466, 1893, 2163, 2356, 2358, 2865, 3005, 3008, 3228, 3341, 3363, 3473, 3600, 3728, 3779, 3789]. Residual-Based [3228, 3341]. Residual-Type [1466]. Residuals [1558]. Resilience [2278]. Resistance [1402]. Resistive [470]. Resolution [57, 94, 154, 176, 188, 347, 378, 427, 512, 727, 901, 995, 1083, 1122, 1289, 1458, 1669, 1928, 1946, 2028, 2548, 2593, 2680, 2719, 3305, 3364, 3393, 3445, 3635, 3794].

Resolvent [2579]. Resolving [275]. Resonance [745, 1721]. Resonances [2156, 3594]. Resonant [779]. Resource [1679]. Respect [2187, 2995]. Response [498, 2003]. Restart [2972, 3263]. Restoration [376, 650, 668, 986, 1121, 1129, 1212, 1290, 1360, 1523, 1628, 2006, 2297, 2309, 2578, 2817, 2853, 3233, 3257, 3422, 3808, 3811]. restoring [146]. Restricted [586]. Restrictions [991, 2962]. Resulting [1829]. Results [30, 38, 110, 980, 1081, 1309, 1450, 2215]. Retracted [3218]. Retraction [3333]. Retrieval [2333]. reversibility [2810]. Reversible [253]. Reverting [2077]. Review [479, 963, 1081]. Revisit [1480]. Revisited [2498, 3332]. Revisiting [872]. Reynolds [151, 157, 240, 537, 1236, 2064, 2160]. Rheological [383]. Riccati [2633, 3563]. Richards [721, 2576]. Richardson [133, 1029, 1196, 1292, 1987]. Rician [2309, 3004]. Riemann [804, 884, 940, 943, 1106, 1147, 1302, 1713, 3184, 3345, 3544, 3717]. Riemannian [2562, 3263, 3580]. Riesz [1820, 2391, 2757, 2949]. Right [2285, 2286]. Right-Hand [2286]. Rigid [355]. Rigorous [1257, 1854, 2598]. Ring [1610]. Ritz [2459, 3732]. River [2496]. RK [2500, 3300]. RK-SAV [3300]. RK-SAV/DG [3300]. RK/Implicit [2500]. RKDG [444, 1160, 1670, 1733, 1951]. RKDG2 [985]. RNG [240]. Roadmap [1804]. Roaming [2918]. Robin [1686, 2748, 2935, 2994, 3573]. Robinson [2070]. Robots [2991]. Robust [91, 184, 312, 337, 736, 909, 1130, 1131, 1311, 1317, 1434, 1441, 1466, 1642, 1706, 1737, 2195, 2228, 2229, 2232, 2421, 2461, 2477, 2520, 2551, 2570, 2643, 2695, 2740, 2749, 2776, 2839, 2914, 2963, 3022, 3110, 3118, 3205, 3211, 3246, 3261, 3339, 3361, 3401, 3448, 3515, 3536, 3567, 3604, 3650, 3712, 3743, 3749, 3793]. Robustly [3072]. Robustness [709, 1928, 3206, 3672]. ROCK [3669]. Rods [1720]. Roe [748, 2229]. Role [1469, 1930, 2219, 2794, 3251]. ROM [2969]. Root [184, 3558]. Roots [276, 2877]. Rosenau [2092]. Rosenbrock [1994]. Rossby [660]. Rotated [1966, 2873, 3765]. Rotated- [3765]. Rotating [605, 734, 903, 2838, 3142]. Rotating-Grid [2838]. Rotational [663]. Rough [1319, 1476, 2795, 3151, 3179]. Round [415]. ROW [2588]. RS [152]. RS/6000 [152]. RT [795, 3347]. RTN [3347]. Rubenchik [1271, 3381]. Rule [83, 360, 1545, 2001, 2422, 3378]. Rules [1101, 3130, 3168, 3644]. Runge [222, 287, 288, 361, 462, 483– 485, 648, 696, 749, 758, 789, 860, 901, 904, 1181, 1220, 1300, 1377, 1392, 1403, 1407, 1522, 1614, 1615, 1778, 1828, 1925, 1931, 1932, 2011, 2072, 2123, 2127, 2139, 2155, 2220, 2386, 2411, 2502, 2608, 2612, 2625, 2631, 2826, 2929, 2962, 3026, 3074, 3086, 3096, 3160, 3248, 3253, 3284, 3408, 3527, 3535, 3611, 3754, 3760, 3782]. Runs [752]. Ruptures [1140].

S [1710, 3669]. S-FDTD [1710]. S-ROCK [3669]. Saddle [1200, 2473, 2616, 2928, 3135, 3137, 3340]. Saddles [2264, 2265]. Saha [165]. Saint [1409]. Saint-Venant [1409]. Sample [3689]. Sampled [3154]. Sampling [3451, 3454, 3568, 3571, 3748]. SAT [2698, 2961]. satisfaction [81]. Satisfying [617, 632, 804, 981, 1378, 1464, 1677, 2368, 2799, 3304, 3465].

Saturable [2668]. Saturation [2591]. Saul [2488, 2489]. SAV [2314, 2548, 2832, 2866, 2891, 2979, 3003, 3220, 3382, 3553, 3671, 3678]. SAV/DG [3300]. Saving [3106]. Saxton [3768]. SBFEM [1061]. SBP [2057, 2498, 2698, 2961]. SBP-SAT [2961]. SCAD [3004]. Scalability [169, 906, 1745]. Scalable [475, 1327, 2094, 2705, 2982, 3134]. Scalar [247, 417, 566, 875, 922, 939, 1228, 1464, 1533, 1621, 1787, 2163, 2252, 2572, 2831, 2866, 3039, 3085, 3253, 3279, 3301, 3346, 3542, 3664]. Scalars [166, 393]. Scale [46, 49, 101, 129, 166, 176, 353, 674, 710, 847, 960, 1188, 1319, 1333, 1395, 1563, 1580, 1654, 1746, 1805, 2095, 2956, 3031, 3109, 3134, 3175, 3364, 3425, 3453, 3558, 3700, 3745]. Scale-dependent [49]. Scale-Invariant [3364]. Scaled [665, 1499, 2340, 2836, 3686]. Scales [965, 1168]. Scalings [1921]. Scanning [2441]. Scans [1008]. Scattered [14, 3065]. Scattering [244, 246, 260, 888, 1092, 1359, 1656, 1744, 1784, 1881, 1987, 2319, 2792, 2939, 3012, 3095, 3579, 3594, 3676, 3726, 3749]. Schatten [2520]. Schatten-p [2520]. Scheme [212, 213, 345, 417, 497, 552, 613, 625, 658, 662, 671, 816, 817, 867, 868, 899, 900, 922, 938, 941, 945, 972, 1014, 1030, 1065, 1076, 1099, 1138, 1150, 1156, 1159, 1170, 1277, 1337, 1421, 1425, 1597, 1674, 1675, 1695, 1700, 1724, 1823, 1843, 1858, 1869, 1882, 1889, 1893, 1928, 2014, 2044, 2088, 2089, 2125, 2133, 2175, 2178, 2185, 2186, 2212, 2213, 2229, 2294, 2337, 2408, 2417, 2442, 2456, 2499, 2568, 2576, 2593, 2652, 2732, 2770, 2839, 2943, 2978, 3030, 3033, 3169, 3185, 3191, 3219, 3225, 3237, 3322, 3373, 3397, 3440, 3521, 3545, 3566, 3588, 3625, 3649, 3660, 3680, 3718, 3724, 3780]. Scheme [239, 281, 299, 368, 389, 430, 566, 576, 599, 605, 610, 634, 656, 670, 705, 707, 728, 742, 846, 903, 950, 1044, 1312, 1448, 1457, 1458, 1460, 1464, 1474, 1497, 1501, 1508, 1560, 1647, 1710, 1726, 1780, 1794, 1796, 1827, 1834, 1892, 1942, 1973, 1981, 1994, 2028, 2124, 2148, 2150, 2177, 2221, 2227, 2299, 2361, 2399, 2413, 2438, 2466, 2516, 2550, 2566, 2572, 2594, 2600, 2623, 2694, 2703, 2775, 2819, 2838, 2843, 2852, 2876, 2884, 2886, 2899, 2958, 2959, 2965, 2968, 2970, 3003, 3025, 3029, 3068, 3112, 3118, 3183, 3206, 3284, 3290, 3297, 3307, 3364, 3388, 3399, 3424, 3439, 3460, 3495, 3602, 3635, 3638, 3671, 3751, 3755, 3798, 3800]. Scheme [71, 113, 191, 752, 795, 798, 890, 893, 1083, 1133, 1174, 1208, 1252, 1260, 1262, 1429, 1442, 1544, 1593, 1616, 1655, 1740, 1767, 1863, 2103, 2109, 2151, 2162, 2469, 2497, 2558, 2604, 2614, 2622, 2738, 2774, 2785, 2800, 2977, 2979, 3002, 3048, 3308, 3334, 3362, 3367, 3428, 3441, 3448, 3552, 3696, 3752, 3783, 3809]. Schemes [206, 215, 274, 286, 296, 318, 356, 362, 373, 378, 405, 433, 482, 485, 504, 525, 529, 532, 554, 581, 591, 596, 659, 660, 697, 721, 748, 762, 782, 923, 936, 939, 1018, 1077, 1107, 1192, 1205, 1311, 1314, 1320, 1350, 1361, 1382, 1411, 1525, 1566, 1568, 1612, 1613, 1645, 1662, 1665, 1687, 1693, 1702, 1708, 1719, 1750, 1776, 1802, 1890, 1902, 1917, 1921, 1926, 2032, 2048, 2127, 2143, 2272, 2275, 2314, 2394, 2407, 2543, 2656, 2696, 2721, 2728, 2730, 2751, 2796, 2805, 2823, 2891, 2902, 3000, 3018, 3057, 3075, 3107, 3223, 3226, 3402, 3517, 3535, 3679, 3682, 3759, 3790]. Schemes [275, 347, 357, 360, 427, 484, 501, 534, 536, 542, 609, 620, 628, 666, 667, 911, 928, 946, 952, 997, 1033, 1059, 1080, 1131, 1177, 1198, 1220, 1228, 1344, 1355, 1370, 1385, 1404, 1407, 1438, 1465, 1495, 1539, 1558, 1578, 1584, 1624, 1715, 1718, 1759, 1772, 1852, 1855, 1913, 1931, 1932, 1974, 1979, 2020, 2070, 2072, 2075, 2076, 2098, 2123, 2153, 2293, 2295, 2386, 2406, 2411, 2444, 2474, 2490, 2542, 2560, 2567, 2599, 2606, 2673, 2739, 2743, 2753, 2767, 2802, 2808, 2831, 2848, 2862, 2885, 2888, 2900, 2942, 3101, 3147, 3215, 3220, 3262, 3313, 3357, 3393, 3442, 3498, 3527, 3529, 3569, 3591, 3620, 3650, 3688, 3806]. Schemes [16, 44, 69, 70, 118, 129, 139, 261, 267, 269, 429, 431, 432, 471, 490, 508, 548, 580, 582, 607, 623, 681, 758, 764, 810, 907, 917, 925, 956, 957, 981, 1110, 1160, 1169, 1181, 1229, 1238, 1244, 1282, 1283, 1331, 1334, 1365, 1376, 1377, 1409, 1477, 1494, 1561, 1603, 1604, 1663, 1678, 1688, 1717, 1722, 1754, 1788, 1799, 1809, 1863, 1884, 1952, 1956, 1993, 2068, 2226, 2274, 2439, 2478, 2492, 2508, 2518, 2546, 2585, 2601, 2650, 2766, 2790, 2814, 2944, 2966, 2986, 3041, 3042, 3054, 3088, 3132, 3172, 3174, 3224, 3300, 3381, 3426, 3434, 3510, 3526, 3530, 3548, 3586, 3590, 3607, 3684, 3727, 3734, 3781, 3794]. Schemes [1095]. Schemings [2231]. Schmit [3266]. Schnirelman [1516]. Scholes [759, 1063, 2545]. Schrödinger [683, 779, 820, 966, 1082, 1098, 1143, 1303, 1446, 1487, 1515, 1528, 1844, 1845, 1873, 1887, 1950, 2099, 2203, 2237, 2287, 2481, 2504, 2509, 2668, 2725, 2811, 2831, 2851, 2891, 2906, 2911, 2944, 2951, 3002, 3044, 3071, 3084, 3335, 3348, 3419, 3447, 3483, 3678, 3682, 3683, 3709]. Schrödinger-Type [1845]. Schur [1485, 2459, 3617]. Schwarz [196, 448, 460, 518, 699, 916, 927, 1526, 1760, 1847, 1877, 1977, 2271, 2367, 2634, 2735, 2935, 2972, 3144, 3432, 3573]. Schwarzschild [2937]. Science [710]. Scientific [131, 2489, 3239]. SDEs [3105]. SDFEM [996, 1681, 3051]. SDG [1766]. Sea [1777]. Seafloor [1946]. Seamless [2983]. Search [807, 2871, 2910, 3372, 3433, 3561]. Second [318, 339, 409, 516, 624, 662, 664, 728, 731, 778, 782, 799, 815, 939, 950, 1038, 1042, 1044, 1155, 1195, 1219, 1261, 1274, 1277, 1299, 1380, 1424, 1476, 1503, 1566, 1649, 1671, 1672, 1724, 1748, 1762, 1765, 1772, 1796, 1840, 1858, 1882, 1899, 1902, 1945, 2012, 2015, 2022, 2045, 2063, 2067, 2069, 2151, 2154, 2177, 2212, 2226, 2227, 2241, 2274, 2377, 2390, 2406, 2434, 2442, 2458, 2465, 2471, 2484, 2501, 2540, 2565, 2626, 2648, 2677, 2722, 2828, 2866, 2886, 2890, 2891, 2900, 2911, 2942, 2961, 3003, 3127, 3279, 3320, 3388, 3390, 3430, 3485, 3486, 3514, 3533, 3545, 3559, 3614, 3629, 3630, 3698, 3704, 3780, 3806]. Second [88, 681, 732, 897, 971, 999, 1477, 1615, 1871, 2005, 2086, 2370, 2384, 2718, 2738, 3020, 3034, 3187, 3276, 3428, 3562, 3607, 3745]. Second-Derivative [3020, 3559]. Second-Derivatives [1038, 1042]. Second-Order [88, 516, 662, 664, 681, 799, 815, 897, 999, 1155, 1277, 1299, 1380, 1424, 1477, 1615, 1649, 1724, 1762, 1796, 1840, 1871, 1882, 2005, 2086, 2151, 2370, 2377, 2458, 2471, 2722, 2890, 2891, 2900, 2911, 2942, 3276, 3320, 3388, 3428, 3430, 3485, 3514, 3533, 3562, 3607, 3614, 3745, 3780, 3806]. Sediment [382, 951, 2692]. Sedimentation [355]. Segel [773, 1105, 1936, 2311, 2445, 2686, 3625]. Segmentation [682, 879, 986, 1130, 1178, 1215, 1266, 1360, 1394, 1685, 1703, 1782, 1897, 2309, 2362, 3271]. Segmenting [1896]. Segregated [2746]. Seidel [1185, 3808]. Seidel-Based [3808]. Seismic [1946, 2917, 3142, 3202, 3716]. Seismo [1965]. Sekerka [3747]. Selecting [452]. Selection [297, 1022, 1065, 1277, 1499, 1570, 1953, 2177, 2442, 2456, 2512, 2958, 2968, 3008, 3122, 3259, 3284, 3498]. Selective [433, 2144]. Self [39, 75, 2206, 2780, 2921, 3416, 3783]. Self-adaptive [2780]. Self-Attention [3416]. Self-contact [3783]. self-focusing [75]. Self-regulating [39]. Self-similar [2206]. selfadjoint [1108, 2385]. SEM [2938]. Semi [279, 325, 346, 392, 491, 492, 525, 605, 690, 743, 754, 761, 789, 798, 804, 971, 1159, 1218, 1244, 1339, 1397, 1519, 1614, 1649, 1665, 1667, 1678, 1748, 1844, 1852, 1919, 1963, 1995, 2006, 2021, 2042, 2059, 2118, 2211, 2212, 2218, 2226, 2327, 2364, 2379, 2423, 2439, 2518, 2568, 2573, 2580, 2600, 2675, 2679, 2694, 2766, 2795, 2828, 2880, 2957, 2980, 3018, 3075, 3114, 3132, 3190, 3273, 3353, 3373, 3396, 3412, 3496, 3504, 3529, 3614, 3639, 3718]. Semi- [3132]. Semi-Analytical [1397, 2580]. Semi-classical [2573]. Semi-Compact [605]. Semi-definite [2327]. Semi-Discrete [525, 804, 1244, 1678, 2957, 3018, 3496]. Semi-discretizations [1963]. Semi-discretized [1844]. Semi-Implicit [279, 325, 392, 491, 690, 1614, 1665, 1667, 1748, 1995, 2042, 2211, 2212, 2218, 2226, 2423, 2439, 2518, 2568, 2600, 2675, 2694, 2766, 2980, 3353, 3373, 3529, 3718]. Semi-infinite [1649]. Semi-Lagrangian [346, 492, 743, 761, 789, 798, 971, 1159, 1339, 1519, 1852, 1919, 2021, 2364, 2679, 2795, 2828, 3075, 3273]. Semi-linear [754, 2880, 3639]. Semi-Markov [3190]. Semi-Separable [1218]. Semi-Smooth [2059, 2118, 3114]. Semi-stationary [3396]. Semi-Supervised [2006, 3412]. Semicirculant [259]. Semiclassical [683, 820, 2099, 3335]. Semiconductor [2403, 2684]. Semidefinite [2722, 2871, 2945, 3150, 3433]. Semidiscrete [290, 2740, 2846]. Semilinear [700, 978, 1064, 1363, 1372, 1516, 1615, 1994, 2383, 2388, 2775, 3296, 3491, 3570]. Semismooth [927, 2655, 3233, 3558, 3610]. SENO [3377]. sense [87, 99]. Sensing [1318, 3412]. Sensitivities [2580]. Sensitivity [28, 1626, 2371, 2392, 2497, 2852, 3394]. Sensor [2469, 3060]. Separable [1218, 1564, 1781, 3087]. separated [181]. Separation [36, 849, 1041, 1323, 2169]. Separation-bubble [36]. sequences [84, 100]. Sequential [149, 1968, 3154, 3476, 3522, 3593]. Series [173, 238, 256, 506, 2544, 2850, 3289, 3628, 3704]. Serre [1324]. Set [374, 377, 380, 384, 392, 394, 589, 615, 618, 679, 681–684, 686, 688–690, 724, 763, 902, 1123, 1125, 1345, 1379, 1478, 1520, 1557, 1732, 1761, 1800, 1933, 2030, 2147, 2262, 2517, 2601, 2713, 2724, 2799, 2967, 2973, 3078, 3268, 3420, 3610, 3657]. Set-Moment [1125]. Sets [100, 387, 512, 828, 995, 1606, 3361, 3438]. Settings [1531]. Setups [2307]. Sevcik [253]. Seven [1712]. Seven-Equation [1712]. Seventh [2970]. Seventh-Order [2970]. Several [858]. Severely [3371]. Shah [1660]. Shallow [439, 446, 447, 527, 552, 701, 777, 849, 928, 936–938, 943, 947, 950, 954, 955, 957, 985, 1033, 1156, 1179, 1197, 1384, 1409, 1593, 1629, 1670, 1823, 1830, 1929, 2407, 2605, 2609, 2692, 2710, 2737, 2812, 2823, 2849, 2876, 2903, 2987, 3030, 3184, 3223, 3237, 3260, 3262, 3318, 3471, 3601, 3684, 3729, 3746, 3755]. Shallow-Water [439, 937, 1384, 1670, 2849, 3030, 3729]. Sham [1151, 2346]. Shannon [2060]. Shannon-Wavelet [2060]. Shape [119, 384, 450, 682, 898, 1126, 1309, 1737, 1758, 1930, 2053, 2195, 2304, 2319, 2864, 3469]. shape-preserving [119]. Shaped [2319]. Shapes [377, 1416]. Sharp [188, 676, 1122, 1349, 1396, 1461, 2011, 3118, 3219, 3637, 3658]. Sharp-Interface [1349, 3637]. Sharpness [3365]. Sharpness-Aware [3365]. Shaw [851, 1560].

Shear [2982, 3684]. Sheet [329, 1379, 3528]. Shell [1930]. Shift [683, 3648]. Shifted [365, 566, 2285, 2286, 2633, 2922, 3158]. Shifts [3504]. Shishkin [1681, 2167, 2707, 3266]. Shock [142, 188, 206, 388, 410, 445, 667, 764, 1783, 2201, 2229, 2338, 2475, 2571, 2721, 3079, 3717]. Shock-Vortex [388]. Shocks [319, 390, 2789]. Shooting [3645]. Short [1081]. Shortley [1980, 2055]. Shrinking [3560]. Shu [386, 1916]. SIAC [887, 1046, 1079, 1254, 1453, 1581, 2409, 2475, 2491]. Sided [559, 735, 1453, 2349, 2993]. Sides [2285, 2286]. Sieving [3150, 3778]. Sign [1627, 2551]. Signal [2144, 2779, 3516, 3786]. Signals [706, 1447, 3578]. Signorini [1055, 3117]. similar [2206]. Simons [3763]. Simple [178, 389, 785, 901, 940, 1240, 1247, 1389, 1511, 1939, 2741]. Simpler [2726]. Simplex [2089, 2765, 3176, 3463, 3468]. Simplices [3204, 3748]. Simplicial [998, 2884, 3609]. Simplified [976, 1586, 1713, 2079, 2619, 3615]. Simply [1506]. Simpson [83]. Simulated [131]. Simulating [46, 130, 575, 823, 1326, 1431, 2000, 2011, 2535, 2596, 2669, 3243, 3671]. Simulation [3, 7, 8, 25, 32, 40, 58, 90, 92, 108, 109, 123, 163, 170, 179, 186, 203, 232, 235, 252, 298, 311, 320, 328, 337, 355, 380, 381, 401, 413, 423, 431, 459, 465, 469, 524, 537, 539, 583, 600, 646, 667, 670, 694, 710, 714, 733, 779, 818, 837, 847, 851, 878, 920, 942, 953, 959, 976, 1140, 1198, 1253, 1284, 1310, 1320, 1396, 1415, 1430, 1446, 1467, 1497, 1511, 1578, 1777, 1791, 1793, 1816, 1825, 1903, 1967, 2042, 2102, 2105, 2152, 2156, 2211, 2262, 2347, 2373, 2441, 2684, 2693, 2742, 2981, 3213, 3467, 3638]. Simulations [55, 57, 169, 250, 285, 286, 289, 293, 312, 329, 346, 352, 414, 424, 576, 610, 618, 734, 751, 928, 963, 1077, 1119, 1123, 1150, 1170, 1183, 1222, 1271, 1557, 1845, 1919, 2021, 2121, 2127, 2342, 2481, 2591, 2800, 2969, 2975, 3188, 3582]. Simultaneous [831, 2033, 2817]. Simultaneous-Approximation-Term [831]. Sinc [2818]. Sine [1100, 2438, 3472, 3624]. Sine-Cosine [1100]. Single [260, 333, 728, 755, 1227, 1379, 1448, 1632, 2134, 2152, 2686, 2704, 2774, 2917, 3108]. Single-Domain [1448]. Single-Parameter [333]. Single-Pass [3108]. Single-Stage [1632]. Single-Step [728, 1632, 2774]. Singly [2825, 3535]. Singular [41, 88, 107, 194, 297, 390, 586, 703, 747, 791, 794, 1292, 1572, 1697, 1709, 1812, 1815, 1817, 1870, 2069, 2204, 2238, 2434, 2565, 2636, 2816, 2893, 3031, 3143, 3224, 3303, 3352, 3384, 3486, 3572, 3628, 3707, 3734, 3812]. Singularities [50, 270, 358, 396, 1509, 1886, 2372, 2414, 3506, 3584]. Singularity [1383, 1449, 1515, 2901, 3213, 3622]. Singularly [996, 1317, 1347, 1401, 1858, 1875, 1908, 2167, 2526, 2953, 3051, 3221, 3235, 3266, 3536]. Singularly-Perturbed [2526]. Sinkhorn [3685]. SISC [3499, 3508]. Sivashinsky [101, 970]. Six [1457]. Sixth [801, 1110, 1457, 2098, 2601, 3307]. Sixth-Order [801, 1110, 2098, 2601, 3307]. Size [177, 656, 665, 911, 931, 1845, 3334]. Size-Structured [656, 911]. Sketch [3245]. Sketch-Based [3245]. Sketching [3475]. Skew [1207, 1341, 1495, 2467, 2473]. Skew-Hamiltonian [1341]. Skew-Symmetric [1495, 2467, 2473]. Skew-Symmetry [1207]. Skewed [3444]. SLaT [1896]. Slightly [1901]. Slip [1235, 1559, 1619, 1640, 1909, 2137, 2462, 2714, 2950].

Slipping [2213]. Slit [1148, 1671, 1672, 1744, 2279]. Slope [1065, 1277, 2177, 2442, 2456, 2958, 2968, 3122, 3284]. Slot [2342]. Slow [752]. Slowness [395]. Slyozov [402]. Small [129, 166, 235, 237, 390, 521, 586, 2325, 2641, 2855, 3061, 3169, 3225, 3647, 3661]. small-scale [129, 166]. Smith [856]. Smooth [123, 209, 271, 319, 384, 506, 564, 606, 654, 716, 732, 870, 1086, 1266, 1345, 1463, 1509, 1782, 1934, 1957, 2059, 2118, 2338, 2402, 2408, 2518, 2642, 2923, 3019, 3114, 3235, 3264, 3277, 3549, 3641, 3811]. Smoothed [578, 1028, 3016]. Smoothed- [578]. Smoother [1005, 2500]. Smoothers [1847, 2294]. Smoothing [472, 1463, 1896, 2194, 2301, 2629, 2708, 2779, 3009, 3314, 3546, 3556]. Smoothness [623, 735, 887, 1046, 1079, 1254, 1453, 1581, 1839, 1940, 2201, 2491, 2566, 3615]. Smoothness-Increasing [735, 887, 1046, 1079, 1254, 1453, 1581, 1940, 2491]. Snapshot [3216]. Snapshot-Based [3216]. Sobolev [802, 1013, 2267, 2322, 3294]. Soil [1061]. Soil-Structure [1061]. Sojourn [3190]. Sojourn-Based [3190]. Sokolov [2833]. Solar [806, 2127]. Solely [3019]. Solid [823, 1337, 1431, 2131]. Solidification [34, 1396]. Solids [715, 1281]. Solitary [393, 973, 1253]. Soliton [3483]. Solitons [878, 1374]. Solute [1600]. Solution [9, 36, 52, 95, 117, 121, 127, 133, 135, 136, 140, 149, 156, 164, 165, 167, 180, 183, 205, 218, 222, 228, 249, 260, 430, 496, 505, 544, 571, 586, 592, 629, 642, 643, 712, 832, 848, 1151, 1304, 1309, 1328, 1334, 1398, 1457, 1538, 1686, 1705, 1717, 1723– 1725, 1831, 1886, 1913, 1958, 1998, 2089, 2112, 2181, 2216, 2328, 2329, 2443, 2483, 2485, 2534, 2561, 2587, 2633, 2733, 2818, 2829, 2923, 2947, 2996, 3083, 3090, 3137, 3277, 3304, 3446, 3479, 3487, 3535, 3737, 3742, 3802]. Solution-Domain-Decomposition [205]. Solutions [23, 53, 222, 237, 270, 292, 324, 362, 395, 399, 437, 589, 614, 617, 623, 685, 731, 783, 844, 854, 870, 897, 925, 961, 985, 1046, 1079, 1193, 1254, 1342, 1352, 1372, 1427, 1452, 1516, 1531, 1572, 1619, 1719, 1745, 1805, 1815, 2168, 2206, 2238, 2402, 2412, 2491, 2503, 2512, 2518, 2856, 2919, 2941, 2995, 3019, 3094, 3143, 3155, 3345, 3376, 3449, 3455, 3549, 3643, 3768, 3775]. Solvability [1416]. Solvable [2178, 2696, 3588]. Solve [625, 817, 1085, 1094, 1279, 1306, 1326, 1646, 1929, 2810, 2844, 3149, 3781]. Solvent [801, 1600]. Solver [217, 251, 257, 263, 336, 421, 583, 741, 790, 940, 943, 972, 1012, 1027, 1053, 1218, 1484, 1629, 1713, 1770, 1829, 2042, 2228, 2382, 2564, 2920, 3544, 3613, 3692, 3717, 3743, 3774, 3788]. Solvers [108, 182, 502, 523, 744, 746, 796, 804, 865, 869, 1147, 1239, 1241, 1302, 1369, 1507, 1690, 1775, 1943, 2039, 2278, 2460, 2874, 3032, 3100, 3178, 3184, 3618, 3632]. Solving [198, 202, 212, 213, 276, 340, 365, 383, 426, 490, 554, 588, 664, 707, 722, 723, 911, 915, 1025, 1051, 1198, 1228, 1269, 1275, 1344, 1423, 1494, 1601, 1653, 1693, 1702, 1736, 1795, 1821, 1824, 1902, 1908, 2002, 2004, 2083, 2100, 2122, 2166, 2171, 2173, 2252, 2272, 2285, 2286, 2295, 2334, 2352, 2384, 2428, 2434, 2458, 2464, 2472, 2490, 2513, 2562, 2597, 2672, 2678, 2690, 2723, 2780, 2819, 2862, 2875, 2887, 2924, 3008, 3050, 3052, 3108, 3133, 3134, 3138, 3183, 3203, 3221, 3273, 3280, 3282, 3292, 3325, 3342, 3357, 3363, 3368, 3395, 3411, 3422, 3436, 3454, 3465, 3505, 3537, 3583, 3590, 3605, 3623, 3627, 3703, 3739, 3789, 3810]. Solving [35, 106, 128, 398, 944, 1319, 1387, 1432, 1616, 1697, 1970, 2659, 2910, 3178, 3354]. Some [139, 159, 182, 222, 292, 295, 627, 747, 840, 1205, 1292, 1498, 1561, 1858, 1907, 1960, 2048, 2113, 2204, 2235, 2266, 2551, 3483, 3497, 3655]. Sonic [3790]. SOR [716]. SORAS [2921]. Sorted [3786]. Sound [311, 1333, 2133, 2230]. Source [246, 267, 274, 548, 747, 941, 948, 955, 956, 1041, 1265, 1608, 1609, 1669, 1688, 2495, 2516, 2603, 2739, 2810, 3019, 3224, 3307, 3465, 3734, 3779, 3805]. Sources [38, 2354, 2893]. SP1 [169]. SP2 [169]. Space [342, 367, 505, 555, 625, 631, 634, 671, 1147, 1169, 1240, 1245, 1286, 1287, 1335, 1533, 1608, 1713, 1749, 1751, 1764, 1871, 1882, 1892, 1895, 1996, 2025, 2075, 2133, 2180, 2267, 2282, 2283, 2324, 2382, 2394, 2421, 2427, 2429, 2443, 2481, 2486, 2506, 2518, 2589, 2590, 2675, 2709, 2757, 2772, 2920, 2927, 2942, 2949, 2971, 2993, 3134, 3143, 3172, 3195, 3242, 3253, 3256, 3376, 3467, 3665, 3683, 3773, 3779]. Space-Fractional [1996, 2429, 2443, 2590, 2772, 2927, 2993]. Space-Time [505, 1147, 2025, 2920, 3134, 3376, 3467, 3683]. Spaces [84, 1072, 1160, 2672, 2723, 2910, 2941, 3144, 3254, 3360, 3392, 3405, 3468, 3537, 3584, 3642]. Spacetime [2841]. Sparse [199, 353, 902, 930, 1041, 1159, 1426, 1570, 1582, 1650, 1711, 1736, 1944, 2016, 2077, 2199, 2306, 2410, 2629, 2695, 2769, 2800, 3065, 3097, 3125, 3176, 3202, 3456, 3457, 3558, 3606, 3686, 3786]. Sparse-Grid [930]. Sparsely [1186]. Sparsity [1002, 1003, 1021, 1483, 1523, 1700, 1990, 2260, 2280]. Spatial [169, 405, 534, 910, 1184, 1276, 1562, 1711, 1849, 1938, 1943, 2002, 2192, 2288, 2319, 2391, 2608, 2755, 2930, 3275, 3493, 3526, 3607, 3805]. Spatially [1239, 1544, 1587, 2050, 2663, 3187, 3213]. Spatially-Adaptive [2663]. Spatio [2895, 3183, 3374, 3470]. Spatio-Temporal [2895, 3183, 3374, 3470]. SPDEs [1467, 1994, 2974]. Special [1000, 1649, 1916, 2231, 2448, 2488, 3463]. Specific [3387]. Spectral [53, 98, 125, 143, 145, 199, 242, 254, 271, 284, 305, 308, 312, 322, 324, 329, 340, 345, 352, 358, 399, 412, 420, 421, 447, 474, 507, 517, 521– 524, 531, 587, 601, 624, 641, 644, 693, 747, 761, 803, 835, 862, 880, 882, 934, 993, 1015, 1063, 1068, 1078, 1214, 1225, 1235, 1246, 1272, 1312, 1341, 1369, 1385, 1488, 1509, 1559, 1573, 1634, 1643, 1656, 1725, 2010, 2019, 2029, 2036, 2066, 2110, 2238, 2434, 2475, 2514, 2574, 2675, 2716, 2834, 2894, 2948, 3039, 3087, 3097, 3144, 3157, 3229, 3315, 3330, 3395, 3419, 3486, 3500, 3551, 3583, 3603, 3661, 3742]. Spectral [22, 30, 47, 153, 171, 200, 226, 266, 279, 309, 321, 325, 328, 349– 351, 394, 404, 460, 467, 491, 513, 518, 537, 539–541, 543, 545, 551, 584, 630, 645, 718, 725, 790, 857, 889, 905, 1007, 1091, 1104, 1166, 1202, 1206, 1207, 1222, 1233, 1280, 1342, 1422, 1541, 1572, 1587, 1623, 1650, 1748, 1801, 1840, 1847, 1856, 1870, 1886, 1934, 2111, 2120, 2204, 2225, 2273, 2282, 2291, 2300, 2301, 2326, 2369, 2384, 2385, 2506, 2529, 2538, 2610, 2693, 2832, 2847, 2848, 2959, 2960, 3008, 3055, 3056, 3082, 3089, 3186, 3272, 3310, 3398, 3411, 3455, 3474, 3511, 3578, 3753]. Spectral [12, 24, 33, 68, 80, 86, 89, 103, 124, 137, 140, 142, 155, 159, 170, 174, 194, 346, 535, 549, 669, 704, 755, 759, 805, 824, 1019, 1067, 1093, 1346, 1532, 1534, 1618, 1724, 1754, 1871, 1909, 1988, 2218, 2298, 2518, 2759, 3034, 3287, 3359, 3572, 3762].

Spectral-Element [143, 805, 1222, 1225]. spectral-element-low-order [86]. Spectral-Galerkin [1643, 3097, 3315, 3486, 3511, 3572]. Spectral/ [447, 1988, 3056]. Spectrally [2360, 3802]. Spectrally-Consistent [2360]. Spectroscopy [1041]. Spectrum [67, 1383, 2220, 2874]. Spectrum-Free [2874]. Speed [232, 907, 1170, 2133, 2861, 2899, 3476]. Speed-Up [907]. Speeds [1154, 2042, 3356]. SPH [1456, 2876]. SPH-ALE [2876]. Sphere [121, 886, 918, 1040, 1204, 1328, 1339, 1701, 2948]. Spheres [183, 3371]. Spherical [422, 551, 2036, 2842, 2878, 3034, 3276, 3377, 3551, 3689, 3713]. Spherically [2471]. Spheroidal [1754, 1822, 2035]. Spill [1777]. Spin [1420]. Spin-1 [1420]. Spiral [1006]. Spirals [1379]. Spline [119, 176, 356, 709, 1012, 1356, 1398, 1512, 1699, 2004, 2028, 2063, 2683, 2792, 3479]. Splines [1571, 2836, 3506]. Split [696, 879, 1583, 2723, 2780, 2938, 3057, 3133, 3669]. Split-form [2938, 3057]. Splits [3469]. Splitting [61, 77, 146, 193, 207, 224, 311, 502, 507, 875, 913, 1031, 1178, 1243, 1405, 1440, 1455, 1508, 1583, 1789, 1840, 2058, 2328, 2329, 2364, 2473, 2481, 2498, 2545, 2725, 2727, 2770, 2845, 2851, 2875, 2930, 3040, 3055, 3240, 3257, 3287, 3348, 3369, 3422, 3493, 3605, 3693, 3702, 3733, 3740]. splitting-integrating [146]. Splitting-up [3287]. Splittings [405, 595]. spots [25]. Spread [1903]. Spurious [3643]. Square [617, 972, 2332, 2866, 2977, 3558]. Square-Root [3558]. Squared [2283]. Squares [301, 321, 322, 420, 468, 531, 541, 725, 1086, 1185, 1552, 1668, 1836, 2342, 2657, 2825, 3125, 3150, 3176, 3277, 3412, 3451, 3533, 3571, 3730]. Squares/Reduced [2452]. Squares/Relaxation [2181]. Stabilities [2791]. Stability [15, 158, 309, 314, 332, 361, 369, 425, 483, 502, 503, 534, 586, 638, 718, 739, 742, 749, 860, 882, 921, 982, 997, 1039, 1078, 1098, 1103, 1205, 1210, 1220, 1244, 1292, 1300, 1313, 1344, 1355, 1440, 1456, 1471, 1486, 1603, 1604, 1612, 1631, 1662, 1663, 1701, 1732, 1757, 1759, 1769, 1776, 1778, 1828, 1925, 1945, 2026, 2048, 2072, 2073, 2075, 2120, 2139, 2188, 2218, 2340, 2401, 2411, 2456, 2465, 2502, 2529, 2545, 2661, 2665, 2685, 2727, 2766, 2785, 2789, 2822, 2848, 2894, 2929, 2962, 2979, 2995, 3001, 3020, 3057, 3122, 3237, 3295, 3402, 3405, 3481, 3510, 3550, 3609, 3704]. stability [43]. Stabilization [80, 226, 301, 549, 653, 841, 932, 1207, 1473, 1504, 1556, 1797, 1982, 2106, 2476, 2563, 2614, 2664, 2698, 2859, 2960, 3398, 3717]. Stabilized [107, 449, 466, 842, 988, 1310, 1343, 1417, 1501, 1537, 1554, 1555, 1692, 1719, 1907, 2169, 2226, 2413, 2423, 2674, 2766, 2805, 3167, 3220, 3222, 3255, 3363, 3782]. Stabilizing [3478]. Stable [419, 517, 552, 554, 571, 691, 788, 798, 799, 803, 810, 833, 917, 987, 1018, 1059, 1065, 1143, 1229, 1239, 1277, 1282, 1283, 1370, 1386, 1431, 1481, 1495, 1560, 1714, 1718, 1724, 1770, 1772, 1789, 1791, 1804, 2057, 2070, 2076, 2172, 2177, 2178, 2185, 2186, 2192, 2205, 2207, 2227, 2394, 2406, 2467, 2474, 2498, 2515, 2528, 2559, 2560, 2605, 2630, 2656, 2696, 2745, 2752, 2753, 2767, 2808, 2884, 2886, 2900, 2931, 2938, 2940, 2946, 2958, 3011, 3029, 3040, 3041, 3057, 3082, 3085, 3172, 3174, 3185, 3206, 3285, 3367, 3379, 3390, 3424, 3460, 3485, 3495, 3530, 3588, 3591, 3625, 3679, 3680, 3714, 3746, 3755, 3780]. Stable [480, 851, 1477, 1556, 1593, 1678, 1740, 1788, 1971, 2131, 2190, 2439, 2968, 3168, 3434, 3684, 3696, 3727, 3752]. stably [180]. Stage [982, 1632, 1721, 1896, 2457, 3035, 3782]. Staggered [262, 360, 542, 1497, 1540, 1550, 1843, 2014, 2048, 2074, 2164, 2198, 2239, 2519, 2600, 2728, 3061, 3373, 3448, 3635]. Staggering [651, 1796]. Standard [182, 412, 1596, 1716, 2100, 2325, 3118]. Standing [257]. Standing-Wave [257]. Star [1294, 2319]. Star-Based [1294]. Start [2323]. State [164, 286, 411, 461, 565, 623, 899, 925, 953, 1420, 1444, 1445, 1542, 1573, 1688, 1753, 2070, 2239, 2318, 2356, 2485, 2580, 2763, 2869, 3027, 3265, 3304, 3565, 3727]. statement [1, 85]. States [734, 1910, 2594, 2668]. Static [418, 579, 622, 890, 1025, 1069, 1984, 2002, 2617, 3197, 3404]. Stationary [91, 734, 1400, 1500, 1528, 1537, 1640, 1723, 1742, 1779, 1843, 1884, 1983, 2137, 2145, 2477, 2581, 2669, 2676, 2712, 2714, 2860, 2908, 2932, 3048, 3165, 3396, 3401, 3467, 3703]. Statistical [191, 713, 1127]. Statistics [910]. Steady [164, 248, 262, 286, 351, 403, 411, 461, 565, 623, 702, 853, 925, 1174, 1310, 1330, 1444, 1445, 1542, 1688, 1753, 2023, 2239, 2356, 2363, 2580, 2697, 2763, 2916, 2995, 3255, 3265, 3304, 3397, 3565]. Steady-State [164, 286, 461, 565, 1542, 1688, 1753, 2580, 2763, 3265]. Stefan [71, 1053, 2262, 3094]. Steklov [2385, 2855, 3095, 3745]. Stem [1232]. Stencil [559, 610, 1025, 2067, 2478, 3020, 3304, 3498]. Stencils [3334]. Step [349, 483, 532, 728, 860, 949, 955, 1010, 1147, 1223, 1367, 1459, 1632, 1698, 1781, 1845, 1862, 1926, 1972, 2276, 2422, 2479, 2502, 2738, 2774, 2819, 2962, 3191, 3288, 3575, 3751, 3780, 3802]. Step-by-Step [3802]. Steplength [1499]. Stepping [534, 576, 594, 609, 1233, 1428, 1662, 1663, 1890, 2034, 2190, 2343, 2398, 2419, 2484, 2500, 2527, 2756, 2892, 2942, 2960, 3224, 3285, 3380, 3398, 3399, 3452, 3570, 3754]. Steps [2118, 2500, 2840, 3049, 3206, 3219, 3388, 3460, 3605, 3702, 3759]. Stepsize [991, 3008]. Stewartson [3764]. Stiefel [3462, 3784]. Stiff [63, 267, 274, 480, 487, 800, 946, 1361, 1397, 1440, 1931, 2010, 2220, 2739, 2805, 3064, 3279, 3535]. Stiffened [3445]. Stiffness [2743, 3647]. Stitching [1637]. Stochastic [324, 713, 761, 824, 830, 993, 1023, 1187, 1250, 1263, 1339, 1429, 1542, 1561, 1582, 1588, 1612, 1621, 1708, 1776, 1921, 1957, 1991, 1994, 2132, 2199, 2219, 2223, 2318, 2370, 2375, 2396, 2404, 2425, 2426, 2487, 2493, 2569, 2632, 2708, 2754, 2813, 2818, 2819, 2845, 2868, 2915, 2936, 2974, 2997, 3076, 3128, 3196, 3292, 3314, 3324, 3362, 3372, 3400, 3402, 3417, 3418, 3441, 3561, 3577, 3585, 3666, 3669, 3700, 3708, 3731]. Stokes [3, 67, 101, 127, 137, 138, 192, 193, 262, 290, 327, 343, 345, 412, 420, 520, 532, 558, 597, 614, 772, 816, 853, 876, 877, 934, 988, 990, 1019, 1032, 1052, 1056, 1058, 1089, 1139, 1141, 1235, 1237, 1285, 1304, 1309, 1355, 1386, 1411, 1514, 1519, 1526, 1542, 1550, 1559, 1723, 1728, 1766, 1768, 1805, 1843, 1853, 1931, 1972, 1992, 2002, 2023, 2048, 2088, 2164, 2185, 2186, 2243, 2246, 2289, 2462, 2499, 2674, 2682, 2696, 2755, 2778, 2820, 2868, 2916, 2932, 2964, 2994, 3001, 3021, 3123, 3182, 3185, 3205, 3217, 3232, 3421, 3489, 3501, 3545, 3575, 3608, 3631, 3663, 3703, 3735]. Stokes [22, 36, 87, 227, 321, 426, 430, 450, 455, 502, 509, 541, 702, 728, 787, 788, 790, 805, 1069, 1146, 1174, 1185, 1203, 1247, 1312, 1343, 1349, 1406, 1410, 1417, 1441, 1474, 1502, 1591, 1600, 1619, 1640, 1756, 1854, 1864, 1909, 1971, 1983, 2137, 2150, 2161, 2255, 2256, 2343, 2360, 2374, 2463, 2477, 2482, 2490, 2550, 2558, 2563, 2623, 2646, 2714, 2763, 2803, 2859, 2890, 2914, 2950, 2979, 3025, 3046, 3120, 3167, 3198, 3227, 3252, 3255, 3267, 3278, 3290, 3326, 3343, 3362, 3366, 3367, 3396, 3401, 3405, 3415, 3434, 3556, 3565, 3582, 3587, 3666, 3670, 3715, 3730, 3765, 3798]. Stokes-Like [3405]. Stokes/Darcy [2150, 3227]. Stopping [672, 881, 1242, 2460]. Storage [1614, 2046]. Storm [906]. Stormer [2005, 2054]. Störmer-Cowell [2005]. Stormer-Numerov [2054]. Straight [1735]. Strain [2882, 3243, 3427]. Strained [714]. Strand [1574]. Strang [2481, 2851, 3493]. Strategies [493, 604, 659, 1175, 1679, 2415, 3577]. Strategy [907, 1194, 1806, 1884, 2730, 2779, 2904, 2967, 3026, 3150, 3263, 3304, 3305, 3610, 3657, 3742, 3778]. Stratified [180, 959]. Stream [1803, 1909, 2369, 2490, 2508, 3147, 3586]. Stream-Function-Vorticity [1803]. Streamfunction [702, 816, 877, 1174, 1652, 2499]. streamfunction-velocity [1174]. Streaming [1306]. Streamline [730, 735]. Streamwise [1026]. Stress [157, 313, 698, 858, 2246, 2682, 3198, 3254]. Stress-Assisted [3254]. Stretch [2310]. Strictly [987, 1449]. Strong [314, 361, 425, 483, 492, 534, 739, 749, 982, 1220, 1300, 1471, 1662, 1663, 1684, 1769, 1828, 1925, 1994, 2073, 2139, 2229, 2249, 2411, 2465, 2502, 2682, 2767, 2780, 2894, 2910, 2962, 3076, 3232, 3324, 3402, 3422, 3704]. Strong-Stability-Preserving [314, 534, 982, 1300, 2962]. Strongly [942, 1145, 1577, 1879, 2243, 2246, 2524, 3315, 3690]. Structural [2078, 3404, 3609]. Structure [34, 166, 177, 335, 512, 522, 694, 711, 817, 898, 926, 1017, 1061, 1153, 1379, 1468, 1565, 1850, 1859, 2014, 2020, 2122, 2299, 2449, 2572, 2592, 2621, 2944, 2981, 2986, 3001, 3037, 3044, 3063, 3088, 3121, 3353, 3488, 3524, 3526, 3671, 3688, 3729, 3790, 3791]. Structure-Preserving [1468, 2449, 2592, 2621, 2986, 3044, 3088, 3121, 3488, 3524, 3526, 3671, 3790, 3791]. Structured [656, 911, 1254, 1570, 2080, 2701, 2797, 3693, 3781]. Structures [235, 252, 1956, 2122, 2378, 2471, 2506, 2592, 2773, 3090, 3697, 3774]. Studies [827, 830, 1071, 1746, 3473]. Study [10, 70, 82, 120, 390, 401, 402, 413, 423, 426, 458, 576, 696, 716, 790, 865, 910, 923, 1016, 1076, 1100, 1106, 1111, 1181, 1259, 1292, 1306, 1331, 1385, 1420, 1579, 1613, 1938, 1949, 2012, 2240, 2615, 3493, 3516]. Sub [775, 1460, 1544, 1562, 1783, 1902, 2141, 2338, 3224, 3467, 3528, 3734]. Sub-cell [2338, 3467]. Sub-cells [1783]. Sub-Diffusion [1460, 1544, 1562, 1902, 2141, 3224, 3734]. Sub-diffusions [3528]. Sub-optimal [775]. Subcell [94, 1083, 2600, 2751, 3026, 3638, 3729, 3785]. Subcells [3297]. Subdiffusion [1260, 1344, 1726, 2135, 2388, 2645, 2683, 2774, 2806, 2840, 2892, 2965, 3177, 3293, 3481, 3507, 3531, 3650, 3652]. Subdivision [1373]. Subdomain [391, 1992, 2555]. Subdomain-Wise [1992]. Subdomains [813, 1275, 2378]. Subfilter [960, 965]. Subfilter-Scale [960]. Subgradient [2431, 2936, 2988, 3049]. Subgrid [1310, 1805, 2795]. Subharmonic [745]. Subject [1854]. Sublinear [1049, 3547]. Submanifolds [2629]. Submatrices [3490]. Submerged [289]. submicron [92]. Suboptimal [1736]. Suboptimality [808]. Subordinate [2830]. Subproblem [2741]. Subsampling [2315, 2967]. Subspace [645, 1532, 1755, 1773, 1848, 2095, 2326, 2705, 2905, 3278, 3383, 3389].

Subspaces [3038, 3651]. Substantial [2019]. Substitution [1269]. Substructured [3144]. Substructuring [708]. Subsurface [770, 1524]. Subsurfaces [2007]. Subtraction [1409]. Subtractional [421]. Successive [2734]. Sufficient [1480]. Sufficiently [965]. Suitable [961, 1805, 3168, 3347]. Suliciu [804]. Sum [1137, 2170, 3302, 3329]. Sum-of-Exponentials [3302]. Sum-of-Poles [1137]. Summation [363, 400, 461, 632, 677, 872, 994, 1038, 1042, 1707, 1798, 1886, 2003, 2033, 2086, 2205, 2394, 2528, 2537, 2559, 2767, 2789, 2815, 3020, 3127, 3238, 3279, 3298, 3344, 3714]. Summation-By-Parts [363, 400, 461, 872, 994, 1798, 2003, 2033, 2086, 2205, 2394, 2528, 2537, 2559, 2815, 3020, 3127, 3238, 3279, 3298, 3344, 3714]. Sums [173, 238, 256, 3031]. sup [1556, 1971]. Super [727, 1289, 1511, 1669, 1949, 2055, 2548, 3528]. Super- [3528]. Super-Convergence [1511, 2055]. Super-Resolution [727, 1289, 1669, 2548]. Supercloseness [2045, 2167, 2380, 2824]. Supercomputers [220]. Superconductivity [3341]. Superconsistent [304]. Superconvergence [437, 649, 669, 754, 786, 822, 891, 996, 1192, 1280, 1286, 1351, 1472, 1474, 1515, 1739, 1743, 1849, 1874, 1880, 1887, 1920, 1950, 1978, 2051, 2235, 2303, 2380, 2491, 2509, 2521, 2542, 2625, 2645, 2761, 2897, 2898, 3020, 3068, 3298, 3484, 3636, 3751]. Superconvergent [635, 771, 864, 1047, 1108, 1234, 1729, 1741, 1774, 1924, 2159, 2246, 2537, 2768, 3022, 3171, 3380]. Superdiffusion [3696]. Superfast [3641]. Superiority [1256]. Superlinear [3547]. Supermesh [3094]. Superposition [685]. Supersingular [1545]. Supersolutions [1049]. Supersonic [1123]. Supervised [2006, 3412]. SUPG [1536, 1613, 2106]. Support [1432, 3560]. Supported [1416, 2338]. suppressing [99]. Supra [613]. Supra-Convergent [613]. Surface [48, 59, 148, 364, 384, 392, 457, 465, 523, 599, 692, 770, 780, 828, 879, 980, 1004, 1009, 1020, 1130, 1732, 1912, 1992, 2156, 2224, 2236, 2464, 2595, 2607, 2709, 2758, 2762, 2878, 2975, 3187, 3262, 3273, 3299, 3349, 3373, 3420, 3602, 3644, 3664, 3671]. Surfaces [244, 673, 686, 986, 1085, 1094, 1191, 1255, 1279, 1373, 1379, 1412, 1496, 1506, 1607, 1761, 1897, 2156, 2217, 2248, 2347, 2405, 3452, 3632, 3810]. Surfactant [2018, 2395, 3174, 3452]. Surge [906, 1062]. Surrogate [1933, 2705, 3360]. SVD [2280, 3309, 3490, 3774]. SVD-Based [3774]. SVDsolver [3628]. SWAN [1062]. Swarm [2507]. Sweep [2954]. Sweeping [579, 622, 890, 1057, 1354, 1444, 1445, 1607, 1939, 2534, 2617, 2651, 2838, 2906, 2911, 3265]. Swift [2474, 2886, 3308, 3780]. switch [66]. Switching [785, 852, 1242, 1889, 2096, 2097, 2610, 2715, 3731]. SWMHD [3752]. Sylvester [554]. Symbolic [90, 214]. Symmetric [454, 626, 767, 775, 861, 916, 1200, 1237, 1278, 1340, 1341, 1495, 1638, 1710, 1784, 1785, 1794, 1949, 1970, 2377, 2467, 2471, 2473, 2505, 2618, 2682, 2717, 2948, 3138, 3263, 3296, 3376, 3430, 3468, 3491]. Symmetrical [3204]. Symmetrized [3195]. Symmetry [663, 858, 933, 1048, 1207, 1224, 1885, 1976, 2246, 2971, 3276]. Symmetry-Based [1224]. Symplectic [482, 1566, 2127, 2870]. Synge [3727]. Synge-Type [3727]. System [46, 72, 139, 237, 280, 402, 435, 507, 602, 705, 745, 851, 877, 937, 943, 945, 953, 959, 967, 1285, 1317, 1409, 1560, 1597, 1629, 1697, 1733, 1766, 1806, 1824, 1879, 1947, 2018, 2052, 2127, 2227, 2234, 2262, 2439, 2466, 2510, 2522, 2538, 2582, 2623, 2656, 2696, 2697, 2731, 2755, 2790, 2822, 2840, 2841, 2845, 2849, 2860, 2889, 2904, 2931, 3006, 3068, 3100, 3110, 3146, 3185, 3200, 3208, 3315, 3396, 3424, 3520, 3525, 3613, 3630, 3631, 3680, 3754, 3764]. Systematic [2012]. Systems [17, 214, 217, 218, 259, 316, 332, 374, 382, 404, 447, 454, 468, 484, 487, 548, 555, 565, 592, 617, 627, 687, 713, 748, 781, 856, 908, 933, 940, 944, 946, 952, 1048, 1099, 1117, 1305, 1339, 1361, 1369, 1415, 1449, 1458, 1468, 1543, 1566, 1571, 1580, 1598, 1642, 1653, 1673, 1693, 1757, 1970, 2095, 2171, 2217, 2285, 2313, 2367, 2389, 2445, 2449, 2460, 2496, 2505, 2525, 2530, 2587, 2650, 2726, 2739, 2788, 2796, 2808, 2813, 2828, 2877, 2938, 2971, 3008, 3014, 3018, 3052, 3091, 3099, 3104, 3123, 3152, 3168, 3180, 3184, 3376, 3413, 3421, 3425, 3488, 3599, 3611, 3738, 3770, 3788]. systems [18, 63, 91, 128, 149, 167, 1181, 2286, 2953].

T [3037, 3309]. T-Product [3037, 3309]. t-SVD [3309]. Tailored [703, 791, 836, 855, 924, 977, 1747, 1908, 3043]. Tails [1183]. Tangent [3662]. Tangential [2095]. TASE [3535]. Tau [815, 2525, 3055]. Taylor [355, 458, 1486, 2720, 3278, 3427, 3704]. TDNNS [3503]. Technique [29, 115, 125, 153, 165, 309, 627, 1117, 1151, 1203, 1546, 1838, 1939, 2115, 2200, 2207, 2582, 2590, 2806, 2840, 2873, 3348, 3624, 3781]. Techniques [62, 80, 149, 187, 396, 523, 1398, 1409, 1428, 1435, 1848, 2300, 2507, 2733, 2801, 2896, 2905, 3195, 3278]. Teichmüller [1297]. Telescopic [2130]. Temam [3326]. Temperature [211, 1402, 2083, 2290, 2694, 2712]. Temperature-Dependent [2712]. Tempered [2008, 2096, 2136, 2391, 2582, 2627, 3143]. Template [3741]. Template-Based [3741]. Temporal [264, 502, 1902, 2274, 2437, 2536, 2711, 2895, 3183, 3374, 3428, 3470, 3806]. Temporally [2846]. Ten [2076, 2516]. Ten-Moment [2516]. TENO [3033, 3042, 3112, 3215, 3498]. Tension [3373]. Tensor [746, 1216, 1678, 1755, 1970, 2031, 2246, 2306, 2443, 2450, 2480, 2520, 2537, 2618, 2660, 2695, 2776, 2863, 2895, 2963, 3037, 3052, 3093, 3153, 3162, 3193, 3234, 3269, 3298, 3309, 3311, 3323, 3355, 3358, 3604, 3640, 3656, 3687, 3700, 3707, 3799]. Tensor-Based [1678]. Tensor-Product [746, 2537, 3298]. Tensor-Train [1755, 2443]. Tensorial [2004]. Tensors [1323, 1653, 1654, 2171, 3327, 3350, 3475]. Term [192, 422, 586, 831, 955, 1313, 1510, 1688, 1902, 2180, 2274, 2292, 2516, 2521, 2739, 3307, 3780, 3805]. Terminal [3645]. Terms [207, 267, 274, 548, 747, 941, 948, 1266, 1782, 2033, 2721, 2869, 3224, 3506, 3734]. Ternary [2656]. Terrain [1333]. Tessellations [843, 1390]. Test [37, 472, 1866, 3251]. Testing [229, 1036, 2652]. Tetrahedra [1282, 1283]. Tetrahedral [96, 1234, 1254, 1351, 1717, 2193, 2791, 3082]. Tetrahedron [587, 3097]. texture [147]. Textures [397]. Their [159, 638, 758, 826, 878, 1123, 1160, 2008, 2219, 2220, 2296, 2891, 3138].

Theoretical [47, 1505, 3110]. Theory [2, 37, 42, 64, 73, 122, 216, 247, 333, 359, 856, 1443, 2476, 2922, 3429, 3500, 3573, 3579]. Thermal [204, 248, 2980]. Thermally [3165, 3437]. Thermistor [1535]. Thermo [2378, 2982, 3197]. Thermo-Mechanical [2378, 2982]. Thermo-Poroelasticity [3197]. Thermocapillary [248]. Thermodynamically [2823, 2856, 3318, 3616]. Thermoelasticity [1171]. thermohydrodynamics [130]. Thick [3783]. Thiele [253]. Thin [329, 547, 692, 711, 849, 1065, 1277, 1398, 1571, 1840, 2177, 2314, 2442, 2456, 2836]. Third [254, 598, 904, 944, 1027, 1354, 1543, 1622, 1651, 1671, 1672, 1927, 2016, 2103, 2361, 2456, 2566, 2826, 2986, 3052, 3088, 3449]. Third-Kind [3449]. Third-Order [254, 944, 1027, 1543, 1651, 1927, 2103, 2361, 2566, 2986, 3052, 3088]. Thomas [1399, 3082, 3141, 3310]. Three [15, 54, 56, 207, 213, 219, 260, 293, 446, 465, 499, 557, 632, 714, 720, 746, 830, 988, 1031, 1110, 1157, 1161, 1256, 1312, 1315, 1457, 1486, 1521, 1574, 1608, 1610, 1623, 1788, 1812, 1896, 1897, 1996, 2077, 2170, 2181, 2429, 2492, 2538, 2592, 2704, 2843, 2917, 2981, 2995, 3013, 3246, 3257, 3310, 3397, 3420, 3608, 3632, 3745]. Three-Dimension [2704]. Three-Dimensional [15, 54, 56, 207, 213, 219, 260, 293, 446, 465, 632, 746, 830, 988, 1031, 1110, 1157, 1161, 1312, 1315, 1521, 1574, 1608, 1610, 1812, 1897, 1996, 2181, 2843, 2981, 3310, 3397, 3420]. Three-field [3246]. Three-Operator [3257]. Three-Phase [1788]. Three-Scale [3745]. Three-Stage [1896]. Threshold [1127, 1845, 1973]. Thresholding [639, 1595, 1896, 2353, 3546, 3574, 3606]. THz [2684]. Tight [1212, 1700, 3064]. Tikhonov [1188, 1389, 1773, 3037]. Tilted [2651]. Time [257, 263, 312, 331, 352, 361, 378, 402, 405, 483, 505, 532, 534, 557, 594, 671, 739, 762, 770, 782, 831, 966, 968, 976, 1014, 1168, 1175, 1257, 1371, 1381, 1401, 1446, 1447, 1521, 1590, 1608, 1613, 1659, 1662, 1663, 1665, 1696, 1702, 1713, 1860, 1890, 1899, 1902, 1923, 1962, 1995, 1998, 2044, 2133, 2136, 2180, 2211, 2231, 2235, 2313, 2324, 2391, 2394, 2403, 2408, 2423, 2433, 2443, 2456, 2540, 2640, 2662, 2699, 2756, 2770, 2783, 2805, 2868, 2911, 2917, 2964, 2990, 2994, 3075, 3134, 3218, 3219, 3270, 3333, 3341, 3376, 3459, 3476, 3519, 3531, 3567, 3583, 3676, 3683, 3759, 3760, 3780]. Time [197, 219, 314, 342, 425, 486, 487, 507, 609, 634, 728, 783, 891, 915, 926, 949, 982, 1012, 1222, 1245, 1300, 1335, 1344, 1355, 1367, 1397, 1428, 1465, 1508, 1541, 1667, 1756, 1792, 1821, 1845, 1878, 1882, 1898, 2025, 2034, 2075, 2083, 2135, 2154, 2197, 2225, 2227, 2263, 2282, 2347, 2373, 2402, 2437, 2446, 2447, 2484, 2500, 2506, 2521, 2527, 2541, 2620, 2628, 2645, 2671, 2709, 2740, 2767, 2942, 2960, 2997, 3003, 3019, 3022, 3089, 3096, 3102, 3118, 3129, 3143, 3285, 3306, 3320, 3398, 3399, 3406, 3452, 3467, 3481, 3485, 3492, 3541, 3549, 3581, 3614, 3652, 3654, 3663, 3665, 3685, 3702, 3751, 3754, 3763, 3788, 3800, 3805, 3806]. Time [46, 61, 77, 117, 126, 144, 193, 232, 246, 306, 426, 576, 584, 651, 707, 738, 742, 861, 1061, 1134, 1147, 1172, 1226, 1233, 1291, 1429, 1469, 1477, 1561, 1615, 1644, 1751, 1752, 1764, 1784, 1871, 1877, 1996, 2051, 2064, 2151, 2158, 2188, 2190, 2257, 2274, 2343, 2398, 2419, 2422, 2486, 2518, 2589, 2616, 2661, 2702, 2739, 2742, 2778, 2840, 2859, 2892, 2920, 2962, 2998, 3005, 3032, 3071, 3131, 3145, 3224, 3249, 3253, 3284, 3308, 3343, 3346, 3354, 3369, 3380, 3382, 3417, 3488, 3507, 3552, 3570, 3666, 3734, 3740, 3743, 3779, 3797, 3801]. Time-Accurate [3485]. Time-Based [257]. Time-Caputo-Tempered [2391]. Time-Continuous [3346]. Time-Dependent [77, 126, 232, 312, 352, 532, 770, 831, 966, 976, 1014, 1355, 1371, 1401, 1613, 1923, 1995, 1996, 2064, 2075, 2197, 2235, 2313, 2347, 2403, 2423, 2443, 2446, 2447, 2620, 2640, 2783, 2859, 2868, 2911, 2964, 2994, 3005, 3019, 3218, 3333, 3341, 3354, 3406, 3417, 3459, 3492, 3583, 3663, 3788]. Time-Discretization [982]. Time-Domain [246, 707, 915, 1381, 1447, 1521, 1644, 2742]. Time-Fractional [1344, 1477, 1752, 1756, 1877, 1878, 1962, 2083, 2151, 2188, 2437, 2521, 2589, 2620, 2671, 2699, 2702, 3019, 3022, 3118, 3129, 3131, 3249, 3270, 3306, 3382, 3507, 3531, 3549, 3567, 3683, 3740, 3743, 3797, 3801, 3806]. Time-Harmonic [861, 891, 2051, 2433, 3676]. Time-Limited [3654]. Time-Limiters [378]. Time-Marching [144, 197, 219]. Time-Parallel [3581]. Time-Space [2518, 2942, 3143]. Time-Spectral [2225, 3089]. Time-Splitting [61, 507]. Time-Stepping [534, 576, 1428, 1662, 1663, 1890, 2190, 2398, 2419, 2484, 2527, 2942, 2960, 3224, 3285, 3398, 3452]. Time-Steps [3702]. Time-Tempered [2136]. Time-Transient [2628]. Time/Space [2133]. Times [1719, 2102, 2130]. Timestepping [655]. Timoshenko [526, 850, 1548, 3480]. Tissue [1182, 1232]. Toeplitz [217, 2930, 3738]. Tokamak [3555]. Tolerance [1194]. tomographic [26, 187]. Tomography [1002, 1008, 1021, 1826, 2663, 3036, 3454, 3733]. Toolbox [242, 689]. Toolkit [966]. Top [578]. Top-Hat [578]. Topographic [3420]. Topography [1965, 2812, 3184]. Topological [1393, 2827]. Topologies [3741]. Topology [1897]. Total [214, 397, 525, 639, 650, 986, 1252, 1290, 1364, 1570, 1637, 1786, 1817, 1851, 2114, 2292, 2480, 2548, 2704, 2817, 3107, 3264, 3323, 3349, 3542, 3617, 3792]. TR [1769]. Trace [631, 1563, 3619]. Trace-Penalty [1563]. Tracer [1328]. Tracing [730, 1725, 1806, 2189, 2284]. Tracking [94, 374, 555, 621, 792, 2813]. Traditional [2415]. Traffic [721, 1392, 2501, 2895]. Train [1755, 2443, 2480]. Training [1933, 2571, 2967, 3189, 3365]. Trajectories [1132, 2869]. Transactional [1119]. Transfer [200, 205, 625, 1371, 2145, 2251, 2269, 2603, 2615, 2794, 2842, 2980, 3072, 3080, 3181, 3356, 3417, 3607]. Transferable [3756]. Transfinite [225]. Transform [78, 150, 1142, 1589, 1811, 1815, 1955, 2097, 2157, 2165, 2930, 3234, 3378, 3416, 3472, 3578]. Transformation [903, 995, 1611, 1812, 3624]. Transformation-free [903]. Transformations [400, 626, 2894]. Transformed [280, 2776, 3309]. Transforms [3202, 3311]. Transient [164, 437, 600, 740, 941, 1103, 2278, 2628, 2794, 3227]. Transients [223]. Transition [57, 214, 1281, 3006]. Transitional [305, 1469]. Transitions [319]. Translating [605, 903]. Transmission [1211, 1225, 1298, 1422, 1639, 1730, 1867, 2036, 2270, 2534, 2549, 2604, 3327, 3482, 3540, 3551, 3579]. Transmittance [806]. Transparent [1656, 1881]. Transplantation [863]. Transport [97, 207, 231, 236, 240, 382, 456, 489, 527, 566, 951, 1328, 1361, 1919, 1921, 2044, 2093, 2182, 2262, 2291, 2435, 2445, 2657, 2685, 3017, 3068, 3113, 3114, 3265, 3321, 3406, 3613, 3633, 3652, 3685]. Transportation [2533, 2896].

Transpose [1770]. Transverse [1659, 2342]. Transversely [2651]. Trapezoidal [1545, 3130]. Travel [2917]. Traveling [2919]. Treat [1337]. Treatment [109, 799, 829, 890, 937, 1209, 1380, 1600, 1759, 1852, 2949, 3291]. Tree [252, 675]. Treecode [2557]. Trees [3757]. Trefftz [2555]. Tri [1515]. Tri-linear [1515]. Trial [2639]. triangles [98]. Triangular [268, 373, 598, 662, 981, 1018, 1161, 1197, 1334, 1411, 1681, 1740, 1824, 1900, 1942, 2167, 3052, 3192, 3331, 3398, 3685]. Triangularized [3325]. Triangulated [986, 1761, 1895, 3349]. Triangulation [673]. Triangulations [1735, 3290]. Tridiagonal [167, 217]. Trigonometric [79, 173, 1160, 1301, 2930]. Triple [1270, 1450, 2895]. Triplets [3628]. Trivariate [3479]. Trivial [323]. Trotter [2845]. Troubled [3283]. Troubled-Cell [3283]. True [3289]. Truncated [2857, 3438, 3457]. Truncation [35, 365, 1202, 1436, 2273, 3654]. Truncations [2174, 2544]. Trust [1479, 2904, 3263, 3455, 3580]. Trust-Region [2904, 3263, 3455, 3580]. Tseng [3605]. Tsunami [1946]. Tube [667]. Tubular [512]. Tucker [2867, 3475]. Tukey [1928]. Tumor [688, 1090, 2507]. Tunable [1993]. Tunneling [779]. Turbulence [2, 9, 23, 37, 49, 65, 68, 73, 75, 82, 103, 104, 157, 166, 177, 179, 181, 186, 230, 235, 243, 247, 346, 572, 610, 880, 2078, 2315, 2670, 3692]. Turbulence-Degraded [2315]. Turbulent [25, 32, 112, 123, 151, 187, 211, 227, 231, 236, 303, 431, 539, 920, 1170, 1320, 1574, 1578, 2240, 2823, 3318]. Turek [2768]. Turing [285, 469]. TV [376, 668, 727, 990, 1032, 1118, 1628, 1864, 2280]. TV-Regularization [727]. TV-Stokes [990, 1032]. TVD [118, 318, 471, 922, 1823, 1855]. Two [204, 212, 249, 371, 448, 496, 552, 581, 615, 625, 627, 649, 729, 748, 791, 792, 860, 916, 937, 943, 944, 975, 979, 983, 1010, 1053, 1077, 1155, 1156, 1165, 1173, 1201, 1223, 1231, 1311, 1362, 1386, 1400, 1411, 1512, 1524, 1568, 1577, 1608, 1659, 1662– 1664, 1683, 1698, 1702, 1712, 1718, 1741, 1768, 1772, 1802, 1873, 1919, 1954, 1972, 1997, 2007, 2011, 2088, 2180, 2328, 2329, 2339, 2349, 2364, 2366, 2391, 2421, 2553, 2570, 2591, 2655, 2662, 2678, 2902, 2916, 2926, 2996, 3011, 3028, 3035, 3066, 3191, 3229, 3274, 3313, 3373, 3493, 3567, 3575, 3608, 3621, 3637, 3679, 3722, 3738, 3746]. Two [3, 54, 191, 240, 243, 293, 314, 349, 353, 367, 494, 571, 618, 633, 647, 676, 730, 734, 759, 760, 809, 827, 849, 927, 1029, 1059, 1066, 1069, 1083, 1125, 1158, 1217, 1232, 1310, 1317, 1370, 1427, 1459, 1500, 1557, 1636, 1652, 1710, 1715, 1721, 1743, 1752, 1764, 1779, 1781, 1878, 1888, 1951, 1977, 2027, 2119, 2156, 2162, 2322, 2352, 2368, 2388, 2479, 2502, 2581, 2594, 2612, 2614, 2634, 2651, 2664, 2670, 2711, 2726, 2879– 2881, 2900, 2950, 2973, 2993, 3002, 3012, 3014, 3041, 3042, 3055, 3109, 3133, 3175, 3224, 3227, 3268, 3288, 3297, 3303, 3310, 3345, 3399, 3432, 3518, 3597, 3616, 3650, 3692, 3803]. two [7, 10, 52, 101, 120, 158, 175, 176, 179, 183]. Two- [54, 293]. Two-Body [647]. Two-Component [734, 2900]. Two-D [3803]. Two-Derivative [1662, 1663, 1715]. Two-Dimensional [3, 7, 10, 52, 101, 158, 179, 191, 212, 243, 249, 349, 367, 496, 552, 627, 649, 729, 730, 748, 944, 1066, 1077, 1165, 1232, 1362, 1427, 1512, 1568, 1636, 1659, 1664, 1702, 1710, 1768, 1919, 1951, 1997, 2162, 2180, 2322, 2364, 2368, 2421, 2496, 2553, 2591, 2612, 2651, 2881, 3002, 3268, 3297, 3637, 3722]. Two-Equation [240, 2670, 3692]. Two-Fluid [581, 633, 1059, 1311, 3616, 3679]. Two-Grid [927, 975, 979, 1069, 1155, 1310, 1577, 1683, 1741, 1878, 1954, 1972, 2662, 2879, 2880, 2916, 2973, 3066, 3175, 3227, 3229, 3567, 3738]. Two-Layer [937, 943, 1156, 2339, 3012, 3746]. Two-Level [175, 448, 760, 916, 1201, 1400, 1500, 1608, 1764, 1779, 1977, 2388, 2581, 2634, 3432, 3493]. Two-Medium [1083, 2678]. Two-Norm [1217]. Two-Parameter [204, 2996]. Two-Phase [371, 615, 618, 676, 792, 809, 849, 1053, 1125, 1231, 1370, 1386, 1524, 1557, 1712, 1718, 1772, 1888, 2614, 2950, 3041, 3345, 3373, 3399, 3597, 3621, 3637]. Two-Point [3011]. Two-Scale [176, 353, 3109]. Two-Sided [2349, 2993]. Two-Stage [1721, 3035]. Two-Step [860, 1010, 1223, 1698, 1781, 2479, 2502, 3191, 3288]. Twogrid [3504]. Type [101, 163, 308, 559, 652, 661, 701, 848, 856, 857, 926, 1045, 1055, 1070, 1091, 1117, 1199, 1280, 1466, 1493, 1510, 1580, 1612, 1742, 1776, 1799, 1816, 1826, 1845, 1927, 1949, 1952, 1972, 2080, 2124, 2126, 2137, 2141, 2151, 2215, 2445, 2462, 2476, 2497, 2536, 2569, 2573, 2579, 2624, 2688, 2723, 2731, 2824, 2844, 2941, 3005, 3037, 3051, 3102, 3220, 3328, 3352, 3502, 3549, 3576, 3622, 3719, 3723, 3727]. Types [2493, 2693].

Ultra [2287, 2509, 2897]. Ultra-Weak [2287, 2509, 2897]. Ultrashort [859]. Ultrasound [1851]. Ultraspherical [3583]. Unbalanced [2094]. Unbounded [703, 805, 978, 1088, 1093, 1299, 1509, 1680, 1871, 2116, 2341, 2603, 2675, 2983, 3107, 3315, 3438, 3474, 3548, 3801]. Uncertain [2016, 2145, 3417, 3577]. Uncertainties [3744]. Uncertainty [761, 1530, 1650, 2319, 2495, 2912, 3788]. Unconditional [1535, 1739, 1887, 1945, 2151, 2388, 2811, 2832, 2979, 3419, 3567]. Unconditionally [691, 851, 1239, 1431, 1560, 1714, 1740, 1769, 1770, 1791, 1879, 2070, 2135, 2227, 2361, 2474, 2753, 2802, 2900, 2986, 3025, 3029, 3088, 3174, 3390, 3424, 3485, 3495, 3588, 3591, 3625, 3696, 3723]. Unconditioned [3172]. Unconstrained [807, 1619]. Under-Expanded [1722]. Under-Sampled [3154]. Underdetermined [2904]. Underlying [269, 759]. Undersampled [1589]. Understanding [386, 3409]. Unequal [3334]. Unequal-Size [3334]. Unfitted [2827, 2950, 3070, 3123, 3141, 3337, 3410, 3539, 3621, 3664, 3703]. unforced [120]. Uniaxial [1456]. Unidirectional [1319]. Unified [602, 730, 858, 892, 915, 1353, 1581, 2163, 2922, 3063, 3283, 3573, 3603, 3607, 3646]. Uniform [176, 403, 606, 889, 941, 1039, 1244, 1302, 1338, 1453, 1474, 1558, 1747, 1760, 1811, 1819, 1858, 1997, 2028, 2029, 2059, 2071, 2410, 2526, 2550, 2636, 2687, 2688, 2927, 2995, 3068, 3379, 3466, 3488, 3665, 3793]. Uniformly [766, 1853, 2387, 2707, 3207]. Unifying [3238]. Unilateral [1454, 2716, 3521, 3538]. Uniquely [2178, 2696, 3588]. Uniqueness [3805]. Unisolvence [3468]. Unisolvency [3204]. Unit [2948, 3342]. Unity [1296, 1433, 1482, 1617, 2088, 2340, 2630, 3779]. Univariate [2907]. Universal [3139, 3647]. Unknown [2397]. Unmodified [2107]. Unrestricted [1147]. Unsplit [782]. Unsteady [34, 224, 281, 328, 345, 558, 1236, 1608, 1932, 2457, 2588, 2703, 3735].

Unstructured [208, 298, 404, 414, 540, 630, 730, 758, 906, 1027, 1062, 1119, 1131, 1197, 1208, 1336, 1525, 1578, 1674, 1675, 1745, 1843, 1858, 2104, 2180, 2368, 2553, 2679, 2844, 2862, 2966, 3026, 3215, 3498]. Unstructured-Mesh [1062]. Unsupervised [3264]. Unsymmetric [18]. Update [332, 523, 1872]. Updating [2001, 2904, 3457]. upon [2697]. Upper [1217, 1249, 1357, 3281]. Upscaled [2250]. Upwind [118, 347, 433, 536, 848, 941, 956, 1025, 1105, 1661, 1693, 2199, 2308, 2469, 2583, 2738, 2838, 2943, 3086, 3147, 3223, 3463, 3517, 3569, 3625, 3671]. Upwind-Biased [2308, 2583, 3086, 3147, 3463]. Upwind-Difference [1105]. Upwind-SAV [3671]. URV [3244]. Use [133, 220, 547, 702, 1657, 1694]. Using [205, 241, 259, 276, 277, 326, 377, 383, 523, 535, 558, 559, 629, 667, 694, 695, 706, 751, 763, 792, 796, 821, 843, 847, 848, 867, 884, 908, 923, 942, 944, 994, 1004, 1005, 1020, 1080, 1095, 1117, 1123, 1127, 1140, 1245, 1284, 1307, 1338, 1405, 1458, 1462, 1467, 1557, 1559, 1572, 1589, 1660, 1705, 1775, 1795, 1819, 1848, 1895, 1917, 1967, 2017, 2053, 2089, 2146, 2248, 2260, 2264, 2265, 2292, 2314, 2342, 2435, 2450, 2480, 2484, 2494, 2708, 2746, 2804, 2831, 2864, 2905, 2991, 3024, 3026, 3097, 3152, 3170, 3297, 3357, 3413, 3477, 3478, 3536, 3556, 3615, 3643, 3706, 3707, 3737, 3770]. Using [36, 117, 162, 173, 346, 461, 681, 704, 823, 1023, 1185, 1215, 1391, 1442, 1483, 1594, 1755, 1767, 1783, 1845, 1976, 1990, 2039, 2109, 2121, 2194, 2362, 2487, 2546, 2595, 2736, 2792, 2966, 3003, 3047, 3059, 3082, 3115, 3278, 3283, 3522, 3593, 3733]. Utility [2715]. Uzawa [143, 313, 509, 1200, 1226].

V [2281]. V-cycle [2281]. Vacuum [2842]. Validation [1777, 2002]. Value [28, 106, 144, 185, 198, 222, 297, 317, 440, 500, 503, 723, 762, 833, 895, 1051, 1093, 1302, 1313, 1346, 1457, 1513, 1538, 1623, 1749, 1753, 1790, 1817, 1913, 1934, 2034, 2111, 2119, 2125, 2352, 2531, 2669, 2670, 2740, 2898, 3149, 3158, 3303, 3352, 3380, 3629, 3645, 3690, 3707, 3803]. Valued [95, 589, 644, 685, 870, 1394, 2182, 2705]. Values [88, 1743, 3384]. van-der-Waals [953]. Vanishing [539, 731, 1747]. Variable [110, 283, 370, 590, 626, 791, 987, 1038, 1042, 1239, 1269, 1350, 1370, 1544, 1842, 2075, 2100, 2133, 2141, 2283, 2308, 2349, 2376, 2576, 2583, 2671, 2708, 2740, 2772, 2831, 2866, 2993, 3030, 3077, 3129, 3131, 3206, 3219, 3301, 3306, 3388, 3426, 3460, 3495, 3531, 3576, 3591, 3612, 3650, 3664, 3702, 3751, 3772]. Variable-Coefficient [987, 1239]. Variable-Flux [2576]. Variable-Order [2740, 3129, 3131, 3306, 3531, 3576, 3612, 3650, 3772]. Variable-Time-Step [3751]. Variables [1323, 2169, 2572]. Variably [2340, 2836, 3306]. Variance [2260, 2936, 3372, 3561]. Variance-Based [2936]. Variant [905, 2424, 2858, 3414]. Variants [2117, 2726]. Variation [397, 525, 639, 650, 986, 1252, 1290, 1364, 1570, 1637, 1786, 1851, 2114, 2292, 2480, 2548, 2704, 2817, 3107, 3264, 3323, 3349, 3542, 3792]. Variation-Based [1252]. Variational [273, 726, 826, 847, 1056, 1266, 1270, 1327, 1335, 1613, 1660, 1676, 1703, 1816, 1960, 2283, 2309, 2315, 2351, 2390, 2431, 2462, 2542, 2663, 2672, 2780, 2875, 2910, 2936, 2941, 2988, 3045, 3049, 3121, 3133, 3134, 3251, 3316, 3537, 3595, 3623, 3715, 3779]. Variational-Hemivariational [3595]. Variations [1070, 1765]. Various [3530]. Varying [560, 609, 1156, 2230]. Vector [870, 888, 930, 1394, 1432, 2182, 2183, 2726, 2889]. Vector-Valued [870, 2182]. Vectorial [3542]. Vectors [1157]. Velocity [247, 313, 702, 934, 1019, 1052, 1174, 1652, 2109, 2161, 2217, 2436, 3171, 3198, 3237, 3502, 3513]. Velocity-Based [3237]. VEM [3348]. Venant [1409]. Verifiability [3316]. Verifying [3392]. Version [500, 608, 808, 1155, 1491, 1870, 1999, 2453, 3286]. Versions [2253]. Versus [399, 1079, 2415, 2652, 3232]. Vertex [1809]. Vertex-Centered [1809]. Very [93, 142, 1843]. Vesicle [3428]. VI [3623]. VI-DGP [3623]. Via [42, 117, 554, 594, 686, 725, 793, 813, 814, 1049, 1087, 1115, 1118, 1120, 1186, 1212, 1269, 1297, 1420, 1470, 1523, 1570, 1611, 1669, 1735, 1820, 1973, 1986, 2004, 2195, 2340, 2410, 2425, 2427, 2498, 2551, 2554, 2561, 2607, 2758, 2836, 2877, 2894, 2895, 2919, 2934, 2963, 2987, 3008, 3016, 3035, 3093, 3104, 3162, 3177, 3178, 3271, 3292, 3309, 3311, 3368, 3456, 3476, 3523, 3589, 3598, 3610, 3633, 3651, 3653, 3656, 3720]. Vibration [1548]. Video [3350]. VIEs [3486]. View [1581, 3409, 3532]. Violating [581]. Virtual [2126, 2253, 2264, 2265, 2289, 2390, 2482, 2619, 2646, 2674, 2714, 2837, 2855, 2924, 3028, 3067, 3098, 3217, 3222, 3246, 3274, 3294, 3405, 3421, 3491, 3539, 3595, 3661, 3670, 3705, 3721, 3722, 3730, 3737, 3777]. Virus [1930]. Viscoelastic [312, 313, 352, 531, 676, 2299, 2342, 3480]. Viscosities [2363]. Viscosity [112, 142, 145, 151, 390, 539, 818, 880, 943, 961, 965, 1531, 2290, 2300, 2413, 2641, 2692, 2841, 3228, 3363, 3379, 3672]. Viscosity-Independent [2413]. Viscous [3, 48, 50, 156, 224, 281, 309, 355, 602, 667, 671, 676, 790, 903, 938, 1078, 1336, 1712, 1715, 1718, 1788, 1961, 2122, 2338, 2487, 2664, 3373, 3474, 3525, 3796]. Visibility [684, 2517]. Visualization [359]. Visualizing [3633]. Vlasov [280, 1181, 1733, 1806, 1947, 1951, 2189, 2359, 2364, 2822, 3525]. Void [720, 1569]. Voigt [3252]. Volatility [824, 3731]. Voltage [3083]. Volterra [1091, 1104, 1491, 1572, 1618, 1709, 1790, 1870, 2069, 2101, 2277, 2384, 2434, 2565, 2636, 2840, 3449]. Volume [287, 389, 404, 433, 488, 511, 527, 587, 657, 666, 700, 790, 800, 834, 893, 905, 907, 936, 949, 1040, 1064, 1102, 1192, 1238, 1259, 1311, 1332, 1424, 1464, 1502, 1525, 1537, 1587, 1593, 1597, 1605, 1645, 1647, 1651, 1674, 1675, 1688, 1695, 1740, 1783, 1796, 1843, 1861, 1862, 1889, 1900, 1920, 1946, 1952, 1978, 1983, 2028, 2032, 2068, 2104, 2143, 2199, 2213, 2259, 2335, 2429, 2530, 2543, 2600, 2615, 2641, 2671, 2679, 2751, 2790, 2823, 2850, 2852, 2862, 2884, 2896, 2966, 2977, 3039, 3064, 3148, 3215, 3266, 3373, 3397, 3467, 3498, 3559, 3566, 3590, 3603, 3701, 3715, 3718, 3762, 3767, 3798, 3809]. volume [60, 96, 190]. Volume-Augmented [1102]. Volume-Complete [893]. Volume-Measure-Preserving [2896]. Volume-Preserving [389]. volume/finite [190]. Volumetric [2747]. Voronoi [843, 1390]. Vortex [59, 311, 388, 458, 2064]. Vortical [2506]. Vortices [393]. Vorticity [934, 1019, 1052, 1803, 1854, 1909, 2240, 2436, 2490, 3513, 3796]. Vorticity-Stream [1909, 2490]. Vries [1633, 1927, 1974, 3186, 3359]. vs [182, 357].

W [2263]. Waals [953]. WAF [701]. Walks [3699]. Wall [597, 1229, 1469, 2615, 3232]. Wang [1916]. Wasserstein [2444, 3689]. Waste [1862]. Water [439, 446, 447, 527, 552, 618, 701, 777, 818, 829, 849, 928, 936–938, 943, 947, 950, 954, 957, 985, 1033, 1156, 1179, 1197, 1384, 1409, 1593, 1629, 1670, 1679, 1823, 1830, 1912, 1929, 2100, 2339, 2407, 2605, 2630, 2692, 2710, 2812, 2823, 2849, 2876, 2903, 2987, 3030, 3184, 3223, 3237, 3260, 3262, 3318, 3471, 3684, 3729, 3746, 3755]. Wave [29, 219, 239, 257, 282, 339, 366, 369, 399, 486, 487, 504, 516, 543, 624, 637, 651, 658, 685, 774, 799, 936, 948, 973, 977, 1044, 1077, 1157, 1184, 1245, 1265, 1268, 1321, 1335, 1380, 1429, 1447, 1487, 1512, 1576, 1629, 1666, 1702, 1770, 1798, 1813, 1822, 1880, 1881, 1891, 1958, 1965, 2003, 2004, 2017, 2133, 2140, 2154, 2179, 2180, 2193, 2208, 2225, 2242, 2254, 2261, 2274, 2336, 2382, 2414, 2418, 2492, 2527, 2571, 2589, 2606, 2628, 2677, 2809, 2827, 2831, 2832, 2848, 2917, 2919, 2938, 2998, 3028, 3127, 3128, 3145, 3226, 3242, 3279, 3315, 3474, 3514, 3665, 3694, 3716, 3726, 3729, 3805]. Wave [102, 113, 161, 246, 1477, 1754, 2573, 2781, 3296, 3426]. Wave-Type [2573]. Wavefield [3202]. Waveform [399, 1877, 2624, 3581]. Waveguides [1103, 3660]. Wavelength [275]. Wavelet [132, 221, 303, 323, 407, 409, 497, 585, 586, 639, 823, 1008, 1088, 1120, 1121, 1216, 1395, 1523, 2060]. Wavelet-Based [1008]. Wavelets [241, 2736, 3757]. Wavenumber [191, 1822, 3370, 3744]. Waves [59, 275, 393, 457, 501, 518, 660, 818, 823, 837, 942, 994, 1062, 1150, 1253, 1912, 2054, 2339, 2709, 2909]. Wavy [2615]. Wavy-Wall [2615]. Way [2336, 3289]. Weak [362, 857, 858, 961, 994, 1187, 1273, 1274, 1304, 1317, 1383, 1442, 1475, 1524, 1673, 1797, 1805, 1826, 1885, 2012, 2062, 2152, 2195, 2287, 2289, 2357, 2509, 2526, 2707, 2718, 2754, 2897, 3232, 3281, 3376, 3443, 3634, 3665, 3740]. Weakly [41, 767, 794, 818, 916, 1237, 1267, 1410, 1435, 1530, 1572, 1709, 1724, 1870, 1987, 2069, 2238, 2434, 2483, 2565, 2636, 2737, 3143, 3153, 3486, 3597, 3601]. Weakly-Intrusive [1530]. Weight [2275]. Weighted [580, 732, 928, 1028, 1030, 1077, 1320, 1448, 1575, 1584, 1587, 1594, 1689, 1726, 1811, 1847, 1884, 1886, 1944, 1979, 2056, 2098, 2455, 2478, 2577, 2720, 2963, 2970, 3004, 3534, 3689, 3779, 3787]. Weighting [1115]. Weights [258, 1248, 1314, 1584, 2046, 2497]. Well [87, 99, 548, 936, 948, 952, 957, 998, 1033, 1102, 1133, 1179, 1197, 1409, 1558, 1592, 1597, 1621, 1804, 1830, 1832, 1993, 2530, 2605, 2710, 2712, 2772, 2812, 2852, 2876, 2903, 2937, 2987, 3030, 3223, 3237, 3262, 3517, 3631, 3718]. Well-Balanced [548, 936, 948, 952, 957, 1033, 1102, 1133, 1179, 1197, 1409, 1558, 1592, 1597, 1621, 1830, 1832, 2530, 2605, 2710, 2812, 2852, 2876, 2903, 2937, 3030, 3223, 3237, 3262, 3517, 3718]. Well-Balancing [2987]. Well-conditioned [998]. Well-Posedness [87, 99, 2712, 2772, 3631]. Weller [1980, 2055]. Wendroff [603, 926, 928, 1472, 1759, 1792, 1799, 2536, 2720, 3442, 3785]. Wendroff-Type [1799]. WENO [287, 318, 391, 533, 548, 565, 610, 623, 656, 758, 833, 890, 907, 921, 923, 925, 1033, 1083, 1131, 1133, 1150, 1160, 1198, 1208, 1314, 1332, 1385, 1404, 1438, 1464, 1525, 1539, 1543, 1616, 1622, 1627, 1632, 1722, 1758, 1794, 1839, 1883, 1928, 1938, 1952, 1981, 2016, 2103, 2275, 2469, 2497, 2516, 2566, 2585, 2599, 2617, 2622, 2719, 2800, 2839, 2852, 2885, 2927, 2978, 3000, 3026, 3126, 3183, 3265, 3297, 3322, 3334, 3364, 3467, 3566, 3615, 3790, 3794]. WENO-Z [2103]. WENO3 [1651, 2221]. WENO5 [2661]. Wet [1179, 2710, 2987, 3262]. Wet/Dry [1179, 2987]. Wetting [985]. Where [3239]. Whitham [721]. Whole [3186, 3242]. Wide [2327, 2722]. Width [1156]. Wiener [546]. Wigner [2024, 2341, 2669]. Willmore [780, 2147]. Wilson [2833]. Wind [1278, 1699]. Wind-Driven [1699]. Wise [209, 1262, 1992, 2852]. Within [586]. Without [340, 502, 755, 1065, 1277, 1463, 1684, 2177, 2197, 2284, 2354, 2364, 2442, 2528, 2924, 3122, 3284, 3307, 3309, 3496, 3647]. WKB [779, 2573]. Wolf [3261]. Work [2489]. Wormhole [2000, 2535, 2933]. Worsey [3469]. WSGD [2162, 3172]. WSGD-OSC [2162].

X [1053, 1120, 1569]. X-HDG [1569]. X-Ray [1120]. XFEM [2262]. XFEM/Level [2262]. Xu [2816].

Yawed [337]. Yee [750]. Yield [2236]. Yuan [3014].

Z [2103]. Zabolotskaya [1922]. Zakharov [507, 1271, 3381]. Zenith [806]. Zero [2549]. Zero-Index [2549]. Zeroes [218]. Zonal [36, 1822]. Zone [521, 1157]. Zones [142, 3521]. ZQ [2617].

References

Orszag:1986:IS

[1] Steven A. Orszag. Introductory statement. Journal of Scientific Computing, 1(1):1–2, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print),

      1573-7691 (electronic). URL http://link.springer.com/article/10.
      1007/BF01061451;   http://link.springer.com/content/pdf/10.
      1007/BF01061451;   http://www.springerlink.com/openurl.asp?
      genre=article&issn=0885-7474&volume=1&issue=1&spage=1-2.

Yakhot:1986:RGA

[2] Victor Yakhot and Steven A. Orszag. Renormalization group analysis of turbulence. I. Basic theory. Journal of Scientific Computing, 1(1):3–51, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print),

      1573-7691 (electronic). URL http://link.springer.com/article/10.
      1007/BF01061452;   http://link.springer.com/content/pdf/10.
      1007/BF01061452;   http://www.springerlink.com/openurl.asp?
      genre=article&issn=0885-7474&volume=1&issue=1&spage=3-51.

Nosenchuck:1986:TDN

[3] Daniel M. Nosenchuck and Michael G. Littman. Two-dimensional nonsteady viscous flow simulation on the Navier–Stokes Computer miniNode. Journal of Scientific Computing, 1(1):53–73, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01061453;
http://link.springer.com/content/pdf/10.1007/BF01061453;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=1&issue=1&spage=53-73.

Orszag:1986:BCI

[4] Steven A. Orszag, Moshe Israeli, and Michel O. Deville. Boundary conditions for incompressible flows. Journal of Scientific Computing, 1(1):75–111, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061454;   http://link.springer.com/content/pdf/10.
1007/BF01061454;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=1&issue=1&spage=75-111.

Anonymous:1986:IA

[5] Anonymous. Important announcement. Journal of Scientific Computing, 1(2):113, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061387;   http://link.springer.com/content/pdf/10.
1007/BF01061387;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=1&issue=2&spage=113.

Metropolis:1986:MPP

[6] N. Metropolis. Massively parallel processing. Journal of Scientific Computing, 1(2):115–116, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061388;   http://link.springer.com/
content/pdf/10.1007/BF01061388;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=1&issue=2&spage=
115-116.

McFadden:1986:BIM

[7] G. B. McFadden, P. W. Voorhees, and R. F. Boisvert. A boundary integral method for the simulation of two-dimensional particle coarsening. Journal of Scientific Computing, 1(2):117–144, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061389;   http://link.springer.com/content/pdf/10.
1007/BF01061389;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=1&issue=2&spage=117-144.

Hannon:1986:MDS

[8] L. Hannon, G. C. Lie, and E. Clementi. Molecular dynamics simulation of flow past a plate. Journal of Scientific Computing, 1(2): 145–150, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061390;   http://link.springer.com/content/pdf/10.
1007/BF01061390;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=1&issue=2&spage=145-150.

Dannevik:1986:ESN

[9] William P. Dannevik. Efficient solution of non-Markovian covariance evolution equations in fluid turbulence. Journal of Scientific Computing, 1(2):151–182, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061391;   http://link.springer.com/content/pdf/10.
1007/BF01061391;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=1&issue=2&spage=151-182.

Boyd:1986:ANS

[10] John P. Boyd. An analytical and numerical study of the twodimensional Bratu equation. Journal of Scientific Computing, 1(2): 183–206, ???? 1986. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061392;   http://link.springer.com/content/pdf/10.
1007/BF01061392;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=1&issue=2&spage=183-206.

Anonymous:1987:IA

[11] Anonymous. Important announcement. Journal of Scientific Computing, 2(1):1, March 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061509;   http://link.springer.com/content/pdf/10.
1007/BF01061509;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=1&spage=1.

Dubiner:1987:AAS

[12] Moshe Dubiner. Asymptotic analysis of spectral methods. Journal of Scientific Computing, 2(1):3–31, March 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061510;   http://link.springer.com/
content/pdf/10.1007/BF01061510;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=2&issue=1&spage=
3-31.

Goldhirsch:1987:EMC

[13] L. Goldhirsch, Steven A. Orszag, and B. K. Maulik. An efficient method for computing leading eigenvalues and eigenvectors of large asymmetric matrices. Journal of Scientific Computing, 2(1):33–58, March 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061511;   http://link.springer.com/content/pdf/10.
1007/BF01061511;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=1&spage=33-58.

Morison:1987:SDF

[14] R. Morison and S. Otto. The scattered decomposition for finite elements. Journal of Scientific Computing, 2(1):59–76, March 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01061512;   http://link.springer.com/content/pdf/10.1007/
BF01061512;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=2&issue=1&spage=59-76.

Malik:1987:LSA

[15] Mujeeb R. Malik and Steven A. Orszag. Linear stability analysis of threedimensional compressible boundary layers. Journal of Scientific Computing, 2(1):77–97, March 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061513;   http://link.springer.com/content/pdf/10.
1007/BF01061513;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=1&spage=77-97.

Boyd:1987:ECF

[16] John P. Boyd. Exponentially convergent Fourier–Chebshev quadrature schemes on bounded and infinite intervals. Journal of Scientific Computing, 2(2):99–109, June 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061480;   http://link.springer.com/content/pdf/10.
1007/BF01061480;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=2&spage=99-109.

Bayly:1987:IDF

[17] B. J. Bayly, I. Goldhirsch, and Steven A. Orszag. Independent degrees of freedom of dynamical systems. Journal of Scientific Computing, 2(2):111–121, June 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061481;   http://link.springer.com/content/pdf/10.
1007/BF01061481;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=2&spage=111-121.

Gustafsson:1987:UHS

[18] Bertil Gustafsson. Unsymmetric hyperbolic systems and the Euler equations at low Mach numbers. Journal of Scientific Computing, 2 (2):123–136, June 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061482;   http://link.springer.com/content/pdf/10.
1007/BF01061482;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=2&spage=123-136.

Thacker:1987:CFA

[19] W. C. Thacker. A cost-function approach to the assimilation of asynoptic data. Journal of Scientific Computing, 2(2):137–158, June 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061483;   http://link.springer.com/content/pdf/10.
1007/BF01061483;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=2&spage=137-158.

Ong:1987:AFR

[20] J. H. Ong. An algorithm for frontwidth reduction. Journal of Scientific Computing, 2(2):159–173, June 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061484;   http://link.springer.com/
content/pdf/10.1007/BF01061484;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=2&issue=2&spage=
159-173.

Fatoohi:1987:IAM

[21] Raad A. Fatoohi and Chester E. Grosch. Implementation of an ADI method on parallel computers. Journal of Scientific Computing, 2 (2):175–193, June 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061485;   http://link.springer.com/content/pdf/10.
1007/BF01061485;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=2&spage=175-193.

Pasquarelli:1987:SAS

[22] F. Pasquarelli, A. Quarteroni, and G. Sacchi-Landriani. Spectral approximations of the Stokes problem by divergence-free functions. Journal of Scientific Computing, 2(3):195–226, September 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061110;   http://link.springer.com/content/pdf/10.
1007/BF01061110;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=3&spage=195-226.

Domaradzki:1987:NSD

[23] J. Andrzej Domaradzki and Steven A. Orszag. Numerical solutions of the direct interaction approximation equations for anisotropic turbulence. Journal of Scientific Computing, 2(3):227–248, September 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061111;   http://link.springer.com/content/pdf/10.
1007/BF01061111;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=3&spage=227-248.

Kopriva:1987:PAS

[24] David A. Kopriva. A practical assessment of spectral accuracy for hyperbolic problems with discontinuities. Journal of Scientific Computing, 2(3):249–262, September 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061112;   http://link.springer.com/
content/pdf/10.1007/BF01061112;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=2&issue=3&spage=
249-262.

Bullister:1987:NST

[25] Edward T. Bullister and Steven A. Orszag. Numerical simulation of turbulent spots in channel and boundary layer flows. Journal of Scientific Computing, 2(3):263–281, September 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061113;   http://link.springer.com/
content/pdf/10.1007/BF01061113;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=2&issue=3&spage=
263-281.

Smith:1987:AFE

[26] Robert T. Smith and Csaba K. Zoltani. An application of the finite element method to maximum entropy tomographic image reconstruction. Journal of Scientific Computing, 2(3):283–295, September 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.

1007/BF01061114;   http://link.springer.com/content/pdf/10.
1007/BF01061114;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=3&spage=283-295.

Derby:1987:FEM

[27] J. J. Derby, L. J. Atherton, P. D. Thomas, and R. A. Brown. Finiteelement methods for analysis of the dynamics and control of Czochralski crystal growth. Journal of Scientific Computing, 2(4):297–343, December 1987. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061294;   http://link.springer.com/content/pdf/10.
1007/BF01061294;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=4&spage=297-343.

Reuven:1987:SAO

[28] Y. Reuven, M. D. Smooke, and H. Rabitz. Sensitivity analysis of onedimensional mixed initial-boundary value problems: Application to freely propagating premixed laminar flames. Journal of Scientific Computing, 2 (4):345–370, December 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061295;   http://link.springer.com/content/pdf/10.
1007/BF01061295;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=4&spage=345-370.

Hariharan:1987:WET

[29] S. I. Hariharan and S. I. Sudharsanan. Wave envelope technique for multimode wave guide problems. Journal of Scientific Computing, 2(4): 371–387, December 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061296;   http://link.springer.com/content/pdf/10.
1007/BF01061296;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=4&spage=371-387.

Ronquist:1987:SEM

[30] Einar M. Rønquist and Anthony T. Patera. Spectral element multigrid. I. Formulation and numerical results. Journal of Scientific Computing, 2 (4):389–406, December 1987. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061297;   http://link.springer.com/content/pdf/10.
1007/BF01061297;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=2&issue=4&spage=389-406.

Reyna:1988:ECC

[31] Luis G. Reyna. L_{2} estimates for Chebyshev collocation. Journal of Scientific Computing, 3(1):1–23, March 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01066480;   http://link.springer.com/
content/pdf/10.1007/BF01066480;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=3&issue=1&spage=
1-23.

Leonard:1988:DNS

[32] Andy D. Leonard and James C. Hill. Direct numerical simulation of turbulent flows with chemical reaction. Journal of Scientific Computing, 3(1):25–43, March 1988. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01066481;   http://link.springer.com/content/pdf/10.
1007/BF01066481;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=1&spage=25-43.

Quarteroni:1988:DDP

[33] Alfio Quarteroni and Giovanni Sacchi-Landriani. Domain decomposition preconditioners for the spectral collocation method. Journal of Scientific Computing, 3(1):45–76, March 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01066482;   http://link.springer.com/
content/pdf/10.1007/BF01066482;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=3&issue=1&spage=
45-76.

Ungar:1988:FEM

[34] L. H. Ungar, N. Ramprasad, and R. A. Brown. Finite element methods for unsteady solidification problems arising in prediction of morphological structure. Journal of Scientific Computing, 3(1):77– 108, March 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01066483;   http://link.springer.com/content/pdf/10.
1007/BF01066483;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=1&spage=77-108.

Boyd:1988:CDT

[35] John P. Boyd. Chebyshev domain truncation is inferior to Fourier domain truncation for solving problems on an infinite interval. Journal of Scientific Computing, 3(2):109–120, June 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061252;   http://link.springer.com/
content/pdf/10.1007/BF01061252;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=3&issue=2&spage=
109-120.

Liu:1988:SBF

[36] C. H. Liu, T. C. Wong, and O. A. Kandil. Separation-bubble flow solution using Euler/Navier–Stokes zonal approach with downstream compatibility conditions. Journal of Scientific Computing, 3(2):121– 137, June 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061253;   http://link.springer.com/content/pdf/10.
1007/BF01061253;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=2&spage=121-137.

Yakhot:1988:CTR

[37] Victor Yakhot, Steven A. Orszag, and Raj Panda. Computational test of the renormalization group theory of turbulence. Journal of Scientific Computing, 3(2):139–147, June 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061254;   http://link.springer.com/
content/pdf/10.1007/BF01061254;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=3&issue=2&spage=
139-147.

Buning:1988:SEG

[38] Pieter G. Buning. Sources of error in the graphical analysis of CFD results. Journal of Scientific Computing, 3(2):149–164, June 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061255;   http://link.springer.com/content/pdf/10.
1007/BF01061255;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=2&spage=149-164.

Greenberg:1988:SRF

[39] J. B. Greenberg. Self-regulating finite-difference methods for the computation of reacting flows with nonlinear kinetics. Journal of Scientific Computing, 3(2):165–187, June 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061256;   http://link.springer.com/
content/pdf/10.1007/BF01061256;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=3&issue=2&spage=
165-187.

Saetre:1988:DSG

[40] Tor O. Saetre and Nils Ryum. Dynamic simulation of grain boundary migration. Journal of Scientific Computing, 3(2):189– 199, June 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061257;   http://link.springer.com/content/pdf/10.
1007/BF01061257;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=2&spage=189-199.

Sidi:1988:QMP

[41] Avram Sidi and Moshe Israeli. Quadrature methods for periodic singular and weakly singular Fredholm integral equations. Journal of Scientific Computing, 3(2):201–231, June 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061258;   http://link.springer.com/
content/pdf/10.1007/BF01061258;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=3&issue=2&spage=
201-231.

Jameson:1988:ADC

[42] Antony Jameson. Aerodynamic design via control theory. Journal of Scientific Computing, 3(3):233–260, September 1988. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01061285;
http://link.springer.com/content/pdf/10.1007/BF01061285;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=3&issue=3&spage=233-260.

Taasan:1988:MMS

[43] Shlomo Ta’asan. Multigrid method for stability problems. Journal of Scientific Computing, 3(3):261–274, September 1988. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01061286;
http://link.springer.com/content/pdf/10.1007/BF01061286;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=3&issue=3&spage=261-274.

Abarbanel:1988:CHO

[44] Saul Abarbanel and Ajay Kumar. Compact high-order schemes for the Euler equations. Journal of Scientific Computing, 3(3):275–288, September 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.

1007/BF01061287;   http://link.springer.com/content/pdf/10.
1007/BF01061287;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=3&spage=275-288.

Chiang:1988:PMC

[45] Yi ling F. Chiang, Ji-Suing Ma, and Kuo-Lin Hu. Parallel multischeme computation. Journal of Scientific Computing, 3(3):289– 306, September 1988. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061288;   http://link.springer.com/content/pdf/10.
1007/BF01061288;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=3&spage=289-306.

Somuah:1988:CMT

[46] Clement B. Somuah and Syed M. Islam. A class of multiple time scale algorithms for simulating power system frequency dynamics. Journal of Scientific Computing, 3(3):307–322, September 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061289;   http://link.springer.com/content/pdf/10.
1007/BF01061289;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=3&spage=307-322.

Maday:1988:SEM

[47] Yvon Maday and Rafael Munoz. Spectral element multigrid. II. Theoretical justification. Journal of Scientific Computing, 3(4):323– 353, December 1988. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01065177;   http://link.springer.com/content/pdf/10.
1007/BF01065177;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=4&spage=323-353.

Christodoulou:1988:FLM

[48] K. N. Christodoulou and L. E. Scriven. Finding leading modes of a viscous free surface flow: An asymmetric generalized eigenproblem. Journal of Scientific Computing, 3(4):355–406, December 1988. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01065178;   http://link.springer.com/content/pdf/10.
1007/BF01065178;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=4&spage=355-406.

She:1988:SDI

[49] Zhen-Su She, Eric Jackson, and Steven A. Orszag. Scale-dependent intermittency and coherence in turbulence. Journal of Scientific Computing, 3 (4):407–434, December 1988. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01065179;   http://link.springer.com/content/pdf/10.
1007/BF01065179;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=3&issue=4&spage=407-434.

Schultz:1989:CPM

[50] W. W. Schultz, N. Y. Lee, and J. P. Boyd. Chebyshev pseudospectral method of viscous flows with corner singularities. Journal of Scientific Computing, 4(1):1–24, March 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061264;   http://link.springer.com/
content/pdf/10.1007/BF01061264;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=4&issue=1&spage=
1-24.

Tal-Ezer:1989:PAF

[51] Hillel Tal-Ezer. Polynomial approximation of functions of matrices and applications. Journal of Scientific Computing, 4(1):25–60, March 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061265;   http://link.springer.com/content/pdf/10.
1007/BF01061265;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=1&spage=25-60.

Succi:1989:FCP

[52] Sauro Succi and Maurizio Benassi. A four-color parallel algorithm for the solution of a two-dimensional advection-diffusion equation with the finite element method. Journal of Scientific Computing, 4(1): 61–70, March 1989. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061266;   http://link.springer.com/content/pdf/10.
1007/BF01061266;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=1&spage=61-70.

Gruberger:1989:SMC

[53] Nira Gruberger. Spectral methods for the computation of discontinuous solutions. Journal of Scientific Computing, 4(1):71–117, March 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061267;   http://link.springer.com/content/pdf/10.
1007/BF01061267;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=1&spage=71-117.

Bisshopp:1989:CLT

[54] Frederic Bisshopp. Characteristic lines for two- and three-dimensional flows. Journal of Scientific Computing, 4(1):119–137, March 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01061268;
http://link.springer.com/content/pdf/10.1007/BF01061268;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=4&issue=1&spage=119-137.

Yakhot:1989:RGF

[55] A. Yakhot, S. A. Orszag, V. Yakhot, and M. Israeli. Renormalization group formulation of large-eddy simulations. Journal of Scientific Computing, 4(2):139–158, June 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061499;   http://link.springer.com/
content/pdf/10.1007/BF01061499;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=4&issue=2&spage=
139-158.

Olsson:1989:BMD

[56] Pelle Olsson and S. Lennart Johnson. Boundary modifications of the dissipation operators for the three-dimensional Euler equations. Journal of Scientific Computing, 4(2):159–195, June 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061500;   http://link.springer.com/
content/pdf/10.1007/BF01061500;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=4&issue=2&spage=
159-195.

Zang:1989:RRN

[57] Thomas A. Zang, Steven E. Krist, and M. Yousuff Hussaini. Resolution requirements for numerical simulations of transition. Journal of Scientific Computing, 4(2):197–217, June 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061501;   http://link.springer.com/
content/pdf/10.1007/BF01061501;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=4&issue=2&spage=
197-217.

Balachandar:1989:NSH

[58] S. Balachandar, M. R. Maxey, and L. Sirovich. Numerical simulation of high Rayleigh number convection. Journal of Scientific Computing, 4(2):219–236, June 1989. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061502;   http://link.springer.com/content/pdf/10.
1007/BF01061502;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=2&spage=219-236.

Baker:1989:GVM

[59] Gregory R. Baker, Daniel I. Meiron, and Steven A. Orszag. Generalized vortex methods for free surface flow problems. II: Radiating waves. Journal of Scientific Computing, 4(3):237–259, September 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061057;   http://link.springer.com/content/pdf/10.
1007/BF01061057;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=3&spage=237-259.

Rose:1989:CFV

[60] Milton E. Rose. Compact finite volume methods for the diffusion equation. Journal of Scientific Computing, 4(3):261–290, September 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061058;   http://link.springer.com/content/pdf/10.
1007/BF01061058;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=3&spage=261-290.

Tomboulides:1989:ERB

[61] A. G. Tomboulides, M. Israeli, and G. E. Karniadakis. Efficient removal of boundary-divergence errors in time-splitting methods. Journal of Scientific Computing, 4(3):291–308, September 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01061059;   http://link.springer.com/content/pdf/10.1007/
BF01061059;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=4&issue=3&spage=291-308.

Gottlieb:1989:PPD

[62] David Gottlieb and Richard S. Hirsh. Parallel pseudospectral domain decomposition techniques. Journal of Scientific Computing, 4(4):309– 325, December 1989. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060991;   http://link.springer.com/content/pdf/10.
1007/BF01060991;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=4&spage=309-325.

Friesner:1989:MEP

[63] Richard A. Friesner and Laurette S. Tuckerman. A method for exponential propagation of large systems of stiff nonlinear differential equations. Journal of Scientific Computing, 4(4):327–354, December 1989. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060992;   http://link.springer.com/content/pdf/10.
1007/BF01060992;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=4&spage=327-354.

Goldenfeld:1989:IAR

[64] Nigel Goldenfeld, Olivier Martin, and Y. Oono. Intermediate asymptotics and renormalization group theory. Journal of Scientific Computing, 4(4): 355–372, December 1989. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060993;   http://link.springer.com/content/pdf/10.
1007/BF01060993;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=4&issue=4&spage=355-372.

Kida:1990:EBD

[65] Shigeo Kida and Steven A. Orszag. Enstrophy budget in decaying compressible turbulence. Journal of Scientific Computing, 5(1): 1–34, March 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01063424;   http://link.springer.com/content/pdf/10.
1007/BF01063424;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=1&spage=1-34.

Chiang:1990:PMC

[66] Yi ling F. Chiang and Jih-Fu Lai. Parallel multischeme computation with the switch of computation history. Journal of Scientific Computing, 5(1):35–53, March 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01063425;   http://link.springer.com/content/pdf/10.
1007/BF01063425;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=1&spage=35-53.

Ehrenstein:1990:SCF

[67] U. Ehrenstein. The spectrum of a Chebyshev–Fourier approximation for the Stokes equations. Journal of Scientific Computing, 5 (1):55–84, March 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01063426;   http://link.springer.com/content/pdf/10.
1007/BF01063426;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=1&spage=55-84.

Kida:1990:ESD

[68] Shigeo Kida and Steven A. Orszag. Energy and spectral dynamics in forced compressible turbulence. Journal of Scientific Computing, 5 (2):85–125, June 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01065580;   http://link.springer.com/content/pdf/10.
1007/BF01065580;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=2&spage=85-125.

Shu:1990:NEA

[69] Chi-Wang Shu. Numerical experiments on the accuracy of ENO and modified ENO schemes. Journal of Scientific Computing, 5(2): 127–149, June 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01065581;   http://link.springer.com/content/pdf/10.
1007/BF01065581;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=2&spage=127-149.

Rogerson:1990:NSC

[70] A. M. Rogerson and E. Meiburg. A numerical study of the convergence properties of ENO schemes. Journal of Scientific Computing, 5(2):151–167, June 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01065582;   http://link.springer.com/content/pdf/10.
1007/BF01065582;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=2&spage=151-167.

Rose:1990:IES

[71] Milton E. Rose. An implicit enthalpy scheme for one-phase Stefan problems. Journal of Scientific Computing, 5(2):169–185, June 1990. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01065583;   http://link.springer.com/content/pdf/10.1007/

BF01065583;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=5&issue=2&spage=169-185.

Parker:1990:SFE

[72] J. R. Parker. A system for fast erosion and dilation of bi-level images. Journal of Scientific Computing, 5(3):187–198, September 1990. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01089163;   http://link.springer.com/content/pdf/10.
1007/BF01089163;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=3&spage=187-198.

Yakhot:1990:PTP

[73] Victor Yakhot, Steven A. Orszag, and S. Balachandar. Phenomenological theory of probability distributions in turbulence. Journal of Scientific Computing, 5(3):199–221, September 1990. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01089164;   http://link.springer.com/
content/pdf/10.1007/BF01089164;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=5&issue=3&spage=
199-221.

Gottlieb:1990:QIC

[74] D. Gottlieb and C. L. Streett. Quadrature imposition of compatibility conditions in Chebyshev methods. Journal of Scientific Computing, 5(3): 223–239, September 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01089165;   http://link.springer.com/content/pdf/10.
1007/BF01089165;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=3&spage=223-239.

Orszag:1990:IMS

[75] Steven A. Orszag and Victor Yakhot. Inhibition of mixing and selffocusing in high Mach number turbulence. Journal of Scientific Computing, 5(3):241–243, September 1990. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01089166;   http://link.springer.com/
content/pdf/10.1007/BF01089166;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=5&issue=3&spage=
241-243.

ONeal:1990:OFE

[76] David C. O’Neal. Optimization of finite element codes. Journal of Scientific Computing, 5(3):245–262, September 1990. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01089167;   http://link.springer.com/
content/pdf/10.1007/BF01089167;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=5&issue=3&spage=
245-262.

Maday:1990:OIF

[77] Y. Maday, Anthony T. Patera, and Einar M. Rønquist. An operatorintegration-factor splitting method for time-dependent problems: Application to incompressible fluid flow. Journal of Scientific Computing, 5 (4):263–292, December 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01063118;   http://link.springer.com/content/pdf/10.
1007/BF01063118;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=4&spage=263-292.

Dubey:1990:CFT

[78] A. Dubey, M. Zubair, and C. E. Grosch. Computing the Fourier transform of real data on a hypercube. Journal of Scientific Computing, 5(4): 293–309, December 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01063119;   http://link.springer.com/content/pdf/10.
1007/BF01063119;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=4&spage=293-309.

Boyd:1990:EET

[79] John P. Boyd. The envelope of the error for trigonometric and Chebyshev interpolation. Journal of Scientific Computing, 5(4):311– 363, December 1990. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01063120;   http://link.springer.com/content/pdf/10.
1007/BF01063120;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=5&issue=4&spage=311-363.

Heinrichs:1991:STS

[80] Wilhelm Heinrichs. Stabilization techniques for spectral methods. Journal of Scientific Computing, 6(1):1–19, March 1991. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01068121;
http://link.springer.com/content/pdf/10.1007/BF01068121;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=6&issue=1&spage=1-19.

Karageorghis:1991:NSB

[81] Andreas Karageorghis. A note on the satisfaction of the boundary conditions for Chebyshev collocation methods in rectangular domains. Journal of Scientific Computing, 6(1):21–26, March 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01068122;
http://link.springer.com/content/pdf/10.1007/BF01068122;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=6&issue=1&spage=21-26.

Jackson:1991:CSP

[82] Eric Jackson, Zhen-Su She, and Steven A. Orszag. A case study in parallel computing: I. Homogeneous turbulence on a hypercube. Journal of Scientific Computing, 6(1):27–45, March 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01068123;   http://link.springer.com/
content/pdf/10.1007/BF01068123;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=6&issue=1&spage=
27-45.

Zhang:1991:SRD

[83] Peisen Zhang. Simpson’s rule of discretized Feynman path integration. Journal of Scientific Computing, 6(1):47–60, March 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01068124;   http://link.springer.com/content/pdf/10.1007/
BF01068124;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=6&issue=1&spage=47-60.

Halton:1991:RSF

[84] John H. Halton. Random sequences in Fréchet spaces. Journal of Scientific Computing, 6(1):61–77, March 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01068125;   http://link.springer.com/
content/pdf/10.1007/BF01068125;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=6&issue=1&spage=
61-77.

Anonymous:1991:ES

[85] Anonymous. Editorial statement. Journal of Scientific Computing, 6(1):I–II, March 1991. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01068120;   http://link.springer.com/content/pdf/10.

1007/BF01068120;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=1&spage=I-II.

Henderson:1991:HSE

[86] Ron Henderson and George Em. Karniadakis. Hybrid spectral-element-low-order methods for incompressible flows. Journal of Scientific Computing, 6(2):79–100, June 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01062115;   http://link.springer.com/
content/pdf/10.1007/BF01062115;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=6&issue=2&spage=
79-100.

Johansson:1991:WPGa

[87] B. Christer V. Johansson. Well-posedness in the generalized sense for the incompressible Navier–Stokes equation. Journal of Scientific Computing, 6(2):101–127, June 1991. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01062116;   http://link.springer.com/content/pdf/10.
1007/BF01062116;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=2&spage=101-127.

Parter:1991:PSO

[88] Seymour V. Parter and Sze-Ping Wong. Preconditioning secondorder elliptic operators: Condition numbers and the distribution of the singular values. Journal of Scientific Computing, 6(2):129– 157, June 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01062117;   http://link.springer.com/content/pdf/10.
1007/BF01062117;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=2&spage=129-157.

Vandeven:1991:FSF

[89] Hervé Vandeven. Family of spectral filters for discontinuous problems. Journal of Scientific Computing, 6(2):159–192, June 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01062118;
http://link.springer.com/content/pdf/10.1007/BF01062118;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=6&issue=2&spage=159-192.

Cook:1991:HSC

[90] Grant O. Cook, Jr., Jeffrey F. Painter, and Stewart A. Brown. How symbolic computation boosts productivity in the simulation of partial differential equations. Journal of Scientific Computing, 6(2): 193–209, June 1991. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01062119;   http://link.springer.com/content/pdf/10.
1007/BF01062119;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=2&spage=193-209.

Cheremensky:1991:RFD

[91] Alexander G. Cheremensky. Robust frequency design of linear stationary systems in aeroautoelastics. Journal of Scientific Computing, 6(2):211–227, June 1991. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01062120;   http://link.springer.com/content/pdf/10.
1007/BF01062120;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=2&spage=211-227.

Barouch:1991:NSS

[92] Eytan Barouch, John W. Cahn, and Uwe Hollerbach. Numerical simulation of submicron photolithographic processing. Journal of Scientific Computing, 6(3):229–250, September 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01062811;   http://link.springer.com/
content/pdf/10.1007/BF01062811;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=6&issue=3&spage=
229-250.

Groh:1991:AME

[93] D. J. Groh, R. A. Marshall, A. B. Kunz, and C. R. Givens. An approximation method for eigenvectors of very large matrices. Journal of Scientific Computing, 6(3):251–267, September 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01062812;   http://link.springer.com/content/pdf/10.1007/
BF01062812;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=6&issue=3&spage=251-267.

Cooke:1991:CFT

[94] Charlie H. Cooke and Tze-Jang Chen. Continuous front tracking with subcell resolution. Journal of Scientific Computing, 6(3):269– 282, September 1991. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01062813;   http://link.springer.com/content/pdf/10.
1007/BF01062813;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=3&spage=269-282.

Chiang:1991:EAS

[95] Yi-Ling F. Chiang. Error approximation in the solution of a linear PDE initial-valued problem. Journal of Scientific Computing, 6(3): 283–303, September 1991. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01062814;   http://link.springer.com/content/pdf/10.
1007/BF01062814;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=3&spage=283-303.

Guerinoni:1991:AMR

[96] Fabio Guerinoni. An appropriate mesh representation for tetrahedral finite volume computations. Journal of Scientific Computing, 6(3): 305–321, September 1991. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01062815;   http://link.springer.com/content/pdf/10.
1007/BF01062815;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=3&spage=305-321.

Book:1991:QRN

[97] David L. Book, Chiping Li, and Gopal Patnaik. Quantifying residual numerical diffusion in flux-corrected transport algorithms. Journal of Scientific Computing, 6(3):323–343, September 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01062816;   http://link.springer.com/
content/pdf/10.1007/BF01062816;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=6&issue=3&spage=
323-343.

Dubiner:1991:SMT

[98] Moshe Dubiner. Spectral methods on triangles and other domains. Journal of Scientific Computing, 6(4):345–390, December 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060030;   http://link.springer.com/content/pdf/10.
1007/BF01060030;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=4&spage=345-390.

Johansson:1991:WPGb

[99] B. Christer V. Johansson. Well-posedness in the generalized sense for boundary layer suppressing boundary conditions. Journal of Scientific Computing, 6(4):391–414, December 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01060031;   http://link.springer.com/
content/pdf/10.1007/BF01060031;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=6&issue=4&spage=
391-414.

Halton:1991:RSG

[100] John H. Halton. Random sequences in generalized Cantor sets. Journal of Scientific Computing, 6(4):415–423, December 1991. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01060032;
http://link.springer.com/content/pdf/10.1007/BF01060032;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=6&issue=4&spage=415-423.

Gama:1991:TDN

[101] S. Gama, U. Frisch, and H. Scholl. The two-dimensional Navier– Stokes equations with a large-scale instability of the Kuramoto– Sivashinsky type: Numerical exploration on the connection machine. Journal of Scientific Computing, 6(4):425–452, December 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060033;   http://link.springer.com/content/pdf/10.
1007/BF01060033;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=4&spage=425-452.

Carcione:1991:DDW

[102] José M. Carcione. Domain decomposition for wave propagation problems. Journal of Scientific Computing, 6(4):453–472, December 1991. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060034;   http://link.springer.com/content/pdf/10.
1007/BF01060034;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=6&issue=4&spage=453-472.

Kida:1992:ESD

[103] Shigeo Kida and Steven A. Orszag. Energy and spectral dynamics in decaying compressible turbulence. Journal of Scientific Computing, 7 (1):1–34, March 1992. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060209;   http://link.springer.com/content/pdf/10.
1007/BF01060209;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=1&spage=1-34.

Yakhot:1992:RGE

[104] Victor Yakhot and Leslie M. Smith. The renormalization group, the ϵ-expansion and derivation of turbulence models. Journal of Scientific Computing, 7(1):35–61, March 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01060210;   http://link.springer.com/
content/pdf/10.1007/BF01060210;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=7&issue=1&spage=
35-61.

Jeng:1992:NAG

[105] Yih Nen Jeng and Yuan Chang Liou. A new adaptive grid generation by elliptic equations with orthogonality at all of the boundaries. Journal of Scientific Computing, 7(1):63–80, March 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01060211;   http://link.springer.com/content/pdf/10.1007/
BF01060211;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=7&issue=1&spage=63-80.

Sanugi:1992:IMF

[106] Bahrom B. Sanugi. An iterative multistep formula for solving initial value problems. Journal of Scientific Computing, 7(1):81–94, March 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060212;   http://link.springer.com/content/pdf/10.
1007/BF01060212;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=1&spage=81-94.

Heinrichs:1992:SMA

[107] Wilhelm Heinrichs. A stabilized multidomain approach for singular perturbation problems. Journal of Scientific Computing, 7(2): 95–125, June 1992. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01059944;   http://link.springer.com/content/pdf/10.
1007/BF01059944;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=2&spage=95-125.

Bhogeswara:1992:DDM

[108] R. Bhogeswara and J. E. Killough. Domain decomposition and multigrid solvers for flow simulation in porous media on distributed memory parallel processors. Journal of Scientific Computing, 7(2):127– 162, June 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01059945;   http://link.springer.com/content/pdf/10.
1007/BF01059945;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=2&spage=127-162.

Babovsky:1992:ETC

[109] Hans Babovsky. Efficient treatment of collisions in simulation codes for heavy-ion reactions. Journal of Scientific Computing, 7(2):163– 174, June 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01059946;   http://link.springer.com/content/pdf/10.
1007/BF01059946;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=2&spage=163-174.

Dattoli:1992:ANR

[110] G. Dattoli, C. Mari, A. Torre, and C. Chiccoli. Analytical and numerical results on M-variable generalized Bessel functions. Journal of Scientific Computing, 7(2):175–196, June 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01059947;   http://link.springer.com/
content/pdf/10.1007/BF01059947;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=7&issue=2&spage=
175-196.

Berger:1992:HGM

[111] M. F. Berger. Hierarchical gradient methods for nonlinear LSQ problems. Journal of Scientific Computing, 7(3):197–228, September 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061328;   http://link.springer.com/content/pdf/10.
1007/BF01061328;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=3&spage=197-228.

Yakhot:1992:AQT

[112] Alexander Yakhot, Omer Kedar, and Steven A. Orszag. An algebraic-Q4 turbulent eddy viscosity model: Boundary layer flow over a flat plate and flow in a pipe. Journal of Scientific Computing, 7(3):229– 239, September 1992. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061329;   http://link.springer.com/content/pdf/10.
1007/BF01061329;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=3&spage=229-239.

Whang:1992:NAC

[113] M. H. Whang, S. P. Kuo, and M. C. Lee. Numerical analysis of coupled hydromagnetic wave equations with a finite difference scheme. Journal of Scientific Computing, 7(3):241–261, September 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061330;   http://link.springer.com/content/pdf/10.
1007/BF01061330;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=3&spage=241-261.

Gupta:1992:MAP

[114] S. N. Gupta, M. Zubair, and C. E. Grosch. A multigrid algorithm for parallel computers: CPMG. Journal of Scientific Computing, 7(3): 263–279, September 1992. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061331;   http://link.springer.com/content/pdf/10.
1007/BF01061331;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=3&spage=263-279.

Halton:1992:RRT

[115] John H. Halton. Reject the rejection technique. Journal of Scientific Computing, 7(3):281–287, September 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061332;   http://link.springer.com/
content/pdf/10.1007/BF01061332;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=7&issue=3&spage=
281-287.

Yang:1992:NCO

[116] Hong Yang and J. Michael Orszag. Numerical computation of H^{∞} optimal performance. Journal of Scientific Computing, 7(4):289–311, December 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01108034;   http://link.springer.com/content/pdf/10.
1007/BF01108034;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=4&spage=289-311.

Black:1992:PCU

[117] Kelly Black. Polynomial collocation using a domain decomposition solution to parabolic PDE’s via the penalty method and explicit/ implicit time marching. Journal of Scientific Computing, 7(4):313– 338, December 1992. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01108035;   http://link.springer.com/content/pdf/10.
1007/BF01108035;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=4&spage=313-338.

Jeng:1992:ALI

[118] Yih Nen Jeng and Uon Jien Payne. The application of the locally implicit method to upwind TVD schemes. Journal of Scientific Computing, 7(4): 339–357, December 1992. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01108036;   http://link.springer.com/content/pdf/10.
1007/BF01108036;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=4&spage=339-357.

Iqbal:1992:OPA

[119] Rafhat Iqbal. A one-pass algorithm for shape-preserving quadratic spline interpolation. Journal of Scientific Computing, 7(4):359–376, December 1992. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01108037;   http://link.springer.com/content/pdf/10.
1007/BF01108037;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=7&issue=4&spage=359-376.

Goldhirsch:1993:MDS

[120] I. Goldhirsch, M.-L. Tan, and G. Zanetti. A molecular dynamical study of granular fluids I: The unforced granular gas in two dimensions. Journal of Scientific Computing, 8(1):1–40, March 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060830;   http://link.springer.com/content/pdf/10.
1007/BF01060830;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=1&spage=1-40.

Zheligovsky:1993:NSK

[121] V. A. Zheligovsky. Numerical solution of the kinematic dynamo problem for Beltrami flows in a sphere. Journal of Scientific Computing, 8(1):41–68, March 1993. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060831;   http://link.springer.com/content/pdf/10.
1007/BF01060831;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=1&spage=41-68.

Dattoli:1993:ATG

[122] G. Dattoli, C. Chiccoli, S. Lorenzutta, and G. Maino. Advances on the theory of generalized Bessel functions and applications to multiphoton processes. Journal of Scientific Computing, 8(1):69–109, March 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060832;   http://link.springer.com/content/pdf/10.
1007/BF01060832;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=1&spage=69-109.

Yakhot:1993:NST

[123] Alexander Yakhot and Steven A. Orszag. Numerical simulation of turbulent flow in the inlet region of a smooth pipe. Journal of Scientific Computing, 8(2):111–121, June 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01060867;   http://link.springer.com/
content/pdf/10.1007/BF01060867;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=8&issue=2&spage=
111-121.

Karageorghis:1993:CCS

[124] Andreas Karageorghis. Conforming Chebyshev spectral methods for Poisson problems in rectangular domains. Journal of Scientific Computing, 8(2):123–133, June 1993. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060868;   http://link.springer.com/content/pdf/10.
1007/BF01060868;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=2&spage=123-133.

Israeli:1993:SMT

[125] M. Israeli, L. Vozovoi, and A. Averbuch. Spectral multidomain technique with local Fourier basis. Journal of Scientific Computing, 8 (2):135–149, June 1993. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060869;   http://link.springer.com/content/pdf/10.
1007/BF01060869;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=2&spage=135-149.

Israeli:1993:PIA

[126] M. Israeli, L. Vozovoi, and A. Averbuch. Parallelizing implicit algorithms for time-dependent problems by parabolic domain decomposition. Journal of Scientific Computing, 8(2):151–166, June 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01060870;   http://link.springer.com/content/pdf/10.1007/
BF01060870;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=8&issue=2&spage=151-166.

Dubois:1993:SIN

[127] T. Dubois, F. Jauberteau, and R. Temam. Solution of the incompressible Navier–Stokes equations by the nonlinear Galerkin method. Journal of Scientific Computing, 8(2):167–194, June 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01060871;   http://link.springer.com/
content/pdf/10.1007/BF01060871;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=8&issue=2&spage=
167-194.

Ioffe:1993:AAP

[128] Ludmila Ioffe and Shlomit S. Pinter. Applying asynchronous parallel characteristic methods for solving systems of hyperbolic PDEs. Journal of Scientific Computing, 8(3):195–218, September 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060930;   http://link.springer.com/content/pdf/10.
1007/BF01060930;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=3&spage=195-218.

Succi:1993:SSD

[129] S. Succi, D. d’Humières, Y. H. Qian, and S. A. Orszag. On the small-scale dynamical behavior of lattice BGK and lattice Boltzmann schemes. Journal of Scientific Computing, 8(3):219–230, September 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060931;   http://link.springer.com/content/pdf/10.
1007/BF01060931;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=3&spage=219-230.

Qian:1993:STL

[130] Y. H. Qian. Simulating thermohydrodynamics with lattice BGK models. Journal of Scientific Computing, 8(3):231–242, September 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060932;   http://link.springer.com/content/pdf/10.
1007/BF01060932;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=3&spage=231-242.

Parker:1993:SAD

[131] J. R. Parker and G. Groisman. Simulated annealing for data modeling in scientific images. Journal of Scientific Computing, 8(3):243– 266, September 1993. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01060933;   http://link.springer.com/content/pdf/10.
1007/BF01060933;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=3&spage=243-266.

Jameson:1993:WBD

[132] Leland Jameson. On the wavelet based differentiation matrix. Journal of Scientific Computing, 8(3):267–305, September 1993. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01060934;
http://link.springer.com/content/pdf/10.1007/BF01060934;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=8&issue=3&spage=267-305.

Christer:1993:URE

[133] B. Christer and V. Johansson. The use of Richardson extrapolation for the numerical solution of low Mach number flow in confined regions. Journal of Scientific Computing, 8(4):307–340, December 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061142;   http://link.springer.com/content/pdf/10.
1007/BF01061142;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=4&spage=307-340.

Liu:1993:MAM

[134] C. Liu, Z. Liu, and S. McCormick. Multilevel adaptive methods for laminar diffusion flames. Journal of Scientific Computing, 8(4):341– 355, December 1993. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061143;   http://link.springer.com/content/pdf/10.
1007/BF01061143;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=4&spage=341-355.

Nguyen:1993:NSA

[135] Hoa D. Nguyen and Seungho Paik. A noniterative solution approach for parallel pseudospectral domain decomposition. Journal of Scientific Computing, 8(4):357–372, December 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01061144;   http://link.springer.com/
content/pdf/10.1007/BF01061144;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=8&issue=4&spage=
357-372.

Mehrabi:1993:IND

[136] M. Reza Mehrabi and Robert A. Brown. An incomplete nested dissection algorithm for parallel direct solution of finite element discretizations of partial differential equations. Journal of Scientific Computing, 8(4): 373–387, December 1993. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061145;   http://link.springer.com/content/pdf/10.
1007/BF01061145;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=4&spage=373-387.

Heinrichs:1993:DRS

[137] Wilhelm Heinrichs. Distributive relaxations for the spectral Stokes operator. Journal of Scientific Computing, 8(4):389–398, December 1993. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01061146;   http://link.springer.com/content/pdf/10.
1007/BF01061146;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=8&issue=4&spage=389-398.

Boffi:1994:AAA

[138] D. Boffi and D. Funaro. An alternative approach to the analysis and the approximation of the Navier–Stokes equations. Journal of Scientific Computing, 9(1):1–16, March 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01573175;   http://link.springer.com/
content/pdf/10.1007/BF01573175;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=1&spage=
1-16.

Cheremensky:1994:PFA

[139] Alexander G. Cheremensky. Principle of feedback with additive system disturbances and some compensator schemes. Journal of Scientific Computing, 9(1):17–37, March 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01573176;   http://link.springer.com/
content/pdf/10.1007/BF01573176;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=1&spage=
17-37.

Lie:1994:MSA

[140] Ivar Lie. Multidomain solution of advection problems by Chebyshev spectral collocation. Journal of Scientific Computing, 9(1):39– 64, March 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01573177;   http://link.springer.com/content/pdf/10.
1007/BF01573177;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=1&spage=39-64.

Lin:1994:CMA

[141] Shun-Shii Lin. A chained-matrices approach for parallel computation of continued fractions and its applications. Journal of Scientific Computing, 9(1):65–80, March 1994. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01573178;   http://link.springer.com/content/pdf/10.
1007/BF01573178;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=1&spage=65-80.

Boyd:1994:HSL

[142] John P. Boyd. Hyperviscous shock layers and diffusion zones: Monotonicity, spectral viscosity, and pseudospectral methods for very high order differential equations. Journal of Scientific Computing, 9(1): 81–106, March 1994. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01573179;   http://link.springer.com/content/pdf/10.
1007/BF01573179;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=1&spage=81-106.

Couzy:1994:SEP

[143] W. Couzy and M. O. Deville. Spectral-element preconditioners for the Uzawa pressure operator applied to incompressible flows. Journal of Scientific Computing, 9(2):107–122, June 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01578382;   http://link.springer.com/
content/pdf/10.1007/BF01578382;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=2&spage=
107-122.

Luo:1994:PTM

[144] Yong Luo and Matthew J. Yedlin. Polynomial time-marching for nonperiodic boundary value problems. Journal of Scientific Computing, 9(2):123–136, June 1994. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01578383;   http://link.springer.com/content/pdf/10.
1007/BF01578383;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=2&spage=123-136.

Heinrichs:1994:SVC

[145] Wilhelm Heinrichs. Spectral viscosity for convection dominated flow. Journal of Scientific Computing, 9(2):137–148, June 1994. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF01578384;
http://link.springer.com/content/pdf/10.1007/BF01578384;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=9&issue=2&spage=137-148.

Li:1994:ASI

[146] Zi-Cai Li. Advanced splitting-integrating methods with high convergence rates for restoring images and patterns. Journal of Scientific Computing, 9(2):149–172, June 1994. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01578385;   http://link.springer.com/content/pdf/10.
1007/BF01578385;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=2&spage=149-172.

Schaeben:1994:ADT

[147] H. Schaeben. Analogy and duality of texture analysis by harmonics or indicators. Journal of Scientific Computing, 9(2):173– 195, June 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01578386;   http://link.springer.com/content/pdf/10.
1007/BF01578386;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=2&spage=173-195.

Iqbal:1994:ACP

[148] Rafhat Iqbal. An algorithm for convexity-preserving surface interpolation. Journal of Scientific Computing, 9(2):197–212, June 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL http://link.springer.com/article/10.1007/
BF01578387;   http://link.springer.com/content/pdf/10.1007/
BF01578387;   http://www.springerlink.com/openurl.asp?genre=
article&issn=0885-7474&volume=9&issue=2&spage=197-212.

Halton:1994:SMC

[149] John H. Halton. Sequential Monte Carlo techniques for the solution of linear systems. Journal of Scientific Computing, 9(2):213– 257, June 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01578388;   http://link.springer.com/content/pdf/10.
1007/BF01578388;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=2&spage=213-257.

Noullez:1994:FLT

[150] A. Noullez and M. Vergassola. A fast Legendre transform algorithm and applications to the adhesion model. Journal of Scientific Computing, 9(3):259–281, September 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575032;   http://link.springer.com/
content/pdf/10.1007/BF01575032;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=3&spage=
259-281.

Yakhot:1994:LRN

[151] A. Yakhot, S. Rakib, and W. S. Flannery. Low-Reynolds number approximation for turbulent eddy viscosity. Journal of Scientific Computing, 9(3):283–292, September 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575033;   http://link.springer.com/
content/pdf/10.1007/BF01575033;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=3&spage=
283-292.

Falsaperla:1994:PEC

[152] P. Falsaperla, S. Motta, and S. Succi. Parallel efficiency of the C.R.F. method on an IBM RS/6000 cluster platform. Journal of Scientific Computing, 9(3):293–309, September 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575034;   http://link.springer.com/
content/pdf/10.1007/BF01575034;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=3&spage=
293-309.

Vozovoi:1994:SMT

[153] L. Vozovoi, M. Israeli, and A. Averbuch. Spectral multidomain technique with local Fourier basis II: Decomposition into cells. Journal of Scientific Computing, 9(3):311–326, September 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575035;   http://link.springer.com/
content/pdf/10.1007/BF01575035;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=3&spage=
311-326.

Cooke:1994:MRA

[154] Charlie H. Cooke and Sang Kyu Yang. Multi-resolution analysis with frontal decomposition. Journal of Scientific Computing, 9(3):327– 340, September 1994. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01575036;   http://link.springer.com/content/pdf/10.
1007/BF01575036;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=3&spage=327-340.

Karageorghis:1994:CSM

[155] Andreas Karageorghis. Conforming spectral methods for Poisson problems in cuboidal domains. Journal of Scientific Computing, 9(3):341– 350, September 1994. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01575037;   http://link.springer.com/content/pdf/10.
1007/BF01575037;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=3&spage=341-350.

Nguyen:1994:SDD

[156] Hoa D. Nguyen and Seungho Paik. Solution domain decomposition method for viscous incompressible flow. Journal of Scientific Computing, 9(3):351–368, September 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575038;   http://link.springer.com/
content/pdf/10.1007/BF01575038;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=3&spage=
351-368.

Speziale:1994:RRS

[157] Charles G. Speziale, Ridha Abid, and Paul A. Durbin. On the realizability of Reynolds stress turbulence closures. Journal of Scientific Computing, 9(4):369–403, December 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575099;   http://link.springer.com/
content/pdf/10.1007/BF01575099;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=4&spage=
369-403.

Marion:1994:SCM

[158] Martine Marion. On the stability of collocation methods for the twodimensional Burgers equation- the Fourier case. Journal of Scientific Computing, 9(4):405–417, December 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575100;   http://link.springer.com/
content/pdf/10.1007/BF01575100;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=4&spage=
405-417.

Zampieri:1994:CNS

[159] E. Zampieri. On the condition number of some spectral collocation operators and their finite element preconditioning. Journal of Scientific Computing, 9(4):419–443, December 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF01575101;   http://link.springer.com/
content/pdf/10.1007/BF01575101;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=9&issue=4&spage=
419-443.

Olsson:1994:NBH

[160] Pelle Olsson. The numerical behavior of high-order finite difference methods. Journal of Scientific Computing, 9(4):445–466, December 1994. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF01575102;   http://link.springer.com/content/pdf/10.
1007/BF01575102;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=9&issue=4&spage=445-466.

Sei:1995:DAN

[161] Alain Sei and William Symes. Dispersion analysis of numerical wave propagation and its computational consequences. Journal of Scientific Computing, 10(1):1–27, March 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02087959;   http://link.springer.com/
content/pdf/10.1007/BF02087959;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=1&
spage=1-27.

Solomonoff:1995:RDF

[162] Alex Solomonoff. Reconstruction of a discontinuous function from a few Fourier coefficients using Bayesian estimation. Journal of Scientific Computing, 10(1):29–80, March 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02087960;   http://link.springer.com/
content/pdf/10.1007/BF02087960;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=1&
spage=29-80.

Croisille:1995:NSM

[163] Jean-Pierre Croisille, Rabia Khanfir, and Gérard Chanteur. Numerical simulation of the MHD equations by a kinetic-type method. Journal of Scientific Computing, 10(1):81–92, March 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02087961;   http://link.springer.com/
content/pdf/10.1007/BF02087961;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=1&
spage=81-92.

Mehrabi:1995:PIF

[164] M. Reza Mehrabi and Robert A. Brown. Parallel implementation of finite-element/Newton method for solution of steady-state and transient nonlinear partial differential equations. Journal of Scientific Computing, 10(1):93–137, March 1995. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02087962;   http://link.springer.com/content/pdf/10.
1007/BF02087962;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=1&spage=93-137.

Trayner:1995:NTS

[165] C. Trayner and M. H. Glowacki. A new technique for the solution of the Saha equation. Journal of Scientific Computing, 10(1): 139–149, March 1995. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02087963;   http://link.springer.com/content/pdf/10.
1007/BF02087963;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=1&spage=139-149.

Miller:1995:SSS

[166] R. S. Miller, F. A. Jaberi, C. K. Madnia, and P. Givi. The structure and the small-scale intermittency of passive scalars in homogeneous turbulence. Journal of Scientific Computing, 10(1):151–180, March 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02087964;   http://link.springer.com/content/pdf/10.

1007/BF02087964;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=1&spage=151-180.

Schumann:1995:PST

[167] U. Schumann and M. Strietzel. Parallel solution of tridiagonal systems for the Poisson equation. Journal of Scientific Computing, 10 (2):181–190, June 1995. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02089949;   http://link.springer.com/content/pdf/10.
1007/BF02089949;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=2&spage=181-190.

Konstantinov:1995:ELP

[168] Alexander B. Konstantinov and Steven A. Orszag. Extended Lagrangian particle-in-cell (ELPIC) code for inhomogeneous compressible flows. Journal of Scientific Computing, 10(2):191–231, June 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02089950;   http://link.springer.com/content/pdf/10.
1007/BF02089950;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=2&spage=191-231.

Joslin:1995:SPS

[169] Ronald D. Joslin, Ulf R. Hanebutte, and Mohammad Zubair. Scalability of parallel spatial direct numerical simulations on Intel Hypercube and IBM SP1 and SP2. Journal of Scientific Computing, 10 (2):233–269, June 1995. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02089951;   http://link.springer.com/content/pdf/10.
1007/BF02089951;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=2&spage=233-269.

Hosokawa:1995:CFS

[170] Iwao Hosokawa and Kiyoshi Yamamoto. Conditional Fourier spectral method for direct numerical simulation of incompressible flows. Journal of Scientific Computing, 10(2):271–287, June 1995. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02089952;
http://link.springer.com/content/pdf/10.1007/BF02089952;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=10&issue=2&spage=271-287.

Pasquarelli:1995:SMA

[171] F. Pasquarelli. Spectral multidomain approximation of elliptic problems with mixed conditions on the interfaces. Journal of Scientific Computing, 10(3):289–304, September 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02091777;   http://link.springer.com/
content/pdf/10.1007/BF02091777;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=3&
spage=289-304.

Gomez-Valdes:1995:MPP

[172] José Gómez-Valdés and Dong-Ping Wang. Massively parallel processing in coastal ocean circulation model. Journal of Scientific Computing, 10(3):305–323, September 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02091778;   http://link.springer.com/
content/pdf/10.1007/BF02091778;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=3&
spage=305-323.

Geer:1995:RTA

[173] James F. Geer. Rational trigonometric approximations using Fourier series partial sums. Journal of Scientific Computing, 10(3):325–356, September 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02091779;   http://link.springer.com/content/pdf/10.
1007/BF02091779;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=3&spage=325-356.

Shu:1995:NAS

[174] Chi-Wang Shu and Peter S. Wong. A note on the accuracy of spectral method applied to nonlinear conservation laws. Journal of Scientific Computing, 10(3):357–369, September 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02091780;   http://link.springer.com/
content/pdf/10.1007/BF02091780;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=3&
spage=357-369.

Dettori:1995:NGM

[175] Lucia Dettori, David Gottlieb, and Roger Temam. A nonlinear Galerkin method: The two-level Fourier-collocation case. Journal of Scientific Computing, 10(4):371–389, December 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02088956;   http://link.springer.com/
content/pdf/10.1007/BF02088956;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=4&
spage=371-389.

Yang:1995:MRA

[176] Sang Kyu Yang and Charlie H. Cooke. Multi-resolution analysis on the interval with natural spline projection and uniform twoscale relation. Journal of Scientific Computing, 10(4):391–407, December 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088957;   http://link.springer.com/content/pdf/10.
1007/BF02088957;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=10&issue=4&spage=391-407.

Kukharkin:1995:CSF

[177] Nikolai N. Kukharkin. Coherent structure formation, phase correlations, and finite-size effects in 2D turbulence. Journal of Scientific Computing, 10(4):409–448, December 1995. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02088958;   http://link.springer.com/
content/pdf/10.1007/BF02088958;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=10&issue=4&
spage=409-448.

Manoranjan:1996:SIP

[178] Valipuram S. Manoranjan and Yafang Song. A simple ILU preconditioning coupled to orthomin method. Journal of Scientific Computing, 11(1):1–12, March 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088947;   http://link.springer.com/content/pdf/10.
1007/BF02088947;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=1&spage=1-12.

Sukoriansky:1996:LES

[179] Semion Sukoriansky, Alexei Chekhlov, and Steven A. Orszag. Large eddy simulation of two-dimensional isotropic turbulence. Journal of Scientific Computing, 11(1):13–45, March 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02088948;   http://link.springer.com/
content/pdf/10.1007/BF02088948;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=11&issue=1&
spage=13-45.

Yavneh:1996:MSS

[180] Irad Yavneh and James C. McWilliams. Multigrid solution of stably stratified flows: The quasigeostrophic equations. Journal of Scientific Computing, 11(1):47–69, March 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02088949;   http://link.springer.com/
content/pdf/10.1007/BF02088949;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=11&issue=1&
spage=47-69.

Yakhot:1996:ATM

[181] A. Yakhot, E. Shalman, O. Igra, and Y. Yadlin. An algebraic-Q_{4} turbulence model for attached and separated airfoil flows. Journal of Scientific Computing, 11(2):71–98, June 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02088818;   http://link.springer.com/
content/pdf/10.1007/BF02088818;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=11&issue=2&
spage=71-98.

Lee:1996:CSS

[182] Chang-Ock Lee. A comparison of some standard elliptic solvers: CM-5 vs. Cray C-90. Journal of Scientific Computing, 11(2):99– 126, June 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088819;   http://link.springer.com/content/pdf/10.
1007/BF02088819;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=2&spage=99-126.

Jerri:1996:SPB

[183] Abdul J. Jerri and Russell L. Herman. The solution of the Poisson–Boltzmann equation between two spheres: Modified iterative method. Journal of Scientific Computing, 11(2):127–153, June 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088820;   http://link.springer.com/content/pdf/10.
1007/BF02088820;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=2&spage=127-153.

Chen:1996:CRR

[184] Jyh-Jia Chen and Chyi Hwang. Characterization of robust root loci of polytopes of polynomials. Journal of Scientific Computing, 11(2): 155–166, June 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088821;   http://link.springer.com/content/pdf/10.
1007/BF02088821;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=2&spage=155-166.

Karageorghis:1996:NCM

[185] A. Karageorghis, N. S. Stylianopoulos, and H. A. Zachariades. A numerical conformal mapping method for harmonic mixed boundary value problems. Journal of Scientific Computing, 11(3):167– 178, September 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088814;   http://link.springer.com/content/pdf/10.
1007/BF02088814;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=3&spage=167-178.

Ladeinde:1996:PEP

[186] F. Ladeinde, E. E. O’Brien, and X. D. Cai. A parallelized ENO procedure for direct numerical simulation of compressible turbulence. Journal of Scientific Computing, 11(3):179–205, September 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088815;   http://link.springer.com/content/pdf/10.
1007/BF02088815;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=3&spage=179-205.

Stricker:1996:ITF

[187] Josef Stricker, Yair Censor, and Boris Zakharin. Inhomogeneous turbulent field diagnostics by distance-dependent tomographic reconstruction techniques. Journal of Scientific Computing, 11(3):207– 227, September 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088816;   http://link.springer.com/content/pdf/10.
1007/BF02088816;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=3&spage=207-227.

Gustafsson:1996:HOC

[188] Bertil Gustafsson and Pelle Olsson. High-order centered difference methods with sharp shock resolution. Journal of Scientific Computing, 11(3):

229–260, September 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088817;   http://link.springer.com/content/pdf/10.
1007/BF02088817;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=3&spage=229-260.

Anonymous:1996:IA

[189] Anonymous. Important announcement. Journal of Scientific Computing, 11(4):297, December 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088950;   http://link.springer.com/content/pdf/10.
1007/BF02088950;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=4&spage=297.

Ciccoli:1996:ADD

[190] M. C. Ciccoli. Adaptive domain decomposition algorithms and finite volume/finite element approximation for advection-diffusion equations. Journal of Scientific Computing, 11(4):299–341, December 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088951;   http://link.springer.com/content/pdf/10.
1007/BF02088951;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=4&spage=299-341.

Bowman:1996:WPS

[191] John C. Bowman. A wavenumber partitioning scheme for twodimensional statistical closures. Journal of Scientific Computing, 11(4): 343–372, December 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088952;   http://link.springer.com/content/pdf/10.
1007/BF02088952;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=4&spage=343-372.

Cao:1996:PML

[192] Wei-Ming Cao and Claudio Carlenzoli. A projection method for longterm computation of Navier–Stokes equations. Journal of Scientific Computing, 11(4):373–395, December 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02088953;   http://link.springer.com/
content/pdf/10.1007/BF02088953;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=11&issue=4&
spage=373-395.

Heinrichs:1996:HOT

[193] Wilhelm Heinrichs. High-order time splitting for the Stokes equations. Journal of Scientific Computing, 11(4):397–410, December 1996. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088954;   http://link.springer.com/content/pdf/10.
1007/BF02088954;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=4&spage=397-410.

Liu:1996:NES

[194] W. B. Liu and Jie Shen. A new efficient spectral Galerkin method for singular perturbation problems. Journal of Scientific Computing, 11(4): 411–437, December 1996. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/BF02088955;   http://link.springer.com/content/pdf/10.
1007/BF02088955;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=11&issue=4&spage=411-437.

Hollerbach:1997:EAE

[195] Uwe Hollerbach. Effects of aliasing errors on microlithographic image computations. Journal of Scientific Computing, 12(1):3–10, March 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025685102440;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=1&spage=3-10.

Brakkee:1997:IIC

[196] Erik Brakkee and Peter Wilders. The influence of interface conditions on convergence of Krylov–Schwarz domain decomposition for the advection– diffusion equation. Journal of Scientific Computing, 12(1):11–30, March 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025602319278;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=1&spage=11-30.

Luo:1997:PTMa

[197] Yong Luo and Matthew J. Yedlin. Polynomial time-marching for nonreflecting boundary problems. Journal of Scientific Computing, 12(1): 31–50, March 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025654303349;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=12&issue=1&spage=31-

Hon:1997:MIM

[198] Y. C. Hon and X. Z. Mao. A multiquadric interpolation method for solving initial value problems. Journal of Scientific Computing, 12(1):51–55, March 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025606420187;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=1&spage=51-55.

Ierley:1997:CSS

[199] Glenn R. Ierley. A class of sparse spectral operators for inversion of powers of the Laplacian in N dimensions. Journal of Scientific Computing, 12(1):57–73, March 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025658404257;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=1&spage=57-73.

Nguyen:1997:LSE

[200] Hoa D. Nguyen, Seungho Paik, and Rod W. Douglass. A Legendre– spectral element method for flow and heat transfer about an accelerating droplet. Journal of Scientific Computing, 12(1):75–97, March 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025610521095;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=1&spage=75-97.

Yang:1997:PGM

[201] Daoqi Yang. A parallel grid modification and domain decomposition algorithm for local phenomena capturing and load balancing. Journal of Scientific Computing, 12(1):99–117, March 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025662505166;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=1&spage=99-117.

Eckhoff:1997:HON

[202] Knut S. Eckhoff. On a high order numerical method for solving partial differential equations in complex geometries. Journal of Scientific Computing, 12(2):119–138, June 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025617731306;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=2&spage=119-138.

Tomboulides:1997:NSL

[203] A. G. Tomboulides, J. C. Y. Lee, and S. A. Orszag. Numerical simulation of low Mach number reactive flows. Journal of Scientific Computing, 12(2):139–167, June 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025669715376;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=2&spage=139-167.

Chen:1997:TPT

[204] Y. Chen, H. Ohashi, and M. Akiyama. Two-parameter thermal lattice BGK model with a controllable Prandtl number. Journal of Scientific Computing, 12(2):169–185, June 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025621832215;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=2&spage=169-185.

Oh:1997:SDD

[205] Seyoung Oh, Seungho Paik, and Hoa D. Nguyen. Solution-domain-decomposition method for heat transfer problem using parallel distributed computing. Journal of Scientific Computing, 12(2):187–204, June 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025673816285;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=2&spage=187-204.

Cunha:1997:SIA

[206] Cristina Cunha and Sônia M. Gomes. A shock indicator for adaptive schemes for conservation laws. Journal of Scientific Computing, 12(2): 205–214, June 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025625900355;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=12&issue=2&spage=205-
214.

vanderHouwen:1997:SMT

[207] P. J. van der Houwen and B. P. Sommeijer. Splitting methods for threedimensional transport models with interaction terms. Journal of Scientific Computing, 12(3):215–231, September 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025645326705;   http:

//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=3&spage=215-231.

Marco:1997:AMO

[208] Nathalie Marco, Bruno Koobus, and Alain Dervieux. An additive multilevel optimization method and its application to unstructured meshes. Journal of Scientific Computing, 12(3):233–251, September 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025697310775;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=3&spage=233-251.

Geer:1997:EAA

[209] James Geer and Nana Saheb Banerjee. Exponentially accurate approximations to piece-wise smooth periodic functions. Journal of Scientific Computing, 12(3):253–287, September 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025649427614;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=3&spage=253-287.

Arsham:1997:AGM

[210] H. Arsham. Affine geometric method for linear programs. Journal of Scientific Computing, 12(3):289–303, September 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025601511684;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=3&spage=289-303.

Borue:1997:TCD

[211] Vadim Borue and Steven A. Orszag. Turbulent convection driven by a constant temperature gradient. Journal of Scientific Computing, 12(3): 305–351, September 1997. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1025653628522;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=12&issue=3&
spage=305-351.

Dai:1997:GPR

[212] Weizhong Dai. A generalized Peaceman–Rachford ADI scheme for solving two-dimensional parabolic differential equations. Journal of Scientific Computing, 12(4):353–360, December 1997. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025631211217;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=4&spage=353-360.

Dai:1997:NAS

[213] Weizhong Dai. A new ADI scheme for solving three-dimensional parabolic differential equations. Journal of Scientific Computing, 12(4): 361–369, December 1997. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1025620828055;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=12&issue=4&
spage=361-369.

Kastner:1997:SET

[214] S. O. Kastner. Symbolic evaluation of total transition probabilities in kinetic systems: Potential applications. Journal of Scientific Computing, 12(4):371–384, December 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025672812125;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=4&spage=371-384.

Funaro:1997:IPI

[215] Daniele Funaro. Improving the performances of implicit schemes for hyperbolic equations. Journal of Scientific Computing, 12(4):385–394, December 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025624928964;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=12&issue=4&spage=385-
394.

Succi:1997:LKT

[216] S. Succi, G. Bella, and F. Papetti. Lattice kinetic theory for numerical combustion. Journal of Scientific Computing, 12(4):395–408, December 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025676913034;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=4&spage=395-408.

Wu:1997:PSC

[217] Jung-Gen Wu, Wen-Ming Yan, and Kuo-Liang Chung. A parallel solver for circulant Toeplitz tridiagonal systems on hypercubes. Journal of Scientific Computing, 12(4):409–431, December 1997. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025629029872;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=4&spage=409-431.

Zheligovsky:1997:OIM

[218] V. A. Zheligovsky and O. M. Podvigina. An optimized iterative method for numerical solution of large systems of equations based on the extremal property of zeroes of Chebyshev polynomials. Journal of Scientific Computing, 12(4):433–464, December 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025681013942;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=12&issue=4&spage=433-464.

Luo:1997:PTMb

[219] Yong Luo. Polynomial time-marching for three-dimensional wave equations. Journal of Scientific Computing, 12(4):465–477, December 1997. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025633130781;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=12&issue=4&spage=465-477.

Birkeland:1998:USQ

[220] C. Birkeland, R. Holmestad, K. Marthinsen, and R. Høier. Use of supercomputers in quantitative electron diffraction. Journal of Scientific Computing, 13(1):1–18, March 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023232026276;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=1&spage=1-18.

Holmstrom:1998:AWM

[221] Mats Holmström and Johan Waldén. Adaptive wavelet methods for hyperbolic PDEs. Journal of Scientific Computing, 13(1):19–49, March 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023252610346;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=13&issue=1&spage=19-49.

Simos:1998:SMR

[222] T. E. Simos. Some modified Runge–Kutta methods for the numerical solution of initial-value problems with oscillating solutions. Journal of Scientific Computing, 13(1):51–63, March 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023204727185;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=1&spage=51-63.

Janardhan:1998:NFM

[223] R. Janardhan and T. Downar. A nested FGMRES method for parallel calculation of nuclear reactor transients. Journal of Scientific Computing, 13(1):65–93, March 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023256711255;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=1&spage=65-93.

Funaro:1998:SMU

[224] D. Funaro, M. Giangi, and D. Mansutti. A splitting method for unsteady incompressible viscous fluids imposing no boundary conditions on pressure. Journal of Scientific Computing, 13(1):95–104, March 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023208828093;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=13&issue=1&spage=95-104.

Liou:1998:TIM

[225] Yuan Chang Liou and Yih Nen Jeng. A transfinite interpolation method of grid generation based on multipoints. Journal of Scientific Computing, 13(1):105–114, March 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023260812163;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=1&spage=105-114.

Puppo:1998:BSS

[226] Gabriella Puppo. Bubble stabilization of spectral methods: The multidimensional case. Journal of Scientific Computing, 13(2):115–149, June 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023239710075;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=13&issue=2&spage=115-149.

Shalman:1998:AAD

[227] E. Shalman, A. Yakhot, S. Shalman, O. Igra, and Y. Yadlin. Attenuating artificial dissipation in the computation of Navier–Stokes turbulent boundary layers. Journal of Scientific Computing, 13(2):151–172, June 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023270026913;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=13&issue=2&spage=151-172.

Gottlieb:1998:MCG

[228] Sigal Gottlieb and Paul F. Fischer. Modified conjugate gradient method for the solution of Ax = b. Journal of Scientific Computing, 13(2):173–183, June 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023222110984;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=2&spage=173-183.

Bickham:1998:TMP

[229] Jerry L. Bickham, Germana Peggion, and Benjamin R. Seyfarth. Testing moderately parallel environments for an ocean modeling application. Journal of Scientific Computing, 13(2):185–200, June 1998. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1023274127822;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=13&issue=2&spage=185-200.

Jaberi:1998:EHR

[230] F. A. Jaberi and C. K. Madnia. Effects of heat of reaction on homogeneous compressible turbulence. Journal of Scientific Computing, 13(2): 201–228, June 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1023226211892;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=13&issue=2&spage=201-
228.

Konstantinov:1998:RGBa

[231] Alexandre Konstantinov and Ilya Staroselsky. Renormalization groupbased transport modeling of premixed turbulent combustion: I. Incompressible deflagration model. Journal of Scientific Computing, 13(3):229– 252, September 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.

1023/A%3A1023229615392;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=13&issue=3&spage=229-
252.

Speziale:1998:CLE

[232] Charles G. Speziale. A combined large-eddy simulation and timedependent RANS capability for high-speed compressible flows. Journal of Scientific Computing, 13(3):253–274, September 1998. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1023266932231;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=13&issue=3&spage=253-274.

Castillo:1998:PME

[233] Paul Castillo and Yousef Saad. Preconditioning the matrix exponential operator with applications. Journal of Scientific Computing, 13(3):275– 302, September 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1023219016301;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=13&issue=3&spage=275-
302.

Chang:1998:MAO

[234] H.-C. Chang, D. Gottlieb, M. Marion, and B. W. Sheldon. Mathematical analysis and optimization of infiltration processes. Journal of Scientific Computing, 13(3):303–321, September 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023271100371;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=3&spage=303-321.

Dubois:1998:DMM

[235] T. Dubois and F. Jauberteau. A dynamic multilevel model for the simulation of the small structures in homogeneous isotropic turbulence. Journal of Scientific Computing, 13(3):323–367, September 1998. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1023223217210;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=13&issue=3&spage=323-367.

Konstantinov:1998:RGBb

[236] Alexandr Konstantinov and Steven A. Orszag. Renormalization groupbased transport modeling of premixed turbulent combustion. II. Finite density gradient and direct heat release. Journal of Scientific Computing, 13(4):369–404, December 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023271100835;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=4&spage=369-404.

Chou:1998:NSN

[237] Chung-Chiang Chou and Yuefan Deng. Numerical solutions of a nonlinear evolution system with small dissipation on parallel processors. Journal of Scientific Computing, 13(4):405–417, December 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1023237317673;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=13&issue=4&spage=405-417.

Banerjee:1998:EAA

[238] Nana S. Banerjee and James F. Geer. Exponentially accurate approximations to periodic Lipschitz functions based on Fourier series partial sums. Journal of Scientific Computing, 13(4):419–460, December 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023289301743;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=13&issue=4&spage=419-460.

Qian:1998:NFO

[239] Yuehong Qian and Jose Jimenez. Note on a fourth order finite difference scheme for the wave equation. Journal of Scientific Computing, 13(4):461–469, December 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023241418582;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=4&spage=461-469.

Zhang:1998:TER

[240] Yansi Zhang and Steven A. Orszag. Two-equation RNG transport modeling of high Reynolds number pipe flow. Journal of Scientific Computing, 13(4):471–483, December 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023293402652;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=4&spage=471-483.

Leu:1998:CBL

[241] Jeng-Fan Leu, Jyh-Cheng Jang, and Chyi Hwang. Computation of Battle–Lemarie wavelets using an FFT–based algorithm. Journal of Scientific Computing, 13(4):485–504, December 1998. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023245519490;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=13&issue=4&spage=485-504.

Dubois-Pelerin:1999:OOT

[242] Yves Dubois-Pelerin and Vincent van Kemenade. An objectoriented toolbox for spectral element analysis. Journal of Scientific Computing, 14(1):1–29, March 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025677921253;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=1&spage=1-29.

Jaberi:1999:ECR

[243] F. A. Jaberi and S. James. Effects of chemical reaction on twodimensional turbulence. Journal of Scientific Computing, 14(1):31–72, March 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025672705324;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=14&issue=1&spage=31-72.

Cai:1999:HOM

[244] Wei Cai. High-order mixed current basis functions for electromagnetic scattering of curved surfaces. Journal of Scientific Computing, 14(1): 73–105, March 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025624822162;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=14&issue=1&spage=73-
105.

Karageorghis:1999:CPM

[245] Andreas Karageorghis and Marcin Paprzycki. Conditioning of pseudospectral matrices for certain domain decompositions. Journal of Scientific Computing, 14(1):107–119, March 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025676806232;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=1&spage=107-119.

Yeung:1999:IWS

[246] Michael S. Yeung. Incident wave source for finite-difference time-domain computation of electromagnetic scattering for objects embedded in layered dispersive media. Journal of Scientific Computing, 14(2):121–145, June 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023278902960;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=14&issue=2&spage=121-145.

Orszag:1999:AET

[247] Steven A. Orszag and Victor Yakhot. Analysis of the ϵ-expansion in turbulence theory: Approximate renormalization group for diffusion of a passive scalar in a random velocity field. Journal of Scientific Computing, 14(2):147–178, June 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023228619798;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=2&spage=147-178.

Giangi:1999:SFC

[248] M. Giangi, D. Mansutti, and G. Richelli. Steady 3D flow configurations for the horizontal thermal convection with thermocapillary effects. Journal of Scientific Computing, 14(2):179–195, June 1999. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1023280703869;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=14&issue=2&spage=179-195.

Guoqiang:1999:ENS

[249] Han Guoqiang and Wang Jiong. Extrapolation of Nyström solution for two-dimensional nonlinear Fredholm integral equations. Journal of Scientific Computing, 14(2):197–209, June 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023232820707;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=2&spage=197-209.

Bernaschi:1999:DPS

[250] M. Bernaschi, S. Succi, G. Bella, and H. Chen. Digital physics simulations of reactive flow in a catalytic converter. Journal of Scientific Computing, 14(3):211–222, September 1999. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023216017749;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=3&spage=211-222.

Gustafsson:1999:FDD

[251] Bertil Gustafsson and Lina Hemmingsson-Frändén. A fast domain decomposition high order Poisson solver. Journal of Scientific Computing, 14(3):223–243, September 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023265301820;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=3&spage=223-243.

Keh:1999:SCB

[252] Huan-Chao Keh and Jen-Chih Lin. Simulation of complete binary tree structures in a faulty flexible hypercube. Journal of Scientific Computing, 14(3):245–258, September 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023217518658;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=3&spage=245-258.

Lether:1999:TRI

[253] Frank G. Lether. Thiele rational interpolation for the numerical computation of the reversible Randles–Sevcik function in electrochemistry. Journal of Scientific Computing, 14(3):259–274, September 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023269502728;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=14&issue=3&spage=259-274.

Heinrichs:1999:SAT

[254] Wilhelm Heinrichs. Spectral approximation of third-order problems. Journal of Scientific Computing, 14(3):275–289, September 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023221619567;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=14&issue=3&spage=275-289.

Guo:1999:FLL

[255] Z. L. Guo, B. C. Shi, and N. C. Wang. Fully Lagrangian and lattice Boltzmann methods for the advection–diffusion equation. Journal of Scientific Computing, 14(3):291–300, September 1999. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1023273603637;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=14&issue=3&spage=291-300.

Kvernadze:1999:LDB

[256] George Kvernadze, Thomas Hagstrom, and Henry Shapiro. Locating discontinuities of a bounded function by the partial sums of its Fourier series. Journal of Scientific Computing, 14(4):301–327, December 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023204330916;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=14&issue=4&spage=301-327.

Ganguly:1999:NTB

[257] Kishor Ganguly and Steven A. Orszag. A new time-based iterative solver for linear standing-wave problems. Journal of Scientific Computing, 14(4):329–346, December 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023252514987;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=4&spage=329-346.

Parter:1999:LGL

[258] Seymour V. Parter. On the Legendre–Gauss–Lobatto points and weights. Journal of Scientific Computing, 14(4):347–355, December 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1023204631825;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=14&issue=4&spage=347-355.

Branden:1999:CAH

[259] Henrik Brandén and Sverker Holmgren. Convergence acceleration for hyperbolic systems using semicirculant approximations. Journal of Scientific Computing, 14(4):357–393, December 1999. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1023256615895;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=14&issue=4&spage=357-393.

Yeung:2000:SES

[260] Michael S. Yeung. Solution of electromagnetic scattering problems involving three-dimensional homogeneous dielectric objects by the single integral equation method. Journal of Scientific Computing, 15(1):1–17, March 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1007575417396;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=15&issue=1&spage=1-17.

Tang:2000:PPA

[261] Hua-Zhong Tang and Kun Xu. Positivity-preserving analysis of explicit and implicit Lax–Friedrichs schemes for compressible Euler equations. Journal of Scientific Computing, 15(1):19–28, March 2000. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1007593601466;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=15&issue=1&spage=19-28.

Gustafsson:2000:BCE

[262] Bertil Gustafsson and Jonas Nilsson. Boundary conditions and estimates for the steady Stokes equations on staggered grids. Journal of Scientific Computing, 15(1):29–59, March 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1007545718304;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=1&spage=29-59.

Edelvik:2000:EHT

[263] Fredrik Edelvik and Gunnar Ledfelt. Explicit hybrid time domain solver for the Maxwell equations in 3D. Journal of Scientific Computing, 15(1):61–78, March 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1007625629485;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=1&spage=61-78.

Abarbanel:2000:EBF

[264] Saul Abarbanel, Adi Ditkowski, and Bertil Gustafsson. On error bounds of finite difference approximations to partial differential equations — temporal behavior and rate of convergence. Journal of Scientific Computing, 15(1):79–116, March 2000. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1007688522777;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=1&spage=79-116.

Guo:2000:RAA

[265] Ben-Yu Guo, Jie Shen, and Zhong-Qing Wang. A rational approximation and its applications to differential equations on the half line. Journal of Scientific Computing, 15(2):117–147, June 2000. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1007698525506;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=15&issue=2&spage=117-147.

Stanescu:2000:DAD

[266] D. Stanescu, D. A. Kopriva, and M. Y. Hussaini. Dispersion analysis for discontinuous spectral element methods. Journal of Scientific Computing, 15(2):149–171, June 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1007629609576;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=2&spage=149-171.

Tang:2000:CMR

[267] Tao Tang and Jinghua Wang. Convergence of MUSCL relaxing schemes to the relaxed schemes for conservation laws with stiff source terms. Journal of Scientific Computing, 15(2):173–195, June 2000. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1007681726414;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=15&issue=2&spage=173-195.

Augoula:2000:HON

[268] Steeve Augoula and Rémi Abgrall. High order numerical discretization for Hamilton–Jacobi equations on triangular meshes. Journal of Scientific Computing, 15(2):197–229, June 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1007633810484;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=2&spage=197-229.

Yang:2000:LEM

[269] Huanan Yang. A local extrapolation method for hyperbolic conservation laws. I. The ENO underlying schemes. Journal of Scientific Computing, 15(2):231–264, June 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1007685827323;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=2&spage=231-264.

Huang:2000:FDS

[270] Huaxiong Huang and Brian R. Seymour. Finite difference solutions of incompressible flow problems with corner singularities. Journal of Scientific Computing, 15(3):265–292, September 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1011138516712;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=3&spage=265-292.

Gelb:2000:HAS

[271] Anne Gelb. A hybrid approach to spectral reconstruction of piecewise smooth functions. Journal of Scientific Computing, 15(3):293–322, September 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1011126400782;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=15&issue=3&spage=293-
322.

Kang:2000:BCC

[272] Myungjoo Kang, Ronald P. Fedkiw, and Xu-Dong Liu. A boundary condition capturing method for multiphase incompressible flow. Journal of Scientific Computing, 15(3):323–360, September 2000. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1011178417620;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=15&issue=3&spage=323-360.

Liu:2000:PEE

[273] Wenbin Liu and Ningning Yan. A posteriori error estimators for a class of variational inequalities. Journal of Scientific Computing, 15(3):361–393, September 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1011130501691;   http://www.springerlink.com/openurl.

asp?genre=article&issn=0885-7474&volume=15&issue=3&spage=361-
393.

Chalabi:2000:RSH

[274] A. Chalabi and Y. Qiu. Relaxation schemes for hyperbolic conservation laws with stiff source terms: Application to reacting Euler equations. Journal of Scientific Computing, 15(4):395–416, December 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1011189729919;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=15&issue=4&spage=395-416.

Jameson:2000:HOS

[275] Leland Jameson. High order schemes for resolving waves: Number of points per wavelength. Journal of Scientific Computing, 15(4):417–439, December 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1011180613990;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=15&issue=4&spage=417-
439.

Fischer:2000:SUM

[276] Paul F. Fischer and Sigal Gottlieb. Solving A\underline{x} = \underline{b} using a modified conjugate gradient method based on roots of A. Journal of Scientific Computing, 15(4):441–456, December 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1011132730828;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=4&spage=441-456.

Ravindran:2000:ROA

[277] S. S. Ravindran. Reduced-order adaptive controllers for fluid flows using POD. Journal of Scientific Computing, 15(4):457–478, December 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1011184714898;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=15&issue=4&spage=457-478.

Lether:2000:AEN

[278] Frank G. Lether. Analytical expansion and numerical approximation of the Fermi–Dirac integrals F_{j}(x) of order j = −1/2 and j = 1/2. Journal of Scientific Computing, 15(4):479–497, December 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1011136831736;

http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=15&issue=4&spage=479-497.

Thomas:2000:PSI

[279] S. J. Thomas and R. D. Loft. Parallel semi-implicit spectral element methods for atmospheric general circulation models. Journal of Scientific Computing, 15(4):499–518, December 2000. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1011188832645;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=15&issue=4&spage=499-518.

Eliasson:2001:OBC

[280] Bengt Eliasson. Outflow boundary conditions for the Fourier transformed one-dimensional Vlasov–Poisson system. Journal of Scientific Computing, 16(1):1–28, March 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1011132312956;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=1&spage=1-28.

Li:2001:CFO

[281] Ming Li and Tao Tang. A compact fourth-order finite difference scheme for unsteady viscous incompressible flows. Journal of Scientific Computing, 16(1):29–45, March 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1011146429794;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=1&spage=29-45.

Fatkullin:2001:AHO

[282] I. Fatkullin and J. S. Hesthaven. Adaptive high-order finitedifference method for nonlinear wave problems. Journal of Scientific Computing, 16(1):47–67, March 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1011198413865;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=1&spage=47-67.

Lether:2001:VPA

[283] F. G. Lether. Variable precision algorithm for the numerical computation of the Fermi–Dirac function F_{j}(x) of order j = −3/2. Journal of Scientific Computing, 16(1):69–79, March 2001. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1011150530703;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=1&spage=69-79.

Black:2001:SHA

[284] Kelly Black and John B. Geddes. Spectral Hermite approximations for the actively mode-locked laser. Journal of Scientific Computing, 16(2):81–120, June 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1012289822006;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=2&spage=81-120.

Li:2001:LST

[285] Qing Li, Chu-Guang Zheng, Neng-Chao Wang, and Bao-Chang Shi. LBGK simulations of Turing patterns in CIMA model. Journal of Scientific Computing, 16(2):121–134, June 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1012278606077;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=2&spage=121-134.

Bernaschi:2001:ALB

[286] M. Bernaschi, S. Succi, and H. Chen. Accelerated lattice Boltzmann schemes for steady-state flow simulations. Journal of Scientific Computing, 16(2):135–144, June 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1012230722915;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=2&spage=135-144.

Zhou:2001:NCW

[287] Tie Zhou, Yinfan Li, and Chi-Wang Shu. Numerical comparison of WENO finite volume and Runge–Kutta discontinuous Galerkin methods. Journal of Scientific Computing, 16(2):145–171, June 2001. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1012282706985;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=2&spage=145-171.

Cockburn:2001:RKD

[288] Bernardo Cockburn and Chi-Wang Shu. Runge–Kutta discontinuous Galerkin methods for convection-dominated problems. Journal of Scientific Computing, 16(3):173–261, September 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1012873910884;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=3&spage=173-261.

Han:2001:LAB

[289] Houde Han and Xin Wen. The local artificial boundary conditions for numerical simulations of the flow around a submerged body. Journal of Scientific Computing, 16(3):263–286, September 2001. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1012817627722;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=3&spage=263-286.

Hou:2001:EES

[290] L. S. Hou. Error estimates for semidiscrete finite element approximations of the Stokes equations under minimal regularity assumptions. Journal of Scientific Computing, 16(3):287–317, September 2001. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1012869611793;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=3&spage=287-317.

Maury:2001:FBM

[291] Bertrand Maury. A fat boundary method for the Poisson problem in a domain with holes. Journal of Scientific Computing, 16(3):319–339, September 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1012821728631;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=16&issue=3&spage=319-
339.

Smyrlis:2001:SAM

[292] Yiorgos-Sokratis Smyrlis and Andreas Karageorghis. Some aspects of the method of fundamental solutions for certain harmonic problems. Journal of Scientific Computing, 16(3):341–371, September 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1012873712701;

http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=3&spage=341-371.

Hollerbach:2001:TTD

[293] Uwe Hollerbach, Duan-Pin Chen, and Robert S. Eisenberg. Two- and three-dimensional Poisson–Nernst–Planck simulations of current flow through Gramicidin A. Journal of Scientific Computing, 16(4):373– 409, December 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013203223798;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=16&issue=4&spage=373-
409.

Perugia:2001:CLD

[294] Ilaria Perugia and Dominik Schötzau. On the coupling of local discontinuous Galerkin and conforming finite element methods. Journal of Scientific Computing, 16(4):411–433, December 2001. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1013294207868;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=4&spage=411-433.

Liu:2001:SPE

[295] Wenbin Liu and Ningning Yan. Some a posteriori error estimators for p-Laplacian based on residual estimation or gradient recovery. Journal of Scientific Computing, 16(4):435–477, December 2001. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1013246424707;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=4&spage=435-477.

Despres:2001:CDC

[296] Bruno Després and Frédéric Lagoutière. Contact discontinuity capturing schemes for linear advection and compressible gas dynamics. Journal of Scientific Computing, 16(4):479–524, December 2001. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1013298408777;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=16&issue=4&spage=479-524.

Budd:2001:MSN

[297] Chris J. Budd, Huaxiong Huang, and Robert D. Russell. Mesh selection for a nearly singular boundary value problem. Journal of Scientific Computing, 16(4):525–552, December 2001. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1013250525615;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=16&issue=4&spage=525-552.

Singh:2001:PUD

[298] Abhinav Singh and Yong Zhao. Parallel unstructured dynamic grid direct Monte Carlo simulation of molecular gas dynamics and its applications. Journal of Scientific Computing, 16(4):553–568, December 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1013202608140;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=16&issue=4&spage=553-568.

Guo:2001:LBS

[299] Zhaoli Guo, Chuguang Zheng, and T. S. Zhao. A lattice BGK scheme with general propagation. Journal of Scientific Computing, 16(4):569– 585, December 2001. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1013280900427;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=16&issue=4&spage=569-
585.

Gustafsson:2002:F

[300] Bertil Gustafsson. Foreword. Journal of Scientific Computing, 17(1–4): 1–2, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015182625909;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=17&issue=1&spage=1-
2.

Houston:2002:DGF

[301] Paul Houston, Max Jensen, and Endre Süli. hp-discontinuous Galerkin finite element methods with least-squares stabilization. Journal of Scientific Computing, 17(1–4):3–25, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015180009979;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=3-25.

Yan:2002:LDG

[302] Jue Yan and Chi-Wang Shu. Local discontinuous Galerkin methods for partial differential equations with higher order derivatives. Journal of Scientific Computing, 17(1–4):27–47, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015132126817;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=27-47.

Albukrek:2002:DFW

[303] Cem M. Albukrek, Karsten Urban, and Dietmar Rempfer. Divergencefree wavelet analysis of turbulent flows. Journal of Scientific Computing, 17(1–4):49–66, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015184110888;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=49-66.

Funaro:2002:SD

[304] Daniele Funaro. Superconsistent discretizations. Journal of Scientific Computing, 17(1–4):67–79, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015136227726;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=67-79.

Fischer:2002:SEM

[305] Paul F. Fischer, Gerald W. Kruse, and Francis Loth. Spectral element methods for transitional flows in complex geometries. Journal of Scientific Computing, 17(1–4):81–98, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015188211796;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=81-98.

Schatzman:2002:TNC

[306] Michelle Schatzman. Toward non commutative numerical analysis: High order integration in time. Journal of Scientific Computing, 17(1–4):99– 116, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015140328635;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=17&issue=1&spage=99-
116.

Demkowicz:2002:FAA

[307] L. Demkowicz, W. Rachowicz, and Ph. Devloo. A fully automatic hpadaptivity. Journal of Scientific Computing, 17(1–4):117–142, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015192312705;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=117-142.

Eskilsson:2002:DSE

[308] C. Eskilsson and S. J. Sherwin. A discontinuous spectral element model for Boussinesq-Type equations. Journal of Scientific Computing, 17(1–4):143–152, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015144429543;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=143-152.

Serre:2002:AAM

[309] Eric Serre and Serge Gauthier. An auto-adaptive multidomain spectral technique for linear stability analysis: Application to viscous compressible flows. Journal of Scientific Computing, 17(1–4):153–165, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015196413614;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=153-165.

Archibald:2002:REN

[310] Rick Archibald and A. Gelb. Reducing the effects of noise in image reconstruction. Journal of Scientific Computing, 17(1–4):167–180, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015148530452;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=167-180.

Muller:2002:ESH

[311] Bernhard Müller and H. C. Yee. Entropy splitting for high order numerical simulation of vortex sound at low Mach numbers. Journal of Scientific Computing, 17(1–4):181–190, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015100614522;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=181-190.

Chauviere:2002:RSE

[312] Cédric Chauvière and Robert G. Owens. A robust spectral element method for simulations of time-dependent viscoelastic flows, derived from the Brownian configuration field method. Journal of Scientific Computing, 17(1–4):191–199, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015152631360;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=191-199.

Escriva:2002:PUA

[313] X. Escriva, E. Leriche, and T. N. Phillips. Preconditioned Uzawa algorithm for the velocity–pressure–stress formulation of viscoelastic flow problems. Journal of Scientific Computing, 17(1–4):201–210, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015104815431;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=201-210.

Ruuth:2002:TBS

[314] Steven J. Ruuth and Raymond J. Spiteri. Two barriers on strongstability-preserving time discretization methods. Journal of Scientific Computing, 17(1–4):211–220, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015156832269;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=211-220.

Schonauer:2002:HOM

[315] Willi Schönauer and Torsten Adolph. Higher order may be better or may not be better: Investigations with the FDEM (finite difference element method). Journal of Scientific Computing, 17(1–4):221–229, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015108916339;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=221-229.

Schwartzkopff:2002:AHO

[316] T. Schwartzkopff, C. D. Munz, and E. F. Toro. ADER: a highorder approach for linear hyperbolic systems in 2D. Journal of Scientific Computing, 17(1–4):231–240, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015160900410;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=231-240.

Kress:2002:DCM

[317] Wendy Kress and Bertil Gustafsson. Deferred correction methods for initial boundary value problems. Journal of Scientific Computing, 17(1–4):241–251, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015113017248;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=241-251.

Furst:2002:ASO

[318] Jirí Fürst and Karel Kozel. Application of second order TVD and WENO schemes in internal aerodynamics. Journal of Scientific Computing, 17(1–4):253–262, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015165001318;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=253-262.

Puppo:2002:NEP

[319] Gabriella Puppo. Numerical entropy production on shocks and smooth transitions. Journal of Scientific Computing, 17(1–4):263–271, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015117118157;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=263-271.

Pasquetti:2002:HOA

[320] R. Pasquetti and C. J. Xu. High-order algorithms for largeeddy simulation of incompressible flows. Journal of Scientific Computing, 17(1–4):273–284, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015169102227;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=273-284.

Proot:2002:LSS

[321] M. M. J. Proot and M. I. Gerritsma. A least-squares spectral element formulation for the Stokes problem. Journal of Scientific Computing, 17(1–4):285–296, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015121219065;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=285-296.

Gerritsma:2002:ADL

[322] Marc I. Gerritsma and Michael M. J. Proot. Analysis of a discontinuous least squares spectral element method. Journal of Scientific Computing, 17(1–4):297–306, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015173203136;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=297-306.

Berrone:2002:TRW

[323] Stefano Berrone and Laurent Emmel. Towards a realization of a wavelet Galerkin method on non-trivial domains. Journal of Scientific Computing, 17(1–4):307–317, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015278419974;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=307-317.

Jardak:2002:SPC

[324] M. Jardak, C.-H. Su, and G. E. Karniadakis. Spectral polynomial chaos solutions of the stochastic advection equation. Journal of Scientific Computing, 17(1–4):319–338, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015125304044;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=319-338.

Thomas:2002:SIS

[325] S. J. Thomas and R. D. Loft. Semi-implicit spectral element atmospheric model. Journal of Scientific Computing, 17(1–4):339–350, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015129420882;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=339-350.

Fibich:2002:CNB

[326] G. Fibich, B. Ilan, and S. Tsynkov. Computation of nonlinear backscattering using a high-order numerical method. Journal of Scientific Computing, 17(1–4):351–364, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015181404953;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=351-364.

Droll:2002:PMD

[327] P. Droll and M. Schäfer. A pseudospectral multi-domain method for the incompressible Navier–Stokes equations. Journal of Scientific Computing, 17(1–4):365–374, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015133521791;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=365-374.

Lozinski:2002:SUF

[328] Alexei Lozinski, Robert G. Owens, and Alfio Quarteroni. On the simulation of unsteady flow of an Oldroyd–B fluid by spectral methods. Journal of Scientific Computing, 17(1–4):375–383, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015185605861;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=375-383.

Feng:2002:ASE

[329] Huiyu Feng and Catherine Mavriplis. Adaptive spectral element simulations of thin premixed flame sheet deformations. Journal of Scientific Computing, 17(1–4):385–395, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015137722700;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=385-395.

Duster:2002:NIH

[330] A. Düster, A. Niggl, V. Nübel, and E. Rank. A numerical investigation of high-order finite elements for problems of elastoplasticity. Journal of Scientific Computing, 17(1–4):397–404, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015189706770;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=397-404.

Abarbanel:2002:LTB

[331] S. Abarbanel, D. Gottlieb, and J. S. Hesthaven. Long time behavior of the perfectly matched layer equations in computational electromagnetics. Journal of Scientific Computing, 17(1–4):405–422, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%

3A1015141823608;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=405-422.

Goldberg:2002:SCF

[332] Moshe Goldberg and Anna Pidgirnyak. Stability criteria for finite difference approximations to parabolic systems — an update. Journal of Scientific Computing, 17(1–4):423–435, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015193807679;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=423-435.

Maday:2002:PCT

[333] Yvon Maday, Anthony T. Patera, and Gabriel Turinici. A priori convergence theory for reduced-basis approximations of singleparameter elliptic partial differential equations. Journal of Scientific Computing, 17(1–4):437–446, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015145924517;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=437-446.

Maday:2002:RBE

[334] Yvon Maday and Einar M. Rønquist. A reduced-basis element method. Journal of Scientific Computing, 17(1–4):447–459, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015197908587;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=447-459.

Cances:2002:TRB

[335] Eric Cancès, Claude LeBris, Yvon Maday, and Gabriel Turinici. Towards reduced basis approaches in ab initio electronic structure computations. Journal of Scientific Computing, 17(1–4):461–469, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015150025426;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=461-469.

Israeli:2002:HDP

[336] Moshe Israeli, Elena Braverman, and Amir Averbuch. A hierarchical 3-D Poisson modified Fourier solver by domain decomposition. Journal of Scientific Computing, 17(1–4):471–479, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015102109496;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=471-479.

Montero:2002:RMA

[337] Ruben S. Montero, Ignacio M. Llorente, and Manuel D. Salas. A robust multigrid algorithm for the simulation of a yawed flat plate. Journal of Scientific Computing, 17(1–4):481–490, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015154126334;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=481-490.

Chertock:2002:PMK

[338] Alina Chertock and Doron Levy. A particle method for the KdV equation. Journal of Scientific Computing, 17(1–4):491–499, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015106210404;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=491-499.

Gelb:2002:ABC

[339] A. Gelb, Z. Jackiewicz, and B. D. Welfert. Absorbing boundary conditions of the second order for the pseudospectral Chebyshev methods for wave propagation. Journal of Scientific Computing, 17(1–4):501– 512, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015158227243;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=17&issue=1&spage=501-
512.

Dijkstra:2002:DDP

[340] D. Dijkstra. Doubling the degree of precision without doubling the grid when solving a differential equation with a pseudo-spectral collocation method. Journal of Scientific Computing, 17(1–4):513–527, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015110311313;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=513-527.

Naess:2002:MFG

[341] Ole F. Næss and Knut S. Eckhoff. A modified Fourier–Galerkin method for the Poisson and Helmholtz equations. Journal of Scientific Computing, 17(1–4):529–539, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015162328151;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=529-539.

Gustafsson:2002:DCS

[342] Bertil Gustafsson and Lina Hemmingsson-Frändén. Deferred correction in space and time. Journal of Scientific Computing, 17(1–4):541–550, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015114412222;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=17&issue=1&spage=541-
550.

Bruger:2002:CHO

[343] Arnim Brüger, Jonas Nilsson, and Wendy Kress. A compact higher order finite difference method for the incompressible Navier–Stokes equations. Journal of Scientific Computing, 17(1–4):551–560, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015166529060;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=551-560.

Perugia:2002:ALD

[344] Ilaria Perugia and Dominik Schötzau. An hp-analysis of the local discontinuous Galerkin method for diffusion problems. Journal of Scientific Computing, 17(1–4):561–571, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015118613130;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=561-571.

Azaiez:2002:SEP

[345] Mejdi Azaiez. A spectral element projection scheme for incompressible flow with application to the unsteady axisymmetric Stokes problem. Journal of Scientific Computing, 17(1–4):573–584, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%

3A1015170629969;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=573-584.

Xu:2002:SLM

[346] Jin Xu, Dongbin Xiu, and George Em Karniadakis. A semi-Lagrangian method for turbulence simulations using mixed spectral discretizations. Journal of Scientific Computing, 17(1–4):585–597, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015122714039;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=585-597.

Schroll:2002:HRR

[347] H. Joachim Schroll. High resolution relaxed upwind schemes in gas dynamics. Journal of Scientific Computing, 17(1–4):599–607, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015174730877;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=599-607.

Titarev:2002:AAH

[348] V. A. Titarev and E. F. Toro. ADER: Arbitrary high order Godunov approach. Journal of Scientific Computing, 17(1–4):609–618, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015126814947;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=609-618.

Wilhelm:2002:ASE

[349] D. Wilhelm and L. Kleiser. Application of a spectral element method to two-dimensional forward-facing step flow. Journal of Scientific Computing, 17(1–4):619–627, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015178831786;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=619-627.

Stefanica:2002:FFD

[350] Dan Stefanica. FETI and FETI–DP methods for spectral and mortar spectral elements: a performance comparison. Journal of Scientific Computing, 17(1–4):629–638, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015130915856;   http:

//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=629-638.

Weill:2002:SGF

[351] D. Weill and M. O. Deville. Steady gap flows by the spectral and mortar element method. Journal of Scientific Computing, 17(1–4):639–648, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015182932694;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=17&issue=1&spage=639-
648.

Fietier:2002:STD

[352] Nicolas Fiétier and Michel O. Deville. Simulations of time-dependent flows of viscoelastic fluids with spectral element methods. Journal of Scientific Computing, 17(1–4):649–657, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015135016765;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=649-657.

Matache:2002:STS

[353] A.-M. Matache. Sparse two-scale FEM for homogenization problems. Journal of Scientific Computing, 17(1–4):659–669, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1015187000835;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=17&issue=1&spage=659-669.

Ruegg:2002:IGF

[354] A. W. Rüegg. Implementation of generalized finite element methods for homogenization problems. Journal of Scientific Computing, 17(1–4):671– 681, December 2002. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1015139117673;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=17&issue=1&spage=671-
681.

Juarez:2002:NSS

[355] L. H. Juárez, R. Glowinski, and T. W. Pan. Numerical simulation of the sedimentation of rigid bodies in an incompressible viscous fluid by Lagrange multiplier/fictitious domain methods combined with the Taylor–Hood finite element approximation. Journal of Scientific Computing, 17(1–4):683–694, December 2002. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1015191101744;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=17&issue=1&spage=683-694.

Coppola:2002:GSI

[356] Gennaro Coppola and Carlo Meola. Generalization of the spline interpolation based on the principle of the compact schemes. Journal of Scientific Computing, 17(1–4):695–706, December 2002. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1015143218582;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=17&issue=1&spage=695-706.

Jameson:2003:AVH

[357] Leland Jameson. AMR vs high order schemes. Journal of Scientific Computing, 18(1):1–24, February 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1020378726919;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=1&spage=1-24.

Chorfi:2003:HGS

[358] Nejmeddine Chorfi. Handling geometric singularities by the mortar spectral element method I. Case of the Laplace equation. Journal of Scientific Computing, 18(1):25–48, February 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1020382010989;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=1&spage=25-48.

Liu:2003:VGA

[359] Xiyu Liu, John Frazer, and Ming Xi Tang. Visualization and genetic algorithms in minimax theory for nonlinear functionals. Journal of Scientific Computing, 18(1):49–68, February 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1020334127827;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=1&spage=49-68.

Lie:2003:IQR

[360] Knut-Andreas Lie and Sebastian Noelle. An improved quadrature rule for the flux-computation in staggered central difference schemes in multidimensions. Journal of Scientific Computing, 18(1):69–81, February 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020386111898;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=18&issue=1&spage=69-81.

Gottlieb:2003:SSP

[361] Sigal Gottlieb and Lee-Ad J. Gottlieb. Strong stability preserving properties of Runge–Kutta time discretization methods for linear constant coefficient operators. Journal of Scientific Computing, 18(1):83–109, February 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1020338228736;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=18&issue=1&spage=83-109.

Carpenter:2003:CCW

[362] Mark H. Carpenter, David Gottlieb, and Chi-Wang Shu. On the conservation and convergence to weak solutions of global schemes. Journal of Scientific Computing, 18(1):111–132, February 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1020390212806;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=18&issue=1&spage=111-132.

Mattsson:2003:BPS

[363] Ken Mattsson. Boundary procedures for summation-by-parts operators. Journal of Scientific Computing, 18(1):133–153, February 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1020342429644;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=18&issue=1&spage=133-153.

Tsynkov:2003:DSP

[364] S. V. Tsynkov. On the definition of surface potentials for finitedifference operators. Journal of Scientific Computing, 18(2):155–189, April 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1021111713715;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=18&issue=2&spage=155-189.

Li:2003:SCC

[365] Huiyuan Li and Heping Ma. Shifted Chebyshev collocation domain truncation for solving problems on an infinite interval. Journal of Scientific Computing, 18(2):191–213, April 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1021146530553;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=2&spage=191-213.

Nordstrom:2003:HOF

[366] Jan Nordström and Rikard Gustafsson. High order finite difference approximations of electromagnetic wave propagation close to material discontinuities. Journal of Scientific Computing, 18(2):215–234, April 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1021149523112;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=18&issue=2&spage=215-234.

Kwon:2003:PET

[367] YongHoon Kwon, Ben yu Guo, and Ki-Hyun Cha. Parameter estimation in two-dimensional space by mixed Chebyshev–Legendre approximations. Journal of Scientific Computing, 18(2):235–251, April 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1021116907182;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=18&issue=2&spage=235-251.

Liu:2003:FOS

[368] Jian-Guo Liu, Cheng Wang, and Hans Johnston. A fourth order scheme for incompressible Boussinesq equations. Journal of Scientific Computing, 18(2):253–285, April 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1021168924020;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=2&spage=253-285.

Jackiewicz:2003:SGR

[369] Z. Jackiewicz and B. D. Welfert. Stability of Gauss–Radau pseudospectral approximations of the one-dimensional wave equation. Journal of Scientific Computing, 18(2):287–313, April 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1021121008091;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=2&spage=287-313.

Herrero:2003:CCC

[370] H. Herrero, S. Hoyas, A. Donoso, and A. M. Mancho. Chebyshev collocation for a convective problem in primitive variable formulation. Journal of Scientific Computing, 18(3):315–328, June 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1022678124929;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=3&spage=315-328.

Chen:2003:DTP

[371] Z. Chen and R. E. Ewing. Degenerate two-phase incompressible flow IV: Local refinement and domain decomposition. Journal of Scientific Computing, 18(3):329–360, June 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1022673427893;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=18&issue=3&spage=329-360.

Shu:2003:F

[372] Chi-Wang Shu. Foreword. Journal of Scientific Computing, 19(1–3): 1–2, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025334204364;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=19&issue=1&spage=1-
2.

Abgrall:2003:HOF

[373] R. Abgrall and P. L. Roe. High order fluctuation schemes on triangular meshes. Journal of Scientific Computing, 19(1–3):3–36, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025335421202;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=19&issue=1&spage=3-36.

Aslam:2003:LSA

[374] Tariq D. Aslam. A level set algorithm for tracking discontinuities in hyperbolic conservation laws II: Systems of equations. Journal of Scientific Computing, 19(1–3):37–62, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025387405273;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=37-62.

Benamou:2003:IEG

[375] Jean-David Benamou. An introduction to Eulerian geometrical optics (1992-2002). Journal of Scientific Computing, 19(1–3):63–93, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025339522111;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=19&issue=1&spage=63-93.

Bertalmio:2003:TBI

[376] M. Bertalmio, V. Caselles, B. Rougé, and A. Solé. TV based image restoration with local constraints. Journal of Scientific Computing, 19(1–3):95–122, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025391506181;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=95-122.

Cheng:2003:CSA

[377] Li-Tien Cheng. Construction of shapes arising from the Minkowski problem using a level set approach. Journal of Scientific Computing, 19(1–3): 123–138, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1023/A%3A1025343723019;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=19&issue=1&
spage=123-138.

Duraisamy:2003:CAT

[378] Karthikeyan Duraisamy, James D. Baeder, and Jian-Guo Liu. Concepts and application of time-limiters to high resolution schemes. Journal of Scientific Computing, 19(1–3):139–162, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025395707090;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=139-162.

Esmaeeli:2003:CEB

[379] Asghar Esmaeeli and Grétar Tryggvason. Computations of explosive boiling in microgravity. Journal of Scientific Computing, 19(1–3):163– 182, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025347823928;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=19&issue=1&spage=163-

Gibou:2003:LSA

[380] Frédéric Gibou, Ronald Fedkiw, and Russel Caflisch. A level set approach for the numerical simulation of dendritic growth. Journal of Scientific Computing, 19(1–3):183–199, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025399807998;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=183-199.

He:2003:NSP

[381] X. He and A. R. Karagozian. Numerical simulation of pulse detonation engine phenomena. Journal of Scientific Computing, 19(1–3):201– 224, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025351924837;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=19&issue=1&spage=201-
224.

Hudson:2003:FNA

[382] Justin Hudson and Peter K. Sweby. Formulations for numerically approximating hyperbolic systems governing sediment transport. Journal of Scientific Computing, 19(1–3):225–252, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025304008907;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=225-252.

Ito:2003:SNP

[383] Kazufumi Ito and Zhilin Li. Solving a nonlinear problem in magnetorheological fluids using the immersed interface method. Journal of Scientific Computing, 19(1–3):253–266, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025356025745;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=253-266.

Jin:2003:ESS

[384] Hailin Jin, Anthony J. Yezzi, and Yen-Hsi Tsai. Estimation of 3D surface shape and smooth radiance from 2D images: a level set approach. Journal of Scientific Computing, 19(1–3):267–292, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025308109816;

http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=267-292.

Kruzik:2003:CMM

[385] Martin Kruzík and Mitchell Luskin. The computation of Martensitic microstructure with piecewise laminates. Journal of Scientific Computing, 19(1–3):293–308, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025360126654;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=293-308.

Merriman:2003:USO

[386] Barry Merriman. Understanding the Shu–Osher conservative finite difference form. Journal of Scientific Computing, 19(1–3):309–322, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025312210724;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=19&issue=1&spage=309-322.

Mitchell:2003:ORS

[387] Ian M. Mitchell and Claire J. Tomlin. Overapproximating reachable sets by Hamilton–Jacobi projections. Journal of Scientific Computing, 19(1–3):323–346, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025364227563;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=323-346.

Rault:2003:SVI

[388] Audrey Rault, Guillaume Chiavassa, and Rosa Donat. Shockvortex interactions at high Mach numbers. Journal of Scientific Computing, 19(1–3):347–371, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025316311633;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=347-371.

Ruuth:2003:SSV

[389] Steven J. Ruuth and Brian T. R. Wetton. A simple scheme for volume-preserving motion by mean curvature. Journal of Scientific Computing, 19(1–3):373–384, December 2003. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025368328471;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=373-384.

Sanders:2003:NSS

[390] Richard Sanders and Michael Sever. The numerical study of singular shocks regularized by small viscosity. Journal of Scientific Computing, 19(1–3):385–404, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025320412541;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=385-404.

Sebastian:2003:MWF

[391] Kurt Sebastian and Chi-Wang Shu. Multidomain WENO finite difference method with interpolation at subdomain interfaces. Journal of Scientific Computing, 19(1–3):405–438, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025372429380;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=405-438.

Smereka:2003:SIL

[392] Peter Smereka. Semi-implicit level set methods for curvature and surface diffusion motion. Journal of Scientific Computing, 19(1–3):439– 456, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1023/A%3A1025324613450;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=19&issue=1&spage=439-
456.

Steinhoff:2003:CCV

[393] John Steinhoff, Meng Fan, Lesong Wang, and William Dietz. Convection of concentrated vortices and passive scalars as solitary waves. Journal of Scientific Computing, 19(1–3):457–478, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025376630288;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=457-478.

Sussman:2003:DSE

[394] Mark Sussman and M. Y. Hussaini. A discontinuous spectral element method for the level set equation. Journal of Scientific Computing, 19(1–3):479–500, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025328714359;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=19&issue=1&spage=479-500.

Symes:2003:SME

[395] William W. Symes and Jianliang Qian. A slowness matching Eulerian method for multivalued solutions of eikonal equations. Journal of Scientific Computing, 19(1–3):501–526, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025380731197;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=501-526.

Tornberg:2003:RTN

[396] Anna-Karin Tornberg and Björn Engquist. Regularization techniques for numerical approximation of PDEs with singularities. Journal of Scientific Computing, 19(1–3):527–552, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025332815267;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=527-552.

Vese:2003:MTT

[397] Luminita A. Vese and Stanley J. Osher. Modeling textures with total variation minimization and oscillating patterns in image processing. Journal of Scientific Computing, 19(1–3):553–572, December 2003. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025384832106;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=19&issue=1&spage=553-572.

Xu:2003:EFS

[398] Jian-Jun Xu and Hong-Kai Zhao. An Eulerian formulation for solving partial differential equations along a moving interface. Journal of Scientific Computing, 19(1–3):573–594, December 2003. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1023/A%3A1025336916176;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=19&issue=1&spage=573-594.

Jackiewicz:2004:SVP

[399] Z. Jackiewicz, B. D. Welfert, and B. Zubik-Kowal. Spectral versus pseudospectral solutions of the wave equation by waveform relaxation methods. Journal of Scientific Computing, 20(1):1–28, February 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025900611963;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=20&issue=1&spage=1-28.

Svard:2004:CTS

[400] M. Svärd. On coordinate transformations for summation-by-parts operators. Journal of Scientific Computing, 20(1):29–42, February 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025881528802;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=20&issue=1&spage=29-42.

Cole:2004:SSAa

[401] D. C. Cole and Y. Zou. Simulation study of aspects of the classical hydrogen atom interacting with electromagnetic radiation: Circular orbits. Journal of Scientific Computing, 20(1):43–68, February 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025846412872;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=20&issue=1&spage=43-68.

Carrillo:2004:NSL

[402] J. A. Carrillo and T. Goudon. A numerical study on largetime asymptotics of the Lifshitz–Slyozov system. Journal of Scientific Computing, 20(1):69–113, February 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1023/A%3A1025898429710;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=20&issue=1&spage=69-113.

Bucchignani:2004:NMR

[403] E. Bucchignani and D. Mansutti. A numerical modeling of Rayleigh-Marangoni steady convection in a non-uniform differentially heated 3D cavity. Journal of Scientific Computing, 20(1):115–136, February 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025850513781;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=20&issue=1&spage=115-136.

Wang:2004:SFV

[404] Z. J. Wang and Y. Liu. Spectral (finite) volume method for conservation laws on unstructured grids III: One dimensional systems and partition optimization. Journal of Scientific Computing, 20(1):137–157, February 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1023/A%
3A1025896119548;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=20&issue=1&spage=137-157.

Chehab:2004:TES

[405] Jean-Paul Chehab and Bruno Costa. Time explicit schemes and spatial finite differences splittings. Journal of Scientific Computing, 20(2):159–189, April 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000008719.48134.4f;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=20&issue=2&spage=159-189.

Pennacchio:2004:MFE

[406] Micol Pennacchio. The mortar finite element method for the cardiac “bidomain” model of extracellular potential. Journal of Scientific Computing, 20(2):191–210, April 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000008720.85771.d0;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=20&issue=2&spage=191-210.

Sjogreen:2004:MWB

[407] B. Sjögreen and H. C. Yee. Multiresolution wavelet based adaptive numerical dissipation control for high order methods. Journal of Scientific Computing, 20(2):211–255, April 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000008721.30071.e4;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=20&issue=2&spage=211-255.

Luong:2004:AMG

[408] Phu Luong, Clay P. Breshears, and Le N. Ly. Application of multiblock grid and dual-level parallelism in coastal ocean circulation modeling. Journal of Scientific Computing, 20(2):257–275, April 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000008722.81924.50;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=20&issue=2&spage=257-

275. See erratum [416].

Fang:2004:IFW

[409] Weifu Fang, Yi Wang, and Yuesheng Xu. An implementation of fast wavelet Galerkin methods for integral equations of the second kind. Journal of Scientific Computing, 20(2):277–302, April 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000008723.85496.ce;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=20&issue=2&spage=277-
302.

Li:2004:NCG

[410] Ke Li and Zi-Niu Wu. Nonet–Cartesian grid method for shock flow computations. Journal of Scientific Computing, 20(3):303–329, June 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000025931.79444.99;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=20&issue=3&spage=303-
329.

Hornell:2004:AIS

[411] Karl Hörnell and Per Lötstedt. Adaptive iteration to steady state of flow problems. Journal of Scientific Computing, 20(3):331–354, June 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000025928.28208.32;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=20&issue=3&spage=331-
354.

Bernard:2004:SDS

[412] J. M. Bernard. Spectral discretizations of the Stokes equations with non standard boundary conditions. Journal of Scientific Computing, 20(3):355–377, June 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000025929.09802.f3;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=20&issue=3&spage=355-377.

Cole:2004:SSAb

[413] Daniel C. Cole and Yi Zou. Simulation study of aspects of the classical hydrogen atom interacting with electromagnetic radiation: Elliptical orbits. Journal of Scientific Computing, 20(3):379–404, June 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000025930.62480.05;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=20&issue=3&spage=379-
404.

Kabakian:2004:UGB

[414] Adour V. Kabakian, Vijaya Shankar, and William F. Hall. Unstructured grid-based discontinuous Galerkin method for broadband electromagnetic simulations. Journal of Scientific Computing, 20(3):405–431, June 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000025932.17082.18;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=20&issue=3&spage=405-
431.

Gelb:2004:POR

[415] Anne Gelb. Parameter optimization and reduction of round off error for the Gegenbauer reconstruction method. Journal of Scientific Computing, 20(3):433–459, June 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000025933.39334.17;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=20&issue=3&spage=433-459.

Anonymous:2004:EAM

[416] Anonymous. Erratum: Application of Multiblock Grid and Dual-Level Parallelism in Coastal Ocean Circulation Modeling. Journal of Scientific Computing, 20(3):461–462, June 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000026109.40507.ff;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=20&issue=3&spage=461-462. See [408].

Bouchut:2004:AES

[417] F. Bouchut. An antidiffusive entropy scheme for monotone scalar conservation laws. Journal of Scientific Computing, 21(1):1–30, August 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000027953.74841.8c;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=1&spage=1-
30.

Li:2004:MMM

[418] R. Li, W. B. Liu, and H. P. Ma. Moving mesh method with errorestimator-based monitor and its applications to static obstacle problem. Journal of Scientific Computing, 21(1):31–55, August 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000027954.83289.00;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=1&spage=31-
55.

Mattsson:2004:SAA

[419] Ken Mattsson, Magnus Svärd, and Jan Nordström. Stable and accurate artificial dissipation. Journal of Scientific Computing, 21 (1):57–79, August 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000027955.75872.3f;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=21&issue=1&spage=57-79.

Heinrichs:2004:LSS

[420] Wilhelm Heinrichs. Least-squares spectral collocation for the Navier– Stokes equations. Journal of Scientific Computing, 21(1):81–90, August 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000027956.13510.5a;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=1&spage=81-
90.

Braverman:2004:FSS

[421] Elena Braverman, Boris Epstein, and Moshe Israeli. A fast spectral subtractional solver for elliptic equations. Journal of Scientific Computing, 21(1):91–128, August 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1023/B%3AJOMP.0000027957.39059.6b;   http://
www.springerlink.com/openurl.asp?genre=article&issn=0885-7474&
volume=21&issue=1&spage=91-128.

Alexander:2004:HOC

[422] P. Alexander. High order computation of the history term in the equation of motion for a spherical particle in a fluid. Journal of Scientific Computing, 21(2):129–143, October 2004. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000030072.32108.d9;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=2&spage=129-
143.

Cole:2004:PAS

[423] Daniel C. Cole and Yi Zou. Perturbation analysis and simulation study of the effects of phase on the classical hydrogen atom interacting with circularly polarized electromagnetic radiation. Journal of Scientific Computing, 21(2):145–172, October 2004. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000030073.54361.ba;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=2&spage=145-
172.

Lewandowski:2004:PCS

[424] J. L. V. Lewandowski. Particle-in-cell simulations with kinetic electrons. Journal of Scientific Computing, 21(2):173–192, October 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000030074.13977.45;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=2&spage=173-
192.

Higueras:2004:SSP

[425] Inmaculada Higueras. On strong stability preserving time discretization methods. Journal of Scientific Computing, 21(2):193–223, October 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000030075.59237.61;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=2&spage=193-
223.

Sengupta:2004:CST

[426] T. K. Sengupta and A. Dipankar. A comparative study of time advancement methods for solving Navier–Stokes equations. Journal of Scientific Computing, 21(2):225–250, October 2004. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000030076.74896.d7;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=2&spage=225-
250.

Schroll:2004:RHR

[427] H. Joachim Schroll. Relaxed high resolution schemes for hyperbolic conservation laws. Journal of Scientific Computing, 21(2):251–279, October 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/content/pdf/10.1023/B%3AJOMP.
0000035624.42048.db;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=21&issue=2&spage=251-
279.

Sagaut:2004:F

[428] Pierre Sagaut. Foreword. Journal of Scientific Computing, 21(3): 251, December 2004. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-1316-3;   http://link.springer.com/content/
pdf/10.1007/s10915-004-1316-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=21&issue=3&
spage=251.

Sengupta:2004:HAC

[429] Tapan K. Sengupta, G. Ganerwal, and Anurag Dipankar. High accuracy compact schemes and Gibbs’ Phenomenon. Journal of Scientific Computing, 21(3):253–268, December 2004. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-1317-2;
 http://link.springer.com/content/pdf/10.1007/s10915-004-1317-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=21&issue=3&spage=253-268.

Sengupta:2004:NSS

[430] Tapan K. Sengupta, A. Guntaka, and S. Dey. Navier–Stokes solution by new compact scheme for incompressible flows. Journal of Scientific Computing, 21(3):269–282, December 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-1318-1;
 http://link.springer.com/content/pdf/10.1007/s10915-004-1318-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=21&issue=3&spage=269-282.

Uzun:2004:ACS

[431] Ali Uzun, Gregory A. Blaisdell, and Anastasios S. Lyrintzis. Application of compact schemes to large eddy simulation of turbulent jets. Journal of Scientific Computing, 21(3):283–319, December 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-1319-0;   http://link.springer.com/content/pdf/10.1007/
s10915-004-1319-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=21&issue=3&spage=283-319.

Zhang:2004:HOC

[432] Xiangang Zhang and Gregory A. Blaisdell. High-order compact schemes with filters on multi-block domains. Journal of Scientific Computing, 21 (3):321–339, December 2004. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-1320-7;   http://link.springer.com/content/
pdf/10.1007/s10915-004-1320-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=21&issue=3&
spage=321-339.

Broeckhoven:2004:FVF

[433] Tim Broeckhoven, Sergey Smirnov, and Jan Ramboer. Finite volume formulation of compact upwind and central schemes with artificial selective damping. Journal of Scientific Computing, 21(3):341–367, December 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-1321-6;   http://link.springer.com/content/pdf/10.1007/
s10915-004-1321-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=21&issue=3&spage=341-367.

Monnier:2004:ALH

[434] J. Monnier and P. Witomski. Analysis of a local hydrodynamic model with Marangoni effect. Journal of Scientific Computing, 21(3):369– 403, December 2004. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4095-y;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4095-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=21&issue=3&
spage=369-403.

Rousseau:2004:BLO

[435] A. Rousseau, R. Temam, and J. Tribbia. Boundary layers in an ocean related system. Journal of Scientific Computing, 21(3):405–432, December 2004. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4096-x;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4096-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=21&issue=3&spage=405-432.

Anonymous:2005:F

[436] Anonymous. Foreword. Journal of Scientific Computing, 22–23(1–3):1–2, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4131-y;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4131-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=1-2.

Adjerid:2005:SDF

[437] Slimane Adjerid and Andreas Klauser. Superconvergence of discontinuous finite element solutions for transient convection-diffusion problems. Journal of Scientific Computing, 22–23(1–3):5–24, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4133-9;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4133-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=5-24.

Arnold:2005:FDG

[438] Douglas N. Arnold, Franco Brezzi, and L. Donatella Marini. A family of discontinuous Galerkin finite elements for the Reissner– Mindlin plate. Journal of Scientific Computing, 22–23(1–3):25–45, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4134-8;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4134-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=25-45.

Bokhove:2005:FDD

[439] Onno Bokhove. Flooding and drying in discontinuous Galerkin finiteelement discretizations of shallow-water equations. Part 1: One dimension. Journal of Scientific Computing, 22–23(1–3):47–82, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4136-6;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4136-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=47-82.

Brenner:2005:IPM

[440] Susanne C. Brenner and Li-Yeng Sung. C^{0} interior penalty methods for fourth order elliptic boundary value problems on polygonal domains. Journal of Scientific Computing, 22–23(1–3):83–118, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4135-7;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4135-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=83-118.

Brezzi:2005:MDG

[441] F. Brezzi, T. J. R. Hughes, L. D. Marini, and A. Masud. Mixed discontinuous Galerkin methods for Darcy flow. Journal of Scientific Computing, 22–23(1–3):119–145, June 2005. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4150-8;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4150-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=119-145.

Bustinza:2005:PEE

[442] Rommel Bustinza, Gabriel N. Gatica, and Bernardo Cockburn. An a posteriori error estimate for the local discontinuous Galerkin method applied to linear and nonlinear diffusion problems. Journal of Scientific Computing, 22–23(1–3):147–185, June 2005. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4137-5;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4137-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=147-185.

Castillo:2005:PEE

[443] Paul Castillo. An a posteriori error estimate for the local discontinuous Galerkin method. Journal of Scientific Computing, 22–23(1–3):187–204, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4151-7;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4151-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=187-204.

Chen:2005:HOR

[444] Min-Hung Chen, Bernardo Cockburn, and Fernando Reitich. Highorder RKDG methods for computational electromagnetics. Journal of Scientific Computing, 22–23(1–3):205–226, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4152-6;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4152-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=205-226.

Chevaugeon:2005:DGM

[445] N. Chevaugeon, J. Xin, P. Hu, X. Li, and D. Cler. Discontinuous Galerkin methods applied to shock and blast problems. Journal of Scientific Computing, 22–23(1–3):227–243, June 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4138-4;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4138-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=227-243.

Dawson:2005:DGM

[446] Clint Dawson and Vadym Aizinger. A discontinuous Galerkin method for three-dimensional shallow water equations. Journal of Scientific Computing, 22–23(1–3):245–267, June 2005. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4139-3;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4139-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=245-267.

Eskilsson:2005:DGS

[447] C. Eskilsson and S. J. Sherwin. Discontinuous Galerkin spectral/ hp element modelling of dispersive shallow water systems. Journal of Scientific Computing, 22–23(1–3):269–288, June 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4140-x;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4140-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=269-288.

Feng:2005:TLN

[448] Xiaobing Feng and Ohannes A. Karakashian. Two-level non-overlapping Schwarz preconditioners for a discontinuous Galerkin approximation of the biharmonic equation. Journal of Scientific Computing, 22–23(1– 3):289–314, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4141-9;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4141-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=22&issue=1&
spage=289-314.

Houston:2005:MDG

[449] Paul Houston, Ilaria Perugia, and Dominik Schötzau. Mixed discontinuous Galerkin approximation of the Maxwell operator: Non-stabilized formulation. Journal of Scientific Computing, 22–23(1–3):315–346, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4142-8;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4142-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=315-346.

Paul:2005:ENS

[450] Houston Paul, Dominik Schötzau, and Thomas P. Wihler. Energy norm shape a posteriori error estimation for mixed discontinuous Galerkin approximations of the Stokes problem. Journal of Scientific Computing, 22– 23(1–3):347–370, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4143-7;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4143-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=22&issue=1&
spage=347-370.

Guido:2005:BPL

[451] Kanschat Guido. Block preconditioners for LDG discretizations of linear incompressible flow problems. Journal of Scientific Computing, 22–23 (1–3):371–384, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4144-6;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4144-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=22&issue=1&
spage=371-384.

Kirby:2005:SNF

[452] Robert M. Kirby and George Em Karniadakis. Selecting the numerical flux in discontinuous Galerkin methods for diffusion problems. Journal of Scientific Computing, 22–23(1–3):385–411, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

004-4145-5;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4145-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=385-411.

Li:2005:LDF

[453] Fengyan Li and Chi-Wang Shu. Locally divergence-free discontinuous Galerkin methods for MHD equations. Journal of Scientific Computing, 22–23(1–3):413–442, June 2005. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4146-4;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4146-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=413-442.

Monk:2005:DGM

[454] Peter Monk and Gerard R. Richter. A discontinuous Galerkin method for linear symmetric hyperbolic systems in inhomogeneous media. Journal of Scientific Computing, 22–23(1–3):443–477, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4132-5;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4132-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=443-477.

Riviere:2005:ADF

[455] Béatrice Riviére. Analysis of a discontinuous finite element method for the coupled Stokes and Darcy problems. Journal of Scientific Computing, 22–23(1–3):479–500, June 2005. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4147-3;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4147-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=22&issue=1&spage=479-500.

Sun:2005:NPE

[456] Shuyu Sun and Mary F. Wheeler. L^{2} (H^{1} norm) a posteriori error estimation for discontinuous Galerkin approximations of reactive transport problems. Journal of Scientific Computing, 22–23(1–3):501–530, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-004-4148-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=22&issue=1&spage=501-530.

vanderVegt:2005:DGM

[457] J. J. W. van der Vegt and S. K. Tomar. Discontinuous Galerkin method for linear free-surface gravity waves. Journal of Scientific Computing, 22–23(1–3):531–567, June 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-004-4149-1;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=22&issue=1&spage=531-567.

Shu:2005:NCS

[458] Chi-Wang Shu, Wai-Sun Don, and David Gottlieb. Numerical convergence study of nearly incompressible, inviscid Taylor–Green vortex flow. Journal of Scientific Computing, 24(1):1–27, July 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-5407-y;   http://link.springer.com/content/pdf/10.1007/
s10915-004-5407-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=1&spage=1-27.

Ha:2005:NSH

[459] Youngsoo Ha, Carl L. Gardner, Anne Gelb, and Chi-Wang Shu. Numerical simulation of high Mach number astrophysical jets with radiative cooling. Journal of Scientific Computing, 24(1):29–44, July 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4786-4;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4786-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=1&spage=29-44.

Lottes:2005:HMS

[460] James W. Lottes and Paul F. Fischer. Hybrid multigrid/Schwarz algorithms for the spectral element method. Journal of Scientific Computing, 24(1):45–78, July 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4787-3;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4787-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=24&issue=1&
spage=45-78.

Svard:2005:SSC

[461] Magnus Svärd, Ken Mattsson, and Jan Nordström. Steady-state computations using summation-by-parts operators. Journal of Scientific Computing, 24(1):79–95, July 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-004-4788-2;   http://link.springer.com/
content/pdf/10.1007/s10915-004-4788-2;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=24&issue=
1&spage=79-95.

Higueras:2005:MRK

[462] Inmaculada Higueras. Monotonicity for Runge–Kutta methods: Inner product norms. Journal of Scientific Computing, 24(1):97–117, July 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4789-1;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4789-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=1&spage=97-117.

Tang:2005:P

[463] Tao Tang and Andy Wathen. Preface. Journal of Scientific Computing, 24(2):119, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-5352-4;   http://link.springer.com/content/
pdf/10.1007/s10915-005-5352-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=24&issue=2&
spage=119.

Feng:2005:PEE

[464] Xiaobing Feng and Hai jun Wu. A posteriori error estimates and an adaptive finite element method for the Allen–Cahn equation and the mean curvature flow. Journal of Scientific Computing, 24(2):121–146, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4610-1;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4610-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=2&spage=121-146.

Walkley:2005:FES

[465] M. A. Walkley, P. H. Gaskell, and P. K. Jimack. Finite element simulation of three-dimensional free-surface flow problems. Journal of Scientific Computing, 24(2):147–162, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4611-0;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4611-

0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=24&issue=2&spage=147-162.

Miller:2005:SMF

[466] Keith Miller. Stabilized moving finite elements for convection dominated problems. Journal of Scientific Computing, 24(2):163–182, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4612-z;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4612-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=2&spage=163-182.

Wang:2005:EAM

[467] Li-Lian Wang and Jie Shen. Error analysis for mapped Jacobi spectral methods. Journal of Scientific Computing, 24(2):183–218, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4613-y;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4613-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=2&spage=183-218.

Tourigny:2005:OMF

[468] Yves Tourigny. The optimisation of the mesh in first-order systems leastsquares methods. Journal of Scientific Computing, 24(2):219–245, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4614-x;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4614-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=2&spage=219-245.

Madzvamuse:2005:MGF

[469] Anotida Madzvamuse, Philip K. Maini, and Andrew J. Wathen. A moving grid finite element method for the simulation of pattern generation by Turing models on growing domains. Journal of Scientific Computing, 24(2):247–262, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4617-7;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4617-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=24&issue=2&
spage=247-262.

Zegeling:2005:RMM

[470] Paul A. Zegeling. On resistive MHD models with adaptive moving meshes. Journal of Scientific Computing, 24(2):263–284, August 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4618-6;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4618-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=2&spage=263-284.

Toro:2005:TFH

[471] E. F. Toro and V. A. Titarev. TVD fluxes for the high-order ADER schemes. Journal of Scientific Computing, 24(3):285–309, September 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4790-8;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4790-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=24&issue=3&spage=285-309.

Cordier:2005:FNT

[472] S. Cordier, B. Lucquin-Desreux, and S. Mancini. Focalization: a numerical test for smoothing effects of collision operators. Journal of Scientific Computing, 24(3):311–320, September 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4791-7;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4791-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=24&issue=3&spage=311-320.

Li:2005:MMA

[473] Ruo Li. On multi-mesh H-adaptive methods. Journal of Scientific Computing, 24(3):321–341, September 2005. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4793-5;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4793-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=24&issue=3&spage=321-341.

Cao:2005:PCS

[474] Weiming Cao, Ronald D. Haynes, and Manfred R. Trummer. Preconditioning for a class of spectral differentiation matrices. Journal of Scientific Computing, 24(3):343–371, September 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4794-4;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4794-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=24&issue=3&spage=343-371.

Orszag:2005:PSD

[475] Steven A. Orszag, Isaac Goldhirsch, and Sudhir Srinivasan. ‘perfectly’ scalable data I/O. Journal of Scientific Computing, 24(3):373– 404, September 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-4811-2;   http://link.springer.com/content/
pdf/10.1007/s10915-005-4811-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=24&issue=3&
spage=373-404.

Carpenter:2005:F

[476] Mark H. Carpenter, David I. Gottlieb, and Jan S. Hesthaven. Foreword. Journal of Scientific Computing, 25(1–2):1–2, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728978;
http://link.springer.com/article/10.1007/s10915-005-4649-7;
 http://link.springer.com/content/pdf/10.1007/BF02728978;
http://link.springer.com/content/pdf/10.1007/s10915-005-4649-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=1-2.

Shampine:2005:EEC

[477] L. F. Shampine. Error estimation and control for ODEs. Journal of Scientific Computing, 25(1–2):3–16, October 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728979;
http://link.springer.com/article/10.1007/s10915-004-4629-3;
 http://link.springer.com/content/pdf/10.1007/BF02728979;
http://link.springer.com/content/pdf/10.1007/s10915-004-4629-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=3-16.

Gear:2005:CDM

[478] C. William Gear and Ioannis G. Kevrekidis. Constraint-defined manifolds: a legacy code approach to low-dimensional computation. Journal of Scientific Computing, 25(1–2):17–28, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728980;

http://link.springer.com/article/10.1007/s10915-004-4630-x;
 http://link.springer.com/content/pdf/10.1007/BF02728980;
http://link.springer.com/content/pdf/10.1007/s10915-004-4630-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=17-28.

Jackiewicz:2005:CIG

[479] Z. Jackiewicz. Construction and implementation of general linear methods for ordinary differential equations: a review. Journal of Scientific Computing, 25(1–2):29–49, October 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728981;
http://link.springer.com/article/10.1007/s10915-004-4631-9;
 http://link.springer.com/content/pdf/10.1007/BF02728981;
http://link.springer.com/content/pdf/10.1007/s10915-004-4631-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=29-49.

Butcher:2005:HOS

[480] J. C. Butcher. High order A-stable numerical methods for stiff problems. Journal of Scientific Computing, 25(1–2):51–66, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728982;
http://link.springer.com/article/10.1007/s10915-004-4632-8;
 http://link.springer.com/content/pdf/10.1007/BF02728982;
http://link.springer.com/content/pdf/10.1007/s10915-004-4632-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=51-66.

Hairer:2005:IAG

[481] Ernst Hairer. Important aspects of geometric numerical integration. Journal of Scientific Computing, 25(1–2):67–81, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728983;
http://link.springer.com/article/10.1007/s10915-004-4633-7;
 http://link.springer.com/content/pdf/10.1007/BF02728983;
http://link.springer.com/content/pdf/10.1007/s10915-004-4633-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=67-81.

Ascher:2005:SMS

[482] U. M. Ascher and R. I. McLachlan. On symplectic and multisymplectic schemes for the KdV equation. Journal of Scientific Computing, 25(1–2):83–104, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02728984;   http://link.springer.com/
article/10.1007/s10915-004-4634-6;   http://link.springer.com/
content/pdf/10.1007/BF02728984;   http://link.springer.com/
content/pdf/10.1007/s10915-004-4634-6;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=25&issue=
1&spage=83-104.

Gottlieb:2005:HOS

[483] Sigal Gottlieb. On high order strong stability preserving Runge–Kutta and multi step time discretizations. Journal of Scientific Computing, 25(1–2):105–128, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/BF02728985;   http://link.springer.com/
article/10.1007/s10915-004-4635-5;   http://link.springer.com/
content/pdf/10.1007/BF02728985;   http://link.springer.com/
content/pdf/10.1007/s10915-004-4635-5;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=25&issue=
1&spage=105-128.

Pareschi:2005:IER

[484] Lorenzo Pareschi and Giovanni Russo. Implicit-explicit Runge–Kutta schemes and applications to hyperbolic systems with relaxation. Journal of Scientific Computing, 25(1–2):129–155, October 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728986;
http://link.springer.com/article/10.1007/s10915-004-4636-4;
 http://link.springer.com/content/pdf/10.1007/BF02728986;
http://link.springer.com/content/pdf/10.1007/s10915-004-4636-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=129-155.

Carpenter:2005:FOR

[485] M. H. Carpenter, C. A. Kennedy, Hester Bijl, and S. A. Viken. Fourthorder Runge–Kutta schemes for fluid mechanics applications. Journal of Scientific Computing, 25(1–2):157–194, October 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL http://link.springer.com/article/10.1007/BF02728987;
http://link.springer.com/article/10.1007/s10915-004-4637-3;
 http://link.springer.com/content/pdf/10.1007/BF02728987;
http://link.springer.com/content/pdf/10.1007/s10915-004-4637-

3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=25&issue=1&spage=157-194.

Gustafsson:2005:TCH

[486] Bertil Gustafsson and Per Wahlund. Time compact high order difference methods for wave propagation, 2D. Journal of Scientific Computing, 25(1–2):195–211, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF02728988;   http://link.
springer.com/content/pdf/10.1007/s10915-004-4639-1;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=25&issue=1&spage=195-211.

Knoll:2005:JFN

[487] D. A. Knoll, V. A. Mousseau, L. Chacón, and J. Reisner. Jacobian-free Newton–Krylov methods for the accurate time integration of stiff wave systems. Journal of Scientific Computing, 25(1–2):213–230, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
BF02728989;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4640-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=25&issue=1&spage=213-230.

Faure:2005:FVD

[488] S. Faure, J. Laminie, and R. Temam. Finite volume discretization and multilevel methods in flow problems. Journal of Scientific Computing, 25(1–2):231–261, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF02728990;   http://link.
springer.com/content/pdf/10.1007/s10915-004-4642-6;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=25&issue=1&spage=231-261.

Najm:2005:MLM

[489] H. N. Najm and O. M. Knio. Modeling low Mach number reacting flow with detailed chemistry and transport. Journal of Scientific Computing, 25(1–2):263–287, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF02728991;   http://link.
springer.com/content/pdf/10.1007/s10915-003-4643-x;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=25&issue=1&spage=263-287.

Shang:2005:SSC

[490] J. S. Shang. Solving schemes for computational magneto-aerodynamics. Journal of Scientific Computing, 25(1–2):289–306, October 2005. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/content/pdf/10.1007/BF02728992;   http:
//link.springer.com/content/pdf/10.1007/s10915-004-4645-3;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=25&issue=1&spage=289-306.

Thomas:2005:NSE

[491] S. J. Thomas and R. D. Loft. The NCAR spectral element climate dynamical core: Semi-implicit Eulerian formulation. Journal of Scientific Computing, 25(1–2):307–322, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/BF02728993;   http://link.
springer.com/content/pdf/10.1007/s10915-004-4646-2;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=25&issue=1&spage=307-322.

Xiu:2005:SAF

[492] D. Xiu, S. J. Sherwin, S. Dong, and G. E. Karniadakis. Strong and auxiliary forms of the semi-Lagrangian method for incompressible flows. Journal of Scientific Computing, 25(1–2):323–346, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/BF02728994;   http:/
/link.springer.com/content/pdf/10.1007/s10915-004-4647-1;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=25&issue=1&spage=323-346.

Elman:2005:PSM

[493] H. C. Elman. Preconditioning strategies for models of incompressible flow. Journal of Scientific Computing, 25(1–2):347–366, October 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
BF02728995;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4648-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=25&issue=1&spage=347-366.

Jung:2005:IPR

[494] Jae-Hun Jung and Bernie D. Shizgal. Inverse polynomial reconstruction of two dimensional Fourier images. Journal of Scientific Computing, 25 (3):367–399, December 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4795-3;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4795-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=25&issue=3&
spage=367-399.

Chrysafinos:2005:MMF

[495] Konstantinos Chrysafinos. Moving mesh finite element methods for an optimal control problem for the advection–diffusion equation. Journal of Scientific Computing, 25(3):401–421, December 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4804-6;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4804-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=25&issue=3&spage=401-421.

Guillaume:2005:NCP

[496] Philippe Guillaume and Vladimir Latocha. Numerical convergence of a parameterisation method for the solution of a highly anisotropic twodimensional elliptic problem. Journal of Scientific Computing, 25(3): 423–444, December 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4805-5;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4805-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=25&issue=3&
spage=423-444.

Debussche:2005:DML

[497] A. Debussche, J. Laminie, and E. Zahrouni. A dynamical multi-level scheme for the Burgers equation: Wavelet and hierarchical finite element. Journal of Scientific Computing, 25(3):445–497, December 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4806-4;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4806-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=25&issue=3&spage=445-497.

Reiser:2005:CLD

[498] P. A. Reiser. Calculation of lossy dielectric multilayer filter response. Journal of Scientific Computing, 25(3):499–513, December 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4808-2;   http://link.springer.com/content/pdf/10.1007/

s10915-004-4808-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=25&issue=3&spage=499-513.

Adams:2005:MFE

[499] Scot Adams and Bernardo Cockburn. A mixed finite element method for elasticity in three dimensions. Journal of Scientific Computing, 25(3): 515–521, December 2005. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4807-3;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4807-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=25&issue=3&
spage=515-521.

Wihler:2005:PEA

[500] Thomas P. Wihler. An a priori error analysis of the hp-version of the continuous Galerkin FEM for nonlinear initial value problems. Journal of Scientific Computing, 25(3):523–549, December 2005. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4796-2;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4796-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=25&issue=3&spage=523-549.

Li:2006:HOC

[501] Jichun Li and Miguel R. Visbal. High-order compact schemes for nonlinear dispersive waves. Journal of Scientific Computing, 26(1):1–23, January 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4797-1;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4797-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=26&issue=1&spage=1-23.

Leriche:2006:NEA

[502] E. Leriche, E. Perchat, G. Labrosse, and M. O. Deville. Numerical evaluation of the accuracy and stability properties of high-order direct Stokes solvers with or without temporal splitting. Journal of Scientific Computing, 26(1):25–43, January 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4798-0;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4798-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=26&issue=1&spage=25-43.

Sousa:2006:SAD

[503] Ercília Sousa. Stability analysis of difference methods for parabolic initial value problems. Journal of Scientific Computing, 26(1):45–66, January 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4799-z;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4799-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=26&issue=1&spage=45-66.

Abarbanel:2006:BES

[504] Saul Abarbanel, Adi Ditkowski, and Amir Yefet. Bounded error schemes for the wave equation on complex domains. Journal of Scientific Computing, 26(1):67–81, January 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4800-x;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4800-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=26&issue=1&spage=67-81.

Ferm:2006:STA

[505] Lars Ferm and Per Lötstedt. Space-time adaptive solution of first order PDES. Journal of Scientific Computing, 26(1):83–110, January 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
004-4801-9;   http://link.springer.com/content/pdf/10.1007/
s10915-004-4801-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=26&issue=1&spage=83-110.

Nersessian:2006:RLA

[506] A. Nersessian and A. Poghosyan. On a rational linear approximation of Fourier series for smooth functions. Journal of Scientific Computing, 26 (1):111–125, January 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4809-1;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4809-1;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=26&issue=1&
spage=111-125.

Jin:2006:TSS

[507] Shi Jin and Chunxiong Zheng. A time-splitting spectral method for the generalized Zakharov system in multi-dimensions. Journal of Scientific Computing, 26(2):127–149, February 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-4929-2;
 http://link.springer.com/content/pdf/10.1007/s10915-005-4929-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=26&issue=2&spage=127-149.

Sengupta:2006:HAS

[508] T. K. Sengupta, S. K. Sircar, and A. Dipankar. High accuracy schemes for DNS and acoustics. Journal of Scientific Computing, 26(2):151– 193, February 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-4928-3;   http://link.springer.com/content/
pdf/10.1007/s10915-005-4928-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=26&issue=2&
spage=151-193.

Koko:2006:UCG

[509] Jonas Koko. Uzawa conjugate gradient domain decomposition methods for coupled Stokes flows. Journal of Scientific Computing, 26(2): 195–216, February 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-4933-6;   http://link.springer.com/content/
pdf/10.1007/s10915-005-4933-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=26&issue=2&
spage=195-216.

Feng:2006:AFE

[510] Tao Feng, Mårten Gulliksson, and Wenbin Liu. Adaptive finite element methods for the identification of elastic constants. Journal of Scientific Computing, 26(2):217–235, February 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-004-4935-9;
 http://link.springer.com/content/pdf/10.1007/s10915-004-4935-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=26&issue=2&spage=217-235.

Schroll:2006:BHF

[511] H. J. Schroll and F. Svensson. A bi-hyperbolic finite volume method on quadrilateral meshes. Journal of Scientific Computing, 26(2):237– 260, February 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-004-4927-9;   http://link.springer.com/content/
pdf/10.1007/s10915-004-4927-9;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=26&issue=2&
spage=237-260.

Nielsen:2006:DTG

[512] Michael B. Nielsen and Ken Museth. Dynamic tubular grid: an efficient data structure and algorithms for high resolution level sets. Journal of Scientific Computing, 26(3):261–299, March 2006. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9062-8;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9062-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=26&issue=3&spage=261-299.

Kopriva:2006:MID

[513] David A. Kopriva. Metric identities and the discontinuous spectral element method on curvilinear meshes. Journal of Scientific Computing, 26(3):301–327, March 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9070-8;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9070-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=26&issue=3&
spage=301-327.

Jakobsson:2006:FOC

[514] Stefan Jakobsson. Frequency optimized computation methods. Journal of Scientific Computing, 26(3):329–362, March 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9071-7;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9071-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=26&issue=3&spage=329-362.

Gottlieb:2006:F

[515] David Gottlieb, Jan S. Hesthaven, and George Karniadakis. Foreword. Journal of Scientific Computing, 27(1–3):1–3, June 2006. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9061-9;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9061-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=1-3.

Ainsworth:2006:DDP

[516] M. Ainsworth, P. Monk, and W. Muniz. Dispersive and dissipative properties of discontinuous Galerkin finite element methods for the secondorder wave equation. Journal of Scientific Computing, 27(1–3):5–40, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9044-x;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9044-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=5-40.

Azaiez:2006:SSM

[517] M. Azaïez, R. Gruber, M. O. Deville, and E. H. Mund. On a stable spectral method for the grad(div) eigenvalue problem. Journal of Scientific Computing, 27(1–3):41–50, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9037-9;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9037-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=41-50.

Pavarino:2006:OSS

[518] Luca F. Pavarino and Elena Zampieri. Overlapping Schwarz and spectral element methods for linear elasticity and elastic waves. Journal of Scientific Computing, 27(1–3):51–73, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9047-7;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9047-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=51-73.

Kashdan:2006:HOA

[519] Eugene Kashdan and Eli Turkel. A high-order accurate method for frequency domain Maxwell equations with discontinuous coefficients. Journal of Scientific Computing, 27(1–3):75–95, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9049-5;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9049-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=75-95.

Belhachmi:2006:EDN

[520] Z. Belhachmi, C. Bernardi, S. Deparis, and F. Hecht. An efficient discretization of the Navier–Stokes equations in an axisymmetric domain. Part 1: The discrete problem and its numerical analysis. Journal of Scientific Computing, 27(1–3):97–110, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9035-y;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9035-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=97-110.

Belhachmi:2006:IRS

[521] Z. Belhachmi, A. Karageorghis, and K. Taous. Identification and reconstruction of a small leak zone in a pipe by a spectral element method. Journal of Scientific Computing, 27(1–3):111–122, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9058-4;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9058-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=111-122.

Bodard:2006:FSI

[522] N. Bodard and M. O. Deville. Fluid-structure interaction by the spectral element method. Journal of Scientific Computing, 27(1–3):123–136, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9031-2;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9031-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=123-136.

Bouffanais:2006:MUT

[523] Roland Bouffanais and Michel O. Deville. Mesh update techniques for free-surface flow solvers using spectral element method. Journal of Scientific Computing, 27(1–3):137–149, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9050-z;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9050-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=137-149.

Bouffanais:2006:LES

[524] Roland Bouffanais, Michel O. Deville, and Paul F. Fischer. Largeeddy simulation of the lid-driven cubic cavity flow by the spectral element method. Journal of Scientific Computing, 27(1–3):151–162, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

005-9039-7;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9039-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=151-162.

Bryson:2006:TVH

[525] Steve Bryson and Doron Levy. On the total variation of highorder semi-discrete central schemes for conservation laws. Journal of Scientific Computing, 27(1–3):163–175, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9046-8;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9046-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=163-175.

Celiker:2006:EEP

[526] Fatih Celiker and Bernardo Cockburn. Element-by-element postprocessing of discontinuous Galerkin methods for Timoshenko beams. Journal of Scientific Computing, 27(1–3):177–187, June 2006. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9057-5;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9057-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=177-187.

Chertock:2006:FVP

[527] Alina Chertock, Alexander Kurganov, and Guergana Petrova. Finitevolume-particle methods for models of transport of pollutant in shallow water. Journal of Scientific Computing, 27(1–3):189–199, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9060-x;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9060-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=189-199.

DeMaerschalck:2006:HOG

[528] Bart De Maerschalck and Marc I. Gerritsma. Higher-order Gauss– Lobatto integration for non-linear hyperbolic equations. Journal of Scientific Computing, 27(1–3):201–214, June 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9052-x;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9052-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=201-214.

Dumbser:2006:BBA

[529] Michael Dumbser and Claus-Dieter Munz. Building blocks for arbitrary high order discontinuous Galerkin schemes. Journal of Scientific Computing, 27(1–3):215–230, June 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9025-0;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9025-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=215-230.

Feng:2006:PME

[530] Huiyu Feng, Catherine Mavriplis, and Rob Van der Feng. Parallel 3D mortar element method for adaptive nonconforming meshes. Journal of Scientific Computing, 27(1–3):231–243, June 2006. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9026-z;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9026-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=231-243.

Gerritsma:2006:DMD

[531] Marc I. Gerritsma. Direct minimization of the discontinuous leastsquares spectral element method for viscoelastic fluids. Journal of Scientific Computing, 27(1–3):245–256, June 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9042-z;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9042-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=245-256.

Gervasio:2006:AFS

[532] Paola Gervasio and Fausto Saleri. Algebraic fractional-step schemes for time-dependent incompressible Navier–Stokes equations. Journal of Scientific Computing, 27(1–3):257–269, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9051-y;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9051-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=27&issue=1&spage=257-269.

Gottlieb:2006:FOF

[533] Sigal Gottlieb, Julia S. Mullen, and Steven J. Ruuth. A fifth order flux implicit WENO method. Journal of Scientific Computing, 27(1– 3):271–287, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9034-z;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9034-z;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=27&issue=1&
spage=271-287.

Gottlieb:2006:OSS

[534] Sigal Gottlieb and Steven J. Ruuth. Optimal strong-stability-preserving time-stepping schemes with fast downwind spatial discretizations. Journal of Scientific Computing, 27(1–3):289–303, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9054-8;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9054-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=289-303.

Guo:2006:OSG

[535] Ben-Yu Guo, Jie Shen, and Li-Lian Wang. Optimal Spectral–Galerkin methods using generalized Jacobi polynomials. Journal of Scientific Computing, 27(1–3):305–322, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-005-9055-7;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=27&issue=1&spage=305-322.

Kurganov:2006:ACU

[536] Alexander Kurganov and Guergana Petrova. Adaptive central-upwind schemes for Hamilton–Jacobi equations with nonconvex Hamiltonians. Journal of Scientific Computing, 27(1–3):323–333, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s10915-005-9033-0;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=27&issue=1&spage=323-333.

Leriche:2006:DNS

[537] Emmanuel Leriche. Direct numerical simulation in a lid-driven cubical cavity at high Reynolds number by a Chebyshev spectral method. Journal of Scientific Computing, 27(1–3):335–345, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s10915-005-9032-1;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=27&issue=1&spage=335-345.

Li:2006:MMD

[538] Ruo Li and Tao Tang. Moving mesh discontinuous Galerkin method for hyperbolic conservation laws. Journal of Scientific Computing, 27(1–3): 347–363, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-005-9045-9;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=27&issue=1&spage=347-
363.

Pasquetti:2006:SVV

[539] Richard Pasquetti. Spectral vanishing viscosity method for large-eddy simulation of turbulent flows. Journal of Scientific Computing, 27(1–3): 365–375, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-005-9029-9;   http://www.springerlink.com/openurl.
asp?genre=article&issn=0885-7474&volume=27&issue=1&spage=365-
375.

Pasquetti:2006:SEM

[540] Richard Pasquetti and Francesca Rapetti. Spectral element methods on unstructured meshes: Comparisons and recent advances. Journal of Scientific Computing, 27(1–3):377–387, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-005-9048-6;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=27&issue=1&spage=377-387.

Proot:2006:MMC

[541] Michael M. J. Proot and Marc I. Gerritsma. Mass- and momentum conservation of the least-squares spectral element method for the Stokes problem. Journal of Scientific Computing, 27(1–3):389–401, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-005-9030-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=389-401.

Puppo:2006:SFD

[542] Gabriella Puppo and Giovanni Russo. Staggered finite difference schemes for conservation laws. Journal of Scientific Computing, 27(1–3):403–418, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-005-9036-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=403-418.

Sofronov:2006:SAG

[543] Ivan L. Sofronov and Olga V. Podgornova. A spectral approach for generating non-local boundary conditions for external wave problems in anisotropic media. Journal of Scientific Computing, 27(1–3):419–430, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-005-9041-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=419-430.

Staalberg:2006:HOA

[544] Erik Stålberg, Arnim Brüger, and Per Lötstedt. High order accurate solution of flow past a Circular cylinder. Journal of Scientific Computing, 27(1–3):431–441, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-005-9043-y;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=27&issue=1&spage=431-441.

Talbot:2006:PSM

[545] Chris Talbot and Andrew Crampton. Pseudo spectral methods applied to problems in elasticity. Journal of Scientific Computing, 27(1–3):443–454, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-005-9028-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=443-454.

Wan:2006:BWA

[546] Xiaoliang Wan and George Em Karniadakis. Beyond Wiener– Askey expansions: Handling arbitrary PDFs. Journal of Scientific Computing, 27(1–3):455–464, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-005-9038-8;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=27&issue=1&spage=455-464.

Xenophontos:2006:UCE

[547] Christos Xenophontos. The use of curved elements in the finite element approximation of thin plates by high order p and hp methods. Journal of Scientific Computing, 27(1–3):465–476, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s10915-005-9053-9;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=27&issue=1&spage=465-476.

Xing:2006:HOW

[548] Yulong Xing and Chi-Wang Shu. High-order well-balanced finite difference WENO schemes for a class of hyperbolic systems with source terms. Journal of Scientific Computing, 27(1–3):477–494, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s10915-005-9027-y;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=27&issue=1&spage=477-494.

Xu:2006:SMS

[549] Chuanju Xu. Stabilization methods for spectral element computations of incompressible flows. Journal of Scientific Computing, 27(1–3):495–505, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-005-9059-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=27&issue=1&spage=495-505.

Yee:2006:NLF

[550] H. C. Yee and B. Sjögreen. Non-linear filtering and limiting in high order methods for ideal and non-ideal MHD. Journal of Scientific Computing, 27(1–3):507–521, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-005-9024-1;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=27&issue=1&spage=507-521.

Xiao-yong:2006:SHG

[551] Zhang Xiao-yong and Guo Ben-yu. Spherical harmonic-generalized Laguerre spectral method for exterior problems. Journal of Scientific Computing, 27(1–3):523–537, June 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-005-9056-6;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=27&issue=1&spage=523-537.

Brown:2006:ESH

[552] Matthew Brown and Margot Gerritsen. An energy-stable highorder central difference scheme for the two-dimensional shallow water equations. Journal of Scientific Computing, 28(1):1–30, July 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9005-4;   http://link.springer.com/content/pdf/10.1007/

s10915-005-9005-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=1&spage=1-30.

Tsangaris:2006:MDM

[553] Th. Tsangaris, Y.-S. Smyrlis, and A. Karageorghis. A matrix decomposition MFS algorithm for problems in hollow axisymmetric domains. Journal of Scientific Computing, 28(1):31–50, July 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9006-3;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9006-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=1&spage=31-50.

Benner:2006:SSS

[554] Peter Benner and Enrique S. Quintana-Ortí. Solving stable Sylvester equations via rational iterative schemes. Journal of Scientific Computing, 28(1):51–83, July 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9007-2;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9007-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=1&
spage=51-83.

Yan:2006:FDC

[555] Liu Yan and Mao De-kang. Further development of a conservative front-tracking method for systems of conservation laws in one space dimension. Journal of Scientific Computing, 28(1):85–119, July 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9008-1;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9008-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=1&spage=85-119.

Abarbanel:2006:I

[556] Saul Abarbanel. Introduction. Journal of Scientific Computing, 28(2–3): 121–123, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9093-9;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9093-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=2&
spage=121-123.

Abarbanel:2006:NLP

[557] S. Abarbanel, D. Gottlieb, and J. S. Hesthaven. Non-linear PML equations for time dependent electromagnetics in three dimensions. Journal of Scientific Computing, 28(2–3):125–137, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9072-1;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9072-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=125-137.

Agoshkov:2006:MAD

[558] Valery Agoshkov, Alfio Quarteroni, and Gianluigi Rozza. A mathematical approach in the design of arterial bypass using unsteady Stokes equations. Journal of Scientific Computing, 28(2–3):139–165, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9077-9;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9077-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=139-165.

Archibald:2006:OSP

[559] Rick Archibald, Anne Gelb, Sigal Gottlieb, and Jennifer Ryan. One-sided post-processing for the discontinuous Galerkin method using ENO type stencil choosing and the local edge detection method. Journal of Scientific Computing, 28(2–3):167–190, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9090-z;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9090-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=167-190.

Banks:2006:HPV

[560] H. T. Banks, V. A. Bokil, D. Cioranescu, and N. L. Gibson. Homogenization of periodically varying coefficients in electromagnetic materials. Journal of Scientific Computing, 28(2–3):191–221, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9091-y;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9091-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=191-221.

Canuto:2006:ECP

[561] Claudio Canuto. Enhanced coercivity for pure advection and advection– diffusion problems. Journal of Scientific Computing, 28(2–3):223– 244, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9081-0;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9081-0;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=2&
spage=223-244.

Chorin:2006:PPD

[562] Alexandre J. Chorin, Ole H. Hald, and Raz Kupferman. Prediction from partial data, renormalization, and averaging. Journal of Scientific Computing, 28(2–3):245–261, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9089-5;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9089-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=28&issue=2&spage=245-261.

Cho:2006:BIE

[563] Min Hyung Cho, Wei Cai, and Tsing-Hua Her. A boundary integral equation method for photonic crystal fibers. Journal of Scientific Computing, 28(2–3):263–278, September 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9080-1;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9080-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=28&issue=2&spage=263-278.

Gelb:2006:AED

[564] A. Gelb and E. Tadmor. Adaptive edge detectors for piecewise smooth data based on the minmod limiter. Journal of Scientific Computing, 28(2–3):279–306, September 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9088-6;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9088-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=28&issue=2&spage=279-306.

Gottlieb:2006:RHO

[565] Sigal Gottlieb, David Gottlieb, and Chi-Wang Shu. Recovering highorder accuracy in WENO computations of steady-state hyperbolic systems. Journal of Scientific Computing, 28(2–3):307–318, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9078-8;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9078-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=307-318.

Gustafsson:2006:SBS

[566] Bertil Gustafsson and Yaser Khalighi. The shifted box scheme for scalar transport problems. Journal of Scientific Computing, 28(2–3): 319–335, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9079-7;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9079-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=2&
spage=319-335.

Hesthaven:2006:PLI

[567] J. S. Hesthaven, S. M. Kaber, and L. Lurati. Padé–Legendre interpolants for Gibbs reconstruction. Journal of Scientific Computing, 28(2– 3):337–359, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9085-9;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9085-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=2&
spage=337-359.

Jung:2006:PBL

[568] Chang-Yeol Jung and Roger Temam. On parabolic boundary layers for convection–diffusion equations in a channel: Analysis and numerical applications. Journal of Scientific Computing, 28(2–3):361–410, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9086-8;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9086-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=361-410.

Kindermann:2006:DBD

[569] Stefan Kindermann, Stanley Osher, and Jinjun Xu. Denoising by BVduality. Journal of Scientific Computing, 28(2–3):411–444, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9074-z;   http://link.springer.com/content/pdf/10.1007/

s10915-006-9074-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=411-444.

Lax:2006:GP

[570] Peter D. Lax. Gibbs phenomena. Journal of Scientific Computing, 28(2– 3):445–449, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9075-y;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9075-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=2&
spage=445-449.

Maday:2006:SAS

[571] Yvon Maday. L^{∞}-stable approximation of a solution to Div(Y) = f for f ∈ L^{2} in two dimensions. Journal of Scientific Computing, 28(2–3): 451–458, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9073-0;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9073-0;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=2&
spage=451-458.

Orszag:2006:TEK

[572] Steven A. Orszag, Hudong Chen, Sauro Succi, and Jonas Latt. Turbulence effects on kinetic equations. Journal of Scientific Computing, 28(2– 3):459–466, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9076-x;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9076-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=28&issue=2&
spage=459-466.

Ryabenkii:2006:MDP

[573] V. Ryaben’kii. On the method of difference potentials. Journal of Scientific Computing, 28(2–3):467–478, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9084-x;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9084-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=28&issue=2&spage=467-478.

Salas:2006:DFL

[574] M. D. Salas. Digital flight: The last CFD aeronautical grand challenge. Journal of Scientific Computing, 28(2–3):479–505, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9087-7;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9087-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=479-505.

Shang:2006:SPM

[575] J. S. Shang. Simulating plasma microwave diagnostics. Journal of Scientific Computing, 28(2–3):507–532, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9092-x;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9092-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=28&issue=2&spage=507-532.

Turek:2006:NSM

[576] Stefan Turek, Ludmila Rivkind, and Jaroslav Hron. Numerical study of a modified time-stepping θ-scheme for incompressible flow simulations. Journal of Scientific Computing, 28(2–3):533–547, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-006-9083-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=28&issue=2&spage=533-547.

Turkel:2006:NMN

[577] E. Turkel. Numerical methods and nature. Journal of Scientific Computing, 28(2–3):549–570, September 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-006-9082-z;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=28&issue=2&spage=549-570.

Boyd:2006:AFC

[578] John P. Boyd. Asymptotic Fourier coefficients for a C_{∞} bell (smoothed-“top-hat”) & the Fourier extension problem. Journal of Scientific Computing, 29(1):1–24, October 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9010-7;

 http://link.springer.com/content/pdf/10.1007/s10915-005-9010-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=29&issue=1&spage=1-24.

Zhang:2006:HOF

[579] Yong-Tao Zhang, Hong-Kai Zhao, and Jianliang Qian. High order fast sweeping methods for static Hamilton–Jacobi equations. Journal of Scientific Computing, 29(1):25–56, October 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9014-3;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9014-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=29&issue=1&spage=25-56.

Serna:2006:FOW

[580] Susana Serna and Jianliang Qian. Fifth-order weighted power–ENO schemes for Hamilton–Jacobi equations. Journal of Scientific Computing, 29(1):57–81, October 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9015-2;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9015-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=29&issue=1&
spage=57-81.

Evje:2006:CVN

[581] Steinar Evje and Tore Flåtten. CFL-violating numerical schemes for a two-fluid model. Journal of Scientific Computing, 29(1):83–114, October 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9000-9;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9000-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=29&issue=1&spage=83-114.

Yee:2006:ELD

[582] H. C. Yee and Björn Sjögreen. Efficient low dissipative high order schemes for multiscale MHD flows, II: Minimization of Δ · B numerical error. Journal of Scientific Computing, 29(1):115–164, October 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9004-5;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9004-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=29&issue=1&spage=115-164.

Morgan:2006:POG

[583] P. Morgan, M. Visbal, and D. Rizzetta. A parallel overset grid high-order flow solver for large eddy simulation. Journal of Scientific Computing, 29 (2):165–200, November 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9011-6;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9011-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=29&issue=2&
spage=165-200.

Suhov:2006:SMT

[584] A. Y. Suhov. A spectral method for the time evolution in parabolic problems. Journal of Scientific Computing, 29(2):201–217, November 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9001-8;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9001-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=29&issue=2&spage=201-217.

Bertoluzza:2006:IMM

[585] Silvia Bertoluzza, Silvia Falletta, and Valérie Perrier. Implementation of the mortar method in the wavelet context. Journal of Scientific Computing, 29(2):219–255, November 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9002-7;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9002-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=29&issue=2&spage=219-255.

Goujot:2006:BWC

[586] Daniel Goujot. Blind wavelet compression of the solution of a nonlinear PDE with singular forcing term within optimal order cost: Stability of restricted approximation to small errors. Journal of Scientific Computing, 29(2):257–297, November 2006. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9003-6;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9003-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=29&issue=2&spage=257-297.

Chen:2006:PSF

[587] Qian-Yong Chen. Partitions for spectral (finite) volume reconstruction in the tetrahedron. Journal of Scientific Computing, 29(3):299–319, December 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9009-0;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9009-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=29&issue=3&spage=299-319.

Greer:2006:IRE

[588] John B. Greer. An improvement of a recent Eulerian method for solving PDEs on general geometries. Journal of Scientific Computing, 29(3): 321–352, December 2006. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9012-5;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9012-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=29&issue=3&
spage=321-352.

Liu:2006:LSF

[589] Hailiang Liu, Li-Tien Cheng, and Stanley Osher. A level set framework for capturing multi-valued solutions of nonlinear first-order equations. Journal of Scientific Computing, 29(3):353–373, December 2006. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9016-1;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9016-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=29&issue=3&spage=353-373.

Nordstrom:2006:CFD

[590] Jan Nordström. Conservative finite difference formulations, variable coefficients, energy estimates and artificial dissipation. Journal of Scientific Computing, 29(3):375–404, December 2006. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9013-4;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9013-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=29&issue=3&spage=375-404.

Bokanowski:2007:ADS

[591] Olivier Bokanowski and Hasnaa Zidani. Anti-dissipative schemes for advection and application to Hamilton–Jacobi–Bellmann equations. Journal of Scientific Computing, 30(1):1–33, January 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9017-0;

 http://link.springer.com/content/pdf/10.1007/s10915-005-9017-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=1&spage=1-33.

Branden:2007:DFS

[592] Henrik Brandén, Sverker Holmgren, and Per Sundqvist. Discrete fundamental solution preconditioning for hyperbolic systems of PDE. Journal of Scientific Computing, 30(1):35–60, January 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9018-z;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9018-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=1&spage=35-60.

Li:2007:FDM

[593] Zhilin Li, C. V. Pao, and Zhonghua Qiao. A finite difference method and analysis for 2D nonlinear Poisson–Boltzmann equations. Journal of Scientific Computing, 30(1):61–81, January 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9019-y;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9019-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=1&spage=61-81.

Amundsen:2007:TSO

[594] David E. Amundsen and Oscar Bruno. Time stepping via onedimensional Padé approximation. Journal of Scientific Computing, 30 (1):83–115, January 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9021-4;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9021-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=30&issue=1&
spage=83-115.

Chen:2007:ECC

[595] Jinhai Chen and Weiguo Li. Equivalent conditions for convergence of splittings of non-Hermitian indefinite matrices. Journal of Scientific Computing, 30(1):117–130, January 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9022-3;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9022-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=1&spage=117-130.

Abgrall:2007:RDS

[596] R. Abgrall and F. Marpeau. Residual distribution schemes on quadrilateral meshes. Journal of Scientific Computing, 30(1):131–175, January 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9023-2;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9023-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=1&spage=131-175.

Jacobs:2007:CIW

[597] G. B. Jacobs, D. A. Kopriva, and F. Mashayek. A conservative isothermal Wall boundary condition for the compressible Navier–Stokes equations. Journal of Scientific Computing, 30(2):177–192, February 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
005-9040-1;   http://link.springer.com/content/pdf/10.1007/
s10915-005-9040-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=2&spage=177-192.

Artebrant:2007:TOA

[598] Robert Artebrant. Third order accurate non-polynomial reconstruction on rectangular and triangular meshes. Journal of Scientific Computing, 30(2):193–221, February 2007. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-005-9063-7;
 http://link.springer.com/content/pdf/10.1007/s10915-005-9063-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=2&spage=193-221.

Hausser:2007:DSP

[599] Frank Haußer and Axel Voigt. A discrete scheme for parametric anisotropic surface diffusion. Journal of Scientific Computing, 30(2): 223–235, February 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9064-6;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9064-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=30&issue=2&
spage=223-235.

Gobbert:2007:GMS

[600] Matthias K. Gobbert, Samuel G. Webster, and Timothy S. Cale. A Galerkin method for the simulation of the transient 2-D/2-D and 3-D/ 3-D linear Boltzmann equation. Journal of Scientific Computing, 30 (2):237–273, February 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-005-9069-1;   http://link.springer.com/content/
pdf/10.1007/s10915-005-9069-1;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=30&issue=2&
spage=237-273.

Belhachmi:2007:SME

[601] Zakaria Belhachmi and Andreas Karageorghis. A spectral mortar element discretization of the Poisson equation with mixed boundary conditions. Journal of Scientific Computing, 30(2):275–299, February 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9095-7;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9095-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=2&spage=275-299.

Hui:2007:PEU

[602] W. H. Hui, Z. N. Wu, and B. Gao. Preliminary extension of the unified coordinate system approach to computation of viscous flows. Journal of Scientific Computing, 30(2):301–344, February 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9096-6;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9096-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=2&spage=301-344.

Qiu:2007:NCL

[603] Jianxian Qiu. A numerical comparison of the Lax–Wendroff discontinuous Galerkin method based on different numerical fluxes. Journal of Scientific Computing, 30(3):345–367, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9109-5;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9109-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=3&spage=345-367.

Tosi:2007:OSE

[604] Francesca Tosi, Stefano Ubertini, S. Succi, and I. V. Karlin. Optimization strategies for the entropic lattice Boltzmann method. Journal of Scientific Computing, 30(3):369–387, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9097-5;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9097-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=3&spage=369-387.

Sanyasiraju:2007:FOS

[605] Y. V. S. S. Sanyasiraju and V. Manjula. Fourth-order semi-compact scheme for flow past a rotating and translating cylinder. Journal of Scientific Computing, 30(3):389–407, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9098-4;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9098-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=3&spage=389-407.

Gelb:2007:RPS

[606] Anne Gelb. Reconstruction of piecewise smooth functions from nonuniform grid point data. Journal of Scientific Computing, 30(3):409–440, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9099-3;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9099-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=3&spage=409-440.

Taube:2007:AHO

[607] Arne Taube, Michael Dumbser, and Dinshaw S. Balsara. Arbitrary high-order discontinuous Galerkin schemes for the magnetohydrodynamic equations. Journal of Scientific Computing, 30(3):441–464, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9101-0;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9101-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=3&spage=441-464.

Mozolevski:2007:VPE

[608] Igor Mozolevski, Endre Süli, and Paulo R. Bösing. hp-version a priori error analysis of interior penalty discontinuous Galerkin finite element approximations to the biharmonic equation. Journal of Scientific Computing, 30(3):465–491, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9100-1;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9100-

1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=3&spage=465-491.

Muller:2007:FAM

[609] Siegfried Müller and Youssef Stiriba. Fully adaptive multiscale schemes for conservation laws employing locally varying time stepping. Journal of Scientific Computing, 30(3):493–531, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9102-z;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9102-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=30&issue=3&spage=493-531.

Taylor:2007:SAP

[610] Ellen M. Taylor and M. Pino Martín. Stencil adaptation properties of a WENO scheme in direct numerical simulations of compressible turbulence. Journal of Scientific Computing, 30(3):533–554, March 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9126-4;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9126-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=30&issue=3&spage=533-554.

Osher:2007:P

[611] Stanley Osher. Preface. Journal of Scientific Computing, 31(1–2):1–3, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9125-5;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9125-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=1&spage=1-3.

Cafferelli:2007:OPP

[612] L. A. Cafferelli and Fang Hua Lin. An optimal partition problem for eigenvalues. Journal of Scientific Computing, 31(1–2):5–18, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9114-8;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9114-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=1&spage=5-18.

Chen:2007:SCF

[613] Han Chen, Chohong Min, and Frédéric Gibou. A supra-convergent finite difference scheme for the Poisson and heat equations on irregular domains and non-graded adaptive Cartesian grids. Journal of Scientific Computing, 31(1–2):19–60, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9122-8;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9122-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=31&issue=1&spage=19-60.

Cockburn:2007:NDG

[614] Bernardo Cockburn, Guido Kanschat, and Dominik Schötzau. A note on discontinuous Galerkin divergence-free solutions of the Navier– Stokes equations. Journal of Scientific Computing, 31(1–2):61–73, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9107-7;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9107-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=1&spage=61-73.

Di:2007:LSC

[615] Yana Di, Ruo Li, Tao Tang, and Pingwen Zhang. Level set calculations for incompressible two-phase flows on a dynamically adaptive grid. Journal of Scientific Computing, 31(1–2):75–98, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9119-3;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9119-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=1&spage=75-98.

Hong:2007:BCC

[616] Jeong-Mo Hong, Tamar Shinar, and Myungjoo Kang. On boundary condition capturing for multiphase interfaces. Journal of Scientific Computing, 31(1–2):99–125, May 2007. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9120-x;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9120-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=31&issue=1&spage=99-125.

Hou:2007:SMD

[617] Songming Hou and Xu-Dong Liu. Solutions of multi-dimensional hyperbolic systems of conservation laws by square entropy condition satisfying discontinuous Galerkin method. Journal of Scientific Computing, 31 (1–2):127–151, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9105-9;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9105-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=31&issue=1&
spage=127-151.

Kang:2007:LSB

[618] Myungjoo Kang, Hyeseon Shim, and Stanley Osher. Level set based simulations of two-phase oil–water flows in pipes. Journal of Scientific Computing, 31(1–2):153–184, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9103-y;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9103-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=31&issue=1&spage=153-184.

Lax:2007:CFD

[619] Peter D. Lax. Computational fluid dynamics. Journal of Scientific Computing, 31(1–2):185–193, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-006-9104-x;   http://link.springer.com/
content/pdf/10.1007/s10915-006-9104-x;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=31&issue=
1&spage=185-193.

Lin:2007:CES

[620] Chi-Tien Lin and Xu-Dong Liu. Convex ENO schemes for Hamilton– Jacobi equations. Journal of Scientific Computing, 31(1–2):195–211, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9121-9;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9121-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=1&spage=195-211.

Liu:2007:CFT

[621] Jinjie Liu, Hyun-Kyung Lim, James Glimm, and Xiaolin Li. A conservative front tracking method in N-dimensions. Journal of Scientific Computing, 31(1–2):213–236, May 2007. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9117-5;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9117-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=31&issue=1&spage=213-236.

Qian:2007:FSM

[622] Jianliang Qian, Yong-Tao Zhang, and Hong-Kai Zhao. A fast sweeping method for static convex Hamilton–Jacobi equations. Journal of Scientific Computing, 31(1–2):237–271, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9124-6;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9124-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=31&issue=1&spage=237-271.

Zhang:2007:NSI

[623] Shuhai Zhang and Chi-Wang Shu. A new smoothness indicator for the WENO schemes and its effect on the convergence to steady state solutions. Journal of Scientific Computing, 31(1–2):273–305, May 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9111-y;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9111-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=1&spage=273-305.

Boulmezaoud:2007:ESS

[624] T. Z. Boulmezaoud and J. M. Urquiza. On the eigenvalues of the spectral second order differentiation operator and application to the boundary observability of the wave equation. Journal of Scientific Computing, 31(3):307–345, June 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-006-9106-8;   http://link.springer.com/content/
pdf/10.1007/s10915-006-9106-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=31&issue=3&
spage=307-345.

Berthon:2007:HSS

[625] Christophe Berthon, Pierre Charrier, and Bruno Dubroca. An HLLC scheme to solve the M_{1} model of radiative transfer in two space dimensions. Journal of Scientific Computing, 31(3):347–389, June 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9108-6;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9108-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=3&spage=347-389.

Sidi:2007:NCS

[626] Avram Sidi. A novel class of symmetric and nonsymmetric periodizing variable transformations for numerical integration. Journal of Scientific Computing, 31(3):391–417, June 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9110-z;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9110-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=31&issue=3&spage=391-417.

Alouges:2007:TDE

[627] François Alouges, Gérard Le Coq, and Emmanuel Lorin. Twodimensional extension of the reservoir technique for some linear advection systems. Journal of Scientific Computing, 31(3):419–458, June 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9115-7;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9115-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=31&issue=3&spage=419-458.

Pieraccini:2007:IES

[628] Sandra Pieraccini and Gabriella Puppo. Implicit-explicit schemes for BGK kinetic equations. Journal of Scientific Computing, 32(1):1–28, July 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9116-6;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9116-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=1&spage=1-28.

Ditkowski:2007:ESU

[629] Adi Ditkowski, Gadi Fibich, and Nir Gavish. Efficient solution of A, x^{(k)} = b^{(k)} using A^{−1}. Journal of Scientific Computing, 32(1):29–44, July 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9112-x;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9112-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=1&spage=29-44.

Wang:2007:SDM

[630] Z. J. Wang, Yen Liu, Georg May, and Antony Jameson. Spectral difference method for unstructured grids II: Extension to the Euler equations. Journal of Scientific Computing, 32(1):45–71, July 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9113-9;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9113-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=1&spage=45-71.

Kloucek:2007:ACB

[631] Petr Kloucek, Danny C. Sorensen, and Jennifer L. Wightman. The approximation and computation of a basis of the trace space H^{1/2}. Journal of Scientific Computing, 32(1):73–108, July 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9118-4;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9118-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=32&issue=1&spage=73-108.

Diener:2007:OHO

[632] Peter Diener, Ernst Nils Dorband, and Erik Schnetter. Optimized high-order derivative and dissipation operators satisfying summation by parts, and applications in three-dimensional multi-block evolutions. Journal of Scientific Computing, 32(1):109–145, July 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
006-9123-7;   http://link.springer.com/content/pdf/10.1007/
s10915-006-9123-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=1&spage=109-145.

Ohmori:2007:FFF

[633] Katsushi Ohmori and Norikazu Saito. Flux-free finite element method with Lagrange multipliers for two-fluid flows. Journal of Scientific Computing, 32(2):147–173, August 2007. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-006-9127-3;
 http://link.springer.com/content/pdf/10.1007/s10915-006-9127-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=32&issue=2&spage=147-173.

Lorcher:2007:DGS

[634] F. Lörcher, G. Gassner, and C.-D. Munz. A discontinuous Galerkin scheme based on a space–time expansion. I. Inviscid compressible flow in one space dimension. Journal of Scientific Computing, 32(2): 175–199, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9128-x;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9128-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=32&issue=2&
spage=175-199.

Cockburn:2007:ARS

[635] Bernardo Cockburn and Ryuhei Ichikawa. Adjoint recovery of superconvergent linear functionals from Galerkin approximations. The onedimensional case. Journal of Scientific Computing, 32(2):201–232, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9129-9;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9129-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=2&spage=201-232.

Cockburn:2007:AMD

[636] Bernardo Cockburn and Bo Dong. An analysis of the minimal dissipation local discontinuous Galerkin method for convection–diffusion problems. Journal of Scientific Computing, 32(2):233–262, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9130-3;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9130-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=2&spage=233-262.

Li:2007:FEA

[637] Jichun Li and Aihua Wood. Finite element analysis for wave propagation in double negative metamaterials. Journal of Scientific Computing, 32(2):263–286, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9131-2;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9131-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=32&issue=2&
spage=263-286.

Martin-Vaquero:2007:ABA

[638] J. Martín-Vaquero and J. Vigo-Aguiar. Adapted BDF algorithms: Higher-order methods and their stability. Journal of Scientific Computing, 32(2):287–313, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-007-9132-1;   http://link.springer.com/
content/pdf/10.1007/s10915-007-9132-1;   http://www.springerlink.

com/openurl.asp?genre=article&issn=0885-7474&volume=32&issue=
2&spage=287-313.

Chan:2007:TVW

[639] Tony F. Chan and Hao-Min Zhou. Total variation wavelet thresholding. Journal of Scientific Computing, 32(2):315–341, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9133-0;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9133-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=2&spage=315-341.

Huang:2007:PDA

[640] Y. Q. Huang, Ruo Li, and Wenbin Liu. Preconditioned descent algorithms for p-Laplacian. Journal of Scientific Computing, 32(2): 343–371, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9134-z;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9134-z;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=32&issue=2&
spage=343-371.

Bialecki:2007:NDD

[641] Bernard Bialecki and Andreas Karageorghis. A nonoverlapping domain decomposition method for Legendre spectral collocation problems. Journal of Scientific Computing, 32(2):373–409, August 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9136-x;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9136-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=2&spage=373-409.

Kitzhofer:2007:ENSa

[642] G. Kitzhofer, O. Koch, P. Lima, and E. Weinmüller. Efficient numerical solution of the density profile equation in hydrodynamics. Journal of Scientific Computing, 32(3):411–424, September 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9141-0;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9141-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=3&spage=411-424.

Kitzhofer:2007:ENSb

[643] G. Kitzhofer, O. Koch, P. Lima, and E. Weinmüller. Efficient numerical solution of the density profile equation in hydrodynamics. Journal of Scientific Computing, 32(3):425, September 2007. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9142-z;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9142-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=32&issue=3&spage=425.

Black:2007:CVS

[644] Kelly Black and John B. Geddes. Complex valued spectral Hermite approximations for the actively mode-locked laser. Journal of Scientific Computing, 32(3):427–448, September 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9138-8;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9138-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=32&issue=3&spage=427-448.

Lambers:2007:PIK

[645] James V. Lambers. Practical implementation of Krylov subspace spectral methods. Journal of Scientific Computing, 32(3):449–476, September 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9140-1;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9140-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=32&issue=3&spage=449-476.

Ballereau:2007:SPL

[646] Ph. Ballereau, M. Casanova, F. Duboc, and D. Dureau. Simulation of the paraxial laser propagation coupled with hydrodynamics in 3D geometry. Journal of Scientific Computing, 33(1):1–24, October 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9135-y;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9135-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=33&issue=1&spage=1-24.

Wohlmuth:2007:PEE

[647] Barbara I. Wohlmuth. An a posteriori error estimator for two-body contact problems on non-matching meshes. Journal of Scientific Computing, 33(1):25–45, October 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9139-7;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9139-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=33&issue=1&
spage=25-45.

Sarmany:2007:DDE

[648] D. Sármány, M. A. Botchev, and J. J. W. van der Vegt. Dispersion and dissipation error in high-order Runge–Kutta discontinuous Galerkin discretisations of the Maxwell equations. Journal of Scientific Computing, 33(1):47–74, October 2007. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9143-y;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9143-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=33&issue=1&
spage=47-74.

Adjerid:2007:DGM

[649] Slimane Adjerid and Mahboub Baccouch. The discontinuous Galerkin method for two-dimensional hyperbolic problems. Part I: Superconvergence error analysis. Journal of Scientific Computing, 33(1):75–113, October 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9144-x;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9144-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=33&issue=1&spage=75-113.

Chen:2007:NMM

[650] Ke Chen and Xue-Cheng Tai. A nonlinear multigrid method for total variation minimization from image restoration. Journal of Scientific Computing, 33(2):115–138, November 2007. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9145-9;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9145-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=33&issue=2&spage=115-138.

Verwer:2007:TSW

[651] J. G. Verwer. On time staggering for wave equations. Journal of Scientific Computing, 33(2):139–154, November 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9146-8;

 http://link.springer.com/content/pdf/10.1007/s10915-007-9146-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=33&issue=2&spage=139-154.

Li:2007:PEE

[652] Ruo Li, Wenbin Liu, and Ningning Yan. A posteriori error estimates of recovery type for distributed convex optimal control problems. Journal of Scientific Computing, 33(2):155–182, November 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9147-7;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9147-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=33&issue=2&spage=155-182.

Burman:2007:MSD

[653] E. Burman and B. Stamm. Minimal stabilization for discontinuous Galerkin finite element methods for hyperbolic problems. Journal of Scientific Computing, 33(2):183–208, November 2007. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9149-5;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9149-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=33&issue=2&spage=183-208.

Kovalov:2007:PMA

[654] Pavlo Kovalov, Vadim Linetsky, and Michael Marcozzi. Pricing multiasset American options: a finite element method-of-lines with smooth penalty. Journal of Scientific Computing, 33(3):209–237, December 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9150-z;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9150-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=33&issue=3&spage=209-237.

Constantinescu:2007:MTM

[655] Emil M. Constantinescu and Adrian Sandu. Multirate timestepping methods for hyperbolic conservation laws. Journal of Scientific Computing, 33(3):239–278, December 2007. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9151-y;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9151-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=33&issue=3&spage=239-278.

Shen:2007:HOW

[656] Jun Shen, Chi-Wang Shu, and Mengping Zhang. A high order WENO scheme for a hierarchical size-structured population model. Journal of Scientific Computing, 33(3):279–291, December 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9152-x;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9152-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=33&issue=3&spage=279-291.

Alessandrini:2007:CVB

[657] Giovanni Alessandrini, Antonio Bilotta, and Antonino Morassi. Computing volume bounds of inclusions by EIT measurements. Journal of Scientific Computing, 33(3):293–312, December 2007. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9153-9;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9153-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=33&issue=3&spage=293-312.

Feng:2007:PPS

[658] Ko-An Feng, Chun-Hao Teng, and Min-Hung Chen. A pseudospectral penalty scheme for 2D isotropic elastic wave computations. Journal of Scientific Computing, 33(3):313–348, December 2007. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9154-8;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9154-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=33&issue=3&spage=313-348.

Arvanitis:2008:MRS

[659] Christos Arvanitis. Mesh redistribution strategies and finite element schemes for hyperbolic conservation laws. Journal of Scientific Computing, 34(1):1–25, January 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-007-9155-7;   http://link.springer.com/
content/pdf/10.1007/s10915-007-9155-7;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=34&issue=
1&spage=1-25.

Bernard:2008:DAD

[660] P.-E. Bernard, E. Deleersnijder, V. Legat, and J.-F. Remacle. Dispersion analysis of discontinuous Galerkin schemes applied to Poincaré, Kelvin and Rossby waves. Journal of Scientific Computing, 34(1):26–47, January 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9156-6;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9156-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=1&spage=26-47.

Yuan:2008:DPG

[661] Juan-Ming Yuan, Jie Shen, and Jiahong Wu. A dual-Petrov– Galerkin method for the Kawahara-type equations. Journal of Scientific Computing, 34(1):48–63, January 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9158-4;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9158-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=34&issue=1&spage=48-63.

Chen:2008:SOA

[662] Guoxian Chen, Huazhong Tang, and Pingwen Zhang. Second-order accurate Godunov scheme for multicomponent flows on moving triangular meshes. Journal of Scientific Computing, 34(1):64–86, January 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9162-8;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9162-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=1&spage=64-86.

Chen:2008:DRS

[663] Hudong Chen, Isaac Goldhirsch, and Steven A. Orszag. Discrete rotational symmetry, moment isotropy, and higher order lattice Boltzmann models. Journal of Scientific Computing, 34(1):87–112, January 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9159-3;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9159-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=1&spage=87-112.

Elbarbary:2008:ECP

[664] Elsayed M. E. Elbarbary. Efficient Chebyshev–Petrov–Galerkin method for solving second-order equations. Journal of Scientific Computing, 34 (2):113–126, February 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9161-9;   http://link.springer.com/content/

pdf/10.1007/s10915-007-9161-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=34&issue=2&
spage=113-126.

Ferm:2008:HAM

[665] Lars Ferm, Per Lötstedt, and Andreas Hellander. A hierarchy of approximations of the master equation scaled by a size parameter. Journal of Scientific Computing, 34(2):127–151, February 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9179-z;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9179-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=2&spage=127-151.

Jameson:2008:CDC

[666] Antony Jameson. The construction of discretely conservative finite volume schemes that also globally conserve energy or entropy. Journal of Scientific Computing, 34(2):152–187, February 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9171-7;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9171-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=2&spage=152-187.

Jameson:2008:FKE

[667] Antony Jameson. Formulation of kinetic energy preserving conservative schemes for gas dynamics and direct numerical simulation of onedimensional viscous compressible flow in a shock tube using entropy and kinetic energy preserving schemes. Journal of Scientific Computing, 34 (2):188–208, February 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9172-6;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9172-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=34&issue=2&
spage=188-208.

Almansa:2008:TBR

[668] A. Almansa, C. Ballester, V. Caselles, and G. Haro. A TV based restoration model with local constraints. Journal of Scientific Computing, 34 (3):209–236, March 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9160-x;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9160-x;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=34&issue=3&
spage=209-236.

Zhang:2008:SCS

[669] Zhimin Zhang. Superconvergence of a Chebyshev spectral collocation method. Journal of Scientific Computing, 34(3):237–246, March 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9163-7;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9163-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=3&spage=237-246.

Ha:2008:PSN

[670] Youngsoo Ha and Carl L. Gardner. Positive scheme numerical simulation of high Mach number astrophysical jets. Journal of Scientific Computing, 34(3):247–259, March 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9165-5;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9165-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=34&issue=3&
spage=247-259.

Gassner:2008:DGS

[671] G. Gassner, F. Lörcher, and C.-D. Munz. A discontinuous Galerkin scheme based on a space–time expansion II. Viscous flow equations in multi dimensions. Journal of Scientific Computing, 34(3):260–286, March 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9169-1;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9169-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=3&spage=260-286.

Marcozzi:2008:AID

[672] Michael D. Marcozzi. On the approximation of infinite dimensional optimal stopping problems with application to mathematical finance. Journal of Scientific Computing, 34(3):287–307, March 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9168-2;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9168-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=3&spage=287-307.

Nordsletten:2008:TOP

[673] D. Nordsletten and N. P. Smith. Triangulation of p-order parametric surfaces. Journal of Scientific Computing, 34(3):308–335, March 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9167-3;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9167-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=34&issue=3&spage=308-335.

Berkels:2008:EGB

[674] Benjamin Berkels, Andreas Rätz, Martin Rumpf, and Axel Voigt. Extracting grain boundaries and macroscopic deformations from images on atomic scale. Journal of Scientific Computing, 35(1):1–23, April 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9157-5;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9157-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=1&spage=1-23.

Cecil:2008:ENO

[675] Thomas C. Cecil, Stanley J. Osher, and Jianliang Qian. Essentially non-oscillatory adaptive tree methods. Journal of Scientific Computing, 35(1):25–41, April 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9164-6;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9164-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=35&issue=1&
spage=25-41.

Stewart:2008:ISI

[676] P. A. Stewart, N. Lay, M. Sussman, and M. Ohta. An improved sharp interface method for viscoelastic and viscous twophase flows. Journal of Scientific Computing, 35(1):43–61, April 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9173-5;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9173-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=1&spage=43-61.

Rzadkowski:2008:GEM

[677] Grzegorz Rzadkowski and Slawomir Lepkowski. A generalization of the Euler–Maclaurin summation formula: an application to numerical computation of the Fermi–Dirac integrals. Journal of Scientific Computing, 35(1):63–74, April 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9175-3;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9175-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=35&issue=1&
spage=63-74.

Osher:2008:P

[678] Stanley Osher. Preface. Journal of Scientific Computing, 35(2–3):75–76, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9213-9;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9213-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=2&spage=75-76.

Burger:2008:FEA

[679] Martin Burger, Christina Stöcker, and Axel Voigt. Finite element–ased level set methods for higher order flows. Journal of Scientific Computing, 35(2–3):77–98, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9204-x;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9204-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=35&issue=2&
spage=77-98.

Candela:2008:NAL

[680] Vicente F. Candela and Antonio Marquina. On the numerical approximation of the length of (implicit) level curves. Journal of Scientific Computing, 35(2–3):99–113, June 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9148-6;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9148-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=35&issue=2&spage=99-113.

duChene:2008:SOA

[681] Antoine du Chéné, Chohong Min, and Frédéric Gibou. Second-order accurate computation of curvatures in a level set framework using novel high-order reinitialization schemes. Journal of Scientific Computing, 35 (2–3):114–131, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9177-1;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9177-1;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=35&issue=2&
spage=114-131.

Cremers:2008:NDS

[682] Daniel Cremers. Nonlinear dynamical shape priors for level set segmentation. Journal of Scientific Computing, 35(2–3):132–143, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9220-x;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9220-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=2&spage=132-143.

Jin:2008:CSL

[683] Shi Jin and Xu Yang. Computation of the semiclassical limit of the Schrödinger equation with phase shift by a level set method. Journal of Scientific Computing, 35(2–3):144–169, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9137-9;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9137-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=35&issue=2&spage=144-169.

Kao:2008:PLS

[684] Chiu-Yen Kao and Richard Tsai. Properties of a level set algorithm for the visibility problems. Journal of Scientific Computing, 35(2– 3):170–191, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9197-5;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9197-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=35&issue=2&
spage=170-191.

Liu:2008:SMV

[685] Hailiang Liu and Zhongming Wang. Superposition of multi-valued solutions in high frequency wave dynamics. Journal of Scientific Computing, 35(2–3):192–218, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9198-4;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9198-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=35&issue=2&
spage=192-218.

Macdonald:2008:LSE

[686] Colin B. Macdonald and Steven J. Ruuth. Level set equations on surfaces via the closest point method. Journal of Scientific Computing, 35 (2–3):219–240, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9196-6;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9196-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=35&issue=2&
spage=219-240.

Xu:2008:DHF

[687] Yan Xu, Jaap J. W. van der Vegt, and Onno Bokhove. Discontinuous Hamiltonian finite element method for linear hyperbolic systems. Journal of Scientific Computing, 35(2–3):241–265, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9191-y;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9191-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=2&spage=241-265.

Macklin:2008:NGC

[688] Paul Macklin and John S. Lowengrub. A new ghost cell/level set method for moving boundary problems: Application to tumor growth. Journal of Scientific Computing, 35(2–3):266–299, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9190-z;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9190-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=2&spage=266-299.

Mitchell:2008:FEE

[689] Ian M. Mitchell. The flexible, extensible and efficient toolbox of level set methods. Journal of Scientific Computing, 35(2–3):300–329, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9174-4;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9174-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=2&spage=300-329.

Salac:2008:LSI

[690] David Salac and Wei Lu. A local semi-implicit level-set method for interface motion. Journal of Scientific Computing, 35(2–3):330–349, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9188-6;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9188-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=35&issue=2&spage=330-349.

Selle:2008:USM

[691] Andrew Selle, Ronald Fedkiw, ByungMoon Kim, and Yingjie Liu. An unconditionally stable MacCormack method. Journal of Scientific Computing, 35(2–3):350–371, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-007-9166-4;   http://link.springer.com/
content/pdf/10.1007/s10915-007-9166-4;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=35&issue=
2&spage=350-371.

Yokoi:2008:NMF

[692] Kensuke Yokoi. A numerical method for free-surface flows and its application to droplet impact on a thin liquid layer. Journal of Scientific Computing, 35(2–3):372–396, June 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9202-z;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9202-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=35&issue=2&spage=372-396.

Gelb:2008:DES

[693] Anne Gelb and Dennis Cates. Detection of edges in spectral data III– refinement of the concentration method. Journal of Scientific Computing, 36(1):1–43, July 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9170-8;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9170-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=36&issue=1&
spage=1-43.

Liu:2008:NSF

[694] T. G. Liu, J. Y. Ho, B. C. Khoo, and A. W. Chowdhury. Numerical simulation of fluid-structure interaction using modified ghost fluid method and Naviers equations. Journal of Scientific Computing, 36(1):45–68, July 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9176-2;   http://link.springer.com/content/pdf/10.1007/

s10915-007-9176-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=1&spage=45-68.

Chen:2008:BPC

[695] Chin-Yun Chen. Bivariate product cubature using Peano kernels for local error estimates. Journal of Scientific Computing, 36(1):69–88, July 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9178-0;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9178-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=1&spage=69-88.

Macdonald:2008:NSD

[696] Colin B. Macdonald, Sigal Gottlieb, and Steven J. Ruuth. A numerical study of diagonally split Runge–Kutta methods for PDEs with discontinuities. Journal of Scientific Computing, 36(1):89–112, July 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9180-6;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9180-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=1&spage=89-112.

Carrillo:2008:NSD

[697] J. A. Carrillo, T. Goudon, P. Lafitte, and F. Vecil. Numerical schemes of diffusion asymptotics and moment closures for kinetic equations. Journal of Scientific Computing, 36(1):113–149, July 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
007-9181-5;   http://link.springer.com/content/pdf/10.1007/
s10915-007-9181-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=1&spage=113-149.

Cockburn:2008:CLC

[698] Bernardo Cockburn and Haiying Wang. The computation of a locally conservative stress for the continuous Galerkin method for compressible linearly elastic materials. Journal of Scientific Computing, 36(2): 151–163, August 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9182-4;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9182-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=36&issue=2&
spage=151-163.

Loisel:2008:HMS

[699] Sébastien Loisel, Reinhard Nabben, and Daniel B. Szyld. On hybrid multigrid–Schwarz algorithms. Journal of Scientific Computing, 36(2): 165–175, August 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-007-9183-3;   http://link.springer.com/content/
pdf/10.1007/s10915-007-9183-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=36&issue=2&
spage=165-175.

Xiong:2008:RFV

[700] Zhiguang Xiong and Yanping Chen. A rectangular finite volume element method for a semilinear elliptic equation. Journal of Scientific Computing, 36(2):177–191, August 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-007-9184-2;
 http://link.springer.com/content/pdf/10.1007/s10915-007-9184-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=36&issue=2&spage=177-191.

Fernandez-Nieto:2008:EWT

[701] E. D. Fernández-Nieto and G. Narbona-Reina. Extension of WAF type methods to non-homogeneous shallow water equations with pollutant. Journal of Scientific Computing, 36(2):193–217, August 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9185-9;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9185-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=2&spage=193-217.

Pandit:2008:UCS

[702] Swapan K. Pandit. On the use of compact streamfunction–velocity formulation of steady Navier–Stokes equations on geometries beyond rectangular. Journal of Scientific Computing, 36(2):219–242, August 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9186-8;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9186-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=2&spage=219-242.

Han:2008:TFP

[703] Houde Han, Zhongyi Huang, and R. Bruce Kellogg. A tailored finite point method for a singular perturbation problem on an unbounded domain. Journal of Scientific Computing, 36(2):243–261, August 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9187-7;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9187-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=2&spage=243-261.

Zhang:2008:MFL

[704] Rong Zhang, Zhong qing Wang, and Ben yu Guo. Mixed Fourier– Laguerre spectral and pseudospectral methods for exterior problems using generalized Laguerre functions. Journal of Scientific Computing, 36 (2):263–283, August 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9189-5;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9189-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=36&issue=2&
spage=263-283.

Li:2008:NND

[705] Hongxia Li, Zhigang Wang, and De kang Mao. Numerically neither dissipative nor compressive scheme for linear advection equation and its application to the Euler system. Journal of Scientific Computing, 36(3): 285–331, September 2008. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9192-x;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9192-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=36&issue=3&
spage=285-331.

Ditkowski:2008:CDN

[706] Adi Ditkowski, Abhinav Bhandari, and Brian W. Sheldon. Computing derivatives of noisy signals using orthogonal functions expansions. Journal of Scientific Computing, 36(3):333–349, September 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9193-9;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9193-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=3&spage=333-349.

Teng:2008:LPP

[707] Chun-Hao Teng, Bang-Yan Lin, and Hung-Chun Chang. A Legendre pseudospectral penalty scheme for solving time-domain Maxwell’s equations. Journal of Scientific Computing, 36(3):351–390, September 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9194-8;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9194-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=3&spage=351-390.

Pennacchio:2008:SPM

[708] Micol Pennacchio and Valeria Simoncini. Substructuring preconditioners for mortar discretization of a degenerate evolution problem. Journal of Scientific Computing, 36(3):391–419, September 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9195-7;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9195-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=3&spage=391-419.

Awanou:2008:RSE

[709] Gerard Awanou. Robustness of a spline element method with constraints. Journal of Scientific Computing, 36(3):421–432, September 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9199-3;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9199-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=36&issue=3&spage=421-432.

Schulze:2008:MSM

[710] Tim Schulze, Vasilios Alexiades, and Xiaobing Feng. Multi-scale modeling and simulation in materials science, preface. Journal of Scientific Computing, 37(1):1–2, October 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9211-y;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9211-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=1&spage=1-2.

Caflisch:2008:GSP

[711] Russel E. Caflisch. Growth, structure and pattern formation for thin films. Journal of Scientific Computing, 37(1):3–17, October 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9206-8;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9206-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=1&spage=3-17.

Dobson:2008:ISQ

[712] Matthew Dobson and Mitchell Luskin. Iterative solution of the quasicontinuum equilibrium equations with continuation. Journal of Scientific Computing, 37(1):19–41, October 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9208-6;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9208-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=1&spage=19-41.

Katsoulakis:2008:NSM

[713] Markos A. Katsoulakis, Petr Plechác, and Luc Rey-Bellet. Numerical and statistical methods for the coarse-graining of many-particle stochastic systems. Journal of Scientific Computing, 37(1):43–71, October 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9216-6;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9216-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=1&spage=43-71.

Lam:2008:FKM

[714] Chi-Hang Lam, M. T. Lung, and Leonard M. Sander. Fast kinetic Monte Carlo simulation of strained heteroepitaxy in three dimensions. Journal of Scientific Computing, 37(1):73–88, October 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9205-9;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9205-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=1&spage=73-88.

Zhang:2008:MNA

[715] Lei Zhang, Long-Qing Chen, and Qiang Du. Mathematical and numerical aspects of a phase-field approach to critical nuclei morphology in solids. Journal of Scientific Computing, 37(1):89–102, October 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9207-7;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9207-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=1&spage=89-102.

Glowinski:2008:NSS

[716] R. Glowinski, T. Kärkkäinen, T. Valkonen, and A. Ivannikov. Nonsmooth SOR for L^{1}-fitting: Convergence study and discussion of related issues. Journal of Scientific Computing, 37(2):103–138, November 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9229-1;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9229-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=2&spage=103-138.

Gudi:2008:MDG

[717] Thirupathi Gudi, Neela Nataraj, and Amiya K. Pani. Mixed discontinuous Galerkin finite element method for the biharmonic equation. Journal of Scientific Computing, 37(2):139–161, November 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9200-1;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9200-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=2&spage=139-161.

VandenAbeele:2008:SAS

[718] Kris Van den Abeele, Chris Lacor, and Z. J. Wang. On the stability and accuracy of the spectral difference method. Journal of Scientific Computing, 37(2):162–188, November 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9201-0;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9201-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=2&spage=162-188.

Chertock:2008:MOP

[719] Alina Chertock, David Gottlieb, and Alex Solomonoff. Modified optimal prediction and its application to a particle-method problem. Journal of Scientific Computing, 37(2):189–201, November 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9242-4;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9242-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=2&spage=189-201.

Banas:2008:FEA

[720] Lubomír Banas and Robert Nürnberg. Finite element approximation of a three dimensional phase field model for void electromigration. Journal of Scientific Computing, 37(2):202–232, November 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

008-9203-y;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9203-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=2&spage=202-232.

Donat:2008:CBS

[721] Rosa Donat and Pep Mulet. Characteristic-based schemes for multiclass Lighthill–Whitham–Richards traffic models. Journal of Scientific Computing, 37(3):233–250, December 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9209-5;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9209-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=3&spage=233-250.

Cai:2008:EAS

[722] Haotao Cai. An efficient algorithm for solving the generalized airfoil equation. Journal of Scientific Computing, 37(3):251–267, December 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9241-5;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9241-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=3&spage=251-267.

Ditkowski:2008:NMS

[723] Adi Ditkowski. Numerical method for solving discontinuous initial/finalvalue problems. Journal of Scientific Computing, 37(3):268–281, December 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9243-3;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9243-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=37&issue=3&spage=268-281.

Brandman:2008:LSM

[724] Jeremy Brandman. A level-set method for computing the eigenvalues of elliptic operators defined on compact hypersurfaces. Journal of Scientific Computing, 37(3):282–315, December 2008. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9210-z;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9210-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=3&spage=282-315.

Parussini:2008:FDA

[725] Lucia Parussini. Fictitious domain approach via Lagrange multipliers with least squares spectral element method. Journal of Scientific Computing, 37(3):316–335, December 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9212-x;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9212-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=3&spage=316-335.

Achdou:2008:PEE

[726] Yves Achdou, Frédéric Hecht, and David Pommier. A posteriori error estimates for parabolic variational inequalities. Journal of Scientific Computing, 37(3):336–366, December 2008. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9215-7;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9215-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=3&spage=336-366.

Marquina:2008:ISR

[727] Antonio Marquina and Stanley J. Osher. Image super-resolution by TV-regularization and Bregman iteration. Journal of Scientific Computing, 37(3):367–382, December 2008. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9214-8;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9214-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=37&issue=3&spage=367-382.

Notsu:2009:SSC

[728] Hirofumi Notsu and Masahisa Tabata. A single-step characteristiccurve finite element scheme of second order in time for the incompressible Navier–Stokes equations. Journal of Scientific Computing, 38 (1):1–14, January 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9217-5;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9217-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=38&issue=1&
spage=1-14.

Adjerid:2009:DGM

[729] Slimane Adjerid and Mahboub Baccouch. The discontinuous Galerkin method for two-dimensional hyperbolic problems. Part II: a posteriori error estimation. Journal of Scientific Computing, 38(1):15–49, January 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9222-8;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9222-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=1&spage=15-49.

Juanes:2009:UFH

[730] Ruben Juanes and Sébastien F. Matringe. Unified formulation for high-order streamline tracing on two-dimensional unstructured grids. Journal of Scientific Computing, 38(1):50–73, January 2009. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9228-2;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9228-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=38&issue=1&spage=50-73.

Feng:2009:VMM

[731] Xiaobing Feng and Michael Neilan. Vanishing moment method and moment solutions for fully nonlinear second order partial differential equations. Journal of Scientific Computing, 38(1):74–98, January 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9221-9;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9221-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=1&spage=74-98.

Zunino:2009:DGM

[732] Paolo Zunino. Discontinuous Galerkin methods based on weighted interior penalties for second order PDEs with non-smooth coefficients. Journal of Scientific Computing, 38(1):99–126, January 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9219-3;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9219-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=1&spage=99-126.

Zhang:2009:NSP

[733] Qiang Zhang and Zi-Long Wu. Numerical simulation for porous medium equation by local discontinuous Galerkin finite element method. Journal of Scientific Computing, 38(2):127–148, February 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9223-7;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9223-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=2&spage=127-148.

Wang:2009:NSS

[734] Hanquan Wang. Numerical simulations on stationary states for rotating two-component Bose–Einstein condensates. Journal of Scientific Computing, 38(2):149–163, February 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9225-5;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9225-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=38&issue=2&spage=149-163.

Walfisch:2009:OSS

[735] David Walfisch, Jennifer K. Ryan, and Robert M. Kirby. Onesided smoothness-increasing accuracy-conserving filtering for enhanced streamline integration through discontinuous fields. Journal of Scientific Computing, 38(2):164–184, February 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9230-8;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9230-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=38&issue=2&spage=164-184.

Zhu:2009:RHO

[736] Qi Zhu, Ying Qiao, and Shaohua Tan. A robust high-order mixed L_{2}–L_{∞} estimation for linear-in-the-parameters models. Journal of Scientific Computing, 38(2):185–206, February 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9231-7;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9231-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=38&issue=2&spage=185-206.

Chniti:2009:IIC

[737] Chokri Chniti, Frédéric Nataf, and Francis Nier. Improved interface conditions for 2 D domain decomposition with corners: Numerical applications. Journal of Scientific Computing, 38(2):207–228, February 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9234-4;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9234-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=2&spage=207-228.

Sandu:2009:MEA

[738] Adrian Sandu and Emil M. Constantinescu. Multirate explicit Adams methods for time integration of conservation laws. Journal of Scientific Computing, 38(2):229–249, February 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9235-3;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9235-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=38&issue=2&spage=229-249.

Gottlieb:2009:HOS

[739] Sigal Gottlieb, David I. Ketcheson, and Chi-Wang Shu. High order strong stability preserving time discretizations. Journal of Scientific Computing, 38(3):251–289, March 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9239-z;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9239-z;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=38&issue=3&
spage=251-289.

Fu:2009:PEE

[740] Hongfei Fu and Hongxing Rui. A priori error estimates for optimal control problems governed by transient advection–diffusion equations. Journal of Scientific Computing, 38(3):290–315, March 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9224-6;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9224-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=3&spage=290-315.

Martinsson:2009:FDS

[741] Per-Gunnar Martinsson. A fast direct solver for a class of elliptic partial differential equations. Journal of Scientific Computing, 38(3):316–330, March 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9240-6;   http://link.springer.com/content/pdf/10.1007/

s10915-008-9240-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=3&spage=316-330.

Tone:2009:LTS

[742] Florentina Tone. On the long-time H^{2}-stability of the implicit Euler scheme for the 2D magnetohydrodynamics equations. Journal of Scientific Computing, 38(3):331–348, March 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9236-2;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9236-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=38&issue=3&spage=331-348.

Klar:2009:SLM

[743] Axel Klar, Philip Reuterswärd, and Mohammed Seaïd. A semi-Lagrangian method for a Fokker–Planck equation describing fiber dynamics. Journal of Scientific Computing, 38(3):349–367, March 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9244-2;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9244-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=38&issue=3&spage=349-367.

Glaser:2009:NCH

[744] Andreas Glaser and Vladimir Rokhlin. A new class of highly accurate solvers for ordinary differential equations. Journal of Scientific Computing, 38(3):368–399, March 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9245-1;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9245-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=38&issue=3&spage=368-399.

Cole:2009:SRB

[745] Daniel C. Cole and Yi Zou. Subharmonic resonance behavior for the classical hydrogen atomic system. Journal of Scientific Computing, 39(1):1–27, April 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9248-y;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9248-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=39&issue=1&
spage=1-27.

Bjontegaard:2009:FTP

[746] Tormod Bjøntegaard, Yvon Maday, and Einar M. Rønquist. Fast tensor-product solvers: Partially deformed three-dimensional domains. Journal of Scientific Computing, 39(1):28–48, April 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9246-0;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9246-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=39&issue=1&spage=28-48.

Jung:2009:NSC

[747] Jae-Hun Jung. A note on the spectral collocation approximation of some differential equations with singular source terms. Journal of Scientific Computing, 39(1):49–66, April 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9249-x;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9249-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=39&issue=1&spage=49-66.

Castro:2009:HOE

[748] M. J. Castro, E. D. Fernández-Nieto, and A. M. Ferreiro. High order extensions of Roe schemes for two-dimensional nonconservative hyperbolic systems. Journal of Scientific Computing, 39(1):67–114, April 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9250-4;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9250-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=1&spage=67-114.

Higueras:2009:CSS

[749] Inmaculada Higueras. Characterizing strong stability preserving additive Runge–Kutta methods. Journal of Scientific Computing, 39(1):115–128, April 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9252-2;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9252-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=1&spage=115-128.

Liu:2009:OYF

[750] Jinjie Liu, Moysey Brio, and Jerome V. Moloney. Overlapping Yee FDTD method on nonorthogonal grids. Journal of Scientific Computing, 39(1):129–143, April 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9253-1;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9253-1;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=39&issue=1&
spage=129-143.

Dooley:2009:PSD

[751] Isaac Dooley, Sandhya Mangala, and Laxmikant Kale. Parallel simulations of dynamic fracture using extrinsic cohesive elements. Journal of Scientific Computing, 39(1):144–165, April 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9254-0;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9254-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=1&spage=144-165.

Vandekerckhove:2009:ENK

[752] Christophe Vandekerckhove and Ioannis Kevrekidis. An efficient Newton–Krylov implementation of the constrained runs scheme for initializing on a slow manifold. Journal of Scientific Computing, 39 (2):167–188, May 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9256-y;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9256-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=39&issue=2&
spage=167-188.

Cristiani:2009:FMM

[753] Emiliano Cristiani. A fast marching method for Hamilton–Jacobi equations modeling monotone front propagations. Journal of Scientific Computing, 39(2):189–205, May 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-008-9257-x;   http://link.springer.com/
content/pdf/10.1007/s10915-008-9257-x;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=39&issue=
2&spage=189-205.

Chen:2009:SOC

[754] Yanping Chen and Yongquan Dai. Superconvergence for optimal control problems governed by semi-linear elliptic equations. Journal of Scientific Computing, 39(2):206–221, May 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9258-9;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9258-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=39&issue=2&spage=206-221.

Zhu:2009:SEM

[755] Wuming Zhu and David A. Kopriva. A spectral element method to price European options. I. Single asset with and without jump diffusion. Journal of Scientific Computing, 39(2):222–243, May 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9267-8;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9267-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=2&spage=222-243.

Platte:2009:HFC

[756] Rodrigo B. Platte and Anne Gelb. A hybrid Fourier–Chebyshev method for partial differential equations. Journal of Scientific Computing, 39 (2):244–264, May 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9264-y;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9264-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=39&issue=2&
spage=244-264.

Binford:2009:ENR

[757] Tommy L. Binford, Jr., David P. Nicholls, and Nilima Nigam. Exact non-reflecting boundary conditions on perturbed domains and hpfinite elements. Journal of Scientific Computing, 39(2):265–292, May 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9263-z;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9263-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=2&spage=265-292.

Zhu:2009:HWS

[758] Jun Zhu and Jianxian Qiu. Hermite WENO schemes and their application as limiters for Runge–Kutta discontinuous Galerkin method, III: Unstructured meshes. Journal of Scientific Computing, 39(2):293–321, May 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9271-7;   http://link.springer.com/content/pdf/10.1007/

s10915-009-9271-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=2&spage=293-321.

Zhu:2009:SEA

[759] Wuming Zhu and David A. Kopriva. A spectral element approximation to price European options. II. The Black–Scholes model with two underlying assets. Journal of Scientific Computing, 39(3):323–339, June 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9270-8;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9270-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=3&spage=323-339.

Tang:2009:CTL

[760] J. M. Tang, R. Nabben, C. Vuik, and Y. A. Erlangga. Comparison of two-level preconditioners derived from deflation, domain decomposition and multigrid methods. Journal of Scientific Computing, 39(3):340–370, June 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9272-6;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9272-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=3&spage=340-370.

El-Amrani:2009:SSS

[761] Mofdi El-Amrani and Mohammed Seaïd. A spectral stochastic semi-Lagrangian method for convection–diffusion equations with uncertainty. Journal of Scientific Computing, 39(3):371–393, June 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9273-5;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9273-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=3&spage=371-393.

Ditkowski:2009:HOE

[762] Adi Ditkowski and Yuval Harness. High-order embedded finite difference schemes for initial boundary value problems on time dependent irregular domains. Journal of Scientific Computing, 39(3):394–440, June 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9277-1;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9277-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=39&issue=3&spage=394-440.

Di:2009:CDG

[763] Yana Di and Ruo Li. Computation of dendritic growth with level set model using a multi-mesh adaptive finite element method. Journal of Scientific Computing, 39(3):441–453, June 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9275-3;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9275-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=39&issue=3&spage=441-453.

Svard:2009:SCA

[764] Magnus Svärd and Siddhartha Mishra. Shock capturing artificial dissipation for high-order finite difference schemes. Journal of Scientific Computing, 39(3):454–484, June 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9285-1;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9285-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=39&issue=3&spage=454-484.

Cockburn:2009:F

[765] Bernardo Cockburn and Chi-Wang Shu. Foreword. Journal of Scientific Computing, 40(1–3):1–3, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-009-9298-9;   http://link.springer.com/
content/pdf/10.1007/s10915-009-9298-9;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=40&issue=
1&spage=1-3.

deDios:2009:UCI

[766] Blanca Ayuso de Dios and Ludmil Zikatanov. Uniformly convergent iterative methods for discontinuous Galerkin discretizations. Journal of Scientific Computing, 40(1–3):4–36, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9293-1;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9293-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=40&issue=1&spage=4-36.

Brenner:2009:PEC

[767] Susanne C. Brenner, Thirupathi Gudi, and Li yeng Sung. A posteriori error control for a weakly over-penalized symmetric interior penalty method. Journal of Scientific Computing, 40(1–3):37–50, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9278-0;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9278-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=37-50.

Brenner:2009:NME

[768] Susanne C. Brenner, Fengyan Li, and Li yeng Sung. Nonconforming Maxwell eigensolvers. Journal of Scientific Computing, 40(1–3):51–85, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9266-9;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9266-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=51-85.

Buffa:2009:MDG

[769] Annalisa Buffa, Ilaria Perugia, and Tim Warburton. The mortardiscontinuous Galerkin method for the 2D Maxwell eigenproblem. Journal of Scientific Computing, 40(1–3):86–114, July 2009. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9238-0;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9238-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=40&issue=1&spage=86-114.

Cesmelioglu:2009:PDG

[770] Ayçil Çesmelioglu and Béatrice Rivière. Primal discontinuous Galerkin methods for time-dependent coupled surface and subsurface flow. Journal of Scientific Computing, 40(1–3):115–140, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9274-4;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9274-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=115-140.

Cockburn:2009:HSD

[771] Bernardo Cockburn, Bo Dong, and Johnny Guzmán. A hybridizable and superconvergent discontinuous Galerkin method for biharmonic problems. Journal of Scientific Computing, 40(1–3):141–187, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9279-z;   http://link.springer.com/content/pdf/10.1007/

s10915-009-9279-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=141-187.

Cockburn:2009:EOD

[772] Bernardo Cockburn, Guido Kanschat, and Dominik Schötzau. An equal-order DG method for the incompressible Navier–Stokes equations. Journal of Scientific Computing, 40(1–3):188–210, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9261-1;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9261-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=188-210.

Epshteyn:2009:FDA

[773] Yekaterina Epshteyn and Ahmet Izmirlioglu. Fully discrete analysis of a discontinuous finite element method for the Keller–Segel chemotaxis model. Journal of Scientific Computing, 40(1–3):211–256, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9281-5;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9281-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=211-256.

Grote:2009:OEE

[774] Marcus J. Grote and Dominik Schötzau. Optimal error estimates for the fully discrete interior penalty DG method for the wave equation. Journal of Scientific Computing, 40(1–3):257–272, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9247-z;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9247-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=257-272.

Guzman:2009:SOC

[775] Johnny Guzmán and Béatrice Rivière. Sub-optimal convergence of non-symmetric discontinuous Galerkin methods for odd polynomial approximations. Journal of Scientific Computing, 40(1–3):273–280, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9255-z;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9255-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=273-280.

Houston:2009:MDM

[776] Paul Houston, Dominik Schötzau, and Xiaoxi Wei. A mixed DG method for linearized incompressible magnetohydrodynamics. Journal of Scientific Computing, 40(1–3):281–314, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9265-x;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9265-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=40&issue=1&spage=281-314.

Kubatko:2009:PCC

[777] Ethan J. Kubatko, Shintaro Bunya, and Clint Dawson. A performance comparison of continuous and discontinuous finite element shallow water models. Journal of Scientific Computing, 40(1–3):315–339, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9268-2;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9268-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=315-339.

Lovadina:2009:PEE

[778] Carlo Lovadina and L. Donatella Marini. A-posteriori error estimates for discontinuous Galerkin approximations of second order elliptic problems. Journal of Scientific Computing, 40(1–3):340–359, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9286-0;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9286-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=340-359.

Wang:2009:WLD

[779] Wei Wang and Chi-Wang Shu. The WKB local discontinuous Galerkin method for the simulation of Schrödinger equation in a resonant tunneling diode. Journal of Scientific Computing, 40(1–3):360–374, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9237-1;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9237-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=360-374.

Xu:2009:LDG

[780] Yan Xu and Chi-Wang Shu. Local discontinuous Galerkin method for surface diffusion and Willmore flow of graphs. Journal of Scientific Computing, 40(1–3):375–390, July 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9262-0;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9262-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=40&issue=1&spage=375-390.

Zhu:2009:ADG

[781] Jianfeng Zhu, Yong-Tao Zhang, and Stuart A. Newman. Application of discontinuous Galerkin methods for reaction–diffusion systems in developmental biology. Journal of Scientific Computing, 40(1–3):391–418, July 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9218-4;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9218-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=40&issue=1&spage=391-418.

Abarbanel:2009:LTP

[782] S. Abarbanel, H. Qasimov, and S. Tsynkov. Long-time performance of unsplit PMLs with explicit second order schemes. Journal of Scientific Computing, 41(1):1–12, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9282-4;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9282-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=1&spage=1-12.

Jung:2009:EPD

[783] Chang-Yeol Jung. Evolution of probability distribution in time for solutions of hyperbolic equations. Journal of Scientific Computing, 41 (1):13–48, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9284-2;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9284-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=41&issue=1&
spage=13-48.

Sarra:2009:EDF

[784] Scott A. Sarra. Edge detection free postprocessing for pseudospectral approximations. Journal of Scientific Computing, 41(1):49–61, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9287-z;   http://link.springer.com/content/pdf/10.1007/

s10915-009-9287-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=1&spage=49-61.

Hughes:2009:BLA

[785] J. Hughes and M. Friedman. A bisection-like algorithm for branch switching at a simple branch point. Journal of Scientific Computing, 41(1):62–69, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9306-0;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9306-0;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=41&issue=1&
spage=62-69.

Zhang:2009:SDG

[786] Zuozheng Zhang, Ziqing Xie, and Zhimin Zhang. Superconvergence of discontinuous Galerkin methods for convection–diffusion problems. Journal of Scientific Computing, 41(1):70–93, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9288-y;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9288-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=1&spage=70-93.

Kupiainen:2009:CEB

[787] Marco Kupiainen and Björn Sjögreen. A Cartesian embedded boundary method for the compressible Navier–Stokes equations. Journal of Scientific Computing, 41(1):94–117, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9289-x;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9289-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=1&spage=94-117.

Rahunanthan:2009:SIC

[788] Arunasalam Rahunanthan and Dan Stanescu. Stable interface conditions for discontinuous Galerkin approximations of Navier–Stokes equations. Journal of Scientific Computing, 41(1):118–138, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9290-4;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9290-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=1&spage=118-138.

Celledoni:2009:SLR

[789] Elena Celledoni and Bawfeh Kingsley Kometa. Semi-Lagrangian Runge–Kutta exponential integrators for convection dominated problems. Journal of Scientific Computing, 41(1):139–164, October 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9291-3;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9291-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=1&spage=139-164.

Kannan:2009:SVF

[790] R. Kannan and Z. J. Wang. A study of viscous flux formulations for a fp-multigrid spectral volume Navier–Stokes solver. Journal of Scientific Computing, 41(2):165–199, November 2009. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9269-1;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9269-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=2&spage=165-199.

Han:2009:TFP

[791] Houde Han and Zhongyi Huang. Tailored finite point method for a singular perturbation problem with variable coefficients in two dimensions. Journal of Scientific Computing, 41(2):200–220, November 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9292-2;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9292-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=2&spage=200-220.

Delage-Santacreu:2009:TFO

[792] Stephanie Delage-Santacreu and Stephane Vincent. Tracking fronts in one and two-phase incompressible flows using an adaptive mesh refinement approach. Journal of Scientific Computing, 41(2):221–237, November 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9294-0;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9294-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=2&spage=221-237.

Ge:2009:AFE

[793] Liang Ge, Wenbin Liu, and Danping Yang. Adaptive finite element approximation for a constrained optimal control problem via multimeshes. Journal of Scientific Computing, 41(2):238–255, November 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9296-y;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9296-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=2&spage=238-255.

Chen:2009:FCM

[794] Zhongying Chen, Gnaneshwar Nelakanti, and Yuesheng Xu. A fast collocation method for eigen-problems of weakly singular integral operators. Journal of Scientific Computing, 41(2):256–272, November 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9295-z;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9295-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=2&spage=256-272.

Yan:2009:RMF

[795] Ningning Yan and Zhaojie Zhou. A RT mixed FEM/DG scheme for optimal control governed by convection diffusion equations. Journal of Scientific Computing, 41(2):273–299, November 2009. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9297-x;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9297-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=2&spage=273-299.

Ng:2009:GPS

[796] Yen Ting Ng, Han Chen, Chohong Min, and Frédéric Gibou. Guidelines for Poisson solvers on irregular domains with Dirichlet boundary conditions using the ghost fluid method. Journal of Scientific Computing, 41 (2):300–320, November 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9299-8;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9299-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=41&issue=2&
spage=300-320.

Huang:2009:IPD

[797] Yunqing Huang and Jichun Li. Interior penalty discontinuous Galerkin method for Maxwell’s equations in Cold plasma. Journal of Scientific Computing, 41(3):321–340, December 2009. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9300-6;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9300-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=3&spage=321-340.

Belaouar:2009:ASS

[798] R. Belaouar, N. Crouseilles, P. Degond, and E. Sonnendrücker. An asymptotically stable semi-Lagrangian scheme in the quasi-neutral limit. Journal of Scientific Computing, 41(3):341–365, December 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9302-4;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9302-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=3&spage=341-365.

Mattsson:2009:SBT

[799] Ken Mattsson, Frank Ham, and Gianluca Iaccarino. Stable boundary treatment for the wave equation on second-order form. Journal of Scientific Computing, 41(3):366–383, December 2009. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9305-1;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9305-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=3&spage=366-383.

Jung:2009:FVA

[800] Chang-Yeol Jung and Roger Temam. Finite volume approximation of one-dimensional stiff convection–diffusion equations. Journal of Scientific Computing, 41(3):384–410, December 2009. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9304-2;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9304-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=3&spage=384-410.

Xiang:2009:SOI

[801] Ming Xiang, Shaozhong Deng, and Wei Cai. A sixth-order image approximation to the ionic solvent induced reaction field. Journal of Scientific Computing, 41(3):411–435, December 2009. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9307-z;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9307-

z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=41&issue=3&spage=411-435.

Gao:2009:LDG

[802] Fuzheng Gao, Jianxian Qiu, and Qiang Zhang. Local discontinuous Galerkin finite element method and error estimates for one class of Sobolev equation. Journal of Scientific Computing, 41(3):436– 460, December 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9308-y;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9308-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=41&issue=3&
spage=436-460.

He:2009:EEC

[803] Li ping He. Error estimation of a class of stable spectral approximation to the Cahn–Hilliard equation. Journal of Scientific Computing, 41(3): 461–482, December 2009. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9309-x;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9309-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=41&issue=3&
spage=461-482.

Bouchut:2009:SDE

[804] F. Bouchut and T. Morales de Luna. Semi-discrete entropy satisfying approximate Riemann solvers. The case of the Suliciu relaxation approximation. Journal of Scientific Computing, 41(3):483–509, December 2009. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9311-3;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9311-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=41&issue=3&spage=483-509.

Zhuang:2010:CLL

[805] Qingqu Zhuang, Jie Shen, and Chuanju Xu. A coupled Legendre– Laguerre spectral-element method for the Navier–Stokes equations in unbounded domains. Journal of Scientific Computing, 42(1):1–22, January 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9313-1;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9313-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=1&spage=1-22.

deAbreu:2010:DDO

[806] Marcos Pimenta de Abreu. On the dependence of discrete ordinates models for layer reflectance and transmittance on relative optical depth and solar zenith angle. Journal of Scientific Computing, 42(1):23–37, January 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9312-2;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9312-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=1&spage=23-37.

Hu:2010:NML

[807] Sheng-Long Hu, Zheng-Hai Huang, and Nan Lu. A non-monotone line search algorithm for unconstrained optimization. Journal of Scientific Computing, 42(1):38–53, January 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9314-0;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9314-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=42&issue=1&spage=38-53.

Georgoulis:2010:SVI

[808] Emmanuil H. Georgoulis, Edward Hall, and Jens Markus Melenk. On the suboptimality of the p-version interior penalty discontinuous Galerkin method. Journal of Scientific Computing, 42(1):54–67, January 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9315-z;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9315-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=1&spage=54-67.

Romenski:2010:CMN

[809] Evgeniy Romenski, Dimitris Drikakis, and Eleuterio Toro. Conservative models and numerical methods for compressible two-phase flow. Journal of Scientific Computing, 42(1):68–95, January 2010. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9316-y;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9316-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=42&issue=1&spage=68-95.

Sescu:2010:CSM

[810] Adrian Sescu, Abdollah A. Afjeh, Ray Hixon, and Carmen Sescu. Conditionally stable multidimensional schemes for advective equations. Journal of Scientific Computing, 42(1):96–117, January 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9317-x;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9317-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=1&spage=96-117.

Brenner:2010:IPF

[811] S. C. Brenner, T. Gudi, L. Owens, and L.-Y. Sung. An intrinsically parallel finite element method. Journal of Scientific Computing, 42(1): 118–121, January 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9318-9;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9318-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=42&issue=1&
spage=118-121.

Dolejsi:2010:OIM

[812] Vít Dolejsí and Oto Havle. The L^{2}-optimality of the IIPG method for odd degrees of polynomial approximation in 1D. Journal of Scientific Computing, 42(1):122–143, January 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9319-8;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9319-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=42&issue=1&spage=122-143.

Cockburn:2010:BCD

[813] Bernardo Cockburn, Deepa Gupta, and Fernando Reitich. Boundaryconforming discontinuous Galerkin methods via extensions from subdomains. Journal of Scientific Computing, 42(1):144–184, January 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9321-1;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9321-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=1&spage=144-184.

Lou:2010:IRN

[814] Yifei Lou, Xiaoqun Zhang, Stanley Osher, and Andrea Bertozzi. Image recovery via nonlocal operators. Journal of Scientific Computing, 42 (2):185–197, February 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9320-2;   http://link.springer.com/content/

pdf/10.1007/s10915-009-9320-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=42&issue=2&
spage=185-197.

Shen:2010:OEE

[815] Ting-Ting Shen, Zhong-Qiang Zhang, and He-Ping Ma. Optimal error estimates of the Legendre Tau method for second-order differential equations. Journal of Scientific Computing, 42(2):198–215, February 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9323-z;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9323-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=2&spage=198-215.

Ben-Artzi:2010:HOC

[816] M. Ben-Artzi, J.-P. Croisille, and D. Fishelov. A high order compact scheme for the pure-streamfunction formulation of the Navier–Stokes equations. Journal of Scientific Computing, 42(2):216–250, February 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9322-0;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9322-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=2&spage=216-250.

Bokanowski:2010:EDS

[817] O. Bokanowski, E. Cristiani, and H. Zidani. An efficient data structure and accurate scheme to solve front propagation problems. Journal of Scientific Computing, 42(2):251–273, February 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9329-6;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9329-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=2&spage=251-273.

Kakleas:2010:NSW

[818] Maria Kakleas and David P. Nicholls. Numerical simulation of a weakly nonlinear model for water waves with viscosity. Journal of Scientific Computing, 42(2):274–290, February 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9324-y;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9324-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=42&issue=2&spage=274-290.

Wang:2010:CME

[819] Zhong-Qing Wang and Li-Lian Wang. A collocation method with exact imposition of mixed boundary conditions. Journal of Scientific Computing, 42(2):291–317, February 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9325-x;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9325-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=42&issue=2&spage=291-317.

Wen:2010:HON

[820] Xin Wen. A high order numerical method for computing physical observables in the semiclassical limit of the one-dimensional linear Schrödinger equation with discontinuous potentials. Journal of Scientific Computing, 42(2):318–344, February 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9326-9;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9326-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=42&issue=2&spage=318-344.

Bergot:2010:HOF

[821] Morgane Bergot, Gary Cohen, and Marc Duruflé. Higher-order finite elements for hybrid meshes using new nodal pyramidal elements. Journal of Scientific Computing, 42(3):345–381, March 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9334-9;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9334-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=3&spage=345-381.

Chen:2010:EES

[822] Yanping Chen, Yunqing Huang, Wenbin Liu, and Ningning Yan. Error estimates and superconvergence of mixed finite element methods for convex optimal control problems. Journal of Scientific Computing, 42 (3):382–403, March 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9327-8;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9327-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=42&issue=3&
spage=382-403.

Yousefi:2010:SWP

[823] H. Yousefi, A. Noorzad, and J. Farjoodi. Simulating 2D waves propagation in elastic solid media using wavelet based adaptive method. Journal of Scientific Computing, 42(3):404–425, March 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9328-7;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9328-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=3&spage=404-425.

Zhu:2010:SEA

[824] Wuming Zhu and David A. Kopriva. A spectral element approximation to price European options with one asset and stochastic volatility. Journal of Scientific Computing, 42(3):426–446, March 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9333-x;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9333-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=42&issue=3&spage=426-446.

Koch:2010:HMC

[825] Othmar Koch, Christopher Ede, and Gerald Jordan. Hierarchical matrices in computations of electron dynamics. Journal of Scientific Computing, 42(3):447–455, March 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9330-0;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9330-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=42&issue=3&spage=447-455.

Amat:2010:FMA

[826] Sergio Amat, Juan Ruiz, and J. Carlos Trillo. Fast multiresolution algorithms and their related variational problems for image denoising. Journal of Scientific Computing, 43(1):1–23, April 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9336-7;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9336-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=1&spage=1-23.

Sun:2010:NSA

[827] Pengtao Sun, Long Chen, and Jinchao Xu. Numerical studies of adaptive finite element methods for two dimensional convectiondominated problems. Journal of Scientific Computing, 43(1):24–43, April 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9337-6;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9337-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=1&spage=24-43.

vandenDoel:2010:MLS

[828] K. van den Doel, U. M. Ascher, and A. Leitão. Multiple level sets for piecewise constant surface reconstruction in highly ill-posed problems. Journal of Scientific Computing, 43(1):44–66, April 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9341-x;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9341-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=1&spage=44-66.

Alvarez-Vazquez:2010:OMB

[829] Lino J. Alvarez-Vázquez and Francisco J. Fernández. Optimal management of a bioreactor for eutrophicated water treatment: a numerical approach. Journal of Scientific Computing, 43(1):67–91, April 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9344-7;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9344-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=1&spage=67-91.

Lin:2010:NST

[830] G. Lin and A. M. Tartakovsky. Numerical studies of three-dimensional stochastic Darcy’s equation and stochastic advection–diffusion-dispersion equation. Journal of Scientific Computing, 43(1):92–117, April 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9346-5;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9346-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=1&spage=92-117.

Bodony:2010:ASA

[831] Daniel J. Bodony. Accuracy of the simultaneous-approximation-term boundary condition for time-dependent problems. Journal of Scientific Computing, 43(1):118–133, April 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9347-4;

 http://link.springer.com/content/pdf/10.1007/s10915-010-9347-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=1&spage=118-133.

Acebron:2010:EPS

[832] Juan A. Acebrón, Ángel Rodríguez-Rozas, and Renato Spigler. Efficient parallel solution of nonlinear parabolic partial differential equations by a probabilistic domain decomposition. Journal of Scientific Computing, 43(2):135–157, May 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9349-2;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9349-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=43&issue=2&
spage=135-157.

Arandiga:2010:PVW

[833] F. Aràndiga, A. M. Belda, and P. Mulet. Point-value WENO multiresolution applications to stable image compression. Journal of Scientific Computing, 43(2):158–182, May 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9351-8;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9351-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=2&spage=158-182.

Afif:2010:PEE

[834] M. Afif, B. Amaziane, G. Kunert, Z. Mghazli, and S. Nicaise. A posteriori error estimation for a finite volume discretization on anisotropic meshes. Journal of Scientific Computing, 43(2):183–200, May 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9352-7;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9352-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=2&spage=183-200.

Guo:2010:GJR

[835] Ben-Yu Guo and Yong-Gang Yi. Generalized Jacobi rational spectral method and its applications. Journal of Scientific Computing, 43 (2):201–238, May 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9353-6;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9353-6;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=43&issue=2&
spage=201-238.

Shih:2010:TFP

[836] Yintzer Shih, R. Bruce Kellogg, and Peishan Tsai. A tailored finite point method for convection–diffusion-reaction problems. Journal of Scientific Computing, 43(2):239–260, May 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9354-5;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9354-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=2&spage=239-260. See erratum

[855].

Burger:2010:AMM

[837] Raimund Bürger, Ricardo Ruiz-Baier, and Kai Schneider. Adaptive multiresolution methods for the simulation of waves in excitable media. Journal of Scientific Computing, 43(2):261–290, May 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9356-3;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9356-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=2&spage=261-290.

Pironneau:2010:P

[838] Olivier Pironneau, Alfio Quarteroni, and Masahisa Tabata. Preface. Journal of Scientific Computing, 43(3):291–292, June 2010. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9370-5;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9370-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=3&spage=291-292.

Burman:2010:IPC

[839] Erik Burman, Alfio Quarteroni, and Benjamin Stamm. Interior penalty continuous and discontinuous finite element approximations of hyperbolic equations. Journal of Scientific Computing, 43(3):293–312, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9232-6;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9232-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=3&spage=293-312.

Flueck:2010:SMN

[840] M. Flueck, A. Janka, C. Laurent, M. Picasso, and J. Rappaz. Some mathematical and numerical aspects in aluminum production. Journal of Scientific Computing, 43(3):313–325, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9227-3;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9227-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=3&spage=313-325.

Ganesan:2010:SLP

[841] Sashikumaar Ganesan and Lutz Tobiska. Stabilization by local projection for convection–diffusion and incompressible flow problems. Journal of Scientific Computing, 43(3):326–342, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9259-8;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9259-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=3&spage=326-342.

Hughes:2010:SMC

[842] Thomas J. R. Hughes and Guglielmo Scovazzi. Stabilized methods for compressible flows. Journal of Scientific Computing, 43(3):343–368, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
008-9233-5;   http://link.springer.com/content/pdf/10.1007/
s10915-008-9233-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=3&spage=343-368.

Lee:2010:BFC

[843] Hyung-Chun Lee and Guang-Ri Piao. Boundary feedback control of the Burgers equations by a reduced-order approach using centroidal Voronoi tessellations. Journal of Scientific Computing, 43(3):369–387, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9310-4;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9310-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=3&spage=369-387.

Nakao:2010:CAP

[844] Mitsuhiro T. Nakao, Yoshitaka Watanabe, and Nobito Yamamoto. Computer assisted proofs of bifurcating solutions for nonlinear heat convection problems. Journal of Scientific Computing, 43(3):388–401, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9303-3;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9303-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=3&spage=388-401.

See erratum [854].

Pironneau:2010:FEC

[845] Olivier Pironneau. Finite element characteristic methods requiring no quadrature. Journal of Scientific Computing, 43(3):402–415, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9276-2;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9276-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=43&issue=3&spage=402-415.

Rui:2010:MCC

[846] Hongxing Rui and Masahisa Tabata. A mass-conservative characteristic finite element scheme for convection–diffusion problems. Journal of Scientific Computing, 43(3):416–432, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9283-3;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9283-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=3&spage=416-432.

Eguchi:2010:DVM

[847] Yuzuru Eguchi. Development of a variational multiscale large-eddy simulation code ‘MISTRAL’ using double-scale finite elements. Journal of Scientific Computing, 43(3):433–453, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-008-9226-4;
 http://link.springer.com/content/pdf/10.1007/s10915-008-9226-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=43&issue=3&spage=433-453.

Knobloch:2010:NSC

[848] Petr Knobloch. Numerical solution of convection–diffusion equations using a nonlinear method of upwind type. Journal of Scientific Computing, 43(3):454–470, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-008-9260-2;   http://link.springer.com/content/
pdf/10.1007/s10915-008-9260-2;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=43&issue=3&
spage=454-470.

Yasuda:2010:TPS

[849] Hidenori Yasuda. Two-phase shallow water equations and phase separation in thin immiscible liquid films. Journal of Scientific Computing, 43(3):471–487, June 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9280-6;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9280-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=43&issue=3&
spage=471-487.

Celiker:2010:HDG

[850] Fatih Celiker, Bernardo Cockburn, and Ke Shi. Hybridizable discontinuous Galerkin methods for Timoshenko beams. Journal of Scientific Computing, 44(1):1–37, July 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9357-2;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9357-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=44&issue=1&spage=1-37.

Wise:2010:USF

[851] S. M. Wise. Unconditionally stable finite difference, nonlinear multigrid simulation of the Cahn–Hilliard–Hele–Shaw system of equations. Journal of Scientific Computing, 44(1):38–68, July 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9363-4;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9363-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=44&issue=1&spage=38-68.

Yang:2010:NAA

[852] Hongtao Yang. A numerical analysis of American options with regime switching. Journal of Scientific Computing, 44(1):69–91, July 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9365-2;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9365-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=1&spage=69-91.

He:2010:NIP

[853] Yinnian He, Liquan Mei, Yueqiang Shang, and Juan Cui. Newton iterative parallel finite element algorithm for the steady Navier– Stokes equations. Journal of Scientific Computing, 44(1):92–106, July 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9371-4;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9371-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=1&spage=92-106.

Nakao:2010:ECA

[854] Mitsuhiro T. Nakao, Yoshitaka Watanabe, and Nobito Yamamoto. Erratum to: Computer Assisted Proofs of Bifurcating Solutions for Nonlinear Heat Convection Problems. Journal of Scientific Computing, 44(1):107, July 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9355-4;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9355-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=1&spage=107. See

[844].

Shih:2010:ETF

[855] Yintzer Shih, R. Bruce Kellogg, and Peishan Tsai. Erratum to: A Tailored Finite Point Method for Convection–Diffusion-Reaction Problems. Journal of Scientific Computing, 44(1):108, July 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9362-5;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9362-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=1&spage=108. See

[836].

Dougalis:2010:BSB

[856] V. A. Dougalis, D. E. Mitsotakis, and J.-C. Saut. Boussinesq systems of Bona–Smith type on plane domains: Theory and numerical analysis. Journal of Scientific Computing, 44(2):109–135, August 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9368-z;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9368-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=2&spage=109-135.

Kopriva:2010:QWF

[857] David A. Kopriva and Gregor Gassner. On the quadrature and weak form choices in collocation type discontinuous Galerkin spectral element methods. Journal of Scientific Computing, 44(2):136–155, August 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9372-3;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9372-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=2&spage=136-155.

Guzman:2010:UAS

[858] J. Guzmán. A unified analysis of several mixed methods for elasticity with weak stress symmetry. Journal of Scientific Computing, 44 (2):156–169, August 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9373-2;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9373-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=44&issue=2&
spage=156-169.

Bourgeade:2010:MEI

[859] Antoine Bourgeade, Candice Mézel, and Olivier Saut. Modeling the early ionization of dielectrics by ultrashort laser pulses. Journal of Scientific Computing, 44(2):170–190, August 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9375-0;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9375-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=44&issue=2&spage=170-190.

Conte:2010:TSR

[860] D. Conte, R. D’Ambrosio, and Z. Jackiewicz. Two-step Runge– Kutta methods with quadratic stability functions. Journal of Scientific Computing, 44(2):191–218, August 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9378-x;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9378-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=44&issue=2&spage=191-218.

Sarmany:2010:OPP

[861] D. Sármány, F. Izsák, and J. J. W. van der Vegt. Optimal penalty parameters for symmetric discontinuous Galerkin discretisations of the timeharmonic Maxwell equations. Journal of Scientific Computing, 44(3): 219–254, September 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9366-1;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9366-1;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=44&issue=3&
spage=219-254.

Guo:2010:CLL

[862] Ben-Yu Guo and Tian-Jun Wang. Composite Laguerre–Legendre spectral method for fourth-order exterior problems. Journal of Scientific Computing, 44(3):255–285, September 2010. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9367-0;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9367-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=44&issue=3&spage=255-285.

Heryudono:2010:RBF

[863] Alfa R. H. Heryudono and Tobin A. Driscoll. Radial basis function interpolation on irregular domain through conformal transplantation. Journal of Scientific Computing, 44(3):286–300, September 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9380-3;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9380-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=3&spage=286-300.

Huang:2010:SCR

[864] Yunqing Huang and Nianyu Yi. The superconvergent cluster recovery method. Journal of Scientific Computing, 44(3):301–322, September 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9379-9;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9379-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=44&issue=3&spage=301-322.

Gao:2010:SCB

[865] Haiyang Gao, Z. J. Wang, and Yen Liu. A study of curved boundary representations for 2D high order Euler solvers. Journal of Scientific Computing, 44(3):323–336, September 2010. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9386-x;

 http://link.springer.com/content/pdf/10.1007/s10915-010-9386-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=44&issue=3&spage=323-336.

Gottlieb:2010:F

[866] Sigal Gottlieb and Chi-Wang Shu. Foreword. Journal of Scientific Computing, 45(1–3):1–2, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-010-9398-6;   http://link.springer.com/
content/pdf/10.1007/s10915-010-9398-6;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=45&issue=
1&spage=1-2.

Abgrall:2010:EHO

[867] R. Abgrall and J. Trefilík. An example of high order residual distribution scheme using non-Lagrange elements. Journal of Scientific Computing, 45(1–3):3–25, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9405-y;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9405-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=45&issue=1&
spage=3-25.

Britt:2010:CFO

[868] S. Britt, S. Tsynkov, and E. Turkel. A compact fourth order scheme for the Helmholtz equation in polar coordinates. Journal of Scientific Computing, 45(1–3):26–47, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9348-3;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9348-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=45&issue=1&spage=26-47.

Brown:2010:ENS

[869] Jed Brown. Efficient nonlinear solvers for nodal high-order finite elements in 3D. Journal of Scientific Computing, 45(1–3):48–63, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9396-8;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9396-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=48-63.

Caboussat:2010:NMV

[870] Alexandre Caboussat and Roland Glowinski. Numerical methods for the vector-valued solutions of non-smooth eigenvalue problems. Journal of Scientific Computing, 45(1–3):64–89, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9383-0;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9383-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=64-89.

Cances:2010:NAN

[871] Eric Cancès, Rachida Chakir, and Yvon Maday. Numerical analysis of nonlinear eigenvalue problems. Journal of Scientific Computing, 45(1– 3):90–117, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9358-1;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9358-1;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=45&issue=1&
spage=90-117.

Carpenter:2010:REI

[872] Mark H. Carpenter, Jan Nordström, and David Gottlieb. Revisiting and extending interface penalties for multi-domain summation-by-parts operators. Journal of Scientific Computing, 45(1–3):118–150, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9301-5;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9301-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=118-150.

Catral:2010:KCF

[873] M. Catral, S. J. Kirkland, M. Neumann, and N.-S. Sze. The Kemeny constant for finite homogeneous ergodic Markov chains. Journal of Scientific Computing, 45(1–3):151–166, October 2010. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9382-1;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9382-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=45&issue=1&spage=151-166.

Chen:2010:BMP

[874] Qingshan Chen, Ming-Cheng Shiue, and Roger Temam. The barotropic mode for the primitive equations. Journal of Scientific Computing, 45(1– 3):167–199, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9343-8;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9343-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=45&issue=1&
spage=167-199.

Chertock:2010:FEO

[875] Alina Chertock, Charles R. Doering, and Eugene Kashdan. A fast explicit operator splitting method for passive scalar advection. Journal of Scientific Computing, 45(1–3):200–214, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9381-2;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9381-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=200-214.

Cockburn:2010:CHM

[876] B. Cockburn, N. C. Nguyen, and J. Peraire. A comparison of HDG methods for Stokes flow. Journal of Scientific Computing, 45(1–3): 215–237, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9359-0;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9359-0;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=45&issue=1&
spage=215-237.

Fishelov:2010:RDP

[877] D. Fishelov, M. Ben-Artzi, and J.-P. Croisille. Recent developments in the pure streamfunction formulation of the Navier–Stokes system. Journal of Scientific Computing, 45(1–3):238–258, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9374-1;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9374-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=238-258.

Funaro:2010:NSE

[878] Daniele Funaro. Numerical simulation of electromagnetic solitons and their interaction with matter. Journal of Scientific Computing, 45(1– 3):259–271, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9338-5;   http://link.springer.com/content/

pdf/10.1007/s10915-009-9338-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=45&issue=1&
spage=259-271.

Goldstein:2010:GAS

[879] Tom Goldstein, Xavier Bresson, and Stanley Osher. Geometric applications of the split Bregman method: Segmentation and surface reconstruction. Journal of Scientific Computing, 45(1–3):272–293, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9331-z;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9331-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=272-293.

Gunzburger:2010:ANS

[880] Max Gunzburger, Eunjung Lee, Yuki Saka, and Catalin Trenchea. Analysis of nonlinear spectral eddy–viscosity models of turbulence. Journal of Scientific Computing, 45(1–3):294–332, October 2010. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-009-9335-8;
 http://link.springer.com/content/pdf/10.1007/s10915-009-9335-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=45&issue=1&spage=294-332.

Ilyevsky:2010:SCA

[881] A. Ilyevsky and E. Turkel. Stopping criteria for anisotropic PDEs in image processing. Journal of Scientific Computing, 45(1–3):333–347, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9361-6;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9361-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=333-347.

Jameson:2010:PSS

[882] Antony Jameson. A proof of the stability of the spectral difference method for all orders of accuracy. Journal of Scientific Computing, 45(1– 3):348–358, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-009-9339-4;   http://link.springer.com/content/
pdf/10.1007/s10915-009-9339-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=45&issue=1&
spage=348-358.

Jung:2010:RHO

[883] Jae-Hun Jung, Sigal Gottlieb, and Saeja Oh Kim. Recovery of high order accuracy in radial basis function approximations of discontinuous problems. Journal of Scientific Computing, 45(1–3):359–381, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9360-7;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9360-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=359-381.

Karni:2010:HAB

[884] Smadar Karni and Gerardo Hernández-Dueñas. A hybrid algorithm for the Baer–Nunziato model using the Riemann invariants. Journal of Scientific Computing, 45(1–3):382–403, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
009-9332-y;   http://link.springer.com/content/pdf/10.1007/
s10915-009-9332-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=382-403.

Li:2010:CDG

[885] Fengyan Li and Sergey Yakovlev. A central discontinuous Galerkin method for Hamilton–Jacobi equations. Journal of Scientific Computing, 45(1–3):404–428, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-009-9340-y;   http://link.springer.com/
content/pdf/10.1007/s10915-009-9340-y;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=45&issue=
1&spage=404-428.

Lin:2010:PMN

[886] Bang-Yan Lin, Chun-Hao Teng, and Hung-Chun Chang. Pseudospectral modeling of nano-optics in Ag sphere arrays. Journal of Scientific Computing, 45(1–3):429–446, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-010-9376-z;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=45&issue=1&spage=429-446.

Mirzaee:2010:QEI

[887] Hanieh Mirzaee, Jennifer K. Ryan, and Robert M. Kirby. Quantification of errors introduced in the numerical approximation and implementation of smoothness-increasing accuracy conserving (SIAC) filtering of discontinuous Galerkin (DG) fields. Journal of Scientific Computing, 45(1–3):447–470, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-009-9342-9;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=45&issue=1&spage=447-470.

Nicholls:2010:BPM

[888] David P. Nicholls and Joseph Orville. A boundary perturbation method for vector electromagnetic scattering from families of doubly periodic gratings. Journal of Scientific Computing, 45(1–3):471–486, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-010-9350-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=45&issue=1&spage=471-486.

Viswanathan:2010:RNU

[889] Adityavikram Viswanathan, Anne Gelb, and Douglas Cochran. On reconstruction from non-uniform spectral data. Journal of Scientific Computing, 45(1–3):487–513, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-010-9364-3;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=45&issue=1&spage=487-513.

Xiong:2010:FSF

[890] Tao Xiong, Mengping Zhang, Yong-Tao Zhang, and Chi-Wang Shu. Fast sweeping fifth order WENO scheme for static Hamilton–Jacobi equations with accurate boundary treatment. Journal of Scientific Computing, 45(1–3):514–536, October 2010. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-010-9345-6;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=45&issue=1&spage=514-536.

Qiao:2011:SEA

[891] Zhonghua Qiao, Changhui Yao, and Shanghui Jia. Superconvergence and extrapolation analysis of a nonconforming mixed finite element approximation for time-harmonic Maxwell’s equations. Journal of Scientific Computing, 46(1):1–19, January 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9406-x;

 http://link.springer.com/content/pdf/10.1007/s10915-010-9406-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=1&spage=1-19.

Zhang:2011:UPD

[892] Xiaoqun Zhang, Martin Burger, and Stanley Osher. A unified primal–dual algorithm framework based on Bregman iteration. Journal of Scientific Computing, 46(1):20–46, January 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9408-8;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9408-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=1&spage=20-46.

tenThijeBoonkkamp:2011:FVC

[893] J. H. M. ten Thije Boonkkamp and M. J. H. Anthonissen. The finite volume-complete flux scheme for advection–diffusion-reaction equations. Journal of Scientific Computing, 46(1):47–70, January 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9388-8;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9388-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=1&spage=47-70.

Pani:2011:LDG

[894] Amiya K. Pani and Sangita Yadav. An hp-local discontinuous Galerkin method for parabolic Integro–Differential equations. Journal of Scientific Computing, 46(1):71–99, January 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9384-z;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9384-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=1&spage=71-99.

Vergara:2011:NMD

[895] Christian Vergara. Nitsche’s method for defective boundary value problems in incompressibile fluid-dynamics. Journal of Scientific Computing, 46(1):100–123, January 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-010-9389-7;   http://link.springer.com/
content/pdf/10.1007/s10915-010-9389-7;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=46&issue=
1&spage=100-123.

Antonietti:2011:CDD

[896] Paola F. Antonietti and Paul Houston. A class of domain decomposition preconditioners for hp-discontinuous Galerkin finite element methods. Journal of Scientific Computing, 46(1):124–149, January 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9390-1;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9390-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=1&spage=124-149.

Wihler:2011:DGM

[897] Thomas P. Wihler and Béatrice Rivière. Discontinuous Galerkin methods for second-order elliptic PDE with low-regularity solutions. Journal of Scientific Computing, 46(2):151–165, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9387-9;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9387-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=2&spage=151-165.

Yokoi:2011:NMI

[898] Kensuke Yokoi. Numerical method for interaction among multi-particle, fluid and arbitrary shape structure. Journal of Scientific Computing, 46 (2):166–181, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9385-y;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9385-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=46&issue=2&
spage=166-181.

Gong:2011:MFE

[899] Wei Gong and Ningning Yan. A mixed finite element scheme for optimal control problems with pointwise state constraints. Journal of Scientific Computing, 46(2):182–203, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9392-z;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9392-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=2&spage=182-203.

Filbet:2011:APS

[900] Francis Filbet and Shi Jin. An asymptotic preserving scheme for the ES–BGK model of the Boltzmann equation. Journal of Scientific Computing, 46(2):204–224, February 2011. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9394-x;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9394-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=2&spage=204-224.

Jung:2011:SRP

[901] Jae-Hun Jung and Wolfgang Stefan. A simple regularization of the polynomial interpolation for the resolution of the Runge phenomenon. Journal of Scientific Computing, 46(2):225–242, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9397-7;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9397-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=2&spage=225-242.

vanderLaan:2011:MCE

[902] Wladimir J. van der Laan and Andrei C. Jalba. A memory and computation efficient sparse level-set method. Journal of Scientific Computing, 46 (2):243–264, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9399-5;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9399-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=46&issue=2&
spage=243-264.

Ray:2011:TFH

[903] Rajendra K. Ray. A transformation-free HOC scheme for incompressible viscous flow past a rotating and translating Circular cylinder. Journal of Scientific Computing, 46(2):265–293, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9401-2;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9401-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=2&spage=265-293.

Zhang:2011:TOE

[904] Qiang Zhang. Third order explicit Runge–Kutta discontinuous Galerkin method for linear conservation law with inflow boundary condition. Journal of Scientific Computing, 46(2):294–313, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

010-9403-0;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9403-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=2&spage=294-313.

Kannan:2011:LVL

[905] R. Kannan and Z. J. Wang. LDG2: a variant of the LDG flux formulation for the spectral volume method. Journal of Scientific Computing, 46 (2):314–328, February 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9391-0;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9391-0;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=46&issue=2&
spage=314-328.

Tanaka:2011:SUG

[906] S. Tanaka, S. Bunya, J. J. Westerink, and C. Dawson. Scalability of an unstructured grid continuous Galerkin based hurricane Storm surge model. Journal of Scientific Computing, 46(3):329–358, March 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9402-1;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9402-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=3&spage=329-358.

Teng:2011:SSF

[907] Fei Teng, Li Yuan, and Tao Tang. A speed-up strategy for finite volume WENO schemes for hyperbolic conservation laws. Journal of Scientific Computing, 46(3):359–378, March 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9407-9;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9407-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=3&spage=359-378.

Lamichhane:2011:MFE

[908] Bishnu P. Lamichhane. A mixed finite element method for the biharmonic problem using biorthogonal or quasi-biorthogonal systems. Journal of Scientific Computing, 46(3):379–396, March 2011. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9409-7;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9409-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=3&spage=379-396.

Vohralik:2011:GFR

[909] Martin Vohralík. Guaranteed and fully robust a posteriori error estimates for conforming discretizations of diffusion problems with discontinuous coefficients. Journal of Scientific Computing, 46(3):397–438, March 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9410-1;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9410-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=3&spage=397-438.

Karunatillake:2011:RSS

[910] Suniti Karunatillake and Steven W. Squyres. Recipes for spatial statistics with global datasets: a Martian case study. Journal of Scientific Computing, 46(3):439–451, March 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9412-z;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9412-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=3&spage=439-451.

Marinov:2011:CLB

[911] Tchavdar T. Marinov and Keng Deng. Characteristic line based schemes for solving a quasilinear hierarchical size-structured model. Journal of Scientific Computing, 46(3):452–469, March 2011. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9413-y;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9413-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=46&issue=3&spage=452-469.

Wang:2011:LBM

[912] Huimin Wang, Guangwu Yan, and Bo Yan. Lattice Boltzmann model based on the rebuilding-divergency method for the Laplace equation and the Poisson equation. Journal of Scientific Computing, 46(3):470–484, March 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9414-x;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9414-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=3&spage=470-484.

Llanas:2011:EDA

[913] Bernardo Llanas and Sagrario Lantarón. Edge detection by adaptive splitting. Journal of Scientific Computing, 46(3):485–518, March 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9416-8;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9416-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=3&spage=485-518.

Karageorghis:2011:EMA

[914] A. Karageorghis. Efficient MFS algorithms for inhomogeneous polyharmonic problems. Journal of Scientific Computing, 46(3):519–541, March 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9418-6;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9418-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=46&issue=3&spage=519-541.

Li:2011:UAL

[915] Jichun Li. Unified analysis of leap-frog methods for solving time-domain Maxwell’s equations in dispersive media. Journal of Scientific Computing, 47(1):1–26, April 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9417-7;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9417-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=47&issue=1&
spage=1-26.

Barker:2011:TLA

[916] A. T. Barker, S. C. Brenner, E.-H. Park, and L.-Y. Sung. Two-level additive Schwarz preconditioners for a weakly over-penalized symmetric interior penalty method. Journal of Scientific Computing, 47(1):27–49, April 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9419-5;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9419-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=1&spage=27-49.

Vincent:2011:NCH

[917] P. E. Vincent, P. Castonguay, and A. Jameson. A new class of high-order energy stable flux reconstruction schemes. Journal of Scientific Computing, 47(1):50–72, April 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9420-z;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9420-z;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=47&issue=1&
spage=50-72.

Zhang:2011:PEM

[918] Jing Zhang, Li-Lian Wang, and Zhijian Rong. A prolate-element method for nonlinear PDEs on the sphere. Journal of Scientific Computing, 47(1):73–92, April 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9421-y;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9421-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=47&issue=1&
spage=73-92.

Chen:2011:CRM

[919] Shao-Chun Chen and Ya-Na Wang. Conforming rectangular mixed finite elements for elasticity. Journal of Scientific Computing, 47(1):93–108, April 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9422-x;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9422-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=1&spage=93-108.

Yilmaz:2011:IPL

[920] S. L. Yilmaz, M. B. Nik, M. R. H. Sheikhi, and P. A. Strakey. An irregularly portioned Lagrangian Monte Carlo method for turbulent flow simulation. Journal of Scientific Computing, 47(1):109–125, April 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9424-8;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9424-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=1&spage=109-125.

Motamed:2011:LSF

[921] Mohammad Motamed, Colin B. Macdonald, and Steven J. Ruuth. On the linear stability of the fifth-order WENO discretization. Journal of Scientific Computing, 47(2):127–149, May 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9423-9;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9423-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=47&issue=2&spage=127-149.

Chen:2011:ETS

[922] Rongsan Chen and De kang Mao. Entropy-TVD scheme for nonlinear scalar conservation laws. Journal of Scientific Computing, 47(2):150–169, May 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9431-9;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9431-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=2&spage=150-169.

Chaudhuri:2011:NSC

[923] Arnab Chaudhuri, Abdellah Hadjadj, and Ashwin Chinnayya. Numerical study of compressible mixing layers using high-order WENO schemes. Journal of Scientific Computing, 47(2):170–197, May 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9429-3;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9429-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=2&spage=170-197.

Shih:2011:CTF

[924] Yintzer Shih, R. Bruce Kellogg, and Yoyo Chang. Characteristic tailored finite point method for convection-dominated convection–diffusionreaction problems. Journal of Scientific Computing, 47(2):198–215, May 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9433-7;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9433-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=2&spage=198-215.

Zhang:2011:ICS

[925] Shuhai Zhang, Shufen Jiang, and Chi-Wang Shu. Improvement of convergence to steady state solutions of Euler equations with the WENO schemes. Journal of Scientific Computing, 47(2):216–238, May 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9435-5;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9435-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=2&spage=216-238.

Guo:2011:LSP

[926] Wei Guo, Fengyan Li, and Jianxian Qiu. Local-structure-preserving discontinuous Galerkin methods with Lax–Wendroff type time discretizations for Hamilton–Jacobi equations. Journal of Scientific Computing, 47(2):239–257, May 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9434-6;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9434-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=47&issue=2&
spage=239-257.

Yang:2011:PTG

[927] Haijian Yang and Xiao-Chuan Cai. Parallel two-grid semismooth Newton–Krylov–Schwarz method for nonlinear complementarity problems. Journal of Scientific Computing, 47(2):258–280, May 2011. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9436-4;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9436-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=47&issue=2&spage=258-280.

Lu:2011:SSW

[928] Changna Lu and Jianxian Qiu. Simulations of shallow water equations with finite difference Lax–Wendroff weighted essentially nonoscillatory schemes. Journal of Scientific Computing, 47(3):281–302, June 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9437-3;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9437-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=3&spage=281-302.

Feng:2011:AGM

[929] Xiaobing Feng and Michael Neilan. Analysis of Galerkin methods for the fully nonlinear Monge–Ampère equation. Journal of Scientific Computing, 47(3):303–327, June 2011. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9439-1;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9439-
1;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=47&issue=3&spage=303-327.

Zeiser:2011:FMV

[930] Andreas Zeiser. Fast matrix-vector multiplication in the sparse-grid Galerkin method. Journal of Scientific Computing, 47(3):328–346, June 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9438-2;   http://link.springer.com/content/pdf/10.1007/

s10915-010-9438-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=3&spage=328-346.

Chaudhry:2011:FEA

[931] Jehanzeb Hameed Chaudhry, Stephen D. Bond, and Luke N. Olson. Finite element approximation to a finite-size modified Poisson– Boltzmann equation. Journal of Scientific Computing, 47(3):347–364, June 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9441-7;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9441-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=3&spage=347-364.

Dedner:2011:GSA

[932] Andreas Dedner and Robert Klöfkorn. A generic stabilization approach for higher order discontinuous Galerkin methods for convection dominated problems. Journal of Scientific Computing, 47(3):365–388, June 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9448-0;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9448-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=47&issue=3&spage=365-388.

Cliffe:2011:APE

[933] K. Andrew Cliffe, Edward J. C. Hall, and Paul Houston. Adaptivity and a posteriori error control for bifurcation problems II: Incompressible fluid flow in open systems with Z_{2} symmetry. Journal of Scientific Computing, 47(3):389–418, June 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9453-3;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9453-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=47&issue=3&
spage=389-418.

Abdellatif:2011:SDA

[934] Nehla Abdellatif, Nejmeddine Chorfi, and Sihem Trabelsi. Spectral discretization of the axisymmetric vorticity, velocity and pressure formulation of the Stokes problem. Journal of Scientific Computing, 47 (3):419–440, June 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9446-2;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9446-2;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=47&issue=3&

spage=419-440. See erratum [1019].

Abgrall:2011:P

[935] Rémi Abgrall, Denise Aregba, and Christophe Berthon. Preface. Journal of Scientific Computing, 48(1–3):1–2, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9497-z;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9497-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=1-2.

Bohorquez:2011:HIW

[936] Patricio Bohorquez and Martin Rentschler. Hydrodynamic instabilities in well-balanced finite volume schemes for frictional shallow water equations. The kinematic wave case. Journal of Scientific Computing, 48(1–3):3–15, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9444-4;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9444-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=48&issue=1&
spage=3-15.

Castro-Diaz:2011:NTL

[937] M. J. Castro-Díaz and E. D. Fernández-Nieto. Numerical treatment of the loss of hyperbolicity of the two-layer shallow-water system. Journal of Scientific Computing, 48(1–3):16–40, July 2011. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9427-5;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9427-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=16-40.

Delestre:2011:NSV

[938] Olivier Delestre and Fabien Marche. A numerical scheme for a viscous shallow water model with friction. Journal of Scientific Computing, 48(1–3):41–51, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9393-y;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9393-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=48&issue=1&
spage=41-51.

Donat:2011:HSO

[939] R. Donat and A. Martinez-Gavara. Hybrid second order schemes for scalar balance laws. Journal of Scientific Computing, 48(1–3):52–69, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9404-z;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9404-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=52-69.

Dumbser:2011:SEO

[940] Michael Dumbser and Eleuterio F. Toro. A simple extension of the Osher Riemann solver to non-conservative hyperbolic systems. Journal of Scientific Computing, 48(1–3):70–88, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9400-3;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9400-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=70-88.

Bourdarias:2011:KST

[941] C. Bourdarias, M. Ersoy, and S. Gerbi. A kinetic scheme for transient mixed flows in non uniform closed pipes: a global manner to upwind all the source terms. Journal of Scientific Computing, 48(1–3):89–104, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9456-0;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9456-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=89-104.

Chazel:2011:NSS

[942] F. Chazel, D. Lannes, and F. Marche. Numerical simulation of strongly nonlinear and dispersive waves using a Green–Naghdi model. Journal of Scientific Computing, 48(1–3):105–116, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9395-9;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9395-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=105-116.

Fernandez-Nieto:2011:IFC

[943] E. D. Fernández-Nieto, M. J. Castro Díaz, and C. Parés. On an intermediate field capturing Riemann solver based on a parabolic viscosity matrix for the two-layer shallow water system. Journal of Scientific Computing, 48(1–3):117–140, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9465-7;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9465-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=117-140.

Gallardo:2011:TDC

[944] José M. Gallardo, Sergio Ortega, and Marc de la Asunción. Twodimensional compact third-order polynomial reconstructions. Solving nonconservative hyperbolic systems using GPUs. Journal of Scientific Computing, 48(1–3):141–163, July 2011. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9470-x;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9470-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=141-163.

Coquel:2011:APS

[945] F. Coquel and E. Godlewski. Asymptotic preserving scheme for Euler system with large friction. Journal of Scientific Computing, 48(1– 3):164–172, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9459-5;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9459-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=48&issue=1&
spage=164-172.

Hidalgo:2011:ASN

[946] Arturo Hidalgo and Michael Dumbser. ADER schemes for nonlinear systems of stiff advection–diffusion-reaction equations. Journal of Scientific Computing, 48(1–3):173–189, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9426-6;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9426-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=173-189.

Hernandez-Duenas:2011:SWF

[947] Gerardo Hernández-Dueñas and Smadar Karni. Shallow water flows in channels. Journal of Scientific Computing, 48(1–3):190–208, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9430-x;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9430-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=190-208.

LeVeque:2011:WBP

[948] Randall J. LeVeque. A well-balanced path-integral f-wave method for hyperbolic problems with source terms. Journal of Scientific Computing, 48(1–3):209–226, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9411-0;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9411-0;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=48&issue=1&
spage=209-226.

Hundertmark-Zauskova:2011:LTS

[949] A. Hundertmark-Zausková and M. Lukácová-Medvid’ová. Large time step finite volume evolution Galerkin methods. Journal of Scientific Computing, 48(1–3):227–240, July 2011. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9443-5;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9443-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=227-240.

Martinez-Gavara:2011:HSO

[950] A. Martinez-Gavara and R. Donat. A hybrid second order scheme for shallow water flows. Journal of Scientific Computing, 48(1–3):241–257, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9440-8;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9440-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=241-257.

deLuna:2011:DMS

[951] T. Morales de Luna, M. J. Castro Díaz, and C. Parés Madroñal. A duality method for sediment transport based on a modified Meyer–Peter & Müller model. Journal of Scientific Computing, 48(1–3):258–273, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9447-1;   http://link.springer.com/content/pdf/10.1007/

s10915-010-9447-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=258-273.

Munoz-Ruiz:2011:CWB

[952] María Luz Muñoz-Ruiz and Carlos Parés. On the convergence and wellbalanced property of path-conservative numerical schemes for systems of balance laws. Journal of Scientific Computing, 48(1–3):274–295, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9425-7;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9425-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=274-295.

Perrier:2011:CMS

[953] Vincent Perrier. A conservative method for the simulation of the isothermal Euler system with the van-der-Waals equation of state. Journal of Scientific Computing, 48(1–3):296–303, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9415-9;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9415-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=296-303.

Ricchiuto:2011:CPG

[954] M. Ricchiuto. On the C-property and generalized C-property of residual distribution for the shallow water equations. Journal of Scientific Computing, 48(1–3):304–318, July 2011. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9369-y;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9369-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=48&issue=1&spage=304-318.

Begnudelli:2011:HSF

[955] L. Begnudelli and G. Rosatti. Hyperconcentrated 1D shallow flows on fixed bed with geometrical source term due to a bottom step. Journal of Scientific Computing, 48(1–3):319–332, July 2011. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/content/pdf/10.1007/s10915-010-9457-z;
http://www.springerlink.com/openurl.asp?genre=article&issn=
0885-7474&volume=48&issue=1&spage=319-332.

Simeoni:2011:RCU

[956] Chiara Simeoni. Remarks on the consistency of upwind source at interface schemes on nonuniform grids. Journal of Scientific Computing, 48(1–3):333–338, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-010-9442-6;   http:
//www.springerlink.com/openurl.asp?genre=article&issn=0885-
7474&volume=48&issue=1&spage=333-338.

Xing:2011:AWB

[957] Yulong Xing, Chi-Wang Shu, and Sebastian Noelle. On the advantage of well-balanced schemes for moving-water equilibria of the shallow water equations. Journal of Scientific Computing, 48(1–3):339–349, July 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-010-9377-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=48&issue=1&spage=339-349.

Salvetti:2011:P

[958] Maria Vittoria Salvetti and Pierre Sagaut. Preface. Journal of Scientific Computing, 49(1):1–2, October 2011. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9510-6;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9510-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=49&issue=1&spage=1-2.

Berselli:2011:HLE

[959] Luigi C. Berselli, Paul F. Fischer, and Traian Iliescu. Horizontal large eddy simulation of stratified mixing in a lock–exchange system. Journal of Scientific Computing, 49(1):3–20, October 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9464-8;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9464-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=1&spage=3-20.

Graham:2011:ESS

[960] Jonathan Pietarila Graham, Darryl D. Holm, and Pablo Mininni. The effect of subfilter-scale physics on regularization models. Journal of Scientific Computing, 49(1):21–34, October 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9428-4;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9428-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=49&issue=1&spage=21-34.

Guermond:2011:SWS

[961] Jean-Luc Guermond, Richard Pasquetti, and Bojan Popov. From suitable weak solutions to entropy viscosity. Journal of Scientific Computing, 49(1):35–50, October 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9445-3;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9445-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=49&issue=1&
spage=35-50.

Kempf:2011:QIC

[962] A. M. Kempf, B. J. Geurts, T. Ma, and M. W. A. Pettit. Quality issues in combustion LES. Journal of Scientific Computing, 49(1):51–64, October 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9481-7;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9481-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=1&spage=51-64.

Meyers:2011:ELA

[963] Johan Meyers. Error-landscape assessment of large-eddy simulations: a review of the methodology. Journal of Scientific Computing, 49(1): 65–77, October 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9449-z;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9449-z;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=49&issue=1&
spage=65-77.

Poinsot:2011:NPI

[964] T. Poinsot, M. Garcia, J. M. Senoner, and L. Gicquel. Numerical and physical instabilities in massively parallel LES of reacting flows. Journal of Scientific Computing, 49(1):78–93, October 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9432-8;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9432-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=1&spage=78-93.

Verstappen:2011:WDE

[965] Roel Verstappen. When does eddy viscosity damp subfilter scales sufficiently? Journal of Scientific Computing, 49(1):94–110, October 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9504-4;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9504-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=1&spage=94-110.

Baudouin:2011:ATM

[966] Lucie Baudouin, Julien Salomon, and Gabriel Turinici. Analysis of the “Toolkit” method for the time-dependent Schrödinger equation. Journal of Scientific Computing, 49(2):111–136, November 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9450-6;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9450-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=2&spage=111-136.

Behrens:2011:NFM

[967] Edwin M. Behrens and J. Guzmán. A new family of mixed methods for the Reissner–Mindlin plate model based on a system of first-order equations. Journal of Scientific Computing, 49(2):137–166, November 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9451-5;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9451-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=2&spage=137-166.

Christlieb:2011:IPT

[968] Andrew Christlieb and Benjamin Ong. Implicit parallel time integrators. Journal of Scientific Computing, 49(2):167–179, November 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9452-4;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9452-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=2&spage=167-179.

Chamard:2011:CME

[969] Jeremy Chamard, Josef Otta, and David J. B. Lloyd. Computation of minimum energy paths for quasi-linear problems. Journal of Scientific Computing, 49(2):180–194, November 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9462-x;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9462-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=49&issue=2&spage=180-194.

Ye:2011:NMB

[970] Lina Ye, Guangwu Yan, and Tingting Li. Numerical method based on the lattice Boltzmann model for the Kuramoto–Sivashinsky equation. Journal of Scientific Computing, 49(2):195–210, November 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
010-9455-1;   http://link.springer.com/content/pdf/10.1007/
s10915-010-9455-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=2&spage=195-210.

delSastre:2011:EAF

[971] Pedro Galán del Sastre and Rodolfo Bermejo. Error analysis for hp-FEM semi-Lagrangian second order BDF method for convectiondominated diffusion problems. Journal of Scientific Computing, 49(2): 211–237, November 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-010-9454-2;   http://link.springer.com/content/
pdf/10.1007/s10915-010-9454-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=49&issue=2&
spage=211-237.

Abbas:2011:FOH

[972] Ali Abbas and Jean-Pierre Croisille. A fourth order Hermitian boxscheme with fast solver for the Poisson problem in a square. Journal of Scientific Computing, 49(3):239–267, December 2011. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-010-9458-y;
 http://link.springer.com/content/pdf/10.1007/s10915-010-9458-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=49&issue=3&spage=239-267.

Simpson:2011:SWB

[973] G. Simpson and M. Spiegelman. Solitary wave benchmarks in magma dynamics. Journal of Scientific Computing, 49(3):268–290, December 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9461-y;   http://link.springer.com/content/pdf/10.1007/

s10915-011-9461-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=3&spage=268-290.

Griesmaier:2011:EAH

[974] Roland Griesmaier and Peter Monk. Error analysis for a hybridizable discontinuous Galerkin method for the Helmholtz equation. Journal of Scientific Computing, 49(3):291–310, December 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9460-z;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9460-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=3&spage=291-310.

Bi:2011:TGD

[975] Chunjia Bi and Victor Ginting. Two-grid discontinuous Galerkin method for quasi-linear elliptic problems. Journal of Scientific Computing, 49(3): 311–331, December 2011. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9463-9;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9463-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=49&issue=3&
spage=311-331.

Frank:2011:AFE

[976] Martin Frank, Jens Lang, and Matthias Schäfer. Adaptive finite element simulation of the time-dependent simplified P_{N} equations. Journal of Scientific Computing, 49(3):332–350, December 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9466-6;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9466-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=3&spage=332-350.

Huang:2011:TFP

[977] Zhongyi Huang and Xu Yang. Tailored finite point method for first order wave equation. Journal of Scientific Computing, 49(3):351–366, December 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9468-4;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9468-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=3&spage=351-366.

Zhang:2011:NBS

[978] Jiwei Zhang, Houde Han, and Hermann Brunner. Numerical blow-up of semilinear parabolic PDEs on unbounded domains in R^{2}. Journal of Scientific Computing, 49(3):367–382, December 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9467-5;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9467-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=3&spage=367-382.

Chen:2011:TGM

[979] Luoping Chen and Yanping Chen. Two-grid method for nonlinear reaction–diffusion equations by mixed finite element methods. Journal of Scientific Computing, 49(3):383–401, December 2011. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9469-3;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9469-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=49&issue=3&spage=383-401.

Constantin:2012:NNR

[980] Peter Constantin, Ming-Chih Lai, and Ramjee Sharma. New numerical results for the surface quasi-geostrophic equation. Journal of Scientific Computing, 50(1):1–28, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9471-9;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9471-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=1&spage=1-28.

Zhang:2012:MPS

[981] Xiangxiong Zhang, Yinhua Xia, and Chi-Wang Shu. Maximumprinciple-satisfying and positivity-preserving high order discontinuous Galerkin schemes for conservation laws on triangular meshes. Journal of Scientific Computing, 50(1):29–62, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9472-8;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9472-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=1&spage=29-62.

Nguyen-Ba:2012:SSP

[982] Truong Nguyen-Ba, Huong Nguyen-Thu, and Thierry Giordano. Strong-stability-preserving 7-stage Hermite–Birkhoff time-discretization methods. Journal of Scientific Computing, 50(1):63–90, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9473-7;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9473-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=1&spage=63-90.

Awanou:2012:TRR

[983] Gerard Awanou. Two remarks on rectangular mixed finite elements for elasticity. Journal of Scientific Computing, 50(1):91–102, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9474-6;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9474-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=1&spage=91-102.

Cai:2012:ENX

[984] Zhenning Cai, Ruo Li, and Yanli Wang. An efficient NRxx method for Boltzmann–BGK equation. Journal of Scientific Computing, 50(1): 103–119, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9475-5;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9475-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=50&issue=1&
spage=103-119.

Kesserwani:2012:LLF

[985] Georges Kesserwani and Qiuhua Liang. Locally limited and fully conserved RKDG2 shallow water solutions with wetting and drying. Journal of Scientific Computing, 50(1):120–144, January 2012. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9476-4;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9476-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=1&spage=120-144.

Wu:2012:ALM

[986] Chunlin Wu, Juyong Zhang, Yuping Duan, and Xue-Cheng Tai. Augmented Lagrangian method for total variation based image restoration and segmentation over triangulated surfaces. Journal of Scientific Computing, 50(1):145–166, January 2012. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9477-3;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9477-
3;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=1&spage=145-166.

Kormann:2012:DSS

[987] Katharina Kormann, Martin Kronbichler, and Bernhard Müller. Derivation of strictly stable high order difference approximations for variablecoefficient PDE. Journal of Scientific Computing, 50(1):167–197, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9479-1;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9479-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=1&spage=167-197.

Deparis:2012:SRB

[988] Simone Deparis and A. Emil Løvgren. Stabilized reduced basis approximation of incompressible three-dimensional Navier–Stokes equations in parametrized deformed domains. Journal of Scientific Computing, 50 (1):198–212, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9478-2;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9478-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=50&issue=1&
spage=198-212.

Laminie:2012:DPP

[989] J. Laminie and P. Poullet. A dynamic penalty or projection method for incompressible fluids. Journal of Scientific Computing, 50(1):213– 234, January 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9480-8;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9480-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=50&issue=1&
spage=213-234.

Hahn:2012:ALM

[990] Jooyoung Hahn, Chunlin Wu, and Xue-Cheng Tai. Augmented Lagrangian method for generalized TV-Stokes model. Journal of Scientific Computing, 50(2):235–264, February 2012. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9482-6;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9482-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=2&spage=235-264.

Hundsdorfer:2012:SRB

[991] W. Hundsdorfer, A. Mozartova, and M. N. Spijker. Stepsize restrictions for boundedness and monotonicity of multistep methods. Journal of Scientific Computing, 50(2):265–286, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9487-1;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9487-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=2&spage=265-286.

Dede:2012:RBM

[992] Luca Dedè. Reduced basis method and error estimation for parametrized optimal control problems with control constraints. Journal of Scientific Computing, 50(2):287–305, February 2012. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9483-5;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9483-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=2&spage=287-305.

Alexanderian:2012:MSP

[993] Alen Alexanderian, Oliver P. Le Maître, and Habib N. Najm. Multiscale stochastic preconditioners in non-intrusive spectral projection. Journal of Scientific Computing, 50(2):306–340, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9486-2;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9486-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=2&spage=306-340.

Kozdon:2012:IWF

[994] Jeremy E. Kozdon, Eric M. Dunham, and Jan Nordström. Interaction of waves with frictional interfaces using summation-by-parts difference operators: Weak enforcement of nonlinear boundary conditions. Journal of Scientific Computing, 50(2):341–367, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9485-3;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9485-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=2&spage=341-367.

Christensen:2012:CCH

[995] Brian B. Christensen, Michael B. Nielsen, and Ken Museth. Out-of-core computations of high-resolution level sets by means of code transformation. Journal of Scientific Computing, 50(2):368–404, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9488-0;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9488-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=2&spage=368-404.

Celiker:2012:NSS

[996] Fatih Celiker, Zhimin Zhang, and Huiqing Zhu. Nodal superconvergence of SDFEM for singularly perturbed problems. Journal of Scientific Computing, 50(2):405–433, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9489-z;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9489-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=2&spage=405-433.

Jameson:2012:NLS

[997] A. Jameson, P. E. Vincent, and P. Castonguay. On the non-linear stability of flux reconstruction schemes. Journal of Scientific Computing, 50 (2):434–445, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9490-6;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9490-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=50&issue=2&
spage=434-445.

Xin:2012:WCO

[998] Jianguo Xin and Wei Cai. Well-conditioned orthonormal hierarchical L_{∈} bases on R^{n} simplicial elements. Journal of Scientific Computing, 50 (2):446–461, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9491-5;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9491-5;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=50&issue=2&
spage=446-461.

vanderPijl:2012:EBM

[999] S. P. van der Pijl and C. W. Oosterlee. An ENO-based method for second-order equations and application to the control of dike levels. Journal of Scientific Computing, 50(2):462–492, February 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9493-3;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9493-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=2&spage=462-492.

Gelb:2012:SIM

[1000] Anne Gelb, Rosemary Renaut, and Svetlana Roudenko. Special issue on mathematical methods in medical imaging. Journal of Scientific Computing, 50(3):493–494, March 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9576-9;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9576-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=3&spage=493-494.

Wang:2012:MDE

[1001] Yang Wang, Guo-Wei Wei, and Siyang Yang. Mode decomposition evolution equations. Journal of Scientific Computing, 50(3):495–518, March 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9509-z;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9509-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=3&spage=495-518.

Zhang:2012:NSR

[1002] Xiaoqun Zhang, Yujie Lu, and Tony Chan. A novel sparsity reconstruction method from Poisson data for 3D bioluminescence tomography. Journal of Scientific Computing, 50(3):519–535, March 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9533-z;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9533-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=3&spage=519-535.

See erratum [1021].

Stefan:2012:SEE

[1003] W. Stefan, A. Viswanathan, A. Gelb, and R. Renaut. Sparsity enforcing edge detection method for blurred and noisy Fourier data. Journal of Scientific Computing, 50(3):536–556, March 2012. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9536-9;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9536-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=3&spage=536-556.

Lui:2012:OSR

[1004] Lok Ming Lui, Tsz Wai Wong, Wei Zeng, and Xianfeng Gu. Optimization of surface registrations using Beltrami holomorphic flow. Journal of Scientific Computing, 50(3):557–585, March 2012. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9506-2;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9506-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=3&spage=557-585. See erratum

[1020].

Lin:2012:GED

[1005] Tungyou Lin, Carole Le Guyader, Ivo Dinov, and Paul Thompson. Gene expression data to mouse Atlas registration using a nonlinear elasticity smoother and landmark points constraints. Journal of Scientific Computing, 50(3):586–609, March 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9563-6;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9563-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=3&spage=586-609.

Turnes:2012:ECD

[1006] Christopher K. Turnes and Justin Romberg. Efficient calculations of 3-D FFTs on spiral contours. Journal of Scientific Computing, 50(3):610–628, March 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9534-y;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9534-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=50&issue=3&spage=610-628.

Wang:2012:IFD

[1007] Yang Wang, Guo-Wei Wei, and Siyang Yang. Iterative filtering decomposition based on local spectral evolution kernel. Journal of Scientific Computing, 50(3):629–664, March 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9496-0;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9496-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=3&spage=629-664.

Stefan:2012:WBN

[1008] Wolfgang Stefan, Kewei Chen, and Hongbin Guo. Wavelet-based de-noising of positron emission tomography scans. Journal of Scientific Computing, 50(3):665–677, March 2012. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9529-8;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9529-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=50&issue=3&spage=665-677.

Ji:2012:OEE

[1009] Liangyue Ji and Yan Xu. Optimal error estimates of the local discontinuous Galerkin method for surface diffusion of graphs on Cartesian meshes. Journal of Scientific Computing, 51(1):1–27, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9492-4;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9492-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=1&spage=1-27.

Eftang:2012:TSC

[1010] J. L. Eftang, D. B. P. Huynh, and D. J. Knezevic. A two-step certified reduced basis method. Journal of Scientific Computing, 51(1):28–58, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9494-2;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9494-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=1&spage=28-58.

Degond:2012:NAE

[1011] P. Degond, H. Liu, D. Savelief, and M.-H. Vignal. Numerical approximation of the Euler–Poisson–Boltzmann model in the quasineutral limit. Journal of Scientific Computing, 51(1):59–86, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9495-1;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9495-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=1&spage=59-86.

Ratnani:2012:AHO

[1012] Ahmed Ratnani and Eric Sonnendrücker. An arbitrary highorder spline finite element solver for the time domain Maxwell equations. Journal of Scientific Computing, 51(1):87–106, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9500-8;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9500-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=1&spage=87-106.

Zhang:2012:FDL

[1013] Qiang Zhang and Fuzheng Gao. A fully-discrete local discontinuous Galerkin method for convection-dominated Sobolev equation. Journal of Scientific Computing, 51(1):107–134, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9498-y;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9498-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=51&issue=1&spage=107-134.

Chabaud:2012:HIE

[1014] B. Chabaud and Q. Du. A hybrid implicit-explicit adaptive multirate numerical scheme for time-dependent equations. Journal of Scientific Computing, 51(1):135–157, April 2012. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9499-x;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9499-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=51&issue=1&spage=135-157.

Gelb:2012:REA

[1015] Anne Gelb and Taylor Hines. Recovering exponential accuracy from non-harmonic Fourier data through spectral reprojection. Journal of Scientific Computing, 51(1):158–182, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9502-6;

 http://link.springer.com/content/pdf/10.1007/s10915-011-9502-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=51&issue=1&spage=158-182.

Kirby:2012:CHC

[1016] Robert M. Kirby, Spencer J. Sherwin, and Bernardo Cockburn. To CG or to HDG: a comparative study. Journal of Scientific Computing, 51(1):183–212, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9501-7;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9501-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=1&
spage=183-212.

Li:2012:NON

[1017] Fengyan Li. On the negative-order norm accuracy of a local-structure-preserving LDG method. Journal of Scientific Computing, 51(1): 213–223, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9503-5;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9503-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=1&
spage=213-223.

Castonguay:2012:NCH

[1018] P. Castonguay, P. E. Vincent, and A. Jameson. A new class of high-order energy stable flux reconstruction schemes for triangular elements. Journal of Scientific Computing, 51(1):224–256, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9505-3;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9505-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=1&spage=224-256.

Abdellatif:2012:ESD

[1019] Nehla Abdellatif, Nejmeddine Chorfi, and Sihem Trabelsi. Erratum to: Spectral Discretization of the Axisymmetric Vorticity, Velocity and Pressure Formulation of the Stokes Problem. Journal of Scientific Computing, 51(1):257, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9484-4;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9484-4;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=51&issue=1&

spage=257. See [934].

Lui:2012:EOS

[1020] Lok Ming Lui, Tsz Wai Wong, Wei Zeng, and Xianfeng Gu. Erratum to: Optimization of Surface Registrations Using Beltrami Holomorphic Flow. Journal of Scientific Computing, 51(1):258, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9541-z;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9541-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=1&spage=258. See

[1004].

Zhang:2012:ENS

[1021] Xiaoqun Zhang, Yujie Lu, and Tony Chan. Erratum to: A Novel Sparsity Reconstruction Method from Poisson Data for 3D Bioluminescence Tomography. Journal of Scientific Computing, 51(1):259, April 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9544-9;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9544-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=1&spage=259. See

[1002].

Yuan:2012:ADM

[1022] Xiaoming Yuan. Alternating direction method for covariance selection models. Journal of Scientific Computing, 51(2):261–273, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9507-1;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9507-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=2&spage=261-273.

Zhou:2012:GMS

[1023] Tao Zhou and Tao Tang. Galerkin methods for stochastic hyperbolic problems using bi-orthogonal polynomials. Journal of Scientific Computing, 51(2):274–292, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-011-9508-0;   http://link.springer.com/
content/pdf/10.1007/s10915-011-9508-0;   http://www.springerlink.
com/openurl.asp?genre=article&issn=0885-7474&volume=51&issue=
2&spage=274-292.

Pulch:2012:GPC

[1024] Roland Pulch and Dongbin Xiu. Generalised polynomial chaos for a class of linear conservation laws. Journal of Scientific Computing, 51 (2):293–312, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9511-5;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9511-5;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=2&
spage=293-312.

Alton:2012:OUM

[1025] Ken Alton and Ian M. Mitchell. An ordered upwind method with precomputed stencil and monotone node acceptance for solving static convex Hamilton–Jacobi equations. Journal of Scientific Computing, 51 (2):313–348, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9512-4;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9512-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=2&
spage=313-348.

Bernardi:2012:PAP

[1026] Christine Bernardi, Tomás Chacón Rebollo, and Marco Restelli. A posteriori analysis of a positive streamwise invariant discretization of a convection-diffusion equation. Journal of Scientific Computing, 51 (2):349–374, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9514-2;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9514-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=2&
spage=349-374.

Katz:2012:ECM

[1027] Aaron Katz and Venkateswaran Sankaran. An efficient correction method to obtain a formally third-order accurate flow solver for node-centered unstructured grids. Journal of Scientific Computing, 51(2):375–393, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9515-1;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9515-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=2&spage=375-393.

Gopalakrishnan:2012:CSP

[1028] J. Gopalakrishnan and M. Oh. Commuting smoothed projectors in weighted norms with an application to axisymmetric Maxwell equations. Journal of Scientific Computing, 51(2):394–420, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9513-3;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9513-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=2&spage=394-420.

Panigrahi:2012:REI

[1029] Bijaya Laxmi Panigrahi and Gnaneshwar Nelakanti. Richardson extrapolation of iterated discrete Galerkin method for eigenvalue problem of a two dimensional compact integral operator. Journal of Scientific Computing, 51(2):421–448, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9516-0;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9516-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=51&issue=2&spage=421-448.

Feng:2012:NMW

[1030] Hui Feng, Fuxing Hu, and Rong Wang. A new mapped weighted essentially non-oscillatory scheme. Journal of Scientific Computing, 51 (2):449–473, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9518-y;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9518-y;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=2&
spage=449-473.

Llanas:2012:EDA

[1031] Bernardo Llanas and Sagrario Lantarón. Edge detection by adaptive splitting II. The three-dimensional case. Journal of Scientific Computing, 51(2):474–503, May 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9517-z;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9517-z;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=2&
spage=474-503.

Chen:2012:NTS

[1032] Dai-Qiang Chen, Li-Zhi Cheng, and Fang Su. A new TV-Stokes model with augmented Lagrangian method for image denoising and deconvolution. Journal of Scientific Computing, 51(3):505–526, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9519-x;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9519-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=3&spage=505-526.

Li:2012:HWB

[1033] Gang Li, Changna Lu, and Jianxian Qiu. Hybrid well-balanced WENO schemes with different indicators for shallow water equations. Journal of Scientific Computing, 51(3):527–559, June 2012. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9520-4;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9520-
4;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=51&issue=3&spage=527-559.

vandenDoel:2012:CNF

[1034] Kees van den Doel and Uri Ascher. The chaotic nature of faster gradient descent methods. Journal of Scientific Computing, 51(3):560–581, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9521-3;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9521-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=3&spage=560-581.

Cockburn:2012:PEE

[1035] Bernardo Cockburn and Wujun Zhang. A posteriori error estimates for HDG methods. Journal of Scientific Computing, 51(3):582–607, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9522-2;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9522-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=3&spage=582-607.

Petersen:2012:HTF

[1036] A. Petersen, A. Gelb, and R. Eubank. Hypothesis testing for Fourier based edge detection methods. Journal of Scientific Computing, 51 (3):608–630, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9523-1;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9523-1;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=51&issue=3&
spage=608-630.

Viswanathan:2012:IDC

[1037] Adityavikram Viswanathan, Anne Gelb, and Douglas Cochran. Iterative design of concentration factors for jump detection. Journal of Scientific Computing, 51(3):631–649, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9524-0;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9524-
0;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=51&issue=3&spage=631-649.

Mattsson:2012:SPO

[1038] Ken Mattsson. Summation by parts operators for finite difference approximations of second-derivatives with variable coefficients. Journal of Scientific Computing, 51(3):650–682, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9525-z;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9525-
z;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=51&issue=3&spage=650-682. See erratum

[1042].

Chen:2012:MSM

[1039] Xinjuan Chen and Jae-Hun Jung. Matrix stability of multiquadric radial basis function methods for hyperbolic equations with uniform centers. Journal of Scientific Computing, 51(3):683–702, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9526-y;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9526-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=3&spage=683-702.

Sjogreen:2012:HOF

[1040] Björn Sjögreen. High order finite difference and finite volume methods for advection on the sphere. Journal of Scientific Computing, 51(3):703–732, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9527-x;   http://link.springer.com/content/pdf/10.1007/

s10915-011-9527-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=3&spage=703-732.

Sun:2012:RSB

[1041] Yuanchang Sun and Jack Xin. A recursive sparse blind source separation method and its application to correlated data in NMR spectroscopy of biofluids. Journal of Scientific Computing, 51(3):733–753, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9528-9;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9528-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=3&spage=733-753.

Mattsson:2012:ESP

[1042] Ken Mattsson. Erratum to: Summation by Parts Operators for Finite Difference Approximations of Second-Derivatives with Variable Coefficients. Journal of Scientific Computing, 51(3):754, June 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9568-1;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9568-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=51&issue=3&spage=754. See

[1038].

Zhang:2012:LBM

[1043] Jianying Zhang and Guangwu Yan. A lattice Boltzmann model for the reaction–diffusion equations with higher-order accuracy. Journal of Scientific Computing, 52(1):1–16, July 2012. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9530-2;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9530-
2;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=1&spage=1-16.

Sjogreen:2012:FOA

[1044] Björn Sjögreen and N. Anders Petersson. A fourth order accurate finite difference scheme for the elastic wave equation in second order formulation. Journal of Scientific Computing, 52(1):17–48, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9531-1;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9531-1;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=1&spage=17-48.

Celiker:2012:LFO

[1045] Fatih Celiker, Li Fan, Sheng Zhang, and Zhimin Zhang. Lockingfree optimal discontinuous Galerkin methods for a Naghdi-type arch model. Journal of Scientific Computing, 52(1):49–84, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9532-0;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9532-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=1&spage=49-84.

Mirzaee:2012:EIS

[1046] Hanieh Mirzaee, Jennifer K. Ryan, and Robert M. Kirby. Efficient implementation of smoothness-increasing accuracy–conserving (SIAC) filters for discontinuous Galerkin solutions. Journal of Scientific Computing, 52(1):85–112, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9535-x;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9535-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=1&
spage=85-112.

Adjerid:2012:SLD

[1047] Slimane Adjerid and Mahboub Baccouch. A superconvergent local discontinuous Galerkin method for elliptic problems. Journal of Scientific Computing, 52(1):113–152, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9537-8;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9537-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=1&spage=113-152.

Cliffe:2012:APE

[1048] K. Andrew Cliffe, Edward J. C. Hall, and Paul Houston. Adaptivity and a posteriori error control for bifurcation problems III: Incompressible fluid flow in open systems with O(2) symmetry. Journal of Scientific Computing, 52(1):153–179, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9545-8;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9545-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=1&spage=153-179.

Biezuner:2012:CFE

[1049] Rodney Josué Biezuner, Jed Brown, and Grey Ercole. Computing the first eigenpair of the p-Laplacian via inverse iteration of sublinear supersolutions. Journal of Scientific Computing, 52(1):180–201, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9540-0;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9540-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=1&spage=180-201.

Aksoylu:2012:GOA

[1050] Burak Aksoylu, Stephen D. Bond, and Eric C. Cyr. Goal-oriented adaptivity and multilevel preconditioning for the Poisson–Boltzmann equation. Journal of Scientific Computing, 52(1):202–225, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9539-6;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9539-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=1&spage=202-225.

Wang:2012:LGR

[1051] Zhong qing Wang and Ben yu Guo. Legendre–Gauss–Radau collocation method for solving initial value problems of first order ordinary differential equations. Journal of Scientific Computing, 52(1):226–255, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9538-7;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9538-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=1&spage=226-255.

Cockburn:2012:DFH

[1052] Bernardo Cockburn and Jintao Cui. Divergence-free HDG methods for the vorticity–velocity formulation of the Stokes problem. Journal of Scientific Computing, 52(1):256–270, July 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9542-y;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9542-
y;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=1&spage=256-270.

Bernauer:2012:IXF

[1053] Martin K. Bernauer and Roland Herzog. Implementation of an X–FEM solver for the classical two-phase Stefan problem. Journal of Scientific Computing, 52(2):271–293, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9543-x;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9543-
x;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=2&spage=271-293.

Goudenege:2012:HOF

[1054] Ludovic Goudenège, Daniel Martin, and Grégory Vial. High order finite element calculations for the Cahn–Hilliard equation. Journal of Scientific Computing, 52(2):294–321, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9546-7;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9546-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=2&spage=294-321.

Bustinza:2012:EEL

[1055] Rommel Bustinza and Francisco-Javier Sayas. Error estimates for an LDG method applied to Signorini type problems. Journal of Scientific Computing, 52(2):322–339, August 2012. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9548-5;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9548-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=2&spage=322-339.

DElia:2012:VDA

[1056] Marta D’Elia, Mauro Perego, and Alessandro Veneziani. A variational data assimilation procedure for the incompressible Navier–Stokes equations in hemodynamics. Journal of Scientific Computing, 52(2): 340–359, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9547-6;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9547-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=2&
spage=340-359.

Luo:2012:FSM

[1057] Songting Luo and Jianliang Qian. Fast sweeping methods for factored anisotropic eikonal equations: Multiplicative and additive factors. Journal of Scientific Computing, 52(2):360–382, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9550-y;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9550-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=2&spage=360-382.

Boffi:2012:LMC

[1058] D. Boffi, N. Cavallini, F. Gardini, and L. Gastaldi. Local mass conservation of Stokes finite elements. Journal of Scientific Computing, 52 (2):383–400, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9549-4;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9549-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=2&
spage=383-400.

Kumar:2012:ESN

[1059] Harish Kumar and Siddhartha Mishra. Entropy stable numerical schemes for two-fluid plasma equations. Journal of Scientific Computing, 52(2): 401–425, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9554-7;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9554-7;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=2&
spage=401-425.

Hu:2012:HAP

[1060] Jun Hu, Yunqing Huang, and Quan Shen. A high accuracy postprocessing algorithm for the eigenvalues of elliptic operators. Journal of Scientific Computing, 52(2):426–445, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-011-9552-9;
 http://link.springer.com/content/pdf/10.1007/s10915-011-9552-
9;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=52&issue=2&spage=426-445.

Schauer:2012:PCD

[1061] Marco Schauer, Jose E. Roman, and Enrique S. Quintana-Ortí. Parallel computation of 3-D soil-structure interaction in time domain with a coupled FEM/SBFEM approach. Journal of Scientific Computing, 52 (2):446–467, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9551-x;   http://link.springer.com/content/

pdf/10.1007/s10915-011-9551-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=2&
spage=446-467.

Dietrich:2012:PUM

[1062] J. C. Dietrich, S. Tanaka, J. J. Westerink, and C. N. Dawson. Performance of the unstructured-mesh, SWAN+ADCIRC model in computing hurricane waves and surge. Journal of Scientific Computing, 52 (2):468–497, August 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9555-6;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9555-6;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=2&
spage=468-497.

Chen:2012:NSE

[1063] Feng Chen, Jie Shen, and Haijun Yu. A new spectral element method for pricing European options under the Black–Scholes and Merton jump diffusion models. Journal of Scientific Computing, 52(3):499–518, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9556-5;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9556-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=3&spage=499-518.

Feng:2012:NQF

[1064] Xinlong Feng, Rongfei Li, Yinnian He, and Demin Liu. P_{1}-nonconforming quadrilateral finite volume methods for the semilinear elliptic equations. Journal of Scientific Computing, 52(3):519–545, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9557-4;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9557-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=3&spage=519-545.

Chen:2012:LES

[1065] Wenbin Chen, Sidafa Conde, Cheng Wang, and Xiaoming Wang. A linear energy stable scheme for a thin film model without slope selection. Journal of Scientific Computing, 52(3):546–562, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9559-2;   http://link.springer.com/content/pdf/10.1007/

s10915-011-9559-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=3&spage=546-562.

Pan:2012:MEM

[1066] Hao Pan and Hongxing Rui. Mixed element method for two-dimensional Darcy–Forchheimer model. Journal of Scientific Computing, 52(3):563– 587, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9558-3;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9558-3;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=3&
spage=563-587.

Zhao:2012:EAC

[1067] Tinggang Zhao, Yujiang Wu, and Heping Ma. Error analysis of Chebyshev–Legendre pseudo-spectral method for a class of nonclassical parabolic equation. Journal of Scientific Computing, 52(3):588– 602, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9560-9;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9560-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=3&
spage=588-602.

Bjontegaard:2012:SAP

[1068] Tormod Bjøntegaard, Einar M. Rønquist, and Øystein Tråsdahl. Spectral approximation of partial differential equations in highly distorted domains. Journal of Scientific Computing, 52(3):603–618, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9561-8;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9561-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=3&spage=603-618.

deFrutos:2012:STG

[1069] Javier de Frutos, Bosco García-Archilla, and Julia Novo. Static twogrid mixed finite-element approximations to the Navier–Stokes equations. Journal of Scientific Computing, 52(3):619–637, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9562-7;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9562-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=3&spage=619-637.

Zhang:2012:FTE

[1070] Mengping Zhang and Jue Yan. Fourier type error analysis of the direct discontinuous Galerkin method and its variations for diffusion equations. Journal of Scientific Computing, 52(3):638–655, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9564-5;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9564-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=3&spage=638-655.

Zhang:2012:NSB

[1071] Jianying Zhang and Guangwu Yan. Numerical studies based on higherorder accuracy lattice Boltzmann model for the complex Ginzburg– Landau equation. Journal of Scientific Computing, 52(3):656–674, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9565-4;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9565-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=52&issue=3&
spage=656-674.

Botti:2012:IRP

[1072] Lorenzo Botti. Influence of reference-to-physical frame mappings on approximation properties of discontinuous piecewise polynomial spaces. Journal of Scientific Computing, 52(3):675–703, September 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9566-3;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9566-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=52&issue=3&spage=675-703.

Gottlieb:2012:P

[1073] Sigal Gottlieb and Adi Ditkowski. Preface. Journal of Scientific Computing, 53(1):1–2, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9613-8;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9613-8;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=1&
spage=1-2.

Chen:2012:APM

[1074] Qingshan Chen, Youngjoon Hong, and Roger Temam. Analysis of a penalty method. Journal of Scientific Computing, 53(1):3–34, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9553-8;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9553-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=3-34.

Corem:2012:NAD

[1075] Neta Corem and Adi Ditkowski. New analysis of the Du Fort– Frankel methods. Journal of Scientific Computing, 53(1):35–54, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9627-2;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9627-2;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=35-54.

Fishelov:2012:RAS

[1076] D. Fishelov, M. Ben-Artzi, and J.-P. Croisille. Recent advances in the study of a fourth-order compact scheme for the one-dimensional biharmonic equation. Journal of Scientific Computing, 53(1):55–79, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9611-x;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9611-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=55-79.

Gao:2012:HOW

[1077] Zhen Gao, Wai Sun Don, and Zhiqiu Li. High order weighted essentially non-oscillation schemes for two-dimensional detonation wave simulations. Journal of Scientific Computing, 53(1):80–101, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9569-0;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9569-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=80-101.

Gottlieb:2012:SCA

[1078] Sigal Gottlieb and Cheng Wang. Stability and convergence analysis of fully discrete Fourier collocation spectral method for 3-D viscous Burgers’ equation. Journal of Scientific Computing, 53(1):102–128, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9621-8;   http://link.springer.com/content/pdf/10.1007/

s10915-012-9621-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=102-128.

King:2012:SIA

[1079] James King, Hanieh Mirzaee, and Jennifer K. Ryan. Smoothnessincreasing accuracy-conserving (SIAC) filtering for discontinuous Galerkin solutions: Improved errors versus higher-order accuracy. Journal of Scientific Computing, 53(1):129–149, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9593-8;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9593-
8;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=53&issue=1&spage=129-149.

Medvinsky:2012:MDP

[1080] M. Medvinsky, S. Tsynkov, and E. Turkel. The method of difference potentials for the Helmholtz equation using compact high order schemes. Journal of Scientific Computing, 53(1):150–193, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9602-y;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9602-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=150-193.

See erratum [1095].

Schuss:2012:NEP

[1081] Z. Schuss. The narrow escape problem — a short review of recent results. Journal of Scientific Computing, 53(1):194–210, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9590-y;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9590-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=194-210.

Tal-Ezer:2012:NHA

[1082] Hillel Tal-Ezer, Ronnie Kosloff, and Ido Schaefer. New, highly accurate propagator for the linear and nonlinear Schrödinger equation. Journal of Scientific Computing, 53(1):211–221, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9583-x;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9583-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=211-221.

Xiong:2012:WSS

[1083] Tao Xiong, Chi-Wang Shu, and Mengping Zhang. WENO scheme with subcell resolution for computing nonconservative Euler equations with applications to one-dimensional compressible two-medium flows. Journal of Scientific Computing, 53(1):222–247, October 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9578-7;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9578-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=1&spage=222-247.

Lin:2012:IAE

[1084] Huimin Lin, Zhenli Xu, Huazhong Tang, and Wei Cai. Image approximations to electrostatic potentials in layered electrolytes/dielectrics and an ion-channel model. Journal of Scientific Computing, 53(2): 249–267, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9567-2;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9567-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=2&
spage=249-267.

Chun:2012:MMF

[1085] Sehun Chun. Method of moving frames to solve conservation laws on curved surfaces. Journal of Scientific Computing, 53(2):268–294, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9570-7;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9570-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=268-294.

See [1094].

Kumar:2012:NLS

[1086] N. Kishore Kumar and G. Naga Raju. Nonconforming least-squares method for elliptic partial differential equations with smooth interfaces. Journal of Scientific Computing, 53(2):295–319, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9572-5;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9572-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=295-319.

Belgacem:2012:LCL

[1087] Faker Ben Belgacem, Duc Thang Du, and Faten Jelassi. Local convergence of the Lavrentiev method for the Cauchy problem via a Carleman inequality. Journal of Scientific Computing, 53(2):320–341, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
011-9571-6;   http://link.springer.com/content/pdf/10.1007/
s10915-011-9571-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=320-341.

Kestler:2012:AWM

[1088] Sebastian Kestler and Karsten Urban. Adaptive wavelet methods on unbounded domains. Journal of Scientific Computing, 53(2):342– 376, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-011-9573-4;   http://link.springer.com/content/
pdf/10.1007/s10915-011-9573-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=2&
spage=342-376.

Huang:2012:MNF

[1089] Peiqi Huang, Jinru Chen, and Mingchao Cai. A mixed and nonconforming FEM with nonmatching meshes for a coupled Stokes–Darcy model. Journal of Scientific Computing, 53(2):377–394, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9574-y;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9574-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=377-394.

Hao:2012:CAB

[1090] Wenrui Hao, Jonathan D. Hauenstein, Bei Hu, and Yuan Liu. Continuation along bifurcation branches for a tumor model with a necrotic core. Journal of Scientific Computing, 53(2):395–413, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9575-x;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9575-x;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=395-413.

Xie:2012:CAS

[1091] Ziqing Xie, Xianjuan Li, and Tao Tang. Convergence analysis of spectral Galerkin methods for Volterra type integral equations. Journal of Scientific Computing, 53(2):414–434, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9577-8;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9577-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=414-434.

Li:2012:FPI

[1092] Chenliang Li and Zhonghua Qiao. A fast preconditioned iterative algorithm for the electromagnetic scattering from a large cavity. Journal of Scientific Computing, 53(2):435–450, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9580-0;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9580-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=435-450.

Zhang:2012:DDS

[1093] Chao Zhang and Ben yu Guo. Domain decomposition spectral method for mixed inhomogeneous boundary value problems of high order differential equations on unbounded domains. Journal of Scientific Computing, 53 (2):451–480, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9581-z;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9581-z;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=2&
spage=451-480.

Chun:2012:EMM

[1094] Sehun Chun. Erratum to: Method of Moving Frames to Solve Conservation Laws on Curved Surfaces. Journal of Scientific Computing, 53(2):481, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9604-9;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9604-9;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=2&

spage=481. See [1085].

Medvinsky:2012:EMD

[1095] M. Medvinsky, S. Tsynkov, and E. Turkel. Erratum to: The Method of Difference Potentials for the Helmholtz Equation Using Compact High Order Schemes. Journal of Scientific Computing, 53(2):482, November 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9638-z;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9638-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=2&spage=482. See

[1080].

Leykekhman:2012:ICP

[1096] Dmitriy Leykekhman. Investigation of commutative properties of discontinuous Galerkin methods in PDE constrained optimal control problems. Journal of Scientific Computing, 53(3):483–511, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9582-y;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9582-y;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=3&spage=483-511.

Hongying:2012:CLD

[1097] Huang Hongying, Yang Ju’e, and Yu Dehao. A coupling of local discontinuous Galerkin and natural boundary element method for exterior problems. Journal of Scientific Computing, 53(3):512–527, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9584-9;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9584-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=3&spage=512-527.

Nissen:2012:SNG

[1098] A. Nissen, G. Kreiss, and M. Gerritsen. Stability at nonconforming grid interfaces for a high order discretization of the Schrödinger equation. Journal of Scientific Computing, 53(3):528–551, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9586-7;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9586-7;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=3&spage=528-551.

Boonkkamp:2012:ECF

[1099] J. H. M. ten Thije Boonkkamp, J. van Dijk, and L. Liu. Extension of the complete flux scheme to systems of conservation laws. Journal of Scientific Computing, 53(3):552–568, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9588-5;   http://link.springer.com/content/pdf/10.1007/

s10915-012-9588-5;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=3&spage=552-568.

Lai:2012:NSN

[1100] Huilin Lai and Changfeng Ma. Numerical study of the nonlinear combined sine-cosine–Gordon equation with the lattice Boltzmann method. Journal of Scientific Computing, 53(3):569–585, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9587-6;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9587-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=53&issue=3&spage=569-585.

Ledoux:2012:IQR

[1101] Veerle Ledoux and Marnix Van Daele. Interpolatory quadrature rules for oscillatory integrals. Journal of Scientific Computing, 53(3):586– 607, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9589-4;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9589-4;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=3&
spage=586-607.

Acary-Robert:2012:WBF

[1102] C. Acary-Robert, E. D. Fernández-Nieto, and G. Narbona-Reina. A well-balanced finite volume-augmented Lagrangian method for an integrated Herschel–Bulkley model. Journal of Scientific Computing, 53(3): 608–641, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9591-x;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9591-x;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=3&
spage=608-641.

Duru:2012:ASP

[1103] Kenneth Duru and Gunilla Kreiss. On the accuracy and stability of the perfectly matched layer in transient waveguides. Journal of Scientific Computing, 53(3):642–671, December 2012. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9594-7;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9594-
7;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=53&issue=3&spage=642-671.

Wei:2012:LSC

[1104] Yunxia Wei and Yanping Chen. Legendre spectral collocation methods for pantograph Volterra delay–integro-differential equations. Journal of Scientific Computing, 53(3):672–688, December 2012. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9595-6;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9595-
6;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=53&issue=3&spage=672-688.

Epshteyn:2012:UDP

[1105] Yekaterina Epshteyn. Upwind-difference potentials method for Patlak– Keller–Segel chemotaxis model. Journal of Scientific Computing, 53(3): 689–713, December 2012. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9599-2;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9599-2;   http://www.springerlink.com/
openurl.asp?genre=article&issn=0885-7474&volume=53&issue=3&
spage=689-713.

Nejat:2013:CSC

[1106] Amir Nejat and Vahid Abdollahi. A critical study of the compressible lattice Boltzmann methods for Riemann problem. Journal of Scientific Computing, 54(1):1–20, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9596-5;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9596-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=54&issue=1&spage=1-20.

Gassner:2013:ADD

[1107] Gregor J. Gassner. An analysis of the dissipation and dispersion errors of the P_{N}P_{M} schemes. Journal of Scientific Computing, 54(1):21–44, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9600-0;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9600-0;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=54&issue=1&spage=21-44.

Yadav:2013:SDG

[1108] Sangita Yadav, Amiya K. Pani, and Neela Nataraj. Superconvergent discontinuous Galerkin methods for linear non-selfadjoint and indefinite elliptic problems. Journal of Scientific Computing, 54(1):45–76, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9601-z;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9601-z;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=54&issue=1&spage=45-76.

Zou:2013:NHE

[1109] Qingsong Zou. A novel hierarchial error estimate for elliptic obstacle problems. Journal of Scientific Computing, 54(1):77–96, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9605-8;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9605-8;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=54&issue=1&spage=77-96.

Zhai:2013:FFO

[1110] Shuying Zhai, Xinlong Feng, and Yinnian He. A family of fourth-order and sixth-order compact difference schemes for the three-dimensional Poisson equation. Journal of Scientific Computing, 54(1):97–120, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9607-6;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9607-6;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=54&issue=1&spage=97-120.

Li:2013:NSP

[1111] Jichun Li, Yunqing Huang, and Wei Yang. Numerical study of the plasma–Lorentz model in metamaterials. Journal of Scientific Computing, 54(1):121–144, January 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9608-5;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9608-
5;   http://www.springerlink.com/openurl.asp?genre=article&
issn=0885-7474&volume=54&issue=1&spage=121-144.

Sidi:2013:CNQ

[1112] Avram Sidi. Compact numerical quadrature formulas for hypersingular integrals and integral equations. Journal of Scientific Computing, 54 (1):145–176, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9610-y;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9610-y;   http://www.springerlink.com/

openurl.asp?genre=article&issn=0885-7474&volume=54&issue=1&
spage=145-176.

Gudi:2013:AIP

[1113] Thirupathi Gudi, Hari Shanker Gupta, and Neela Nataraj. Analysis of an interior penalty method for fourth order problems on polygonal domains. Journal of Scientific Computing, 54(1):177–199, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9612-9;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9612-9;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=54&issue=1&spage=177-199.

Neilan:2013:QFE

[1114] Michael Neilan. Quadratic finite element approximations of the Monge– Ampère equation. Journal of Scientific Computing, 54(1):200–226, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9617-4;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9617-4;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=54&issue=1&spage=200-226.

Yu:2013:CBC

[1115] Meilin Yu and Z. J. Wang. On the connection between the correction and weighting functions in the correction procedure via reconstruction method. Journal of Scientific Computing, 54(1):227–244, January 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9618-3;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9618-3;   http://www.springerlink.com/openurl.asp?
genre=article&issn=0885-7474&volume=54&issue=1&spage=227-244.

Shu:2013:P

[1116] Chi-Wang Shu. Preface. Journal of Scientific Computing, 54(2–3): 245–246, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9673-9;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9673-9.pdf.

Ariel:2013:MMH

[1117] G. Ariel, B. Engquist, S. Kim, Y. Lee, and R. Tsai. A multiscale method for highly oscillatory dynamical systems using a Poincaré map type technique. Journal of Scientific Computing, 54(2–3):247–268, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9656-x;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9656-x.pdf.

Benning:2013:HOT

[1118] Martin Benning, Christoph Brune, and Martin Burger …. Higherorder TV methods — enhancement via Bregman iteration. Journal of Scientific Computing, 54(2–3):269–310, February 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9650-3;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9650-
3.pdf.

Bihari:2013:TMU

[1119] Barna L. Bihari. Transactional memory for unstructured mesh simulations. Journal of Scientific Computing, 54(2–3):311–332, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9643-2;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9643-2.pdf.

Dong:2013:XRC

[1120] Bin Dong, Jia Li, and Zuowei Shen. X-ray CT image reconstruction via wavelet frame based regularization and Radon domain inpainting. Journal of Scientific Computing, 54(2–3):333–349, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9579-6;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9579-6.pdf.

Dong:2013:EAM

[1121] Bin Dong and Yong Zhang. An efficient algorithm for l_{0} minimization in wavelet frame based image restoration. Journal of Scientific Computing, 54(2–3):350–368, February 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9597-4;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9597-
4.pdf.

Gibou:2013:HRS

[1122] Frédéric Gibou, Chohong Min, and Ron Fedkiw. High resolution sharp computational methods for elliptic and parabolic problems in complex geometries. Journal of Scientific Computing, 54(2–3):369–413, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9660-1;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9660-1.pdf.

Ha:2013:SSA

[1123] Youngsoo Ha, Chang Ho Kim, and Myungjoo Kang. Simulations of supersonic astrophysical jets and their environments using level set methods. Journal of Scientific Computing, 54(2–3):414–427, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9606-7;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9606-7.pdf.

Huang:2013:ALB

[1124] Bo Huang, Shiqian Ma, and Donald Goldfarb. Accelerated linearized Bregman method. Journal of Scientific Computing, 54(2–3):428–453, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9592-9;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9592-9.pdf.

Jemison:2013:CLS

[1125] Matthew Jemison, Eva Loch, and Mark Sussman …. A coupled level set-moment of fluid method for incompressible two-phase flows. Journal of Scientific Computing, 54(2–3):454–491, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9614-7;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9614-7.pdf.

Kao:2013:ERA

[1126] Chiu-Yen Kao and Shu Su. Efficient rearrangement algorithms for shape optimization on elliptic eigenvalue problems. Journal of Scientific Computing, 54(2–3):492–512, February 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9629-0;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9629-
0.pdf.

Kostic:2013:SDE

[1127] Tijana Kostić and Andrea Bertozzi. Statistical density estimation using threshold dynamics for geometric motion. Journal of Scientific Computing, 54(2–3):513–530, February 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9615-6;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9615-
6.pdf.

Ji:2013:NON

[1128] Liangyue Ji, Yan Xu, and Jennifer K. Ryan. Negative-order norm estimates for nonlinear hyperbolic conservation laws. Journal of Scientific Computing, 54(2–3):531–548, February 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9668-6;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9668-
6.pdf.

Langer:2013:BDD

[1129] Andreas Langer, Stanley Osher, and Carola-Bibiane Schönlieb. Bregmanized domain decomposition for image restoration. Journal of Scientific Computing, 54(2–3):549–576, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9603-x;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9603-
x.pdf.

Liang:2013:REI

[1130] Jian Liang, Frederick Park, and Hongkai Zhao. Robust and efficient implicit surface reconstruction for point clouds based on convexified image segmentation. Journal of Scientific Computing, 54(2–3):577–602, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9674-8;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9674-8.pdf.

Liu:2013:RRU

[1131] Yuan Liu and Yong-Tao Zhang. A robust reconstruction for unstructured WENO schemes. Journal of Scientific Computing, 54(2–3):603– 621, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9598-3;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9598-3.pdf.

Takei:2013:OTC

[1132] Ryo Takei and Richard Tsai. Optimal trajectories of curvature constrained motion in the Hamilton–Jacobi formulation. Journal of Scientific Computing, 54(2–3):622–644, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9671-y;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9671-
y.pdf.

Xing:2013:HOW

[1133] Yulong Xing and Chi-Wang Shu. High order well-balanced WENO scheme for the gas dynamics equations under gravitational fields. Journal of Scientific Computing, 54(2–3):645–662, February 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9585-8;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9585-
8.pdf.

Yan:2013:NND

[1134] Jue Yan. A new nonsymmetric discontinuous Galerkin method for time dependent convection diffusion equations. Journal of Scientific Computing, 54(2–3):663–683, February 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9637-0;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9637-
0.pdf.

Yin:2013:EFB

[1135] Wotao Yin and Stanley Osher. Error forgetting of Bregman iteration. Journal of Scientific Computing, 54(2–3):684–695, February 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9616-5;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9616-5.pdf.

Theillard:2013:MMN

[1136] Maxime Theillard, Chris H. Rycroft, and Frédéric Gibou. A multigrid method on non-graded adaptive octree and quadtree Cartesian grids. Journal of Scientific Computing, 55(1):1–15, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9619-2;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9619-2.pdf.

Xu:2013:BMS

[1137] Kuan Xu and Shidong Jiang. A bootstrap method for sum-of-poles approximations. Journal of Scientific Computing, 55(1):16–39, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9620-9;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9620-9.pdf.

Arun:2013:CBG

[1138] K. R. Arun and M. Lukácová-Medvidová. A characteristics based genuinely multidimensional discrete Kinetic scheme for the Euler equations. Journal of Scientific Computing, 55(1):40–64, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9623-6;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9623-6.pdf.

Hammouch:2013:ASE

[1139] Zohra Hammouch, Gérard Labrosse, and Serge Gauthier. Anelastic Stokes eigenmodes in infinite channel. Journal of Scientific Computing, 55(1):65–91, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9622-7;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9622-7.pdf.

Kozdon:2013:SDE

[1140] Jeremy E. Kozdon, Eric M. Dunham, and Jan Nordström. Simulation of dynamic earthquake ruptures in complex geometries using high-order finite difference methods. Journal of Scientific Computing, 55(1):92–124, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9624-5;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9624-5.pdf.

Hu:2013:COA

[1141] Jun Hu and Jinchao Xu. Convergence and optimality of the adaptive nonconforming linear element method for the Stokes problem. Journal of Scientific Computing, 55(1):125–148, April 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9625-4;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9625-
4.pdf.

Griebel:2013:FAD

[1142] Michael Griebel and Daniel Wissel. Fast approximation of the discrete Gauss transform in higher dimensions. Journal of Scientific Computing, 55(1):149–172, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9626-3;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9626-3.pdf.

Nissen:2013:HOS

[1143] A. Nissen, G. Kreiss, and M. Gerritsen. High order stable finite difference methods for the Schrödinger equation. Journal of Scientific Computing, 55(1):173–199, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9628-1;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9628-1.pdf.

Davis:2013:MPM

[1144] Christopher Davis, June G. Kim, and Hae-Soo Oh …. Meshfree particle methods in the framework of boundary element methods for the Helmholtz equation. Journal of Scientific Computing, 55(1):200–230, April 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9645-0;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9645-0.pdf.

Besse:2013:ENM

[1145] Christophe Besse, Fabrice Deluzet, and Claudia Negulescu …. Efficient numerical methods for strongly anisotropic elliptic equations. Journal of Scientific Computing, 55(1):231–254, April 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9630-7;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9630-
7.pdf.

Song:2013:ALP

[1146] Lina Song, Yanren Hou, and Haibiao Zheng. Adaptive local postprocessing finite element method for the Navier–Stokes equations. Journal of Scientific Computing, 55(2):255–267, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9631-6;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9631-
6.pdf.

Berthon:2013:STG

[1147] Christophe Berthon, Céline Sarazin, and Rodolphe Turpault. Space-time generalized Riemann problem solvers of order k for linear advection with unrestricted time step. Journal of Scientific Computing, 55(2):268–308, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9632-5;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9632-5.pdf.

Sangawi:2013:RSM

[1148] Ali W. K. Sangawi, Ali H. M. Murid, and M. M. S. Nasser. Radial slit maps of bounded multiply connected regions. Journal of Scientific Computing, 55(2):309–326, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9634-3;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9634-
3.pdf.

Sun:2013:GJA

[1149] Tao Sun and Ben yu Guo. Generalized Jacobi approximation in multiple dimensions and its applications. Journal of Scientific Computing, 55(2):327–350, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9633-4;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9633-4.pdf.

Gao:2013:MHC

[1150] Zhen Gao and Wai Sun Don. Mapped hybrid central-WENO finite difference scheme for detonation waves simulations. Journal of Scientific Computing, 55(2):351–371, May 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9635-2;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9635-
2.pdf.

Bao:2013:NSK

[1151] Gang Bao, Guanghui Hu, and Di Liu. Numerical solution of the Kohn– Sham equation by finite element methods with an adaptive mesh redistribution technique. Journal of Scientific Computing, 55(2):372–391, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9636-1;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9636-1.pdf.

Cesmelioglu:2013:AHM

[1152] Aycil Cesmelioglu, Bernardo Cockburn, and Ngoc Cuong Nguyen …. Analysis of HDG methods for Oseen equations. Journal of Scientific Computing, 55(2):392–431, May 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9639-y;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9639-
y.pdf.

Fang:2013:FER

[1153] Jun Fang, Xingyu Gao, and Aihui Zhou. A finite element recovery approach to eigenvalue approximations with applications to electronic structure calculations. Journal of Scientific Computing, 55(2):432–454, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9640-5;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9640-5.pdf.

Shen:2013:FEC

[1154] Lihua Shen, Jack Xin, and Aihui Zhou. Finite element computation of KPP front speeds in cellular and cat’s eye flows. Journal of Scientific Computing, 55(2):455–470, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9641-4;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9641-
4.pdf.

Congreve:2013:TGH

[1155] Scott Congreve, Paul Houston, and Thomas P. Wihler. Two-grid hp-version discontinuous Galerkin finite element methods for secondorder quasilinear elliptic PDEs. Journal of Scientific Computing, 55 (2):471–497, May 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9644-1;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9644-1.pdf.

Balbas:2013:NOC

[1156] Jorge Balbás and Smadar Karni. A non-oscillatory central scheme for one-dimensional two-layer shallow water flows along channels with varying width. Journal of Scientific Computing, 55(3):499–528, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

012-9642-3;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9642-3.pdf.

Huang:2013:CEE

[1157] Tsung-Ming Huang, Yueh-Cheng Kuo, and Weichung Wang. Computing extremal eigenvalues for three-dimensional photonic crystals with wave vectors near the Brillouin zone center. Journal of Scientific Computing, 55(3):529–551, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9646-z;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9646-z.pdf.

Yang:2013:DDE

[1158] He Yang, Fengyan Li, and Jianxian Qiu. Dispersion and dissipation errors of two fully discrete discontinuous Galerkin methods. Journal of Scientific Computing, 55(3):552–574, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9647-y;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9647-
y.pdf.

Bokanowski:2013:ASG

[1159] Olivier Bokanowski, Jochen Garcke, and Michael Griebel …. An adaptive sparse grid semi-Lagrangian scheme for first order Hamilton–Jacobi Bellman equations. Journal of Scientific Computing, 55(3):575–605, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9648-x;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9648-x.pdf.

Zhu:2013:WST

[1160] Jun Zhu and Jianxian Qiu. WENO schemes and their application as limiters for RKDG methods based on trigonometric approximation spaces. Journal of Scientific Computing, 55(3):606–644, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9649-9;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9649-9.pdf.

Chen:2013:NTP

[1161] Hong-Ru Chen, Shao-Chun Chen, and Zhong-Hua Qiao. C^{0}- nonconforming triangular prism elements for the three-dimensional fourth order elliptic problem. Journal of Scientific Computing, 55 (3):645–658, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9652-1;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9652-1.pdf.

Bi:2013:PEE

[1162] Chunjia Bi and Victor Ginting. A posteriori error estimates of discontinuous Galerkin method for nonmonotone quasi-linear elliptic problems. Journal of Scientific Computing, 55(3):659–687, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9651-2;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9651-2.pdf.

Abreu:2013:NMD

[1163] Eduardo Abreu and Duilio Conceição. Numerical modeling of degenerate equations in porous media flow. Journal of Scientific Computing, 55(3):688–717, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9653-0;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9653-0.pdf.

Chen:2013:RCM

[1164] Yanlai Chen and Sigal Gottlieb. Reduced collocation methods: Reduced basis methods in the collocation framework. Journal of Scientific Computing, 55(3):718–737, June 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9654-z;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9654-
z.pdf.

Brenner:2013:AFE

[1165] S. C. Brenner, J. Gedicke, and L.-Y. Sung. An adaptive P_{1} finite element method for two-dimensional Maxwell’s equations. Journal of Scientific Computing, 55(3):738–754, June 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9658-8;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9658-
8.pdf.

Tang:2013:HOC

[1166] Tao Tang, Hehu Xie, and Xiaobo Yin. High-order convergence of spectral deferred correction methods on general quadrature nodes.

Journal of Scientific Computing, 56(1):1–13, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9657-9;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9657-
9.pdf.

Antonietti:2013:MDE

[1167] Paola F. Antonietti, Nadia Bigoni, and Marco Verani. Mimetic discretizations of elliptic control problems. Journal of Scientific Computing, 56(1):14–27, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9659-7;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9659-7.pdf.

Constantinescu:2013:EMM

[1168] Emil M. Constantinescu and Adrian Sandu. Extrapolated multirate methods for differential equations with multiple time scales. Journal of Scientific Computing, 56(1):28–44, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9662-z;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9662-
z.pdf.

Zhou:2013:QCF

[1169] Han Zhou, WenYi Tian, and Weihua Deng. Quasi-compact finite difference schemes for space fractional diffusion equations. Journal of Scientific Computing, 56(1):45–66, July 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9661-0;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9661-
0.pdf.

Fang:2013:OLD

[1170] Jian Fang, Zhaorui Li, and Lipeng Lu. An optimized low-dissipation monotonicity-preserving scheme for numerical simulations of high-speed turbulent flows. Journal of Scientific Computing, 56(1):67–95, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9663-y;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9663-y.pdf.

Karageorghis:2013:EMA

[1171] A. Karageorghis and L. Marin. Efficient MFS algorithms for problems in thermoelasticity. Journal of Scientific Computing, 56(1):96–121, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9664-x;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9664-x.pdf.

Du:2013:NMR

[1172] Rui Du and Wenwen Liu. A new multiple-relaxation-time Lattice Boltzmann Method for natural convection. Journal of Scientific Computing, 56(1):122–130, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9665-9;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9665-9.pdf.

Goswami:2013:OEE

[1173] Deepjyoti Goswami, Amiya K. Pani, and Sangita Yadav. Optimal error estimates of two mixed finite element methods for parabolic integro-differential equations with nonsmooth initial data. Journal of Scientific Computing, 56(1):131–164, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9666-8;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9666-
8.pdf.

Yu:2013:CSS

[1174] P. X. Yu and Zhen F. Tian. A compact scheme for the streamfunctionvelocity formulation of the 2D steady incompressible Navier–Stokes equations in polar coordinates. Journal of Scientific Computing, 56 (1):165–189, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9667-7;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9667-7.pdf.

Descombes:2013:LIT

[1175] Stéphane Descombes, Stéphane Lanteri, and Ludovic Moya. Locally implicit time integration strategies in a discontinuous Galerkin method for Maxwell’s equations. Journal of Scientific Computing, 56(1):190–218, July 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9669-5;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9669-5.pdf.

Guzman:2013:NLB

[1176] Johnny Guzmán, Abner J. Salgado, and Francisco-Javier Sayas. A note on the Ladyzenskaja–Babuska–Brezzi condition. Journal of Scientific Computing, 56(2):219–229, August 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9670-z;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9670-
z.pdf.

Hegele:2013:RLB

[1177] L. A. Hegele, Jr., K. Mattila, and P. C. Philippi. Rectangular lattice-Boltzmann schemes with BGK-collision operator. Journal of Scientific Computing, 56(2):230–242, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9672-x;
 http://link.springer.com/content/pdf/10.1007/s10915-012-9672-
x.pdf.

Liu:2013:SAI

[1178] Jun Liu and Shingyu Leung. A splitting algorithm for image segmentation on manifolds represented by the Grid based particle method. Journal of Scientific Computing, 56(2):243–266, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9675-7;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9675-7.pdf.

Bollermann:2013:WBR

[1179] Andreas Bollermann, Guoxian Chen, Alexander Kurganov, and Sebastian Noelle. A well-balanced reconstruction of wet/dry fronts for the shallow water equations. Journal of Scientific Computing, 56(2):267–290, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9677-5;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9677-5.pdf.

Huang:2013:CDG

[1180] Xuehai Huang and Jianguo Huang. The compact discontinuous Galerkin method for nearly incompressible linear elasticity. Journal of Scientific Computing, 56(2):291–318, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-012-9676-6;

 http://link.springer.com/content/pdf/10.1007/s10915-012-9676-
6.pdf.

Cheng:2013:SCR

[1181] Yingda Cheng, Irene M. Gamba, and Philip J. Morrison. Study of conservation and recurrence of Runge–Kutta discontinuous Galerkin schemes for Vlasov–Poisson systems. Journal of Scientific Computing, 56(2): 319–349, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9680-x;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9680-x.pdf.

Hao:2013:CCC

[1182] Wenrui Hao, Bei Hu, and Andrew J. Sommese. Cell cycle control and bifurcation for a free boundary problem modeling tissue growth. Journal of Scientific Computing, 56(2):350–365, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9678-4;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9678-4.pdf.

Khanna:2013:HPN

[1183] Gaurav Khanna. High-precision numerical simulations on a CUDA GPU: Kerr black hole tails. Journal of Scientific Computing, 56(2): 366–380, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-012-9679-3;   http://link.springer.com/content/
pdf/10.1007/s10915-012-9679-3.pdf.

Ren:2013:NAH

[1184] Jincheng Ren and Zhi zhong Sun. Numerical algorithm with high spatial accuracy for the fractional diffusion-wave equation with Neumann boundary conditions. Journal of Scientific Computing, 56(2):381–408, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
012-9681-9;   http://link.springer.com/content/pdf/10.1007/
s10915-012-9681-9.pdf.

Wang:2013:MMS

[1185] Ming Wang and Long Chen. Multigrid methods for the Stokes equations using distributive Gauss–Seidel relaxations based on the least squares commutator. Journal of Scientific Computing, 56(2):409–431, August 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9684-1;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9684-1.pdf.

Yan:2013:ELR

[1186] Ming Yan, Yi Yang, and Stanley Osher. Exact low-rank matrix completion from sparsely corrupted entries via adaptive outlier pursuit. Journal of Scientific Computing, 56(3):433–449, September 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9682-3;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9682-3.pdf.

Qi:2013:WCF

[1187] Ruisheng Qi and Xiaoyuan Yang. Weak convergence of finite element method for stochastic elastic equation driven by additive noise. Journal of Scientific Computing, 56(3):450–470, September 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9683-2;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9683-2.pdf.

Martin:2013:PTR

[1188] David R. Martin and Lothar Reichel. Projected Tikhonov regularization of large-scale discrete ill-posed problems. Journal of Scientific Computing, 56(3):471–493, September 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9685-0;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9685-
0.pdf.

Chen:2013:FFG

[1189] Xiangling Chen, Rui Wang, and Yuesheng Xu. Fast Fourier– Galerkin methods for nonlinear boundary integral equations. Journal of Scientific Computing, 56(3):494–514, September 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9687-y;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9687-
y.pdf.

Kang:2013:ABM

[1190] Myeongmin Kang, Sangwoon Yun, Hyenkyun Woo, and Myungjoo Kang. Accelerated Bregman method for linearly constrained ℓ_{1}–ℓ_{2} minimization. Journal of Scientific Computing, 56(3):515–534, September 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9686-z;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9686-z.pdf.

Fuselier:2013:HOK

[1191] Edward J. Fuselier and Grady B. Wright. A high-order kernel method for diffusion and reaction–diffusion equations on surfaces. Journal of Scientific Computing, 56(3):535–565, September 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9688-x;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9688-x.pdf.

Cao:2013:SOF

[1192] Waixiang Cao, Zhimin Zhang, and Qingsong Zou. Superconvergence of any order finite volume schemes for 1D general elliptic equations. Journal of Scientific Computing, 56(3):566–590, September 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9691-2;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9691-2.pdf.

Le:2013:FMS

[1193] An Le, Zhi-Qiang Wang, and Jianxin Zhou. Finding multiple solutions to elliptic PDE with nonlinear boundary conditions. Journal of Scientific Computing, 56(3):591–615, September 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9689-9;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9689-
9.pdf.

Li:2013:AMM

[1194] Ruo Li and Shuonan Wu. H-adaptive mesh method with double tolerance adaptive strategy for hyperbolic conservation laws. Journal of Scientific Computing, 56(3):616–636, September 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9692-1;

 http://link.springer.com/content/pdf/10.1007/s10915-013-9692-
1.pdf.

Shi:2013:ANQ

[1195] Dong yang Shi, Chao Xu, and Jin huan Chen. Anisotropic nonconforming EQ_{1}^{rot} quadrilateral finite element approximation to second order elliptic problems. Journal of Scientific Computing, 56(3):637–653, September 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9690-3;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9690-3.pdf.

Banks:2013:REL

[1196] J. W. Banks and T. D. Aslam. Richardson extrapolation for linearly degenerate discontinuities. Journal of Scientific Computing, 57 (1):1–18, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9693-0;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9693-0.pdf.

Xing:2013:PPW

[1197] Yulong Xing and Xiangxiong Zhang. Positivity-preserving well-balanced discontinuous Galerkin methods for the shallow water equations on unstructured triangular meshes. Journal of Scientific Computing, 57(1): 19–41, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9695-y;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9695-y.pdf.

Kudryavtsev:2013:NMS

[1198] A. N. Kudryavtsev and A. A. Shershnev. A numerical method for simulation of microflows by solving directly kinetic equations with WENO schemes. Journal of Scientific Computing, 57(1):42–73, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9694-z;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9694-z.pdf.

Wu:2013:AAT

[1199] Gang Wu, Ying Zhang, and Yimin Wei. Accelerating the Arnolditype algorithm for the PageRank problem and the ProteinRank problem. Journal of Scientific Computing, 57(1):74–104, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9696-x;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9696-x.pdf.

Bacuta:2013:MGU

[1200] Constantin Bacuta and Lu Shu. Multilevel gradient Uzawa algorithms for symmetric saddle point problems. Journal of Scientific Computing, 57 (1):105–123, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9697-9;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9697-9.pdf.

He:2013:TLN

[1201] Yinnian He, Yan Zhang, and Hui Xu. Two-level Newton’s method for nonlinear elliptic PDEs. Journal of Scientific Computing, 57(1): 124–145, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9699-7;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9699-7.pdf.

Rubio:2013:ETE

[1202] G. Rubio, F. Fraysse, J. de Vicente, and E. Valero. The estimation of truncation error by τ-estimation for Chebyshev spectral collocation method. Journal of Scientific Computing, 57(1):146–173, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9698-8;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9698-8.pdf.

Marquez:2013:DPT

[1203] Antonio Márquez, Salim Meddahi, and Francisco-Javier Sayas. A decoupled preconditioning technique for a mixed Stokes–Darcy model. Journal of Scientific Computing, 57(1):174–192, October 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9700-5;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9700-
5.pdf.

Croisille:2013:HCI

[1204] Jean-Pierre Croisille. Hermitian compact interpolation on the cubedsphere grid. Journal of Scientific Computing, 57(1):193–212, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9702-3;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9702-3.pdf.

Ferretti:2013:SSG

[1205] Roberto Ferretti. Stability of some generalized Godunov schemes with linear high-order reconstructions. Journal of Scientific Computing, 57 (1):213–228, October 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9701-4;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9701-4.pdf.

Xiang:2013:GHA

[1206] Xin min Xiang and Zhong qing Wang. Generalized Hermite approximations and spectral method for partial differential equations in multiple dimensions. Journal of Scientific Computing, 57(2):229–253, November 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9703-2;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9703-2.pdf.

Malm:2013:SSE

[1207] Johan Malm, Philipp Schlatter, Paul F. Fischer, and Dan S. Henningson. Stabilization of the spectral element method in convection dominated flows by recovery of skew-symmetry. Journal of Scientific Computing, 57 (2):254–277, November 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9704-1;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9704-1.pdf.

Xu:2013:HOM

[1208] Xihua Xu, Guoxi Ni, and Song Jiang. A high-order moving mesh kinetic scheme based on WENO reconstruction for compressible flows on unstructured meshes. Journal of Scientific Computing, 57(2):278– 299, November 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9705-0;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9705-0.pdf.

Calabro:2013:NTE

[1209] Francesco Calabrò. Numerical treatment of elliptic problems nonlinearly coupled through the interface. Journal of Scientific Computing, 57(2):

300–312, November 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9706-z;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9706-z.pdf.

Cangiani:2013:SCD

[1210] Andrea Cangiani, John Chapman, Emmanuil Georgoulis, and Max Jensen. On the stability of continuous–discontinuous Galerkin methods for advection–diffusion-reaction problems. Journal of Scientific Computing, 57(2):313–330, November 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9707-y;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9707-
y.pdf.

Wu:2013:EEF

[1211] Xinming Wu and Wenbin Chen. Error estimates of the finite element method for interior transmission problems. Journal of Scientific Computing, 57(2):331–348, November 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9708-x;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9708-
x.pdf.

Li:2013:IRT

[1212] Fang Li and Tieyong Zeng. Image restoration via tight frame regularization and local constraints. Journal of Scientific Computing, 57(2): 349–371, November 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9709-9;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9709-9.pdf.

Boyd:2013:NRP

[1213] John P. Boyd, Gregor Gassner, and Burhan A. Sadiq. The nonconvergence of h-refinement in prolate elements. Journal of Scientific Computing, 57(2):372–389, November 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9711-2;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9711-
2.pdf.

Garreau:2013:SEF

[1214] Pierre Garreau and David Kopriva. A spectral element framework for option pricing under general exponential Lévy processes. Journal of Scientific Computing, 57(2):390–413, November 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9713-0;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9713-
0.pdf.

Zhu:2013:ISU

[1215] Wei Zhu, Xue-Cheng Tai, and Tony Chan. Image segmentation using Euler’s elastica as the regularization. Journal of Scientific Computing, 57 (2):414–438, November 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9710-3;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9710-3.pdf.

Kestler:2013:EAR

[1216] Sebastian Kestler and Rob Stevenson. An efficient approximate residual evaluation in the adaptive tensor product wavelet method. Journal of Scientific Computing, 57(3):439–463, December 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9712-1;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9712-
1.pdf.

Hochstenbach:2013:PUB

[1217] Michiel E. Hochstenbach. Probabilistic upper bounds for the matrix two-norm. Journal of Scientific Computing, 57(3):464–476, December 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9716-x;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9716-x.pdf.

Ambikasaran:2013:FDS

[1218] Sivaram Ambikasaran and Eric Darve. An O(N log N) fast direct solver for partial hierarchically semi-separable matrices. Journal of Scientific Computing, 57(3):477–501, December 2013. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9714-z;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9714-
z.pdf.

Chen:2013:IFC

[1219] Zhongying Chen, Guangqing Long, Gnaneshwar Nelakanti, and Yongdong Zhang. Iterated fast collocation methods for integral equations of the second kind. Journal of Scientific Computing, 57(3):502–517, December 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9717-9;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9717-9.pdf.

Higueras:2013:SSR

[1220] Inmaculada Higueras. Strong stability for Runge–Kutta schemes on a class of nonlinear problems. Journal of Scientific Computing, 57(3): 518–535, December 2013. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9715-y;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9715-y.pdf.

Melenk:2013:GDM

[1221] J. M. Melenk, A. Parsania, and S. Sauter. General DG-methods for highly indefinite Helmholtz problems. Journal of Scientific Computing, 57(3):536–581, December 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-013-9726-8;   http://link.springer.com/
content/pdf/10.1007/s10915-013-9726-8.pdf.

Min:2013:EHO

[1222] Misun Min and Paul Fischer. An efficient high-order time integration method for spectral-element discontinuous Galerkin simulations in electromagnetics. Journal of Scientific Computing, 57(3):582–603, December 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9718-8;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9718-8.pdf.

Antil:2013:TSG

[1223] Harbir Antil, Scott E. Field, Frank Herrmann, Ricardo H. Nochetto, and Manuel Tiglio. Two-step greedy algorithm for reduced order quadratures. Journal of Scientific Computing, 57(3):604–637, December 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9722-z;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9722-z.pdf.

Fang:2013:SBD

[1224] Jun Fang, Xingyu Gao, and Aihui Zhou. A symmetry-based decomposition approach to eigenvalue problems. Journal of Scientific Computing, 57(3):638–669, December 2013. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9719-7;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9719-
7.pdf.

An:2013:SEM

[1225] Jing An and Jie Shen. A spectral-element method for transmission eigenvalue problems. Journal of Scientific Computing, 57(3):670–688, December 2013. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9720-1;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9720-1.pdf.

Si:2014:MCG

[1226] Zhiyong Si, Xiaogang Song, and Pengzhan Huang. Modified characteristics gauge–Uzawa finite element method for time dependent conduction– convection problems. Journal of Scientific Computing, 58(1):1–24, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9721-0;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9721-0.pdf.

Fabreges:2014:ASL

[1227] B. Fabrèges and B. Maury. Approximation of single layer distributions by Dirac masses in finite element computations. Journal of Scientific Computing, 58(1):25–40, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9723-y;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9723-
y.pdf.

Liang:2014:PMP

[1228] Chao Liang and Zhengfu Xu. Parametrized maximum principle preserving flux limiters for high order schemes solving multi-dimensional scalar hyperbolic conservation laws. Journal of Scientific Computing, 58 (1):41–60, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9724-x;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9724-x.pdf.

Svard:2014:ESS

[1229] Magnus Svärd and Hatice Özcan. Entropy-stable schemes for the Euler equations with far-field and wall boundary conditions. Journal of Scientific Computing, 58(1):61–89, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9727-7;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9727-
7.pdf.

Hu:2014:AEL

[1230] Xiaozhe Hu, Young-Ju Lee, Jinchao Xu, and Chen-Song Zhang. On adaptive Eulerian–Lagrangian method for linear convection–diffusion problems. Journal of Scientific Computing, 58(1):90–114, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9731-y;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9731-y.pdf.

Cordier:2014:PAD

[1231] Floraine Cordier, Pierre Degond, and Anela Kumbaro. Phase appearance or disappearance in two-phase flows. Journal of Scientific Computing, 58 (1):115–148, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9725-9;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9725-9.pdf.

Ovadia:2014:NMT

[1232] Jeremy Ovadia and Qing Nie. Numerical methods for two-dimensional stem cell tissue growth. Journal of Scientific Computing, 58(1):149– 175, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9728-6;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9728-6.pdf.

Winters:2014:HOL

[1233] Andrew R. Winters and David A. Kopriva. High-order local time stepping on moving DG spectral element meshes. Journal of Scientific Computing, 58(1):176–202, January 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9730-z;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9730-
z.pdf.

Adjerid:2014:SDG

[1234] Slimane Adjerid and Idir Mechai. A superconvergent discontinuous Galerkin method for hyperbolic problems on tetrahedral meshes. Journal of Scientific Computing, 58(1):203–248, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9735-7;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9735-7.pdf.

Guo:2014:SMN

[1235] Ben yu Guo and Yu jian Jiao. Spectral method for Navier–Stokes equations with slip boundary conditions. Journal of Scientific Computing, 58 (1):249–274, January 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9729-5;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9729-5.pdf.

Chen:2014:PDD

[1236] Rongliang Chen, Yuqi Wu, Zhengzheng Yan, Yubo Zhao, and Xiao-Chuan Cai. A parallel domain decomposition method for 3D unsteady incompressible flows at high Reynolds number. Journal of Scientific Computing, 58(2):275–289, February 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9732-x;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9732-
x.pdf.

Barker:2014:MFE

[1237] Andrew T. Barker and Susanne C. Brenner. A mixed finite element method for the Stokes equations based on a weakly over-penalized symmetric interior penalty approach. Journal of Scientific Computing, 58 (2):290–307, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9733-9;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9733-9.pdf.

Zhang:2014:FFV

[1238] Zhimin Zhang and Qingsong Zou. A family of finite volume schemes of arbitrary order on rectangular meshes. Journal of Scientific Computing, 58 (2):308–330, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9737-5;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9737-5.pdf.

Bruno:2014:SDU

[1239] O. P. Bruno and A. Prieto. Spatially dispersionless, unconditionally stable FC–AD solvers for variable-coefficient PDEs. Journal of Scientific Computing, 58(2):331–366, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9734-8;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9734-
8.pdf.

Hu:2014:SCM

[1240] Jun Hu, Hongying Man, and Shangyou Zhang. A simple conforming mixed finite element for linear elasticity on rectangular grids in any space dimension. Journal of Scientific Computing, 58(2):367–379, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9736-6;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9736-6.pdf.

Guo:2014:ESD

[1241] Ruihan Guo and Yan Xu. Efficient solvers of discontinuous Galerkin discretization for the Cahn–Hilliard equations. Journal of Scientific Computing, 58(2):380–408, February 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9738-4;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9738-
4.pdf.

Babbin:2014:CIO

[1242] J. Babbin, P. A. Forsyth, and G. Labahn. A comparison of iterated optimal stopping and local policy iteration for American options under regime switching. Journal of Scientific Computing, 58(2):409–430, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9739-3;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9739-3.pdf.

Lai:2014:SMO

[1243] Rongjie Lai and Stanley Osher. A splitting method for orthogonality constrained problems. Journal of Scientific Computing, 58(2):431– 449, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9740-x;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9740-x.pdf.

Wu:2014:SCS

[1244] Conghai Wu and Ning Zhao. A stability criterion for semidiscrete difference schemes of hyperbolic conservation laws on uniform grids. Journal of Scientific Computing, 58(2):450–471, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9742-8;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9742-8.pdf.

Griesmaier:2014:DWE

[1245] Roland Griesmaier and Peter Monk. Discretization of the wave equation using continuous elements in time and a hybridizable discontinuous Galerkin method in space. Journal of Scientific Computing, 58 (2):472–498, February 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9741-9;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9741-9.pdf.

Huang:2014:SCM

[1246] Can Huang and Zhimin Zhang. Spectral collocation methods for differential-algebraic equations with arbitrary index. Journal of Scientific Computing, 58(3):499–516, March 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9755-3;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9755-
3.pdf.

deDios:2014:SPD

[1247] Blanca Ayuso de Dios, Franco Brezzi, L. Donatella Marini, Jinchao Xu, and Ludmil Zikatanov. A simple preconditioner for a discontinuous Galerkin method for the Stokes problem. Journal of Scientific Computing, 58(3):517–547, March 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9758-0;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9758-
0.pdf.

Jin:2014:NPN

[1248] Qiyu Jin, Ion Grama, and Quansheng Liu. A new Poisson noise filter based on weights optimization. Journal of Scientific Computing, 58 (3):548–573, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9743-7;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9743-7.pdf.

Hu:2014:LUB

[1249] Jun Hu, Yunqing Huang, and Quan Shen. The lower/upper bound property of approximate eigenvalues by nonconforming finite element methods for elliptic operators. Journal of Scientific Computing, 58(3):574–591, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9744-6;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9744-6.pdf.

Alexanderian:2014:PBR

[1250] Alen Alexanderian, Francesco Rizzi, Muruhan Rathinam, Olivier P. Le Maître, and Omar M. Knio. Preconditioned Bayesian regression for stochastic chemical kinetics. Journal of Scientific Computing, 58(3): 592–626, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9745-5;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9745-5.pdf.

Gao:2014:OEA

[1251] Huadong Gao. Optimal error analysis of Galerkin FEMs for nonlinear Joule heating equations. Journal of Scientific Computing, 58(3):627–647, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9746-4;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9746-4.pdf.

Yi:2014:IST

[1252] Dokkyun Yi. An iterative scheme for total variation-based image denoising. Journal of Scientific Computing, 58(3):648–671, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9750-8;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9750-8.pdf.

deLuna:2014:NSC

[1253] Manuel Quezada de Luna and David I. Ketcheson. Numerical simulation of cylindrical solitary waves in periodic media. Journal of Scientific Computing, 58(3):672–689, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9747-3;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9747-
3.pdf.

Mirzaee:2014:SIA

[1254] Hanieh Mirzaee, Jennifer K. Ryan, and Robert M. Kirby. Smoothness-Increasing Accuracy-Conserving (SIAC) filters for discontinuous Galerkin solutions: Application to structured tetrahedral meshes. Journal of Scientific Computing, 58(3):690–704, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9748-2;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9748-
2.pdf.

Lai:2014:FFG

[1255] Rongjie Lai, Zaiwen Wen, Wotao Yin, Xianfeng Gu, and Lok Ming Lui. Folding-free global conformal mapping for genus-0 surfaces by harmonic energy minimization. Journal of Scientific Computing, 58(3):705–725, March 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9752-6;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9752-6.pdf.

Boyd:2014:RBC

[1256] John P. Boyd and Rolfe Petschek. The relationships between Chebyshev, Legendre and Jacobi polynomials: The generic superiority of Chebyshev polynomials and three important exceptions. Journal of Scientific Computing, 59(1):1–27, April 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9751-7;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9751-
7.pdf.

Cyranka:2014:EGA

[1257] Jacek Cyranka. Efficient and generic algorithm for rigorous integration forward in time of dPDEs: Part I. Journal of Scientific Computing, 59(1):28–52, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9749-1;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9749-1.pdf.

Zhao:2014:CFA

[1258] Xuying Zhao, Zhong-Ci Shi, and Qiang Du. Constraint-free adaptive FEMs on quadrilateral nonconforming meshes. Journal of Scientific Computing, 59(1):53–79, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-013-9753-5;   http://link.springer.com/
content/pdf/10.1007/s10915-013-9753-5.pdf.

Zarghami:2014:HPM

[1259] Ahad Zarghami, Chiara Biscarini, Sauro Succi, and Stefano Ubertini. Hydrodynamics in porous media: a finite volume lattice Boltzmann study. Journal of Scientific Computing, 59(1):80–103, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9754-4;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9754-4.pdf.

Zhang:2014:EAC

[1260] Ya nan Zhang and Zhi zhong Sun. Error analysis of a compact ADI scheme for the 2D fractional subdiffusion equation. Journal of Scientific Computing, 59(1):104–128, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9756-2;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9756-
2.pdf.

Feng:2014:LDG

[1261] Xiaobing Feng and Thomas Lewis. Local discontinuous Galerkin methods for one-dimensional second order fully nonlinear elliptic and parabolic equations. Journal of Scientific Computing, 59(1):129–157, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9763-3;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9763-3.pdf.

Wang:2014:OPW

[1262] Tingchun Wang. Optimal point-wise error estimate of a compact difference scheme for the coupled Gross–Pitaevskii equations in one dimension. Journal of Scientific Computing, 59(1):158–186, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9757-1;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9757-1.pdf.

Chen:2014:CBR

[1263] Peng Chen, Alfio Quarteroni, and Gianluigi Rozza. Comparison between reduced basis and stochastic collocation methods for elliptic problems. Journal of Scientific Computing, 59(1):187–216, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9764-2;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9764-2.pdf.

Celiker:2014:HMN

[1264] Fatih Celiker and Li Fan. HDG methods for Naghdi arches. Journal of Scientific Computing, 59(1):217–246, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9759-z;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9759-
z.pdf.

Sjogreen:2014:SEF

[1265] Björn Sjögreen and N. Anders Petersson. Source estimation by full wave form inversion. Journal of Scientific Computing, 59(1):247–276, April 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9760-6;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9760-6.pdf.

Jung:2014:VIS

[1266] Miyoun Jung, Myeongmin Kang, and Myungjoo Kang. Variational image segmentation models involving non-smooth data-fidelity terms. Journal of Scientific Computing, 59(2):277–308, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9766-0;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9766-
0.pdf.

Owens:2014:QOC

[1267] Luke Owens. Quasi-optimal convergence rate of an adaptive weakly overpenalized interior penalty method. Journal of Scientific Computing, 59 (2):309–333, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9765-1;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9765-1.pdf.

Meixner:2014:DGC

[1268] Jessica Meixner, J. Casey Dietrich, Clint Dawson, Marcel Zijlema, and Leo H. Holthuijsen. A discontinuous Galerkin coupled wave propagation/circulation model. Journal of Scientific Computing, 59(2):334–370, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9761-5;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9761-5.pdf.

Hsiao:2014:SNO

[1269] George C. Hsiao, Liwei Xu, and Shangyou Zhang. Solving negative order equations by the multigrid method via variable substitution. Journal of Scientific Computing, 59(2):371–385, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9762-4;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9762-4.pdf.

Schaeffer:2014:VDF

[1270] Hayden Schaeffer and Luminita Vese. Variational dynamics of free triple junctions. Journal of Scientific Computing, 59(2):386–411, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9767-z;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9767-z.pdf.

Zhao:2014:NMS

[1271] Xiaofei Zhao and Ziyi Li. Numerical methods and simulations for the dynamics of one-dimensional Zakharov–Rubenchik equations. Journal of Scientific Computing, 59(2):412–438, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9768-y;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9768-
y.pdf.

Guo:2014:CSM

[1272] Ben-Yu Guo and Xu-Hong Yu. Composite spectral method for exterior problems with polygonal obstacles. Journal of Scientific Computing, 59(2):439–472, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9769-x;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9769-x.pdf.

Mu:2014:WGF

[1273] Lin Mu, Junping Wang, Xiu Ye, and Shangyou Zhang. A C^{0}- weak Galerkin finite element method for the biharmonic equation. Journal of Scientific Computing, 59(2):473–495, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9770-4;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9770-
4.pdf.

Chen:2014:PEE

[1274] Long Chen, Junping Wang, and Xiu Ye. A posteriori error estimates for weak Galerkin finite element methods for second order elliptic problems. Journal of Scientific Computing, 59(2):496–511, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9771-3;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9771-3.pdf.

Cockburn:2014:SCD

[1275] Bernardo Cockburn and Manuel Solano. Solving convection–diffusion problems on curved domains by extensions from subdomains. Journal of Scientific Computing, 59(2):512–543, May 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9776-y;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9776-
y.pdf.

Leclaire:2014:HOS

[1276] Sébastien Leclaire, Maud El-Hachem, Jean-Yves Trépanier, and Marcelo Reggio. High order spatial generalization of 2D and 3D isotropic discrete gradient operators with fast evaluation on GPUs. Journal of Scientific Computing, 59(3):545–573, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9772-2;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9772-
2.pdf.

Chen:2014:LIA

[1277] Wenbin Chen, Cheng Wang, Xiaoming Wang, and Steven M. Wise. A linear iteration algorithm for a second-order energy stable scheme for a thin film model without slope selection. Journal of Scientific Computing, 59(3):574–601, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9774-0;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9774-0.pdf.

Lewis:2014:CAS

[1278] Thomas Lewis and Michael Neilan. Convergence analysis of a symmetric dual-wind discontinuous Galerkin method. Journal of Scientific Computing, 59(3):602–625, June 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9773-1;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9773-
1.pdf.

Chun:2014:MMF

[1279] Sehun Chun. Method of moving frames to solve (an)isotropic diffusion equations on curved surfaces. Journal of Scientific Computing, 59(3):626–666, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9775-z;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9775-z.pdf.

Wang:2014:SJG

[1280] Li-Lian Wang, Xiaodan Zhao, and Zhimin Zhang. Superconvergence of Jacobi–Gauss-type spectral interpolation. Journal of Scientific Computing, 59(3):667–687, June 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9777-x;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9777-
x.pdf.

Tian:2014:LDG

[1281] Lulu Tian, Yan Xu, J. G. M. Kuerten, and J. J. W. Van der Vegt. A local discontinuous Galerkin method for the propagation of phase transition in solids and fluids. Journal of Scientific Computing, 59(3):688–720, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9778-9;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9778-9.pdf.

Williams:2014:ESF

[1282] D. M. Williams and A. Jameson. Energy stable flux reconstruction schemes for advection–diffusion problems on tetrahedra. Journal of Scientific Computing, 59(3):721–759, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9780-2;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9780-

2.pdf. See erratum [1283].

Williams:2014:EES

[1283] D. M. Williams and A. Jameson. Erratum to: Energy stable flux reconstruction schemes for advection–diffusion problems on tetrahedra. Journal of Scientific Computing, 59(3):760, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9789-6;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9789-6.pdf. See [1282].

Hu:2014:SNN

[1284] Zhicheng Hu, Ruo Li, Tiao Lu, Yanli Wang, and Wenqi Yao. Simulation of an n^{+}-n-n^{+} diode by using globally-hyperbolically-closed highorder moment models. Journal of Scientific Computing, 59(3):761–774, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9781-1;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9781-1.pdf.

Ervin:2014:ASD

[1285] V. J. Ervin, E. W. Jenkins, and Hyesuk Lee. Approximation of the Stokes–Darcy system by optimization. Journal of Scientific Computing, 59(3):775–794, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9779-8;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9779-8.pdf.

Baccouch:2014:LDGa

[1286] Mahboub Baccouch. The local discontinuous Galerkin method for the fourth-order Euler–Bernoulli partial differential equation in one space dimension. Part I: Superconvergence error analysis. Journal of Scientific Computing, 59(3):795–840, June 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9782-0;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9782-
0.pdf.

Baccouch:2014:LDGb

[1287] Mahboub Baccouch. The local discontinuous Galerkin method for the fourth-order Euler–Bernoulli partial differential equation in one space dimension. Part II: A posteriori error estimation. Journal of Scientific Computing, 60(1):1–34, July 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9783-z;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9783-
z.pdf.

Gotschel:2014:LCO

[1288] S. Götschel, N. Chamakuri, K. Kunisch, and M. Weiser. Lossy compression in optimal control of cardiac defibrillation. Journal of Scientific Computing, 60(1):35–59, July 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9785-x;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9785-
x.pdf.

Shao:2014:KLD

[1289] Wen-Ze Shao, Hai-Song Deng, and Zhi-Hui Wei. Kullback–Leibler divergence based composite prior modeling for Bayesian superresolution. Journal of Scientific Computing, 60(1):60–78, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9784-y;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9784-y.pdf.

Chang:2014:DDM

[1290] Huibin Chang, Xiaoqun Zhang, Xue-Cheng Tai, and Danping Yang. Domain decomposition methods for nonlocal total variation image restoration. Journal of Scientific Computing, 60(1):79–100, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9786-9;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9786-9.pdf.

Seal:2014:HOM

[1291] David C. Seal, Yaman Güçlü, and Andrew J. Christlieb. Highorder multiderivative time integrators for hyperbolic conservation laws. Journal of Scientific Computing, 60(1):101–140, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9787-8;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9787-
8.pdf.

Sidi:2014:RES

[1292] Avram Sidi. Richardson extrapolation on some recent numerical quadrature formulas for singular and hypersingular integrals and its study of stability. Journal of Scientific Computing, 60(1):141–159, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9788-7;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9788-7.pdf.

Du:2014:AFE

[1293] Ning Du, Liang Ge, and Wenbin Liu. Adaptive finite element approximation for an elliptic optimal control problem with both pointwise and integral control constraints. Journal of Scientific Computing, 60(1):160–183, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9790-0;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9790-0.pdf.

Achchab:2014:SBP

[1294] B. Achchab, A. Agouzal, N. Debit, and K. Bouihat. Star-based a posteriori error estimates for elliptic problems. Journal of Scientific Computing, 60(1):184–202, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9793-x;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9793-x.pdf.

Antonietti:2014:DDP

[1295] Paola F. Antonietti, Stefano Giani, and Paul Houston. Domain decomposition preconditioners for discontinuous Galerkin methods for elliptic problems on complicated domains. Journal of Scientific Computing, 60(1):203–227, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9792-y;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9792-y.pdf.

Davis:2014:PUM

[1296] Christopher B. Davis. A partition of unity method with penalty for fourth order problems. Journal of Scientific Computing, 60(1):228–248, July 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9795-8;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9795-8.pdf.

Ng:2014:CET

[1297] Tsz Ching Ng, Xianfeng Gu, and Lok Ming Lui. Computing extremal Teichmüller map of multiply-connected domains via Beltrami holomorphic flow. Journal of Scientific Computing, 60(2):249–275, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9791-z;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9791-z.pdf.

Ji:2014:MMH

[1298] Xia Ji, Jiguang Sun, and Hehu Xie. A multigrid method for Helmholtz transmission eigenvalue problems. Journal of Scientific Computing, 60 (2):276–294, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9794-9;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9794-9.pdf.

Boulmezaoud:2014:NAS

[1299] Tahar Z. Boulmezaoud, Samy Mziou, and Tahar Boudjedaa. Numerical approximation of second-order elliptic problems in unbounded domains. Journal of Scientific Computing, 60(2):295–312, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9798-5;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9798-5.pdf.

Kubatko:2014:OSS

[1300] Ethan J. Kubatko, Benjamin A. Yeager, and David I. Ketcheson. Optimal strong-stability-preserving Runge–Kutta time discretizations for discontinuous Galerkin methods. Journal of Scientific Computing, 60 (2):313–344, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9796-7;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9796-7.pdf.

Tal-Ezer:2014:NTP

[1301] Hillel Tal-Ezer. Nonperiodic trigonometric polynomial approximation. Journal of Scientific Computing, 60(2):345–362, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9797-6;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9797-6.pdf.

Castro:2014:CIR

[1302] Manuel J. Castro, José M. Gallardo, and Antonio Marquina. A class of incomplete Riemann solvers based on uniform rational approximations to the absolute value function. Journal of Scientific Computing, 60 (2):363–389, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9800-2;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9800-2.pdf.

Wang:2014:NEA

[1303] Jilu Wang. A new error analysis of Crank–Nicolson Galerkin FEMs for a generalized nonlinear Schrödinger equation. Journal of Scientific Computing, 60(2):390–407, August 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9799-4;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9799-
4.pdf.

Fernandez-Nieto:2014:MMH

[1304] E. D. Fernández-Nieto, E. H. Koné, and T. Chacón Rebollo. A multilayer method for the hydrostatic Navier–Stokes equations: a particular weak solution. Journal of Scientific Computing, 60(2):408–437, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9802-0;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9802-0.pdf.

Schutz:2014:APM

[1305] Jochen Schütz. An asymptotic preserving method for linear systems of balance laws based on Galerkin’s method. Journal of Scientific Computing, 60(2):438–456, August 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9801-1;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9801-
1.pdf.

King:2014:EBP

[1306] James King, Sergey Yakovlev, Zhisong Fu, Robert M. Kirby, and Spencer J. Sherwin. Exploiting batch processing on streaming architectures to solve 2D elliptic finite element problems: A hybridized discontinuous Galerkin (HDG) case study. Journal of Scientific Computing, 60(2):457–482, August 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9805-x;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9805-x.pdf.

Chen:2014:MDB

[1307] Dai-Qiang Chen and Yan Zhou. Multiplicative denoising based on linearized alternating direction method using discrepancy function constraint. Journal of Scientific Computing, 60(3):483–504, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9803-z;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9803-z.pdf.

Perotto:2014:CMG

[1308] Simona Perotto and Alessandro Veneziani. Coupled model and grid adaptivity in hierarchical reduction of elliptic problems. Journal of Scientific Computing, 60(3):505–536, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9804-y;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9804-
y.pdf.

Ballarin:2014:SOF

[1309] Francesco Ballarin, Andrea Manzoni, Gianluigi Rozza, and Sandro Salsa. Shape optimization by free-form deformation: Existence results and numerical solution for Stokes flows. Journal of Scientific Computing, 60(3): 537–563, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9807-8;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9807-8.pdf.

Shang:2014:PSS

[1310] Yueqiang Shang and Shumei Huang. A parallel subgrid stabilized finite element method based on two-grid discretization for simulation of 2D/ 3D steady incompressible flows. Journal of Scientific Computing, 60(3): 564–583, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9806-9;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9806-9.pdf.

Abgrall:2014:RFV

[1311] Remi Abgrall and Harish Kumar. Robust finite volume schemes for two-fluid plasma equations. Journal of Scientific Computing, 60(3): 584–611, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9809-6;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9809-6.pdf.

Johnston:2014:LPB

[1312] Hans Johnston, Cheng Wang, and Jian-Guo Liu. A local pressure boundary condition spectral collocation scheme for the three-dimensional Navier–Stokes equations. Journal of Scientific Computing, 60(3):612– 626, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9808-7;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9808-7.pdf.

DAmbrosio:2014:LTS

[1313] R. D’Ambrosio and E. Hairer. Long-term stability of multivalue methods for ordinary differential equations. Journal of Scientific Computing, 60(3):627–640, September 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9812-y;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9812-
y.pdf.

Arandiga:2014:WDM

[1314] F. Aràndiga, M. C. Martí, and P. Mulet. Weights design for maximal order WENO schemes. Journal of Scientific Computing, 60(3): 641–659, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9810-0;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9810-0.pdf.

Zhang:2014:TDL

[1315] Jianying Zhang and Guangwu Yan. Three-dimensional lattice Boltzmann model for the complex Ginzburg–Landau equation. Journal of Scientific Computing, 60(3):660–683, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9811-z;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9811-
z.pdf.

Suhov:2014:APA

[1316] A. Y. Suhov. An accurate polynomial approximation of exponential integrators. Journal of Scientific Computing, 60(3):684–698, September 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9813-x;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9813-x.pdf.

Liu:2014:RAG

[1317] Li-Bin Liu and Yanping Chen. A robust adaptive grid method for a system of two singularly perturbed convection–diffusion equations with weak coupling. Journal of Scientific Computing, 61(1):1–16, October 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
013-9814-9;   http://link.springer.com/content/pdf/10.1007/
s10915-013-9814-9.pdf.

Xiao:2014:NBB

[1318] Yunhai Xiao, Soon-Yi Wu, and Liqun Qi. Nonmonotone Barzilai– Borwein gradient algorithm for ℓ_{1}-regularized nonsmooth minimization in compressive sensing. Journal of Scientific Computing, 61(1): 17–41, October 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-013-9815-8;   http://link.springer.com/content/
pdf/10.1007/s10915-013-9815-8.pdf.

Zhang:2014:MSD

[1319] Yifan Zhang, Wei Wang, Johnny Guzmán, and Chi-Wang Shu. Multi-scale discontinuous Galerkin method for solving elliptic problems with curvilinear unidirectional rough coefficients. Journal of Scientific Computing, 61(1):42–60, October 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-013-9816-7;
 http://link.springer.com/content/pdf/10.1007/s10915-013-9816-
7.pdf.

Ghosh:2014:WNL

[1320] Debojyoti Ghosh and James D. Baeder. Weighted non-linear compact schemes for the direct numerical simulation of compressible, turbulent flows. Journal of Scientific Computing, 61(1):61–89, October 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9818-0;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9818-0.pdf.

Virta:2014:AWP

[1321] Kristoffer Virta and Ken Mattsson. Acoustic wave propagation in complicated geometries and heterogeneous media. Journal of Scientific Computing, 61(1):90–118, October 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9817-1;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9817-
1.pdf.

Zhang:2014:PIE

[1322] Hong Zhang, Adrian Sandu, and Sebastien Blaise. Partitioned and implicit–explicit general linear methods for ordinary differential equations. Journal of Scientific Computing, 61(1):119–144, October 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9819-z;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9819-z.pdf.

Bebendorf:2014:SVF

[1323] M. Bebendorf and C. Kuske. Separation of variables for function generated high-order tensors. Journal of Scientific Computing, 61(1): 145–165, October 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9822-4;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9822-4.pdf.

Mitsotakis:2014:GFE

[1324] Dimitrios Mitsotakis, Boaz Ilan, and Denys Dutykh. On the Galerkin/finite-element method for the Serre equations. Journal of Scientific Computing, 61(1):166–195, October 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9823-3;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9823-
3.pdf.

Hu:2014:LBE

[1325] Jun Hu, Yunqing Huang, and Qun Lin. Lower bounds for eigenvalues of elliptic operators: By nonconforming finite element methods. Journal of Scientific Computing, 61(1):196–221, October 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9821-5;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9821-5.pdf.

Yang:2014:CLB

[1326] Xuguang Yang, Baochang Shi, Zhenhua Chai, and Zhaoli Guo. A coupled lattice Boltzmann method to solve Nernst–Planck model for simulating electro-osmotic flows. Journal of Scientific Computing, 61(1): 222–238, October 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9820-6;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9820-6.pdf.

DAmore:2014:SAV

[1327] Luisa D’Amore, Rossella Arcucci, Luisa Carracciuolo, and Almerico Murli. A scalable approach for variational data assimilation. Journal of Scientific Computing, 61(2):239–257, November 2014. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9824-2;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9824-
2.pdf.

Yang:2014:PDD

[1328] Haijian Yang, Chao Yang, and Xiao-Chuan Cai. Parallel domain decomposition methods with mixed order discretization for fully implicit solution of tracer transport problems on the cubedsphere. Journal of Scientific Computing, 61(2):258–280, November 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9828-y;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9828-y.pdf.

Shi:2014:FHC

[1329] Lei Shi, Z. J. Wang, L. P. Zhang, Wei Liu, and Song Fu. A P_{N}P_{M}-CPR framework for hyperbolic conservation laws. Journal of Scientific Computing, 61(2):281–307, November 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9829-x;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9829-
x.pdf.

Li:2014:ELB

[1330] Qianhuan Li, Zhenhua Chai, and Baochang Shi. An efficient lattice Boltzmann model for steady convection–diffusion equation. Journal of Scientific Computing, 61(2):308–326, November 2014. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9827-z;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9827-
z.pdf.

Sescu:2014:NAS

[1331] Adrian Sescu and Ray Hixon. Numerical anisotropy study of a class of compact schemes. Journal of Scientific Computing, 61(2):327– 342, November 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9826-0;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9826-0.pdf.

Buchmuller:2014:IAH

[1332] Pawel Buchmüller and Christiane Helzel. Improved accuracy of highorder WENO finite volume methods on Cartesian grids. Journal of Scientific Computing, 61(2):343–368, November 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9825-1;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9825-
1.pdf.

Almquist:2014:ASP

[1333] Martin Almquist, Ilkka Karasalo, and Ken Mattsson. Atmospheric sound propagation over large-scale irregular terrain. Journal of Scientific Computing, 61(2):369–397, November 2014. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9830-4;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9830-
4.pdf.

Witherden:2014:ASP

[1334] F. D. Witherden and P. E. Vincent. An analysis of solution point coordinates for flux reconstruction schemes on triangular elements. Journal of Scientific Computing, 61(2):398–423, November 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9832-2;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9832-2.pdf.

Kocher:2014:VST

[1335] Uwe Köcher and Markus Bause. Variational space–time methods for the wave equation. Journal of Scientific Computing, 61(2):424–453, November 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9831-3;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9831-3.pdf.

Pincock:2014:HOF

[1336] B. Pincock and A. Katz. High-order flux correction for viscous flows on arbitrary unstructured grids. Journal of Scientific Computing, 61(2): 454–476, November 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9833-1;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9833-1.pdf.

DeRosis:2014:CBI

[1337] Alessandro De Rosis, Stefano Ubertini, and Francesco Ubertini. A comparison between the interpolated bounce-back scheme and the immersed boundary method to treat solid boundary conditions for laminar flows in the lattice Boltzmann framework. Journal of Scientific Computing, 61 (3):477–489, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9834-0;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9834-0.pdf.

Martinez:2014:EDN

[1338] Adam Martinez, Anne Gelb, and Alexander Gutierrez. Edge detection from non-uniform Fourier data using the convolutional gridding algorithm. Journal of Scientific Computing, 61(3):490–512, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9836-y;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9836-y.pdf.

Roth:2014:SLM

[1339] A. Roth, A. Klar, B. Simeon, and E. Zharovsky. A semi-Lagrangian method for 3-D Fokker Planck equations for stochastic dynamical systems on the sphere. Journal of Scientific Computing, 61(3):513– 532, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9835-z;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9835-z.pdf.

Ji:2014:SCI

[1340] Haifeng Ji, Jinru Chen, and Zhilin Li. A symmetric and consistent immersed finite element method for interface problems. Journal of Scientific Computing, 61(3):533–557, December 2014. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9837-x;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9837-
x.pdf.

Huang:2014:NSS

[1341] Wei-Qiang Huang, Xianfeng David Gu, Wen-Wei Lin, and Shing-Tung Yau. A novel symmetric skew-Hamiltonian isotropic Lanczos algorithm for spectral conformal parameterizations. Journal of Scientific Computing, 61(3):558–583, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9840-2;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9840-
2.pdf.

Xia:2014:FSM

[1342] Yinhua Xia. Fourier spectral methods for Degasperis–Procesi equation with discontinuous solutions. Journal of Scientific Computing, 61(3): 584–603, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9839-8;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9839-8.pdf.

Massing:2014:SNF

[1343] André Massing, Mats G. Larson, Anders Logg, and Marie E. Rognes. A stabilized Nitsche fictitious domain method for the Stokes problem. Journal of Scientific Computing, 61(3):604–628, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9838-9;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9838-9.pdf.

Liao:2014:SCM

[1344] Hong lin Liao, Ya nan Zhang, Ying Zhao, and Han sheng Shi. Stability and convergence of modified Du Fort–Frankel schemes for solving timefractional subdiffusion equations. Journal of Scientific Computing, 61(3): 629–648, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9841-1;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9841-1.pdf.

Starinshak:2014:NLS

[1345] David P. Starinshak, Smadar Karni, and Philip L. Roe. A new level-set model for the representation of non-smooth geometries. Journal of Scientific Computing, 61(3):649–672, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9842-0;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9842-0.pdf.

Yu:2014:SEM

[1346] Xu-Hong Yu and Ben-Yu Guo. Spectral element method for mixed inhomogeneous boundary value problems of fourth order. Journal of Scientific Computing, 61(3):673–701, December 2014. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9844-y;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9844-
y.pdf.

Hong:2015:NAS

[1347] Youngjoon Hong. Numerical approximation of the singularly perturbed heat equation in a circle. Journal of Scientific Computing, 62 (1):1–24, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9845-x;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9845-x.pdf.

Gerhard:2015:HOD

[1348] Nils Gerhard, Francesca Iacono, Georg May, Siegfried Müller, and Roland Schäfer. A high-order discontinuous Galerkin discretization with multiwavelet-based grid adaptation for compressible flows. Journal of Scientific Computing, 62(1):25–52, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9846-9;

 http://link.springer.com/content/pdf/10.1007/s10915-014-9846-
9.pdf.

Shirokoff:2015:SIA

[1349] D. Shirokoff and J.-C. Nave. A sharp-interface active penalty method for the incompressible Navier–Stokes equations. Journal of Scientific Computing, 62(1):53–77, January 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9849-6;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9849-
6.pdf.

Bialecki:2015:FDS

[1350] Bernard Bialecki and Andreas Karageorghis. Finite difference schemes for the Cauchy–Navier equations of elasticity with variable coefficients. Journal of Scientific Computing, 62(1):78–121, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9847-8;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9847-8.pdf.

Huang:2015:SAL

[1351] Yunqing Huang, Jichun Li, Chao Wu, and Wei Yang. Superconvergence analysis for linear tetrahedral edge elements. Journal of Scientific Computing, 62(1):122–145, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9848-7;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9848-
7.pdf.

Xu:2015:DPN

[1352] Da Xu. Decay properties for the numerical solutions of a partial differential equation with memory. Journal of Scientific Computing, 62 (1):146–178, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9850-0;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9850-0.pdf.

Wang:2015:UMC

[1353] Chunmei Wang, Shangyou Zhang, and Jinru Chen. A unified mortar condition for nonconforming finite elements. Journal of Scientific Computing, 62(1):179–197, January 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9852-y;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9852-
y.pdf.

Wu:2015:TOF

[1354] Liang Wu and Yong-Tao Zhang. A third order fast sweeping method with linear computational complexity for eikonal equations. Journal of Scientific Computing, 62(1):198–229, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9856-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9856-7.pdf.

He:2015:SFO

[1355] Yinnian He, Pengzhan Huang, and Xinlong Feng. H^{2}-stability of the first order fully discrete schemes for the time-dependent Navier–Stokes equations. Journal of Scientific Computing, 62(1):230–264, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9854-9;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9854-9.pdf.

Bialecki:2015:ECN

[1356] B. Bialecki, G. Fairweather, and J. C. López-Marcos. The extrapolated Crank–Nicolson orthogonal spline collocation method for a quasilinear parabolic problem with nonlocal boundary conditions. Journal of Scientific Computing, 62(1):265–283, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9853-x;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9853-x.pdf.

Yang:2015:LUB

[1357] Yidu Yang, Jiayu Han, Hai Bi, and Yuanyuan Yu. The lower/upper bound property of the Crouzeix–Raviart element eigenvalues on adaptive meshes. Journal of Scientific Computing, 62(1):284–299, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9855-8;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9855-8.pdf.

Li:2015:NCD

[1358] Ming Li, Yujiao Wang, and Leevan Ling. Numerical Caputo differentiation by radial basis functions. Journal of Scientific Computing, 62 (1):300–315, January 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9857-6;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9857-6.pdf.

Li:2015:DKP

[1359] Ruo Li, Tiao Lu, and Wenqi Yao. Discrete kernel preserving model for 1D electron–optical phonon scattering. Journal of Scientific Computing, 62 (2):317–335, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9858-5;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9858-5.pdf.

Jung:2015:ENN

[1360] Miyoun Jung and Myungjoo Kang. Efficient nonsmooth nonconvex optimization for image restoration and segmentation. Journal of Scientific Computing, 62(2):336–370, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9860-y;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9860-
y.pdf.

Buet:2015:APS

[1361] Christophe Buet, Bruno Després, and Emmanuel Franck. Asymptotic preserving schemes on distorted meshes for Friedrichs systems with stiff relaxation: Application to angular models in linear transport. Journal of Scientific Computing, 62(2):371–398, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9859-4;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9859-4.pdf.

Baccouch:2015:PLD

[1362] Mahboub Baccouch and Slimane Adjerid. A posteriori local discontinuous Galerkin error estimation for two-dimensional convection–diffusion problems. Journal of Scientific Computing, 62(2):399–430, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9861-x;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9861-x.pdf.

Ju:2015:FEI

[1363] Lili Ju, Jian Zhang, Liyong Zhu, and Qiang Du. Fast explicit integration factor methods for semilinear parabolic equations. Journal of Scientific Computing, 62(2):431–455, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9862-9;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9862-9.pdf.

Hintermuller:2015:NOD

[1364] Michael Hintermüller and Andreas Langer. Non-overlapping domain decomposition methods for dual total variation based image denoising. Journal of Scientific Computing, 62(2):456–481, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9863-8;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9863-8.pdf.

Yu:2015:DRP

[1365] C. H. Yu, Yogesh G. Bhumkar, and Tony W. H. Sheu. Dispersion relation preserving combined compact difference schemes for flow problems. Journal of Scientific Computing, 62(2):482–516, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9864-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9864-7.pdf.

Skelton:2015:PRR

[1366] Andrew Skelton and Allan R. Willms. Parameter range reduction in ordinary differential equation models. Journal of Scientific Computing, 62 (2):517–531, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9865-6;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9865-6.pdf.

Reyna:2015:OBT

[1367] Matthew A. Reyna and Fengyan Li. Operator bounds and time step conditions for the DG and central DG methods. Journal of Scientific Computing, 62(2):532–554, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9866-5;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9866-
5.pdf.

Hu:2015:APE

[1368] Jingwei Hu, Qin Li, and Lorenzo Pareschi. Asymptotic-preserving exponential methods for the quantum Boltzmann equation with high-order accuracy. Journal of Scientific Computing, 62(2):555–574, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9869-2;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9869-2.pdf.

Chen:2015:NFG

[1369] Feng Chen. A new framework of GPU-accelerated spectral solvers: Collocation and glerkin methods for systems of coupled elliptic equations. Journal of Scientific Computing, 62(2):575–600, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9868-3;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9868-3.pdf.

Liu:2015:DES

[1370] Chun Liu, Jie Shen, and Xiaofeng Yang. Decoupled energy stable schemes for a phase-field model of two-phase incompressible flows with variable density. Journal of Scientific Computing, 62(2):601–622, February 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9867-4;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9867-4.pdf.

Addam:2015:FDA

[1371] Mohamed Addam, Abderrahman Bouhamidi, and Mohammed Seaid. A frequency-domain approach for the P_{1} approximation of time-dependent radiative transfer. Journal of Scientific Computing, 62(3):623–651, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9870-9;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9870-9.pdf.

Yao:2015:CAM

[1372] Xudong Yao. Convergence analysis of a minimax method for finding multiple solutions of semilinear elliptic equation: Part I — on polyhedral domain. Journal of Scientific Computing, 62(3):652–673, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

014-9871-8;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9871-8.pdf.

Xu:2015:DLS

[1373] Guoliang Xu and Qing Pan. Design of loop’s subdivision surfaces by fourth-order geometric PDEs with G^{1} boundary conditions. Journal of Scientific Computing, 62(3):674–692, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9872-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9872-7.pdf.

Ashwin:2015:KEC

[1374] V. M. Ashwin, K. Saurabh, M. Sriramkrishnan, P. M. Bagade, M. K. Parvathi, and Tapan K. Sengupta. KdV equation and computations of solitons: Nonlinear error dynamics. Journal of Scientific Computing, 62 (3):693–717, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9875-4;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9875-4.pdf.

Deng:2015:NAF

[1375] Weihua Deng, Minghua Chen, and Eli Barkai. Numerical algorithms for the forward and backward fractional Feynman–Kac equations. Journal of Scientific Computing, 62(3):718–746, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9873-6;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9873-6.pdf.

Zhao:2015:CCN

[1376] Xuan Zhao and Zhi-Zhong Sun. Compact Crank–Nicolson schemes for a class of fractional Cattaneo equation in inhomogeneous medium. Journal of Scientific Computing, 62(3):747–771, March 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9874-5;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9874-
5.pdf.

Warzynski:2015:RKR

[1377] Andrzej Warzyński, Matthew E. Hubbard, and Mario Ricchiuto. Runge– Kutta residual distribution schemes. Journal of Scientific Computing, 62(3):772–802, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9879-0;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9879-0.pdf.

Liu:2015:ESD

[1378] Hailiang Liu and Hui Yu. The entropy satisfying discontinuous Galerkin method for Fokker–Planck equations. Journal of Scientific Computing, 62(3):803–830, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9878-1;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9878-1.pdf.

Ohtsuka:2015:LSA

[1379] T. Ohtsuka, Y.-H. R. Tsai, and Y. Giga. A level set approach reflecting sheet structure with single auxiliary function for evolving spirals on crystal surfaces. Journal of Scientific Computing, 62(3):831–874, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9877-2;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9877-2.pdf.

Parisi:2015:NTI

[1380] Florencia Parisi, Mariana Cécere, Mirta Iriondo, and Oscar Reula. Numerical treatment of interfaces for second-order wave equations. Journal of Scientific Computing, 62(3):875–897, March 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9880-7;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9880-
7.pdf.

Ala:2015:NII

[1381] Guido Ala and Elisa Francomano. Numerical investigations of an implicit leapfrog time-domain meshless method. Journal of Scientific Computing, 62(3):898–912, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9881-6;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9881-6.pdf.

Asthana:2015:HOF

[1382] Kartikey Asthana and Antony Jameson. High-order flux reconstruction schemes with minimal dispersion and dissipation. Journal of Scientific Computing, 62(3):913–944, March 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9882-5;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9882-
5.pdf.

Huang:2015:SCN

[1383] Can Huang and Zhimin Zhang. On the spectrum computation of non-oscillatory and highly oscillatory kernel with weak singularity. Journal of Scientific Computing, 63(1):1–22, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9884-3;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9884-
3.pdf.

Saetra:2015:EGI

[1384] Martin L. Sætra, André R. Brodtkorb, and Knut-Andreas Lie. Efficient GPU-implementation of adaptive mesh refinement for the shallowwater equations. Journal of Scientific Computing, 63(1):23–48, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9883-4;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9883-4.pdf.

Jia:2015:SSD

[1385] Feilin Jia, Zhen Gao, and Wai Sun Don. A spectral study on the dissipation and dispersion of the WENO schemes. Journal of Scientific Computing, 63(1):49–77, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9886-1;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9886-1.pdf.

Barrett:2015:SPF

[1386] John W. Barrett, Harald Garcke, and Robert Nürnberg. A stable parametric finite element discretization of two-phase Navier– Stokes flow. Journal of Scientific Computing, 63(1):78–117, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9885-2;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9885-2.pdf.

Zhao:2015:MAD

[1387] Shan Zhao. A matched alternating direction implicit (ADI) method for solving the heat equation with interfaces. Journal of Scientific Computing, 63(1):118–137, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-014-9887-0;   http://link.springer.com/
content/pdf/10.1007/s10915-014-9887-0.pdf.

Wilcox:2015:DED

[1388] Lucas C. Wilcox, Georg Stadler, Tan Bui-Thanh, and Omar Ghattas. Discretely exact derivatives for hyperbolic PDE-constrained optimization problems discretized by the discontinuous Galerkin method. Journal of Scientific Computing, 63(1):138–162, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9890-5;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9890-5.pdf.

Bazan:2015:SED

[1389] Fermín S. Viloche Bazán. Simple and efficient determination of the Tikhonov regularization parameter chosen by the generalized discrepancy principle for discrete ill–posed problems. Journal of Scientific Computing, 63(1):163–184, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9888-z;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9888-
z.pdf.

Hateley:2015:FMC

[1390] James C. Hateley, Huayi Wei, and Long Chen. Fast methods for computing centroidal Voronoi tessellations. Journal of Scientific Computing, 63(1):185–212, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9894-1;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9894-1.pdf.

Oh:2015:IDU

[1391] Seungmi Oh and Myungjoo Kang. Image decomposition using a local gradient constraint. Journal of Scientific Computing, 63(1):213–232, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9889-y;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9889-y.pdf.

Canic:2015:RKD

[1392] Suncica Canic, Benedetto Piccoli, Jing-Mei Qiu, and Tan Ren. Runge–Kutta discontinuous Galerkin method for traffic flow model on networks. Journal of Scientific Computing, 63(1):233–255, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9896-z;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9896-z.pdf.

deFaria:2015:TDI

[1393] J. Rocha de Faria and D. Lesnic. Topological derivative for the inverse conductivity problem: a Bayesian approach. Journal of Scientific Computing, 63(1):256–278, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9891-4;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9891-
4.pdf.

Elsey:2015:RDV

[1394] Matt Elsey and Benedikt Wirth. Redistancing dynamics for vector– valued multilabel segmentation with costly fidelity: Grain identification in polycrystal images. Journal of Scientific Computing, 63(1):279–306, April 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9892-3;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9892-3.pdf.

Quan:2015:DDM

[1395] Yuhui Quan, Hui Ji, and Zuowei Shen. Data-driven multi-scale non-local wavelet frame construction and image recovery. Journal of Scientific Computing, 63(2):307–329, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9893-2;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9893-
2.pdf.

Theillard:2015:SCM

[1396] Maxime Theillard, Frédéric Gibou, and Tresa Pollock. A sharp computational method for the simulation of the solidification of binary alloys. Journal of Scientific Computing, 63(2):330–354, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9895-0;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9895-0.pdf.

Jung:2015:SAT

[1397] Chang-Yeol Jung and Thien Binh Nguyen. Semi-analytical time differencing methods for stiff problems. Journal of Scientific Computing, 63(2):355–373, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9897-y;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9897-y.pdf.

Stals:2015:EST

[1398] Linda Stals. Efficient solution techniques for a finite element thin plate spline formulation. Journal of Scientific Computing, 63(2):374–409, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9898-x;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9898-x.pdf.

Hu:2015:ECR

[1399] Jun Hu and Rui Ma. The enriched Crouzeix–Raviart elements are equivalent to the Raviart–Thomas elements. Journal of Scientific Computing, 63(2):410–425, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9899-9;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9899-9.pdf.

Dong:2015:TLN

[1400] Xiaojing Dong and Yinnian He. Two-level Newton iterative method for the 2D /3D stationary incompressible magnetohydrodynamics. Journal of Scientific Computing, 63(2):426–451, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9900-7;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9900-
7.pdf.

Cheng:2015:LAL

[1401] Yao Cheng, Feng Zhang, and Qiang Zhang. Local analysis of local discontinuous Galerkin method for the time-dependent singularly perturbed problem. Journal of Scientific Computing, 63(2):452–477, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9901-6;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9901-6.pdf.

Belgacem:2015:FEM

[1402] Faker Ben Belgacem, Christine Bernardi, Faten Jelassi, and Maimouna Mint Brahim. Finite element methods for the temperature in composite media with contact resistance. Journal of Scientific Computing, 63(2):478–501, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9907-0;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9907-0.pdf.

Arpaia:2015:AFE

[1403] L. Arpaia, M. Ricchiuto, and R. Abgrall. An ALE formulation for explicit Runge–Kutta residual distribution. Journal of Scientific Computing, 63 (2):502–547, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9910-5;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9910-5.pdf.

Liu:2015:FDH

[1404] Hongxia Liu and Jianxian Qiu. Finite difference Hermite WENO schemes for hyperbolic conservation laws. Journal of Scientific Computing, 63 (2):548–572, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9905-2;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9905-2.pdf.

Lui:2015:SMD

[1405] Lok Ming Lui and Tsz Ching Ng. A splitting method for diffeomorphism optimization problem using Beltrami coefficients. Journal of Scientific Computing, 63(2):573–611, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9903-4;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9903-
4.pdf.

Mohammed:2015:ECM

[1406] Akmal Nizam Mohammed and Farzad Ismail. Entropy consistent methods for the Navier–Stokes equations. Journal of Scientific Computing, 63(2):612–631, May 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9904-3;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9904-3.pdf.

Hundsdorfer:2015:EAE

[1407] Willem Hundsdorfer, David I. Ketcheson, and Igor Savostianov. Error analysis of explicit partitioned Runge–Kutta schemes for conservation laws. Journal of Scientific Computing, 63(3):633–653, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9906-1;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9906-1.pdf.

Natalini:2015:NCB

[1408] Roberto Natalini, Magali Ribot, and Monika Twarogowska. A numerical comparison between degenerate parabolic and quasilinear hyperbolic models of cell movements under chemotaxis. Journal of Scientific Computing, 63(3):654–677, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9909-y;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9909-
y.pdf.

Yang:2015:WBC

[1409] Suo Yang, Alexander Kurganov, and Yingjie Liu. Well-balanced central schemes on overlapping cells with constant subtraction techniques for the Saint-Venant shallow water system. Journal of Scientific Computing, 63(3):678–698, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9908-z;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9908-z.pdf.

Yang:2015:PRW

[1410] Min Yang, Jiangguo Liu, and Yanping Lin. Pressure recovery for weakly over–penalized discontinuous Galerkin methods for the Stokes problem. Journal of Scientific Computing, 63(3):699–715, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9911-4;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9911-4.pdf.

Chen:2015:CAT

[1411] Long Chen, Ming Wang, and Lin Zhong. Convergence analysis of triangular MAC schemes for two dimensional Stokes equations. Journal of Scientific Computing, 63(3):716–744, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9916-z;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9916-
z.pdf.

Shankar:2015:RBF

[1412] Varun Shankar, Grady B. Wright, Robert M. Kirby, and Aaron L. Fogelson. A radial basis function (RBF)-finite difference (FD) method for diffusion and reaction–diffusion equations on surfaces. Journal of Scientific Computing, 63(3):745–768, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9914-1;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9914-
1.pdf.

Canuto:2015:COP

[1413] Claudio Canuto, Valeria Simoncini, and Marco Verani. Contraction and optimality properties of an adaptive Legendre–Galerkin method: The multi-dimensional case. Journal of Scientific Computing, 63(3):769–798, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9912-3;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9912-3.pdf.

Rossides:2015:CIM

[1414] Tasos Rossides, David J. B. Lloyd, and Sergey Zelik. Computing interacting multi-fronts in one dimensional real Ginzburg Landau equations. Journal of Scientific Computing, 63(3):799–819, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9917-y;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9917-y.pdf.

Mathis:2015:DMA

[1415] Hélène Mathis, Clément Cancès, Edwige Godlewski, and Nicolas Seguin. Dynamic model adaptation for multiscale simulation of hyperbolic systems with relaxation. Journal of Scientific Computing, 63(3):820–861, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9915-0;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9915-0.pdf.

Zhu:2015:CSC

[1416] Shengxin Zhu and Andrew J. Wathen. Convexity and solvability for compactly supported radial basis functions with different shapes. Journal of Scientific Computing, 63(3):862–884, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9919-9;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9919-9.pdf.

Rebollo:2015:NAP

[1417] T. Chacón Rebollo, M. Gómez Mármol, and M. Restelli. Numerical analysis of penalty stabilized finite element discretizations of evolution Navier–Stokes equations. Journal of Scientific Computing, 63 (3):885–912, June 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9918-x;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9918-x.pdf.

Guo:2015:LDG

[1418] Ruihan Guo, Yan Xu, and Zhengfu Xu. Local discontinuous Galerkin methods for the functionalized Cahn–Hilliard equation. Journal of Scientific Computing, 63(3):913–937, June 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9920-3;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9920-
3.pdf.

Dolejsi:2015:ADE

[1419] Vít Dolejsí, Ivana Sebestová, and Martin Vohralík. Algebraic and discretization error estimation by equilibrated fluxes for discontinuous Galerkin methods on nonmatching grids. Journal of Scientific Computing, 64(1):1–34, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9921-2;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9921-2.pdf.

Chen:2015:CSG

[1420] Jen-Hao Chen, I-Liang Chern, and Weichung Wang. A complete study of the ground state phase diagrams of spin-1 Bose–Einstein condensates in a magnetic field via continuation methods. Journal of Scientific Computing, 64(1):35–54, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9924-z;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9924-z.pdf.

Bose:2015:ADN

[1421] Rikhi Bose and Tapan K. Sengupta. Analysis and design of a new dispersion relation preserving alternate direction bidiagonal compact scheme. Journal of Scientific Computing, 64(1):55–82, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9922-1;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9922-1.pdf.

Li:2015:SAN

[1422] Tiexiang Li, Wei-Qiang Huang, Wen-Wei Lin, and Jijun Liu. On spectral analysis and a novel algorithm for transmission eigenvalue problems. Journal of Scientific Computing, 64(1):83–108, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9923-0;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9923-0.pdf.

Wang:2015:NAN

[1423] Changchun Wang and Jianxin Zhou. A new approach for numerically solving nonlinear eigensolution problems. Journal of Scientific Computing, 64(1):109–129, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-014-9925-y;   http://link.springer.com/
content/pdf/10.1007/s10915-014-9925-y.pdf.

Chen:2015:ABE

[1424] Yanli Chen, Yonghai Li, Zhiqiang Sheng, and Guangwei Yuan. Adaptive bilinear element finite volume methods for second-order elliptic problems on nonmatching grids. Journal of Scientific Computing, 64(1):130–150, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9927-9;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9927-9.pdf.

Bowman:2015:FLA

[1425] John C. Bowman, Mohammad Ali Yassaei, and Anup Basu. A fully Lagrangian advection scheme. Journal of Scientific Computing, 64 (1):151–177, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9928-8;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9928-8.pdf.

Lou:2015:CSR

[1426] Yifei Lou, Penghang Yin, Qi He, and Jack Xin. Computing sparse representation in a highly coherent dictionary based on difference of L_{1} and L_{2}. Journal of Scientific Computing, 64(1):178–196, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9930-1;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9930-1.pdf.

Sun:2015:IMF

[1427] Yao Sun, Fuming Ma, and Xu Zhou. An invariant method of fundamental solutions for the Cauchy problem in two-dimensional isotropic linear elasticity. Journal of Scientific Computing, 64(1):197–215, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9929-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9929-7.pdf.

Salgado:2015:CAF

[1428] Abner J. Salgado. Convergence analysis of fractional time-stepping techniques for incompressible fluids with microstructure. Journal of Scientific Computing, 64(1):216–233, July 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9926-x;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9926-
x.pdf.

Wang:2015:EIS

[1429] Xiaojie Wang. An exponential integrator scheme for time discretization of nonlinear stochastic wave equation. Journal of Scientific Computing, 64(1):234–263, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9931-0;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9931-0.pdf.

Jiang:2015:HOE

[1430] Nan Jiang. A higher order ensemble simulation algorithm for fluid flows. Journal of Scientific Computing, 64(1):264–288, July 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9932-z;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9932-z.pdf.

Hsieh:2015:UES

[1431] Po-Wen Hsieh, Ming-Chih Lai, Suh-Yuh Yang, and Cheng-Shu You. An unconditionally energy stable penalty immersed boundary method for simulating the dynamics of an inextensible interface interacting with a solid particle. Journal of Scientific Computing, 64(2):289–316, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9933-y;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9933-y.pdf.

Zhong:2015:MSV

[1432] Min Zhong, Yiu Chung Hon, and Shuai Lu. Multiscale support vector approach for solving ill-posed problems. Journal of Scientific Computing, 64(2):317–340, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9934-x;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9934-x.pdf.

Safdari-Vaighani:2015:RBF

[1433] Ali Safdari-Vaighani, Alfa Heryudono, and Elisabeth Larsson. A radial basis function partition of unity collocation method for convection– diffusion equations arising in financial applications. Journal of Scientific Computing, 64(2):341–367, August 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9935-9;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9935-
9.pdf.

Zhao:2015:RPE

[1434] Jikun Zhao, Shaochun Chen, Bei Zhang, and Shipeng Mao. Robust a posteriori error estimates for conforming discretizations of diffusion problems with discontinuous coefficients on anisotropic meshes. Journal of Scientific Computing, 64(2):368–400, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9937-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9937-7.pdf.

Bosing:2015:DGW

[1435] Paulo Rafael Bösing and Carsten Carstensen. Discontinuous Galerkin with weakly over-penalized techniques for Reissner–Mindlin plates. Journal of Scientific Computing, 64(2):401–424, August 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9936-8;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9936-
8.pdf.

Rubio:2015:QPT

[1436] Gonzalo Rubio, François Fraysse, David A. Kopriva, and Eusebio Valero. Quasi-a priori truncation error estimation in the DGSEM. Journal of Scientific Computing, 64(2):425–455, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9938-6;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9938-6.pdf.

Guo:2015:GRC

[1437] Hailong Guo and Zhimin Zhang. Gradient recovery for the Crouzeix– Raviart element. Journal of Scientific Computing, 64(2):456–476, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9939-5;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9939-5.pdf.

Huang:2015:FOM

[1438] Chieh-Sen Huang, Feng Xiao, and Todd Arbogast. Fifth order multimoment WENO schemes for hyperbolic conservation laws. Journal of Scientific Computing, 64(2):477–507, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9940-z;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9940-
z.pdf.

Jeon:2015:HDM

[1439] Youngmok Jeon. Hybrid difference methods for PDEs. Journal of Scientific Computing, 64(2):508–521, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9941-y;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9941-
y.pdf.

Schutz:2015:FSS

[1440] Jochen Schütz and Sebastian Noelle. Flux splitting for stiff equations: a notion on stability. Journal of Scientific Computing, 64(2):522–540, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9942-x;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9942-x.pdf.

Linke:2015:GEE

[1441] A. Linke and C. Merdon. Guaranteed energy error estimators for a modified robust Crouzeix–Raviart Stokes element. Journal of Scientific Computing, 64(2):541–558, August 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9943-9;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9943-
9.pdf.

Zhang:2015:WGF

[1442] Ran Zhang and Qilong Zhai. A weak Galerkin finite element scheme for the biharmonic equations by using polynomials of reduced order. Journal of Scientific Computing, 64(2):559–585, August 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9945-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9945-7.pdf.

Hesthaven:2015:HPT

[1443] Jan S. Hesthaven, Jae-Hun Jung, and Allen Tesdall. Hyperbolic problems: Theory and computation. Journal of Scientific Computing, 64(3): 587–590, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0065-9;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0065-9.pdf.

Chen:2015:LFM

[1444] Weitao Chen, Ching-Shan Chou, and Chiu-Yen Kao. Lax–Friedrichs multigrid fast sweeping methods for steady state problems for hyperbolic conservation laws. Journal of Scientific Computing, 64(3):591– 618, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0006-7;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0006-7.pdf. See erratum [1445].

Chen:2015:ELF

[1445] Weitao Chen, Ching-Shan Chou, and Chiu-Yen Kao. Erratum to: Lax– Friedrichs multigrid fast sweeping methods for steady state problems for hyperbolic conservation laws. Journal of Scientific Computing, 64 (3):619, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0025-4;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0025-4.pdf. See [1444].

Antoine:2015:DDM

[1446] X. Antoine, E. Lorin, and A. D. Bandrauk. Domain decomposition method and high-order absorbing boundary conditions for the numerical simulation of the time dependent Schrödinger equation with ionization and recombination by intense electric field. Journal of Scientific Computing, 64(3):620–646, September 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9902-5;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9902-
5.pdf.

Field:2015:FEF

[1447] Scott E. Field and Stephen R. Lau. Fast evaluation of far-field signals for time-domain wave propagation. Journal of Scientific Computing, 64(3): 647–669, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-9995-5;   http://link.springer.com/content/
pdf/10.1007/s10915-015-9995-5.pdf.

Li:2015:HFC

[1448] Peng Li, Zhen Gao, Wai-Sun Don, and Shusen Xie. Hybrid Fouriercontinuation method and weighted essentially non-oscillatory finite difference scheme for hyperbolic conservation laws in a single-domain framework. Journal of Scientific Computing, 64(3):670–695, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9913-2;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9913-2.pdf.

Li:2015:NSH

[1449] Xuefeng Li and Katarzyna Saxton. Non-strictly hyperbolic systems, singularity and bifurcation. Journal of Scientific Computing, 64(3): 696–720, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9876-3;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9876-3.pdf.

Tesdall:2015:FRG

[1450] Allen M. Tesdall, Richard Sanders, and Nedyu Popivanov. Further results on Guderley Mach reflection and the triple point paradox. Journal of Scientific Computing, 64(3):721–744, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0028-1;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0028-1.pdf.

Jia:2015:IAC

[1451] Junxiong Jia and Ronghua Pan. On isentropic approximations for compressible Euler equations. Journal of Scientific Computing, 64(3):745– 760, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9843-z;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9843-z.pdf.

Temple:2015:NMF

[1452] Blake Temple and Robin Young. A Nash–Moser framework for finding periodic solutions of the compressible Euler equations. Journal of Scientific Computing, 64(3):761–772, September 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9851-z;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9851-
z.pdf.

Ryan:2015:OSP

[1453] Jennifer K. Ryan, Xiaozhou Li, Robert M. Kirby, and Kees Vuik. Onesided position-dependent smoothness-increasing accuracy-conserving (SIAC) filtering over uniform and non-uniform meshes. Journal of Scientific Computing, 64(3):773–817, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9946-6;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9946-
6.pdf.

Choi:2015:PEN

[1454] D. Choï, L. Gallimard, and T. Sassi. A posteriori estimates for a natural Neumann–Neumann domain decomposition algorithm on a unilateral contact problem. Journal of Scientific Computing, 64(3):818– 836, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9944-8;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9944-8.pdf.

Kao:2015:FEO

[1455] Chiu-Yen Kao, Alexander Kurganov, Zhuolin Qu, and Ying Wang. A fast explicit operator splitting method for modified Buckley–Leverett equations. Journal of Scientific Computing, 64(3):837–857, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9950-x;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9950-x.pdf.

Das:2015:EAS

[1456] R. Das and P. W. Cleary. Evaluation of accuracy and stability of the classical SPH method under uniaxial compression. Journal of Scientific Computing, 64(3):858–897, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9948-4;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9948-
4.pdf.

Jha:2015:FSO

[1457] Navnit Jha and Leslaw K. Bieniasz. A fifth (six) order accurate, three–point compact finite difference scheme for the numerical solution of sixth order boundary value problems on geometric meshes. Journal of Scientific Computing, 64(3):898–913, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9947-5;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9947-5.pdf.

Liu:2015:HRE

[1458] Youqiong Liu, Jianhu Feng, and Jiong Ren. High resolution, entropy– consistent scheme using flux limiter for hyperbolic systems of conservation laws. Journal of Scientific Computing, 64(3):914–937, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9949-3;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9949-3.pdf.

Montijano:2015:FFE

[1459] J. I. Montijano, L. Rández, M. Van Daele, and M. Calvo. Functionally fitted explicit two step peer methods. Journal of Scientific Computing, 64(3):938–958, September 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9951-9;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9951-
9.pdf.

Ji:2015:HOC

[1460] Cui cui Ji and Zhi zhong Sun. A high-order compact finite difference scheme for the fractional sub-diffusion equation. Journal of Scientific Computing, 64(3):959–985, September 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9956-4;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9956-
4.pdf.

Shen:2015:SPE

[1461] Wanfang Shen, Liang Ge, Danping Yang, and Wenbin Liu. Sharp a posteriori error estimates for optimal control governed by parabolic integro– differential equations. Journal of Scientific Computing, 65(1):1–33, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9957-3;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9957-3.pdf.

Lemoine:2015:DHH

[1462] A. Lemoine, J.-P. Caltagirone, M. Azaïez, and S. Vincent. Discrete Helmholtz–Hodge decomposition on polyhedral meshes using compatible discrete operators. Journal of Scientific Computing, 65(1):34–53, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9952-8;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9952-8.pdf.

Pinto:2015:TSP

[1463] Martin Campos Pinto. Towards smooth particle methods without smoothing. Journal of Scientific Computing, 65(1):54–82, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9953-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9953-7.pdf.

Guo:2015:MPS

[1464] Yan Guo, Tao Xiong, and Yufeng Shi. A maximum–principlesatisfying high-order finite volume compact WENO scheme for scalar conservation laws with applications in incompressible flows. Journal of Scientific Computing, 65(1):83–109, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9954-6;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9954-
6.pdf.

Kim:2015:CCH

[1465] Kwangil Kim and Yonghai Li. Construction of convergent high order schemes for time dependent Hamilton–Jacobi equations. Journal of Scientific Computing, 65(1):110–137, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9955-5;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9955-
5.pdf.

Du:2015:RRT

[1466] Shaohong Du and Zhimin Zhang. A robust residual-type a posteriori error estimator for convection–diffusion equations. Journal of Scientific Computing, 65(1):138–170, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9972-4;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9972-
4.pdf.

Boulakia:2015:SSE

[1467] Muriel Boulakia, Alexandre Genadot, and Michèle Thieullen. Simulation of SPDEs for excitable media using finite elements. Journal of Scientific Computing, 65(1):171–195, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9960-8;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9960-
8.pdf.

Su:2015:SPN

[1468] Hongling Su and Shengtai Li. Structure-preserving numerical methods for infinite-dimensional Birkhoffian systems. Journal of Scientific Computing, 65(1):196–223, October 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9958-2;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9958-
2.pdf.

Sengupta:2015:RTI

[1469] Tapan K. Sengupta and V. K. Sathyanarayanan. Role of time integration in computing transitional flows caused by wall excitation. Journal of Scientific Computing, 65(1):224–248, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9967-1;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9967-1.pdf.

Gatto:2015:NAF

[1470] Paolo Gatto and Jan S. Hesthaven. Numerical approximation of the fractional Laplacian via hp-finite elements, with an application to image denoising. Journal of Scientific Computing, 65(1):249–270, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9959-1;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9959-1.pdf.

Izzo:2015:SSP

[1471] Giuseppe Izzo and Zdzisław Jackiewicz. Strong stability preserving general linear methods. Journal of Scientific Computing, 65(1):271–298, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9961-7;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9961-7.pdf.

Guo:2015:NLW

[1472] Wei Guo, Jing-Mei Qiu, and Jianxian Qiu. A new Lax–Wendroff discontinuous Galerkin method with superconvergence. Journal of Scientific Computing, 65(1):299–326, October 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9968-0;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9968-
0.pdf.

Oikawa:2015:HDG

[1473] Issei Oikawa. A hybridized discontinuous Galerkin method with reduced stabilization. Journal of Scientific Computing, 65(1):327–340, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9962-6;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9962-6.pdf.

Li:2015:SPP

[1474] Jichun Li and Shuyu Sun. The superconvergence phenomenon and proof of the MAC scheme for the Stokes equations on non-uniform rectangular meshes. Journal of Scientific Computing, 65(1):341–362, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9963-5;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9963-5.pdf.

Mu:2015:WGF

[1475] Lin Mu, Junping Wang, Xiu Ye, and Shangyou Zhang. A weak Galerkin finite element method for the Maxwell equations. Journal of Scientific Computing, 65(1):363–386, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9964-4;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9964-
4.pdf.

Dolz:2015:MHM

[1476] J. Dölz, H. Harbrecht, and M. Peters. H-matrix accelerated second moment analysis for potentials with rough correlation. Journal of Scientific Computing, 65(1):387–410, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9965-3;
 http://link.springer.com/content/pdf/10.1007/s10915-014-9965-
3.pdf.

Zeng:2015:SOS

[1477] Fanhai Zeng. Second-order stable finite difference schemes for the timefractional diffusion-wave equation. Journal of Scientific Computing, 65 (1):411–430, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9966-2;   http://link.springer.com/content/
pdf/10.1007/s10915-014-9966-2.pdf.

Kim:2015:EPR

[1478] Ki-Hong Kim, Min cheol Gwak, and Jack J. Yoh. An enhanced particle reseeding algorithm for the hybrid particle level set method in compressible flows. Journal of Scientific Computing, 65(1):431–453, October 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9970-6;   http://link.springer.com/content/pdf/10.1007/
s10915-014-9970-6.pdf.

Zhang:2015:HMN

[1479] Zhimin Zhang. How many numerical eigenvalues can we trust? Journal of Scientific Computing, 65(2):455–466, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-014-9971-5.

Chen:2015:RIB

[1480] Jeng-Tzong Chen, Yu-Lung Chang, Shing-Kai Kao, and Jie Jian. Revisit of the indirect boundary element method: Necessary and sufficient formulation. Journal of Scientific Computing, 65(2):467–485, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
014-9974-2.

Nissen:2015:SDM

[1481] Anna Nissen, Katharina Kormann, Magnus Grandin, and Kristoffer Virta. Stable difference methods for block-oriented adaptive grids. Journal of Scientific Computing, 65(2):486–511, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-014-9969-z.

Zheng:2015:LPF

[1482] Haibiao Zheng, Jiaping Yu, and Feng Shi. Local and parallel finite element algorithm based on the partition of unity for incompressible flows. Journal of Scientific Computing, 65(2):512–532, November 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9979-x.

Wasserman:2015:IRF

[1483] Gabriel Wasserman, Rick Archibald, and Anne Gelb. Image reconstruction from Fourier data using sparsity of edges. Journal of Scientific Computing, 65(2):533–552, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-014-9973-3.

Wang:2015:FSB

[1484] Rui Wang and Xiangling Chen. A fast solver for boundary integral equations of the modified Helmholtz equation. Journal of Scientific Computing, 65(2):553–575, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-014-9975-1.

Choi:2015:PFS

[1485] Youngsoo Choi, Charbel Farhat, Walter Murray, and Michael Saunders. A practical factorization of a Schur complement for PDE-constrained distributed optimal control. Journal of Scientific Computing, 65(2): 576–597, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-014-9976-0.

Guzman:2015:MNS

[1486] Johnny Guzmán and Manuel A. Sánchez. Max-norm stability of low order Taylor–Hood elements in three dimensions. Journal of Scientific Computing, 65(2):598–621, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-014-9978-y.

Guo:2015:ECL

[1487] Li Guo and Yan Xu. Energy conserving local discontinuous Galerkin methods for the nonlinear Schrödinger equation with wave operator. Journal of Scientific Computing, 65(2):622–647, November 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-014-9977-z.

Guo:2015:GHS

[1488] Ben yu Guo and Chao Zhang. Generalized Hermite spectral method matching different algebraic decay at infinities. Journal of Scientific Computing, 65(2):648–671, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-014-9981-3.

Bernardi:2015:PAI

[1489] Christine Bernardi, Jad Dakroub, Gihane Mansour, and Toni Sayah. A posteriori analysis of iterative algorithms for a nonlinear problem. Journal of Scientific Computing, 65(2):672–697, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-014-9980-4.

Caliari:2015:IPM

[1490] M. Caliari and S. Zuccher. The inverse power method for the p(x)-Laplacian problem. Journal of Scientific Computing, 65(2):698– 714, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-9982-x.

Yi:2015:VCP

[1491] Lijun Yi. An h-p version of the continuous Petrov–Galerkin finite element method for nonlinear Volterra integro-differential equations. Journal of Scientific Computing, 65(2):715–734, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-9983-9.

Bernard:2015:AAP

[1492] Florian Bernard, Angelo Iollo, and Gabriella Puppo. Accurate asymptotic preserving boundary conditions for kinetic equations on Cartesian grids. Journal of Scientific Computing, 65(2):735–766, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-9984-8.

Hochstenbach:2015:GKT

[1493] Michiel E. Hochstenbach, Lothar Reichel, and Xuebo Yu. A Golub– Kahan-type reduction method for matrix pairs. Journal of Scientific Computing, 65(2):767–789, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-9990-x.

Sun:2015:HOB

[1494] Zhen sheng Sun, Yu xin Ren, Bai lin Zha, and Shi ying Zhang. High order boundary conditions for high order finite difference schemes on curvilinear coordinates solving compressible flows. Journal of Scientific Computing, 65(2):790–820, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-9988-4.

Reiss:2015:FES

[1495] Julius Reiss. A family of energy stable, skew-symmetric finite difference schemes on collocated grids. Journal of Scientific Computing, 65(2): 821–838, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-9985-7.

Chen:2015:HOA

[1496] Sheng-Gwo Chen, Mei-Hsiu Chi, and Jyh-Yang Wu. High-order algorithms for Laplace–Beltrami operators and geometric invariants over curved surfaces. Journal of Scientific Computing, 65(2):839–865, November 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-9986-6.

Magdalena:2015:SMC

[1497] I. Magdalena, N. Erwina, and S. R. Pudjaprasetya. Staggered momentum conservative scheme for radial dam break simulation. Journal of Scientific Computing, 65(3):867–874, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/article/10.1007/s10915-015-9987-5.

Lin:2015:PEE

[1498] Tao Lin, Qing Yang, and Xu Zhang. A priori error estimates for some discontinuous Galerkin immersed finite element methods. Journal of Scientific Computing, 65(3):875–894, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/article/10.1007/s10915-015-9989-3.

Porta:2015:NSS

[1499] Federica Porta, Marco Prato, and Luca Zanni. A new steplength selection for scaled gradient methods with application to image deblurring. Journal of Scientific Computing, 65(3):895–919, December 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-9991-9.

Zhang:2015:TLC

[1500] Guo-Dong Zhang, Yan Zhang, and Yinnian He. Two-level coupled and decoupled parallel correction methods for stationary incompressible magnetohydrodynamics. Journal of Scientific Computing, 65(3): 920–939, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-9994-6.

Notsu:2015:EEP

[1501] Hirofumi Notsu and Masahisa Tabata. Error estimates of a pressurestabilized characteristics finite element scheme for the Oseen equations. Journal of Scientific Computing, 65(3):940–955, December 2015. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-9992-8.

Kumar:2015:EOD

[1502] Sarvesh Kumar and Ricardo Ruiz-Baier. Equal order discontinuous finite volume element methods for the Stokes problem. Journal of Scientific Computing, 65(3):956–978, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-9993-7.

Kadalbajoo:2015:SOA

[1503] Mohan K. Kadalbajoo, Lok Pati Tripathi, and Alpesh Kumar. Second order accurate IMEX methods for option pricing under Merton and Kou jump-diffusion models. Journal of Scientific Computing, 65(3): 979–1024, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0001-z.

Shi:2015:NSM

[1504] Dong yang Shi, Ming hao Li, and Chao Xu. A new stabilization method for the elasticity problem. Journal of Scientific Computing, 65(3):1025– 1038, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-9996-4.

Dauge:2015:TNI

[1505] Monique Dauge, Alexander Düster, and Ernst Rank. Theoretical and numerical investigation of the finite cell method. Journal of Scientific Computing, 65(3):1039–1064, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-9997-3.

Choi:2015:FDC

[1506] Pui Tung Choi and Lok Ming Lui. Fast disk conformal parameterization of simply–connected open surfaces. Journal of Scientific Computing, 65(3):1065–1090, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-9998-2.

Crouseilles:2015:CNS

[1507] Nicolas Crouseilles, Matthieu Kuhn, and Guillaume Latu. Comparison of numerical solvers for anisotropic diffusion equations arising in plasma physics. Journal of Scientific Computing, 65(3):1091–1128, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-9999-1.

Guillen-Gonzalez:2015:STS

[1508] F. Guillén-González and J. Koko. A splitting in time scheme and augmented Lagrangian method for a nematic liquid crystal problem. Journal of Scientific Computing, 65(3):1129–1144, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0002-y.

Chen:2015:REA

[1509] Zheng Chen and Chi-Wang Shu. Recovering exponential accuracy in Fourier spectral methods involving piecewise smooth functions with unbounded derivative singularities. Journal of Scientific Computing, 65(3): 1145–1165, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0011-x.

Khebchareon:2015:ADI

[1510] Morrakot Khebchareon, Amiya K. Pani, and Graeme Fairweather. Alternating direction implicit Galerkin methods for an evolution equation with a positive-type memory term. Journal of Scientific Computing, 65(3): 1166–1188, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0004-9.

Xue:2015:SFE

[1511] Yunhua Xue, Cheng Wang, and Jian-Guo Liu. Simple finite element numerical simulation of incompressible flow over non-rectangular domains and the super-convergence analysis. Journal of Scientific Computing, 65(3):1189–1216, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0005-8.

Fairweather:2015:ACN

[1512] Graeme Fairweather, Xuehua Yang, Da Xu, and Haixiang Zhang. An ADI Crank–Nicolson orthogonal spline collocation method for the twodimensional fractional diffusion-wave equation. Journal of Scientific Computing, 65(3):1217–1239, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-0003-x.

Liu:2015:KRB

[1513] Xiao-Yan Liu, Andreas Karageorghis, and C. S. Chen. A Kansa-radial basis function method for elliptic boundary value problems in annular domains. Journal of Scientific Computing, 65(3):1240–1269, December 2015. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0009-4.

Gatica:2015:AAH

[1514] Gabriel N. Gatica and Filánder A. Sequeira. Analysis of an augmented HDG method for a class of quasi-Newtonian Stokes flows. Journal of Scientific Computing, 65(3):1270–1308, December 2015. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/article/10.1007/s10915-015-0008-5.

He:2016:HST

[1515] Wenming He, Zhimin Zhang, and Ren Zhao. The highest superconvergence of the tri-linear element for Schrödinger operator with singularity. Journal of Scientific Computing, 66(1):1–18, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0007-6.

Yao:2016:LSM

[1516] Xudong Yao. Ljusternik–Schnirelman minimax algorithms and an application for finding multiple negative energy solutions of semilinear elliptic Dirichlet problem involving concave and convex nonlinearities: Part I. Algorithms and convergence. Journal of Scientific Computing, 66(1):19–40, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0010-y.

Pang:2016:FNC

[1517] Hong-Kui Pang and Hai-Wei Sun. Fast numerical contour integral method for fractional diffusion equations. Journal of Scientific Computing, 66(1):41–66, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0012-9.

Chen:2016:RRB

[1518] Yanlai Chen, Sigal Gottlieb, Alfa Heryudono, and Akil Narayan. A reduced radial basis function method for partial differential equations on irregular domains. Journal of Scientific Computing, 66(1):67–90, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0013-8.

Celledoni:2016:HOS

[1519] Elena Celledoni, Bawfeh Kingsley Kometa, and Olivier Verdier. High order semi-Lagrangian methods for the incompressible Navier–Stokes equations. Journal of Scientific Computing, 66(1):91–115, January 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0015-6.

Kolomenskiy:2016:AGA

[1520] Dmitry Kolomenskiy, Jean-Christophe Nave, and Kai Schneider. Adaptive gradient-augmented level set method with multiresolution error estimation. Journal of Scientific Computing, 66(1):116–140, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0014-7.

Cai:2016:NAA

[1521] Jiaxiang Cai, Yushun Wang, and Yuezheng Gong. Numerical analysis of AVF methods for three-dimensional time-domain Maxwell’s equations. Journal of Scientific Computing, 66(1):141–176, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0016-5.

Fok:2016:LFO

[1522] Pak-Wing Fok. A linearly fourth order multirate Runge–Kutta method with error control. Journal of Scientific Computing, 66(1):177–195, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0017-4.

Chen:2016:WFB

[1523] Dai-Qiang Chen and Yan Zhou. Wavelet frame based image restoration via combined sparsity and nonlocal prior of coefficients. Journal of Scientific Computing, 66(1):196–224, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-0018-3.

Ginting:2016:AWG

[1524] Victor Ginting, Guang Lin, and Jiangguo Liu. On application of the weak Galerkin finite element method to a two-phase model for subsurface flow. Journal of Scientific Computing, 66(1):225–239, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0021-8.

Boscheri:2016:EQF

[1525] W. Boscheri and M. Dumbser. An efficient quadrature-free formulation for high order arbitrary-Lagrangian–Eulerian ADER–WENO finite volume schemes on unstructured meshes. Journal of Scientific Computing, 66(1):240–274, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0019-2.

Blayo:2016:TOS

[1526] Eric Blayo, David Cherel, and Antoine Rousseau. Towards optimized Schwarz methods for the Navier–Stokes equations. Journal of Scientific Computing, 66(1):275–295, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0020-9.

Olson:2016:PMD

[1527] D. Olson, S. Shukla, G. Simpson, and D. Spirn. Petviashvilli’s method for the Dirichlet problem. Journal of Scientific Computing, 66(1): 296–320, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0023-6.

Dong:2016:NMD

[1528] Bo Dong, Chi-Wang Shu, and Wei Wang. A new multiscale discontinuous Galerkin method for the one-dimensional stationary Schrödinger equation. Journal of Scientific Computing, 66(1):321–345, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0022-7.

Qiu:2016:HMC

[1529] Weifeng Qiu and Ke Shi. An HDG method for convection diffusion equation. Journal of Scientific Computing, 66(1):346–357, January 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0024-5.

Geraci:2016:NWI

[1530] Gianluca Geraci, Pietro Marco Congedo, Rémi Abgrall, and Gianluca Iaccarino. A novel weakly-intrusive non-linear multiresolution framework for uncertainty quantification in hyperbolic partial differential equations. Journal of Scientific Computing, 66(1):358–405, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0026-3.

Michoski:2016:CAV

[1531] C. Michoski, C. Dawson, E. J. Kubatko, D. Wirasaet, S. Brus, and J. J. Westerink. A comparison of artificial viscosity, limiters, and filters, for high order discontinuous Galerkin solutions in nonlinear settings. Journal of Scientific Computing, 66(1):406–434, January 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/article/10.1007/s10915-015-0027-2.

Zhao:2016:SAS

[1532] Tao Zhao. A spectral analysis of subspace enhanced preconditioners. Journal of Scientific Computing, 66(1):435–457, January 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0029-0.

Baccouch:2016:RBE

[1533] Mahboub Baccouch. Recovery-based error estimator for the discontinuous Galerkin method for nonlinear scalar conservation laws in one space dimension. Journal of Scientific Computing, 66(2):459–476, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0030-7;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0030-7.pdf.

Yu:2016:SMF

[1534] Xu hong Yu and Ben yu Guo. Spectral method for fourth-order problems on quadrilaterals. Journal of Scientific Computing, 66(2):477–503, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0031-6;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0031-6.pdf.

Gao:2016:UOE

[1535] Huadong Gao. Unconditional optimal error estimates of BDF– Galerkin FEMs for nonlinear thermistor equations. Journal of Scientific Computing, 66(2):504–527, February 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0032-5;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0032-
5.pdf.

deFrutos:2016:LEE

[1536] Javier de Frutos, Bosco García-Archilla, and Julia Novo. Local error estimates for the SUPG method applied to evolutionary convection– reaction-diffusion equations. Journal of Scientific Computing, 66(2): 528–554, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0035-2;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0035-2.pdf.

Luo:2016:SCN

[1537] Zhen Dong Luo. A stabilized Crank–Nicolson mixed finite volume element formulation for the non-stationary incompressible Boussinesq equations. Journal of Scientific Computing, 66(2):555–576, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0034-3;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0034-3.pdf.

Mancini:2016:EAA

[1538] Sara Mancini, Francisco Bernal, and Juan A. Acebrón. An efficient algorithm for accelerating Monte Carlo approximations of the solution to boundary value problems. Journal of Scientific Computing, 66(2): 577–597, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0033-4;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0033-4.pdf.

Liu:2016:FDH

[1539] Hongxia Liu and Jianxian Qiu. Finite difference Hermite WENO schemes for conservation laws, II: an alternative approach. Journal of Scientific Computing, 66(2):598–624, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0041-4;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0041-
4.pdf.

Lee:2016:ASD

[1540] Jeonghun J. Lee and Hyea Hyun Kim. Analysis of a staggered discontinuous Galerkin method for linear elasticity. Journal of Scientific Computing, 66(2):625–649, February 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0036-1;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0036-
1.pdf.

Jung:2016:NTD

[1541] Chang-Yeol Jung and Thien Binh Nguyen. New time differencing methods for spectral methods. Journal of Scientific Computing, 66(2): 650–671, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0037-0;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0037-0.pdf.

Gunzburger:2016:SSS

[1542] Max D. Gunzburger, Lisheng Hou, and Ju Ming. Stochastic steadystate Navier–Stokes equations with additive random noise. Journal of Scientific Computing, 66(2):672–691, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0039-y;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0039-
y.pdf.

Semplice:2016:AMR

[1543] M. Semplice, A. Coco, and G. Russo. Adaptive mesh refinement for hyperbolic systems based on third-order compact WENO reconstruction. Journal of Scientific Computing, 66(2):692–724, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0038-z;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0038-z.pdf.

Vong:2016:CDS

[1544] Seakweng Vong, Pin Lyu, and Zhibo Wang. A compact difference scheme for fractional sub-diffusion equations with the spatially variable coefficient under Neumann boundary conditions. Journal of Scientific Computing, 66(2):725–739, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0040-5;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0040-
5.pdf.

Li:2016:TRC

[1545] Jin Li, Hongxing Rui, and Dehao Yu. Trapezoidal rule for computing supersingular integral on a circle. Journal of Scientific Computing, 66 (2):740–760, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0042-3;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0042-3.pdf.

Baeza:2016:HOB

[1546] A. Baeza, P. Mulet, and D. Zorío. High order boundary extrapolation technique for finite difference methods on complex domains with Cartesian meshes. Journal of Scientific Computing, 66(2):761–791, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

015-0043-2;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0043-2.pdf.

Berthelin:2016:MFK

[1547] Florent Berthelin, Thierry Goudon, and Sebastian Minjeaud. Multifluid flows: a kinetic approach. Journal of Scientific Computing, 66 (2):792–824, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0044-1;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0044-1.pdf.

Lepe:2016:FEA

[1548] F. Lepe, D. Mora, and R. Rodríguez. Finite element analysis of a bending moment formulation for the vibration problem of a nonhomogeneous Timoshenko beam. Journal of Scientific Computing, 66 (2):825–848, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0046-z;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0046-z.pdf.

Jin:2016:ADM

[1549] Zheng-Fen Jin, Zhongping Wan, Yuling Jiao, and Xiliang Lu. An alternating direction method with continuation for nonconvex low rank minimization. Journal of Scientific Computing, 66(2):849–869, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0045-0;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0045-0.pdf.

Chung:2016:SDM

[1550] Eric Chung, Bernardo Cockburn, and Guosheng Fu. The staggered DG method is the limit of a hybridizable DG method. Part II: The Stokes flow. Journal of Scientific Computing, 66(2):870–887, February 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0047-y;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0047-y.pdf.

Deng:2016:GLC

[1551] Wei Deng and Wotao Yin. On the global and linear convergence of the generalized alternating direction method of multipliers. Journal of Scientific Computing, 66(3):889–916, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-0048-x.

Carr:2016:MFP

[1552] L. E. Carr III, C. F. Borges, and F. X. Giraldo. Matrix-free polynomialbased nonlinear least squares optimized preconditioning and its application to discontinuous Galerkin discretizations of the Euler equations. Journal of Scientific Computing, 66(3):917–940, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0049-9.

Gong:2016:FEM

[1553] Wei Gong, Michael Hinze, and Zhaojie Zhou. Finite element method and a priori error estimates for Dirichlet boundary control problems governed by parabolic PDEs. Journal of Scientific Computing, 66(3):941–967, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0051-2.

Fu:2016:SMF

[1554] Hongfei Fu, Hongxing Rui, Jian Hou, and Haihong Li. A stabilized mixed finite element method for elliptic optimal control problems. Journal of Scientific Computing, 66(3):968–986, March 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/article/10.1007/s10915-015-0050-3. See erra-

tum [1555].

Fu:2016:ESM

[1555] Hongfei Fu, Hongxing Rui, Jian Hou, and Haihong Li. Erratum to: A Stabilized Mixed Finite Element Method for Elliptic Optimal Control Problems. Journal of Scientific Computing, 66(3):987–990, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0086-4;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0086-4.pdf. See [1554].

deFrutos:2016:GDS

[1556] Javier de Frutos, Bosco García-Archilla, Volker John, and Julia Novo. Grad-div stabilization for the evolutionary Oseen problem with inf-sup stable finite elements. Journal of Scientific Computing, 66(3):991–1024, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0052-1.

Lee:2016:FST

[1557] Byungjoon Lee and Myungjoo Kang. Full 3D simulations of two-phase core-annular flow in horizontal pipe using level set method. Journal of Scientific Computing, 66(3):1025–1051, March 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0053-0.

Puppo:2016:WBH

[1558] G. Puppo and M. Semplice. Well-balanced high order 1D schemes on nonuniform grids and entropy residuals. Journal of Scientific Computing, 66 (3):1052–1076, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0056-x.

Guo:2016:SMN

[1559] Ben yu Guo and Yu jian Jiao. Spectral method for Navier–Stokes equations with non-slip boundary conditions by using divergence-free base functions. Journal of Scientific Computing, 66(3):1077–1101, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0054-z.

Han:2016:DUS

[1560] Daozhi Han. A decoupled unconditionally stable numerical scheme for the Cahn–Hilliard–Hele–Shaw system. Journal of Scientific Computing, 66(3):1102–1121, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0055-y.

Warin:2016:SNM

[1561] Xavier Warin. Some non-monotone schemes for time dependent Hamilton–Jacobi–Bellman equations in stochastic control. Journal of Scientific Computing, 66(3):1122–1147, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/article/10.1007/s10915-015-0057-9.

Ji:2016:NAH

[1562] Cui cui Ji, Zhi zhong Sun, and Zhao peng Hao. Numerical algorithms with high spatial accuracy for the fourth-order fractional sub-diffusion equations with the first Dirichlet boundary conditions. Journal of Scientific Computing, 66(3):1148–1174, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-0059-7.

Wen:2016:TPM

[1563] Zaiwen Wen, Chao Yang, Xin Liu, and Yin Zhang. Trace-penalty minimization for large-scale eigenspace computation. Journal of Scientific Computing, 66(3):1175–1203, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-0061-0.

He:2016:PJD

[1564] Bingsheng He, Hong-Kun Xu, and Xiaoming Yuan. On the proximal Jacobian decomposition of ALM for multiple-block separable convex minimization problems and its relationship to ADMM. Journal of Scientific Computing, 66(3):1204–1217, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-0060-1.

Baarman:2016:DME

[1565] K. Baarman, V. Havu, and T. Eirola. Direct minimization for ensemble electronic structure calculations. Journal of Scientific Computing, 66 (3):1218–1233, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0058-8.

Bhatt:2016:SOC

[1566] Ashish Bhatt, Dwayne Floyd, and Brian E. Moore. Second order conformal symplectic schemes for damped Hamiltonian systems. Journal of Scientific Computing, 66(3):1234–1259, March 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0062-z.

Antonietti:2016:DGA

[1567] P. F. Antonietti, M. Grasselli, S. Stangalino, and M. Verani. Discontinuous Galerkin approximation of linear parabolic problems with dynamic boundary conditions. Journal of Scientific Computing, 66(3):1260–1280, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0063-y.

Gao:2016:TADa

[1568] Guang hua Gao and Zhi zhong Sun. Two alternating direction implicit difference schemes for two-dimensional distributed-order fractional diffusion equations. Journal of Scientific Computing, 66(3):1281–1312, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0064-x.

Gurkan:2016:EHD

[1569] Ceren Gürkan, Esther Sala-Lardies, Martin Kronbichler, and Sonia Fernández-Méndez. eXtended hybridizable discontinuous Galerkin (X-HDG) for void problems. Journal of Scientific Computing, 66(3): 1313–1333, March 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0066-8.

Ma:2016:IDT

[1570] Liyan Ma and Tieyong Zeng. Image deblurring via total variation based structured sparse model selection. Journal of Scientific Computing, 67(1):1–19, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0067-7.

Lamichhane:2016:MFE

[1571] Bishnu P. Lamichhane, Stephen G. Roberts, and Linda Stals. A mixed finite element discretisation of thin plate splines based on biorthogonal systems. Journal of Scientific Computing, 67(1):20–42, April 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0068-6.

Li:2016:NSW

[1572] Xianjuan Li, Tao Tang, and Chuanju Xu. Numerical solutions for weakly singular Volterra integral equations using Chebyshev and Legendre pseudo-spectral Galerkin methods. Journal of Scientific Computing, 67(1):43–64, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0069-5.

Chen:2016:GSA

[1573] Yanping Chen and Fenglin Huang. Galerkin spectral approximation of elliptic optimal control problems with H^{1}-norm state constraint. Journal of Scientific Computing, 67(1):65–83, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/article/10.1007/s10915-015-0071-y.

Tong:2016:HOM

[1574] Oisin Tong, Aaron Katz, Yushi Yanagita, Alex Casey, and Robert Schaap. High-order methods for turbulent flows on three-dimensional strand grids. Journal of Scientific Computing, 67(1):84–102, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0070-z.

Hu:2016:RMG

[1575] Haijuan Hu, Bing Li, and Quansheng Liu. Removing mixture of Gaussian and impulse noise by patch-based weighted means. Journal of Scientific Computing, 67(1):103–129, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0073-9.

Du:2016:PEA

[1576] Yu Du and Lingxue Zhu. Preasymptotic error analysis of high order interior penalty discontinuous Galerkin methods for the Helmholtz equation with high wave number. Journal of Scientific Computing, 67(1):130–152, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0074-8.

Bi:2016:PEE

[1577] Chunjia Bi, Cheng Wang, and Yanping Lin. Pointwise error estimates and two-grid algorithms of discontinuous Galerkin method for strongly nonlinear elliptic problems. Journal of Scientific Computing, 67(1):153–175, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0072-x.

Kozelkov:2016:CCF

[1578] Andrey Kozelkov, Vadim Kurulin, Vladislav Emelyanov, Elena Tyatyushkina, and Konstantin Volkov. Comparison of convective flux discretization schemes in detached-eddy simulation of turbulent flows on unstructured meshes. Journal of Scientific Computing, 67(1):176–191, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0075-7.

Yakovlev:2016:CHC

[1579] Sergey Yakovlev, David Moxey, Robert M. Kirby, and Spencer J. Sherwin. To CG or to HDG: a comparative study in 3D. Journal of Scientific Computing, 67(1):192–220, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0076-6.

Barkouki:2016:ARB

[1580] H. Barkouki, A. H. Bentbib, and K. Jbilou. An adaptive rational block Lanczos-type algorithm for model reduction of large scale dynamical systems. Journal of Scientific Computing, 67(1):221–236, April 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0077-5.

Mirzargar:2016:SIA

[1581] Mahsa Mirzargar, Jennifer K. Ryan, and Robert M. Kirby. Smoothness-Increasing Accuracy-Conserving (SIAC) filtering and quasi-interpolation: a unified view. Journal of Scientific Computing, 67(1):237–261, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0081-9.

Zhang:2016:SGS

[1582] Qian Zhang, Zhilin Li, and Zhiyue Zhang. A sparse grid stochastic collocation method for elliptic interface problems with random input. Journal of Scientific Computing, 67(1):262–280, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0080-x.

He:2016:ISA

[1583] Hongjin He, Chen Ling, and Hong-Kun Xu. An implementable splitting algorithm for the ℓ_{1}-norm regularized split feasibility problem. Journal of Scientific Computing, 67(1):281–298, April 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/article/10.1007/s10915-015-0078-4.

Kim:2016:MNL

[1584] Chang Ho Kim, Youngsoo Ha, and Jungho Yoon. Modified non-linear weights for fifth-order weighted essentially non-oscillatory schemes. Journal of Scientific Computing, 67(1):299–323, April 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/article/10.1007/s10915-015-0079-3.

Kolev:2016:MPR

[1585] Tzanio V. Kolev, Jinchao Xu, and Yunrong Zhu. Multilevel preconditioners for reaction–diffusion problems with discontinuous coefficients. Journal of Scientific Computing, 67(1):324–350, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http:/

/link.springer.com/article/10.1007/s10915-015-0083-7.

Romero:2016:SFF

[1586] J. Romero, K. Asthana, and A. Jameson. A simplified formulation of the flux reconstruction method. Journal of Scientific Computing, 67 (1):351–374, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0085-5.

Lamouroux:2016:HOC

[1587] Raphaël Lamouroux, Jérémie Gressier, and Gilles Grondin. A highorder compact limiter based on spatially weighted projections for the spectral volume and the spectral differences method. Journal of Scientific Computing, 67(1):375–403, April 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

article/10.1007/s10915-015-0084-6.

Sun:2016:PEE

[1588] Tongjun Sun, Wanfang Shen, Benxue Gong, and Wenbin Liu. A priori error estimate of stochastic Galerkin method for optimal control problem governed by stochastic elliptic PDE with constrained control. Journal of Scientific Computing, 67(2):405–431, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0091-7;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0091-7.pdf.

Archibald:2016:IRU

[1589] Rick Archibald, Anne Gelb, and Rodrigo B. Platte. Image reconstruction from undersampled Fourier data using the polynomial annihilation transform. Journal of Scientific Computing, 67(2):432–452, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0088-2;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0088-2.pdf.

Alvarado:2016:CPL

[1590] A. Alvarado and P. Castillo. Computational performance of LDG methods applied to time harmonic Maxwell equation in polyhedral domains. Journal of Scientific Computing, 67(2):453–474, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0087-3;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0087-3.pdf.

Oikawa:2016:ARO

[1591] Issei Oikawa. Analysis of a reduced-order HDG method for the Stokes equations. Journal of Scientific Computing, 67(2):475–492, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0090-8;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0090-8.pdf.

Li:2016:WBD

[1592] Gang Li and Yulong Xing. Well-balanced discontinuous Galerkin methods for the Euler equations under gravitational fields. Journal of Scientific Computing, 67(2):493–513, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0093-5;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0093-
5.pdf.

Winters:2016:ESF

[1593] Andrew R. Winters and Gregor J. Gassner. An entropy stable finite volume scheme for the equations of shallow water magnetohydrodynamics. Journal of Scientific Computing, 67(2):514–539, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0092-6;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0092-6.pdf.

Wang:2016:NMW

[1594] Rong Wang, Hui Feng, and Cong Huang. A new mapped weighted essentially non-oscillatory method using rational mapping function. Journal of Scientific Computing, 67(2):540–580, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0095-3;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0095-
3.pdf.

Wu:2016:BBL

[1595] Lei Wu, Zhe Sun, and Dong-Hui Li. A Barzilai–Borwein-like iterative half thresholding algorithm for the L_{1/2} regularized problem. Journal of Scientific Computing, 67(2):581–601, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0094-4;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0094-
4.pdf.

Yoon:2016:CAS

[1596] Gangjoon Yoon and Chohong Min. Convergence analysis of the standard central finite difference method for Poisson equation. Journal of Scientific Computing, 67(2):602–617, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0096-2;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0096-
2.pdf.

Berthon:2016:WBF

[1597] Christophe Berthon, Anaïs Crestetto, and Françoise Foucher. A wellbalanced finite volume scheme for a mixed hyperbolic/parabolic system to model chemotaxis. Journal of Scientific Computing, 67(2):618–643, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0097-1;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0097-1.pdf.

Wu:2016:CAP

[1598] Shu-Lin Wu. Convergence analysis of the parareal-Euler algorithm for systems of ODEs with complex eigenvalues. Journal of Scientific Computing, 67(2):644–668, May 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0100-x;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0100-
x.pdf.

Duan:2016:AFM

[1599] Huoyuan Duan, Fengjuan Qiu, Roger C. E. Tan, and Weiying Zheng. An adaptive FEM for a Maxwell interface problem. Journal of Scientific Computing, 67(2):669–704, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0098-0;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0098-
0.pdf.

Sun:2016:NTS

[1600] Hui Sun, Shenggao Zhou, David K. Moore, Li-Tien Cheng, and Bo Li. Numerical treatment of Stokes solvent flow and solute–solvent interfacial dynamics for nonpolar molecules. Journal of Scientific Computing, 67(2):705–723, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0099-z;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0099-z.pdf.

Wang:2016:PMF

[1601] Wansheng Wang, Long Chen, and Jie Zhou. Postprocessing mixed finite element methods for solving Cahn–Hilliard equation: Methods and error analysis. Journal of Scientific Computing, 67(2):724–746, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0101-9;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0101-9.pdf.

Eigel:2016:EPE

[1602] M. Eigel and C. Merdon. Equilibration a posteriori error estimation for convection–diffusion-reaction problems. Journal of Scientific Computing, 67(2):747–768, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0108-2;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0108-2.pdf.

Sheshadri:2016:SFR

[1603] Abhishek Sheshadri and Antony Jameson. On the stability of the flux reconstruction schemes on quadrilateral elements for the linear advection equation. Journal of Scientific Computing, 67(2):769–790, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0102-8;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0102-8.pdf. See erratum [1604].

Sheshadri:2016:ESF

[1604] Abhishek Sheshadri and Antony Jameson. Erratum to: On the stability of the flux reconstruction schemes on quadrilateral elements for the linear advection equation. Journal of Scientific Computing, 67(2):791–794, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0197-6;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0197-6.pdf. See [1603].

Yang:2016:HOM

[1605] Pei Yang, Tao Xiong, Jing-Mei Qiu, and Zhengfu Xu. High order maximum principle preserving finite volume method for convection dominated problems. Journal of Scientific Computing, 67(2):795–820, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0104-6;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0104-6.pdf.

Liu:2016:NMG

[1606] Xinguo Liu and Jianping You. Numerical methods for the Genvar criterion of multiple-sets canonical analysis. Journal of Scientific Computing, 67(2):821–835, May 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0103-7;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0103-7.pdf.

Wong:2016:FSM

[1607] Tony Wong and Shingyu Leung. A fast sweeping method for eikonal equations on implicit surfaces. Journal of Scientific Computing, 67 (3):837–859, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0105-5;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0105-5.pdf.

Deng:2016:TLS

[1608] Xiaomao Deng, Xiao-Chuan Cai, and Jun Zou. Two-level space–time domain decomposition methods for three-dimensional unsteady inverse source problems. Journal of Scientific Computing, 67(3):860–882, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0109-1;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0109-1.pdf.

Qian:2016:BEH

[1609] Jianliang Qian, Lijun Yuan, Yuan Liu, Songting Luo, and Robert Burridge. Babich’s expansion and high-order Eulerian asymptotics for point-source Helmholtz equations. Journal of Scientific Computing, 67 (3):883–908, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0111-7;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0111-7.pdf.

Lee:2016:IBM

[1610] Seunggyu Lee, Darae Jeong, Wanho Lee, and Junseok Kim. An immersed boundary method for a contractile elastic ring in a three-dimensional Newtonian fluid. Journal of Scientific Computing, 67(3):909–925, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0110-8;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0110-8.pdf.

Lee:2016:LMT

[1611] Yin Tat Lee, Ka Chun Lam, and Lok Ming Lui. Landmarkmatching transformation with large deformation via n-dimensional quasiconformal maps. Journal of Scientific Computing, 67(3):926–954, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0113-5;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0113-5.pdf.

Beyn:2016:SSC

[1612] Wolf-Jürgen Beyn, Elena Isaak, and Raphael Kruse. Stochastic C-stability and B-consistency of explicit and implicit Eulertype schemes. Journal of Scientific Computing, 67(3):955–987, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0114-4;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0114-4.pdf.

Ahmed:2016:NSS

[1613] Naveed Ahmed and Gunar Matthies. Numerical study of SUPG and LPS methods combined with higher order variational time discretization schemes applied to time-dependent linear convection–diffusionreaction equations. Journal of Scientific Computing, 67(3):988–1018, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0115-3;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0115-3.pdf.

Higueras:2016:CAS

[1614] Inmaculada Higueras and Teo Roldán. Construction of additive semi-implicit Runge–Kutta methods with low-storage requirements. Journal of Scientific Computing, 67(3):1019–1042, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0116-2;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0116-2.pdf.

Zhu:2016:FHO

[1615] Liyong Zhu, Lili Ju, and Weidong Zhao. Fast high-order compact exponential time differencing Runge–Kutta methods for second-order semilinear parabolic equations. Journal of Scientific Computing, 67(3): 1043–1065, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0117-1;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0117-1.pdf.

Xiong:2016:PPP

[1616] Tao Xiong, Jing-Mei Qiu, and Zhengfu Xu. Parametrized positivity preserving flux limiters for the high order finite difference WENO scheme solving compressible Euler equations. Journal of Scientific Computing, 67(3):1066–1088, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0118-0;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0118-0.pdf.

Heryudono:2016:PRB

[1617] Alfa Heryudono, Elisabeth Larsson, Alison Ramage, and Lina von Sydow. Preconditioning for radial basis function partition of unity methods. Journal of Scientific Computing, 67(3):1089–1109, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0120-6;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0120-6.pdf.

Zhao:2016:LSC

[1618] Jingjun Zhao, Yang Cao, and Yang Xu. Legendre spectral collocation methods for Volterra delay–integro-differential equations. Journal of Scientific Computing, 67(3):1110–1133, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0121-5;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0121-
5.pdf.

Zhang:2016:GGP

[1619] Qinghai Zhang. GePUP: Generic projection and unconstrained PPE for fourth-order solutions of the incompressible Navier–Stokes equations with no-slip boundary conditions. Journal of Scientific Computing, 67 (3):1134–1180, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0122-4;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0122-4.pdf.

Hu:2016:GLB

[1620] Jun Hu, Yunqing Huang, and Rui Ma. Guaranteed lower bounds for eigenvalues of elliptic operators. Journal of Scientific Computing, 67 (3):1181–1197, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0126-0;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0126-0.pdf.

Jin:2016:WBS

[1621] Shi Jin, Dongbin Xiu, and Xueyu Zhu. A well-balanced stochastic Galerkin method for scalar hyperbolic balance laws with random inputs. Journal of Scientific Computing, 67(3):1198–1218, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0124-2;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0124-2.pdf.

Cravero:2016:AWC

[1622] I. Cravero and M. Semplice. On the accuracy of WENO and CWENO reconstructions of third order on nonuniform meshes. Journal of Scientific Computing, 67(3):1219–1246, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0123-3;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0123-
3.pdf.

Wang:2016:SMF

[1623] Tian jun Wang. A spectral method for fourth-order mixed inhomogeneous boundary value problem in three dimensions. Journal of Scientific Computing, 67(3):1247–1271, June 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0106-4;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0106-
4.pdf.

Mengaldo:2016:CBD

[1624] G. Mengaldo, D. De Grazia, P. E. Vincent, and S. J. Sherwin. On the connections between discontinuous Galerkin and flux reconstruction schemes: Extension to curvilinear meshes. Journal of Scientific Computing, 67(3):1272–1292, June 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0119-z;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0119-
z.pdf.

Du:2016:FGP

[1625] Ning Du, Hong Wang, and Wenbin Liu. A fast gradient projection method for a constrained fractional optimal control. Journal of Scientific Computing, 68(1):1–20, July 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0125-1;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0125-
1.pdf.

Janon:2016:GOE

[1626] Alexandre Janon, Maëlle Nodet, and Clémentine Prieur. Goal-oriented error estimation for the reduced basis method, with application to sensitivity analysis. Journal of Scientific Computing, 68(1):21–41, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0127-z;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0127-z.pdf.

Fjordholm:2016:SPW

[1627] Ulrik S. Fjordholm and Deep Ray. A sign preserving WENO reconstruction method. Journal of Scientific Computing, 68(1):42–63, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0128-y;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0128-y.pdf.

Lanza:2016:CTM

[1628] Alessandro Lanza, Serena Morigi, and Fiorella Sgallari. Constrained TV_{p}–ℓ_{2} model for image restoration. Journal of Scientific Computing, 68(1):64–91, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0129-x;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0129-x.pdf.

Bouchut:2016:WRS

[1629] François Bouchut and Xavier Lhébrard. A 5-wave relaxation solver for the shallow water MHD system. Journal of Scientific Computing, 68 (1):92–115, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0130-4;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0130-4.pdf.

Sun:2016:AAE

[1630] Pengtao Sun and Yuzhou Sun. Asymptotic analysis and error estimates of mixed finite element method for Brinkman model. Journal of Scientific Computing, 68(1):116–142, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0131-3;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0131-
3.pdf.

Antonietti:2016:SAD

[1631] Paola F. Antonietti, Blanca Ayuso de Dios, Ilario Mazzieri, and Alfio Quarteroni. Stability analysis of discontinuous Galerkin approximations to the elastodynamics problem. Journal of Scientific Computing, 68(1):143–170, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0132-2;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0132-2.pdf.

Seal:2016:EHO

[1632] David C. Seal, Qi Tang, Zhengfu Xu, and Andrew J. Christlieb. An explicit high-order single-stage single-step positivity-preserving finite difference WENO method for the compressible Euler equations. Journal of Scientific Computing, 68(1):171–190, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0134-0;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0134-
0.pdf.

Samii:2016:HDG

[1633] Ali Samii, Nishant Panda, Craig Michoski, and Clint Dawson. A hybridized discontinuous Galerkin method for the nonlinear Korteweg–de Vries equation. Journal of Scientific Computing, 68(1):191–212, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0133-1;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0133-1.pdf.

Das:2016:LSP

[1634] Payel Das, Mitali Madhumita Sahani, Gnaneshwar Nelakanti, and Guangqing Long. Legendre spectral projection methods for Fredholm– Hammerstein integral equations. Journal of Scientific Computing, 68 (1):213–230, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0135-z;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0135-z.pdf.

Oberman:2016:AFD

[1635] Adam M. Oberman and Ian Zwiers. Adaptive finite difference methods for nonlinear elliptic and parabolic partial differential equations with free boundaries. Journal of Scientific Computing, 68(1):231–251, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0137-x;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0137-x.pdf.

Yu:2016:NIF

[1636] Bo Yu and Xiaoyun Jiang. Numerical identification of the fractional derivatives in the two-dimensional fractional cable equation. Journal of Scientific Computing, 68(1):252–272, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0136-y;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0136-
y.pdf.

Lee:2016:BDM

[1637] Chang-Ock Lee, Jong Ho Lee, Hyenkyun Woo, and Sangwoon Yun. Block decomposition methods for total variation by primaldual stitching. Journal of Scientific Computing, 68(1):273–302, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0138-9;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0138-9.pdf.

Heuer:2016:SCL

[1638] Norbert Heuer, Salim Meddahi, and Francisco-Javier Sayas. Symmetric coupling of LDG–FEM and DG–BEM. Journal of Scientific Computing, 68(1):303–325, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0139-8;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0139-8.pdf.

Geng:2016:IMT

[1639] Hongrui Geng, Xia Ji, Jiguang Sun, and Liwei Xu. C^{0} IP methods for the transmission eigenvalue problem. Journal of Scientific Computing, 68(1):326–338, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0140-2;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0140-2.pdf.

Zhou:2016:PMS

[1640] Guanyu Zhou, Takahito Kashiwabara, and Issei Oikawa. Penalty method for the stationary Navier–Stokes problems under the slip boundary condition. Journal of Scientific Computing, 68(1):339–374, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0142-0;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0142-0.pdf.

Chen:2016:RBL

[1641] Hongtao Chen, Zhimin Zhang, and Qingsong Zou. A recovery based linear finite element method for 1D bi-harmonic problems. Journal of Scientific Computing, 68(1):375–394, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0141-1;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0141-
1.pdf.

Cavoretto:2016:RAA

[1642] Roberto Cavoretto, Alessandra De Rossi, Emma Perracchione, and Ezio Venturino. Robust approximation algorithms for the detection of attraction basins in dynamical systems. Journal of Scientific Computing, 68(1):395–415, July 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0143-z;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0143-z.pdf.

Hu:2016:ESG

[1643] Lina Hu, Lina Ma, and Jie Shen. Efficient spectral-Galerkin method and analysis for elliptic PDEs with non-local boundary conditions. Journal of Scientific Computing, 68(2):417–437, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0145-x;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0145-x.pdf.

Yang:2016:DTD

[1644] Wei Yang, Yunqing Huang, and Jichun Li. Developing a time-domain finite element method for the Lorentz metamaterial model and applications. Journal of Scientific Computing, 68(2):438–463, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0144-y;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0144-y.pdf.

Cai:2016:PPH

[1645] Xiaofeng Cai, Xiangxiong Zhang, and Jianxian Qiu. Positivity-preserving high order finite volume HWENO schemes for compressible Euler equations. Journal of Scientific Computing, 68(2):464–483, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0147-8;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0147-8.pdf.

Qu:2016:NFI

[1646] Wenzhen Qu, Namdi Brandon, Dangxing Chen, Jingfang Huang, and Tyler Kress. A numerical framework for integrating deferred correction methods to solve high order collocation formulations of ODEs. Journal of Scientific Computing, 68(2):484–520, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0146-9;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0146-9.pdf.

Sheng:2016:CCN

[1647] Zhiqiang Sheng and Guangwei Yuan. A cell-centered nonlinear finite volume scheme preserving fully positivity for diffusion equation. Journal of Scientific Computing, 68(2):521–545, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0148-7;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0148-7.pdf.

Ma:2016:APG

[1648] Shiqian Ma. Alternating proximal gradient method for convex minimization. Journal of Scientific Computing, 68(2):546–572, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-

015-0150-0;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0150-0.pdf.

Hayashi:2016:CAC

[1649] Shunsuke Hayashi, Soon-Yi Wu, and Liping Zhang. Computation algorithm for convex semi-infinite program with second-order cones: Special analyses for affine and quadratic case. Journal of Scientific Computing, 68(2):573–595, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0149-6;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0149-6.pdf.

Winokur:2016:SPS

[1650] J. Winokur, D. Kim, F. Bisetti, O. P. Le Maître, and O. M. Knio. Sparse pseudo spectral projection methods with directional adaptation for uncertainty quantification. Journal of Scientific Computing, 68(2): 596–623, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0153-x;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0153-x.pdf.

Schmidtmann:2016:RBW

[1651] B. Schmidtmann, B. Seibold, and M. Torrilhon. Relations between WENO3 and third-order limiting in finite volume methods. Journal of Scientific Computing, 68(2):624–652, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0151-z;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0151-
z.pdf.

Pandit:2016:EIC

[1652] Swapan K. Pandit and Hemanta Karmakar. An efficient implicit compact streamfunction velocity formulation of two dimensional flows. Journal of Scientific Computing, 68(2):653–688, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0154-9;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0154-9.pdf.

Ding:2016:SML

[1653] Weiyang Ding and Yimin Wei. Solving multi-linear systems with M-tensors. Journal of Scientific Computing, 68(2):689–715, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0156-7;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0156-7.pdf.

Chen:2016:CEE

[1654] Yannan Chen, Liqun Qi, and Qun Wang. Computing extreme eigenvalues of large scale Hankel tensors. Journal of Scientific Computing, 68 (2):716–738, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0155-8;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0155-8.pdf.

vanLith:2016:NSL

[1655] Bart S. van Lith, Jan H. M. ten Thije Boonkkamp, Wilbert L. IJzerman, and Teus W. Tukker. A novel scheme for Liouville’s equation with a discontinuous Hamiltonian and applications to geometrical optics. Journal of Scientific Computing, 68(2):739–771, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0157-6;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0157-6.pdf.

He:2016:SEM

[1656] Ying He, Misun Min, and David P. Nicholls. A spectral element method with transparent boundary condition for periodic layered media scattering. Journal of Scientific Computing, 68(2):772–802, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0158-5;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0158-5.pdf.

Chen:2016:APU

[1657] Jie Chen, Lois C. McInnes, and Hong Zhang. Analysis and practical use of flexible BiCGStab. Journal of Scientific Computing, 68(2): 803–825, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0159-4;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0159-4.pdf.

Wang:2016:HDG

[1658] Shuqin Wang, Jinyun Yuan, Weihua Deng, and Yujiang Wu. A hybridized discontinuous Galerkin method for 2D fractional convectiondiffusion equations. Journal of Scientific Computing, 68(2):826–847, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0160-y;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0160-y.pdf.

Brenner:2016:AFE

[1659] S. C. Brenner, J. Gedicke, and L.-Y. Sung. An adaptive P_{1} finite element method for two-dimensional transverse magnetic time harmonic Maxwell’s equations with general material properties and general boundary conditions. Journal of Scientific Computing, 68(2):848–863, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0161-x;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0161-x.pdf.

Jung:2016:VIC

[1660] Miyoun Jung and Myungjoo Kang. Variational image colorization models using higher-order Mumford–Shah regularizers. Journal of Scientific Computing, 68(2):864–888, August 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0162-9;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0162-
9.pdf.

Shum:2016:CRO

[1661] Alex Shum, Kirsten Morris, and Amir Khajepour. Convergence rate for the ordered upwind method. Journal of Scientific Computing, 68(3): 889–913, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0163-3;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0163-3.pdf.

Christlieb:2016:ESS

[1662] Andrew J. Christlieb, Sigal Gottlieb, Zachary Grant, and David C. Seal. Explicit strong stability preserving multistage two-derivative time-stepping schemes. Journal of Scientific Computing, 68(3):914– 942, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0164-2;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0164-2.pdf. See erratum [1663].

Christlieb:2016:EES

[1663] Andrew J. Christlieb, Sigal Gottlieb, Zachary Grant, and David C. Seal. Erratum to: Explicit Strong Stability Preserving Multistage Two-Derivative Time-Stepping Schemes. Journal of Scientific Computing, 68(3):943–944, September 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0195-8;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0195-

8.pdf. See [1662].

Baccouch:2016:PEA

[1664] Mahboub Baccouch. A posteriori error analysis of the discontinuous Galerkin method for two-dimensional linear hyperbolic conservation laws on Cartesian grids. Journal of Scientific Computing, 68(3):945– 974, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0166-0;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0166-0.pdf.

Boscarino:2016:HOS

[1665] Sebastiano Boscarino, Francis Filbet, and Giovanni Russo. High order semi-implicit schemes for time dependent partial differential equations. Journal of Scientific Computing, 68(3):975–1001, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0168-y;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0168-y.pdf.

Wang:2016:HOF

[1666] Siyang Wang, Kristoffer Virta, and Gunilla Kreiss. High order finite difference methods for the wave equation with non-conforming grid interfaces. Journal of Scientific Computing, 68(3):1002–1028, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0165-1;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0165-1.pdf.

Guo:2016:EHO

[1667] Ruihan Guo, Francis Filbet, and Yan Xu. Efficient high order semiimplicit time discretization and local discontinuous Galerkin methods for highly nonlinear PDEs. Journal of Scientific Computing, 68(3):1029– 1054, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0170-4;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0170-4.pdf.

Xu:2016:JFI

[1668] Wei Xu, Ning Zheng, and Ken Hayami. Jacobian-free implicit inneriteration preconditioner for nonlinear least squares problems. Journal of Scientific Computing, 68(3):1055–1081, September 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0167-z;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0167-
z.pdf.

Lou:2016:PSS

[1669] Yifei Lou, Penghang Yin, and Jack Xin. Point source super-resolution via non-convex L_{1} based methods. Journal of Scientific Computing, 68(3): 1082–1100, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0169-x;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0169-x.pdf.

Briani:2016:NRM

[1670] Maya Briani, Benedetto Piccoli, and Jing-Mei Qiu. Notes on RKDG methods for shallow-water equations in canal networks. Journal of Scientific Computing, 68(3):1101–1123, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0172-2;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0172-
2.pdf.

Sangawi:2016:FCC

[1671] Ali W. K. Sangawi, Ali H. M. Murid, and Lee Khiy Wei. Fast computing of conformal mapping and its inverse of bounded multiply connected regions onto second, third and fourth categories of Koebe’s canonical slit regions. Journal of Scientific Computing, 68(3):1124–1141, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0171-3;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0171-3.pdf. See erratum [1672].

Sangawi:2016:EFC

[1672] Ali W. K. Sangawi, Ali H. M. Murid, and Lee Khiy Wei. Erratum to: Fast Computing of Conformal Mapping and Its Inverse of Bounded Multiply Connected Regions onto Second, Third and Fourth Categories of Koebe’s Canonical Slit Regions. Journal of Scientific Computing, 68(3): 1142–1143, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0192-y;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0192-y.pdf. See [1671].

Wang:2016:DDC

[1673] Chunmei Wang and Junping Wang. Discretization of div-curl systems by weak Galerkin finite element methods on polyhedral partitions. Journal of Scientific Computing, 68(3):1144–1171, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0176-y;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0176-y.pdf.

Chen:2016:ILR

[1674] Li Chen and Ruo Li. An integrated linear reconstruction for finite volume scheme on unstructured grids. Journal of Scientific Computing, 68(3): 1172–1197, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0173-1;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0173-1.pdf. See erratum [1675].

Chen:2016:EIL

[1675] Li Chen and Ruo Li. Erratum to: An Integrated Linear Reconstruction for Finite Volume Scheme on Unstructured Grids. Journal of Scientific Computing, 68(3):1198–1199, September 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0202-0;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0202-

0.pdf. See [1674].

Li:2016:VMN

[1676] Zhi Li, Yifei Lou, and Tieyong Zeng. Variational multiplicative noise removal by DC programming. Journal of Scientific Computing, 68(3): 1200–1216, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0175-z;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0175-z.pdf.

Liu:2016:ESD

[1677] Hailiang Liu and Zhongming Wang. An entropy satisfying discontinuous Galerkin method for nonlinear Fokker–Planck equations. Journal of Scientific Computing, 68(3):1217–1240, September 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0174-0;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0174-
0.pdf.

Zhao:2016:SDE

[1678] Jia Zhao and Qi Wang. Semi-discrete energy-stable schemes for a tensor-based hydrodynamic model of nematic liquid crystal flows. Journal of Scientific Computing, 68(3):1241–1266, September 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0177-x;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0177-
x.pdf.

Barbier:2016:MBD

[1679] S. Barbier, A. Rapaport, and A. Rousseau. Modelling of biological decontamination of a water resource in natural environment and related feedback strategies. Journal of Scientific Computing, 68(3):1267– 1280, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0178-9;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0178-9.pdf.

Brunner:2016:NBN

[1680] Hermann Brunner, Tao Tang, and Jiwei Zhang. Numerical blow-up of nonlinear parabolic integro-differential equations on unbounded domain. Journal of Scientific Computing, 68(3):1281–1298, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0179-8;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0179-8.pdf.

Zhang:2016:ASS

[1681] Jin Zhang and Xiaowei Liu. Analysis of SDFEM on Shishkin triangular meshes and hybrid meshes for problems with characteristic layers. Journal of Scientific Computing, 68(3):1299–1316, September 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-016-0180-2.pdf.

An:2016:OEE

[1682] Rong An. Optimal error estimates of linearized Crank–Nicolson Galerkin method for Landau–Lifshitz equation. Journal of Scientific Computing, 69(1):1–27, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0181-1;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0181-1.pdf.

Chen:2016:EET

[1683] Yanping Chen, Jiaoyan Zeng, and Jie Zhou. L^{p} error estimates of two-grid method for miscible displacement problem. Journal of Scientific Computing, 69(1):28–51, October 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0187-8;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0187-
8.pdf.

Lin:2016:ICA

[1684] Tianyi Lin, Shiqian Ma, and Shuzhong Zhang. Iteration complexity analysis of multi-block ADMM for a family of convex minimization without strong convexity. Journal of Scientific Computing, 69(1):52–81, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0182-0;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0182-0.pdf.

Li:2016:MIS

[1685] Fang Li, Stanley Osher, Jing Qin, and Ming Yan. A multiphase image segmentation based on fuzzy membership functions and L_{1}- norm fidelity. Journal of Scientific Computing, 69(1):82–106, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0183-z;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0183-z.pdf.

Zhou:2016:NSR

[1686] Yijing Zhou and Wei Cai. Numerical solution of the Robin problem of Laplace equations with a Feynman–Kac formula and reflecting Brownian motions. Journal of Scientific Computing, 69(1):107–121, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0184-y;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0184-y.pdf.

Banda:2016:NDC

[1687] Mapundi K. Banda, Axel-Stefan Häck, and Michael Herty. Numerical discretization of coupling conditions by high-order schemes. Journal of Scientific Computing, 69(1):122–145, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0185-x;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0185-
x.pdf.

Thorne:2016:STD

[1688] Jonathan Thorne and Aaron Katz. Source term discretization effects on the steady-state accuracy of finite volume schemes. Journal of Scientific Computing, 69(1):146–169, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0186-9;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0186-
9.pdf.

Baeza:2016:HOW

[1689] A. Baeza, P. Mulet, and D. Zorío. High order weighted extrapolation for boundary conditions for finite difference methods on complex domains with Cartesian meshes. Journal of Scientific Computing, 69(1): 170–200, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0188-7;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0188-7.pdf.

Zheng:2016:FMS

[1690] Bin Zheng, Luoping Chen, Xiaozhe Hu, Long Chen, Ricardo H. Nochetto, and Jinchao Xu. Fast multilevel solvers for a class of discrete fourth order parabolic problems. Journal of Scientific Computing, 69 (1):201–226, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0189-6;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0189-6.pdf.

Conte:2016:MGL

[1691] Dajana Conte and Beatrice Paternoster. Modified Gauss–Laguerre exponential fitting based formulae. Journal of Scientific Computing, 69 (1):227–243, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0190-0;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0190-0.pdf.

Dallmann:2016:SFE

[1692] Helene Dallmann and Daniel Arndt. Stabilized finite element methods for the Oberbeck–Boussinesq model. Journal of Scientific Computing, 69 (1):244–273, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0191-z;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0191-z.pdf.

Do:2016:AMD

[1693] Seongju Do, Youngsoo Ha, M. Kang, and Chang Ho Kim. Application of a multi-dimensional limiting process to central-upwind schemes for solving hyperbolic systems of conservation laws. Journal of Scientific Computing, 69(1):274–291, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0193-x;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0193-
x.pdf.

Hesthaven:2016:UAE

[1694] Jan S. Hesthaven and Shun Zhang. On the use of ANOVA expansions in reduced basis methods for parametric partial differential equations. Journal of Scientific Computing, 69(1):292–313, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0194-9;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0194-9.pdf.

Franck:2016:FVS

[1695] Emmanuel Franck and Laura S. Mendoza. Finite volume scheme with local high order discretization of the hydrostatic equilibrium for the Euler equations with external forces. Journal of Scientific Computing, 69 (1):314–354, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0199-4;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0199-4.pdf.

Chai:2016:MRT

[1696] Zhenhua Chai, Baochang Shi, and Zhaoli Guo. A multiple-relaxation-time lattice Boltzmann model for general nonlinear anisotropic convection–diffusion equations. Journal of Scientific Computing, 69(1): 355–390, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0198-5;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0198-5.pdf.

Yoon:2016:SSS

[1697] Myoungho Yoon, Gangjoon Yoon, and Chohong Min. On solving the singular system arisen from Poisson equation with Neumann boundary condition. Journal of Scientific Computing, 69(1):391–405, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0200-2;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0200-2.pdf.

Gupta:2016:PEA

[1698] Jhuma Sen Gupta, Rajen K. Sinha, G. M. M. Reddy, and Jinank Jain. A posteriori error analysis of two-step backward differentiation formula finite element approximation for parabolic interface problems. Journal of Scientific Computing, 69(1):406–429, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0203-z;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0203-z.pdf.

Rotundo:2016:EAB

[1699] Nella Rotundo, Tae-Yeon Kim, Wen Jiang, Luca Heltai, and Eliot Fried. Error analysis of a B-spline based finite-element method for modeling wind-driven ocean circulation. Journal of Scientific Computing, 69(1): 430–459, October 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0201-1;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0201-1.pdf.

Chen:2016:DDT

[1700] Dai-Qiang Chen. Data-driven tight frame learning scheme based on local and non-local sparsity with application to image recovery. Journal of Scientific Computing, 69(2):461–486, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0205-x;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0205-x.pdf.

Martel:2016:SRB

[1701] Jordan M. Martel and Rodrigo B. Platte. Stability of radial basis function methods for convection problems on the circle and sphere. Journal of Scientific Computing, 69(2):487–505, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0206-9;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0206-9.pdf.

Gao:2016:TADb

[1702] Guang hua Gao and Zhi zhong Sun. Two alternating direction implicit difference schemes for solving the two-dimensional time distributed-order wave equations. Journal of Scientific Computing, 69(2):506–531, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0208-7;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0208-7.pdf.

Duan:2016:CNO

[1703] Yuping Duan, Huibin Chang, and Xue-Cheng Tai. Convergent nonoverlapping domain decomposition methods for variational image segmentation. Journal of Scientific Computing, 69(2):532–555, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0207-8;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0207-8.pdf.

Schaeffer:2016:AMN

[1704] Hayden Schaeffer and Thomas Y. Hou. An accelerated method for nonlinear elliptic PDE. Journal of Scientific Computing, 69(2):556– 580, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0215-8;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0215-8.pdf.

Razi:2016:NSM

[1705] Mani Razi, Peter Attar, and Prakash Vedula. Numerical solution of multidimensional hyperbolic PDEs using defect correction on adaptive grids. Journal of Scientific Computing, 69(2):581–609, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0209-6;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0209-6.pdf.

Lee:2016:REA

[1706] Jeonghun J. Lee. Robust error analysis of coupled mixed methods for Biot’s consolidation model. Journal of Scientific Computing, 69(2): 610–632, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0210-0;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0210-0.pdf.

Albin:2016:AEE

[1707] Nathan Albin and Joshua Klarmann. An algorithmic exploration of the existence of high-order summation by parts operators with diagonal norm. Journal of Scientific Computing, 69(2):633–650, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0211-z;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0211-z.pdf.

Fu:2016:MSF

[1708] Yu Fu, Weidong Zhao, and Tao Zhou. Multistep schemes for forward backward stochastic differential equations with jumps. Journal of Scientific Computing, 69(2):651–672, November 2016. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0212-y;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0212-
y.pdf.

Allaei:2016:JCM

[1709] Sonia Seyed Allaei, Teresa Diogo, and Magda Rebelo. The Jacobi collocation method for a class of nonlinear Volterra integral equations with weakly singular kernel. Journal of Scientific Computing, 69(2): 673–695, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0213-x;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0213-x.pdf.

Li:2016:SEC

[1710] Wanshan Li and Dong Liang. Symmetric energy–conserved S-FDTD scheme for two-dimensional Maxwell’s equations in negative index metamaterials. Journal of Scientific Computing, 69(2):696–735, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0214-9;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0214-9.pdf.

Lu:2016:KIF

[1711] Dong Lu and Yong-Tao Zhang. Krylov integration factor method on sparse grids for high spatial dimension convection–diffusion equations. Journal of Scientific Computing, 69(2):736–763, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0216-7;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0216-7.pdf.

Delchini:2016:VRN

[1712] Marc O. Delchini, Jean C. Ragusa, and Ray A. Berry. Viscous regularization for the non-equilibrium seven-equation two-phase flow model. Journal of Scientific Computing, 69(2):764–804, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0217-6;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0217-6.pdf.

Goetz:2016:NSG

[1713] Claus R. Goetz and Michael Dumbser. A novel solver for the generalized Riemann problem based on a simplified LeFloch–Raviart expansion and a local space–time discontinuous Galerkin formulation. Journal of Scientific Computing, 69(2):805–840, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0218-5;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0218-5.pdf.

Feng:2016:USD

[1714] Xiaobing Feng and Cody Lorton. An unconditionally stable discontinuous Galerkin method for the elastic Helmholtz equations with large frequency. Journal of Scientific Computing, 69(2):841–865, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0219-4;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0219-4.pdf.

Jaust:2016:IMT

[1715] Alexander Jaust, Jochen Schütz, and David C. Seal. Implicit multistage two-derivative discontinuous Galerkin schemes for viscous conservation laws. Journal of Scientific Computing, 69(2):866–891, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0221-x;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0221-x.pdf.

Awanou:2016:SFD

[1716] Gerard Awanou. On standard finite difference discretizations of the elliptic Monge–Ampère equation. Journal of Scientific Computing, 69(2): 892–904, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0220-y;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0220-y.pdf.

Witherden:2016:ASP

[1717] F. D. Witherden, J. S. Park, and P. E. Vincent. An analysis of solution point coordinates for flux reconstruction schemes on tetrahedral elements. Journal of Scientific Computing, 69(2):905–920, November 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0204-y;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0204-y.pdf.

Gong:2016:FDE

[1718] Yuezheng Gong, Xinfeng Liu, and Qi Wang. Fully discretized energy stable schemes for hydrodynamic equations governing two-phase viscous fluid flows. Journal of Scientific Computing, 69(3):921–945, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0224-7;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0224-7.pdf.

Brachet:2016:STS

[1719] Matthieu Brachet and Jean-Paul Chehab. Stabilized times schemes for high accurate finite differences solutions of nonlinear parabolic equations. Journal of Scientific Computing, 69(3):946–982, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0223-8;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0223-8.pdf.

Chen:2016:MEI

[1720] Weitao Chen, Ching-Shan Chou, and Chiu-Yen Kao. Minimizing eigenvalues for inhomogeneous rods and plates. Journal of Scientific Computing, 69(3):983–1013, December 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0222-9;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0222-
9.pdf.

Hou:2016:TSL

[1721] Likun Hou, Hao Gao, and Xiaoqun Zhang. A two-stage low rank approach for calibrationless dynamic parallel magnetic resonance image reconstruction. Journal of Scientific Computing, 69(3):1014–1032, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0225-6;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0225-6.pdf.

Zaghi:2016:AWP

[1722] Stefano Zaghi, Andrea Di Mascio, and Bernardo Favini. Application of WENO–positivity-preserving schemes to highly under-expanded jets. Journal of Scientific Computing, 69(3):1033–1057, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0226-5;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0226-5.pdf.

Djoko:2016:NSS

[1723] J. K. Djoko, J. M. Lubuma, and M. Mbehou. On the numerical solution of the stationary power-law Stokes equations: a penalty finite element approach. Journal of Scientific Computing, 69(3):1058–1082, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0227-4;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0227-4.pdf.

Cheng:2016:SOW

[1724] Kelong Cheng, Cheng Wang, Steven M. Wise, and Xingye Yue. A second-order, weakly energy-stable pseudo-spectral scheme for the Cahn– Hilliard equation and its solution by the homogeneous linear iteration method. Journal of Scientific Computing, 69(3):1083–1114, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0228-3;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0228-3.pdf.

Boyd:2016:TMS

[1725] John P. Boyd. Tracing multiple solution branches for nonlinear ordinary differential equations: Chebyshev and Fourier spectral methods and a degree-increasing spectral homotopy [DISH]. Journal of Scientific Computing, 69(3):1115–1143, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0229-2;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0229-
2.pdf.

Liao:2016:WAS

[1726] Hong lin Liao, Ying Zhao, and Xing hu Teng. A weighted ADI scheme for subdiffusion equations. Journal of Scientific Computing, 69(3):1144– 1164, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0230-9;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0230-9.pdf.

Tcheng:2016:LCA

[1727] Alexandra Tcheng and Jean-Christophe Nave. A low complexity algorithm for non-monotonically evolving fronts. Journal of Scientific Computing, 69(3):1165–1191, December 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0231-8;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0231-
8.pdf.

Gatica:2016:PPE

[1728] Gabriel N. Gatica and Filánder A. Sequeira. A priori and a posteriori error analyses of an augmented HDG method for a class of quasi-Newtonian Stokes flows. Journal of Scientific Computing, 69(3):1192–1250, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0233-6;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0233-6.pdf.

Huang:2016:SDG

[1729] Xuehai Huang and Jianguo Huang. A superconvergent C^{0} discontinuous Galerkin method for Kirchhoff plates: Error estimates, hybridization and postprocessing. Journal of Scientific Computing, 69(3):1251– 1278, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0232-7;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0232-7.pdf.

Han:2016:AFE

[1730] Jiayu Han and Yidu Yang. An adaptive finite element method for the transmission eigenvalue problem. Journal of Scientific Computing, 69(3): 1279–1300, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0234-5;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0234-5.pdf.

Sun:2016:LLC

[1731] Tao Sun, Roberto Barrio, Hao Jiang, and Lizhi Cheng. Local linear convergence of a primal–dual algorithm for the augmented convex models. Journal of Scientific Computing, 69(3):1301–1315, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0235-4;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0235-4.pdf.

Zhang:2016:HOL

[1732] Yujie Zhang and Wenjing Ye. A high-order level-set method with enhanced stability for curvature driven flows and surface diffusion motion. Journal of Scientific Computing, 69(3):1316–1345, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0236-3;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0236-3.pdf.

Zhu:2016:ARM

[1733] Hongqiang Zhu, Jianxian Qiu, and Jing-Mei Qiu. An h-adaptive RKDG method for the Vlasov–Poisson system. Journal of Scientific Computing, 69(3):1346–1365, December 2016. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0238-1;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0238-
1.pdf.

Hu:2016:CAE

[1734] Jun Hu, Xueqin Yang, and Shuo Zhang. Capacity of the Adini element for biharmonic equations. Journal of Scientific Computing, 69(3): 1366–1383, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0237-2;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0237-2.pdf.

Qiu:2016:HOH

[1735] Weifeng Qiu, Manuel Solano, and Patrick Vega. A high order HDG method for curved-interface problems via approximations from straight triangulations. Journal of Scientific Computing, 69(3):1384–1407, December 2016. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0239-0;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0239-0.pdf.

Garcke:2017:SFC

[1736] Jochen Garcke and Axel Kröner. Suboptimal feedback control of PDEs by solving HJB equations on adaptive sparse grids. Journal of Scientific Computing, 70(1):1–28, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0240-7;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0240-
7.pdf.

Haddar:2017:RIM

[1737] Houssem Haddar, Zixian Jiang, and Mohamed Kamel Riahi. A robust inversion method for quantitative 3D shape reconstruction from coaxial eddy current measurements. Journal of Scientific Computing, 70 (1):29–59, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0241-6;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0241-6.pdf.

Li:2017:OQF

[1738] Hengguang Li and Qinghui Zhang. Optimal quadrilateral finite elements on polygonal domains. Journal of Scientific Computing, 70(1): 60–84, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0242-5;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0242-5.pdf.

Shi:2017:USAa

[1739] Dongyang Shi, Junjun Wang, and Fengna Yan. Unconditional superconvergence analysis for nonlinear parabolic equation with EQ_{1}^{rot} nonconforming finite element. Journal of Scientific Computing, 70(1):

85–111, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0243-4;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0243-4.pdf.

Zou:2017:USQ

[1740] Qingsong Zou. An unconditionally stable quadratic finite volume scheme over triangular meshes for elliptic equations. Journal of Scientific Computing, 70(1):112–124, January 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0244-3;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0244-
3.pdf.

Guo:2017:STG

[1741] Hailong Guo, Zhimin Zhang, and Ren Zhao. Superconvergent twogrid methods for elliptic eigenvalue problems. Journal of Scientific Computing, 70(1):125–148, January 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0245-2;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0245-
2.pdf.

Tang:2017:LPF

[1742] Qili Tang and Yunqing Huang. Local and parallel finite element algorithm based on Oseen-type iteration for the stationary incompressible MHD flow. Journal of Scientific Computing, 70(1):149–174, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0246-1;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0246-1.pdf.

Wang:2017:HSA

[1743] Jiangxing Wang, Chuanmiao Chen, and Ziqing Xie. The highest superconvergence analysis of ADG method for two point boundary values problem. Journal of Scientific Computing, 70(1):175–191, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0247-0;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0247-0.pdf.

Wang:2017:GMS

[1744] Yujie Wang, Fuming Ma, and Enxi Zheng. Galerkin method for the scattering problem of a slit. Journal of Scientific Computing, 70(1): 192–209, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0248-z;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0248-z.pdf.

Brus:2017:PSI

[1745] S. R. Brus, D. Wirasaet, J. J. Westerink, and C. Dawson. Performance and scalability improvements for discontinuous Galerkin solutions to conservation laws on unstructured grids. Journal of Scientific Computing, 70 (1):210–242, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0249-y;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0249-y.pdf.

Chang:2017:LSO

[1746] J. Chang, S. Karra, and K. B. Nakshatrala. Large-scale optimizationbased non-negative computational framework for diffusion equations: Parallel implementation and performance studies. Journal of Scientific Computing, 70(1):243–271, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0250-5;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0250-
5.pdf.

Tang:2017:UCT

[1747] Min Tang and Yihong Wang. Uniform convergent tailored finite point method for advection–diffusion equation with discontinuous, anisotropic and vanishing diffusivity. Journal of Scientific Computing, 70(1): 272–300, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0254-1;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0254-1.pdf.

Li:2017:SOS

[1748] Dong Li and Zhonghua Qiao. On second order semi-implicit Fourier spectral methods for 2D Cahn–Hilliard equations. Journal of Scientific Computing, 70(1):301–341, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0251-4;

 http://link.springer.com/content/pdf/10.1007/s10915-016-0251-
4.pdf.

Ma:2017:NFE

[1749] Jingtang Ma. A new finite element analysis for inhomogeneous boundary-value problems of space fractional differential equations. Journal of Scientific Computing, 70(1):342–354, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-015-0082-8;
 http://link.springer.com/content/pdf/10.1007/s10915-015-0082-
8.pdf.

Baffet:2017:HOAa

[1750] Daniel Baffet and Jan S. Hesthaven. High-order accurate local schemes for fractional differential equations. Journal of Scientific Computing, 70 (1):355–385, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0089-1;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0089-1.pdf.

Zhao:2017:GFE

[1751] Zhengang Zhao, Yunying Zheng, and Peng Guo. A Galerkin finite element method for a class of time–space fractional differential equation with nonsmooth data. Journal of Scientific Computing, 70(1): 386–406, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-015-0107-3;   http://link.springer.com/content/
pdf/10.1007/s10915-015-0107-3.pdf.

Zhao:2017:TMF

[1752] Yanmin Zhao, Pan Chen, Weiping Bu, Xiangtao Liu, and Yifa Tang. Two mixed finite element methods for time-fractional diffusion equations. Journal of Scientific Computing, 70(1):407–428, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
015-0152-y;   http://link.springer.com/content/pdf/10.1007/
s10915-015-0152-y.pdf.

Wang:2017:AFE

[1753] Hong Wang, Danping Yang, and Shengfeng Zhu. Accuracy of finite element methods for boundary-value problems of steady-state fractional diffusion equations. Journal of Scientific Computing, 70(1):429–449, January 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0196-7;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0196-7.pdf.

Zhang:2017:OSS

[1754] Jing Zhang, Li-Lian Wang, Huiyuan Li, and Zhimin Zhang. Optimal spectral schemes based on generalized prolate spheroidal wave functions of order −1. Journal of Scientific Computing, 70(2):451–477, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0253-2;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0253-2.pdf.

Zhang:2017:SML

[1755] Junyu Zhang, Zaiwen Wen, and Yin Zhang. Subspace methods with local refinements for eigenvalue computation using low-rank tensortrain format. Journal of Scientific Computing, 70(2):478–499, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0255-0;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0255-0.pdf.

Li:2017:EEM

[1756] Xiaocui Li, Xiaoyuan Yang, and Yinghan Zhang. Error estimates of mixed finite element methods for time-fractional Navier–Stokes equations. Journal of Scientific Computing, 70(2):500–515, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0252-3;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0252-3.pdf.

Michoski:2017:SNC

[1757] C. Michoski, A. Alexanderian, C. Paillet, E. J. Kubatko, and C. Dawson. Stability of nonlinear convection–diffusion–reaction systems in discontinuous Galerkin methods. Journal of Scientific Computing, 70(2): 516–550, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0256-z;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0256-z.pdf.

Guo:2017:RBF

[1758] Jingyang Guo and Jae-Hun Jung. Radial basis function ENO and WENO finite difference methods based on the optimization of shape parameters. Journal of Scientific Computing, 70(2):551–575, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0257-y;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0257-y.pdf.

Li:2017:SAI

[1759] Tingting Li, Chi-Wang Shu, and Mengping Zhang. Stability analysis of the inverse Lax–Wendroff boundary treatment for high order central difference schemes for diffusion equations. Journal of Scientific Computing, 70(2):576–607, February 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0258-x;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0258-
x.pdf.

Antonietti:2017:UAS

[1760] Paola F. Antonietti, Marco Sarti, Marco Verani, and Ludmil T. Zikatanov. A uniform additive Schwarz preconditioner for high-order discontinuous Galerkin approximations of elliptic problems. Journal of Scientific Computing, 70(2):608–630, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0259-9;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0259-
9.pdf.

Liu:2017:GCF

[1761] Zheng Liu, Huayan Zhang, and Chunlin Wu. On geodesic curvature flow with level set formulation over triangulated surfaces. Journal of Scientific Computing, 70(2):631–661, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0260-3;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0260-
3.pdf.

Song:2017:CSO

[1762] Lunji Song and Chaoxia Yang. Convergence of a second-order linearized BDF–IPDG for nonlinear parabolic equations with discontinuous coefficients. Journal of Scientific Computing, 70(2):662–685, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0261-2;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0261-2.pdf.

Wang:2017:FDM

[1763] Cheng Wang and Pengtao Sun. A fictitious domain method with distributed Lagrange multiplier for parabolic problems with moving interfaces. Journal of Scientific Computing, 70(2):686–716, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0262-1;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0262-1.pdf.

Yang:2017:TLS

[1764] Haijian Yang and Xiao-Chuan Cai. Two-level space–time domain decomposition methods for flow control problems. Journal of Scientific Computing, 70(2):717–743, February 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0263-0;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0263-
0.pdf.

Huang:2017:DDG

[1765] Hongying Huang, Zheng Chen, Jin Li, and Jue Yan. Direct discontinuous Galerkin method and its variations for second order elliptic equations. Journal of Scientific Computing, 70(2):744–765, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0264-z;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0264-z.pdf.

Chung:2017:ASM

[1766] Eric T. Chung, Jie Du, and Man Chun Yuen. An adaptive SDG method for the Stokes system. Journal of Scientific Computing, 70(2): 766–792, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0265-y;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0265-y.pdf.

Zhang:2017:NCS

[1767] Chao Zhang, Wenjie Liu, and Li-Lian Wang. A new collocation scheme using non-polynomial basis functions. Journal of Scientific Computing, 70(2):793–818, February 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0269-7;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0269-
7.pdf.

Fernandez-Cara:2017:NCT

[1768] Enrique Fernández-Cara, Arnaud Münch, and Diego A. Souza. On the numerical controllability of the two-dimensional heat, Stokes and Navier– Stokes equations. Journal of Scientific Computing, 70(2):819–858, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0266-x;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0266-x.pdf.

Bonaventura:2017:USS

[1769] L. Bonaventura and A. Della Rocca. Unconditionally strong stability preserving extensions of the TR–BDF2 method. Journal of Scientific Computing, 70(2):859–895, February 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0267-9;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0267-
9.pdf.

Causley:2017:MLT

[1770] Matthew Causley, Andrew Christlieb, and Eric Wolf. Method of lines transpose: an efficient unconditionally stable solver for wave propagation. Journal of Scientific Computing, 70(2):896–921, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0268-8;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0268-8.pdf.

Lu:2017:DGA

[1771] Zhongjie Lu, A. Cesmelioglu, J. J. W. Van der Vegt, and Yan Xu. Discontinuous Galerkin approximations for computing electromagnetic Bloch modes in photonic crystals. Journal of Scientific Computing, 70(2): 922–964, February 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0270-1;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0270-1.pdf.

Han:2017:NAS

[1772] Daozhi Han, Alex Brylev, Xiaofeng Yang, and Zhijun Tan. Numerical analysis of second order, fully discrete energy stable schemes for phase field models of two-phase incompressible flows. Journal of Scientific Computing, 70(3):965–989, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0279-5;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0279-
5.pdf.

Zwaan:2017:MSE

[1773] Ian N. Zwaan and Michiel E. Hochstenbach. Multidirectional subspace expansion for one-parameter and multiparameter Tikhonov regularization. Journal of Scientific Computing, 70(3):990–1009, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0271-0;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0271-0.pdf.

Chen:2017:SHM

[1774] Huangxin Chen, Weifeng Qiu, Ke Shi, and Manuel Solano. A superconvergent HDG method for the Maxwell equations. Journal of Scientific Computing, 70(3):1010–1029, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0272-z;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0272-
z.pdf.

Chen:2017:DSB

[1775] Lizhen Chen, Jing An, and Qingqu Zhuang. Direct solvers for the biharmonic eigenvalue problems using Legendre polynomials. Journal of Scientific Computing, 70(3):1030–1041, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0277-7;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0277-
7.pdf.

Beyn:2017:SCS

[1776] Wolf-Jürgen Beyn, Elena Isaak, and Raphael Kruse. Stochastic C-stability and B-consistency of explicit and implicit Milstein-type schemes. Journal of Scientific Computing, 70(3):1042–1077, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0290-x;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0290-x.pdf.

Ivorra:2017:NAD

[1777] Benjamin Ivorra, Susana Gomez, Roland Glowinski, and Angel Manuel Ramos. Nonlinear advection–diffusion–reaction phenomena involved in the evolution and pumping of oil in open sea: Modeling, numerical simulation and validation considering the Prestige and Oleg Naydenov oil spill cases. Journal of Scientific Computing, 70(3):1078–1104, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0274-x;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0274-x.pdf.

Bras:2017:AIE

[1778] Michal Braś, Giuseppe Izzo, and Zdzislaw Jackiewicz. Accurate implicit–explicit general linear methods with inherent Runge–Kutta stability. Journal of Scientific Computing, 70(3):1105–1143, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0273-y;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0273-y.pdf.

Su:2017:TLP

[1779] Haiyan Su, Xinlong Feng, and Jianping Zhao. Two-level penalty Newton iterative method for the 2D/3D stationary incompressible magnetohydrodynamics equations. Journal of Scientific Computing, 70(3):1144–1179, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0276-8;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0276-8.pdf.

Pan:2017:ECM

[1780] Kejia Pan, Dongdong He, and Hongling Hu. An extrapolation cascadic multigrid method combined with a fourth-order compact scheme for 3D Poisson equation. Journal of Scientific Computing, 70(3):1180–1203, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0275-9;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0275-9.pdf.

Li:2017:TSF

[1781] Qia Li, Yuesheng Xu, and Na Zhang. Two-step fixed-point proximity algorithms for multi-block separable convex problems. Journal of Scientific Computing, 70(3):1204–1228, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0278-6;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0278-
6.pdf.

Jung:2017:PSI

[1782] Miyoun Jung. Piecewise-smooth image segmentation models with L^{1} data-fidelity terms. Journal of Scientific Computing, 70(3):1229–1261, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0280-z;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0280-z.pdf.

Sonntag:2017:EPS

[1783] Matthias Sonntag and Claus-Dieter Munz. Efficient parallelization of a shock capturing for discontinuous Galerkin methods using finite volume sub-cells. Journal of Scientific Computing, 70(3):1262–1289, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0287-5;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0287-5.pdf.

Sanchez-Vizuet:2017:SBF

[1784] Tonatiuh Sánchez-Vizuet and Francisco-Javier Sayas. Symmetric boundary-finite element discretization of time dependent acoustic scattering by elastic obstacles with piezoelectric behavior. Journal of Scientific Computing, 70(3):1290–1315, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0281-y;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0281-
y.pdf.

Zhang:2017:SHM

[1785] Xuping Zhang, Jintao Zhang, and Bo Yu. Symmetric homotopy method for discretized elliptic equations with cubic and quintic nonlinearities. Journal of Scientific Computing, 70(3):1316–1335, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0284-8;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0284-8.pdf.

Ma:2017:LRP

[1786] Liyan Ma, Li Xu, and Tieyong Zeng. Low rank prior and total variation regularization for image deblurring. Journal of Scientific Computing, 70 (3):1336–1357, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/article/10.
1007/s10915-016-0282-x;   http://link.springer.com/content/
pdf/10.1007/s10915-016-0282-x.pdf.

Guermond:2017:ECM

[1787] Jean-Luc Guermond, Bojan Popov, and Yong Yang. The effect of the consistent mass matrix on the maximum-principle for scalar conservation equations. Journal of Scientific Computing, 70(3):1358–1366, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0285-7;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0285-7.pdf.

Zhao:2017:DES

[1788] Jia Zhao, Huiyuan Li, Qi Wang, and Xiaofeng Yang. Decoupled energy stable schemes for a phase field model of three-phase incompressible viscous fluid flow. Journal of Scientific Computing, 70(3):1367–1389, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/article/10.1007/s10915-
016-0283-9;   http://link.springer.com/content/pdf/10.1007/
s10915-016-0283-9.pdf.

Kaiser:2017:NSS

[1789] Klaus Kaiser, Jochen Schütz, Ruth Schöbel, and Sebastian Noelle. A new stable splitting for the isentropic Euler equations. Journal of Scientific Computing, 70(3):1390–1407, March 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

http://link.springer.com/article/10.1007/s10915-016-0286-6;
 http://link.springer.com/content/pdf/10.1007/s10915-016-0286-
6.pdf.

Ma:2017:CBV

[1790] Junjie Ma and Shuhuang Xiang. A collocation boundary value method for linear Volterra integral equations. Journal of Scientific Computing, 71(1):1–20, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0289-3.pdf.

Heister:2017:DUS

[1791] Timo Heister, Muhammad Mohebujjaman, and Leo G. Rebholz. Decoupled, unconditionally stable, higher order discretizations for MHD flow simulation. Journal of Scientific Computing, 71(1):21–43, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0288-4.pdf.

Moe:2017:PPD

[1792] Scott A. Moe, James A. Rossmanith, and David C. Seal. Positivitypreserving discontinuous Galerkin methods with Lax–Wendroff time discretizations. Journal of Scientific Computing, 71(1):44–70, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0291-9.pdf.

Pichard:2017:ACA

[1793] T. Pichard, G. W. Alldredge, S. Brull, B. Dubroca, and M. Frank. An approximation of the M_{2} closure: Application to radiotherapy dose simulation. Journal of Scientific Computing, 71(1):71–108, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0292-8.pdf.

Li:2017:FOS

[1794] Qin Li, Qilong Guo, Dong Sun, Pengxin Liu, and Hanxin Zhang. A fourth-order symmetric WENO scheme with improved performance by new linear and nonlinear optimizations. Journal of Scientific Computing, 71(1):109–143, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0293-7.pdf.

Kang:2017:SCL

[1795] Wei Kang and Lucas C. Wilcox. Solving 1D conservation laws using Pontryagin’s minimum principle. Journal of Scientific Computing, 71 (1):144–165, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0294-6.pdf.

Perrotta:2017:SOF

[1796] Andrea Perrotta and Bernardo Favini. A second-order finite-volume scheme for Euler equations: Kinetic energy preserving and staggering effects. Journal of Scientific Computing, 71(1):166–194, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0295-5.pdf.

Song:2017:WGM

[1797] Lunji Song, Kaifang Liu, and Shan Zhao. A weak Galerkin method with an over–relaxed stabilization for low regularity elliptic problems. Journal of Scientific Computing, 71(1):195–218, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0296-4.pdf.

Wang:2017:CSP

[1798] Siyang Wang and Gunilla Kreiss. Convergence of summation-by-parts finite difference methods for the wave equation. Journal of Scientific Computing, 71(1):219–245, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

content/pdf/10.1007/s10915-016-0297-3.pdf. See response [2003].

Zorio:2017:ALW

[1799] D. Zorío, A. Baeza, and P. Mulet. An approximate Lax–Wendrofftype procedure for high order accurate schemes for hyperbolic conservation laws. Journal of Scientific Computing, 71(1):246–273, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0298-2.pdf.

Li:2017:LGP

[1800] Lei Li, Xiaoqian Xu, and Saverio E. Spagnolie. A locally gradientpreserving reinitialization for level set functions. Journal of Scientific Computing, 71(1):274–302, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/content/pdf/10.1007/s10915-016-0299-1.pdf.

Khan:2017:EAS

[1801] Arbaz Khan and Akhlaq Husain. Exponentially accurate spectral element method for fourth order elliptic problems. Journal of Scientific Computing, 71(1):303–328, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/content/pdf/10.1007/s10915-016-0300-z.pdf.

Bernardi:2017:CAT

[1802] Christine Bernardi, Jad Dakroub, Gihane Mansour, Farah Rafei, and Toni Sayah. Convergence analysis of two numerical schemes applied to a nonlinear elliptic problem. Journal of Scientific Computing, 71 (1):329–347, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0301-y.pdf.

Anaya:2017:MMS

[1803] Verónica Anaya, David Mora, Carlos Reales, and Ricardo Ruiz-Baier. Mixed methods for a stream-function-vorticity formulation of the axisymmetric Brinkman equations. Journal of Scientific Computing, 71 (1):348–364, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0302-x.pdf.

Nordstrom:2017:RWP

[1804] Jan Nordström. A roadmap to well posed and stable problems in computational physics. Journal of Scientific Computing, 71(1):365– 385, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0303-9.pdf.

Badia:2017:CSW

[1805] Santiago Badia and Juan Vicente Gutiérrez-Santacreu. Convergence to suitable weak solutions for a finite element approximation of the Navier– Stokes equations with numerical subgrid scale modeling. Journal of Scientific Computing, 71(1):386–413, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0304-8.pdf.

Qiu:2017:HOM

[1806] Jing-Mei Qiu and Giovanni Russo. A high order multi-dimensional characteristic tracing strategy for the Vlasov–Poisson system. Journal of Scientific Computing, 71(1):414–434, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0305-7.pdf.

Robinson:2017:FAA

[1807] Daniel P. Robinson and Rachael Tappenden. A flexible ADMM algorithm for big data applications. Journal of Scientific Computing, 71 (1):435–467, April 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0306-6.pdf.

Sun:2017:IBI

[1808] Yao Sun. Indirect boundary integral equation method for the Cauchy problem of the Laplace equation. Journal of Scientific Computing, 71 (2):469–498, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0308-4.pdf.

Wu:2017:VCL

[1809] Jiming Wu. Vertex-centered linearity-preserving schemes for nonlinear parabolic problems on polygonal grids. Journal of Scientific Computing, 71(2):499–524, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0309-3.pdf.

Zhang:2017:DGB

[1810] Futao Zhang, Yan Xu, and Falai Chen. Discontinuous Galerkin based isogeometric analysis for geometric flows and applications in geometric modeling. Journal of Scientific Computing, 71(2):525–546, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0307-5.pdf.

Karoui:2017:WFF

[1811] Abderrazek Karoui and Ahmed Souabni. Weighted finite Fourier transform operator: Uniform approximations of the eigenfunctions, eigenvalues decay and behaviour. Journal of Scientific Computing, 71(2): 547–570, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0310-x.pdf.

Cano:2017:TMN

[1812] Alfredo Cano and Carlos Moreno. Transformation methods for the numerical integration of three-dimensional singular functions. Journal of Scientific Computing, 71(2):571–593, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0311-9.pdf.

Guo:2017:PPR

[1813] Hailong Guo and Xu Yang. Polynomial preserving recovery for high frequency wave propagation. Journal of Scientific Computing, 71(2): 594–614, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0312-8.pdf.

Guo:2017:LDG

[1814] Hui Guo, Fan Yu, and Yang Yang. Local discontinuous Galerkin method for incompressible miscible displacement problem in porous media. Journal of Scientific Computing, 71(2):615–633, May 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/content/pdf/10.1007/s10915-016-0313-7.pdf.

Li:2017:NAS

[1815] Zhaoxiang Li, Zhi-Qiang Wang, and Jianxin Zhou. A new augmented singular transform and its partial Newton-correction method for finding more solutions. Journal of Scientific Computing, 71(2):634–659, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0314-6.pdf.

Jia:2017:MTG

[1816] Lueling Jia, Huanzhen Chen, and Hong Wang. Mixed-type Galerkin variational principle and numerical simulation for a generalized nonlocal elastic model. Journal of Scientific Computing, 71(2):660–681, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0316-4.pdf.

Huang:2017:CSV

[1817] Rong Huang and Delin Chu. Computing singular value decompositions of parameterized matrices with total nonpositivity to high relative accuracy. Journal of Scientific Computing, 71(2):682–711, May 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/content/pdf/10.1007/s10915-016-0315-5.pdf.

Deng:2017:PMB

[1818] Wei Deng, Ming-Jun Lai, Zhimin Peng, and Wotao Yin. Parallel multiblock ADMM with o(1/k) convergence. Journal of Scientific Computing, 71(2):712–736, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0318-2.pdf.

Gelb:2017:DEN

[1819] Anne Gelb and Guohui Song. Detecting edges from non-uniform Fourier data using Fourier frames. Journal of Scientific Computing, 71(2):737– 758, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0320-8.pdf.

Ding:2017:HON

[1820] Hengfei Ding and Changpin Li. High-order numerical algorithms for Riesz derivatives via constructing new generating functions. Journal of Scientific Computing, 71(2):759–784, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0317-3.pdf.

Li:2017:HON

[1821] Zhiqiang Li, Zongqi Liang, and Yubin Yan. High-order numerical methods for solving time fractional partial differential equations. Journal of Scientific Computing, 71(2):785–803, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0319-1.pdf.

Alici:2017:HAP

[1822] H. Alici and J. Shen. Highly accurate pseudospectral approximations of the prolate spheroidal wave equation for any bandwidth parameter and zonal wavenumber. Journal of Scientific Computing, 71(2):804– 821, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0321-7.pdf.

Chen:2017:ETS

[1823] Rongsan Chen, Min Zou, and Li Xiao. Entropy-TVD scheme for the shallow water equations in one dimension. Journal of Scientific Computing, 71(2):822–838, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0322-6.pdf.

Chizari:2017:GIL

[1824] Hossain Chizari and Farzad Ismail. A grid-insensitive LDA method on triangular grids solving the system of Euler equations. Journal of Scientific Computing, 71(2):839–874, May 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0323-5.pdf.

Lee:2017:ENO

[1825] B. Lee, M. Kang, and S. Kim. An essentially non-oscillatory Crank– Nicolson procedure for the simulation of convection-dominated flows. Journal of Scientific Computing, 71(2):875–895, May 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/content/pdf/10.1007/s10915-016-0324-4.pdf.

Wang:2017:WGF

[1826] Chunmei Wang and Haomin Zhou. A weak Galerkin finite element method for a type of fourth order problem arising from fluorescence tomography. Journal of Scientific Computing, 71(3):897–918, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0325-3.pdf.

May:2017:EIS

[1827] Sandra May and Marsha Berger. An explicit implicit scheme for cut cells in embedded boundary meshes. Journal of Scientific Computing, 71(3):919–943, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0326-2.pdf.

Ketcheson:2017:DOS

[1828] David I. Ketcheson, Lajos Lóczi, Aliya Jangabylova, and Adil Kusmanov. Dense output for strong stability preserving Runge–Kutta methods. Journal of Scientific Computing, 71(3):944–958, June 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/content/pdf/10.1007/s10915-016-0331-5.pdf.

Cheng:2017:APM

[1829] Yingda Cheng, Andrew J. Christlieb, Wei Guo, and Benjamin Ong. An asymptotic preserving Maxwell solver resulting in the Darwin limit of electrodynamics. Journal of Scientific Computing, 71(3):959–993, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0328-0.pdf.

Li:2017:PPW

[1830] Maojun Li, Philippe Guyenne, Fengyan Li, and Liwei Xu. A positivitypreserving well-balanced central discontinuous Galerkin method for the nonlinear shallow water equations. Journal of Scientific Computing, 71 (3):994–1034, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0329-z.pdf.

Liu:2017:CME

[1831] Xiao-Yan Liu, C. S. Chen, and Andreas Karageorghis. Conformal mapping for the efficient solution of Poisson problems with the Kansa– RBF method. Journal of Scientific Computing, 71(3):1035–1061, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0330-6.pdf.

Chandrashekar:2017:WBN

[1832] Praveen Chandrashekar and Markus Zenk. Well-balanced nodal discontinuous Galerkin method for Euler equations with gravity. Journal of Scientific Computing, 71(3):1062–1093, June 2017. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0339-x.pdf.

Bao:2017:NMC

[1833] Weizhu Bao, Yongyong Cai, Xiaowei Jia, and Qinglin Tang. Numerical methods and comparison for the Dirac equation in the nonrelativistic limit regime. Journal of Scientific Computing, 71(3):1094– 1134, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0333-3.pdf.

Ortleb:2017:KEP

[1834] Sigrun Ortleb. A kinetic energy preserving DG scheme based on Gauss– Legendre points. Journal of Scientific Computing, 71(3):1135–1168, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0334-2.pdf.

Carraro:2017:GOE

[1835] Thomas Carraro and Christian Goll. A goal-oriented error estimator for a class of homogenization problems. Journal of Scientific Computing, 71 (3):1169–1196, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0338-y.pdf.

Chowdhury:2017:EFG

[1836] Habibullah Amin Chowdhury, Adam Wittek, Karol Miller, and Grand Roman Joldes. An element free Galerkin method based on the modified moving least squares approximation. Journal of Scientific Computing, 71(3):1197–1211, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/content/pdf/10.1007/s10915-016-0337-z.pdf.

Ong:2017:NFR

[1837] Benjamin W. Ong, Andrew J. Christlieb, and Bryan D. Quaife. A new family of regularized kernels for the harmonic oscillator. Journal of Scientific Computing, 71(3):1212–1237, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0336-0.pdf.

Wang:2017:BEM

[1838] Qiao Wang, Wei Zhou, Yonggang Cheng, Gang Ma, Xiaolin Chang, and Qiang Huang. The boundary element method with a fast multipole accelerated integration technique for 3D elastostatic problems with arbitrary body forces. Journal of Scientific Computing, 71(3):1238– 1264, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0335-1.pdf.

Amat:2017:NWS

[1839] Sergio Amat and Juan Ruiz. New WENO smoothness indicators computationally efficient in the presence of corner discontinuities. Journal of Scientific Computing, 71(3):1265–1302, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0342-2.pdf.

Lee:2017:SOO

[1840] Hyun Geun Lee, Jaemin Shin, and June-Yub Lee. A second-order operator splitting Fourier spectral method for models of epitaxial thin film growth. Journal of Scientific Computing, 71(3):1303–1318, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0340-4.pdf.

Cho:2017:EAC

[1841] Min Hyung Cho and Wei Cai. Efficient and accurate computation of electric field dyadic Green’s function in layered media. Journal of Scientific Computing, 71(3):1319–1350, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/content/pdf/10.1007/s10915-016-0341-3.pdf.

Moghaddam:2017:EAA

[1842] Behrouz Parsa Moghaddam and José António Tenreiro Machado. Extended algorithms for approximating variable order fractional derivatives with applications. Journal of Scientific Computing, 71(3):1351– 1374, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0343-1.pdf.

Costa:2017:VHO

[1843] Ricardo Costa, Stéphane Clain, Gaspar J. Machado, and Raphaël Loubère. A very high-order accurate staggered finite volume scheme for the stationary incompressible Navier–Stokes and Euler equations on unstructured meshes. Journal of Scientific Computing, 71(3):1375– 1411, June 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-016-0348-9.pdf.

Pang:2017:EBC

[1844] Gang Pang, Yibo Yang, and Shaoqiang Tang. Exact boundary condition for semi-discretized Schrödinger equation and heat equation in a rectangular domain. Journal of Scientific Computing, 72(1):1–13, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0344-0.pdf.

Lakoba:2017:LTS

[1845] T. I. Lakoba. Long-time simulations of nonlinear Schrödinger-type equations using step size exceeding threshold of numerical instability. Journal of Scientific Computing, 72(1):14–48, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0346-y.pdf.

Gatto:2017:EPH

[1846] P. Gatto and J. S. Hesthaven. Efficient preconditioning of hp-FEM matrices by hierarchical low-rank approximations. Journal of Scientific Computing, 72(1):49–80, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/content/pdf/10.1007/s10915-016-0347-x.pdf.

Stiller:2017:NWS

[1847] Jörg Stiller. Nonuniformly weighted Schwarz smoothers for spectral element multigrid. Journal of Scientific Computing, 72(1):81–96, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-016-0345-z.pdf.

Hakula:2017:PEU

[1848] Harri Hakula, Michael Neilan, and Jeffrey S. Ovall. A posteriori estimates using auxiliary subspace techniques. Journal of Scientific Computing, 72(1):97–127, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0352-0.pdf.

Chalmers:2017:SMS

[1849] N. Chalmers and L. Krivodonova. Spatial and modal superconvergence of the discontinuous Galerkin method for linear equations. Journal of Scientific Computing, 72(1):128–146, July 2017. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0349-8.pdf.

DellAcqua:2017:SPP

[1850] Pietro Dell’Acqua, Marco Donatelli, Claudio Estatico, and Mariarosa Mazza. Structure preserving preconditioners for image deblurring. Journal of Scientific Computing, 72(1):147–171, July 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/content/pdf/10.1007/s10915-016-0350-2.pdf.

Kang:2017:TGV

[1851] Myeongmin Kang, Myungjoo Kang, and Miyoun Jung. Total generalized variation based denoising models for ultrasound images. Journal of Scientific Computing, 72(1):172–197, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-017-0357-3.pdf.

Reisinger:2017:BTM

[1852] Christoph Reisinger and Julen Rotaetxe Arto. Boundary treatment and multigrid preconditioning for semi-Lagrangian schemes applied to Hamilton–Jacobi–Bellman equations. Journal of Scientific Computing, 72(1):198–230, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0351-1.pdf.

Chen:2017:UCC

[1853] Shao chun Chen, Li na Dong, and Ji kun Zhao. Uniformly convergent cubic nonconforming element for Darcy–Stokes problem. Journal of Scientific Computing, 72(1):231–251, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-016-0353-z.pdf.

Sen:2017:DNN

[1854] Shuvam Sen and Tony W. H. Sheu. On the development of a nonprimitive Navier–Stokes formulation subject to rigorous implementation of a new vorticity integral condition. Journal of Scientific Computing, 72 (1):252–290, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-016-0355-x.pdf.

Liu:2017:EIT

[1855] Jian-Guo Liu, Zheng Ma, and Zhennan Zhou. Explicit and implicit TVD schemes for conservation laws with Caputo derivatives. Journal of Scientific Computing, 72(1):291–313, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-017-0356-4.pdf.

Kopriva:2017:EBD

[1856] David A. Kopriva, Jan Nordström, and Gregor J. Gassner. Error boundedness of discontinuous Galerkin spectral element approximations of hyperbolic problems. Journal of Scientific Computing, 72(1):314– 330, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/content/pdf/
10.1007/s10915-017-0358-2.pdf.

Zhu:2017:NNO

[1857] Hong Zhu, Xiaowei Zhang, Delin Chu, and Li-Zhi Liao. Nonconvex and nonsmooth optimization with generalized orthogonality constraints: an approximate augmented Lagrangian method. Journal of Scientific Computing, 72(1):331–372, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/content/pdf/10.1007/s10915-017-0359-1.pdf.

Farrell:2017:USO

[1858] Patricio Farrell and Alexander Linke. Uniform second order convergence of a complete flux scheme on unstructured 1D grids for a singularly perturbed advection–diffusion equation and some multidimensional extensions. Journal of Scientific Computing, 72(1):373–395, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-017-0361-7.pdf.

Forti:2017:MAF

[1859] Davide Forti, Martina Bukac, Annalisa Quaini, Suncica Canic, and Simone Deparis. A monolithic approach to fluid–composite structure interaction. Journal of Scientific Computing, 72(1):396–421, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL https://link.springer.com/content/pdf/10.1007/
s10915-017-0363-5.pdf.

Bu:2017:FDF

[1860] Weiping Bu, Aiguo Xiao, and Wei Zeng. Finite difference/finite element methods for distributed-order time fractional diffusion equations. Journal of Scientific Computing, 72(1):422–441, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/content/pdf/10.1007/s10915-017-0360-8.pdf.

Hahn:2017:IBG

[1861] Jooyoung Hahn, Karol Mikula, Peter Frolkovic, and Branislav Basara. Inflow-based gradient finite volume method for a propagation in a normal direction in a polyhedron mesh. Journal of Scientific Computing, 72 (1):442–465, July 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/content/
pdf/10.1007/s10915-017-0364-4.pdf.

Li:2017:MVE

[1862] Changfeng Li, Yirang Yuan, Tongjun Sun, and Qing Yang. Mixed volume element-characteristic fractional step difference method for contamination from nuclear waste disposal. Journal of Scientific Computing, 72(2):467–499, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0365-3;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0365-3.pdf.

Zhang:2017:IAS

[1863] Geng Zhang, Dan Cai, and Yong Du. Improved ADI scheme for linear hyperbolic equations: Extension to nonlinear cases and compact ADI schemes. Journal of Scientific Computing, 72(2):500–521, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0366-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0366-2.pdf.

Jia:2017:NTS

[1864] Zhi-Gang Jia and Musheng Wei. A new TV–Stokes model for image deblurring and denoising with fast algorithms. Journal of Scientific Computing, 72(2):522–541, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0368-0;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0368-0.pdf.

Gurkan:2017:EHD

[1865] Ceren Gürkan, Martin Kronbichler, and Sonia Fernández-Méndez. eXtended hybridizable discontinuous Galerkin with Heaviside enrichment for heat bimaterial problems. Journal of Scientific Computing, 72(2):542– 567, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0370-6;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0370-6.pdf.

Ervin:2017:DMO

[1866] Vincent J. Ervin, Thomas Führer, Norbert Heuer, and Michael Karkulik. DPG method with optimal test functions for a fractional advection diffusion equation. Journal of Scientific Computing, 72(2):568–585, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0369-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0369-z.pdf.

Xie:2017:MCM

[1867] Hehu Xie and Xinming Wu. A multilevel correction method for interior transmission eigenvalue problem. Journal of Scientific Computing, 72(2):586–604, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0367-1;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0367-1.pdf.

Song:2017:EPA

[1868] Yifu Song and Yushun Wang. Energy-preserving algorithms for the Benjamin equation. Journal of Scientific Computing, 72(2):605–622, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0371-5;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0371-5.pdf.

Barsukow:2017:NSC

[1869] Wasilij Barsukow, Philipp V. F. Edelmann, Christian Klingenberg, Fabian Miczek, and Friedrich K. Röpke. A numerical scheme for the compressible low-Mach number regime of ideal fluid dynamics. Journal of Scientific Computing, 72(2):623–646, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0372-4;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0372-4.pdf.

Wang:2017:HVS

[1870] Chuan-Li Wang, Zhong-Qing Wang, and Hong-Li Jia. An hp-version spectral collocation method for nonlinear Volterra integro-differential equation with weakly singular kernels. Journal of Scientific Computing, 72(2):647–678, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0373-3;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0373-3.pdf.

Zhang:2017:EST

[1871] Chao Zhang, Dong qin Gu, Zhong qing Wang, and Hui yuan Li. Efficient space–time spectral methods for second-order problems on unbounded domains. Journal of Scientific Computing, 72(2):679–699, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0374-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0374-2.pdf.

Xu:2017:GCA

[1872] Yangyang Xu and Wotao Yin. A globally convergent algorithm for nonconvex optimization based on block coordinate update. Journal of Scientific Computing, 72(2):700–734, August 2017. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0376-0;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0376-0.pdf.

Besse:2017:DDA

[1873] Christophe Besse and Feng Xing. Domain decomposition algorithms for two dimensional linear Schrödinger equation. Journal of Scientific Computing, 72(2):735–760, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0375-1;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0375-1.pdf.

Cao:2017:SLD

[1874] Waixiang Cao and Qiumei Huang. Superconvergence of local discontinuous Galerkin methods for partial differential equations with higher order derivatives. Journal of Scientific Computing, 72(2):761–791, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0377-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0377-z.pdf.

Cheng:2017:LAL

[1875] Yao Cheng and Qiang Zhang. Local analysis of the local discontinuous Galerkin method with generalized alternating numerical flux for onedimensional singularly perturbed problem. Journal of Scientific Computing, 72(2):792–819, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0378-y;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0378-y.pdf.

Gong:2017:AMC

[1876] Wei Gong, Hehu Xie, and Ningning Yan. Adaptive multilevel correction method for finite element approximations of elliptic optimal control problems. Journal of Scientific Computing, 72(2):820–841, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0386-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0386-y.pdf.

Wu:2017:OOS

[1877] Shu-Lin Wu. Optimized overlapping Schwarz waveform relaxation for a class of time-fractional diffusion problems. Journal of Scientific Computing, 72(2):842–862, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0379-x;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0379-x.pdf.

Li:2017:TGB

[1878] Xiaoli Li and Hongxing Rui. A two-grid block-centered finite difference method for the nonlinear time-fractional parabolic equation. Journal of Scientific Computing, 72(2):863–891, August 2017. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0380-4;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0380-4.pdf.

Li:2017:UOE

[1879] Dongfang Li and Jilu Wang. Unconditionally optimal error analysis of Crank–Nicolson Galerkin FEMs for a strongly nonlinear parabolic system. Journal of Scientific Computing, 72(2):892–915, August 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0381-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0381-3.pdf.

Ren:2017:SFE

[1880] Jincheng Ren, Xiaonian Long, Shipeng Mao, and Jiwei Zhang. Superconvergence of finite element approximations for the fractional diffusionwave equation. Journal of Scientific Computing, 72(3):917–935, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0385-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0385-z.pdf.

Jiang:2017:AFE

[1881] Xue Jiang, Peijun Li, Junliang Lv, and Weiying Zheng. An adaptive finite element method for the wave scattering with transparent boundary condition. Journal of Scientific Computing, 72(3):936–956, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0382-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0382-2.pdf.

Gu:2017:FIM

[1882] Xian-Ming Gu, Ting-Zhu Huang, Cui-Cui Ji, Bruno Carpentieri, and Anatoly A. Alikhanov. Fast iterative method with a second-order implicit difference scheme for time–space fractional convection–diffusion equation. Journal of Scientific Computing, 72(3):957–985, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0388-9;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0388-9.pdf.

Bigoni:2017:AWM

[1883] Caterina Bigoni and Jan S. Hesthaven. Adaptive WENO methods based on radial basis function reconstruction. Journal of Scientific Computing, 72(3):986–1020, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0383-1;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0383-1.pdf.

Zhu:2017:NSF

[1884] Yujie Zhu, Zhensheng Sun, Yuxin Ren, Yu Hu, and Shiying Zhang. A numerical strategy for freestream preservation of the high order weighted essentially non-oscillatory schemes on stationary curvilinear grids. Journal of Scientific Computing, 72(3):1021–1048, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0387-x;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0387-x.pdf.

Garcia:2017:NMF

[1885] Carlos García, Gabriel N. Gatica, and Salim Meddahi. A new mixed finite element method for elastodynamics with weak symmetry. Journal of Scientific Computing, 72(3):1049–1079, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0384-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0384-0.pdf.

Suarez:2017:RSW

[1886] Jean-Piero Suarez and Gustaaf B. Jacobs. Regularization of singularities in the weighted summation of Dirac-delta functions for the spectral solution of hyperbolic conservation laws. Journal of Scientific Computing, 72(3):1080–1092, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0389-8;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0389-8.pdf.

Shi:2017:USAb

[1887] Dongyang Shi and Junjun Wang. Unconditional superconvergence analysis of a Crank–Nicolson Galerkin FEM for nonlinear Schrödinger equation. Journal of Scientific Computing, 72(3):1093–1118, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0390-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0390-2.pdf.

Luo:2017:PFE

[1888] Li Luo, Qian Zhang, Xiao-Ping Wang, and Xiao-Chuan Cai. A parallel finite element method for 3D two-phase moving contact line problems in complex domains. Journal of Scientific Computing, 72(3):1119–1145, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0391-1;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0391-1.pdf.

Deng:2017:NOM

[1889] Xi Deng, Ziyao Sun, Bin Xie, Kensuke Yokoi, Chungang Chen, and Feng Xiao. A non-oscillatory multi-moment finite volume scheme with boundary gradient switching. Journal of Scientific Computing, 72 (3):1146–1168, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0392-0;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0392-0.pdf.

Baffet:2017:HOAb

[1890] Daniel Baffet and Jan S. Hesthaven. High-order accurate adaptive kernel compression time-stepping schemes for fractional differential equations. Journal of Scientific Computing, 72(3):1169–1195, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0393-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0393-z.pdf.

Peterseim:2017:RCC

[1891] Daniel Peterseim and Mira Schedensack. Relaxing the CFL condition for the wave equation on adaptive meshes. Journal of Scientific Computing, 72(3):1196–1213, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0394-y;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0394-y.pdf.

Hou:2017:NAF

[1892] Tianliang Hou, Tao Tang, and Jiang Yang. Numerical analysis of fully discretized Crank–Nicolson scheme for fractional-in-space Allen– Cahn equations. Journal of Scientific Computing, 72(3):1214–1231, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0396-9;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0396-9.pdf.

Abgrall:2017:CAC

[1893] R. Abgrall, Q. Viville, H. Beaugendre, and C. Dobrzynski. Construction of a p-adaptive continuous residual distribution scheme. Journal of Scientific Computing, 72(3):1232–1268, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0399-6;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0399-6.pdf.

Gao:2017:LCF

[1894] Huadong Gao and Dongdong He. Linearized conservative finite element methods for the Nernst–Planck–Poisson equations. Journal of Scientific Computing, 72(3):1269–1289, September 2017. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0400-4;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0400-4.pdf.

Bajars:2017:IAP

[1895] Janis Bajars, David J. Chappell, Timo Hartmann, and Gregor Tanner. Improved approximation of phase-space densities on triangulated domains using discrete flow mapping with p-refinement. Journal of Scientific Computing, 72(3):1290–1312, September 2017. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0397-8;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0397-8.pdf.

Cai:2017:TSA

[1896] Xiaohao Cai, Raymond Chan, Mila Nikolova, and Tieyong Zeng. A three-stage approach for segmenting degraded color images: Smoothing, lifting and thresholding (SLaT). Journal of Scientific Computing, 72 (3):1313–1332, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0402-2;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0402-2.pdf.

Benninghoff:2017:STD

[1897] Heike Benninghoff and Harald Garcke. Segmentation of threedimensional images with parametric active surfaces and topology changes. Journal of Scientific Computing, 72(3):1333–1367, September 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0401-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0401-3.pdf.

Lu:2017:CTO

[1898] Xiliang Lu, Lijuan Wang, and Qishu Yan. Computation of time optimal control problems governed by linear ordinary differential equations. Journal of Scientific Computing, 73(1):1–25, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0403-1;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0403-1.pdf.

Fu:2017:TSO

[1899] Kai Fu and Dong Liang. The time second order mass conservative characteristic FDM for advection–diffusion equations in high dimensions. Journal of Scientific Computing, 73(1):26–49, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0404-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0404-0.pdf.

Gao:2017:MMF

[1900] Yanni Gao and Yonghai Li. A mortar mixed finite volume method for elliptic problems on non-matching multi-block triangular grids. Journal of Scientific Computing, 73(1):50–69, October 2017. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0405-z;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0405-z.pdf.

Rui:2017:BCF

[1901] Hongxing Rui and Hao Pan. A block–centered finite difference method for slightly compressible Darcy–Forchheimer flow in porous media. Journal of Scientific Computing, 73(1):70–92, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0406-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0406-y.pdf.

Gao:2017:TSO

[1902] Guang hua Gao, Anatoly A. Alikhanov, and Zhi zhong Sun. The temporal second order difference schemes based on the interpolation approximation for solving the time multi-term and distributed-order fractional sub-diffusion equations. Journal of Scientific Computing, 73(1):93–121, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0407-x;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0407-x.pdf.

Yano:2017:KML

[1903] Ryosuke Yano. Kinetic modeling of local epidemic spread and its simulation. Journal of Scientific Computing, 73(1):122–156, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0408-9;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0408-9.pdf.

Pitton:2017:ARB

[1904] Giuseppe Pitton and Gianluigi Rozza. On the application of reduced basis methods to bifurcation problems in incompressible fluid dynamics. Journal of Scientific Computing, 73(1):157–177, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0419-6;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0419-6.pdf.

Roy:2017:NIC

[1905] Souvik Roy and Alfio Borzì. Numerical investigation of a class of Liouville control problems. Journal of Scientific Computing, 73(1):178–202, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0410-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0410-2.pdf.

Yueh:2017:EEM

[1906] Mei-Heng Yueh, Wen-Wei Lin, Chin-Tien Wu, and Shing-Tung Yau. An efficient energy minimization for conformal parameterizations. Journal of Scientific Computing, 73(1):203–227, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0414-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0414-y.pdf.

Liu:2017:SFS

[1907] D. J. Liu, D. D. Jiang, Y. Liu, and Q. Q. Xia. Stabilized FEM for some optimal design problem. Journal of Scientific Computing, 73(1):228–241, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0409-8;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0409-8.pdf.

Han:2017:TFP

[1908] Houde Han, Yintzer Shih, and Dongsheng Yin. Tailored finite point methods for solving singularly perturbed eigenvalue problems with higher eigenvalues. Journal of Scientific Computing, 73(1):242–282, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0411-1;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0411-1.pdf.

Yu:2017:SMV

[1909] Xu hong Yu and Ben yu Guo. Spectral method for vorticity-stream function form of Navier–Stokes equations in an infinite channel with slip boundary conditions. Journal of Scientific Computing, 73(1):283–302, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0413-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0413-z.pdf.

Wu:2017:RNM

[1910] Xinming Wu, Zaiwen Wen, and Weizhu Bao. A regularized Newton method for computing ground states of Bose–Einstein condensates. Journal of Scientific Computing, 73(1):303–329, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0412-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0412-0.pdf.

Guzman:2017:FEM

[1911] Johnny Guzmán, Manuel A. Sánchez, and Marcus Sarkis. A finite element method for high-contrast interface problems with error estimates independent of contrast. Journal of Scientific Computing, 73(1):330–365, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0415-x;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0415-x.pdf.

Brink:2017:HFE

[1912] Freekjan Brink, Ferenc Izsák, and J. J. W. van der Vegt. Hamiltonian finite element discretization for nonlinear free surface water waves. Journal of Scientific Computing, 73(1):366–394, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0416-9;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0416-9.pdf.

Hao:2017:IAB

[1913] Zhaopeng Hao and Wanrong Cao. An improved algorithm based on finite difference schemes for fractional boundary value problems with nonsmooth solution. Journal of Scientific Computing, 73(1):395–415, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0417-8;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0417-8.pdf.

Zosso:2017:EPD

[1914] Dominique Zosso, Braxton Osting, Mandy (Mengqi) Xia, and Stanley J. Osher. An efficient primal–dual method for the obstacle problem. Journal of Scientific Computing, 73(1):416–437, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0420-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0420-0.pdf.

Leng:2017:CQO

[1915] Haitao Leng and Yanping Chen. Convergence and quasi-optimality of an adaptive finite element method for optimal control problems on L^{2} errors. Journal of Scientific Computing, 73(1):438–458, October 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0425-8;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0425-8.pdf.

Gottlieb:2017:SIH

[1916] Sigal Gottlieb, Johnny Guzmán, Fengyan Li, and Jennifer K. Ryan. Special issue in honor of Professor Chi-Wang Shu. Journal of Scientific Computing, 73(2–3):459–460, December 2017. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0566-9;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0566-9.pdf.

Abgrall:2017:HOS

[1917] R. Abgrall. High order schemes for hyperbolic problems using globally continuous approximation and avoiding mass matrices. Journal of Scientific Computing, 73(2–3):461–494, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0498-4;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0498-4.pdf.

Brenner:2017:HDM

[1918] Susanne C. Brenner, Jiguang Sun, and Li yeng Sung. Hodge decomposition methods for a quad-curl problem on planar domains. Journal of Scientific Computing, 73(2–3):495–513, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0449-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0449-0.pdf.

Cai:2017:HOC

[1919] Xiaofeng Cai, Wei Guo, and Jing-Mei Qiu. A high order conservative semi-Lagrangian discontinuous Galerkin method for two-dimensional transport simulations. Journal of Scientific Computing, 73(2–3):514–542, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0554-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0554-0.pdf.

Cao:2017:SIF

[1920] Waixiang Cao, Xu Zhang, Zhimin Zhang, and Qingsong Zou. Superconvergence of immersed finite volume methods for one-dimensional interface problems. Journal of Scientific Computing, 73(2–3):543–565, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0532-6;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0532-6.pdf.

Chen:2017:DIS

[1921] Zheng Chen, Liu Liu, and Lin Mu. DG-IMEX stochastic Galerkin schemes for linear transport equation with random inputs and diffusive scalings. Journal of Scientific Computing, 73(2–3):566–592, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0439-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0439-2.pdf.

Chou:2017:LDG

[1922] Ching-Shan Chou, Weizhou Sun, Yulong Xing, and He Yang. Local discontinuous Galerkin methods for the Khokhlov–Zabolotskaya– Kuznetzov equation. Journal of Scientific Computing, 73(2–3):593–616, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0502-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0502-z.pdf.

Chow:2017:AOC

[1923] Yat Tin Chow, Jérôme Darbon, Stanley Osher, and Wotao Yin. Algorithm for overcoming the curse of dimensionality for time-dependent non-convex Hamilton–Jacobi equations arising from optimal control and differential games problems. Journal of Scientific Computing, 73 (2–3):617–643, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0436-5;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0436-5.pdf.

Cockburn:2017:ABS

[1924] Bernardo Cockburn and Zhu Wang. Adjoint-based, superconvergent Galerkin approximations of linear functionals. Journal of Scientific Computing, 73(2–3):644–666, December 2017. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0507-7;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0507-7.pdf.

Conde:2017:IIE

[1925] Sidafa Conde, Sigal Gottlieb, Zachary J. Grant, and John N. Shadid. Implicit and implicit–explicit strong stability preserving Runge–Kutta methods with high linear order. Journal of Scientific Computing, 73 (2–3):667–690, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0560-2;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0560-2.pdf.

Ditkowski:2017:EIB

[1926] A. Ditkowski and S. Gottlieb. Error inhibiting block one-step schemes for ordinary differential equations. Journal of Scientific Computing, 73 (2–3):691–711, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0441-8;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0441-8.pdf.

Dong:2017:OCH

[1927] Bo Dong. Optimally convergent HDG method for third-order Korteweg– de Vries type equations. Journal of Scientific Computing, 73(2– 3):712–735, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0437-4;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0437-4.pdf.

Gao:2017:ERH

[1928] Zhen Gao, Xiao Wen, and Wai Sun Don. Enhanced robustness of the hybrid compact–WENO finite difference scheme for hyperbolic conservation laws with multi-resolution analysis and Tukey’s boxplot method. Journal of Scientific Computing, 73(2–3):736–752, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0465-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0465-0.pdf.

Hernandez-Duenas:2017:HMS

[1929] Gerardo Hernandez-Duenas. A hybrid method to solve shallow water flows with horizontal density gradients. Journal of Scientific Computing, 73(2–3):753–782, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0553-1;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0553-1.pdf.

Hou:2017:AMI

[1930] Yifeng Hou and Roger Temam. About the modeling of the indentation of a virus shell: The role of the shape of the probe. Journal of Scientific Computing, 73(2–3):783–796, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0481-0;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0481-0.pdf.

Hu:2017:CIE

[1931] Jingwei Hu and Xiangxiong Zhang. On a class of implicit–explicit Runge–Kutta schemes for stiff kinetic equations preserving the Navier– Stokes limit. Journal of Scientific Computing, 73(2–3):797–818, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0499-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0499-3.pdf.

Jameson:2017:EFI

[1932] Antony Jameson. Evaluation of fully implicit Runge–Kutta schemes for unsteady flow calculations. Journal of Scientific Computing, 73 (2–3):819–852, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0476-x;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0476-x.pdf.

Jiang:2017:OER

[1933] Jiahua Jiang, Yanlai Chen, and Akil Narayan. Offline-enhanced reduced basis method through adaptive construction of the surrogate training set. Journal of Scientific Computing, 73(2–3):853–875, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0551-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0551-3.pdf.

Khan:2017:SEM

[1934] Arbaz Khan, Pravir Dutt, and Chandra Shekhar Upadhyay. Spectral element method for parabolic initial value problem with non-smooth data: Analysis and application. Journal of Scientific Computing, 73 (2–3):876–905, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0457-0;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0457-0.pdf.

Klingenberg:2017:ALE

[1935] Christian Klingenberg, Gero Schnücke, and Yinhua Xia. An arbitrary Lagrangian–Eulerian local discontinuous Galerkin method for Hamilton– Jacobi equations. Journal of Scientific Computing, 73(2–3):906–942, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0471-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0471-2.pdf.

Li:2017:LDG

[1936] Xingjie Helen Li, Chi-Wang Shu, and Yang Yang. Local discontinuous Galerkin method for the Keller–Segel chemotaxis model. Journal of Scientific Computing, 73(2–3):943–967, December 2017. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-016-0354-y.pdf.

Li:2017:AMO

[1937] Zhilin Li and Fangfang Qin. An augmented method for 4th order PDEs with discontinuous coefficients. Journal of Scientific Computing, 73(2–3):968–979, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0487-7.pdf.

Lu:2017:CCS

[1938] Dong Lu and Yong-Tao Zhang. Computational complexity study on Krylov integration factor WENO method for high spatial dimension convection–diffusion problems. Journal of Scientific Computing, 73(2–3): 980–1027, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10915-017-0398-7.pdf.

Liu:2017:SBP

[1939] Yuan Liu, Yingda Cheng, and Chi-Wang Shu. A simple bound-preserving sweeping technique for conservative numerical approximations. Journal of Scientific Computing, 73(2–3):1028–1071, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-017-0395-x.pdf.

Mirzargar:2017:HSI

[1940] Mahsa Mirzargar, Ashok Jallepalli, Jennifer K. Ryan, and Robert M. Kirby. Hexagonal smoothness-increasing accuracy-conserving filtering. Journal of Scientific Computing, 73(2–3):1072–1093, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-017-0517-5.pdf.

Roe:2017:DRR

[1941] Philip Roe. Is discontinuous reconstruction really a good idea? Journal of Scientific Computing, 73(2–3):1094–1114, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-017-0555-z.pdf.

Romero:2017:DFR

[1942] J. Romero, F. D. Witherden, and A. Jameson. A direct flux reconstruction scheme for advection–diffusion problems on triangular grids. Journal of Scientific Computing, 73(2–3):1115–1144, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-017-0472-1.pdf.

Schutz:2017:IMC

[1943] Jochen Schütz, David C. Seal, and Alexander Jaust. Implicit multiderivative collocation solvers for linear partial differential equations with discontinuous Galerkin spatial discretizations. Journal of Scientific Computing, 73(2–3):1145–1163, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-017-0485-9.pdf.

Shi:2017:WNL

[1944] Zuoqiang Shi, Stanley Osher, and Wei Zhu. Weighted nonlocal Laplacian on interpolation from sparse data. Journal of Scientific Computing, 73(2–3):1164–1177, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0421-z.pdf.

Song:2017:UES

[1945] Huailing Song and Chi-Wang Shu. Unconditional energy stability analysis of a second order implicit–explicit local discontinuous Galerkin method for the Cahn–Hilliard equation. Journal of Scientific Computing, 73(2–3):1178–1203, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0497-5.pdf.

Vogl:2017:HRF

[1946] Christopher J. Vogl and Randall J. LeVeque. A high-resolution finite volume seismic model to generate seafloor deformation for tsunami modeling. Journal of Scientific Computing, 73(2–3):1204–1215, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0459-y.pdf.

Yang:2017:DGM

[1947] He Yang and Fengyan Li. Discontinuous Galerkin methods for relativistic Vlasov–Maxwell system. Journal of Scientific Computing, 73(2–3):1216– 1248, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-016-0332-4.pdf.

Yu:2017:CLD

[1948] Fan Yu, Hui Guo, Nattaporn Chuenjarern, and Yang Yang. Conservative local discontinuous Galerkin method for compressible miscible displacements in porous media. Journal of Scientific Computing, 73(2–3):1249– 1275, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0571-z.pdf.

Zhang:2017:FTS

[1949] Mengping Zhang and Jue Yan. Fourier type super convergence study on DDGIC and symmetric DDG methods. Journal of Scientific Computing, 73(2–3):1276–1289, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-017-0438-3.pdf.

Zhou:2017:SLD

[1950] Lingling Zhou, Yan Xu, Zhimin Zhang, and Waixiang Cao. Superconvergence of local discontinuous Galerkin method for one-dimensional linear Schrödinger equations. Journal of Scientific Computing, 73(2–3):1290– 1315, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0362-6.pdf.

Zhu:2017:ARM

[1951] Hongqiang Zhu, Jianxian Qiu, and Jing-Mei Qiu. An h-adaptive RKDG method for the two-dimensional incompressible Euler equations and the guiding center Vlasov model. Journal of Scientific Computing, 73(2–3): 1316–1337, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10915-017-0440-9.pdf.

Zhu:2017:NTF

[1952] Jun Zhu and Jianxian Qiu. A new type of finite volume WENO schemes for hyperbolic conservation laws. Journal of Scientific Computing, 73(2–3):1338–1359, December 2017. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-017-0486-8.pdf.

Cavoretto:2018:OSL

[1953] Roberto Cavoretto, Alessandra De Rossi, and Emma Perracchione. Optimal selection of local approximants in RBF–PU interpolation. Journal of Scientific Computing, 74(1):1–22, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0418-7.

Bi:2018:PEE

[1954] Chunjia Bi, Cheng Wang, and Yanping Lin. A posteriori error estimates of two-grid finite element methods for nonlinear elliptic problems. Journal of Scientific Computing, 74(1):23–48, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-017-0422-y.

Zhou:2018:FLT

[1955] Zhiqiang Zhou, Jingtang Ma, and Hai wei Sun. Fast Laplace transform methods for free-boundary problems of fractional diffusion equations.

Journal of Scientific Computing, 74(1):49–69, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-017-0423-x.

You:2018:EBI

[1956] Guoqiao You and Shingyu Leung. Eulerian based interpolation schemes for flow map construction and line integral computation with applications to Lagrangian coherent structures extraction. Journal of Scientific Computing, 74(1):70–96, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0424-9.

Jin:2018:DSG

[1957] Shi Jin and Zheng Ma. The discrete stochastic Galerkin method for hyperbolic equations with non-smooth and random coefficients. Journal of Scientific Computing, 74(1):97–121, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0426-7.

Lu:2018:AMM

[1958] Changna Lu, Weizhang Huang, and Jianxian Qiu. An adaptive moving mesh finite element solution of the regularized long wave equation. Journal of Scientific Computing, 74(1):122–144, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-017-0427-6.

Smears:2018:NDD

[1959] Iain Smears. Nonoverlapping domain decomposition preconditioners for discontinuous Galerkin approximations of Hamilton–Jacobi– Bellman equations. Journal of Scientific Computing, 74(1):145–174, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0428-5.

Yin:2018:ERF

[1960] Ke Yin and Xue-Cheng Tai. An effective region force for some variational models for learning and clustering. Journal of Scientific Computing, 74 (1):175–196, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0429-4.

Martini:2018:CRB

[1961] Immanuel Martini, Bernard Haasdonk, and Gianluigi Rozza. Certified reduced basis approximation for the coupling of viscous and inviscid parametrized flow models. Journal of Scientific Computing, 74(1): 197–219, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0430-y.

Fu:2018:PDR

[1962] Hongfei Fu, Hong Wang, and Zhu Wang. POD/DEIM reduced-order modeling of time-fractional partial differential equations with applications in parameter identification. Journal of Scientific Computing, 74 (1):220–243, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0433-8.

Fekete:2018:PCS

[1963] Imre Fekete, David I. Ketcheson, and Lajos Lóczi. Positivity for convective semi-discretizations. Journal of Scientific Computing, 74(1): 244–266, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0432-9.

Cavoretto:2018:OBP

[1964] Roberto Cavoretto, Teseo Schneider, and Patrick Zulian. OpenCL based parallel algorithm for RBF–PUM interpolation. Journal of Scientific Computing, 74(1):267–289, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0431-x.

Petersson:2018:HOA

[1965] N. Anders Petersson and Björn Sjögreen. High order accurate finite difference modeling of seismo–acoustic wave propagation in a moving atmosphere and a heterogeneous Earth model coupled across a realistic topography. Journal of Scientific Computing, 74(1):290–323, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-017-0434-7.

Meng:2018:NRN

[1966] Zhaoliang Meng, Jintao Cui, and Zhongxuan Luo. A new rotated nonconforming quadrilateral element. Journal of Scientific Computing, 74 (1):324–335, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0435-6.

Cai:2018:NSM

[1967] Zhenning Cai and Manuel Torrilhon. Numerical simulation of microflows using moment methods with linearized collision operator. Journal of Scientific Computing, 74(1):336–374, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0442-7.

Chaabane:2018:SDG

[1968] Nabil Chaabane and Béatrice Rivière. A sequential discontinuous Galerkin method for the coupling of flow and geomechanics. Journal of Scientific Computing, 74(1):375–395, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0443-6.

Huang:2018:MMM

[1969] Jian Huang, Long Chen, and Hongxing Rui. Multigrid methods for a mixed finite element method of the Darcy–Forchheimer model. Journal of Scientific Computing, 74(1):396–411, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0466-z.

Xie:2018:TMS

[1970] Ze-Jia Xie, Xiao-Qing Jin, and Yi-Min Wei. Tensor methods for solving symmetric ⇕-tensor systems. Journal of Scientific Computing, 74(1): 412–425, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0444-5.

deFrutos:2018:EAP

[1971] Javier de Frutos, Bosco García-Archilla, and Julia Novo. Error analysis of projection methods for non inf-sup stable mixed finite elements: The Navier–Stokes equations. Journal of Scientific Computing, 74(1): 426–455, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0446-3.

Durango:2018:TGM

[1972] Francisco Durango and Julia Novo. Two-grid mixed finite-element approximations to the Navier–Stokes equations based on a Newton-type step. Journal of Scientific Computing, 74(1):456–473, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-017-0447-2.

Jiang:2018:EBI

[1973] Shidong Jiang, Dong Wang, and Xiao-Ping Wang. An efficient boundary integral scheme for the MBO threshold dynamics method via the NUFFT. Journal of Scientific Computing, 74(1):474–490, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Minjeaud:2018:HOC

[1974] Sebastian Minjeaud and Richard Pasquetti. High order C^{0}-continuous Galerkin schemes for high order PDEs, conservation of quadratic invariants and application to the Korteweg–de Vries model. Journal of Scientific Computing, 74(1):491–518, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Karaa:2018:OEA

[1975] Samir Karaa, Kassem Mustapha, and Amiya K. Pani. Optimal error analysis of a FEM for fractional diffusion problems by energy arguments. Journal of Scientific Computing, 74(1):519–535, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Yi:2018:FDH

[1976] Nianyu Yi, Yunqing Huang, and Wei Yang. Function, derivative and high-order derivatives recovery methods using the local symmetry projection. Journal of Scientific Computing, 74(1):536–572, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Karakashian:2018:TLA

[1977] O. Karakashian and C. Collins. Two-level additive Schwarz methods for discontinuous Galerkin approximations of the biharmonic equation. Journal of Scientific Computing, 74(1):573–604, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Cao:2018:CSI

[1978] Waixiang Cao, Xu Zhang, Zhimin Zhang, and Qingsong Zou. Correction to: Superconvergence of immersed finite volume methods for onedimensional interface problems. Journal of Scientific Computing, 74(1): 605, January 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0609-2.pdf.

Kim:2018:HFD

[1979] Chang Ho Kim, Kwang-Il You, and Youngsoo Ha. Hybrid finite difference weighted essentially non-oscillatory schemes for the compressible ideal magnetohydrodynamics equation. Journal of Scientific Computing, 74 (2):607–630, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0462-3.

Seo:2018:CAM

[1980] Jiwon Seo, Seung yeal Ha, and Chohong Min. Convergence analysis in the maximum norm of the numerical gradient of the Shortley–Weller method. Journal of Scientific Computing, 74(2):631–639, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-017-0458-z.

Li:2018:HOP

[1981] Peng Li, Wai-Sun Don, Cheng Wang, and Zhen Gao. High order positivity- and bound-preserving hybrid compact-WENO finite difference scheme for the compressible Euler equations. Journal of Scientific Computing, 74(2):640–666, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0452-5.

Rebollo:2018:HOL

[1982] Tomás Chacón Rebollo, Macarena Gómez Mármol, Frédéric Hecht, Samuele Rubino, and Isabel Sánchez Muñoz. A high-order local projection stabilization method for natural convection problems. Journal of Scientific Computing, 74(2):667–692, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0469-9.

Li:2018:DFV

[1983] Rui Li, Yali Gao, Jian Li, and Zhangxin Chen. Discontinuous finite volume element method for a coupled non-stationary Stokes–Darcy problem. Journal of Scientific Computing, 74(2):693–727, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-017-0454-3.

Liu:2018:FFD

[1984] Zhengguang Liu and Xiaoli Li. A fast finite difference method for a continuous static linear bond-based peridynamics model of mechanics. Journal of Scientific Computing, 74(2):728–742, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-017-0456-1.

Mei:2018:CNR

[1985] Jin-Jin Mei, Yiqiu Dong, Ting-Zhu Huang, and Wotao Yin. Cauchy noise removal by nonconvex ADMM with convergence guarantees. Journal of Scientific Computing, 74(2):743–766, February 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0460-5;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0460-5.pdf.

Lou:2018:FLL

[1986] Yifei Lou and Ming Yan. Fast L_{1}–L_{2} minimization via a proximal operator. Journal of Scientific Computing, 74(2):767–785, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-017-0463-2.

Buccini:2018:WCL

[1987] Alessandro Buccini, Marco Donatelli, and Fabio Ferri. Weakly constrained Lucy–Richardson with applications to inversion of light scattering data. Journal of Scientific Computing, 74(2):786–804, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-017-0461-4.

Zamolo:2018:DLM

[1988] Riccardo Zamolo, Lucia Parussini, and Valentino Pediroda. Distributed Lagrange multiplier functions for fictitious domain method with spectral/hp element discretization. Journal of Scientific Computing, 74(2): 805–825, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0464-1.

Tao:2018:OLC

[1989] Min Tao and Xiaoming Yuan. On the optimal linear convergence rate of a generalized proximal point algorithm. Journal of Scientific Computing, 74(2):826–850, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0477-9.

Shi:2018:DDF

[1990] Ruonan Shi and Jae-Hun Jung. A domain decomposition Fourier continuation method for enhanced L_{1} regularization using sparsity of edges in reconstructing Fourier data. Journal of Scientific Computing, 74(2): 851–871, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0467-y.

Dunst:2018:SOC

[1991] Thomas Dunst and Andreas Prohl. Stochastic optimal control of finite ensembles of nanomagnets. Journal of Scientific Computing, 74(2): 872–894, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0474-z.

Huber:2018:SCS

[1992] Markus Huber, Ulrich Rüde, Christian Waluga, and Barbara Wohlmuth. Surface couplings for subdomain-wise isoviscous gradient based Stokes finite element discretizations. Journal of Scientific Computing, 74(2): 895–919, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0470-3.

Tang:2018:WCP

[1993] Xiaojun Tang, Yang Shi, and Heyong Xu. Well conditioned pseudospectral schemes with tunable basis for fractional delay differential equations. Journal of Scientific Computing, 74(2):920–936, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-017-0473-0.

Mukam:2018:SCA

[1994] Jean Daniel Mukam and Antoine Tambue. Strong convergence analysis of the stochastic exponential Rosenbrock scheme for the finite element discretization of semilinear SPDEs driven by multiplicative and additive noise. Journal of Scientific Computing, 74(2):937–978, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-017-0475-y.

An:2018:OEE

[1995] Rong An and Jian Su. Optimal error estimates of semi-implicit Galerkin method for time-dependent nematic liquid crystal flows. Journal of Scientific Computing, 74(2):979–1008, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0479-7.

Zhao:2018:FFD

[1996] Meng Zhao, Hong Wang, and Aijie Cheng. A fast finite difference method for three-dimensional time-dependent space-fractional diffusion equations with fractional derivative boundary conditions. Journal of Scientific Computing, 74(2):1009–1033, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-017-0478-8.

Chen:2018:UCC

[1997] Minghua Chen, Weihua Deng, and Stefano Serra-Capizzano. Uniform convergence of V -cycle multigrid algorithms for two-dimensional fractional Feynman–Kac equation. Journal of Scientific Computing, 74(2): 1034–1059, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0480-1.

Ghasemi:2018:TAN

[1998] Maryam Ghasemi and Hermann J. Eberl. Time adaptive numerical solution of a highly degenerate diffusion–reaction biofilm model based on regularisation. Journal of Scientific Computing, 74(2):1060– 1090, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-017-0483-y.

Meng:2018:VCP

[1999] Tingting Meng and Lijun Yi. An h–p version of the continuous Petrov– Galerkin method for nonlinear delay differential equations. Journal of Scientific Computing, 74(2):1091–1114, February 2018. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Li:2018:BCF

[2000] Xiaoli Li and Hongxing Rui. Block-centered finite difference method for simulating compressible wormhole propagation. Journal of Scientific Computing, 74(2):1115–1145, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Zhao:2018:NUR

[2001] Ruixue Zhao and Jinyan Fan. On a new updating rule of the Levenberg– Marquardt parameter. Journal of Scientific Computing, 74(2):1146– 1162, February 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Ivorra:2018:ALN

[2002] Benjamin Ivorra. Application of the laminar Navier–Stokes equations for solving 2D and 4D pathfinding problems with static and dynamic spatial constraints: Implementation and validation in Comsol multiphysics.

Journal of Scientific Computing, 74(2):1163–1187, February 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

Svard:2018:RCS

[2003] Magnus Svärd and Jan Nordström. Response to “Convergence of Summation-by-Parts Finite Difference Methods for the Wave Equation”. Journal of Scientific Computing, 74(2):1188–1192, February 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). See [1798].

Addam:2018:SAW

[2004] Mohamed Addam, Abderrahman Bouhamidi, and Mohammed Heyouni. On solving an acoustic wave problem via frequency–domain approach and tensorial spline Galerkin method. Journal of Scientific Computing, 74(3):1193–1220, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0490-z;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0490-z.pdf.

Zhang:2018:GSC

[2005] Chengjian Zhang and Cui Li. Generalized Störmer-Cowell methods for nonlinear BVPs of second-order delay–integro-differential equations. Journal of Scientific Computing, 74(3):1221–1240, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0491-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0491-y.pdf.

Lai:2018:MBL

[2006] Rongjie Lai and Jia Li. Manifold based low-rank regularization for image restoration and semi-supervised learning. Journal of Scientific Computing, 74(3):1241–1263, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0492-x;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0492-x.pdf.

Deng:2018:LCC

[2007] Quanling Deng and Victor Ginting. Locally conservative continuous Galerkin FEM for pressure equation in two–phase flow model in subsurfaces. Journal of Scientific Computing, 74(3):1264–1285, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/

s10915-017-0493-9;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0493-9.pdf.

Chen:2018:LFT

[2008] Sheng Chen, Jie Shen, and Li-Lian Wang. Laguerre functions and their applications to tempered fractional differential equations on infinite intervals. Journal of Scientific Computing, 74(3):1286–1313, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0495-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0495-7.pdf.

Hao:2018:HMP

[2009] Wenrui Hao. A homotopy method for parameter estimation of nonlinear differential equations with multiple optima. Journal of Scientific Computing, 74(3):1314–1324, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0518-4;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0518-4.pdf.

Hong:2018:ESM

[2010] Youngjoon Hong and Chang-Yeol Jung. Enriched spectral method for stiff convection–dominated equations. Journal of Scientific Computing, 74(3):1325–1346, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0494-8;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0494-8.pdf.

Deng:2018:SCT

[2011] Xiao-Long Deng and Maojun Li. Simulating compressible two– medium flows with sharp–interface adaptive Runge–Kutta discontinuous Galerkin methods. Journal of Scientific Computing, 74(3):1347–1368, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0511-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0511-y.pdf.

Wang:2018:SSW

[2012] Junping Wang, Ruishu Wang, Qilong Zhai, and Ran Zhang. A systematic study on weak Galerkin finite element methods for second order elliptic problems. Journal of Scientific Computing, 74(3):1369–1396, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0496-6;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0496-6.pdf.

Guo:2018:LFE

[2013] Hailong Guo, Zhimin Zhang, and Qingsong Zou. A C^{0} linear finite element method for biharmonic problems. Journal of Scientific Computing, 74(3):1397–1422, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0501-0;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0501-0.pdf.

Cheung:2018:MCS

[2014] Siu Wun Cheung, Eric Chung, and Hyea Hyun Kim. A mass conservative scheme for fluid–structure interaction problems by the staggered discontinuous Galerkin method. Journal of Scientific Computing, 74(3):1423–1456, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0500-1;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0500-1.pdf.

Hu:2018:PCR

[2015] Jun Hu and Limin Ma. A penalized Crouzeix–Raviart element method for second order elliptic eigenvalue problems. Journal of Scientific Computing, 74(3):1457–1479, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0505-9;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0505-9.pdf.

Kolb:2018:TOH

[2016] Oliver Kolb. A third order hierarchical basis WENO interpolation for sparse grids with application to conservation laws with uncertain data. Journal of Scientific Computing, 74(3):1480–1503, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0503-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0503-y.pdf.

Jang:2018:NFI

[2017] T. S. Jang. A new functional iterative algorithm for the regularized long–wave equation using an integral equation formalism. Journal of Scientific Computing, 74(3):1504–1532, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0533-5;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0533-5.pdf.

Yang:2018:NAC

[2018] Xiaofeng Yang. Numerical approximations for the Cahn–Hilliard phase field model of the binary fluid–surfactant system. Journal of Scientific Computing, 74(3):1533–1553, March 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0508-6;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0508-6.pdf.

Huang:2018:SMS

[2019] Can Huang, Zhimin Zhang, and Qingshuo Song. Spectral methods for substantial fractional differential equations. Journal of Scientific Computing, 74(3):1554–1574, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0506-8;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0506-8.pdf.

Pareschi:2018:SPS

[2020] Lorenzo Pareschi and Mattia Zanella. Structure preserving schemes for nonlinear Fokker–Planck equations and applications. Journal of Scientific Computing, 74(3):1575–1600, March 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0510-z;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0510-z.pdf.

Latu:2018:FAI

[2021] Guillaume Latu, Michel Mehrenberger, Yaman Güçlü, Maurizio Ottaviani, and Eric Sonnendrücker. Field-aligned interpolation for semi-Lagrangian gyrokinetic simulations. Journal of Scientific Computing, 74(3):1601–1650, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0509-5;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0509-5.pdf.

Feng:2018:IPD

[2022] Xiaobing Feng, Michael Neilan, and Stefan Schnake. Interior penalty discontinuous Galerkin methods for second order linear non-divergence form elliptic PDEs. Journal of Scientific Computing, 74(3):1651–1676, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0519-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0519-3.pdf.

DiPietro:2018:HHO

[2023] Daniele A. Di Pietro and Stella Krell. A hybrid high-order method for the steady incompressible Navier–Stokes problem. Journal of Scientific Computing, 74(3):1677–1705, March 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0512-x;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0512-x.pdf.

Shen:2018:AWM

[2024] Jie Shen, Jie Xu, and Pingwen Zhang. Approximations on SO(3) by Wigner D-matrix and applications. Journal of Scientific Computing, 74 (3):1706–1724, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10915-017-0515-7.pdf.

Kim:2018:STA

[2025] Dongho Kim, Eun-Jae Park, and Boyoon Seo. Space-time adaptive methods for the mixed formulation of a linear parabolic problem. Journal of Scientific Computing, 74(3):1725–1756, March 2018. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-017-0514-8.pdf.

Sheshadri:2018:ASF

[2026] Abhishek Sheshadri and Antony Jameson. An analysis of stability of the flux reconstruction formulation on quadrilateral elements for the linear advection–diffusion equation. Journal of Scientific Computing, 74(3): 1757–1785, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10915-017-0513-9.pdf.

Liu:2018:TGB

[2027] Wei Liu and Jintao Cui. A two–grid block–centered finite difference algorithm for nonlinear compressible Darcy–Forchheimer model in porous media. Journal of Scientific Computing, 74(3):1786–1815, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0516-6.pdf.

Huang:2018:HRF

[2028] Wen-Feng Huang, Yu-Xin Ren, Qiuju Wang, and Xiong Jiang. High resolution finite volume scheme based on the quintic spline reconstruction on non-uniform grids. Journal of Scientific Computing, 74(3):1816– 1852, March 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0524-6.pdf.

Chang:2018:CSR

[2029] Jingya Chang, Weiyang Ding, Liqun Qi, and Hong Yan. Computing the p-spectral radii of uniform hypergraphs with applications. Journal of Scientific Computing, 75(1):1–25, April 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0520-x;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0520-x.pdf.

Abbasi:2018:CLS

[2030] Bilal Abbasi and Adam M. Oberman. Computing the level set convex hull. Journal of Scientific Computing, 75(1):26–42, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0522-8;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0522-8.pdf.

Gao:2018:MNN

[2031] Shangqi Gao and Qibin Fan. A mixture of nuclear norm and matrix factorization for tensor completion. Journal of Scientific Computing, 75(1): 43–64, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0521-9;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0521-9.pdf.

Cai:2018:FVH

[2032] Xiaofeng Cai, Jianxian Qiu, and Jingmei Qiu. Finite volume HWENO schemes for nonconvex conservation laws. Journal of Scientific Computing, 75(1):65–82, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0525-5;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0525-5.pdf.

DelReyFernandez:2018:SAT

[2033] David C. Del Rey Fernández, Jason E. Hicken, and David W. Zingg. Simultaneous approximation terms for multi-dimensional summationby-parts operators. Journal of Scientific Computing, 75(1):83–110, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0523-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0523-7.pdf.

Kyza:2018:HAG

[2034] Irene Kyza, Stephen Metcalfe, and Thomas P. Wihler. hp-adaptive Galerkin time stepping methods for nonlinear initial value problems. Journal of Scientific Computing, 75(1):111–127, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0565-x;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0565-x.pdf.

Xue:2018:RCL

[2035] Changfeng Xue and Shaozhong Deng. Recursive computation of logarithmic derivatives, ratios, and products of spheroidal harmonics and modified Bessel functions and applications. Journal of Scientific Computing, 75(1):128–156, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0527-3;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0527-3.pdf.

An:2018:ESG

[2036] Jing An and Zhimin Zhang. An efficient spectral–Galerkin approximation and error analysis for Maxwell transmission eigenvalue problems in spherical geometries. Journal of Scientific Computing, 75(1):157–181, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0528-2;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0528-2.pdf.

Li:2018:PME

[2037] Wuchen Li, Ernest K. Ryu, Stanley Osher, Wotao Yin, and Wilfrid Gangbo. A parallel method for Earth Mover’s distance. Journal of Scientific Computing, 75(1):182–197, April 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0529-1;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0529-1.pdf.

Dong:2018:IMA

[2038] Bo Dong, Matthew M. Lin, and Haesun Park. Integer matrix approximation and data mining. Journal of Scientific Computing, 75(1):198– 224, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0531-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0531-7.pdf.

Scarnati:2018:URI

[2039] Theresa Scarnati, Anne Gelb, and Rodrigo B. Platte. Using ℓ_{1} regularization to improve numerical partial differential equation solvers. Journal of Scientific Computing, 75(1):225–252, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0530-8;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0530-8.pdf.

Wang:2018:CDF

[2040] Lai Wang and Meilin Yu. Compact direct flux reconstruction for conservation laws. Journal of Scientific Computing, 75(1):253–275, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0535-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0535-3.pdf.

Karcher:2018:CRB

[2041] Mark Kärcher, Zoi Tokoutsi, Martin A. Grepl, and Karen Veroy. Certified reduced basis methods for parametrized elliptic optimal control problems with distributed controls. Journal of Scientific Computing, 75(1):276– 307, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0539-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0539-z.pdf.

Modesti:2018:ESI

[2042] Davide Modesti and Sergio Pirozzoli. An efficient semi-implicit solver for direct numerical simulation of compressible flows at all speeds. Journal of Scientific Computing, 75(1):308–331, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0534-4;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0534-4.pdf.

Jeong:2018:PMC

[2043] Darae Jeong and Junseok Kim. A projection method for the conservative discretizations of parabolic partial differential equations. Journal of Scientific Computing, 75(1):332–349, April 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0536-2;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0536-2.pdf.

Dubuis:2018:AAT

[2044] Samuel Dubuis and Marco Picasso. An adaptive algorithm for the time dependent transport equation with anisotropic finite elements and the Crank–Nicolson scheme. Journal of Scientific Computing, 75(1):350– 375, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0537-1;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0537-1.pdf.

Cockburn:2018:SPD

[2045] Bernardo Cockburn, Manuel A. Sánchez, and Chunguang Xiong. Supercloseness of primal–dual Galerkin approximations for second order elliptic problems. Journal of Scientific Computing, 75(1):376–394, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0538-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0538-0.pdf.

Higueras:2018:OBL

[2046] Inmaculada Higueras and Teo Roldán. Order barrier for low-storage DIRK methods with positive weights. Journal of Scientific Computing, 75(1):395–404, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0540-6;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0540-6.pdf.

Fatone:2018:IDD

[2047] Lorella Fatone and Daniele Funaro. Isospectral domains for discrete elliptic operators. Journal of Scientific Computing, 75(1):405–426, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0541-5;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0541-5.pdf.

Chen:2018:ESA

[2048] Huangxin Chen, Shuyu Sun, and Tao Zhang. Energy stability analysis of some fully discrete numerical schemes for incompressible Navier–Stokes equations on staggered grids. Journal of Scientific Computing, 75(1):427– 456, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0543-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0543-3.pdf.

Wang:2018:JPB

[2049] Haiyong Wang and Lun Zhang. Jacobi polynomials on the Bernstein ellipse. Journal of Scientific Computing, 75(1):457–477, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0542-4.pdf.

Na:2018:EMS

[2050] Hanwool Na, Myeongmin Kang, Miyoun Jung, and Myungjoo Kang. An exp model with spatially adaptive regularization parameters for multiplicative noise removal. Journal of Scientific Computing, 75(1):478– 509, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0550-4.pdf.

Sun:2018:SAH

[2051] Ming Sun, Jichun Li, Peizhen Wang, and Zhimin Zhang. Superconvergence analysis of high-order rectangular edge elements for time-harmonic Maxwell’s equations. Journal of Scientific Computing, 75(1):510–535, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0544-2.pdf.

Buli:2018:LDG

[2052] Joshua Buli and Yulong Xing. Local discontinuous Galerkin methods for the Boussinesq coupled BBM system. Journal of Scientific Computing, 75(1):536–559, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0546-0.pdf.

Giacomini:2018:EFA

[2053] Matteo Giacomini. An equilibrated fluxes approach to the certified descent algorithm for shape optimization using conforming finite element and discontinuous Galerkin discretizations. Journal of Scientific Computing, 75(1):560–595, April 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0545-1.pdf.

Cockburn:2018:SNH

[2054] Bernardo Cockburn, Zhixing Fu, Allan Hungria, Liangyue Ji, Manuel A. Sánchez, and Francisco-Javier Sayas. Stormer-Numerov HDG methods for acoustic waves. Journal of Scientific Computing, 75(2):597– 624, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0547-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0547-z.pdf.

Weynans:2018:SCM

[2055] Lisl Weynans. Super-convergence in maximum norm of the gradient for the Shortley–Weller method. Journal of Scientific Computing, 75(2):625– 637, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0548-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0548-y.pdf.

Shi:2018:GWN

[2056] Zuoqiang Shi, Stanley Osher, and Wei Zhu. Generalization of the weighted nonlocal Laplacian in low dimensional manifold model. Journal of Scientific Computing, 75(2):638–656, May 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0549-x;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0549-x.pdf.

Friedrich:2018:CSD

[2057] Lucas Friedrich, David C. Del Rey Fernández, Andrew R. Winters, Gregor J. Gassner, David W. Zingg, and Jason Hicken. Conservative and stable degree preserving SBP operators for nonconforming meshes. Journal of Scientific Computing, 75(2):657–686, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0563-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0563-z.pdf.

Zhang:2018:OSI

[2058] Cheng Zhang, Jingfang Huang, Cheng Wang, and Xingye Yue. On the operator splitting and integral equation preconditioned deferred correction methods for the “good” Boussinesq equation. Journal of Scientific Computing, 75(2):687–712, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0552-2;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0552-2.pdf.

Wen:2018:SSN

[2059] You-Wei Wen, Wai-Ki Ching, and Michael Ng. A semi-smooth Newton method for inverse problem with uniform noise. Journal of Scientific Computing, 75(2):713–732, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0557-x;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0557-x.pdf.

Dang:2018:DRS

[2060] Duy-Minh Dang and Luis Ortiz-Gracia. A dimension reduction Shannonwavelet based method for option pricing. Journal of Scientific Computing, 75(2):733–761, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0556-y;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0556-y.pdf.

Rachh:2018:IEF

[2061] M. Rachh and T. Askham. Integral equation formulation of the biharmonic Dirichlet problem. Journal of Scientific Computing, 75(2):762– 781, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0559-8;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0559-8.pdf.

Mu:2018:WGM

[2062] Lin Mu, Junping Wang, and Xiu Ye. A weak Galerkin method for the Reissner–Mindlin plate in primary form. Journal of Scientific Computing, 75(2):782–802, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0564-y;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0564-y.pdf.

Lai:2018:BSM

[2063] Ming-Jun Lai and Chunmei Wang. A bivariate spline method for second order elliptic equations in non-divergence form. Journal of Scientific Computing, 75(2):803–829, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0562-0;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0562-0.pdf.

Schroeder:2018:DFD

[2064] Philipp W. Schroeder and Gert Lube. Divergence-free H (div)-FEM for time-dependent incompressible flows with applications to high Reynolds number vortex dynamics. Journal of Scientific Computing, 75(2):830– 858, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0561-1;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0561-1.pdf.

Nossek:2018:FGN

[2065] Raz Z. Nossek and Guy Gilboa. Flows generating nonlinear eigenfunctions. Journal of Scientific Computing, 75(2):859–888, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0577-6;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0577-6.pdf.

Jeon:2018:HSD

[2066] Youngmok Jeon. Hybrid spectral difference methods for elliptic equations on exterior domains with the discrete radial absorbing boundary condition. Journal of Scientific Computing, 75(2):889–905, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0570-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0570-0.pdf.

Eriksson:2018:DCF

[2067] Sofia Eriksson. A dual consistent finite difference method with narrow stencil second derivative operators. Journal of Scientific Computing, 75(2):906–940, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0569-6;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0569-6.pdf.

Verma:2018:HOF

[2068] Prabal Singh Verma and Wolf-Christian Müller. Higher order finite volume central schemes for multi-dimensional hyperbolic problems. Journal of Scientific Computing, 75(2):941–969, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0567-8;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0567-8.pdf.

Cai:2018:FOC

[2069] Haotao Cai and Yanping Chen. A fractional order collocation method for second kind Volterra integral equations with weakly singular kernels. Journal of Scientific Computing, 75(2):970–992, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0568-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0568-7.pdf.

Li:2018:UES

[2070] Hongwei Li, Lili Ju, Chenfei Zhang, and Qiujin Peng. Unconditionally energy stable linear schemes for the diffuse interface model with peng– Robinson equation of state. Journal of Scientific Computing, 75(2):993– 1015, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0576-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0576-7.pdf.

Don:2018:FIA

[2071] Wai Sun Don, Bao-Shan Wang, and Zhen Gao. Fast iterative adaptive multi-quadric radial basis function method for edges detection of piecewise functions-i: Uniform mesh. Journal of Scientific Computing, 75(2):1016–1039, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0572-y;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0572-y.pdf.

Ranocha:2018:SER

[2072] Hendrik Ranocha and Philipp Öffner. L_{2} stability of explicit Runge– Kutta schemes. Journal of Scientific Computing, 75(2):1040–1056, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/

s10915-017-0595-4;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0595-4.pdf.

Klinge:2018:SSP

[2073] Marcel Klinge and Rüdiger Weiner. Strong stability preserving explicit peer methods for discontinuous Galerkin discretizations. Journal of Scientific Computing, 75(2):1057–1078, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0573-x;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0573-x.pdf.

Chung:2018:GPE

[2074] Eric T. Chung, Eun-Jae Park, and Lina Zhao. Guaranteed a posteriori error estimates for a staggered discontinuous Galerkin method. Journal of Scientific Computing, 75(2):1079–1101, May 2018. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-017-0575-8.pdf.

Lin:2018:SCA

[2075] Xue lei Lin, Michael K. Ng, and Hai-Wei Sun. Stability and convergence analysis of finite difference schemes for time-dependent space–fractional diffusion equations with variable diffusion coefficients. Journal of Scientific Computing, 75(2):1102–1127, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-017-0581-x.pdf.

Sen:2018:ESS

[2076] Chhanda Sen and Harish Kumar. Entropy stable schemes for ten– moment Gaussian closure equations. Journal of Scientific Computing, 75(2):1128–1155, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10915-017-0579-4.pdf.

Long:2018:TBN

[2077] Xiaolong Long, Knut Solna, and Jack Xin. Three l_{1} based nonconvex methods in constructing sparse mean reverting portfolios. Journal of Scientific Computing, 75(2):1156–1186, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-017-0578-5.pdf.

San:2018:GDP

[2078] Omer San and Prakash Vedula. Generalized deconvolution procedure for structural modeling of turbulence. Journal of Scientific Computing, 75 (2):1187–1206, May 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10915-017-0583-8.pdf.

Zhao:2018:CAI

[2079] Tao Zhao. A convergence analysis of the inexact simplified Jacobi– Davidson algorithm for polynomial eigenvalue problems. Journal of Scientific Computing, 75(3):1207–1228, June 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0582-9;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0582-9.pdf.

Huang:2018:PTR

[2080] Rong Huang. A periodic qd-type reduction for computing eigenvalues of structured matrix products to high relative accuracy. Journal of Scientific Computing, 75(3):1229–1261, June 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0584-7;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0584-7.pdf.

Manzanero:2018:IAD

[2081] Juan Manzanero, Gonzalo Rubio, Esteban Ferrer, Eusebio Valero, and David A. Kopriva. Insights on aliasing driven instabilities for advection equations with application to Gauss–Lobatto discontinuous Galerkin methods. Journal of Scientific Computing, 75(3):1262–1281, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0585-6;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0585-6.pdf.

Hamfeldt:2018:HOA

[2082] Brittany Froese Hamfeldt and Tiago Salvador. Higher-order adaptive finite difference methods for fully nonlinear elliptic equations. Journal of Scientific Computing, 75(3):1282–1306, June 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0586-5;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0586-5.pdf.

Ji:2018:NMS

[2083] Cui cui Ji, Weizhong Dai, and Zhi zhong Sun. Numerical method for solving the time-fractional dual–phase-lagging heat conduction equation with the temperature–jump boundary condition. Journal of Scientific Computing, 75(3):1307–1336, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0588-3;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0588-3.pdf.

Luo:2018:FAM

[2084] Yixiang Luo, Jie Xu, and Pingwen Zhang. A fast algorithm for the moments of Bingham distribution. Journal of Scientific Computing, 75(3):1337–1350, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0589-2;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0589-2.pdf.

Dahiya:2018:OLI

[2085] Daisy Dahiya and Maria Cameron. Ordered line integral methods for computing the quasi–potential. Journal of Scientific Computing, 75(3):1351–1384, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0590-9;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0590-9.pdf.

Yan:2018:IPS

[2086] Jianfeng Yan, Jared Crean, and Jason E. Hicken. Interior penalties for summation-by-parts discretizations of linear second-order differential equations. Journal of Scientific Computing, 75(3):1385–1414, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0591-8;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0591-8.pdf.

Zhang:2018:MLM

[2087] Shuo Zhang, Yingxia Xi, and Xia Ji. A multi-level mixed element method for the eigenvalue problem of biharmonic equation. Journal of Scientific Computing, 75(3):1415–1444, June 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0592-7;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0592-7.pdf.

Du:2018:PPU

[2088] Guangzhi Du and Liyun Zuo. A parallel partition of unity scheme based on two–grid discretizations for the Navier–Stokes problem. Journal of Scientific Computing, 75(3):1445–1462, June 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0593-6;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0593-6.pdf.

Carrillo:2018:LSS

[2089] José A. Carrillo, Bertram Düring, Daniel Matthes, and David S. Mc-Cormick. A Lagrangian scheme for the solution of nonlinear diffusion equations using moving simplex meshes. Journal of Scientific Computing, 75(3):1463–1499, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0594-5;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0594-5.pdf.

Amat:2018:NAF

[2090] Sergio Amat, Zhilin Li, and Juan Ruiz. On an new algorithm for function approximation with full accuracy in the presence of discontinuities based on the immersed interface method. Journal of Scientific Computing, 75(3):1500–1534, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0596-3;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0596-3.pdf.

Zhang:2018:FAD

[2091] Xiongjun Zhang, Michael K. Ng, and Minru Bai. A fast algorithm for deconvolution and Poisson noise removal. Journal of Scientific Computing, 75(3):1535–1554, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0597-2;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0597-2.pdf.

Safdari-Vaighani:2018:RBF

[2092] Ali Safdari-Vaighani, Elisabeth Larsson, and Alfa Heryudono. Radial basis function methods for the Rosenau equation and other higher order PDEs. Journal of Scientific Computing, 75(3):1555–1580, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0598-1;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0598-1.pdf.

Li:2018:COT

[2093] Wuchen Li, Penghang Yin, and Stanley Osher. Computations of optimal transport distance with Fisher information regularization. Journal of Scientific Computing, 75(3):1581–1595, June 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0599-0;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0599-0.pdf.

Ryu:2018:UPM

[2094] Ernest K. Ryu, Wuchen Li, Penghang Yin, and Stanley Osher. Unbalanced and partial L_{1} Monge–Kantorovich problem: a scalable parallel first-order method. Journal of Scientific Computing, 75(3):1596– 1613, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0600-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0600-y.pdf.

Bentbib:2018:CGT

[2095] A. H. Bentbib, K. Jbilou, and Y. Kaouane. A computational global tangential Krylov subspace method for model reduction of large–scale MIMO dynamical systems. Journal of Scientific Computing, 75(3):1614– 1632, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0601-x;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0601-x.pdf.

Lei:2018:FPP

[2096] Siu-Long Lei, Wenfei Wang, Xu Chen, and Deng Ding. A fast preconditioned penalty method for American options pricing under regime–switching tempered fractional diffusion models. Journal of Scientific Computing, 75(3):1633–1655, June 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0602-9;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0602-9.pdf.

Ma:2018:CAI

[2097] Jingtang Ma and Zhiqiang Zhou. Convergence analysis of iterative Laplace transform methods for the coupled PDEs from regime–switching option pricing. Journal of Scientific Computing, 75(3):1656–1674, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0604-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0604-7.pdf.

Ha:2018:SOW

[2098] Youngsoo Ha, Chang Ho Kim, Hyoseon Yang, and Jungho Yoon. A sixthorder weighted essentially non-oscillatory schemes based on exponential polynomials for Hamilton–Jacobi equations. Journal of Scientific Computing, 75(3):1675–1700, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0603-8;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0603-8.pdf.

Delgadillo:2018:FGA

[2099] Ricardo Delgadillo, Xu Yang, and Jiwei Zhang. Frozen Gaussian approximation-based artificial boundary conditions for one-dimensional nonlinear Schrödinger equation in the semiclassical regime. Journal of Scientific Computing, 75(3):1701–1720, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-017-0606-5.pdf.

Jang:2018:RIE

[2100] T. S. Jang. A regular integral equation formalism for solving the standard Boussinesq’s equations for variable water depth. Journal of Scientific Computing, 75(3):1721–1756, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0605-6.pdf.

Abdi:2018:LBR

[2101] Ali Abdi, Jean-Paul Berrut, and Seyyed Ahmad Hosseini. The linear barycentric rational method for a class of delay Volterra integro– differential equations. Journal of Scientific Computing, 75(3):1757– 1775, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0608-3.pdf.

Lemaire:2018:ESJ

[2102] Vincent Lemaire, Michèle Thieullen, and Nicolas Thomas. Exact simulation of the jump times of a class of piecewise deterministic Markov processes. Journal of Scientific Computing, 75(3):1776– 1807, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0607-4.pdf.

Xu:2018:ITO

[2103] Weizheng Xu and Weiguo Wu. An improved third-order WENO-Z scheme. Journal of Scientific Computing, 75(3):1808–1841, June 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-017-0587-4.pdf.

Sharbatdar:2018:ABF

[2104] Mahkame Sharbatdar and Carl Ollivier-Gooch. Adjoint-based functional correction for unstructured mesh finite volume methods. Journal of Scientific Computing, 76(1):1–23, July 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0611-8;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0611-8.pdf.

Kornelus:2018:FCH

[2105] Adeline Kornelus and Daniel Appelö. Flux-conservative Hermite methods for simulation of nonlinear conservation laws. Journal of Scientific Computing, 76(1):24–47, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0613-6;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0613-6.pdf.

Di:2018:AMS

[2106] Yana Di, Hehu Xie, and Xiaobo Yin. Anisotropic meshes and stabilization parameter design of linear SUPG method for 2D convection– dominated convection–diffusion equations. Journal of Scientific Computing, 76(1):48–68, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0610-9;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0610-9.pdf.

Lin:2018:GCU

[2107] Tianyi Lin, Shiqian Ma, and Shuzhong Zhang. Global convergence of unmodified 3-block ADMM for a class of convex minimization problems. Journal of Scientific Computing, 76(1):69–88, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0612-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0612-7.pdf.

Aletti:2018:HRA

[2108] Matteo C. Aletti, Simona Perotto, and Alessandro Veneziani. Hi-Mod reduction of advection–diffusion–reaction problems with general boundary conditions. Journal of Scientific Computing, 76(1):89–119, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0614-5;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0614-5.pdf.

You:2018:ELI

[2109] Guoqiao You, Renkun Shi, and Yuhua Xu. An efficient Lagrangian interpolation scheme for computing flow maps and line integrals using discrete velocity data. Journal of Scientific Computing, 76(1):120– 144, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0620-7;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0620-7.pdf.

Gross:2018:SNE

[2110] B. Gross and P. J. Atzberger. Spectral numerical exterior calculus methods for differential equations on radial manifolds. Journal of Scientific Computing, 76(1):145–165, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-017-0617-2;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0617-2.pdf.

Wang:2018:SCM

[2111] Chuanli Wang, Zhongqing Wang, and Lilian Wang. A spectral collocation method for nonlinear fractional boundary value problems with a Caputo derivative. Journal of Scientific Computing, 76(1):166–188, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0616-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0616-3.pdf.

Nicponski:2018:NHP

[2112] John Nicponski and Jae-Hun Jung. A note on high–precision approximation of asymptotically decaying solution and orthogonal decomposition. Journal of Scientific Computing, 76(1):189–215, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0619-0;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0619-0.pdf.

Ranocha:2018:CSE

[2113] Hendrik Ranocha. Comparison of some entropy conservative numerical fluxes for the Euler equations. Journal of Scientific Computing, 76(1):216–242, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0618-1;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0618-1.pdf.

Tian:2018:CAP

[2114] WenYi Tian and Xiaoming Yuan. Convergence analysis of primal– dual based methods for total variation minimization with finite element approximation. Journal of Scientific Computing, 76(1):243–274, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0623-4;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0623-4.pdf.

Sunny:2018:MIT

[2115] Linia Anie Sunny and V. Antony Vijesh. A monotone iterative technique for nonlinear fourth order elliptic equations with nonlocal boundary conditions. Journal of Scientific Computing, 76(1):275–298, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0615-4;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0615-4.pdf.

Hannah:2018:UDA

[2116] Robert Hannah and Wotao Yin. On unbounded delays in asynchronous parallel fixed–point algorithms. Journal of Scientific Computing, 76(1):299–326, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0628-z;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0628-z.pdf.

Gao:2018:IAV

[2117] Xiang Gao, Bo Jiang, and Shuzhong Zhang. On the information– adaptive variants of the ADMM: An iteration complexity perspective. Journal of Scientific Computing, 76(1):327–363, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0621-6;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0621-6.pdf.

Xiao:2018:RSS

[2118] Xiantao Xiao, Yongfeng Li, Zaiwen Wen, and Liwei Zhang. A regularized semi-smooth Newton method with projection steps for composite convex programs. Journal of Scientific Computing, 76(1):364– 389, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0624-3;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0624-3.pdf.

Liang:2018:CMG

[2119] Hui Liang and Martin Stynes. Collocation methods for general Caputo two–point boundary value problems. Journal of Scientific Computing, 76(1):390–425, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0622-5;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0622-5.pdf.

Kopriva:2018:SOM

[2120] David A. Kopriva. Stability of overintegration methods for nodal discontinuous Galerkin spectral element methods. Journal of Scientific Computing, 76(1):426–442, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-017-0626-1;   https://link.springer.
com/content/pdf/10.1007/s10915-017-0626-1.pdf.

Reyes:2018:NCH

[2121] Adam Reyes, Dongwook Lee, Carlo Graziani, and Petros Tzeferacos. A new class of high-order methods for fluid dynamics simulations using Gaussian process modeling: One-dimensional case. Journal of Scientific Computing, 76(1):443–480, July 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0625-2;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0625-2.pdf.

Wang:2018:HOD

[2122] Yifan Wang, Annalisa Quaini, and Suncica Canić. A higher-order discontinuous Galerkin/arbitrary Lagrangian Eulerian partitioned approach to solving fluid–structure interaction problems with incompressible, viscous fluids and elastic structures. Journal of Scientific Computing, 76(1):481– 520, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0629-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0629-y.pdf.

Lozano:2018:EPE

[2123] Carlos Lozano. Entropy production by explicit Runge–Kutta schemes. Journal of Scientific Computing, 76(1):521–564, July 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0627-0;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0627-0.pdf.

Qiao:2018:CAD

[2124] Leijie Qiao and Da Xu. Compact alternating direction implicit scheme for integro–differential equations of parabolic type. Journal of Scientific Computing, 76(1):565–582, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0630-5.pdf.

Gracia:2018:FSC

[2125] J. L. Gracia, E. O’Riordan, and M. Stynes. A fitted scheme for a Caputo initial–boundary value problem. Journal of Scientific Computing, 76(1): 583–609, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-017-0631-4.pdf.

Zhao:2018:MTV

[2126] Jikun Zhao, Bei Zhang, Shaochun Chen, and Shipeng Mao. The Morley– type virtual element for plate bending problems. Journal of Scientific Computing, 76(1):610–629, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0632-3.pdf.

Antonana:2018:NIM

[2127] Mikel Antoñana, Joseba Makazaga, and Ander Murua. New integration methods for perturbed ODEs based on symplectic implicit Runge–Kutta schemes with application to solar system simulations. Journal of Scientific Computing, 76(1):630–650, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-017-0634-1.pdf.

Aceto:2018:EIR

[2128] Lidia Aceto and Paolo Novati. Efficient implementation of rational approximations to fractional differential operators. Journal of Scientific Computing, 76(1):651–671, July 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-017-0633-2.pdf.

Chen:2018:MMH

[2129] Long Chen, Jun Hu, and Xuehai Huang. Multigrid methods for hellan–herrmann-Johnson mixed method of Kirchhoff plate bending problems. Journal of Scientific Computing, 76(2):673–696, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0636-z;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0636-z.pdf.

Melis:2018:TPI

[2130] Ward Melis and Giovanni Samaey. Telescopic projective integration for linear kinetic equations with multiple relaxation times. Journal of Scientific Computing, 76(2):697–726, August 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0635-0;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0635-0.pdf.

Yoon:2018:SCH

[2131] Gangjoon Yoon, Chohong Min, and Seick Kim. A stable and convergent Hodge decomposition method for fluid–solid interaction. Journal of Scientific Computing, 76(2):727–758, August 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-017-
0638-x;   https://link.springer.com/content/pdf/10.1007/s10915-
017-0638-x.pdf.

Meng:2018:CNB

[2132] Jie Meng, Sang-Hyup Seo, and Hyun-Min Kim. Condition numbers and backward error of a matrix polynomial equation arising in stochastic models. Journal of Scientific Computing, 76(2):759–776, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0641-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0641-x.pdf.

Britt:2018:HOC

[2133] Steven Britt, Eli Turkel, and Semyon Tsynkov. A high order compact time/space finite difference scheme for the wave equation with variable speed of sound. Journal of Scientific Computing, 76(2):777–811, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0639-9;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0639-9.pdf.

Ludvigsson:2018:HON

[2134] Gustav Ludvigsson, Kyle R. Steffen, Simon Sticko, Siyang Wang, Qing Xia, Yekaterina Epshteyn, and Gunilla Kreiss. High-order numerical methods for 2D parabolic problems in single and composite domains. Journal of Scientific Computing, 76(2):812–847, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0637-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0637-y.pdf.

Li:2018:UOE

[2135] Dongfang Li, Jiwei Zhang, and Zhimin Zhang. Unconditionally optimal error estimates of a linearized Galerkin method for nonlinear time fractional reaction–subdiffusion equations. Journal of Scientific Computing, 76(2):848–866, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0642-9;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0642-9.pdf.

Chen:2018:HOA

[2136] Minghua Chen and Weihua Deng. High order algorithm for the timetempered fractional Feynman–Kac equation. Journal of Scientific Computing, 76(2):867–887, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0640-y;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0640-y.pdf.

Jing:2018:DGM

[2137] Feifei Jing, Weimin Han, Wenjing Yan, and Fei Wang. Discontinuous Galerkin methods for a stationary Navier–Stokes problem with a nonlinear slip boundary condition of friction type. Journal of Scientific Computing, 76(2):888–912, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0644-7;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0644-7.pdf.

Liu:2018:FDG

[2138] Huan Liu, Aijie Cheng, and Hong Wang. A fast discontinuous Galerkin method for a bond–based linear peridynamic model discretized on a locally refined composite mesh. Journal of Scientific Computing, 76(2):913–942, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0645-6;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0645-6.pdf.

Califano:2018:SSP

[2139] Giovanna Califano, Giuseppe Izzo, and Zdzislaw Jackiewicz. Strong stability preserving general linear methods with Runge–Kutta stability. Journal of Scientific Computing, 76(2):943–968, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0646-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0646-5.pdf.

Schoeder:2018:AHO

[2140] Svenja Schoeder, Martin Kronbichler, and Wolfgang A. Wall. Arbitrary high-order explicit hybridizable discontinuous Galerkin methods for the acoustic wave equation. Journal of Scientific Computing, 76(2):969–1006, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0649-2;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0649-2.pdf.

Wang:2018:HOC

[2141] Yuan-Ming Wang and Lei Ren. High-order compact difference methods for Caputo–type variable coefficient fractional sub-diffusion equations in conservative form. Journal of Scientific Computing, 76(2):1007–1043, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0647-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0647-4.pdf.

Oikawa:2018:HMO

[2142] Issei Oikawa. An HDG method with orthogonal projections in facet integrals. Journal of Scientific Computing, 76(2):1044–1054, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0648-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0648-3.pdf.

Cao:2018:MFV

[2143] Fujun Cao, Zhiqiang Sheng, and Guangwei Yuan. Monotone finite volume schemes for diffusion equation with imperfect interface on distorted meshes. Journal of Scientific Computing, 76(2):1055–1077, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0651-8;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0651-8.pdf.

Wang:2018:GSA

[2144] Weina Wang, Chunlin Wu, and Jiansong Deng. A general selective averaging method for piecewise constant signal and image processing. Journal of Scientific Computing, 76(2):1078–1104, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0650-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0650-9.pdf.

Zhong:2018:GMS

[2145] Xinghui Zhong and Qin Li. Galerkin methods for stationary radiative transfer equations with uncertain coefficients. Journal of Scientific Computing, 76(2):1105–1126, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0652-7;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0652-7.pdf.

Chen:2018:FFF

[2146] Xinjuan Chen, Jae-Hun Jung, and Anne Gelb. Finite Fourier frame approximation using the inverse polynomial reconstruction method. Journal of Scientific Computing, 76(2):1127–1147, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0655-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0655-4.pdf.

Guo:2018:ALD

[2147] Ruihan Guo and Francis Filbet. A p-adaptive local discontinuous Galerkin level set method for Willmore flow. Journal of Scientific Computing, 76(2):1148–1167, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0656-3;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0656-3.pdf.

Sun:2018:MCS

[2148] Zhengjie Sun and Zongmin Wu. Meshless conservative scheme for multivariate nonlinear Hamiltonian PDEs. Journal of Scientific Computing, 76(2):1168–1187, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0658-1;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0658-1.pdf.

Huang:2018:CGL

[2149] Yufang Huang, Jianfeng Lu, and Pingbing Ming. A concurrent global– local numerical method for multiscale PDEs. Journal of Scientific Computing, 76(2):1188–1215, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0662-5.pdf.

Shiue:2018:CMS

[2150] Ming-Cheng Shiue, Kian Chuan Ong, and Ming-Chih Lai. Convergence of the MAC scheme for the Stokes/Darcy coupling problem. Journal of Scientific Computing, 76(2):1216–1251, August 2018. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-018-0660-7.pdf.

Vong:2018:UCM

[2151] Seakweng Vong and Pin Lyu. Unconditional convergence in maximum– norm of a second-order linearized scheme for a time-fractional Burgers– type equation. Journal of Scientific Computing, 76(2):1252–1273, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0659-0.pdf.

Wang:2018:WGF

[2152] Gang Wang, Yinnian He, and Jinjin Yang. Weak Galerkin finite element methods for the simulation of single–phase flow in fractured porous media. Journal of Scientific Computing, 76(2):1274–1300, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0673-2.pdf.

Ma:2018:HSB

[2153] Zhan Ma and Song-Ping Wu. HWENO schemes based on compact differencefor hyperbolic conservation laws. Journal of Scientific Computing, 76(2):1301–1325, August 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0663-4.pdf.

Nordstrom:2018:LTE

[2154] Jan Nordström and Hannes Frenander. On long time error bounds for the wave equation on second order form. Journal of Scientific Computing, 76(3):1327–1336, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0667-0;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0667-0.pdf.

Higueras:2018:OMP

[2155] Inmaculada Higueras, David I. Ketcheson, and Tihamér A. Kocsis. Optimal monotonicity–preserving perturbations of a given Runge–Kutta method. Journal of Scientific Computing, 76(3):1337–1369, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0664-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0664-3.pdf.

Nicholls:2018:HOP

[2156] David P. Nicholls and Xin Tong. High-order perturbation of surfaces algorithms for the simulation of localized surface plasmon resonances in two dimensions. Journal of Scientific Computing, 76(3):1370–1395, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0665-2;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0665-2.pdf.

Yin:2018:LFT

[2157] Penghang Yin, Jack Xin, and Yingyong Qi. Linear feature transform and enhancement of classification on deep neural network. Journal of Scientific Computing, 76(3):1396–1406, September 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0666-1;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0666-1.pdf.

You:2018:IEA

[2158] Guoqiao You and Shingyu Leung. An improved Eulerian approach for the finite time Lyapunov exponent. Journal of Scientific Computing, 76 (3):1407–1435, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0669-y;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0669-y.pdf.

Hu:2018:SHM

[2159] Weiwei Hu, Jiguang Shen, John R. Singler, Yangwen Zhang, and Xiaobo Zheng. A superconvergent HDG method for distributed control of convection diffusion PDEs. Journal of Scientific Computing, 76 (3):1436–1457, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0668-z;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0668-z.pdf.

Rong:2018:NAA

[2160] Yao Rong, William Layton, and Haiyun Zhao. Numerical analysis of an artificial compression method for magnetohydrodynamic flows at low magnetic Reynolds numbers. Journal of Scientific Computing, 76 (3):1458–1483, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0670-5;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0670-5.pdf.

Rhebergen:2018:HDG

[2161] Sander Rhebergen and Garth N. Wells. A hybridizable discontinuous Galerkin method for the Navier–Stokes equations with pointwise divergence–free velocity field. Journal of Scientific Computing, 76 (3):1484–1501, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0671-4;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0671-4.pdf.

Yang:2018:WOS

[2162] Xuehua Yang, Haixiang Zhang, and Da Xu. WSGD-OSC scheme for two-dimensional distributed order fractional reaction–diffusion equation. Journal of Scientific Computing, 76(3):1502–1520, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0672-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0672-3.pdf.

Singh:2018:NUC

[2163] Vishal Singh, Hossain Chizari, and Farzad Ismail. Non-unified compact residual–distribution methods for scalar advection–diffusion problems. Journal of Scientific Computing, 76(3):1521–1546, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0674-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0674-1.pdf.

Du:2018:DGM

[2164] Jie Du, Eric T. Chung, Ming Fai Lam, and Xiao-Ping Wang. Discontinuous Galerkin method with staggered hybridization for a class of nonlinear Stokes equations. Journal of Scientific Computing, 76 (3):1547–1577, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0676-z;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0676-z.pdf.

Bowman:2018:PFF

[2165] John C. Bowman and Zayd Ghoggali. The partial fast Fourier transform. Journal of Scientific Computing, 76(3):1578–1593, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0675-0;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0675-0.pdf.

Jiang:2018:FDF

[2166] Ying Jiang, Bo Wang, and Yuesheng Xu. A fully discrete fast Fourier– Galerkin method solving a boundary integral equation for the biharmonic equation. Journal of Scientific Computing, 76(3):1594–1632, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0688-8;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0688-8.pdf.

Zhang:2018:SCI

[2167] Jin Zhang and Xiaowei Liu. Supercloseness of continuous interior penalty methods on Shishkin triangular meshes and hybrid meshes for singularly perturbed problems with characteristic layers. Journal of Scientific Computing, 76(3):1633–1656, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0677-y;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0677-y.pdf.

Miyashita:2018:HDG

[2168] Masaru Miyashita and Norikazu Saito. Hybridized discontinuous Galerkin method for elliptic interface problems: Error estimates under low regularity assumptions of solutions. Journal of Scientific Computing, 76(3):1657–1673, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0678-x;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0678-x.pdf.

Askham:2018:SSV

[2169] T. Askham. A stabilized separation of variables method for the modified biharmonic equation. Journal of Scientific Computing, 76(3):1674–1697, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0679-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0679-9.pdf.

Yan:2018:NPD

[2170] Ming Yan. A new primal–dual algorithm for minimizing the sum of three functions with a linear operator. Journal of Scientific Computing, 76(3):1698–1717, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0680-3;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0680-3.pdf.

He:2018:GQC

[2171] Hongjin He, Chen Ling, Liqun Qi, and Guanglu Zhou. A globally and quadratically convergent algorithm for solving multilinear systems with M-tensors. Journal of Scientific Computing, 76(3):1718–1741, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0689-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0689-7.pdf.

Kozdon:2018:ESA

[2172] Jeremy E. Kozdon and Lucas C. Wilcox. An energy stable approach for discretizing hyperbolic equations with nonconforming discontinuous Galerkin methods. Journal of Scientific Computing, 76(3):1742–1784, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0682-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0682-1.pdf.

Feng:2018:HAF

[2173] Renzhong Feng and Junna Duan. High accurate finite differences based on RBF interpolation and its application in solving differential equations. Journal of Scientific Computing, 76(3):1785–1812, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0684-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0684-z.pdf.

Zhou:2018:AFE

[2174] Weiqi Zhou and Haijun Wu. An adaptive finite element method for the diffraction grating problem with PML and few–mode DtN truncations. Journal of Scientific Computing, 76(3):1813–1838, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-018-0683-0.pdf.

Chen:2018:MMF

[2175] Yangang Chen, Justin W. L. Wan, and Jessey Lin. Monotone mixed finite difference scheme for Monge–Ampère equation. Journal of Scientific Computing, 76(3):1839–1867, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-018-0685-y.pdf.

Baccouch:2018:AEP

[2176] Mahboub Baccouch. Asymptotically exact posteriori error estimates for the local discontinuous Galerkin method applied to nonlinear convection– diffusion problems. Journal of Scientific Computing, 76(3):1868–1904, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0687-9.pdf.

Li:2018:SOE

[2177] Weijia Li, Wenbin Chen, Cheng Wang, Yue Yan, and Ruijian He. A second order energy stable linear scheme for a thin film model without slope selection. Journal of Scientific Computing, 76(3):1905–1937, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0693-y.pdf.

Feng:2018:USE

[2178] Wenqiang Feng, Zhen Guan, John Lowengrub, Cheng Wang, Steven M. Wise, and Ying Chen. A uniquely solvable, energy stable numerical scheme for the functionalized Cahn–Hilliard equation and its convergence analysis. Journal of Scientific Computing, 76(3):1938–1967, September 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0690-1.pdf.

Macias-Diaz:2018:NED

[2179] Jorge E. Macías-Díaz. A numerically efficient dissipation–preserving implicit method for a nonlinear multidimensional fractional wave equation. Journal of Scientific Computing, 77(1):1–26, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0692-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0692-z.pdf.

Fan:2018:UMF

[2180] Wenping Fan, Xiaoyun Jiang, Fawang Liu, and Vo Anh. The unstructured mesh finite element method for the two-dimensional multiterm time–space fractional diffusion–wave equation on an irregular convex domain. Journal of Scientific Computing, 77(1):27–52, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0694-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0694-x.pdf.

Caboussat:2018:LSR

[2181] Alexandre Caboussat, Roland Glowinski, and Dimitrios Gourzoulidis. A least-squares/relaxation method for the numerical solution of the threedimensional elliptic Monge–Ampère equation. Journal of Scientific Computing, 77(1):53–78, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0698-6;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0698-6.pdf.

Chen:2018:EAM

[2182] Yongxin Chen, Eldad Haber, Kaoru Yamamoto, Tryphon T. Georgiou, and Allen Tannenbaum. An efficient algorithm for matrix-valued and vector-valued optimal mass transport. Journal of Scientific Computing, 77(1):79–100, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0696-8;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0696-8.pdf.

Chen:2018:MPM

[2183] Long Chen, Yongke Wu, Lin Zhong, and Jie Zhou. MultiGrid preconditioners for mixed finite element methods of the vector Laplacian. Journal of Scientific Computing, 77(1):101–128, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0697-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0697-7.pdf.

Garrappa:2018:CMM

[2184] Roberto Garrappa and Marina Popolizio. Computing the matrix Mittag-Leffler function with applications to fractional calculus. Journal of Scientific Computing, 77(1):129–153, October 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0699-5;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0699-5.pdf.

Gassner:2018:BSS

[2185] Gregor J. Gassner, Andrew R. Winters, Florian J. Hindenlang, and David A. Kopriva. The BR1 scheme is stable for the compressible Navier– Stokes equations. Journal of Scientific Computing, 77(1):154–200, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0702-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0702-1.pdf.

Gassner:2018:CBS

[2186] Gregor J. Gassner, Andrew R. Winters, Florian J. Hindenlang, and David A. Kopriva. Correction to: The BR1 scheme is stable for the compressible Navier–Stokes equations. Journal of Scientific Computing, 77(1):201–203, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0758-y;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0758-y.pdf.

Antil:2018:ORO

[2187] Harbir Antil, Enrique Otárola, and Abner J. Salgado. Optimization with respect to order in a fractional diffusion model: Analysis, approximation and algorithmic aspects. Journal of Scientific Computing, 77(1):204–224, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0703-0;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0703-0.pdf.

Zhou:2018:SCA

[2188] Han Zhou and Paul Andries Zegeling. Stability and convergence analysis of a class of continuous piecewise polynomial approximations for time-fractional differential equations. Journal of Scientific Computing, 77(1):225–262, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0704-z;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0704-z.pdf.

Xiong:2018:HOM

[2189] Tao Xiong, Giovanni Russo, and Jing-Mei Qiu. High order multidimensional characteristics tracing for the incompressible Euler equation and the guiding–center Vlasov equation. Journal of Scientific Computing, 77(1):263–282, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0705-y;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0705-y.pdf.

Zeng:2018:SFT

[2190] Fanhai Zeng, Ian Turner, and Kevin Burrage. A stable fast time-stepping method for fractional integral and derivative operators. Journal of Scientific Computing, 77(1):283–307, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0707-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0707-9.pdf.

Borthagaray:2018:FEA

[2191] Juan Pablo Borthagaray, Leandro M. Del Pezzo, and Sandra Martínez. Finite element approximation for the fractional eigenvalue problem. Journal of Scientific Computing, 77(1):308–329, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0710-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0710-1.pdf.

Bokil:2018:HSO

[2192] Vrushali A. Bokil, Yingda Cheng, Yan Jiang, Fengyan Li, and Puttha Sakkaplangkul. High spatial order energy stable FDTD methods for Maxwell’s equations in nonlinear optical media in one dimension. Journal of Scientific Computing, 77(1):330–371, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0716-8;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0716-8.pdf.

Geevers:2018:DPE

[2193] S. Geevers, W. A. Mulder, and J. J. W. van der Vegt. Dispersion properties of explicit finite element methods for wave propagation modelling on tetrahedral meshes. Journal of Scientific Computing, 77(1):372–396, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0709-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0709-7.pdf.

Zala:2018:CMA

[2194] Vidhi Zala, Varun Shankar, Shankar P. Sastry, and Robert M. Kirby. Curvilinear mesh adaptation using radial basis function interpolation and smoothing. Journal of Scientific Computing, 77(1):397–418, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0711-0;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0711-0.pdf.

Bao:2018:RNM

[2195] Gang Bao, Yanzhao Cao, Yongle Hao, and Kai Zhang. A robust numerical method for the random interface grating problem via shape calculus, weak Galerkin method, and low-rank approximation. Journal of Scientific Computing, 77(1):419–442, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0712-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0712-z.pdf.

Ainsworth:2018:FCP

[2196] Mark Ainsworth and Guosheng Fu. Fully computable a posteriori error bounds for hybridizable discontinuous Galerkin finite element approximations. Journal of Scientific Computing, 77(1):443–466, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0715-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0715-9.pdf.

Liu:2018:MDG

[2197] Hailiang Liu and Peimeng Yin. A mixed discontinuous Galerkin method without interior penalty for time-dependent fourth order problems. Journal of Scientific Computing, 77(1):467–501, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0756-0;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0756-0.pdf.

Yang:2018:HSD

[2198] Hee Jun Yang and Hyea Hyun Kim. A hybrid staggered discontinuous Galerkin method for KdV equations. Journal of Scientific Computing, 77(1):502–523, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0714-x;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0714-x.pdf.

Ge:2018:SGS

[2199] Liang Ge, Lianhai Wang, and Yanzhen Chang. A sparse grid stochastic collocation upwind finite volume element method for the constrained optimal control problem governed by random convection diffusion equations. Journal of Scientific Computing, 77(1):524–551, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0713-y.pdf.

Wu:2018:HOL

[2200] Jie Wu, Yuan yuan He, Guo hao Ding, and Yi yu Han. Hybrid optimized low-dissipation and adaptive MUSCL reconstruction technique for hyperbolic conservation laws. Journal of Scientific Computing, 77(1):552– 578, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0717-7.pdf.

Wissink:2018:SRS

[2201] B. W. Wissink, G. B. Jacobs, J. K. Ryan, W. S. Don, and E. T. A. van der Weide. Shock regularization with smoothness–increasing accuracy– conserving Dirac–delta polynomial kernels. Journal of Scientific Computing, 77(1):579–596, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0719-5.pdf.

Meng:2018:PQP

[2202] Tingwei Meng and Lok Ming Lui. PCBC: Quasiconformality of point cloud mappings. Journal of Scientific Computing, 77(1):597– 633, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0724-8.pdf.

Zouraris:2018:LIF

[2203] Georgios E. Zouraris. A linear implicit finite difference discretization of the Schrödinger–Hirota equation. Journal of Scientific Computing, 77 (1):634–656, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL http://link.springer.com/content/
pdf/10.1007/s10915-018-0718-6.pdf.

Ma:2018:ESM

[2204] Suna Ma, Huiyuan Li, and Zhimin Zhang. Efficient spectral methods for some singular eigenvalue problems. Journal of Scientific Computing, 77(1):657–688, October 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0721-y.pdf.

Friedrich:2018:ESN

[2205] Lucas Friedrich, Andrew R. Winters, David C. Del Rey Fernández, Gregor J. Gassner, Matteo Parsani, and Mark H. Carpenter. An entropy stable h/p non-conforming discontinuous Galerkin method with the summation-by-parts property. Journal of Scientific Computing, 77(2):689–725, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0733-7;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0733-7.pdf.

Dong:2018:NME

[2206] Haitao Dong and Fujun Liu. Numerical methods for Euler equations with self-similar and quasi self-similar solutions. Journal of Scientific Computing, 77(2):726–754, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0720-z;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0720-z.pdf.

Aalund:2018:SDD

[2207] Oskar Ålund and Jan Nordström. A stable domain decomposition technique for advection–diffusion problems. Journal of Scientific Computing, 77(2):755–774, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0722-x;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0722-x.pdf.

Wang:2018:IHO

[2208] Siyang Wang. An improved high order finite difference method for non-conforming grid interfaces for the wave equation. Journal of Scientific Computing, 77(2):775–792, November 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0723-9;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0723-9.pdf.

Gao:2018:LLC

[2209] Huadong Gao and Pengtao Sun. A linearized local conservative mixed finite element method for Poisson–Nernst–Planck equations. Journal of Scientific Computing, 77(2):793–817, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0727-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0727-5.pdf. See correction [2210].

Gao:2018:CLL

[2210] Huadong Gao and Pengtao Sun. Correction to: A Linearized Local Conservative Mixed Finite Element Method for Poisson–Nernst– Planck equations. Journal of Scientific Computing, 77(2):818, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0763-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0763-1.pdf. See [2209].

Ghosh:2018:KSC

[2211] Debojyoti Ghosh, Mikhail A. Dorf, Milo R. Dorr, and Jeffrey A. F. Hittinger. Kinetic simulation of collisional magnetized plasmas with semi-implicit time integration. Journal of Scientific Computing, 77 (2):819–849, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0726-6;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0726-6.pdf.

Boscarino:2018:AMN

[2212] S. Boscarino, G. Russo, and L. Scandurra. All Mach number second order semi-implicit scheme for the Euler equations of gas dynamics. Journal of Scientific Computing, 77(2):850–884, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0731-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0731-9.pdf.

Fumagalli:2018:DLR

[2213] Alessio Fumagalli and Isabelle Faille. A double–layer reduced model for fault flow on slipping domains with an hybrid finite volume scheme. Journal of Scientific Computing, 77(2):885–910, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0740-8;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0740-8.pdf.

Kang:2018:ICB

[2214] Myeongmin Kang, Myungjoo Kang, and Miyoun Jung. Image colorization based on a generalization of the low dimensional manifold model. Journal of Scientific Computing, 77(2):911–935, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0732-8;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0732-8.pdf.

Finlay:2018:AHF

[2215] Chris Finlay and Adam M. Oberman. Approximate homogenization of fully nonlinear elliptic PDEs: Estimates and numerical results for pucci type equations. Journal of Scientific Computing, 77(2):936–949, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0730-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0730-x.pdf.

Bertaccini:2018:LMB

[2216] Daniele Bertaccini and Fabio Durastante. Limited memory block preconditioners for fast solution of fractional partial differential equations. Journal of Scientific Computing, 77(2):950–970, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0729-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0729-3.pdf.

Frittelli:2018:NPV

[2217] Massimo Frittelli, Anotida Madzvamuse, Ivonne Sgura, and Chandrasekhar Venkataraman. Numerical preservation of velocity induced invariant regions for reaction–diffusion systems on evolving surfaces. Journal of Scientific Computing, 77(2):971–1000, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0741-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0741-7.pdf.

Zhou:2018:SAE

[2218] Lingling Zhou and Yan Xu. Stability analysis and error estimates of semi-implicit spectral deferred correction coupled with local discontinuous Galerkin method for linear convection–diffusion equations. Journal of Scientific Computing, 77(2):1001–1029, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0735-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0735-5.pdf.

Crowder:2018:CCA

[2219] Adam J. Crowder and Catherine E. Powell. CBS constants & their role in error estimation for stochastic Galerkin finite element methods. Journal of Scientific Computing, 77(2):1030–1054, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0736-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0736-4.pdf.

Bocher:2018:ERK

[2220] Philippe Bocher, Juan I. Montijano, Luis Rández, and Marnix Van Daele. Explicit Runge–Kutta methods for stiff problems with a gap in their eigenvalue spectrum. Journal of Scientific Computing, 77(2):1055–1083, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0737-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0737-3.pdf.

Huang:2018:IEL

[2221] Chieh-Sen Huang and Todd Arbogast. An implicit Eulerian– Lagrangian WENO3 scheme for nonlinear conservation laws. Journal of Scientific Computing, 77(2):1084–1114, November 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0738-2;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0738-2.pdf.

Alvarez:2018:CFF

[2222] Diego Álvarez, Pedro González-Rodríguez, and Miguel Moscoso. A closed–form formula for the RBF–Based approximation of the Laplace– Beltrami operator. Journal of Scientific Computing, 77(2):1115–1132, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0739-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0739-1.pdf.

Yin:2018:SBE

[2223] Penghang Yin, Minh Pham, Adam Oberman, and Stanley Osher. Stochastic backward Euler: an implicit gradient descent algorithm for k-means clustering. Journal of Scientific Computing, 77(2):1133–1146, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0744-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0744-4.pdf.

Li:2018:DDM

[2224] Yibao Li, Xuelin Qi, and Junseok Kim. Direct discretization method for the Cahn–Hilliard equation on an evolving surface. Journal of Scientific Computing, 77(2):1147–1163, November 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0742-6;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0742-6.pdf.

Li:2018:TSA

[2225] Binjie Li, Hao Luo, and Xiaoping Xie. A time-spectral algorithm for fractional wave problems. Journal of Scientific Computing, 77(2):1164– 1184, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0743-5.pdf.

Wang:2018:ESO

[2226] Lin Wang and Haijun Yu. On efficient second order stabilized semiimplicit schemes for the Cahn–Hilliard phase–field equation. Journal of Scientific Computing, 77(2):1185–1209, November 2018. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-018-0746-2.pdf.

Han:2018:SOT

[2227] Daozhi Han and Xiaoming Wang. A second order in time, decoupled, unconditionally stable numerical scheme for the Cahn–Hilliard-Darcy system. Journal of Scientific Computing, 77(2):1210–1233, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0748-0.pdf.

Brenner:2018:RSM

[2228] Susanne C. Brenner, Amanda E. Diegel, and Li-Yeng Sung. A robust solver for a mixed finite element method for the Cahn–Hilliard equation. Journal of Scientific Computing, 77(2):1234–1249, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-018-0753-3.pdf.

Chen:2018:NRC

[2229] Shu sheng Chen, Chao Yan, Bo xi Lin, and Yan su Li. A new robust carbuncle–free Roe scheme for strong shock. Journal of Scientific Computing, 77(2):1250–1277, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0747-1.pdf.

Rydin:2018:HFS

[2230] Ylva Rydin, Ken Mattsson, and Jonatan Werpers. High-fidelity sound propagation in a varying 3D atmosphere. Journal of Scientific Computing, 77(2):1278–1302, November 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0751-5.pdf.

Chen:2018:FSI

[2231] Yanlai Chen, Bo Dong, and Chi-Wang Shu. A foreword to the special issue in honor of Professor Bernardo Cockburn on his 60th birthday: a life time of discontinuous schemings. Journal of Scientific Computing, 77 (3):1303–1309, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0845-0;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0845-0.pdf.

Aghili:2018:ARH

[2232] Joubine Aghili and Daniele A. Di Pietro. An advection-robust hybrid high-order method for the Oseen problem. Journal of Scientific Computing, 77(3):1310–1338, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0681-2;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0681-2.pdf.

Antonietti:2018:FNI

[2233] Paola F. Antonietti, Paul Houston, and Giorgio Pennesi. Fast numerical integration on polytopic meshes with applications to discontinuous Galerkin finite element methods. Journal of Scientific Computing, 77 (3):1339–1370, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0802-y;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0802-y.pdf.

Bona:2018:FEM

[2234] Jerry L. Bona, Hongqiu Chen, Ohannes Karakashian, and Michael M. Wise. Finite element methods for a system of dispersive equations. Journal of Scientific Computing, 77(3):1371–1401, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0767-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0767-x.pdf.

Cao:2018:SRD

[2235] Waixiang Cao and Zhimin Zhang. Some recent developments in superconvergence of discontinuous Galerkin methods for time-dependent partial differential equations. Journal of Scientific Computing, 77 (3):1402–1423, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0762-2;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0762-2.pdf.

Cascavita:2018:HDM

[2236] Karol L. Cascavita, Jérémy Bleyer, Xavier Chateau, and Alexandre Ern. Hybrid discretization methods with adaptive yield surface detection for Bingham pipe flows. Journal of Scientific Computing, 77(3):1424–1443, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0745-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0745-3.pdf.

Castillo:2018:CFN

[2237] P. Castillo and S. Gómez. On the conservation of fractional nonlinear Schrödinger Equation’s invariants by the local discontinuous Galerkin method. Journal of Scientific Computing, 77(3):1444–1467, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0708-8;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0708-8.pdf.

Chen:2018:ESM

[2238] Sheng Chen and Jie Shen. Enriched spectral methods and applications to problems with weakly singular solutions. Journal of Scientific Computing, 77(3):1468–1489, December 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0862-z;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0862-z.pdf.

Du:2018:ASD

[2239] Jie Du and Eric Chung. An adaptive staggered discontinuous Galerkin method for the steady state convection–diffusion equation. Journal of Scientific Computing, 77(3):1490–1518, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0695-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0695-9.pdf.

Farhat:2018:ANT

[2240] Aseel Farhat, Hans Johnston, Michael Jolly, and Edriss S. Titi. Assimilation of nearly turbulent Rayleigh–Bénard flow through vorticity or local circulation measurements: a computational study. Journal of Scientific Computing, 77(3):1519–1533, December 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0686-x;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0686-x.pdf.

Feng:2018:NLD

[2241] Xiaobing Feng and Thomas Lewis. Nonstandard local discontinuous Galerkin methods for fully nonlinear second order elliptic and parabolic equations in high dimensions. Journal of Scientific Computing, 77 (3):1534–1565, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0765-z;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0765-z.pdf.

Fernandez:2018:HDG

[2242] P. Fernandez, A. Christophe, S. Terrana, N. C. Nguyen, and J. Peraire. Hybridized discontinuous Galerkin methods for wave propagation. Journal of Scientific Computing, 77(3):1566–1604, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0811-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0811-x.pdf.

Fu:2018:SCH

[2243] Guosheng Fu and Christoph Lehrenfeld. A strongly conservative hybrid DG/mixed FEM for the coupling of Stokes and Darcy flow. Journal of Scientific Computing, 77(3):1605–1620, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0691-0;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0691-0.pdf.

Fu:2018:GDF

[2244] Pei Fu, Fengyan Li, and Yan Xu. Globally divergence–free discontinuous Galerkin methods for ideal magnetohydrodynamic equations. Journal of Scientific Computing, 77(3):1621–1659, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0750-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0750-6.pdf.

Gao:2018:EAM

[2245] Huadong Gao and Weifeng Qiu. Error analysis of mixed finite element methods for nonlinear parabolic equations. Journal of Scientific Computing, 77(3):1660–1678, December 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0643-8;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0643-8.pdf.

Giacomini:2018:SHM

[2246] Matteo Giacomini, Alexandros Karkoulias, Ruben Sevilla, and Antonio Huerta. A superconvergent HDG method for Stokes flow with strongly enforced symmetry of the stress tensor. Journal of Scientific Computing, 77(3):1679–1702, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0855-y;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0855-y.pdf.

Gopalakrishnan:2018:DAH

[2247] Jay Gopalakrishnan, Manuel Solano, and Felipe Vargas. Dispersion analysis of HDG methods. Journal of Scientific Computing, 77(3):1703–1735, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0781-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0781-z.pdf.

Guzman:2018:AFE

[2248] Johnny Guzman, Alexandre Madureira, Marcus Sarkis, and Shawn Walker. Analysis of the finite element method for the Laplace– Beltrami equation on surfaces with regions of high curvature using graded meshes. Journal of Scientific Computing, 77(3):1736–1761, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-017-0580-y;   https://link.springer.com/content/pdf/
10.1007/s10915-017-0580-y.pdf.

Kanschat:2018:FEM

[2249] Guido Kanschat and Beatrice Riviere. A finite element method with strong mass conservation for Biot’s linear consolidation model. Journal of Scientific Computing, 77(3):1762–1779, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0843-2;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0843-2.pdf.

Li:2018:UHM

[2250] Guanglian Li and Ke Shi. Upscaled HDG methods for Brinkman equations with high–contrast heterogeneous coefficient. Journal of Scientific Computing, 77(3):1780–1800, December 2018. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0725-7;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0725-7.pdf.

Ling:2018:CHO

[2251] Dan Ling, Juan Cheng, and Chi-Wang Shu. Conservative high order positivity–preserving discontinuous Galerkin methods for linear hyperbolic and radiative transfer equations. Journal of Scientific Computing, 77(3):1801–1831, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0700-3.pdf.

AlbellaMartinez:2018:SLI

[2252] Jorge Albella Martínez, Sébastien Imperiale, Patrick Joly, and Jerónimo Rodríguez. Solving 2D linear isotropic elastodynamics by means of scalar potentials: a new challenge for finite elements. Journal of Scientific Computing, 77(3):1832–1873, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.

com/content/pdf/10.1007/s10915-018-0768-9.pdf.

Mascotto:2018:NCH

[2253] Lorenzo Mascotto, Ilaria Perugia, and Alexander Pichler. Nonconforming harmonic virtual element method: h- and p-versions. Journal of Scientific Computing, 77(3):1874–1908, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-018-0797-4.pdf.

Muller:2018:DGM

[2254] Fabian Müller, Dominik Schötzau, and Christoph Schwab. Discontinuous Galerkin methods for acoustic wave propagation in polygons. Journal of Scientific Computing, 77(3):1909–1935, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-018-0706-x.pdf.

Rhebergen:2018:PHD

[2255] Sander Rhebergen and Garth N. Wells. Preconditioning of a hybridized discontinuous Galerkin finite element method for the Stokes equations. Journal of Scientific Computing, 77(3):1936–1952, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0760-4.pdf.

Sevilla:2018:HND

[2256] Ruben Sevilla and Antonio Huerta. HDG-NEFEM with degree adaptivity for Stokes flows. Journal of Scientific Computing, 77(3):1953– 1980, December 2018. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0657-2.pdf.

Wang:2019:LDG

[2257] Haijin Wang, Jingjing Zheng, Fan Yu, Hui Guo, and Qiang Zhang. Local discontinuous Galerkin method with implicit–explicit time marching for incompressible miscible displacement problem in porous media. Journal of Scientific Computing, 78(1):1–28, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0752-4;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0752-4.pdf.

Wang:2019:GCA

[2258] Yu Wang, Wotao Yin, and Jinshan Zeng. Global convergence of ADMM in nonconvex nonsmooth optimization. Journal of Scientific Computing, 78(1):29–63, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0757-z;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0757-z.pdf.

Kumar:2019:EBD

[2259] S. Kumar, R. Ruiz-Baier, and R. Sandilya. Error bounds for discontinuous finite volume discretisations of Brinkman optimal control problems. Journal of Scientific Computing, 78(1):64–93, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0749-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0749-z.pdf.

Gelb:2019:REB

[2260] Anne Gelb and Theresa Scarnati. Reducing effects of bad data using variance based joint sparsity recovery. Journal of Scientific Computing, 78(1):94–120, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0754-2;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0754-2.pdf.

Han:2019:OOE

[2261] Weimin Han, Limin He, and Fei Wang. Optimal order error estimates for discontinuous Galerkin methods for the wave equation. Journal of Scientific Computing, 78(1):121–144, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0755-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0755-1.pdf.

Li:2019:NSS

[2262] Min Li, Hicham Chaouki, Jean-Loup Robert, Donald Ziegler, and Mario Fafard. Numerical simulation of Stefan problem coupled with mass transport in a binary system through XFEM/level set method. Journal of Scientific Computing, 78(1):145–166, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0759-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0759-x.pdf.

Narayanamurthi:2019:EET

[2263] Mahesh Narayanamurthi, Paul Tranquilli, Adrian Sandu, and Mayya Tokman. EPIRK-W and EPIRK-K time discretization methods. Journal of Scientific Computing, 78(1):167–201, January 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0761-3;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0761-3.pdf.

Li:2019:LMMa

[2264] Zhaoxiang Li, Bingbing Ji, and Jianxin Zhou. A local minimax method using virtual geometric objects: Part I — for finding saddles. Journal of Scientific Computing, 78(1):202–225, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0774-y;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0774-y.pdf.

Li:2019:LMMb

[2265] Zhaoxiang Li and Jianxin Zhou. A local minimax method using virtual geometric objects: Part II — for finding equality constrained saddles. Journal of Scientific Computing, 78(1):226–245, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0775-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0775-x.pdf.

Carvalho:2019:CNP

[2266] Pitágoras P. Carvalho and Enrique Fernández-Cara. On the computation of Nash and Pareto equilibria for some bi-objective control problems. Journal of Scientific Computing, 78(1):246–273, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0764-0;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0764-0.pdf.

Zhang:2019:OQF

[2267] Shun Zhang and Erich Novak. Optimal quadrature formulas for the Sobolev space H^{1}. Journal of Scientific Computing, 78(1):274–289, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0766-y;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0766-y.pdf.

Huang:2019:HDG

[2268] Jianguo Huang and Xuehai Huang. A hybridizable discontinuous Galerkin method for Kirchhoff plates. Journal of Scientific Computing, 78(1):290–320, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0780-0;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0780-0.pdf.

Paipuri:2019:CCH

[2269] Mahendra Paipuri, Carlos Tiago, and Sonia Fernández-Méndez. Coupling of continuous and hybridizable discontinuous Galerkin methods: Application to conjugate heat transfer problem. Journal of Scientific Computing, 78(1):321–350, January 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0769-8;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0769-8.pdf.

Bi:2019:LPF

[2270] Hai Bi, Jiayu Han, and Yidu Yang. Local and parallel finite element algorithms for the transmission eigenvalue problem. Journal of Scientific Computing, 78(1):351–375, January 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0770-2;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0770-2.pdf.

Herrero:2019:SMR

[2271] H. Herrero, F. Pla, and M. Ruiz-Ferrández. A Schwarz method for a Rayleigh–Bénard problem. Journal of Scientific Computing, 78(1):376–392, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0771-1;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0771-1.pdf.

Dimarco:2019:CLD

[2272] Giacomo Dimarco, Cory Hauck, and Raphaël Loubère. A class of low dissipative schemes for solving kinetic equations. Journal of Scientific Computing, 78(1):393–432, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0776-9;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0776-9.pdf.

Rueda-Ramirez:2019:TEE

[2273] Andrés M. Rueda-Ramírez, Gonzalo Rubio, Esteban Ferrer, and Eusebio Valero. Truncation error estimation in the p-anisotropic discontinuous Galerkin spectral element method. Journal of Scientific Computing, 78(1):433–466, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0772-0;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0772-0.pdf.

Sun:2019:TSO

[2274] Hong Sun, Xuan Zhao, and Zhi zhong Sun. The temporal second order difference schemes based on the interpolation approximation for the time multi-term fractional wave equation. Journal of Scientific Computing, 78(1):467–498, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0820-9;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0820-9.pdf.

Baeza:2019:CWS

[2275] Antonio Baeza, Raimund Bürger, Pep Mulet, and David Zorío. Central WENO schemes through a global average weight. Journal of Scientific Computing, 78(1):499–530, January 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0773-z;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0773-z.pdf.

Fan:2019:AMS

[2276] Jinyan Fan, Jianchao Huang, and Jianyu Pan. An adaptive multistep Levenberg–Marquardt method. Journal of Scientific Computing, 78(1):531–548, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0777-8;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0777-8.pdf.

Li:2019:LBR

[2277] Min Li and Chengming Huang. The linear barycentric rational quadrature method for auto–convolution Volterra integral equations. Journal of Scientific Computing, 78(1):549–564, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-018-0779-6.pdf.

Cantwell:2019:MIL

[2278] Chris D. Cantwell and Allan S. Nielsen. A minimally intrusive lowmemory approach to resilience for existing transient solvers. Journal of Scientific Computing, 78(1):565–581, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.

springer.com/content/pdf/10.1007/s10915-018-0778-7.pdf.

Nasser:2019:NCP

[2279] Mohamed M. S. Nasser. Numerical computing of preimage domains for bounded multiply connected slit domains. Journal of Scientific Computing, 78(1):582–606, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0784-9.pdf.

Liu:2019:BML

[2280] Jun Liu and Stanley Osher. Block matching local SVD operator based sparsity and TV regularization for image denoising. Journal of Scientific Computing, 78(1):607–624, January 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://link.springer.com/

content/pdf/10.1007/s10915-018-0785-8.pdf.

Antonietti:2019:VCM

[2281] P. F. Antonietti and G. Pennesi. V-cycle multigrid algorithms for discontinuous Galerkin methods on non-nested polytopic meshes. Journal of Scientific Computing, 78(1):625–652, January 2019. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-018-0783-x.pdf.

Pei:2019:NMI

[2282] Chaoxu Pei, Mark Sussman, and M. Yousuff Hussaini. New multiimplicit space–time spectral element methods for advection–diffusionreaction problems. Journal of Scientific Computing, 78(2):653–686, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0654-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0654-5.pdf.

Yang:2019:LSM

[2283] Suxiang Yang, Huanzhen Chen, and Hong Wang. Least-squared mixed variational formulation based on space decomposition for a kind of variable–coefficient fractional diffusion problems. Journal of Scientific Computing, 78(2):687–709, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0782-y;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0782-y.pdf.

Huybrechs:2019:HFA

[2284] Daan Huybrechs and Peter Opsomer. High-frequency asymptotic compression of dense BEM matrices for general geometries without ray tracing. Journal of Scientific Computing, 78(2):710–745, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0786-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0786-7.pdf.

Sun:2019:NSB

[2285] Dong-Lin Sun, Ting-Zhu Huang, Bruno Carpentieri, and Yan-Fei Jing. A new shifted block GMRES method with inexact breakdowns for solving multi-shifted and multiple right–hand sides linear systems. Journal of Scientific Computing, 78(2):746–769, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0787-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0787-6.pdf.

Sun:2019:CNS

[2286] Dong-Lin Sun, Ting-Zhu Huang, Bruno Carpentieri, and Yan-Fei Jing. Correction to: A New Shifted Block GMRES Method with Inexact Breakdowns for Solving Multi-Shifted and Multiple Right-Hand Sides Linear Systems. Journal of Scientific Computing, 78(2):770–771, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0804-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0804-9.pdf.

Chen:2019:UWD

[2287] Anqi Chen, Fengyan Li, and Yingda Cheng. An ultra-weak discontinuous Galerkin method for Schrödinger equation in one dimension. Journal of Scientific Computing, 78(2):772–815, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0789-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0789-4.pdf.

Fu:2019:DGM

[2288] Pei Fu, Yingda Cheng, Fengyan Li, and Yan Xu. Discontinuous Galerkin methods with optimal L^{2} accuracy for one dimensional linear PDEs with high order spatial derivatives. Journal of Scientific Computing, 78(2):816–863, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0788-5;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0788-5.pdf.

Chen:2019:DFW

[2289] Long Chen and Feng Wang. A divergence free weak virtual element method for the Stokes problem on polytopal meshes. Journal of Scientific Computing, 78(2):864–886, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0796-5;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0796-5.pdf.

Almonacid:2019:PEA

[2290] Javier A. Almonacid, Gabriel N. Gatica, and Ricardo Oyarzúa. A posteriori error analysis of a mixed–primal finite element method for the Boussinesq problem with temperature–dependent viscosity. Journal of Scientific Computing, 78(2):887–917, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0810-y;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0810-y.pdf.

Laiu:2019:PLF

[2291] M. Paul Laiu and Cory D. Hauck. Positivity limiters for filtered spectral approximations of linear kinetic transport equations. Journal of Scientific Computing, 78(2):918–950, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0790-y;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0790-y.pdf.

Lee:2019:DDM

[2292] Chang-Ock Lee, Changmin Nam, and Jongho Park. Domain decomposition methods using dual conversion for the total variation minimization with L^{1} fidelity term. Journal of Scientific Computing, 78(2):951–970, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0791-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0791-x.pdf.

Gouasmi:2019:ECS

[2293] Ayoub Gouasmi, Scott M. Murman, and Karthik Duraisamy. Entropy conservative schemes and the receding flow problem. Journal of Scientific Computing, 78(2):971–994, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0793-8;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0793-8.pdf.

Birken:2019:PSF

[2294] Philipp Birken, Jonathan Bull, and Antony Jameson. Preconditioned smoothers for the full approximation scheme for the RANS equations. Journal of Scientific Computing, 78(2):995–1022, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0792-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0792-9.pdf.

Ke:2019:AFD

[2295] Guoyi Ke and Wei Guo. An alternative formulation of discontinous Galerkin schemes for solving Hamilton–Jacobi equations. Journal of Scientific Computing, 78(2):1023–1044, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0794-7;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0794-7.pdf.

Moxey:2019:IEB

[2296] David Moxey, Shankar P. Sastry, and Robert M. Kirby. Interpolation error bounds for curvilinear finite elements and their implications on adaptive mesh refinement. Journal of Scientific Computing, 78(2):1045–1062, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0795-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0795-6.pdf.

You:2019:NMM

[2297] Juntao You, Yuling Jiao, Xiliang Lu, and Tieyong Zeng. A nonconvex model with minimax concave penalty for image restoration. Journal of Scientific Computing, 78(2):1063–1086, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0801-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0801-z.pdf.

Zhang:2019:EES

[2298] Zhongqiang Zhang. Error estimates of spectral Galerkin methods for a linear fractional reaction–diffusion equation. Journal of Scientific Computing, 78(2):1087–1110, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0800-0;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0800-0.pdf.

Kimura:2019:GFS

[2299] Masato Kimura, Hirofumi Notsu, Yoshimi Tanaka, and Hiroki Yamamoto. The gradient flow structure of an extended Maxwell viscoelastic model and a structure–preserving finite element scheme. Journal of Scientific Computing, 78(2):1111–1131, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0799-2;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0799-2.pdf.

Wolf:2019:AMF

[2300] Eric M. Wolf and Christopher R. Schrock. Adaptive modal filters based on artificial and spectral viscosity techniques. Journal of Scientific Computing, 78(2):1132–1151, February 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0798-3;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0798-3.pdf.

Mittal:2019:MSS

[2301] Ketan Mittal and Paul Fischer. Mesh smoothing for the spectral element method. Journal of Scientific Computing, 78(2):1152–1173, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0812-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0812-9.pdf.

Zaspel:2019:APM

[2302] Peter Zaspel. Algorithmic patterns for H-matrices on many–core processors. Journal of Scientific Computing, 78(2):1174–1206, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-018-0809-4.pdf.

Wang:2019:SAP

[2303] Lixiu Wang, Qian Zhang, and Zhimin Zhang. Superconvergence analysis and PPR recovery of arbitrary order edge elements for Maxwell’s equations. Journal of Scientific Computing, 78(2):1207–1230, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-018-0805-8.pdf.

Mohammadi:2019:OSD

[2304] Seyyed Abbas Mohammadi, Farid Bozorgnia, and Heinrich Voss. Optimal shape design for the p-Laplacian eigenvalue problem. Journal of Scientific Computing, 78(2):1231–1249, February 2019. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-018-0806-7.pdf.

Boffi:2019:PEE

[2305] Daniele Boffi, Lucia Gastaldi, Rodolfo Rodríguez, and Ivana Šebestová. A posteriori error estimates for Maxwell’s eigenvalue problem. Journal of Scientific Computing, 78(2):1250–1271, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL http://

link.springer.com/content/pdf/10.1007/s10915-018-0808-5.pdf.

Harbrecht:2019:ACS

[2306] Helmut Harbrecht and Peter Zaspel. On the algebraic construction of sparse multilevel approximations of elliptic tensor product problems. Journal of Scientific Computing, 78(2):1272–1290, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (elec-

tronic). URL http://link.springer.com/content/pdf/10.1007/
s10915-018-0807-6.pdf.

Vevek:2019:ASD

[2307] U. S. Vevek, B. Zang, and T. H. New. On alternative setups of the double Mach reflection problem. Journal of Scientific Computing, 78(2):1291– 1303, February 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL http://link.springer.com/content/pdf/10.
1007/s10915-018-0803-x.pdf.

Li:2019:ADG

[2308] Jia Li, Dazhi Zhang, Xiong Meng, and Boying Wu. Analysis of discontinuous Galerkin methods with upwind-biased fluxes for one dimensional linear hyperbolic equations with degenerate variable coefficients. Journal of Scientific Computing, 78(3):1305–1328, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0831-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0831-6.pdf.

Chen:2019:VIR

[2309] Liyuan Chen, Yutian Li, and Tieyong Zeng. Variational image restoration and segmentation with Rician noise. Journal of Scientific Computing, 78(3):1329–1352, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0826-3;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0826-3.pdf.

Yueh:2019:NSE

[2310] Mei-Heng Yueh, Wen-Wei Lin, Chin-Tien Wu, and Shing-Tung Yau. A novel stretch energy minimization algorithm for equiareal parameterizations. Journal of Scientific Computing, 78(3):1353–1386, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0822-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0822-7.pdf.

Guo:2019:EDL

[2311] Li Guo, Xingjie Helen Li, and Yang Yang. Energy dissipative local discontinuous Galerkin methods for Keller–Segel chemotaxis model. Journal of Scientific Computing, 78(3):1387–1404, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0813-8;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0813-8.pdf.

Droniou:2019:HDM

[2312] Jérôme Droniou, Bishnu P. Lamichhane, and Devika Shylaja. The Hessian discretisation method for fourth order linear elliptic equations. Journal of Scientific Computing, 78(3):1405–1437, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0814-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0814-7.pdf.

Du:2019:AFD

[2313] Qiang Du, Lili Ju, and Jianfang Lu. Analysis of fully discrete approximations for dissipative systems and application to time-dependent nonlocal diffusion problems. Journal of Scientific Computing, 78(3):1438–1466, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0815-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0815-6.pdf.

Cheng:2019:HEA

[2314] Qing Cheng, Jie Shen, and Xiaofeng Yang. Highly efficient and accurate numerical schemes for the epitaxial thin film growth models by using the SAV approach. Journal of Scientific Computing, 78(3):1467–1487, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0832-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0832-5.pdf.

Lau:2019:VMJ

[2315] Chun Pong Lau, Yu Hin Lai, and Lok Ming Lui. Variational models for joint subsampling and reconstruction of turbulence-degraded images. Journal of Scientific Computing, 78(3):1488–1525, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0833-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0833-4.pdf.

Buccini:2019:RML

[2316] Alessandro Buccini and Lothar Reichel. An ℓ^{2}–ℓ^{q} regularization method for large discrete ill-posed problems. Journal of Scientific Computing, 78(3):1526–1549, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0816-5;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0816-5.pdf.

Pronzato:2019:MMR

[2317] Luc Pronzato and Anatoly Zhigljavsky. Measures minimizing regularized dispersion. Journal of Scientific Computing, 78(3):1550–1570, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0817-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0817-4.pdf.

Shen:2019:SGM

[2318] Wanfang Shen, Liang Ge, and Wenbin Liu. Stochastic Galerkin method for optimal control problem governed by random elliptic PDE with state constraints. Journal of Scientific Computing, 78(3):1571–1600, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0823-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0823-6.pdf.

Harness:2019:LDS

[2319] Yuval Harness. Low-dimensional spatial embedding method for shape uncertainty quantification in acoustic scattering by 2D star shaped obstacles. Journal of Scientific Computing, 78(3):1601–1631, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0818-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0818-3.pdf.

Xie:2019:FOK

[2320] Yaning Xie and Wenjun Ying. A fourth-order kernel-free boundary integral method for the modified Helmholtz equation. Journal of Scientific Computing, 78(3):1632–1658, March 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0821-8;   https://link.springer.com/content/pdf/10.1007/s10915-

018-0821-8.pdf. See correction [2321].

Xie:2019:CFO

[2321] Yaning Xie and Wenjun Ying. Correction to: a fourth-order kernel-free boundary integral method for the modified Helmholtz equation. Journal of Scientific Computing, 78(3):1659, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0836-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0836-1.pdf. See [2320].

Zhao:2019:LDG

[2322] Di Zhao and Qiang Zhang. Local discontinuous Galerkin methods with generalized alternating numerical fluxes for two-dimensional linear Sobolev equation. Journal of Scientific Computing, 78(3):1660–1690, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0819-2;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0819-2.pdf.

Berkhout:2019:JSP

[2323] Joost Berkhout and Bernd F. Heidergott. The jump start power method: a new approach for computing the ergodic projector of a finite Markov chain. Journal of Scientific Computing, 78(3):1691–1723, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0828-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0828-1.pdf.

Fu:2019:PFP

[2324] Hongfei Fu and Hong Wang. A preconditioned fast parareal finite difference method for space–time fractional partial differential equation. Journal of Scientific Computing, 78(3):1724–1743, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0835-2;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0835-2.pdf.

Ma:2019:CII

[2325] Yunyun Ma and Yuesheng Xu. Computing integrals involved the Gaussian function with a small standard deviation. Journal of Scientific Computing, 78(3):1744–1767, March 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0825-4;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0825-4.pdf.

Sheikholeslami:2019:CAK

[2326] Somayyeh Sheikholeslami, James V. Lambers, and Carley Walker. Convergence analysis of Krylov subspace spectral methods for reaction– diffusion equations. Journal of Scientific Computing, 78(3):1768–1789, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0824-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0824-5.pdf.

Yang:2019:AII

[2327] Ximei Yang and Yanqin Bai. An adaptive infeasible–interior-point method with the one-norm wide neighborhood for semi-definite programming. Journal of Scientific Computing, 78(3):1790–1810, March 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0827-2;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0827-2.pdf.

Glowinski:2019:FEO

[2328] Roland Glowinski, Hao Liu, Shingyu Leung, and Jianliang Qian. A finite element/operator-splitting method for the numerical solution of the two dimensional elliptic Monge–Ampère equation. Journal of Scientific Computing, 79(1):1–47, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0839-y;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0839-y.pdf. See correction

Glowinski:2019:CFE

[2329] Roland Glowinski, Hao Liu, Shingyu Leung, and Jianliang Qian. Correction to: A Finite Element/Operator-Splitting Method for the Numerical Solution of the Two Dimensional Elliptic Monge– Ampère Equation. Journal of Scientific Computing, 79(1):48, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0854-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0854-z.pdf. See [2328].

Oyarzua:2019:HOM

[2330] Ricardo Oyarzúa, Manuel Solano, and Paulo Zúñiga. A high order mixed-FEM for diffusion problems on curved domains. Journal of Scientific Computing, 79(1):49–78, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0840-5;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0840-5.pdf.

Chandrashekar:2019:GDC

[2331] Praveen Chandrashekar. A global divergence conforming DG method for hyperbolic conservation laws with divergence constraint. Journal of Scientific Computing, 79(1):79–102, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-0841-4;   https://link.
springer.com/content/pdf/10.1007/s10915-018-0841-4.pdf.

Fermo:2019:NMC

[2332] Luisa Fermo, Maria Grazia Russo, and Giada Serafini. Numerical methods for Cauchy bisingular integral equations of the first kind on the square. Journal of Scientific Computing, 79(1):103–127, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0842-3;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0842-3.pdf.

Cai:2019:FRO

[2333] Jian-Feng Cai, Haixia Liu, and Yang Wang. Fast rank-one alternating minimization algorithm for phase retrieval. Journal of Scientific Computing, 79(1):128–147, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0857-9;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0857-9.pdf.

Guo:2019:FMM

[2334] Ruchi Guo, Tao Lin, and Yanping Lin. A fixed mesh method with immersed finite elements for solving interface inverse problems. Journal of Scientific Computing, 79(1):148–175, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0847-y;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0847-y.pdf.

Freret:2019:HOF

[2335] L. Freret, L. Ivan, H. De Sterck, and C. P. T. Groth. Highorder finite-volume method with block-based AMR for magnetohydrodynamics flows. Journal of Scientific Computing, 79(1):176–208, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0844-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0844-1.pdf.

Ainsworth:2019:DBE

[2336] Mark Ainsworth and Guosheng Fu. Dispersive behavior of an energy-conserving discontinuous Galerkin method for the one-way wave equation. Journal of Scientific Computing, 79(1):209–226, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0846-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0846-z.pdf.

Baffet:2019:GJK

[2337] Daniel Baffet. A Gauss–Jacobi kernel compression scheme for fractional differential equations. Journal of Scientific Computing, 79(1):227–248, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0848-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0848-x.pdf.

Glaubitz:2019:SCS

[2338] J. Glaubitz, A. C. Nogueira, Jr., J. L. S. Almeida, R. F. Cantão, and C. A. C. Silva. Smooth and compactly supported viscous sub-cell shock capturing for discontinuous Galerkin methods. Journal of Scientific Computing, 79(1):249–272, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0850-3;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0850-3.pdf.

Escalante:2019:ETL

[2339] C. Escalante, E. D. Fernández-Nieto, T. Morales de Luna, and M. J. Castro. An efficient two-layer non-hydrostatic approach for dispersive water waves. Journal of Scientific Computing, 79(1):273–320, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0849-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0849-9.pdf.

Marchi:2019:RBP

[2340] S. De Marchi, A. Martínez, E. Perracchione, and M. Rossini. RBF-Based partition of unity methods for elliptic PDEs: Adaptivity and stability issues via variably scaled kernels. Journal of Scientific Computing, 79(1):321–344, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0851-2;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0851-2.pdf.

Chen:2019:NMW

[2341] Zhenzhu Chen, Yunfeng Xiong, and Sihong Shao. Numerical methods for the Wigner equation with unbounded potential. Journal of Scientific Computing, 79(1):345–368, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0853-0;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0853-0.pdf.

Lee:2019:NSV

[2342] Hsueh-Chen Lee and Hyesuk Lee. Numerical simulations of viscoelastic fluid flows past a transverse slot using least-squares finite element methods. Journal of Scientific Computing, 79(1):369–388, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0856-x;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0856-x.pdf.

Shan:2019:PTS

[2343] Li Shan, Jiangyong Hou, Wenjing Yan, and Jie Chen. Partitioned time stepping method for a dual-porosity-Stokes model. Journal of Scientific Computing, 79(1):389–413, April 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0879-3;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0879-3.pdf.

Reddy:2019:PEA

[2344] G. Murali Mohan Reddy, Rajen Kumar Sinha, and José Alberto Cuminato. A posteriori error analysis of the Crank–Nicolson finite element method for parabolic integro–differential equations. Journal of Scientific Computing, 79(1):414–441, April 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0860-1;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0860-1.pdf.

Lin:2019:NIF

[2345] Tao Lin, Dongwoo Sheen, and Xu Zhang. A nonconforming immersed finite element method for elliptic interface problems. Journal of Scientific Computing, 79(1):442–463, April 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0865-9;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0865-9.pdf.

Shen:2019:ABA

[2346] Yedan Shen, Yang Kuang, and Guanghui Hu. An asymptoticsbased adaptive finite element method for Kohn–Sham equation. Journal of Scientific Computing, 79(1):464–492, April 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0861-0;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0861-0.pdf.

Mohammadi:2019:NSE

[2347] Vahid Mohammadi, Davoud Mirzaei, and Mehdi Dehghan. Numerical simulation and error estimation of the time-dependent Allen–Cahn equation on surfaces with radial basis functions. Journal of Scientific Computing, 79(1):493–516, April 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0859-7;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0859-7.pdf.

Li:2019:AAD

[2348] Huan Li and Zhouchen Lin. Accelerated alternating direction method of multipliers: an optimal O (1 40/40 K ) nonergodic analysis. Journal of Scientific Computing, 79(2):671–699, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0893-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0893-5.pdf.

Hao:2019:FEM

[2349] Zhaopeng Hao, Moongyu Park, Guang Lin, and Zhiqiang Cai. Finite element method for two-sided fractional differential equations with variable coefficients: Galerkin approach. Journal of Scientific Computing, 79(2):700–717, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0869-5;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0869-5.pdf.

Badia:2019:PBB

[2350] Santiago Badia, Alberto F. Martín, and Hieu Nguyen. Physicsbased balancing domain decomposition by constraints for multimaterial problems. Journal of Scientific Computing, 79(2):718–747, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0870-z;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0870-z.pdf.

Korn:2019:RAA

[2351] Peter Korn. A regularity-aware algorithm for variational data assimilation of an idealized coupled atmosphere–ocean model. Journal of Scientific Computing, 79(2):748–786, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-0871-y;   https://link.
springer.com/content/pdf/10.1007/s10915-018-0871-y.pdf.

Wang:2019:FHO

[2352] Shaobo Wang, Shidong Jiang, and Jing Wang. Fast high-order integral equation methods for solving boundary value problems of two dimensional heat equation in complex geometry. Journal of Scientific Computing, 79(2):787–808, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0872-x;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0872-x.pdf.

Zhang:2019:NPI

[2353] Xue Zhang and Xiaoqun Zhang. A new proximal iterative hard thresholding method with extrapolation for ℓ_{0} minimization. Journal of Scientific Computing, 79(2):809–826, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0874-8;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0874-8.pdf.

Oltean:2019:PPP

[2354] Marius Oltean, Carlos F. Sopuerta, and Alessandro D. A. M. Spallicci. Particle-without-particle: a practical pseudospectral collocation method for linear partial differential equations with distributional sources. Journal of Scientific Computing, 79(2):827–866, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0873-9;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0873-9.pdf.

Bernal:2019:IMM

[2355] Francisco Bernal. An implementation of Milstein’s method for general bounded diffusions. Journal of Scientific Computing, 79(2):867– 890, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0884-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0884-6.pdf.

Lin:2019:HOR

[2356] Jianfang Lin, Rémi Abgrall, and Jianxian Qiu. High order residual distribution for steady state problems for hyperbolic conservation laws. Journal of Scientific Computing, 79(2):891–913, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0878-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0878-4.pdf.

Zhai:2019:AWG

[2357] Qilong Zhai, Hehu Xie, Ran Zhang, and Zhimin Zhang. Acceleration of weak Galerkin methods for the Laplacian eigenvalue problem. Journal of Scientific Computing, 79(2):914–934, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0877-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0877-5.pdf.

Hecht:2019:RPE

[2358] F. Hecht, Z. Mghazli, I. Naji, and J. E. Roberts. A residual a posteriori error estimators for a model for flow in porous media with fractures. Journal of Scientific Computing, 79(2):935–968, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0875-7;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0875-7.pdf.

Di:2019:FHM

[2359] Yana Di, Yuwei Fan, Zhenzhong Kou, Ruo Li, and Yanli Wang. Filtered hyperbolic moment method for the Vlasov equation. Journal of Scientific Computing, 79(2):969–991, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0882-8;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0882-8.pdf.

Trias:2019:SCR

[2360] F. X. Trias, D. Folch, A. Gorobets, and A. Oliva. Spectrally-consistent regularization of Navier–Stokes equations. Journal of Scientific Computing, 79(2):992–1014, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0880-x;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0880-x.pdf.

Huang:2019:TOU

[2361] Juntao Huang, Weifeng Zhao, and Chi-Wang Shu. A thirdorder unconditionally positivity–preserving scheme for production– destruction equations with applications to non-equilibrium flows. Journal of Scientific Computing, 79(2):1015–1056, May 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0881-9;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0881-9.pdf.

Yang:2019:ISU

[2362] Wenli Yang, Zhongyi Huang, and Wei Zhu. Image segmentation using the Cahn–Hilliard equation. Journal of Scientific Computing, 79(2):1057–1077, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-00899-7;   https://
link.springer.com/content/pdf/10.1007/s10915-018-00899-7.pdf.

Su:2019:OEE

[2363] Haiyan Su, Shipeng Mao, and Xinlong Feng. Optimal error estimates of penalty based iterative methods for steady incompressible magnetohydrodynamics equations with different viscosities. Journal of Scientific Computing, 79(2):1078–1110, May 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0883-7;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0883-7.pdf.

Cai:2019:HOS

[2364] Xiaofeng Cai, Wei Guo, and Jing-Mei Qiu. A high order semi-Lagrangian discontinuous Galerkin method for the two-dimensional incompressible Euler equations and the guiding center Vlasov model without operator splitting. Journal of Scientific Computing, 79(2):1111–1134, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0889-1;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0889-1.pdf.

Ayala:2019:AAL

[2365] Alan Ayala, Xavier Claeys, and Laura Grigori. ALORA: Affine lowrank approximations. Journal of Scientific Computing, 79(2):1135–1160, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0885-5;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0885-5.pdf.

Fuhrer:2019:OQD

[2366] Thomas Führer and Norbert Heuer. Optimal quasi-diagonal preconditioners for pseudodifferential operators of order minus two. Journal of Scientific Computing, 79(2):1161–1181, May 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
0887-3;   https://link.springer.com/content/pdf/10.1007/s10915-
018-0887-3.pdf.

Xu:2019:OSM

[2367] Yingxiang Xu and Xin Chen. Optimized Schwarz methods for the optimal control of systems governed by elliptic partial differential equations. Journal of Scientific Computing, 79(2):1182–1213, May 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0886-4;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0886-4.pdf.

Li:2019:MPS

[2368] Maojun Li, Haiyun Dong, Binbin Hu, and Liwei Xu. Maximum-principle-satisfying and positivity-preserving high order central DG methods on unstructured overlapping meshes for two-dimensional hyperbolic conservation laws. Journal of Scientific Computing, 79(3):1361–1388, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
00895-x;   https://link.springer.com/content/pdf/10.1007/s10915-
018-00895-x.pdf.

Kopriva:2019:FSP

[2369] David A. Kopriva, Florian J. Hindenlang, Thomas Bolemann, and Gregor J. Gassner. Free-stream preservation for curved geometrically nonconforming discontinuous Galerkin spectral elements. Journal of Scientific Computing, 79(3):1389–1408, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-00897-9;   https://
link.springer.com/content/pdf/10.1007/s10915-018-00897-9.pdf.

Yang:2019:EDC

[2370] Jie Yang, Weidong Zhao, and Tao Zhou. Explicit deferred correction methods for second-order forward backward stochastic differential equations. Journal of Scientific Computing, 79(3):1409–1432, June 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-018-00896-
w;   https://link.springer.com/content/pdf/10.1007/s10915-018-
00896-w.pdf.

Perrin:2019:EER

[2371] G. Perrin and G. Defaux. Efficient evaluation of reliability–oriented sensitivity indices. Journal of Scientific Computing, 79(3):1433–1455, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-018-
00898-8;   https://link.springer.com/content/pdf/10.1007/s10915-
018-00898-8.pdf.

Qi:2019:CCS

[2372] Dongping Qi and Alexander Vladimirsky. Corner cases, singularities, and dynamic factoring. Journal of Scientific Computing, 79(3):1456– 1476, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-019-00905-6;   https://link.springer.com/content/pdf/
10.1007/s10915-019-00905-6.pdf.

Herrmann:2019:ESF

[2373] S. Herrmann and C. Zucca. Exact simulation of the first-passage time of diffusions. Journal of Scientific Computing, 79(3):1477–1504, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-018-00900-
3;   https://link.springer.com/content/pdf/10.1007/s10915-018-
00900-3.pdf.

Solano:2019:HOH

[2374] Manuel Solano and Felipe Vargas. A high order HDG method for Stokes flow in curved domains. Journal of Scientific Computing, 79(3):1505–1533, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-00901-2;   https://
link.springer.com/content/pdf/10.1007/s10915-018-00901-2.pdf.

E:2019:MPN

[2375] Weinan E, Martin Hutzenthaler, Arnulf Jentzen, and Thomas Kruse. On multilevel Picard numerical approximations for high-dimensional nonlinear parabolic partial differential equations and high-dimensional nonlinear backward stochastic differential equations. Journal of Scientific Computing, 79(3):1534–1571, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-00903-0;   https://
link.springer.com/content/pdf/10.1007/s10915-018-00903-0.pdf.

Offner:2019:EBD

[2376] Philipp Öffner and Hendrik Ranocha. Error boundedness of discontinuous Galerkin methods with variable coefficients. Journal of Scientific Computing, 79(3):1572–1607, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-00902-1;   https://
link.springer.com/content/pdf/10.1007/s10915-018-00902-1.pdf.

Wang:2019:ENM

[2377] Xiang Wang, Xing Li, Lei-Hong Zhang, and Ren-Cang Li. An efficient numerical method for the symmetric positive definite secondorder cone linear complementarity problem. Journal of Scientific Computing, 79(3):1608–1629, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00907-4;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00907-4.pdf.

Dong:2019:MCM

[2378] Hao Dong, Xiaojing Zheng, Junzhi Cui, Yufeng Nie, Zhiqiang Yang, and Zihao Yang. Multiscale computational method for dynamic thermo-mechanical problems of composite structures with diverse periodic configurations in different subdomains. Journal of Scientific Computing, 79(3):1630–1666, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-018-00904-z;   https://
link.springer.com/content/pdf/10.1007/s10915-018-00904-z.pdf.

Chan-Wai-Nam:2019:MLS

[2379] Quentin Chan-Wai-Nam, Joseph Mikael, and Xavier Warin. Machine learning for semi linear PDEs. Journal of Scientific Computing, 79(3):1667–1712, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00908-3;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00908-3.pdf.

Du:2019:SLD

[2380] Yu Du and Zhimin Zhang. Supercloseness of linear DG–FEM and its superconvergence based on the polynomial preserving recovery for Helmholtz equation. Journal of Scientific Computing, 79(3):1713–1736, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-019-00906-5;   https://link.springer.com/content/pdf/
10.1007/s10915-019-00906-5.pdf.

Fu:2019:EDF

[2381] Guosheng Fu. An explicit divergence-free DG method for incompressible magnetohydrodynamics. Journal of Scientific Computing, 79(3):1737–1752, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00909-2;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00909-2.pdf.

Xie:2019:EDP

[2382] Jianqiang Xie and Zhiyue Zhang. An effective dissipation-preserving fourth-order difference solver for fractional-in-space nonlinear wave equations. Journal of Scientific Computing, 79(3):1753–1776, June 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00921-
6;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00921-6.pdf.

Cockburn:2019:IHM

[2383] Bernardo Cockburn, John R. Singler, and Yangwen Zhang. Interpolatory HDG method for parabolic semilinear PDEs. Journal of Scientific Computing, 79(3):1777–1800, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00911-8;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00911-8.pdf.

Wei:2019:JSM

[2384] Yunxia Wei and Yanping Chen. A Jacobi spectral method for solving multidimensional linear Volterra integral equation of the second kind. Journal of Scientific Computing, 79(3):1801–1813, June 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00912-
7;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00912-7.pdf.

Liu:2019:SIM

[2385] J. Liu, J. Sun, and T. Turner. Spectral indicator method for a non-selfadjoint Steklov eigenvalue problem. Journal of Scientific Computing, 79(3):1814–1831, June 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00913-6;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00913-6.pdf.

Lozano:2019:EPI

[2386] Carlos Lozano. Entropy production by implicit Runge–Kutta schemes. Journal of Scientific Computing, 79(3):1832–1853, June 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00914-
5;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00914-5.pdf.

Ghadimi:2019:GUO

[2387] Saeed Ghadimi, Guanghui Lan, and Hongchao Zhang. Generalized uniformly optimal methods for nonlinear programming. Journal of Scientific Computing, 79(3):1854–1881, June 2019. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00915-4;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00915-4.pdf.

Liao:2019:UCF

[2388] Hong lin Liao, Yonggui Yan, and Jiwei Zhang. Unconditional convergence of a fast two-level linearized algorithm for semilinear subdiffusion equations. Journal of Scientific Computing, 80(1):1–25, July 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00927-
0;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00927-0.pdf.

Epshteyn:2019:ENA

[2389] Yekaterina Epshteyn and Qing Xia. Efficient numerical algorithms based on difference potentials for chemotaxis systems in 3D. Journal of Scientific Computing, 80(1):26–59, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00928-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00928-z.pdf.

Feng:2019:VEM

[2390] Fang Feng, Weimin Han, and Jianguo Huang. Virtual element methods for elliptic variational inequalities of the second kind. Journal of Scientific Computing, 80(1):60–80, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00929-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00929-y.pdf.

Ding:2019:HOA

[2391] Hengfei Ding and Changpin Li. A high-0rder algorithm for time-Caputo-tempered partial differential equation with Riesz derivatives in two spatial dimensions. Journal of Scientific Computing, 80(1):81–109, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
00930-5;   https://link.springer.com/content/pdf/10.1007/s10915-
019-00930-5.pdf.

Kharazmi:2019:FSE

[2392] Ehsan Kharazmi and Mohsen Zayernouri. Fractional sensitivity equation method: Application to fractional model construction. Journal of Scientific Computing, 80(1):110–140, July 2019. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00935-0;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00935-0.pdf.

Alvarez:2019:NMF

[2393] Mario Alvarez, Gabriel N. Gatica, Bryan Gomez-Vargas, and Ricardo Ruiz-Baier. New mixed finite element methods for natural convection with phase-change in porous media. Journal of Scientific Computing, 80(1):141–174, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00931-4;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00931-4.pdf.

Friedrich:2019:ESS

[2394] Lucas Friedrich, Gero Schnücke, Andrew R. Winters, David C. Del Rey Fernández, Gregor J. Gassner, and Mark H. Carpenter. Entropy stable space–time discontinuous Galerkin schemes with summationby-parts property for hyperbolic conservation laws. Journal of Scientific Computing, 80(1):175–222, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00933-2;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00933-2.pdf.

Zhu:2019:NAP

[2395] Guangpu Zhu, Jisheng Kou, Shuyu Sun, Jun Yao, and Aifen Li. Numerical approximation of a phase-field surfactant model with fluid flow. Journal of Scientific Computing, 80(1):223–247, July 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00934-
1;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00934-1.pdf.

Fang:2019:ESG

[2396] Zhiwei Fang, Jichun Li, Tao Tang, and Tao Zhou. Efficient stochastic Galerkin methods for Maxwell’s equations with random inputs. Journal of Scientific Computing, 80(1):248–267, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00936-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00936-z.pdf.

Li:2019:AOD

[2397] Ruo Li, Pingbing Ming, Ziyuan Sun, and Zhijian Yang. An arbitraryorder discontinuous Galerkin method with one unknown per element.

Journal of Scientific Computing, 80(1):268–288, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00937-
y;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00937-y.pdf.

Vargas:2019:LTS

[2398] Arturo Vargas, Thomas Hagstrom, Jesse Chan, and Tim Warburton. Leapfrog time-stepping for Hermite methods. Journal of Scientific Computing, 80(1):289–314, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00938-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00938-x.pdf.

Jiang:2019:PCE

[2399] Nan Jiang. A pressure-correction ensemble scheme for computing evolutionary Boussinesq equations. Journal of Scientific Computing, 80(1): 315–350, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-019-00939-w;   https://link.springer.com/content/
pdf/10.1007/s10915-019-00939-w.pdf.

Jensen:2019:DPF

[2400] Max Jensen, Ananta K. Majee, Andreas Prohl, and Christian Schellnegger. Dynamic programming for finite ensembles of nanomagnetic particles. Journal of Scientific Computing, 80(1):351–375, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00940-
3;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00940-3.pdf.

Pazner:2019:AES

[2401] Will Pazner and Per-Olof Persson. Analysis and entropy stability of the line-based discontinuous Galerkin method. Journal of Scientific Computing, 80(1):376–402, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00942-1;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00942-1.pdf.

Li:2019:LGF

[2402] Dongfang Li, Chengda Wu, and Zhimin Zhang. Linearized Galerkin FEMs for nonlinear time fractional parabolic problems with non-smooth solutions in time direction. Journal of Scientific Computing, 80(1):403– 419, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-019-00943-0;   https://link.springer.com/content/pdf/
10.1007/s10915-019-00943-0.pdf.

Chen:2019:HMT

[2403] Gang Chen, Peter Monk, and Yangwen Zhang. An HDG method for the time-dependent drift–diffusion model of semiconductor devices. Journal of Scientific Computing, 80(1):420–443, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00945-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00945-y.pdf.

Prohl:2019:ACS

[2404] Andreas Prohl and Christian Schellnegger. Adaptive concepts for stochastic partial differential equations. Journal of Scientific Computing, 80(1):444–474, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00944-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00944-z.pdf.

Yadav:2019:MAP

[2405] N. K. Yadav, J. H. M. ten Thije Boonkkamp, and W. L. IJzerman. A Monge–Ampère problem with non-quadratic cost function to compute freeform lens surfaces. Journal of Scientific Computing, 80(1):475–499, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00948-9;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00948-9.pdf.

Jing:2019:SOL

[2406] Xiaobo Jing, Jun Li, Xueping Zhao, and Qi Wang. Second order linear energy stable schemes for Allen–Cahn equations with nonlocal constraints. Journal of Scientific Computing, 80(1):500–537, July 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00946-
x;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00946-x.pdf.

Cheng:2019:NAD

[2407] Yuanzhen Cheng, Alina Chertock, Michael Herty, Alexander Kurganov, and Tong Wu. A new approach for designing moving-water equilibria preserving schemes for the shallow water equations. Journal of Scientific Computing, 80(1):538–554, July 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00947-w;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00947-w.pdf.

Duan:2019:ECS

[2408] Beiping Duan and Zhoushun Zheng. An exponentially convergent scheme in time for time fractional diffusion equations with non-smooth initial data. Journal of Scientific Computing, 80(2):717–742, August 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00953-
y;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00953-y.pdf.

Jallepalli:2019:EAL

[2409] Ashok Jallepalli and Robert M. Kirby. Efficient algorithms for the line-SIAC filter. Journal of Scientific Computing, 80(2):743–761, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
00954-x;   https://link.springer.com/content/pdf/10.1007/s10915-
019-00954-x.pdf.

Churchill:2019:DEN

[2410] Victor Churchill and Anne Gelb. Detecting edges from non-uniform Fourier data via sparse Bayesian learning. Journal of Scientific Computing, 80(2):762–783, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00955-w;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00955-w.pdf.

Higueras:2019:SSP

[2411] I. Higueras and T. Roldán. Strong stability preserving properties of composition Runge–Kutta schemes. Journal of Scientific Computing, 80(2):784–807, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00956-9;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00956-9.pdf.

Jacobs:2019:ASH

[2412] Matthew Jacobs and Songting Luo. Asymptotic solutions for high frequency Helmholtz equations in anisotropic media with Hankel functions.

Journal of Scientific Computing, 80(2):808–833, August 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00957-
8;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00957-8.pdf.

Uchiumi:2019:VIE

[2413] Shinya Uchiumi. A viscosity-independent error estimate of a pressurestabilized Lagrange–Galerkin scheme for the Oseen problem. Journal of Scientific Computing, 80(2):834–858, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00958-7;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00958-7.pdf.

Dupuy:2019:DEC

[2414] Mi-Song Dupuy. Discretization error cancellation in the plane-wave approximation of periodic Hamiltonians with Coulomb singularities. Journal of Scientific Computing, 80(2):859–877, August 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00959-
6;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00959-6.pdf.

Bremer:2019:PCH

[2415] Maximilian Bremer, Kazbek Kazhyken, Hartmut Kaiser, Craig Michoski, and Clint Dawson. Performance comparison of HPX versus traditional parallelization strategies for the discontinuous Galerkin method. Journal of Scientific Computing, 80(2):878–902, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00960-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00960-z.pdf.

Wu:2019:NAG

[2416] Chengda Wu and Weiwei Sun. New analysis of Galerkin FEMs for miscible displacement in porous media. Journal of Scientific Computing, 80(2):903–923, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00963-w;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00963-w.pdf.

Berthon:2019:IHR

[2417] Christophe Berthon, Arnaud Duran, Françoise Foucher, Khaled Saleh, and Jean De Dieu Zabsonré. Improvement of the hydrostatic reconstruction scheme to get fully discrete entropy inequalities. Journal of Scientific Computing, 80(2):924–956, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00961-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00961-y.pdf.

Luo:2019:CAP

[2418] Hao Luo, Binjie Li, and Xiaoping Xie. Convergence analysis of a Petrov– Galerkin method for fractional wave problems with nonsmooth data. Journal of Scientific Computing, 80(2):957–992, August 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00962-
x;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00962-x.pdf.

Zhang:2019:PEA

[2419] Chenyang Zhang, Huipo Liu, and Zhaojie Zhou. A priori error analysis for time-stepping discontinuous Galerkin finite element approximation of time fractional optimal control problem. Journal of Scientific Computing, 80(2):993–1018, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00964-9;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00964-9.pdf.

Qiu:2019:AHM

[2420] Weifeng Qiu and Ke Shi. Analysis on an HDG method for the p-Laplacian equations. Journal of Scientific Computing, 80(2):1019–1032, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
00967-6;   https://link.springer.com/content/pdf/10.1007/s10915-
019-00967-6.pdf.

Chen:2019:RPT

[2421] Xu Chen, Si-Wen Deng, and Siu-Long Lei. A robust preconditioner for two-dimensional conservative space–fractional diffusion equations on convex domains. Journal of Scientific Computing, 80(2):1033–1057, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
00966-7;   https://link.springer.com/content/pdf/10.1007/s10915-
019-00966-7.pdf.

Shepherd:2019:ASI

[2422] David Shepherd, James Miles, Matthias Heil, and Milan Mihajlović. An adaptive step implicit midpoint rule for the time integration of Newton’s linearisations of non-linear problems with applications in micromagnetics. Journal of Scientific Computing, 80(2):1058–1082, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00965-8;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00965-8.pdf.

Gao:2019:SSI

[2423] Huadong Gao, Lili Ju, and Wen Xie. A stabilized semi-implicit Euler gauge-invariant method for the time-dependent Ginzburg–Landau equations. Journal of Scientific Computing, 80(2):1083–1115, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00968-
5;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00968-5.pdf.

Xiang:2019:NPV

[2424] Yan-Fei Xiang, Yan-Fei Jing, and Ting-Zhu Huang. A new projected variant of the deflated block conjugate gradient method. Journal of Scientific Computing, 80(2):1116–1138, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00969-4;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00969-4.pdf.

Yuan:2019:GOO

[2425] Honglin Yuan, Xiaoyi Gu, Rongjie Lai, and Zaiwen Wen. Global optimization with orthogonality constraints via stochastic diffusion on manifold. Journal of Scientific Computing, 80(2):1139–1170, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00971-
w;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00971-w.pdf.

Qi:2019:OEE

[2426] Ruisheng Qi and Xiaojie Wang. Optimal error estimates of Galerkin finite element methods for stochastic Allen–Cahn equation with additive noise. Journal of Scientific Computing, 80(2):1171–1194, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00973-
8;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00973-8.pdf.

Chow:2019:AHJ

[2427] Yat Tin Chow, Wuchen Li, Stanley Osher, and Wotao Yin. Algorithm for Hamilton–Jacobi equations in density space via a generalized Hopf formula. Journal of Scientific Computing, 80(2):1195–1239, August 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00972-
9;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00972-9.pdf.

Schumaker:2019:SEP

[2428] Larry L. Schumaker. Solving elliptic PDE’s on domains with curved boundaries with an immersed penalized boundary method. Journal of Scientific Computing, 80(3):1369–1394, September 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00978-
3;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00978-3.pdf.

Zheng:2019:EFV

[2429] Xiangcheng Zheng, Huan Liu, Hong Wang, and Hongfei Fu. An efficient finite volume method for nonlinear distributed-order space-fractional diffusion equations in three space dimensions. Journal of Scientific Computing, 80(3):1395–1418, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00979-2;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00979-2.pdf.

Arrigo:2019:NBP

[2430] Francesca Arrigo, Desmond J. Higham, and Vanni Noferini. Nonbacktracking PageRank. Journal of Scientific Computing, 80(3):1419– 1437, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-019-00981-8;   https://link.springer.com/content/
pdf/10.1007/s10915-019-00981-8.pdf.

Thong:2019:ASE

[2431] Duong Viet Thong, Nguyen The Vinh, and Yeol Je Cho. Accelerated subgradient extragradient methods for variational inequality problems.

Journal of Scientific Computing, 80(3):1438–1462, September 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00984-
5;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00984-5.pdf.

Helzel:2019:NAM

[2432] Christiane Helzel, David Kerkmann, and Leonardo Scandurra. A new ADER method inspired by the active flux method. Journal of Scientific Computing, 80(3):1463–1497, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00988-1;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00988-1.pdf.

Feng:2019:MMM

[2433] Xiaobing Feng, Junshan Lin, and Cody Lorton. A multi-modes Monte Carlo interior penalty discontinuous Galerkin method for the timeharmonic Maxwell’s equations with random coefficients. Journal of Scientific Computing, 80(3):1498–1528, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00986-3;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00986-3.pdf.

Cai:2019:FSC

[2434] Haotao Cai. A fractional spectral collocation for solving second kind nonlinear Volterra integral equations with weakly singular kernels. Journal of Scientific Computing, 80(3):1529–1548, September 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00987-
2;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00987-2.pdf.

Puthawala:2019:DFO

[2435] Michael A. Puthawala, Cory D. Hauck, and Stanley J. Osher. Diagnosing forward operator error using optimal transport. Journal of Scientific Computing, 80(3):1549–1576, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00989-0;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00989-0.pdf.

Anaya:2019:AAV

[2436] V. Anaya, A. Bouharguane, D. Mora, C. Reales, R. Ruiz-Baier, N. Seloula, and H. Torres. Analysis and approximation of a vorticity– velocity–pressure formulation for the Oseen equations. Journal of Scientific Computing, 80(3):1577–1606, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00990-7;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00990-7.pdf.

Lyu:2019:HOM

[2437] Pin Lyu and Seakweng Vong. A high-order method with a temporal nonuniform mesh for a time-fractional Benjamin–Bona–Mahony equation. Journal of Scientific Computing, 80(3):1607–1628, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00991-
6;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00991-6.pdf.

Jiang:2019:LIL

[2438] Chaolong Jiang, Wenjun Cai, and Yushun Wang. A linearly implicit and local energy-preserving scheme for the sine–Gordon equation based on the invariant energy quadratization approach. Journal of Scientific Computing, 80(3):1629–1655, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01001-5;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01001-5.pdf.

Xu:2019:ESS

[2439] Xiang Xu and Yanxiang Zhao. Energy stable semi-implicit schemes for Allen–Cahn–Ohta–Kawasaki model in binary system. Journal of Scientific Computing, 80(3):1656–1680, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00993-4;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00993-4.pdf.

Xie:2019:HOK

[2440] Yaning Xie, Wenjun Ying, and Wei-Cheng Wang. A high-order kernelfree boundary integral method for the biharmonic equation on irregular domains. Journal of Scientific Computing, 80(3):1681–1699, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01000-6;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01000-6.pdf.

Mitchell:2019:PEE

[2441] William F. Mitchell and John S. Villarrubia. An a posteriori error estimate for scanning electron microscope simulation with adaptive mesh refinement. Journal of Scientific Computing, 80(3):1700–1715, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
00995-2;   https://link.springer.com/content/pdf/10.1007/s10915-
019-00995-2.pdf.

Chen:2019:OCA

[2442] Wenbin Chen, Yichao Zhang, Weijia Li, Yanqiu Wang, and Yue Yan. Optimal convergence analysis of a second order scheme for a thin film model without slope selection. Journal of Scientific Computing, 80(3):1716–1730, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00999-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00999-y.pdf.

Chou:2019:TTF

[2443] Lot-Kei Chou and Siu-Long Lei. Tensor-train format solution with preconditioned iterative method for high dimensional time-dependent space-fractional diffusion equations with error analysis. Journal of Scientific Computing, 80(3):1731–1763, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00994-3;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00994-3.pdf.

Ruf:2019:OCR

[2444] Adrian M. Ruf, Espen Sande, and Susanne Solem. The optimal convergence rate of monotone schemes for conservation laws in the Wasserstein distance. Journal of Scientific Computing, 80(3):1764–1776, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
00996-1;   https://link.springer.com/content/pdf/10.1007/s10915-
019-00996-1.pdf.

Carrillo:2019:HMT

[2445] J. A. Carrillo, N. Kolbe, and M. Lukácová-Medvid’ová. A hybrid mass transport finite element method for Keller–Segel type systems. Journal of Scientific Computing, 80(3):1777–1804, September 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00997-
0;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00997-0.pdf.

Manzini:2019:HOM

[2446] Gianmarco Manzini, Gianluca Maguolo, and Mario Putti. The high-order mixed mimetic finite difference method for time-dependent diffusion problems. Journal of Scientific Computing, 80(3):1805–1830, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01002-4;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01002-4.pdf.

Manzini:2019:CHO

[2447] Gianmarco Manzini, Gianluca Maguolo, and Mario Putti. Correction to: The high-order mixed mimetic finite difference method for time-dependent diffusion problems. Journal of Scientific Computing, 80(3):1831, September 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01022-0;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01022-0.pdf.

Lelievre:2019:PSI

[2448] Tony Lelièvre, Simona Perotto, Gianluigi Rozza, Daniele A. Di Pietro, Alexandre Ern, and Luca Formaggia. Preface: Special issue on model reduction. Journal of Scientific Computing, 81(1):1–2, October 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01037-
7;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01037-7.pdf.

Afkham:2019:SPM

[2449] Babak Maboudi Afkham and Jan S. Hesthaven. Structure-preserving model-reduction of dissipative Hamiltonian systems. Journal of Scientific Computing, 81(1):3–21, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-018-0653-6;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0653-6.pdf.

Aguado:2019:TRN

[2450] José V. Aguado, Domenico Borzacchiello, Kiran S. Kollepara, Francisco Chinesta, and Antonio Huerta. Tensor representation of nonlinear models using cross approximations. Journal of Scientific Computing, 81(1):22–47, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00917-2;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00917-2.pdf.

Freya:2019:PBM

[2451] Bachmann Freya, Beermann Dennis, Lu Jianjie, and Volkwein Stefan. POD-based mixed-integer optimal control of the heat equation. Journal of Scientific Computing, 81(1):48–75, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00924-3;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00924-3.pdf.

Gunzburger:2019:IDL

[2452] Max Gunzburger, Michael Schneier, Clayton Webster, and Guannan Zhang. An improved discrete least-squares/reduced-basis method for parameterized elliptic PDEs. Journal of Scientific Computing, 81(1):76– 91, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-018-0661-6;   https://link.springer.com/content/pdf/
10.1007/s10915-018-0661-6.pdf.

Kergrene:2019:GOV

[2453] Kenan Kergrene, Ludovic Chamoin, Marc Laforest, and Serge Prudhomme. On a goal-oriented version of the proper generalized decomposition method. Journal of Scientific Computing, 81(1):92–111, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
00918-1;   https://link.springer.com/content/pdf/10.1007/s10915-
019-00918-1.pdf.

Pichi:2019:RBA

[2454] Federico Pichi and Gianluigi Rozza. Reduced basis approaches for parametrized bifurcation problems held by non-linear von Kármán equations. Journal of Scientific Computing, 81(1):112–135, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01003-
3;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01003-3.pdf.

Venturi:2019:WPM

[2455] Luca Venturi, Francesco Ballarin, and Gianluigi Rozza. A weighted POD method for elliptic PDEs with random inputs. Journal of Scientific Computing, 81(1):136–153, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-018-0830-7;   https://link.springer.
com/content/pdf/10.1007/s10915-018-0830-7.pdf.

Cheng:2019:TOE

[2456] Kelong Cheng, Zhonghua Qiao, and Cheng Wang. A third order exponential time differencing numerical scheme for no-slope-selection epitaxial thin film model with energy stability. Journal of Scientific Computing, 81(1):154–185, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01008-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01008-y.pdf.

Ren:2019:GCD

[2457] Jiong Ren, Gang Wang, and Mingsheng Ma. A group of CFL-dependent flux-limiters to control the numerical dissipation in multi-stage unsteady calculation. Journal of Scientific Computing, 81(1):186–216, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01010-4;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01010-4.pdf.

Cho:2019:SOB

[2458] Hyuntae Cho, Heejae Han, Byungjoon Lee, Youngsoo Ha, and Myungjoo Kang. A second-order boundary condition capturing method for solving the elliptic interface problems on irregular domains. Journal of Scientific Computing, 81(1):217–251, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01016-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01016-y.pdf.

Yin:2019:CIB

[2459] Guojian Yin. A contour-integral based method with Schur–Rayleigh– Ritz procedure for generalized eigenvalue problems. Journal of Scientific Computing, 81(1):252–270, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01014-0;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01014-0.pdf.

Pranjal:2019:BIS

[2460] Pranjal and David Silvester. Balanced iterative solvers for linear nonsymmetric systems and nonlinear systems with PDE origins:

Efficient black–Box stopping criteria. Journal of Scientific Computing, 81(1):271–290, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01018-w;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01018-w.pdf.

Lu:2019:RMP

[2461] Peipei Lu and Xuejun Xu. A robust multilevel preconditioner based on a domain decomposition method for the Helmholtz equation. Journal of Scientific Computing, 81(1):291–311, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01015-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01015-z.pdf.

Ayadi:2019:SPS

[2462] Mekki Ayadi, Hela Ayed, Leonardo Baffico, and Taoufik Sassi. Stokes problem with slip boundary conditions of friction type: Error analysis of a four-field mixed variational formulation. Journal of Scientific Computing, 81(1):312–341, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01017-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01017-x.pdf.

Wang:2019:NNE

[2463] Nan Wang and Jinru Chen. A nonconforming Nitsche’s extended finite element method for Stokes interface problems. Journal of Scientific Computing, 81(1):342–374, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01019-9;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01019-9.pdf.

Chen:2019:KBM

[2464] Meng Chen and Leevan Ling. Kernel-based meshless collocation methods for solving coupled bulk-surface partial differential equations. Journal of Scientific Computing, 81(1):375–391, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01020-2;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01020-2.pdf.

Moradi:2019:SSP

[2465] Afsaneh Moradi, Javad Farzi, and Ali Abdi. Strong stability preserving second derivative general linear methods. Journal of Scientific Computing, 81(1):392–435, October 2019. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01021-1;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01021-1.pdf.

He:2019:PPF

[2466] Dongdong He, Kejia Pan, and Xiaoqiang Yue. A positivity preserving and free energy dissipative difference scheme for the Poisson–Nernst–Planck system. Journal of Scientific Computing, 81(1):436–458, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01025-
x;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01025-x.pdf.

Chan:2019:SSE

[2467] Jesse Chan. Skew-symmetric entropy stable modal discontinuous Galerkin formulations. Journal of Scientific Computing, 81(1):459–485, October 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-019-01026-w;   https://link.springer.com/content/pdf/
10.1007/s10915-019-01026-w.pdf.

Chen:2019:CED

[2468] Gang Chen, Guosheng Fu, John R. Singler, and Yangwen Zhang. A class of embedded DG methods for Dirichlet boundary control of convection diffusion PDEs. Journal of Scientific Computing, 81(2):623–648, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01043-9;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01043-9.pdf.

Zhao:2019:EAC

[2469] Guo-Yan Zhao, Ming-Bo Sun, Yong Mei, Liang Li, Hong-Bo Wang, Guang-Xin Li, Yuan Liu, Yong-Chao Sun, and Chang-Hai Liang. An efficient adaptive central-upwind WENO–CU6 numerical scheme with a new sensor. Journal of Scientific Computing, 81(2):649–670, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01035-
9;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01035-9.pdf.

Chuenjarern:2019:FAL

[2470] Nattaporn Chuenjarern and Yang Yang. Fourier analysis of local discontinuous Galerkin methods for linear parabolic equations on overlapping meshes. Journal of Scientific Computing, 81(2):671–688, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01030-
0;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01030-0.pdf.

Ma:2019:SOA

[2471] Qiang Ma, Hao Wang, Zhiqiang Yang, Zhihui Li, and Junzhi Cui. Secondorder asymptotic analysis and computations of axially and spherically symmetric piezoelectric problems for composite structures. Journal of Scientific Computing, 81(2):689–731, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01041-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01041-x.pdf.

Vevek:2019:EHM

[2472] U. S. Vevek, B. Zang, and T. H. New. An efficient hybrid method for solving Euler equations. Journal of Scientific Computing, 81(2):732–762, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01033-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01033-x.pdf.

Ding:2019:PRS

[2473] Weiyang Ding, Michael K. Ng, and Wenxing Zhang. A Peaceman– Rachford splitting method with monotone plus skew-symmetric splitting for nonlinear saddle point problems. Journal of Scientific Computing, 81(2):763–788, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01034-w;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01034-w.pdf.

Liu:2019:UES

[2474] Hailiang Liu and Peimeng Yin. Unconditionally energy stable DG schemes for the Swift–Hohenberg equation. Journal of Scientific Computing, 81(2):789–819, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01038-6;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01038-6.pdf.

Bohm:2019:MES

[2475] Marvin Bohm, Sven Schermeng, Andrew R. Winters, Gregor J. Gassner, and Gustaaf B. Jacobs. Multi-element SIAC filter for shock capturing applied to high-order discontinuous Galerkin spectral element methods. Journal of Scientific Computing, 81(2):820–844, November 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01036-
8;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01036-8.pdf.

Kundu:2019:GSB

[2476] Sudeep Kundu and Amiya Kumar Pani. Global stabilization of BBM–Burgers’ type equations by nonlinear boundary feedback control laws: Theory and finite element error analysis. Journal of Scientific Computing, 81(2):845–880, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01039-5;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01039-5.pdf.

Kirk:2019:APR

[2477] Keegan L. A. Kirk and Sander Rhebergen. Analysis of a pressure-robust hybridized discontinuous Galerkin method for the stationary Navier– Stokes equations. Journal of Scientific Computing, 81(2):881–897, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-019-01040-y;   https://link.springer.com/content/pdf/
10.1007/s10915-019-01040-y.pdf.

Wang:2019:MSA

[2478] Yahui Wang, Yulong Du, Kunlei Zhao, and Li Yuan. Modified stencil approximations for fifth-order weighted essentially non-oscillatory schemes. Journal of Scientific Computing, 81(2):898–922, November 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01042-
w;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01042-w.pdf.

Li:2019:TSF

[2479] Zheng Li, Guohui Song, and Yuesheng Xu. A two-step fixedpoint proximity algorithm for a class of non-differentiable optimization models in machine learning. Journal of Scientific Computing, 81(2):923–940, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01045-7;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01045-7.pdf.

Ding:2019:LRT

[2480] Meng Ding, Ting-Zhu Huang, Teng-Yu Ji, Xi-Le Zhao, and Jing-Hua Yang. Low-rank tensor completion using matrix factorization based on tensor train rank and total variation. Journal of Scientific Computing, 81(2):941–964, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01044-8;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01044-8.pdf.

Zhai:2019:EAN

[2481] Shuying Zhai, Dongling Wang, Zhifeng Weng, and Xuan Zhao. Error analysis and numerical simulations of Strang splitting method for space fractional nonlinear Schrödinger equation. Journal of Scientific Computing, 81(2):965–989, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01050-w;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01050-w.pdf.

daVeiga:2019:SCV

[2482] L. Beirão da Veiga, D. Mora, and G. Vacca. The Stokes complex for virtual elements with application to Navier–Stokes flows. Journal of Scientific Computing, 81(2):990–1018, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01049-3;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01049-3.pdf.

Bertoluzza:2019:FSE

[2483] S. Bertoluzza and S. Falletta. FEM solution of exterior elliptic problems with weakly enforced integral non reflecting boundary conditions. Journal of Scientific Computing, 81(2):1019–1049, November 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01048-
4;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01048-4.pdf.

Nordstrom:2019:DTS

[2484] Jan Nordström and Andrea A. Ruggiu. Dual time-stepping using second derivatives. Journal of Scientific Computing, 81(2):1050–1071, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01047-5;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01047-5.pdf.

Xie:2019:CEE

[2485] Hehu Xie and Manting Xie. Computable error estimates for ground state solution of Bose–Einstein condensates. Journal of Scientific Computing, 81(2):1072–1087, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01051-9;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01051-9.pdf.

Shen:2019:EST

[2486] Jie Shen and Chang-Tao Sheng. An efficient space–time method for time fractional diffusion equation. Journal of Scientific Computing, 81(2):1088–1110, November 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01052-8;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01052-8.pdf.

Wahlsten:2019:SIF

[2487] Markus Wahlsten and Jan Nordström. On stochastic investigation of flow problems using the viscous Burgers’ equation as an example. Journal of Scientific Computing, 81(2):1111–1117, November 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01053-
7;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01053-7.pdf.

Chertock:2019:PSI

[2488] Alina Chertock, Adi Ditkowski, Anne Gelb, Sigal Gottlieb, and Semyon Tsynkov. Preface to the special issue in memory of Professor saul abarbanel. Journal of Scientific Computing, 81(3):1119–1123, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01084-0;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01084-0.pdf.

Gustafsson:2019:SAH

[2489] Bertil Gustafsson. Saul abarbanel; half a century of scientific work. Journal of Scientific Computing, 81(3):1124–1135, December 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00932-
3;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00932-3.pdf.

Ryabenkii:2019:IFA

[2490] V. S. Ryaben’kii and V. A. Torgashov. An iteration-free approach to solving the Navier–Stokes equations by implicit finite difference schemes in the vorticity-stream function formulation. Journal of Scientific Computing, 81(3):1136–1149, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00926-1;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00926-1.pdf.

Li:2019:SIA

[2491] Xiaozhou Li, Jennifer K. Ryan, Robert M. Kirby, and Kees Vuik. Smoothness-increasing accuracy-conserving (SIAC) filtering for discontinuous Galerkin solutions over nonuniform meshes: Superconvergence and optimal accuracy. Journal of Scientific Computing, 81(3):1150–1180, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00920-7;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00920-7.pdf.

Smith:2019:CHO

[2492] F. Smith, S. Tsynkov, and E. Turkel. Compact high order accurate schemes for the three dimensional wave equation. Journal of Scientific Computing, 81(3):1181–1209, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00970-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00970-x.pdf.

Liu:2019:COF

[2493] Yong Liu, Tianheng Chen, Yanlai Chen, and Chi-Wang Shu. Certified offline-free reduced basis (COFRB) methods for stochastic differential equations driven by arbitrary types of noise. Journal of Scientific Computing, 81(3):1210–1239, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00976-5;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00976-5.pdf.

Gelb:2019:NAC

[2494] Anne Gelb, X. Hou, and Q. Li. Numerical analysis for conservation laws using l_{1} minimization. Journal of Scientific Computing, 81(3):1240–1265, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00982-7;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00982-7.pdf.

Sever:2019:SUC

[2495] Michael Sever. A source of uncertainty in computed discontinuous flows. Journal of Scientific Computing, 81(3):1266–1296, December 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-00992-
5;   https://link.springer.com/content/pdf/10.1007/s10915-019-
00992-5.pdf.

Liu:2019:ODT

[2496] Xin Liu, Alina Chertock, Alexander Kurganov, and Karlan Wolfkill. One-dimensional/two-dimensional coupling approach with quadrilateral confluence region for modeling river systems. Journal of Scientific Computing, 81(3):1297–1328, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00985-4;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00985-4.pdf.

Wang:2019:GSP

[2497] Yinghua Wang, Bao-Shan Wang, and Wai Sun Don. Generalized sensitivity parameter free fifth order WENO finite difference scheme with Z-type weights. Journal of Scientific Computing, 81(3):1329–1358, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-00998-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-00998-z.pdf.

Sjogreen:2019:ESM

[2498] Björn Sjögreen and H. C. Yee. Entropy stable method for the Euler equations revisited: Central differencing via entropy splitting and SBP. Journal of Scientific Computing, 81(3):1359–1385, December 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01013-
1;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01013-1.pdf.

Ben-Artzi:2019:CCS

[2499] Matania Ben-Artzi, Jean-Pierre Croisille, and Dalia Fishelov. A Cartesian compact scheme for the Navier–Stokes equations in streamfunction formulation in irregular domains. Journal of Scientific Computing, 81(3):1386–1408, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01012-2;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01012-2.pdf.

Peles:2019:ATS

[2500] O. Peles and E. Turkel. Adaptive time steps for compressible flows based on dual-time stepping and a RK/implicit smoother. Journal of Scientific Computing, 81(3):1409–1428, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01024-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01024-y.pdf.

Song:2019:SOT

[2501] Jiah Song and Smadar Karni. A second order traffic flow model with lane changing. Journal of Scientific Computing, 81(3):1429–1445, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01023-z;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01023-z.pdf.

Isherwood:2019:SSP

[2502] Leah Isherwood, Zachary J. Grant, and Sigal Gottlieb. Strong stability preserving integrating factor two-step Runge–Kutta methods. Journal of Scientific Computing, 81(3):1446–1471, December 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01046-
6;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01046-6.pdf.

Harari:2019:CSN

[2503] Isaac Harari and Uri Albocher. Complementary solutions of Nitsche’s method. Journal of Scientific Computing, 81(3):1472–1492, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01066-2;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01066-2.pdf.

Meltzer:2019:AAE

[2504] A. Y. Meltzer. An accurate approximation of exponential integrators for the Schrödinger equation. Journal of Scientific Computing, 81(3):1493–1508, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01075-1;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01075-1.pdf.

Hagstrom:2019:DGD

[2505] T. Hagstrom, J. W. Banks, B. B. Buckner, and K. Juhnke. Discontinuous Galerkin difference methods for symmetric hyperbolic systems. Journal of Scientific Computing, 81(3):1509–1526, December 2019. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01070-
6;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01070-6.pdf.

Pei:2019:HST

[2506] Chaoxu Pei, Mehdi Vahab, Mark Sussman, and M. Yousuff Hussaini. A hierarchical space–time spectral element and moment-of-fluid method for improved capturing of vortical structures in incompressible multi-phase/multi-material flows. Journal of Scientific Computing, 81(3):1527–1566, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01087-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01087-x.pdf.

Chang:2019:CGD

[2507] Kao-Pu Chang, Jessica M. Libertini, and Steven Seay. Comparing gradient descent with automatic differentiation and particle swarm optimization techniques for estimating tumor blood flow parameters in contrast-enhanced imaging. Journal of Scientific Computing, 81(3):1567–1576, December 2019. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01099-7;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01099-7.pdf.

Yu:2020:FSP

[2508] Yize Yu, Yan Jiang, and Mengping Zhang. Free-stream preserving finite difference schemes for ideal magnetohydrodynamics on curvilinear meshes. Journal of Scientific Computing, 82(1):??, January 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01125-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01125-z.pdf.

Chen:2020:SUW

[2509] Anqi Chen, Yingda Cheng, Yong Liu, and Mengping Zhang. Superconvergence of ultra-weak discontinuous Galerkin methods for the linear Schrödinger equation in one dimension. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01124-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01124-0.pdf.

Ding:2020:CFE

[2510] Qianqian Ding and Shipeng Mao. A convergent finite element method for the compressible magnetohydrodynamics system. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01129-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01129-9.pdf.

Xu:2020:LPM

[2511] Fei Xu and Qiumei Huang. Local and parallel multigrid method for nonlinear eigenvalue problems. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01128-w;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01128-w.pdf.

Hao:2020:HMA

[2512] Wenrui Hao, Jan Hesthaven, Guang Lin, and Bin Zheng. A homotopy method with adaptive basis selection for computing multiple solutions of differential equations. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01123-1;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01123-1.pdf.

Li:2020:HOC

[2513] Xue-Li Li and Yu-Xin Ren. High order compact generalized finite difference methods for solving inviscid compressible flows. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01105-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01105-y.pdf.

Chen:2020:EAN

[2514] Sheng Chen and Jie Shen. An efficient and accurate numerical method for the spectral fractional Laplacian equation. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01122-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01122-x.pdf.

Lundquist:2020:SAF

[2515] Tomas Lundquist and Jan Nordström. Stable and accurate filtering procedures. Journal of Scientific Computing, 82(1):??, January 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01116-
9;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01116-9.pdf.

Meena:2020:PPF

[2516] Asha Kumari Meena, Rakesh Kumar, and Praveen Chandrashekar. Positivity-preserving finite difference WENO scheme for ten-moment equations with source term. Journal of Scientific Computing, 82(1): ??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-019-01110-1;   https://link.springer.com/content/pdf/
10.1007/s10915-019-01110-1.pdf.

Oberman:2020:PDE

[2517] Adam Oberman and Tiago Salvador. A partial differential equation obstacle problem for the level set approach to visibility. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01106-x;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01106-x.pdf.

Zaky:2020:SIG

[2518] Mahmoud A. Zaky, Ahmed S. Hendy, and Jorge E. Macías-Díaz. Semiimplicit Galerkin–Legendre spectral schemes for nonlinear time-space fractional diffusion–reaction equations with smooth and nonsmooth solutions. Journal of Scientific Computing, 82(1):??, January 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01117-
8;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01117-8.pdf.

Zhao:2020:NHS

[2519] Lina Zhao and Eun-Jae Park. A new hybrid staggered discontinuous Galerkin method on general meshes. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01119-6;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01119-6.pdf.

Gao:2020:RSP

[2520] Shangqi Gao and Qibin Fan. Robust Schatten-p norm based approach for tensor completion. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01108-9;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01108-9.pdf.

Huang:2020:SFE

[2521] Chaobao Huang and Martin Stynes. Superconvergence of a finite element method for the multi-term time-fractional diffusion problem. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01115-w;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01115-w.pdf.

Liu:2020:NNM

[2522] Wei Liu, Jintao Cui, and Zhifeng Wang. A new numerical method for an asymptotic coupled model of fractured media aquifer system. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01112-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01112-z.pdf.

Du:2020:MMM

[2523] Zhijie Du and Huoyuan Duan. A mixed method for Maxwell eigenproblem. Journal of Scientific Computing, 82(1):??, January 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01111-
0;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01111-0.pdf.

Olatunji:2020:SRG

[2524] P. O. Olatunji and M. N. O. Ikhile. Strongly regular general linear methods. Journal of Scientific Computing, 82(1):??, January 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01107-
w;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01107-w.pdf.

Faghih:2020:EFC

[2525] A. Faghih and P. Mokhtary. An efficient formulation of Chebyshev Tau method for constant coefficients systems of multi-order FDEs. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01104-z;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01104-z.pdf.

Cui:2020:UCW

[2526] Ming Cui and Shangyou Zhang. On the uniform convergence of the weak Galerkin finite element method for a singularly-perturbed biharmonic equation. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-019-01120-
z;   https://link.springer.com/content/pdf/10.1007/s10915-019-
01120-z.pdf.

Li:2020:ATS

[2527] Binjie Li, Tao Wang, and Xiaoping Xie. Analysis of a time-stepping discontinuous Galerkin method for fractional diffusion-wave equations with nonsmooth data. Journal of Scientific Computing, 82(1):??, January 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-019-
01118-7;   https://link.springer.com/content/pdf/10.1007/s10915-
019-01118-7.pdf.

Hicken:2020:ESH

[2528] Jason E. Hicken. Entropy-stable, high-order summation-by-parts discretizations without interface penalties. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01154-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01154-8.pdf.

Wihler:2020:SCS

[2529] Thomas P. Wihler and Marcel Wirz. Stability and convergence of spectral mixed discontinuous Galerkin methods for 3D linear elasticity on anisotropic geometric meshes. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01153-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01153-9.pdf.

Castro:2020:WBH

[2530] Manuel J. Castro and Carlos Parés. Well-balanced high-order finite volume methods for systems of balance laws. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01149-5;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01149-5.pdf.

Gu:2020:ARB

[2531] Jiaxi Gu and Jae-Hun Jung. Adaptive radial basis function methods for initial value problems. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01140-0;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01140-0.pdf.

Sarumi:2020:HAG

[2532] Ibrahim O. Sarumi, Khaled M. Furati, and Abdul Q. M. Khaliq. Highly accurate global Padé approximations of generalized Mittag-Leffler function and its inverse. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01150-y;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01150-y.pdf.

Nurbekyan:2020:NCT

[2533] Levon Nurbekyan, Alexander Iannantuono, and Adam M. Oberman. No-collision transportation maps. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01143-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01143-x.pdf.

Nicholls:2020:SPI

[2534] David P. Nicholls, Carlos Pérez-Arancibia, and Catalin Turc. Sweeping preconditioners for the iterative solution of quasiperiodic Helmholtz transmission problems in layered media. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01133-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01133-z.pdf.

Tian:2020:ALD

[2535] Lulu Tian, Hui Guo, Rui Jia, and Yang Yang. An h-adaptive local discontinuous Galerkin method for simulating wormhole propagation with Darcy–Forcheiner model. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01135-x;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01135-x.pdf.

Lou:2020:FRM

[2536] Shuai Lou, Chao Yan, Li-Bin Ma, and Zhen-Hua Jiang. The flux reconstruction method with Lax–Wendroff type temporal discretization for hyperbolic conservation laws. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01146-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01146-8.pdf.

Penner:2020:SFE

[2537] David A. Craig Penner and David W. Zingg. Superconvergent functional estimates from tensor-product generalized summation-by-parts discretizations in curvilinear coordinates. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01147-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01147-7.pdf.

Khan:2020:SMS

[2538] Arbaz Khan. Spectral method and spectral element method for three dimensional linear elliptic system: Analysis and application. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01145-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01145-9.pdf.

Su:2020:IDG

[2539] Wei Su, Peng Wang, Yonghao Zhang, and Lei Wu. Implicit discontinuous Galerkin method for the Boltzmann equation. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01139-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01139-7.pdf.

Cibik:2020:ASO

[2540] Aytekin Cibik, Fatma G. Eroglu, and Songül Kaya. Analysis of second order time filtered backward Euler method for MHD equations. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01142-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01142-y.pdf.

Hoang:2020:NLE

[2541] Thi-Thao-Phuong Hoang, Lili Ju, and Zhu Wang. Nonoverlapping localized exponential time differencing methods for diffusion problems. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01136-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01136-w.pdf.

Li:2020:SHO

[2542] Hao Li and Xiangxiong Zhang. Superconvergence of high order finite difference schemes based on variational formulation for elliptic equations. Journal of Scientific Computing, 82(2):??, February 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01144-
w;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01144-w.pdf.

Arun:2020:APL

[2543] K. R. Arun and S. Samantaray. Asymptotic preserving low Mach number accurate IMEX finite volume schemes for the isentropic Euler equations. Journal of Scientific Computing, 82(2):??, February 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01138-
8;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01138-8.pdf.

Zhang:2020:OTM

[2544] Xiaolong Zhang and John P. Boyd. Optimal truncations for multivariate Fourier and Chebyshev series: Mysteries of the hyperbolic cross: Part i: Bivariate case. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01131-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01131-1.pdf.

Chen:2020:SEA

[2545] Feng Chen and Jie Shen. Stability and error analysis of operator splitting methods for American options under the black–Scholes model. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01137-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01137-9.pdf.

You:2020:FCF

[2546] Guoqiao You and Shingyu Leung. Fast construction of forward flow maps using Eulerian based interpolation schemes. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01141-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01141-z.pdf.

Yang:2020:OMM

[2547] Yanfang Yang, Shubin Fu, and Eric T. Chung. Online mixed multiscale finite element method with oversampling and its applications. Journal of Scientific Computing, 82(2):??, February 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-019-01121-y;   https://
link.springer.com/content/pdf/10.1007/s10915-019-01121-y.pdf.

Yao:2020:TFO

[2548] Wenjuan Yao, Jie Shen, Zhichang Guo, Jiebao Sun, and Boying Wu. A total fractional-order variation model for image super-resolution and its SAV algorithm. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01185-
1;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01185-1.pdf.

Harris:2020:AZI

[2549] Isaac Harris. Approximation of the zero-index transmission eigenvalues with a conductive boundary and parameter estimation. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01183-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01183-3.pdf.

Rui:2020:MSC

[2550] Hongxing Rui and Yue Sun. A MAC scheme for coupled Stokes–Darcy equations on non-uniform grids. Journal of Scientific Computing, 82(3): ??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01181-5;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01181-5.pdf.

Huang:2020:AES

[2551] Rong Huang. Accurate eigenvalues of some generalized sign regular matrices via relatively robust representations. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01182-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01182-4.pdf.

Trojak:2020:EMQ

[2552] Will Trojak, Rob Watson, Ashley Scillitoe, and Paul G. Tucker. Effect of mesh quality on flux reconstruction in multi-dimensions. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01184-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01184-2.pdf.

Hesthaven:2020:TDR

[2553] Jan S. Hesthaven and Fabian Mönkeberg. Two-dimensional RBF–ENO method on unstructured grids. Journal of Scientific Computing, 82(3): ??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01176-2;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01176-2.pdf.

Huynh:2020:DGI

[2554] H. T. Huynh. Discontinuous Galerkin via interpolation: The direct flux reconstruction method. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01175-3;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01175-3.pdf.

Casati:2020:CFM

[2555] Daniele Casati and Ralf Hiptmair. Coupling FEM with a multiplesubdomain Trefftz method. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01179-z;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01179-z.pdf.

Kent:2020:PDL

[2556] James Kent. A positive definite limiter for advection problems. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01178-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01178-0.pdf.

Ong:2020:TAT

[2557] Benjamin W. Ong and Satyen Dhamankar. Towards an adaptive treecode for N-body problems. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01177-1;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01177-1.pdf.

Zhao:2020:BSD

[2558] Weifeng Zhao and Wen-An Yong. Boundary scheme for a discrete kinetic approximation of the Navier–Stokes equations. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01180-6;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01180-6.pdf.

Shadpey:2020:ESM

[2559] Siavosh Shadpey and David W. Zingg. Entropy-stable multidimensional summation-by-parts discretizations on hp-adaptive curvilinear grids for hyperbolic conservation laws. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01169-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01169-1.pdf.

Schnucke:2020:ESD

[2560] Gero Schnücke, Nico Krais, Thomas Bolemann, and Gregor J. Gassner. Entropy stable discontinuous Galerkin schemes on moving meshes for hyperbolic conservation laws. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01171-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01171-7.pdf.

Facca:2020:NSM

[2561] Enrico Facca, Sara Daneri, Franco Cardin, and Mario Putti. Numerical solution of Monge–Kantorovich equations via a dynamic formulation. Journal of Scientific Computing, 82(3):??, March 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01170-
8;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01170-8.pdf.

Li:2020:ROA

[2562] Jiao fen Li, Wen Li, Seak-Weng Vong, Qi-Lun Luo, and MingQing Xiao. A Riemannian optimization approach for solving the generalized eigenvalue problem for nonsquare matrix pencils. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01173-5;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01173-5.pdf.

Rong:2020:NAB

[2563] Y. Rong and J. A. Fiordilino. Numerical analysis of a BDF2 modular grad–div stabilization method for the Navier–Stokes equations. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01165-5;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01165-5.pdf.

Acebron:2020:PLS

[2564] Juan A. Acebrón. A probabilistic linear solver based on a multilevel Monte Carlo method. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01168-2;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01168-2.pdf.

Wang:2020:PEN

[2565] Tongke Wang, Meng Qin, and Zhiyue Zhang. The Puiseux expansion and numerical integration to nonlinear weakly singular Volterra integral equations of the second kind. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01167-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01167-3.pdf.

Ha:2020:CIT

[2566] Youngsoo Ha, Chang Ho Kim, Hyoseon Yang, and Jungho Yoon. Construction of an improved third-order WENO scheme with a new smoothness indicator. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01164-
6;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01164-6.pdf.

Ruggiu:2020:MSH

[2567] Andrea A. Ruggiu and Jan Nordström. Multigrid schemes for high order discretizations of hyperbolic problems. Journal of Scientific Computing, 82(3):??, March 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01166-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01166-4.pdf.

Bouchut:2020:LCS

[2568] François Bouchut, Emmanuel Franck, and Laurent Navoret. A low cost semi-implicit low-Mach relaxation scheme for the full Euler equations. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01206-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01206-z.pdf.

Feng:2020:FDM

[2569] Xiaobing Feng, Yukun Li, and Yi Zhang. A fully discrete mixed finite element method for the stochastic Cahn–Hilliard equation with gradienttype multiplicative noise. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01202-3;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01202-3.pdf.

Cai:2020:REE

[2570] Difeng Cai, Zhiqiang Cai, and Shun Zhang. Robust equilibrated error estimator for diffusion problems: Mixed finite elements in two dimensions. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01199-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01199-9.pdf.

Feng:2020:CFS

[2571] Yiwei Feng, Tiegang Liu, and Kun Wang. A characteristic-featured shock wave indicator for conservation laws based on training an artificial neuron. Journal of Scientific Computing, 83(1):??, April 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01200-
5;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01200-5.pdf.

Jiang:2020:LIS

[2572] Chaolong Jiang, Yuezheng Gong, Wenjun Cai, and Yushun Wang. A linearly implicit structure–preserving scheme for the Camassa–Holm equation based on multiple scalar auxiliary variables approach. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01201-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01201-4.pdf.

Zheng:2020:OEE

[2573] Chunxiong Zheng and Jiashun Hu. Optimal error estimate of the extended–WKB approximation to the high frequency wavetype equation in the semi-classical regime. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01208-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01208-x.pdf.

Gervasio:2020:CCB

[2574] Paola Gervasio, Luca Dedè, Ondine Chanon, and Alfio Quarteroni. A computational comparison between isogeometric analysis and spectral element methods: Accuracy and spectral properties. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01204-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01204-1.pdf.

Li:2020:OEC

[2575] Xiaole Li, Yulong Xing, and Ching-Shan Chou. Optimal energy conserving and energy dissipative local discontinuous Galerkin methods for the Benjamin–Bona–Mahony equation. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01172-6;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01172-6.pdf.

Fengnan:2020:LFD

[2576] Liu Fengnan, Yasuhide Fukumoto, and Xiaopeng Zhao. A linearized finite difference scheme for the Richards equation under variable-flux boundary conditions. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01196-y;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01196-y.pdf.

Kim:2020:CNR

[2577] Geonwoo Kim, Junghee Cho, and Myungjoo Kang. Cauchy noise removal by weighted nuclear norm minimization. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01203-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01203-2.pdf.

Wang:2020:GCA

[2578] Weina Wang, Chunlin Wu, and Xue-Cheng Tai. A globally convergent algorithm for a constrained non–Lipschitz image restoration model. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01190-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01190-4.pdf.

Aceto:2020:PTA

[2579] Lidia Aceto and Paolo Novati. Padé-type approximations to the resolvent of fractional powers of operators. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01198-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01198-w.pdf.

Hassanzadeh:2020:CDS

[2580] Mahdi Hassanzadeh and Mahmood Mazare. Computation of design sensitivities in steady-state incompressible laminar flows based on new semianalytical method. Journal of Scientific Computing, 83(1):??, April 2020.

CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01205-
0;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01205-0.pdf.

Su:2020:TLO

[2581] Haiyan Su, Xinlong Feng, and Jianping Zhao. On two-level Oseen penalty iteration methods for the 2D/3D stationary incompressible magnetohydronamics. Journal of Scientific Computing, 83(1):??, April 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01186-
0;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01186-0.pdf.

Zhao:2020:PTA

[2582] Yong-Liang Zhao, Pei-Yong Zhu, Xian-Ming Gu, Xi-Le Zhao, and Huan-Yan Jian. A preconditioning technique for all-at-once system from the nonlinear tempered fractional diffusion equation. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01193-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01193-1.pdf.

Liu:2020:OEE

[2583] Minghui Liu, Boying Wu, and Xiong Meng. Optimal error estimates of the discontinuous Galerkin method with upwind-biased fluxes for 2D linear variable coefficients hyperbolic equations. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01197-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01197-x.pdf.

Cicuttin:2020:HHO

[2584] Matteo Cicuttin, Alexandre Ern, and Thirupathi Gudi. Hybrid highorder methods for the elliptic obstacle problem. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01195-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01195-z.pdf.

Zhu:2020:NFD

[2585] Jun Zhu, Feng Zheng, and Jianxian Qiu. New finite difference Hermite WENO schemes for Hamilton–Jacobi equations. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01174-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01174-4.pdf.

Semplice:2020:EIA

[2586] M. Semplice and G. Visconti. Efficient implementation of adaptive order reconstructions. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01156-
6;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01156-6.pdf.

Dolejsi:2020:ENS

[2587] Vít Dolejsí and Petr Tichý. On efficient numerical solution of linear algebraic systems arising in goal-oriented error estimates. Journal of Scientific Computing, 83(1):??, April 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01188-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01188-y.pdf.

Wang:2020:CRE

[2588] Lai Wang and Meilin Yu. Comparison of ROW, ESDIRK, and BDF2 for unsteady flows with the high-order flux reconstruction formulation. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01222-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01222-z.pdf.

Shen:2020:HIC

[2589] Jinye Shen, Changpin Li, and Zhi zhong Sun. An H2N2 interpolation for Caputo derivative with order in (1, 2) and its application to time-fractional wave equations in more than one space dimension. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01219-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01219-8.pdf.

Karimi:2020:RTI

[2590] Milad Karimi, Fridoun Moradlou, and Mojtaba Hajipour. Regularization technique for an inverse space-fractional backward heat conduction problem. Journal of Scientific Computing, 83(2):??, May 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01211-
2;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01211-2.pdf.

Bousquet:2020:NST

[2591] Arthur Bousquet, Youngjoon Hong, Roger Temam, and Joseph Tribbia. Numerical simulations of the two-dimensional inviscid hydrostatic primitive equations with humidity and saturation. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01215-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01215-y.pdf.

Huang:2020:SPM

[2592] Tsung-Ming Huang, Tiexiang Li, Jia-Wei Lin, Wen-Wei Lin, and Heng Tian. Structure-preserving methods for computing complex band structures of three dimensional photonic crystals. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01220-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01220-1.pdf.

Ahn:2020:MMP

[2593] Myeong-Hwan Ahn and Duck-Joo Lee. Modified monotonicity preserving constraints for high-resolution optimized compact scheme. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01221-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01221-0.pdf.

Nie:2020:NSF

[2594] Daxin Nie, Jing Sun, and Weihua Deng. Numerical scheme for the Fokker–Planck equations describing anomalous diffusions with two internal states. Journal of Scientific Computing, 83(2):??, May 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01218-
9;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01218-9.pdf.

Zhu:2020:IDU

[2595] Wei Zhu. Image denoising using L^{p}-norm of mean curvature of image surface. Journal of Scientific Computing, 83(2):??, May 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01216-
x;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01216-x.pdf.

Zhang:2020:EMS

[2596] Chao Zhang and Igor Menshov. Eulerian model for simulating multi-fluid flows with an arbitrary number of immiscible compressible components. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01214-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01214-z.pdf.

Hassani:2020:GBP

[2597] H. Hassani, J. A. Tenreiro Machado, Z. Avazzadeh, E. Naraghirad, and M. Sh. Dahaghin. Generalized Bernoulli polynomials: Solving nonlinear 2D fractional optimal control problems. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01213-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01213-0.pdf.

Wang:2020:RCN

[2598] Saihua Wang, Feng Wang, and Xuejun Xu. A rigorous condition number estimate of an immersed finite element method. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01212-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01212-1.pdf.

Guo:2020:NDI

[2599] Qilong Guo, Dong Sun, Chen Li, Pengxin Liu, and Hanxin Zhang. A new discontinuity indicator for hybrid WENO schemes. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01217-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01217-w.pdf.

Ioriatti:2020:SSI

[2600] Matteo Ioriatti, Michael Dumbser, and Raphaël Loubère. A staggered semi-implicit discontinuous Galerkin scheme with a posteriori subcell finite volume limiter for the Euler equations of gasdynamics. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01209-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01209-w.pdf.

Zhang:2020:SOA

[2601] Tiankui Zhang and Charles W. Wolgemuth. Sixth-order accurate schemes for reinitialization and extrapolation in the level set framework. Journal of Scientific Computing, 83(2):??, May 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01210-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01210-3.pdf.

Cai:2020:CAC

[2602] Wentao Cai, Jilu Wang, and Kai Wang. Convergence analysis of Crank– Nicolson Galerkin–Galerkin FEMs for miscible displacement in porous media. Journal of Scientific Computing, 83(2):??, May 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01194-
0;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01194-0.pdf.

Du:2020:PST

[2603] Yu Du and Haijun Wu. A pure source transfer domain decomposition method for Helmholtz equations in unbounded domain. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01249-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01249-2.pdf.

Xi:2020:HAN

[2604] Yingxia Xi, Xia Ji, and Shuo Zhang. A high accuracy nonconforming finite element scheme for Helmholtz transmission eigenvalue problem. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01247-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01247-4.pdf.

Wen:2020:ESW

[2605] Xiao Wen, Wai Sun Don, Zhen Gao, and Yulong Xing. Entropy stable and well-balanced discontinuous Galerkin methods for the nonlinear shallow water equations. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01248-3;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01248-3.pdf.

Li:2020:LIH

[2606] Dongfang Li and Weiwei Sun. Linearly implicit and high-order energy-conserving schemes for nonlinear wave equations. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01245-6;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01245-6.pdf.

Qiu:2020:ISR

[2607] Di Qiu and Lok Ming Lui. Inconsistent surface registration via optimization of mapping distortions. Journal of Scientific Computing, 83(3): ??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01246-5;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01246-5.pdf.

Pereira:2020:FDA

[2608] Carlos A. Pereira and Brian C. Vermeire. Fully-discrete analysis of highorder spatial discretizations with optimal explicit Runge–Kutta methods. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01243-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01243-8.pdf.

Escalante:2020:GNH

[2609] C. Escalante and T. Morales de Luna. A general non-hydrostatic hyperbolic formulation for Boussinesq dispersive shallow flows and its numerical approximation. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01244-
7;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01244-7.pdf.

Tour:2020:SEM

[2610] Geraldine Tour, Nawdha Thakoor, Jingtang Ma, and Désiré Yannick Tangman. A spectral element method for option pricing under regimeswitching with jumps. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01252-7;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01252-7.pdf.

Ong:2020:DCM

[2611] Benjamin W. Ong and Raymond J. Spiteri. Deferred correction methods for ordinary differential equations. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01235-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01235-8.pdf.

Zhang:2020:ERK

[2612] Lu Zhang, Qifeng Zhang, and Hai-Wei Sun. Exponential Runge–Kutta method for two-dimensional nonlinear fractional complex Ginzburg– Landau equations. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01240-
x;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01240-x.pdf.

Gedicke:2020:EPE

[2613] Joscha Gedicke, Sjoerd Geevers, and Ilaria Perugia. An equilibrated a posteriori error estimator for arbitrary-order Nédélec elements for magnetostatic problems. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01224-
x;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01224-x.pdf.

Xu:2020:EAD

[2614] Zhen Xu, Xiaofeng Yang, and Hui Zhang. Error analysis of a decoupled, linear stabilization scheme for the Cahn–Hilliard model of two-phase incompressible flows. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01241-
w;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01241-w.pdf.

DiIlio:2020:NHT

[2615] Giovanni Di Ilio, Stefano Ubertini, Sauro Succi, and Giacomo Falcucci. Nanofluid heat transfer in wavy-wall channels with different geometries: a finite-volume lattice Boltzmann study. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01234-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01234-9.pdf.

Strazzullo:2020:PGM

[2616] Maria Strazzullo, Francesco Ballarin, and Gianluigi Rozza. POD– Galerkin model order reduction for parametrized time dependent linear quadratic optimal control problems in saddle point formulation. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01232-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01232-x.pdf.

Ren:2020:HFD

[2617] Yupeng Ren, Tao Xiong, and Jianxian Qiu. A hybrid finite difference WENO–ZQ fast sweeping method for static Hamilton–Jacobi equations. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01228-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01228-7.pdf.

Che:2020:MAS

[2618] Maolin Che and Yimin Wei. Multiplicative algorithms for symmetric nonnegative tensor factorizations and its applications. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01233-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01233-w.pdf.

Deng:2020:PEE

[2619] Yanling Deng, Fei Wang, and Huayi Wei. A posteriori error estimates of virtual element method for a simplified friction problem. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01242-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01242-9.pdf.

Karaa:2020:MFT

[2620] Samir Karaa and Amiya K. Pani. Mixed FEM for time-fractional diffusion problems with time-dependent coefficients. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01236-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01236-7.pdf.

Sun:2020:SPN

[2621] Shouwen Sun, Jun Li, Jia Zhao, and Qi Wang. Structure-preserving numerical approximations to a non-isothermal hydrodynamic model of binary fluid flows. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01229-
6;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01229-6.pdf.

Wen:2020:EDB

[2622] Xiao Wen, Wai Sun Don, Zhen Gao, and Jan S. Hesthaven. An edge detector based on artificial neural network with application to hybrid Compact–WENO finite difference scheme. Journal of Scientific Computing, 83(3):??, June 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01237-6;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01237-6.pdf.

Mizerova:2020:CEE

[2623] Hana Mizerová and Bangwei She. Convergence and error estimates for a finite difference scheme for the multi-dimensional compressible Navier– Stokes system. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01278-
x;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01278-x.pdf.

Gander:2020:AAO

[2624] Martin J. Gander, Pratik M. Kumbhar, and Albert E. Ruehli. Asymptotic analysis for overlap in waveform relaxation methods for RC type circuits. Journal of Scientific Computing, 84(1):??, July 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01270-
5;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01270-5.pdf.

Xu:2020:SAR

[2625] Yuan Xu, Xiong Meng, Chi-Wang Shu, and Qiang Zhang. Superconvergence analysis of the Runge–Kutta discontinuous Galerkin methods for a linear hyperbolic equation. Journal of Scientific Computing, 84(1): ??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01274-1;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01274-1.pdf.

Han:2020:EHM

[2626] Yihui Han, Huangxin Chen, Xiao-Ping Wang, and Xiaoping Xie. EXtended HDG methods for second order elliptic interface problems. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01272-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01272-3.pdf.

Liu:2020:NAF

[2627] Xing Liu and Weihua Deng. Numerical approximation for fractional diffusion equation forced by a tempered fractional Gaussian noise. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01271-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01271-4.pdf.

Lee:2020:MMT

[2628] Jeonghun J. Lee. A mixed method for time-transient acoustic wave propagation in metamaterials. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01275-0;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01275-0.pdf.

Maier:2020:SDI

[2629] L.-B. Maier. Sparse data interpolation and smoothing on embedded submanifolds. Journal of Scientific Computing, 84(1):??, July 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01268-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01268-z.pdf.

Ben-Ahmed:2020:SRB

[2630] El Hassan Ben-Ahmed, Mohamed Sadik, and Mohamed Wakrim. A stable radial basis function partition of unity method with d - rectangular patches for modelling water flow in porous media. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01273-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01273-2.pdf.

Ranocha:2020:RRK

[2631] Hendrik Ranocha and David I. Ketcheson. Relaxation Runge–Kutta methods for Hamiltonian problems. Journal of Scientific Computing, 84 (1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01277-y;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01277-y.pdf.

Zhu:2020:SPD

[2632] Ya-Nan Zhu and Xiaoqun Zhang. Stochastic primal dual fixed point method for composite optimization. Journal of Scientific Computing, 84 (1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01265-2;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01265-2.pdf.

Liu:2020:ANS

[2633] Changli Liu, Jungong Xue, and Ren-Cang Li. Accurate numerical solution for shifted M-matrix algebraic Riccati equations. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01263-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01263-4.pdf.

Liu:2020:TLS

[2634] Yingzhi Liu and Yinnian He. Two-level Schwarz methods for a discontinuous Galerkin approximation of elliptic problems with jump coefficients. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01257-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01257-2.pdf.

Wang:2020:PEEa

[2635] Wansheng Wang, Lijun Yi, and Aiguo Xiao. A posteriori error estimates for fully discrete finite element method for generalized diffusion equation with delay. Journal of Scientific Computing, 84(1):??, July 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01262-

5;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01262-5.pdf.

Liang:2020:FEE

[2636] Hui Liang and Hermann Brunner. The fine error estimation of collocation methods on uniform meshes for weakly singular Volterra integral equations. Journal of Scientific Computing, 84(1):??, July 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01266-
1;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01266-1.pdf.

Deng:2020:MAC

[2637] Chongyang Deng, Xiali Fan, and Ming-Jun Lai. A minimization approach for constructing generalized barycentric coordinates and its computation. Journal of Scientific Computing, 84(1):??, July 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01267-
0;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01267-0.pdf.

Tian:2020:COO

[2638] Zhong qi Kyle Tian, Jennifer Crodelle, and Douglas Zhou. A combined offline–online algorithm for Hodgkin–Huxley neural networks. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01261-6;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01261-6.pdf.

Rupp:2020:CGE

[2639] Andreas Rupp and Sanghyun Lee. Continuous Galerkin and enriched Galerkin methods with arbitrary order discontinuous trial functions for the elliptic and parabolic problems with jump conditions. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01255-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01255-4.pdf.

Cao:2020:FHO

[2640] Rongjun Cao, Minghua Chen, Michael K. Ng, and Yu-Jiang Wu. Fast and high-order accuracy numerical methods for time-dependent nonlocal problems in R^{2}. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01260-
7;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01260-7.pdf.

Gie:2020:EFV

[2641] Gung-Min Gie, Chang-Yeol Jung, and Hoyeon Lee. Enriched finite volume approximations of the plane–parallel flow at a small viscosity. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01259-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01259-0.pdf.

Sun:2020:EEB

[2642] Jing Sun, Daxin Nie, and Weihua Deng. Error estimates for backward fractional Feynman–Kac equation with non-smooth initial data. Journal of Scientific Computing, 84(1):??, July 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01256-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01256-3.pdf.

Zhang:2020:RLO

[2643] Shun Zhang. Robust and local optimal a priori error estimates for interface problems with low regularity: Mixed finite element approximations. Journal of Scientific Computing, 84(2):??, August 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01284-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01284-z.pdf.

Chandrashekar:2020:CPD

[2644] Praveen Chandrashekar and Rakesh Kumar. Constraint preserving discontinuous Galerkin method for ideal compressible MHD on 2-D Cartesian grids. Journal of Scientific Computing, 84(2):??, August 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01289-
8;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01289-8.pdf.

Ren:2020:SEE

[2645] Jincheng Ren, Hong lin Liao, and Zhimin Zhang. Superconvergence error estimate of a finite element method on nonuniform time meshes for reaction–subdiffusion equations. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01290-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01290-1.pdf.

Wang:2020:PEEb

[2646] Gang Wang, Ying Wang, and Yinnian He. A posteriori error estimates for the virtual element method for the Stokes problem. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01281-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01281-2.pdf.

Yong:2020:NAL

[2647] Wen-An Yong and Weifeng Zhao. Numerical analysis of the lattice Boltzmann method for the Boussinesq equations. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01291-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01291-0.pdf.

Adjerid:2020:EEI

[2648] Slimane Adjerid, Tao Lin, and Qiao Zhuang. Error estimates for an immersed finite element method for second order hyperbolic equations in inhomogeneous media. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01283-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01283-0.pdf.

Kim:2020:CML

[2649] Dongho Kim, Eun-Jae Park, and Boyoon Seo. Convergence of multi-level algorithms for a class of nonlinear problems. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01287-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01287-w.pdf.

Xu:2020:MPP

[2650] Xiang Xu and Yanxiang Zhao. Maximum principle preserving schemes for binary systems with long–range interactions. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01286-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01286-x.pdf.

Huang:2020:HFS

[2651] Guangnan Huang and Songting Luo. Hybrid fast sweeping methods for anisotropic eikonal equation in two-dimensional tilted transversely isotropic media. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01280-
3;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01280-3.pdf.

Bourriaud:2020:PNN

[2652] Alexandre Bourriaud, Raphaël Loubère, and Rodolphe Turpault. A priori neural networks versus a posteriori MOOD loop: a high accurate 1D FV scheme testing bed. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01282-1;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01282-1.pdf.

Minakowski:2020:FEE

[2653] Piotr Minakowski and Thomas Richter. Finite element error estimates on geometrically perturbed domains. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01285-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01285-y.pdf.

Lambert:2020:MNB

[2654] Wanderson Lambert, Amaury Alvarez, Ismael Ledoino, Duilio Tadeu, Dan Marchesin, and Johannes Bruining. Mathematics and numerics for balance partial differential–algebraic equations (PDAEs). Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01279-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01279-w.pdf.

Dabaghi:2020:AIS

[2655] Jad Dabaghi, Vincent Martin, and Martin Vohralík. Adaptive inexact semismooth Newton methods for the contact problem between two membranes. Journal of Scientific Computing, 84(2):??, August 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01264-
3;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01264-3.pdf.

Chen:2020:ESN

[2656] Wenbin Chen, Cheng Wang, Shufen Wang, Xiaoming Wang, and Steven M. Wise. Energy stable numerical schemes for ternary Cahn– Hilliard system. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01276-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01276-z.pdf.

Liu:2020:AFO

[2657] Qinjie Liu and Shun Zhang. Adaptive flux–only least–squares finite element methods for linear transport equations. Journal of Scientific Computing, 84(2):??, August 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01269-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01269-y.pdf.

Ying:2020:MDA

[2658] Lexing Ying. Mirror descent algorithms for minimizing interacting free energy. Journal of Scientific Computing, 84(3):??, September 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01303-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01303-z.pdf.

Yang:2020:IMS

[2659] Jun Yang. The iterative methods for solving pseudomontone equilibrium problems. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01298-
7;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01298-7.pdf.

Dai:2020:FPI

[2660] Ping-Fan Dai. A fixed point iterative method for tensor complementarity problems. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01299-

6;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01299-6.pdf.

Zhang:2020:SLM

[2661] Jianying Zhang. Stability of linear multistep time iterations with the WENO5 discretization at discontinuities. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01300-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01300-2.pdf.

Dai:2020:TGB

[2662] Xiaoying Dai, Xiong Kuang, Jack Xin, and Aihui Zhou. Twogrid based adaptive proper orthogonal decomposition method for time dependent partial differential equations. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01288-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01288-9.pdf.

Huska:2020:SAV

[2663] M. Huska, D. Lazzaro, S. Morigi, A. Samorè, and G. Scrivanti. Spatially-adaptive variational reconstructions for linear inverse electrical impedance tomography. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01295-w;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01295-w.pdf.

Kundu:2020:GST

[2664] Sudeep Kundu and Amiya Kumar Pani. Global stabilization of two dimensional viscous Burgers’ equation by nonlinear Neumann boundary feedback control and its finite element analysis. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01294-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01294-x.pdf.

Zhang:2020:SAI

[2665] Hong Zhang, Zhengyu Liu, Emil Constantinescu, and Robert Jacob. Stability analysis of interface conditions for ocean–atmosphere coupling. Journal of Scientific Computing, 84(3):??, September 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01293-
y;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01293-y.pdf.

Martin:2020:EEH

[2666] Lindsay Martin and Yen-Hsi Richard Tsai. Equivalent extensions of Hamilton–Jacobi-Bellman equations on hypersurfaces. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01292-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01292-z.pdf.

Reshniak:2020:MGP

[2667] Viktor Reshniak and Yuri Melnikov. Method of Green’s potentials for elliptic PDEs in domains with random apertures. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01296-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01296-9.pdf.

Liu:2020:PPI

[2668] Ching-Sung Liu. A positivity preserving iterative method for finding the ground states of saturable nonlinear Schrödinger equations. Journal of Scientific Computing, 84(3):??, September 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01297-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01297-8.pdf.

Sun:2020:OMS

[2669] Zhangpeng Sun, Wenqi Yao, and Tiao Lu. Optimization modeling and simulating of the stationary Wigner inflow boundary value problem. Journal of Scientific Computing, 85(1):??, October 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01338-
2;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01338-2.pdf.

Langer:2020:BVP

[2670] Stefan Langer and R. C. Swanson. On boundary-value problems for RANS equations and two-equation turbulence models. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01323-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01323-9.pdf.

Liu:2020:PFV

[2671] Huan Liu, Aijie Cheng, and Hong Wang. A parareal finite volume method for variable–order time-fractional diffusion equations. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01321-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01321-x.pdf.

Chen:2020:NMS

[2672] Xue Chen, Zhong bao Wang, and Zhang you Chen. A new method for solving variational inequalities and fixed points problems of demi–contractive mappings in Hilbert spaces. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01327-5;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01327-5.pdf.

Jin:2020:CSM

[2673] Yao Jin, Fei Liao, and Jinsheng Cai. Compact schemes for multiscale flows with cell–centered finite difference method. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01314-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01314-w.pdf.

Guo:2020:NPB

[2674] Jun Guo and Minfu Feng. A new projection–based stabilized virtual element method for the Stokes problem. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01301-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01301-1.pdf.

Guo:2020:SIH

[2675] Shimin Guo, Liquan Mei, Can Li, Zhengqiang Zhang, and Ying Li. Semi-implicit Hermite–Galerkin spectral method for distributed–order fractional-in-space nonlinear reaction–diffusion equations in multidimensional unbounded domains. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01320-y;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01320-y.pdf.

Oyarzua:2020:DCD

[2676] Ricardo Oyarzúa and Miguel Serón. A divergence-conforming DG-Mixed finite element method for the stationary Boussinesq problem. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01317-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01317-7.pdf.

Huang:2020:AMI

[2677] Juntao Huang, Yuan Liu, Wei Guo, Zhanjing Tao, and Yingda Cheng. An adaptive multiresolution interior penalty discontinuous Galerkin method for wave equations in second order form. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01322-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01322-w.pdf.

Cheng:2020:QCD

[2678] Jian Cheng, Fan Zhang, and Tiegang Liu. A quasi-conservative discontinuous Galerkin method for solving five equation model of compressible two-medium flows. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01319-5;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01319-5.pdf.

Asmouh:2020:CSL

[2679] Ilham Asmouh, Mofdi El-Amrani, Mohammed Seaid, and Naji Yebari. A conservative semi-Lagrangian finite volume method for convection–diffusion problems on unstructured grids. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01316-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01316-8.pdf.

Caiazzo:2020:RQL

[2680] A. Caiazzo, R. Maier, and D. Peterseim. Reconstruction of quasi-local numerical effective models from low-resolution measurements. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01304-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01304-y.pdf.

Lan:2020:NAL

[2681] Rihui Lan and Pengtao Sun. A novel arbitrary Lagrangian–Eulerian finite element method for a mixed parabolic problem in a moving domain. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01315-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01315-9.pdf.

Du:2020:HMS

[2682] Shukai Du. HDG methods for Stokes equation based on strong symmetric stress formulations. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01309-7;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01309-7.pdf.

Zhang:2020:ESC

[2683] Haixiang Zhang, Xuehua Yang, and Da Xu. An efficient spline collocation method for a nonlinear fourth-order reaction subdiffusion equation. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01308-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01308-8.pdf.

Linn:2020:STO

[2684] Tobias Linn, Kai Bittner, Hans Georg Brachtendorf, and Christoph Jungemann. Simulation of THz oscillations in semiconductor devices based on balance equations. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01311-z;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01311-z.pdf.

Deriaz:2020:NLC

[2685] Erwan Deriaz and Pierre Haldenwang. Non-linear CFL conditions issued from the von Neumann stability analysis for the transport equation. Journal of Scientific Computing, 85(1):??, October 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01302-
0;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01302-0.pdf.

Amine:2020:LIM

[2686] M. Benzakour Amine. Linearized implicit methods based on a single– layer neural network: Application to Keller–Segel models. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01310-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01310-0.pdf.

Song:2020:ACF

[2687] Guohui Song, Gabe Tucker, and Congzhi Xia. Admissible concentration factors for edge detection from non-uniform Fourier data. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01307-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01307-9.pdf.

Zhang:2020:OOU

[2688] Jin Zhang and Xiaowei Liu. Optimal order of uniform convergence for finite element method on bakhvalov-type meshes. Journal of Scientific Computing, 85(1):??, October 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01312-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01312-y.pdf.

Farnham:2020:PCP

[2689] Stephen D. Farnham, Lixin Shen, and Bruce W. Suter. Principal component projection with low-degree polynomials. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01336-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01336-4.pdf.

Archibald:2020:ENA

[2690] Richard Archibald, Feng Bao, Jiongmin Yong, and Tao Zhou. An efficient numerical algorithm for solving data driven feedback control problems. Journal of Scientific Computing, 85(2):??, November 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01358-

y;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01358-y.pdf.

Feliu-Faba:2020:HIF

[2691] Jordi Feliu-Fabà and Lexing Ying. Hierarchical interpolative factorization preconditioner for parabolic equations. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01343-5;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01343-5.pdf.

Burger:2020:MSW

[2692] Raimund Bürger, Enrique D. Fernández-Nieto, and Víctor Osores. A multilayer shallow water approach for polydisperse 40Sedimentation with sediment compressibility and mixture viscosity. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01334-6;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01334-6.pdf.

Pereira:2020:SPH

[2693] Carlos A. Pereira and Brian C. Vermeire. Spectral properties of highorder element types for implicit large eddy simulation. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01329-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01329-3.pdf.

Qiu:2020:EAE

[2694] Hailong Qiu. Error analysis of Euler semi-implicit scheme for the nonstationary magneto-hydrodynamics problem with temperature dependent parameters. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01357-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01357-z.pdf.

Zhao:2020:NOR

[2695] Xueying Zhao, Minru Bai, and Michael K. Ng. Nonconvex optimization for robust tensor completion from grossly sparse observations. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01356-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01356-0.pdf.

Chen:2020:USE

[2696] Wenbin Chen, Daozhi Han, Xiaoming Wang, and Yichao Zhang. Uniquely solvable and energy stable decoupled numerical schemes for the Cahn–Hilliard–Navier–Stokes–Darcy–Boussinesq system. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01341-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01341-7.pdf.

Caucao:2020:FMF

[2697] Sergio Caucao, Gabriel N. Gatica, Ricardo Oyarzúa, and Nestor Sánchez. A fully–mixed formulation for the steady double–diffusive convection system based upon Brinkman–Forchheimer equations. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01305-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01305-x.pdf.

Abgrall:2020:ASS

[2698] R. Abgrall, J. Nordström, P. Öffner, and S. Tokareva. Analysis of the SBP–SAT stabilization for finite element methods. Part I: Linear problems. Journal of Scientific Computing, 85(2):??, November 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01349-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01349-z.pdf.

Du:2020:TFA

[2699] Qiang Du, Jiang Yang, and Zhi Zhou. Time-fractional Allen– Cahn equations: Analysis and numerical methods. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01351-5;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01351-5.pdf.

Li:2020:MAL

[2700] Changpin Li, Zhiqiang Li, and Zhen Wang. Mathematical analysis and the local discontinuous Galerkin method for Caputo– Hadamard fractional partial differential equation. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01353-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01353-3.pdf.

Breda:2020:CNG

[2701] Dimitri Breda, Toshikazu Kuniya, Jordi Ripoll, and Rossana Vermiglio. Collocation of next-generation operators for computing the basic reproduction number of structured populations. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01339-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01339-1.pdf.

Zhou:2020:NAL

[2702] Boya Zhou, Xiaoli Chen, and Dongfang Li. Nonuniform Alikhanov linearized Galerkin finite element methods for nonlinear time-fractional parabolic equations. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01350-6;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01350-6.pdf.

Khouzani:2020:NAU

[2703] Hamed Jalali Khouzani and Ramin Kamali Moghadam. A novel approach of unsteady adjoint lattice Boltzmann method based on circular function scheme. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01318-
6;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01318-6.pdf.

Pan:2020:CTV

[2704] Huan Pan, You-Wei Wen, and Tieyong Zeng. Constrained total variation based three-dimension single particle reconstruction in cryogenic electron microscopy. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01344-
4;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01344-4.pdf.

Guy:2020:DAS

[2705] Hayley Guy, Alen Alexanderian, and Meilin Yu. A distributed active subspace method for scalable surrogate modeling of function valued outputs. Journal of Scientific Computing, 85(2):??, November 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01346-
2;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01346-2.pdf.

Du:2020:PIF

[2706] Guangzhi Du and Liyun Zuo. A parallel iterative finite element method for the linear elliptic equations. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01348-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01348-0.pdf.

Zhu:2020:UCW

[2707] Peng Zhu and Shenglan Xie. A uniformly convergent weak Galerkin finite element method on Shishkin mesh for 1 d convection–diffusion problem. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01345-
3;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01345-3.pdf.

Bot:2020:VSC

[2708] Radu Ioan Bot and Axel Böhm. Variable smoothing for convex optimization problems using stochastic gradients. Journal of Scientific Computing, 85(2):??, November 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01332-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01332-8.pdf.

Jones:2020:STH

[2709] Giselle Sosa Jones, Jeonghun J. Lee, and Sander Rhebergen. A space– time hybridizable discontinuous Galerkin method for linear free-surface waves. Journal of Scientific Computing, 85(3):??, December 2020. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01340-
8;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01340-8.pdf.

Liu:2020:WBP

[2710] Xin Liu. A well-balanced and positivity-preserving numerical model for shallow water flows in channels with wet–dry fronts. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01362-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01362-2.pdf.

Li:2020:NAT

[2711] Binjie Li, Tao Wang, and Xiaoping Xie. Numerical analysis of two Galerkin discretizations with graded temporal grids for fractional evolution equations. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01365-z;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01365-z.pdf.

Qiu:2020:WPF

[2712] Hailong Qiu. Well-posedness and finite element approximation for the stationary magneto-hydrodynamics problem with temperaturedependent parameters. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01361-3;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01361-3.pdf.

Shen:2020:ASP

[2713] Chungen Shen, Wenjuan Xue, Lei-Hong Zhang, and Baiyun Wang. An active-set proximal-Newton algorithm for ℓ_{1} regularized optimization problems with box constraints. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01364-0;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01364-0.pdf.

Ling:2020:NVE

[2714] Min Ling, Fei Wang, and Weimin Han. The nonconforming virtual element method for a stationary Stokes hemivariational inequality with slip boundary condition. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01333-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01333-7.pdf.

Ma:2020:FDM

[2715] Jingtang Ma and Jianjun Ma. Finite difference methods for the Hamilton–Jacobi-Bellman equations arising in regime switching utility maximization. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01352-4;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01352-4.pdf.

Djeridi:2020:SEM

[2716] Bochra Djeridi, Radouen Ghanem, and Hocine Sissaoui. Spectral element methods a priori and a posteriori error estimates for penalized unilateral obstacle problem. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01355-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01355-1.pdf.

Miao:2020:CDM

[2717] Cun-Qiang Miao. On Chebyshev–Davidson method for symmetric generalized eigenvalue problems. Journal of Scientific Computing, 85(3):??, December 2020. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-020-01360-4;   https://link.springer.com/content/pdf/
10.1007/s10915-020-01360-4.pdf.

Xie:2021:CAW

[2718] Yingying Xie and Liuqiang Zhong. Convergence of adaptive weak Galerkin finite element methods for second order elliptic problems. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01387-7;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01387-7.pdf.

Jiang:2021:HOF

[2719] Yan Jiang. High order finite difference multi-resolution WENO method for nonlinear degenerate parabolic equations. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01382-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01382-y.pdf.

Carrillo:2021:LWA

[2720] H. Carrillo, C. Parés, and D. Zorío. Lax–Wendroff approximate Taylor methods with fast and optimized weighted essentially non-oscillatory reconstructions. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01380-
0;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01380-0.pdf.

Abreu:2021:CLE

[2721] E. Abreu, V. Matos, J. Pérez, and P. Rodríguez-Bermúdez. A class of Lagrangian–Eulerian shock–capturing schemes for first-order hyperbolic problems with forcing terms. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01392-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01392-w.pdf.

Kheirfam:2021:SOC

[2722] B. Kheirfam, A. Nasrollahi, and M. Mohammadi. A secondorder corrector infeasible interior–point method for semidefinite optimization based on a wide neighborhood. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01384-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01384-w.pdf.

Taiwo:2021:ITA

[2723] A. Taiwo, L. O. Jolaoso, and O. T. Mewomo. Inertial-type algorithm for solving split common fixed point problems in Banach spaces. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01385-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01385-9.pdf.

Chu:2021:LSM

[2724] Kwunlun Chu and Shingyu Leung. A level set method for the Dirichlet k -partition problem. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01368-w;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01368-w.pdf.

Bernier:2021:ESM

[2725] Joackim Bernier, Nicolas Crouseilles, and Yingzhe Li. Exact splitting methods for kinetic and Schrödinger equations. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01369-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01369-9.pdf.

Tajaddini:2021:TNV

[2726] Azita Tajaddini, Gang Wu, Farid Saberi-Movahed, and Najmeh Azizizadeh. Two new variants of the simpler block GMRES method with vector deflation and eigenvalue deflation for multiple linear systems. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01376-w;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01376-w.pdf.

Zhai:2021:SEE

[2727] Shuying Zhai, Zhifeng Weng, Xinlong Feng, and Yinnian He. Stability and error estimate of the operator splitting method for the phase field crystal equation. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01386-8;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01386-8.pdf.

Feng:2021:FBH

[2728] Hongsong Feng, Guangqing Long, and Shan Zhao. FFT-Based high order central difference schemes for Poisson’s equation with staggered boundaries. Journal of Scientific Computing, 86(1):??, January 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01379-
7;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01379-7.pdf.

Kao:2021:LCR

[2729] Chiu-Yen Kao, Seyyed Abbas Mohammadi, and Braxton Osting. Linear convergence of a rearrangement method for the one-dimensional Poisson equation. Journal of Scientific Computing, 86(1):??, January 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01389-
5;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01389-5.pdf.

Chen:2021:NPC

[2730] Jinqiang Chen, Peixiang Yu, Hua Ouyang, and Zhen F. Tian. A novel parallel computing strategy for compact difference schemes with consistent accuracy and dispersion. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01383-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01383-x.pdf.

Zhang:2021:LDG

[2731] Chao Zhang, Yan Xu, and Yinhua Xia. Local discontinuous Galerkin methods to a dispersive system of KdV-Type equations. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01370-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01370-2.pdf.

Barsukow:2021:AFS

[2732] Wasilij Barsukow. The active flux scheme for nonlinear problems. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01381-z;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01381-z.pdf.

Antonietti:2021:PTN

[2733] Paola F. Antonietti, Jacopo De Ponti, Luca Formaggia, and Anna Scotti. Preconditioning techniques for the numerical solution of flow in fractured porous media. Journal of Scientific Computing, 86(1):??, January 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01372-
0;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01372-0.pdf.

Christlieb:2021:PAS

[2734] Andrew J. Christlieb, Pierson T. Guthrey, William A. Sands, and Mathialakan Thavappiragasm. Parallel algorithms for successive convolution. Journal of Scientific Computing, 86(1):??, January 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01359-
x;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01359-x.pdf.

Chen:2021:OSM

[2735] Xin Chen, Martin J. Gander, and Yingxiang Xu. Optimized Schwarz methods with elliptical domain decompositions. Journal of Scientific Computing, 86(2):??, February 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01394-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01394-8.pdf.

Ying:2021:NGC

[2736] Lexing Ying. Natural gradient for combined loss using wavelets. Journal of Scientific Computing, 86(2):??, February 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01367-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01367-x.pdf.

Garres-Diaz:2021:WNH

[2737] J. Garres-Díaz, E. D. Fernández-Nieto, A. Mangeney, and T. Morales de Luna. A weakly non-hydrostatic shallow model for dry granular flows. Journal of Scientific Computing, 86(2):??, February 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01377-
9;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01377-9.pdf.

Yuan:2021:MFE

[2738] Yirang Yuan, Changfeng Li, and Qing Yang. Mixed finite element– second order upwind fractional step difference scheme of Darcy– Forchheimer miscible displacement and its numerical analysis. Journal of Scientific Computing, 86(2):??, February 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01393-9;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01393-9.pdf.

Sarumi:2021:GET

[2739] Ibrahim O. Sarumi, Khaled M. Furati, Abdul Q. M. Khaliq, and Kassem Mustapha. Generalized exponential time differencing schemes for stiff fractional systems with nonsmooth source term. Journal of Scientific Computing, 86(2):??, February 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01374-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01374-y.pdf.

Le:2021:RSF

[2740] Kim-Ngan Le and Martin Stynes. An α-robust semidiscrete finite element method for a Fokker–Planck initial-boundary value problem with variable-order fractional time derivative. Journal of Scientific Computing, 86(2):??, February 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01375-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01375-x.pdf.

Liu:2021:APD

[2741] Yanli Liu, Yunbei Xu, and Wotao Yin. Acceleration of primal–dual methods by preconditioning and simple subproblem procedures. Journal of Scientific Computing, 86(2):??, February 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01371-1;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01371-1.pdf.

Wang:2021:TDN

[2742] Bo Wang, Zhiguo Yang, Li-Lian Wang, and Shidong Jiang. On timedomain NRBC for Maxwell’s equations and its application in accurate simulation of electromagnetic invisibility cloaks. Journal of Scientific Computing, 86(2):??, February 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01354-2;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01354-2.pdf.

Liu:2021:DDF

[2743] Hongyan Liu, Changtao Sheng, Li-Lian Wang, and Huifang Yuan. On diagonal dominance of FEM stiffness matrix of fractional Laplacian and maximum principle preserving schemes for the fractional Allen–Cahn equation. Journal of Scientific Computing, 86(2):??, February 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01363-
1;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01363-1.pdf.

Yarmohammadi:2021:PFI

[2744] Mohammadreza Yarmohammadi and Shahnam Javadi. Piecewise fractional interpolation with application to fractional differential equation. Journal of Scientific Computing, 86(2):??, February 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01373-
z;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01373-z.pdf.

Lundquist:2021:SDA

[2745] Tomas Lundquist, Jan Nordström, and Arnaud Malan. Stable dynamical adaptive mesh refinement. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01414-1;   https://link.springer.com/content/pdf/
10.1007/s10915-021-01414-1.pdf.

Allen:2021:SAM

[2746] Jeffery M. Allen, Justin Chang, FrançoisL. E. Usseglio-Viretta, Peter Graf, and Kandler Smith. A segregated approach for modeling the electrochemistry in the 3-D microstructure of Li-ion batteries and its acceleration using block preconditioners. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01410-5;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01410-5.pdf.

Choi:2021:VDE

[2747] Gary P. T. Choi and Chris H. Rycroft. Volumetric densityequalizing reference map with applications. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01411-4;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01411-4.pdf.

Kao:2021:ERP

[2748] Chiu-Yen Kao and Seyyed Abbas Mohammadi. Extremal rearrangement problems involving Poisson’s equation with Robin boundary conditions. Journal of Scientific Computing, 86(3):??, March 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01413-
2;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01413-2.pdf.

Fu:2021:LFG

[2749] Guosheng Fu, Christoph Lehrenfeld, Alexander Linke, and Timo Streckenbach. Locking-free and gradient-robust H⃗(div)-conforming HDG methods for linear elasticity. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01396-6;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01396-6.pdf.

Xie:2021:CPA

[2750] Yue Xie and Stephen J. Wright. Complexity of proximal augmented Lagrangian for nonconvex optimization with nonlinear equality constraints. Journal of Scientific Computing, 86(3):??, March 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01409-
y;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01409-y.pdf.

Gaburro:2021:PSF

[2751] Elena Gaburro and Michael Dumbser. A posteriori subcell finite volume limiter for general P_{N}P_{M} schemes: Applications from gasdynamics to relativistic magnetohydrodynamics. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01405-8;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01405-8.pdf.

Hicken:2021:ESD

[2752] Jason E. Hicken. On entropy-stable discretizations and the entropy adjoint. Journal of Scientific Computing, 86(3):??, March 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01395-
7;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01395-7.pdf.

Huang:2021:DAN

[2753] Yunqing Huang, Meng Chen, and Jichun Li. Developing and analyzing new unconditionally stable finite element schemes for Maxwell’s equations in complex media. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-020-
01406-7;   https://link.springer.com/content/pdf/10.1007/s10915-
020-01406-7.pdf.

Cai:2021:WCR

[2754] Meng Cai, Siqing Gan, and Xiaojie Wang. Weak convergence rates for an explicit full-discretization of stochastic Allen–Cahn equation with additive noise. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01378-
8;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01378-8.pdf.

Balashov:2021:NSD

[2755] Vladislav Balashov and Alexander Zlotnik. On a new spatial discretization for a regularized 3D compressible isothermal Navier–Stokes– Cahn–Hilliard system of equations with boundary conditions. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01388-6;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01388-6.pdf.

Cheng:2021:ABT

[2756] Xinyu Cheng, Dong Li, Keith Promislow, and Brian Wetton. Asymptotic behaviour of time stepping methods for phase field models. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01391-x;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01391-x.pdf.

She:2021:BPR

[2757] Zi-Hang She, Cheng-Xue Lao, Hong Yang, and Fu-Rong Lin. Banded preconditioners for Riesz space fractional diffusion equations. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01398-4;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01398-4.pdf.

Schonsheck:2021:NSR

[2758] Stefan C. Schonsheck, Michael M. Bronstein, and Rongjie Lai. Nonisometric surface registration via conformal Laplace–Beltrami basis pursuit. Journal of Scientific Computing, 86(3):??, March 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-020-01390-
y;   https://link.springer.com/content/pdf/10.1007/s10915-020-
01390-y.pdf.

Zheng:2021:OPG

[2759] Xiangcheng Zheng, V. J. Ervin, and Hong Wang. Optimal Petrov– Galerkin spectral approximation method for the fractional diffusion, advection, reaction equation on a bounded interval. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01366-y;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01366-y.pdf.

Alvarez:2021:LRB

[2760] Diego Álvarez, Pedro González-Rodríguez, and Manuel Kindelan. A local radial basis function method for the Laplace–Beltrami operator. Journal of Scientific Computing, 86(3):??, March 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-020-01399-3;   https://
link.springer.com/content/pdf/10.1007/s10915-020-01399-3.pdf.

Liu:2021:SLD

[2761] Xiaobin Liu, Dazhi Zhang, Xiong Meng, and Boying Wu. Superconvergence of the local discontinuous Galerkin method for one dimensional nonlinear convection–diffusion equations. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01446-7;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01446-7.pdf.

Wang:2021:EIM

[2762] Dong Wang. An efficient iterative method for reconstructing surface from point clouds. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01457-
4;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01457-4.pdf.

Leng:2021:AHMa

[2763] Haitao Leng. Adaptive HDG methods for the steady-state incompressible Navier–Stokes equations. Journal of Scientific Computing, 87(1): ??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01456-5;   https://link.springer.com/content/pdf/
10.1007/s10915-021-01456-5.pdf.

Beisiegel:2021:MPQ

[2764] Nicole Beisiegel, Cristóbal E. Castro, and Jörn Behrens. Metrics for performance quantification of adaptive mesh refinement. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01423-0;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01423-0.pdf.

Hao:2021:CIE

[2765] Tingting Hao, Xiaofei Guan, Shipeng Mao, and Shaochun Chen. Computable interpolation error constants for the geometric simplex finite elements. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01449-
4;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01449-4.pdf.

Zhang:2021:AES

[2766] Chenhui Zhang, Jie Ouyang, Xiaodong Wang, Yong Chai, and Mengxia Ma. Analysis of the energy stability for stabilized semi-implicit schemes of the functionalized Cahn–Hilliard mass-conserving gradient flow equation. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01430-
1;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01430-1.pdf.

Ranocha:2021:NCS

[2767] Hendrik Ranocha and Jan Nordström. A new class of A stable summation by parts time integration schemes with strong initial conditions. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01454-7;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01454-7.pdf.

Li:2021:SFR

[2768] Yuwen Li. Superconvergent flux recovery of the Rannacher–Turek nonconforming element. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-021-
01445-8;   https://link.springer.com/content/pdf/10.1007/s10915-
021-01445-8.pdf.

Guo:2021:NRB

[2769] Weihong Guo, Yifei Lou, Jing Qin, and Ming Yan. A novel regularization based on the error function for sparse recovery. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01443-w;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01443-w.pdf.

Chen:2021:DSE

[2770] Hao Chen and Hai-Wei Sun. A dimensional splitting exponential time differencing scheme for multidimensional fractional Allen–Cahn equations. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01431-0;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01431-0.pdf.

Costa-Sole:2021:HOH

[2771] Albert Costa-Solé, Eloi Ruiz-Gironés, and Josep Sarrate. High-order hybridizable discontinuous Galerkin formulation for one-phase flow through porous media. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01436-
9;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01436-9.pdf.

Li:2021:SFD

[2772] Xue-Yang Li and Ai-Guo Xiao. Space-fractional diffusion equation with variable coefficients: Well-posedness and Fourier pseudospectral approximation. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01435-
w;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01435-w.pdf.

Xiao:2021:FEC

[2773] Wenqiang Xiao, Bo Gong, Jiguang Sun, and Zhimin Zhang. Finite element calculation of photonic band structures for frequency dependent materials. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01439-
6;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01439-6.pdf.

Wang:2021:SSC

[2774] Jilu Wang, Jungang Wang, and Lihong Yin. A single-step correction scheme of Crank–Nicolson convolution quadrature for the subdiffusion equation. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01419-
w;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01419-w.pdf.

Qin:2021:LCA

[2775] Hongyu Qin, Fengyan Wu, Jiwei Zhang, and Chunlai Mu. A linearized compact ADI scheme for semilinear parabolic problems with distributed delay. Journal of Scientific Computing, 87(1):??, April 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01441-
y;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01441-y.pdf.

Qiu:2021:RLT

[2776] Duo Qiu, Minru Bai, Michael K. Ng, and Xiongjun Zhang. Robust low transformed multi-rank tensor methods for image alignment. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01437-8;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01437-8.pdf.

Li:2021:HOE

[2777] Buyang Li and Shu Ma. A high-order exponential integrator for nonlinear parabolic equations with nonsmooth initial data. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01438-7;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01438-7.pdf.

Hoang:2021:GTD

[2778] Thi-Thao-Phuong Hoang and Hyesuk Lee. A global-in-time domain decomposition method for the coupled nonlinear Stokes and Darcy flows. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01422-1;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01422-1.pdf.

Wu:2021:SSA

[2779] Caiying Wu, Jing Wang, Jan Harold Alcantara, and Jein-Shan Chen. Smoothing strategy along with conjugate gradient algorithm for signal reconstruction. Journal of Scientific Computing, 87(1):??, April 2021.

CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01440-
z;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01440-z.pdf.

Tan:2021:SCS

[2780] Bing Tan, Xiaolong Qin, and Jen-Chih Yao. Strong convergence of self-adaptive inertial algorithms for solving split variational inclusion problems with applications. Journal of Scientific Computing, 87(1):??, April 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01428-9;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01428-9.pdf.

Zhu:2021:PEA

[2781] Bingxin Zhu and Haijun Wu. Preasymptotic error analysis of the HDG method for Helmholtz equation with large wave number. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01473-4;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01473-4.pdf.

Miyajima:2021:CEM

[2782] Shinya Miyajima. Computing enclosures for the matrix Mittag-Leffler function. Journal of Scientific Computing, 87(2):??, May 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01447-
6;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01447-6.pdf.

Fresca:2021:CDL

[2783] Stefania Fresca, Luca Dede’, and Andrea Manzoni. A comprehensive deep learning-based approach to reduced order modeling of nonlinear time-dependent parametrized PDEs. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01462-7;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01462-7.pdf.

Jo:2021:LCI

[2784] Gwanghyun Jo, Do Y. Kwak, and Young-Ju Lee. Locally conservative immersed finite element method for elliptic interface problems. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01476-1;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01476-1.pdf.

Zhang:2021:CSM

[2785] Qifeng Zhang and Lingling Liu. Convergence and stability in maximum norms of linearized fourth-order conservative compact scheme for Benjamin–Bona–Mahony–Burgers’ equation. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01474-3;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01474-3.pdf.

Dutta:2021:OPE

[2786] Jogen Dutta and Bhupen Deka. Optimal a priori error estimates for the finite element approximation of dual-phase-lag bio heat model in heterogeneous medium. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01460-
9;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01460-9.pdf.

Ying:2021:HIP

[2787] Lexing Ying. A heuristic independent particle approximation to determinantal point processes. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-021-
01472-5;   https://link.springer.com/content/pdf/10.1007/s10915-
021-01472-5.pdf.

Moore:2021:EIB

[2788] Brian E. Moore. Exponential integrators based on discrete gradients for linearly damped/driven Poisson systems. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01468-1;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01468-1.pdf.

LeFloch:2021:KFN

[2789] Philippe G. LeFloch and Hendrik Ranocha. Kinetic functions for nonclassical shocks, entropy stability, and discrete summation by parts. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01463-6;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01463-6.pdf.

Zhou:2021:AFV

[2790] Guanyu Zhou. An analysis on the finite volume schemes and the discrete Lyapunov inequalities for the chemotaxis system. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01466-3;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01466-3.pdf.

Gao:2021:PSP

[2791] Huadong Gao and Weifeng Qiu. The pointwise stabilities of piecewise linear finite element method on non-obtuse tetrahedral meshes of nonconvex polyhedra. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01465-
4;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01465-4.pdf.

Li:2021:NIR

[2792] Kun Li, Ting-Zhu Huang, Liang Li, and Stéphane Lanteri. Nonintrusive reduced-order modeling of parameterized electromagnetic scattering problems using cubic spline interpolation. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01467-2;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01467-2.pdf.

Qiao:2021:NLB

[2793] Zhonghua Qiao, Xuguang Yang, and Yuze Zhang. A novel lattice Boltzmann model for fourth order nonlinear partial differential equations. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01471-6;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01471-6.pdf.

Mei:2021:RBC

[2794] Tie Mei, Jiangming Zhao, Zian Liu, Xuhao Peng, Ziguang Chen, and Florin Bobaru. The role of boundary conditions on convergence properties of peridynamic model for transient heat transfer. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01469-0;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01469-0.pdf.

Simon:2021:SLS

[2795] Konrad Simon and Jörn Behrens. Semi-Lagrangian subgrid reconstruction for advection-dominant multiscale problems with rough data. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01451-w;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01451-w.pdf.

Busto:2021:HOA

[2796] Saray Busto, Michael Dumbser, Cipriano Escalante, Nicolas Favrie, and Sergey Gavrilyuk. On high order ADER discontinuous Galerkin schemes for first order hyperbolic reformulations of nonlinear dispersive systems. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01429-8;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01429-8.pdf.

Chang:2021:GRP

[2797] Xiaokai Chang and Junfeng Yang. A golden ratio primal–dual algorithm for structured convex optimization. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01452-9;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01452-9.pdf.

Leng:2021:AHMb

[2798] Haitao Leng and Huangxin Chen. Adaptive HDG methods for the Brinkman equations with application to optimal control. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01450-x;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01450-x.pdf.

Zhang:2021:HOM

[2799] Fan Zhang, Tiegang Liu, and Moubin Liu. A high-order maximumprinciple-satisfying discontinuous Galerkin method for the level set problem. Journal of Scientific Computing, 87(2):??, May 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01459-
2;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01459-2.pdf.

Zhu:2021:FSG

[2800] Xiaozhi Zhu and Yong-Tao Zhang. Fast sparse grid simulations of fifth order WENO scheme for high dimensional hyperbolic PDEs. Journal of Scientific Computing, 87(2):??, May 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01444-9;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01444-9.pdf.

Palitta:2021:MET

[2801] Davide Palitta. Matrix equation techniques for certain evolutionary partial differential equations. Journal of Scientific Computing, 87(3): ??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01515-x;   https://link.springer.com/content/pdf/
10.1007/s10915-021-01515-x.pdf.

Li:2021:UMB

[2802] Jingwei Li, Lili Ju, Yongyong Cai, and Xinlong Feng. Unconditionally maximum bound principle preserving linear schemes for the conservative Allen–Cahn equation with nonlocal constraint. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01512-0;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01512-0.pdf.

Lepe:2021:EEF

[2803] Felipe Lepe, Enrique Otárola, and Daniel Quero. Error estimates for FEM discretizations of the Navier–Stokes equations with Dirac measures. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01496-x;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01496-x.pdf.

Luo:2021:QCD

[2804] Dongmi Luo, Jianxian Qiu, Jun Zhu, and Yibing Chen. A quasiconservative discontinuous Galerkin method for multi-component flows using the non-oscillatory kinetic flux. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01494-z;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01494-z.pdf.

Chow:2021:LSS

[2805] Kevin Chow and Steven J. Ruuth. Linearly stabilized schemes for the time integration of stiff nonlinear PDEs. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01477-0;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01477-0.pdf.

Chen:2021:BDF

[2806] Minghua Chen, Fan Yu, and Zhi Zhou. Backward difference formulae: The energy technique for subdiffusion equation. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01509-9;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01509-9.pdf.

Hien:2021:ANM

[2807] Le Thi Khanh Hien and Nicolas Gillis. Algorithms for nonnegative matrix factorization with the Kullback–Leibler divergence. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01504-0;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01504-0.pdf.

Liu:2021:EPE

[2808] Hailiang Liu and Wumaier Maimaitiyiming. Efficient, positive, and energy stable schemes for multi-D Poisson–Nernst–Planck systems. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01503-1;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01503-1.pdf.

Burman:2021:CAH

[2809] Erik Burman, Omar Duran, Alexandre Ern, and Morgane Steins. Convergence analysis of hybrid high-order methods for the wave equation. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01492-1;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01492-1.pdf.

Le:2021:QRM

[2810] Thuy T. Le, Loc H. Nguyen, Thi-Phong Nguyen, and William Powell. The quasi-reversibility method to numerically solve an inverse source problem for hyperbolic equations. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01501-3;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01501-3.pdf.

Yang:2021:UOE

[2811] Yun-Bo Yang, Yao-Lin Jiang, and Bo-Hao Yu. Unconditional optimal error estimates of linearized, decoupled and conservative Galerkin FEMs for the Klein–Gordon–Schrödinger equation. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01510-2;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01510-2.pdf.

Zhang:2021:HOW

[2812] Min Zhang, Weizhang Huang, and Jianxian Qiu. A high-order well-balanced positivity-preserving moving mesh DG method for the shallow water equations with non-flat bottom topography. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01490-3;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01490-3.pdf.

Hao:2021:SHT

[2813] Wenrui Hao and Chunyue Zheng. A stochastic homotopy tracking algorithm for parametric systems of nonlinear equations. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01506-y;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01506-y.pdf.

Shi:2021:HOC

[2814] Yilei Shi, Shusen Xie, Dong Liang, and Kai Fu. High order compact block-centered finite difference schemes for elliptic and parabolic problems. Journal of Scientific Computing, 87(3):??, June 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01507-
x;   https://link.springer.com/content/pdf/10.1007/s10915-021-
01507-x.pdf.

Kozdon:2021:HSP

[2815] Jeremy E. Kozdon, Brittany A. Erickson, and Lucas C. Wilcox. Hybridized summation-by-parts finite difference methods. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01448-5;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01448-5.pdf.

Huang:2021:MWX

[2816] Xuehai Huang, Yuling Shi, and Wenqing Wang. A Morley–Wang–Xu element method for a fourth order elliptic singular perturbation problem. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01483-2;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01483-2.pdf.

Kang:2021:SIE

[2817] Myeongmin Kang and Miyoun Jung. Simultaneous image enhancement and restoration with non-convex total variation. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01488-x;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01488-x.pdf.

Youssef:2021:PSS

[2818] Maha Youssef and Roland Pulch. Poly-sinc solution of stochastic elliptic differential equations. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic).

URL https://link.springer.com/article/10.1007/s10915-021-
01498-9;   https://link.springer.com/content/pdf/10.1007/s10915-
021-01498-9.pdf.

Teng:2021:HOC

[2819] Long Teng and Weidong Zhao. High-order combined multi-step scheme for solving forward backward stochastic differential equations. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01505-z;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01505-z.pdf.

He:2021:MFE

[2820] Mingyan He and Pengtao Sun. Mixed finite element method for modified Poisson–Nernst–Planck/Navier–Stokes equations. Journal of Scientific Computing, 87(3):??, June 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01478-z;   https://
link.springer.com/content/pdf/10.1007/s10915-021-01478-z.pdf.

Jaramillo:2021:NMD

[2821] Gabriela Jaramillo, Loic Cappanera, and Cory Ward. Numerical methods for a diffusive class of nonlocal operators. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01543-7.

Funaro:2021:SCP

[2822] Daniele Funaro and Gianmarco Manzini. Stability and conservation properties of Hermite-based approximations of the Vlasov–Poisson system. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01537-5.

Busto:2021:TCF

[2823] Saray Busto, Michael Dumbser, and Kseniya Ivanova. On thermodynamically compatible finite volume methods and path-conservative ADER discontinuous Galerkin schemes for turbulent shallow water flows. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01521-z.

Zhang:2021:CSB

[2824] Jin Zhang and Xiaowei Liu. Convergence and supercloseness in a balanced norm of finite element methods on Bakhvalov-type meshes for reaction–diffusion problems. Journal of Scientific Computing, 88(1): ??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01542-8.

Lin:2021:EHB

[2825] Lanyu Lin and Yong-Jin Liu. An efficient Hessian based algorithm for singly linearly and box constrained least squares regression. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01541-9.

Kometa:2021:OPT

[2826] Bawfeh Kingsley Kometa, Naveed Iqbal, and Adel A. Attiya. Optimal parameters for third order Runge–Kutta exponential integrators for convection–diffusion problems. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01523-x.

Guo:2021:UNM

[2827] Hailong Guo, Xu Yang, and Yi Zhu. Unfitted Nitsche’s method for computing wave modes in topological materials. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01540-w.

Bonaventura:2021:SOF

[2828] Luca Bonaventura, Elisa Calzola, and Roberto Ferretti. Second order fully semi-Lagrangian discretizations of advection–diffusion–reaction systems. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01518-8.

Geist:2021:NSP

[2829] Moritz Geist, Philipp Petersen, and Gitta Kutyniok. Numerical solution of the parametric diffusion equation by deep neural networks. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01532-w.

Gautier:2021:GCN

[2830] Antoine Gautier, Matthias Hein, and Francesco Tudisco. The global convergence of the nonlinear power method for mixed-subordinate matrix norms. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01524-w.

Li:2021:LHO

[2831] Xin Li, Yuezheng Gong, and Luming Zhang. Linear high-order energypreserving schemes for the nonlinear Schrödinger equation with wave operator using the scalar auxiliary variable approach. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01533-9.

Wang:2021:UED

[2832] Nan Wang, Meng Li, and Chengming Huang. Unconditional energy dissipation and error estimates of the SAV Fourier spectral method for nonlinear fractional generalized wave equation. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01534-8.

Lindeberg:2021:HOA

[2833] Ludvig Lindeberg, Tuan Dao, and Ken Mattsson. A high order accurate finite difference method for the Drinfel’d–Sokolov–Wilson equation. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01481-4.

Gu:2021:OJR

[2834] Dong qin Gu and Zhong qing Wang. Orthogonal Jacobi rational functions and spectral methods on the half line. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01535-7.

Hu:2021:OMA

[2835] Jun Hu and Hua Wang. An optimal multigrid algorithm for the combining P_{1}–Q_{1} finite element approximations of interface problems based on local anisotropic fitting meshes. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01536-6.

Campagna:2021:DDE

[2836] Rosanna Campagna and Emma Perracchione. Data-driven extrapolation via feature augmentation based on variably scaled thin plate splines. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01526-8.

Wang:2021:AVE

[2837] Qiming Wang and Zhaojie Zhou. Adaptive virtual element method for optimal control problem governed by general elliptic equation. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01528-6.

Parkinson:2021:RGU

[2838] Christian Parkinson. A rotating-grid upwind fast sweeping scheme for a class of Hamilton–Jacobi equations. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01531-x.

Gu:2021:RHO

[2839] Yaguang Gu, Zhen Gao, and Lifeng Wang. A robust high order alternative WENO scheme for the five-equation model. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01529-5.

Zhao:2021:PTA

[2840] Yong-Liang Zhao, Xian-Ming Gu, and Alexander Ostermann. A preconditioning technique for an all-at-once system from Volterra subdiffusion equations with graded time steps. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01527-7.

Chen:2021:AVJ

[2841] Yangang Chen and Justin W. L. Wan. Artificial viscosity joint spacetime multigrid method for Hamilton–Jacobi–Bellman and Kolmogorov– Fokker–Planck system arising from mean field games. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01520-0.

Sheng:2021:SHD

[2842] Qiwei Sheng and Cheng Wang. A spherical harmonic discontinuous Galerkin method for radiative transfer equations with vacuum boundary conditions. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01530-y.

Pan:2021:ALE

[2843] Liang Pan and Kun Xu. An arbitrary-Lagrangian–Eulerian high-order gas-kinetic scheme for three-dimensional computations. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01525-9.

Chiapolino:2021:MSH

[2844] Alexandre Chiapolino, François Fraysse, and Richard Saurel. A method to solve Hamilton–Jacobi type equation on unstructured meshes. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01517-9.

Berg:2021:LTS

[2845] André Berg, David Cohen, and Guillaume Dujardin. Lie–Trotter splitting for the nonlinear stochastic Manakov system. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01514-y.

Zhou:2021:TSA

[2846] Qin Zhou and Binjie Li. Temporally semidiscrete approximation of a Dirichlet boundary control for a fractional/normal evolution equation with a final observation. Journal of Scientific Computing, 88(1): ??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01522-y.

Sheng:2021:ESM

[2847] Changtao Sheng, Duo Cao, and Jie Shen. Efficient spectral methods for PDEs with spectral fractional Laplacian. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01491-2.

Kopriva:2021:SDG

[2848] David A. Kopriva, Gregor J. Gassner, and Jan Nordström. Stability of discontinuous Galerkin spectral element schemes for wave propagation when the coefficient matrices have jumps. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01516-w.

Figueiredo:2021:MMH

[2849] J. Figueiredo and S. Clain. A MOOD–MUSCL hybrid formulation for the non-conservative shallow-water system. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01513-z.

Li:2021:HOA

[2850] Yetong Li, Tengjin Zhao, and Tongke Wang. The high order augmented finite volume methods based on series expansion for nonlinear degenerate parabolic equations. Journal of Scientific Computing, 88(1):??, July 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01519-7.

Feng:2021:IEB

[2851] Yue Feng. Improved error bounds of the Strang splitting method for the highly oscillatory fractional nonlinear Schrödinger equation. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01558-0.

Li:2021:SPI

[2852] Peng Li, Bao-Shan Wang, and Wai-Sun Don. Sensitivity parameterindependent characteristic-wise well-balanced finite volume WENO scheme for the Euler equations under gravitational fields. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01562-4.

Zhu:2021:FOI

[2853] Wei Zhu. A first-order image restoration model that promotes image contrast preservation. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01557-1.

Zhu:2021:AAL

[2854] Hong Zhu, Michael K. Ng, and Guang-Jing Song. An approximate augmented Lagrangian method for nonnegative low-rank matrix approximation. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01556-2. See

correction [3015].

Lepe:2021:VEM

[2855] Felipe Lepe, David Mora, and Iván Velásquez. A virtual element method for the Steklov eigenvalue problem allowing small edges. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01555-3.

Shen:2021:TCA

[2856] Xiaowen Shen and Qi Wang. Thermodynamically consistent algorithms for models of diblock copolymer solutions interacting with electric and magnetic fields. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01470-7.

Deng:2021:PDC

[2857] Shengxiang Deng and Hongpeng Sun. A preconditioned difference of convex algorithm for truncated quadratic regularization with application to imaging. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01547-3.

Wang:2021:FCA

[2858] Haifeng Wang, Jian-Feng Cai, and Ke Wei. Fast Cadzow’s algorithm and a gradient variant. Journal of Scientific Computing, 88 (2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01550-8.

deFrutos:2021:CFD

[2859] Javier de Frutos, Bosco García-Archilla, and Julia Novo. Corrigenda: Fully discrete approximations to the time-dependent Navier–Stokes equations with a projection method in time and grad-div stabilization. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01551-7.

Zhang:2021:CIA

[2860] Xiaodi Zhang and Qianqian Ding. Coupled iterative analysis for stationary inductionless magnetohydrodynamic system based on chargeconservative finite element method. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01553-5.

Wang:2021:ALC

[2861] Dawei Wang, Yunhui He, and Hans De Sterck. On the asymptotic linear convergence speed of Anderson acceleration applied to ADMM. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01548-2.

Huang:2021:HOF

[2862] Qian-Min Huang, Yu-Xin Ren, and Qian Wang. High order finite volume schemes for solving the non-conservative convection equations on the unstructured grids. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01538-4.

Dektor:2021:RAT

[2863] Alec Dektor, Abram Rodgers, and Daniele Venturi. Rank-adaptive tensor methods for high-dimensional nonlinear PDEs. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01539-3.

Burman:2021:CSD

[2864] Erik Burman, Cuiyu He, and Mats G. Larson. Comparison of shape derivatives using CutFEM for ill-posed Bernoulli free boundary problem. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01544-6.

Dedner:2021:REP

[2865] Andreas Dedner, Jan Giesselmann, and Jennifer K Ryan. Residual estimates for post-processors in elliptic problems. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01502-2.

Wang:2021:SOA

[2866] Min Wang, Qiumei Huang, and Cheng Wang. A second order accurate scalar auxiliary variable (SAV) numerical method for the square phase field crystal equation. Journal of Scientific Computing, 88(2): ??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-021-01487-y.

Che:2021:ERA

[2867] Maolin Che, Yimin Wei, and Hong Yan. An efficient randomized algorithm for computing the approximate Tucker decomposition. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01545-5.

Feng:2021:AFD

[2868] Xiaobing Feng and Hailong Qiu. Analysis of fully discrete mixed finite element methods for time-dependent stochastic Stokes equations with multiplicative noise. Journal of Scientific Computing, 88(2):??, August 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01546-4.

Uy:2021:OIN

[2869] Wayne Isaac Tan Uy and Benjamin Peherstorfer. Operator inference of non-Markovian terms for learning reduced models from partially observed state trajectories. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01580-2.

Zhang:2021:MSM

[2870] Shengliang Zhang. Meshless symplectic and multi-symplectic local RBF collocation methods for Hamiltonian PDEs. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-021-01605-w.

Kheirfam:2021:PII

[2871] B. Kheirfam. A polynomial-iteration infeasible interior-point algorithm with arc-search for semidefinite optimization. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-021-01609-6.

Ventura:2021:CIP

[2872] Jordi Ventura, David Codony, and Sonia Fernández-Méndez. A C0 interior penalty finite element method for flexoelectricity. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01613-w.

Shen:2021:RCD

[2873] Hua Shen and Matteo Parsani. A rotated characteristic decomposition technique for high-order reconstructions in multi-dimensions. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01602-z.

Patane:2021:SFM

[2874] Giuseppe Patané. Spectrum-free and meshless solvers of parabolic PDEs. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01604-x.

Cholamjiak:2021:RFB

[2875] Prasit Cholamjiak, Dang Van Hieu, and Yeol Je Cho. Relaxed forward-backward splitting methods for solving variational inclusions and applications. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01608-7.

Prieto-Arranz:2021:WBS

[2876] Alberto Prieto-Arranz, Luis Ramírez, and Xesús Nogueira. A wellbalanced SPH-ALE scheme for shallow water applications. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01600-1.

Xu:2021:CCR

[2877] Qi Xu, Zaihua Wang, and Li Cheng. Calculating characteristic roots of multi-delayed systems with accumulation points via a definite integral method. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01599-5.

Yu:2021:LPM

[2878] Unjong Yu, Young-Min Lee, and Chi-Ok Hwang. Last-passage Monte Carlo algorithm for charge density on a conducting spherical surface. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01594-w.

Yang:2021:TGC

[2879] Jiming Yang and Yifan Su. A two-grid combined mixed finite element and discontinuous Galerkin method for an incompressible miscible displacement problem in porous media. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01596-8.

Wang:2021:FTG

[2880] Yang Wang, Yanping Chen, and Huaming Yi. A family of two-grid partially penalized immersed finite element methods for semi-linear parabolic interface problems. Journal of Scientific Computing, 88(3): ??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01575-z.

Taddei:2021:RBM

[2881] Tommaso Taddei and Lei Zhang. Registration-based model reduction in complex two-dimensional geometries. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01584-y.

Li:2021:HKI

[2882] Hongliang Li, Pingbing Ming, and Huiyu Wang. H^{2}–Korn’s inequality and the nonconforming elements for the strain gradient elastic model. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01597-7.

Li:2021:RBP

[2883] Yuwen Li. Recovery-based a posteriori error analysis for plate bending problems. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01595-9.

Quenjel:2021:AAS

[2884] El Houssaine Quenjel. Analysis of accurate and stable nonlinear finite volume scheme for anisotropic diffusion equations with drift on simplicial meshes. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01577-x.

Li:2021:DPR

[2885] Qin Li, Pan Yan, and Yancheng You. On developing piecewise rational mapping with fine regulation capability for WENO schemes. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01559-z.

Qi:2021:SOE

[2886] Longzhao Qi and Yanren Hou. A second order energy stable BDF numerical scheme for the Swift–Hohenberg equation. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01593-x.

Beck:2021:SKP

[2887] Christian Beck, Sebastian Becker, and Arnulf Jentzen. Solving the Kolmogorov PDE by means of deep learning. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-021-01590-0.

Law:2021:FSM

[2888] Yann-Meing Law and Jean-Christophe Nave. FDTD schemes for Maxwell’s equations with embedded perfect electric conductors based on the correction function method. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01591-z.

Ding:2021:EEF

[2889] Qianqian Ding, Xiaonian Long, and Shipeng Mao. Error estimate of a fully discrete finite element method for incompressible vector potential magnetohydrodynamic system. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01571-3.

Li:2021:SOC

[2890] Buyang Li, Shu Ma, and Na Wang. Second-order convergence of the linearly extrapolated Crank–Nicolson method for the Navier–Stokes equations with H^{1} initial data. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01588-8.

Deng:2021:SOS

[2891] Beichuan Deng, Jie Shen, and Qingqu Zhuang. Second-order SAV schemes for the nonlinear Schrödinger equation and their error analysis. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01576-y.

Yan:2021:EEC

[2892] Yuyuan Yan, Bernard A. Egwu, and Yubin Yan. Error estimates of a continuous Galerkin time stepping method for subdiffusion problem. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01587-9.

Pan:2021:HOC

[2893] Kejia Pan, Dongdong He, and Zhilin Li. A high order compact FD framework for elliptic BVPs involving singular sources, interfaces, and irregular domains. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01570-4.

Ait-Haddou:2021:COL

[2894] Rachid Ait-Haddou. Computation of optimal linear strong stability preserving methods via adaptive spectral transformations of Poisson– Charlier measures. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01582-0.

Chen:2021:BBG

[2895] Yannan Chen, Xinzhen Zhang, and Yanwei Xu. A Barzilai–Borwein gradient algorithm for spatio-temporal Internet traffic data completion via tensor triple decomposition. Journal of Scientific Computing, 88(3): 65:1–65:24, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01574-0.

Yueh:2021:PGM

[2896] Mei-Heng Yueh, Tsung-Ming Huang, and Shing-Tung Yau. Projected gradient method combined with homotopy techniques for volumemeasure-preserving optimal mass transportation problems. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01583-z.

Tao:2021:SAU

[2897] Qi Tao, Waixiang Cao, and Zhimin Zhang. Superconvergence analysis of the ultra-weak local discontinuous Galerkin method for one dimensional linear fifth order equations. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01579-9.

Ma:2021:SDG

[2898] Limin Ma. Superconvergence of discontinuous Galerkin methods for elliptic boundary value problems. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01589-7.

Qu:2021:LSM

[2899] Feng Qu, Di Sun, and Junqiang Bai. Low-speed modification for the genuinely multidimensional Harten, Lax, van Leer and Einfeldt scheme in curvilinear coordinates. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01561-5.

Li:2021:NAT

[2900] Qi Li and Liquan Mei. Numerical approximation of the two-component PFC models for binary colloidal crystals: Efficient, decoupled, and second-order unconditionally energy stable schemes. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01564-2.

Fu:2021:EFE

[2901] Taibai Fu, Beiping Duan, and Zhoushun Zheng. An effective finite element method with singularity reconstruction for fractional convectiondiffusion equation. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01573-1.

Chen:2021:TES

[2902] Minghua Chen, Suzhen Jiang, and Weiping Bu. Two L_{1} schemes on graded meshes for fractional Feynman–Kac equation. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01581-1.

Zhang:2021:PPW

[2903] Weijie Zhang, Yinhua Xia, and Yan Xu. Positivity-preserving wellbalanced arbitrary Lagrangian–Eulerian discontinuous Galerkin methods for the shallow water equations. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01578-w.

Luo:2021:GCN

[2904] Xin long Luo and Hang Xiao. Generalized continuation Newton methods and the trust-region updating strategy for the underdetermined system. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01566-0.

Giani:2021:PEE

[2905] Stefano Giani, Luka Grubisić, and Jeffrey S. Ovall. A posteriori error estimates for elliptic eigenvalue problems using auxiliary subspace techniques. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01572-2.

Ho:2021:FHS

[2906] Kaho Ho, Shingyu Leung, and Jianliang Qian. Fast Huygens sweeping methods for a class of nonlocal Schrödinger equations. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01549-1.

Liu:2021:CRC

[2907] Zexin Liu and Akil Narayan. On the computation of recurrence coefficients for univariate orthogonal polynomials. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01586-w.

Khan:2021:CND

[2908] Arbaz Khan, Manil T. Mohan, and Ricardo Ruiz-Baier. Conforming, nonconforming and DG methods for the stationary generalized Burgers–Huxley equation. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-021-01563-3.

Duru:2021:NDG

[2909] Kenneth Duru, Leonhard Rannabauer, and Heiner Igel. A new discontinuous Galerkin method for elastic waves with physically motivated numerical fluxes. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01565-1.

Xie:2021:SCM

[2910] Zhongbing Xie, Gang Cai, and Qiao-Li Dong. Strong convergence of the modified inertial extragradient method with line-search process for solving variational inequality problems in Hilbert spaces. Journal of Scientific Computing, 88(3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01585-x.

Gao:2021:SOF

[2911] Yijin Gao, Jay Mayfield, and Songting Luo. A second-order fast Huygens sweeping method for time-dependent Schrödinger equations with perfectly matched layers. Journal of Scientific Computing, 88 (3):??, September 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01560-6.

Tosi:2021:PDA

[2912] Riccardo Tosi, Ramon Amela, and Riccardo Rossi. A parallel dynamic asynchronous framework for uncertainty quantification by hierarchical Monte Carlo algorithms. Journal of Scientific Computing, 89 (1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01598-6.

vanGestel:2021:ECE

[2913] R. A. M. van Gestel, M. J. H. Anthonissen, and W. L. IJzerman. An energy conservative hp-method for Liouville’s equation of geometrical optics. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01612-x. See

correction [3058].

Mu:2021:DPR

[2914] Lin Mu, Xiu Ye, and Shangyou Zhang. Development of pressure-robust discontinuous Galerkin finite element methods for the Stokes problem. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01634-5.

Prazeres:2021:SGD

[2915] Mariana Prazeres and Adam M. Oberman. Stochastic gradient descent with Polyak’s learning rate. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01628-3.

Du:2021:TGA

[2916] Binbin Du, Jianguo Huang, and Haibiao Zheng. Two-grid arrow-Hurwicz methods for the steady incompressible Navier–Stokes equations. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01627-4.

Desquilbet:2021:SPC

[2917] François Desquilbet, Jian Cao, and Jean-Marie Mirebeau. Single pass computation of first seismic wave travel time in three dimensional heterogeneous media with general anisotropy. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01607-8.

Guglielmi:2021:PRC

[2918] Nicola Guglielmi, María López-Fernández, and Mattia Manucci. Pseudospectral roaming contour integral methods for convection–diffusion equations. Journal of Scientific Computing, 89(1):??, October 2021. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01601-
0.

Cho:2021:TWS

[2919] Sung Woong Cho, Hyung Ju Hwang, and Hwijae Son. Traveling wave solutions of partial differential equations via neural networks. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01621-w.

Benedusi:2021:FPS

[2920] Pietro Benedusi, Paola Ferrari, and Stefano Serra-Capizzano. Fast parallel solver for the space-time IgA-DG discretization of the diffusion equation. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01567-z.

Bonazzoli:2021:ASD

[2921] Marcella Bonazzoli, Xavier Claeys, and Pierre-Henri Tournier. Analysis of the SORAS domain decomposition preconditioner for non-self-adjoint or indefinite problems. Journal of Scientific Computing, 89 (1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01631-8.

Liu:2021:UTS

[2922] Yang Liu, Baoli Yin, and Zhimin Zhang. The unified theory of shifted convolution quadrature for fractional calculus. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01630-9.

Park:2021:PAG

[2923] Jea-Hyun Park, Abner J. Salgado, and Steven M. Wise. Preconditioned accelerated gradient descent methods for locally Lipschitz smooth objectives with applications to the solution of nonlinear PDEs. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01615-8.

Dehghan:2021:VEM

[2924] Mehdi Dehghan and Zeinab Gharibi. Virtual element method for solving an inhomogeneous Brusselator model with and without crossdiffusion in pattern formation. Journal of Scientific Computing, 89 (1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01626-5.

Chen:2021:BEI

[2925] Hongjia Chen and Kuan Xu. On the backward error incurred by the compact rational Krylov linearization. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01625-6.

Feireisl:2021:NAM

[2926] Eduard Feireisl, Madalina Petcu, and Bangwei She. Numerical analysis of a model of two phase compressible fluid flow. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01624-7.

Jian:2021:FIW

[2927] Huan-Yan Jian, Ting-Zhu Huang, and Yong-Liang Zhao. Fast IIF-WENO method on non-uniform meshes for nonlinear space-fractional convection–diffusion–reaction equations. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01622-9.

Gu:2021:PMS

[2928] Shuting Gu, Ling Lin, and Xiang Zhou. Projection method for saddle points of energy functional in H^{−1} metric. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01592-y.

Nasab:2021:ORK

[2929] Siavash Hedayati Nasab, Carlos A. Pereira, and Brian C. Vermeire. Optimal Runge–Kutta stability polynomials for multidimensional highorder methods. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01620-x.

Shao:2021:QTT

[2930] Xin-Hui Shao, Yu-Han Li, and Hai-Long Shen. Quasi-Toeplitz trigonometric transform splitting methods for spatial fractional diffusion equations. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01610-z.

Karatzas:2021:ROM

[2931] Efthymios N. Karatzas and Gianluigi Rozza. A reduced order model for a stable embedded boundary parametrized Cahn–Hilliard phase-field system based on cut finite elements. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01623-8.

Han:2021:MFE

[2932] Weimin Han, Kenneth Czuprynski, and Feifei Jing. Mixed finite element method for a hemivariational inequality of stationary Navier– Stokes equations. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01614-9.

Liu:2021:HOB

[2933] Xinyuan Liu, Yang Yang, and Hui Guo. High-order bound-preserving finite difference methods for incompressible wormhole propagation. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01619-4.

Law:2021:DLD

[2934] Ho Law, Chun Yin Siu, and Lok Ming LUI. Decomposition of longitudinal deformations via Beltrami descriptors. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01569-x.

Gu:2021:CRP

[2935] Yaguang Gu and Felix Kwok. On the choice of Robin parameters for the optimized Schwarz method for domains with non-conforming heterogeneities. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01617-6.

Yang:2021:VBS

[2936] Zhen-Ping Yang, Jin Zhang, and Gui-Hua Lin. Variance-based subgradient extragradient method for stochastic variational inequality problems. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01603-y.

LeFloch:2021:CWB

[2937] Philippe G. LeFloch, Carlos Parés, and Ernesto Pimentel-García. A class of well-balanced algorithms for relativistic fluids on a Schwarzschild background. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01611-
y.

Kopriva:2021:SFS

[2938] David A. Kopriva and Gregor J. Gassner. A split-form, stable CG/DG-SEM for wave propagation modeled by linear hyperbolic systems. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01618-5.

Borcea:2021:ROM

[2939] Liliana Borcea, Vladimir Druskin, and Jörn Zimmerling. A reduced order model approach to inverse scattering in lossy layered media. Journal of Scientific Computing, 89(1):??, October 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01616-7.

Chan:2021:MBE

[2940] Jesse Chan, Mario J. Bencomo, and David C. Del Rey Fernández. Mortar-based entropy-stable discontinuous Galerkin methods on nonconforming quadrilateral and hexahedral meshes. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01652-3.

Copur:2021:EIT

[2941] Aysegül Keten Çopur, Emirhan Hacioglu, and Müzeyyen Ertürk. An efficient inertial type iterative algorithm to approximate the solutions of quasi variational inequalities in real Hilbert spaces. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01657-y.

Lyu:2021:SON

[2942] Pin Lyu and Seakweng Vong. Second-order and nonuniform timestepping schemes for time fractional evolution equations with timespace dependent coefficients. Journal of Scientific Computing, 89(2): ??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01661-2.

Chertock:2021:HMA

[2943] Alina Chertock, Shaoshuai Chu, and Alexander Kurganov. Hybrid multifluid algorithms based on the path-conservative central-upwind scheme. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01656-z.

Zhang:2021:ESP

[2944] Yanrong Zhang and Jie Shen. Efficient structure preserving schemes for the Klein–Gordon-Schrödinger equations. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-021-01649-y.

Bellavia:2021:RIP

[2945] Stefania Bellavia, Jacek Gondzio, and Margherita Porcelli. A relaxed interior point method for low-rank semidefinite programming problems with applications to matrix completion. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01654-1.

Lyu:2021:ESN

[2946] Maohui Lyu, Vrushali A. Bokil, and Fengyan Li. Energy stable nodal discontinuous Galerkin methods for nonlinear Maxwell’s equations in multi-dimensions. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01651-4.

Fabiani:2021:NSB

[2947] Gianluca Fabiani, Francesco Calabrò, and Constantinos Siettos. Numerical solution and bifurcation analysis of nonlinear partial differential equations with extreme learning machines. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01650-5.

Jiang:2021:ESM

[2948] Shan Jiang and Haijun Yu. Efficient spectral methods for quasiequilibrium closure approximations of symmetric problems on unit circle and sphere. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01646-1.

Tang:2021:ABT

[2949] Shaoqiang Tang and Gang Pang. Accurate boundary treatment for Riesz space fractional diffusion equations. Journal of Scientific Computing, 89 (2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01655-0.

Olshanskii:2021:UFE

[2950] Maxim Olshanskii, Annalisa Quaini, and Qi Sun. An unfitted finite element method for two-phase Stokes problems with slip between phases. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01658-x.

Zhao:2021:NIC

[2951] Xiaofei Zhao. Numerical integrators for continuous disordered nonlinear Schrödinger equation. Journal of Scientific Computing, 89(2): ??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01653-2.

Manohar:2021:PEE

[2952] Ram Manohar and Rajen Kumar Sinha. A posteriori error estimates for parabolic optimal control problems with controls acting on lower dimensional manifolds. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01645-2.

Franz:2021:SPR

[2953] Sebastian Franz. Singularly perturbed reaction–diffusion problems as first order systems. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01638-1.

Kim:2021:CAC

[2954] Seungil Kim and Hui Zhang. Convergence analysis of the continuous and discrete non-overlapping double sweep domain decomposition method based on PMLs for the Helmholtz equation. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-021-01640-7.

Du:2021:AFE

[2955] Shaohong Du and Zhiqiang Cai. Adaptive finite element method for Dirichlet boundary control of elliptic partial differential equations. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01644-3.

Li:2021:HMM

[2956] Weiming Li, Peng Song, and Yanli Wang. A hybrid moment method for multi-scale kinetic equations based on maximum entropy principle. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01639-0.

Mizuguchi:2021:ECS

[2957] Makoto Mizuguchi, Mitsuhiro T. Nakao, and Shin’ichi Oishi. Error constants for the semi-discrete Galerkin approximation of the linear heat equation. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01636-3.

Li:2021:BES

[2958] Dong Li, Chaoyu Quan, and Wen Yang. The BDF3/EP3 scheme for MBE with no slope selection is stable. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01642-5. See correction [2968].

Sun:2021:OFD

[2959] Zhensheng Sun, Yu Hu, and Kai Mao. An optimal finite difference scheme with minimized dispersion and adaptive dissipation considering the spectral properties of the fully discrete scheme. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01637-2.

Michel:2021:SAC

[2960] Sixtine Michel, Davide Torlo, and Rémi Abgrall. Spectral analysis of continuous FEM for hyperbolic PDEs: Influence of approximation, stabilization, and time-stepping. Journal of Scientific Computing, 89 (2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01632-7.

Eriksson:2021:ISS

[2961] Sofia Eriksson. Inverses of SBP-SAT finite difference operators approximating the first and second derivative. Journal of Scientific Computing, 89(2):??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01606-9.

Yeager:2021:TSR

[2962] Benjamin Yeager, Ethan Kubatko, and Dylan Wood. Time step restrictions for strong-stability-preserving multistep Runge–Kutta discontinuous Galerkin methods. Journal of Scientific Computing, 89(2): ??, November 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01635-4.

Li:2021:NTR

[2963] Minghui Li, Wen Li, and Mingqing Xiao. The nonconvex tensor robust principal component analysis approximation model via the weighted ℓ_{p}-norm regularization. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01679-6.

Fishelov:2021:OCT

[2964] D. Fishelov and J.-P. Croisille. Optimal convergence for time-dependent Stokes equation: a new approach. Journal of Scientific Computing, 89 (3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01684-9.

Qin:2021:NSC

[2965] Hongyu Qin, Dongfang Li, and Zhimin Zhang. A novel scheme to capture the initial dramatic evolutions of nonlinear subdiffusion equations. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01672-z.

Tsoutsanis:2021:CFV

[2966] Panagiotis Tsoutsanis, Ebenezer Mayowa Adebayo, and Martin Skote. CWENO finite-volume interface capturing schemes for multicomponent flows using unstructured meshes. Journal of Scientific Computing, 89 (3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01673-y.

Chellappa:2021:TSS

[2967] Sridhar Chellappa, Lihong Feng, and Peter Benner. A training set subsampling strategy for the reduced basis method. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01665-y.

Li:2021:CBE

[2968] Dong Li, Chaoyu Quan, and Wen Yang. Correction to: The BDF3/EP3 Scheme for MBE with No Slope Selection is Stable. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01666-x. See [2958].

Feng:2021:FRH

[2969] Lihong Feng, Guosheng Fu, and Zhu Wang. A FOM/ROM hybrid approach for accelerating numerical simulations. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01668-9.

Shi:2021:ESO

[2970] YuFeng Shi and Yan Guo. An essential seventh-order weighted compact adaptive scheme for hyperbolic conservation laws. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01664-z.

Yin:2021:SRL

[2971] Jia Yin and Chunxiong Zheng. Space reduction for linear systems with local symmetry. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01663-0.

Park:2021:AAS

[2972] Jongho Park. Accelerated additive Schwarz methods for convex optimization with adaptive restart. Journal of Scientific Computing, 89 (3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01648-z.

Zhang:2021:TGB

[2973] Jing Zhang, Shengfeng Zhu, and Xiaoqin Shen. A two-grid binary level set method for eigenvalue optimization. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01662-1.

Kamm:2021:SME

[2974] Kevin Kamm, Stefano Pagliarani, and Andrea Pascucci. On the stochastic Magnus expansion and its application to SPDEs. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01633-6.

Sanchez:2021:NSD

[2975] Cipriano Escalante Sánchez, Enrique D. Fernández-Nieto, and Jacques Sainte-Marie. Numerical simulations of a dispersive model approximating free-surface Euler equations. Journal of Scientific Computing, 89 (3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01552-6.

Lin:2021:ICB

[2976] Xue lei Lin and Michael K. Ng. Image completion and blind deconvolution: Model and algorithm. Journal of Scientific Computing, 89 (3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01554-4.

Xie:2021:PPF

[2977] Hui Xie, Xuejun Xu, and Heng Yong. A positivity-preserving finite volume scheme with least square interpolation for 3D anisotropic diffusion equation. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01629-2.

Gu:2021:HOF

[2978] Yaguang Gu, Zhen Gao, and Lifeng Wang. High order finite difference alternative WENO scheme for multi-component flows. Journal of Scientific Computing, 89(3):??, December 2021. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01659-w.

Zhang:2022:USO

[2979] Tong Zhang and JinYun Yuan. Unconditional stability and optimal error estimates of Euler implicit/explicit-SAV scheme for the Navier– Stokes equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01681-y.

McClarren:2022:SIH

[2980] Ryan G. McClarren, James A. Rossmanith, and Minwoo Shin. Semiimplicit hybrid discrete H_{N}^{T} approximation of thermal radiative transfer. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01686-7.

Qin:2022:TDC

[2981] Xiaoxue Qin, Luchan Zhang, and Yang Xiang. A three-dimensional continuum simulation method for grain boundary motion incorporating dislocation structure. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01694-7.

An:2022:SDP

[2982] Hengbin An, Zeyao Mo, and Xiaowen Xu. Shear decoupled parallel scalable preconditioners for nonlinear thermo-mechanical coupled contact applications. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01643-4.

Vismara:2022:SED

[2983] Federico Vismara, Tommaso Benacchio, and Luca Bonaventura. A seamless, extended DG approach for advection–diffusion problems on unbounded domains. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01669-8.

Lu:2022:LED

[2984] Lin Lu, Qi Wang, and Yushun Wang. Local energy dissipation rate preserving approximations to driven gradient flows with applications to graphene growth. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01676-9.

Huang:2022:ACA

[2985] Yakui Huang, Yu-Hong Dai, and Hongchao Zhang. On the asymptotic convergence and acceleration of gradient methods. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01685-8.

Zhang:2022:ETO

[2986] Hong Zhang, Jingye Yan, and Songhe Song. Explicit third-order unconditionally structure-preserving schemes for conservative Allen–Cahn equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01691-w. See

correction [3088].

Chertock:2022:WBF

[2987] Alina Chertock, Alexander Kurganov, and Tong Wu. Well-balancing via flux globalization: Applications to shallow water equations with Wet/Dry fronts. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01680-z.

Ogwo:2022:CRI

[2988] G. N. Ogwo, C. Izuchukwu, and O. T. Mewomo. Convergence of relaxed inertial subgradient extragradient methods for quasimonotone variational inequality problems. Journal of Scientific Computing, 90 (1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01670-1.

Wang:2022:AIG

[2989] Yifei Wang and Wuchen Li. Accelerated information gradient flow. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01709-3.

Alla:2022:TAM

[2990] Alessandro Alla, Carmen Gräßle, and Michael Hinze. Time adaptivity in model predictive control. Journal of Scientific Computing, 90 (1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01647-0.

Frederick:2022:CMP

[2991] Christina Frederick, Magnus Egerstedt, and Haomin Zhou. Collective motion planning for a group of robots using intermittent diffusion. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01700-y.

Choi:2022:FBC

[2992] Gary P. T. Choi, Yechen Liu, and Lok Ming Lui. Free-boundary conformal parameterization of point clouds. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01641-6.

Liu:2022:EEF

[2993] Huan Liu, Xiangcheng Zheng, and Hongfei Fu. Error estimate of finite element approximation for two-sided space-fractional evolution equation with variable coefficient. Journal of Scientific Computing, 90 (1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01698-3.

Cao:2022:PRR

[2994] Luling Cao, Yinnian He, and Jian Li. A parallel Robin–Robin domain decomposition method based on modified characteristic FEMs for the time-dependent dual-porosity-Navier–Stokes model with the Beavers– Joseph interface condition. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01674-x.

He:2022:USC

[2995] Yinnian He, Xiaojing Dong, and Xinlong Feng. Uniform stability and convergence with respect to (ν, μ, s, 1 − σ) of the three iterative finite element solutions for the 3D steady MHD equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01671-0.

Dong:2022:HMC

[2996] Bo Dong. The homotopy method for the complete solution of quadratic two-parameter eigenvalue problems. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01693-8.

Liu:2022:HAT

[2997] Xing Liu. High-accuracy time discretization of stochastic fractional diffusion equation. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01710-w.

Sun:2022:PTI

[2998] Yafei Sun, Shu-Lin Wu, and Yingxiang Xu. A parallel-in-time implementation of the Numerov method for wave equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01701-x. See correction [3145].

Zhou:2022:ALE

[2999] Lingling Zhou and Yinhua Xia. Arbitrary Lagrangian–Eulerian local discontinuous Galerkin method for linear convection–diffusion equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01697-4.

Diaz-Adame:2022:FOW

[3000] R. Diaz-Adame, S. Jerez, and H. Carrillo. Fast and optimal WENO schemes for degenerate parabolic conservation laws. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01689-4.

Antonietti:2022:SAP

[3001] Paola F. Antonietti, Lorenzo Mascotto, and Stefano Zonca. Stability analysis of polytopic discontinuous Galerkin approximations of the Stokes problem with applications to fluid-structure interaction problems. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01695-6.

Wang:2022:LFD

[3002] Yanyan Wang, Zhaopeng Hao, and Rui Du. A linear finite difference scheme for the two-dimensional nonlinear Schrödinger equation with fractional Laplacian. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01703-9.

Hou:2022:SOE

[3003] Dianming Hou and Chuanju Xu. A second order energy dissipative scheme for time fractional L^{2} gradient flows using SAV approach. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01667-w.

Li:2022:PBW

[3004] Fang Li, Yamin Ru, and Xiao-Guang Lv. Patch-based weighted SCAD prior for Rician noise removal. Journal of Scientific Computing, 90 (1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01688-5.

Zhu:2022:RTP

[3005] Wanwan Zhu, Ying Yang, and Benzhuo Lu. Residual type A posteriori error estimates for the time-dependent Poisson–Nernst–Planck equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01702-w.

Bacigaluppi:2022:ANC

[3006] Paola Bacigaluppi, Julien Carlier, and Rémi Abgrall. Assessment of a non-conservative four-equation multiphase system with phase transition. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01706-6.

Cao:2022:EAD

[3007] Shuhao Cao, Long Chen, and Xuehai Huang. Error analysis of a decoupled finite element method for quad-curl problems. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01705-7.

Meli:2022:SNS

[3008] Enrico Meli, Benedetta Morini, and Cristina Sgattoni. Solving nonlinear systems of equations via spectral residual methods: Stepsize selection and applications. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01690-x.

Wang:2022:ASG

[3009] Weina Wang and Yunmei Chen. An accelerated smoothing gradient method for nonconvex nonsmooth minimization in image processing. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01677-8.

Druskin:2022:DPM

[3010] Vladimir Druskin, Alexander V. Mamonov, and Mikhail Zaslavsky. Distance preserving model order reduction of graph-Laplacians and cluster analysis. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01660-3.

Aryeni:2022:ASG

[3011] T. Aryeni, Q. Deng, and V. Ginting. On the application of stable generalized finite element method for quasilinear elliptic two-point BVP. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01675-w.

Jiang:2022:PFE

[3012] Xue Jiang and Xiaoqi Duan. A PML finite element method for electromagnetic scattering problems in a two-layer medium. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01678-7.

Hamfeldt:2022:CFD

[3013] Brittany Froese Hamfeldt and Jacob Lesniewski. Convergent finite difference methods for fully nonlinear elliptic equations in three dimensions. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01714-6.

Waziri:2022:TDD

[3014] Mohammed Yusuf Waziri and Kabiru Ahmed. Two descent Dai–Yuan conjugate gradient methods for systems of monotone nonlinear equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01713-7.

Zhu:2022:CAA

[3015] Hong Zhu, Michael K. Ng, and Guang-Jing Song. Correction to: An Approximate Augmented Lagrangian Method for Nonnegative Low-Rank Matrix Approximation. Journal of Scientific Computing, 90 (1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01729-z. See [2854].

Lu:2022:FLE

[3016] Jianfeng Lu, Cody Murphey, and Stefan Steinerberger. Fast localization of eigenfunctions via smoothed potentials. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01682-x.

Jin:2022:NNN

[3017] Pengfei Jin, Tianhao Lai, and Bin Dong. NPTC-net: Narrow-band parallel transport convolutional neural networks on point clouds. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01699-2.

Abreu:2022:CPS

[3018] Eduardo Abreu, Jean François, and John Pérez. A class of positive semi-discrete Lagrangian–Eulerian schemes for multidimensional systems of hyperbolic conservation laws. Journal of Scientific Computing, 90 (1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01712-8.

Hendy:2022:RST

[3019] A. S. Hendy and K. Van Bockstal. On a reconstruction of a solely time-dependent source in a time-fractional diffusion equation with nonsmooth solutions. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01704-8.

Worku:2022:SFS

[3020] Zelalem Arega Worku and David W. Zingg. Stability and functional superconvergence of narrow-stencil second-derivative generalized summation-by-parts discretizations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01707-5.

Du:2022:NLP

[3021] Guangzhi Du, Liyun Zuo, and Yuhong Zhang. A new local and parallel finite element method for the coupled Stokes–Darcy model. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01723-5.

Huang:2022:BRS

[3022] Chaobao Huang, Hu Chen, and Na An. β -robust superconvergent analysis of a finite element method for the distributed order timefractional diffusion equation. Journal of Scientific Computing, 90(1): ??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-021-01726-2.

Ferro:2022:ARB

[3023] Nicola Ferro, Simona Perotto, and Andrea Cangiani. An anisotropic recovery-based error estimator for adaptive discontinuous Galerkin methods. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01724-4.

Luo:2022:QCD

[3024] Dongmi Luo, Shiyi Li, and Yibing Chen. A quasi-conservative discontinuous Galerkin method for multi-component flows using the nonoscillatory kinetic flux II: ALE framework. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01732-4.

Li:2022:UOE

[3025] Yuan Li and Rong An. Unconditionally optimal error analysis of a linear Euler FEM scheme for the Navier–Stokes equations with mass diffusion. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01730-6.

Kochi:2022:CSW

[3026] S. R. Siva Prasad Kochi and M. Ramakrishna. A compact subcell WENO limiting strategy using immediate neighbors for Runge–Kutta discontinuous Galerkin methods for unstructured meshes. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01725-3.

Huang:2022:NBM

[3027] Pengfei Huang and Qingzhi Yang. Newton-based methods for finding the positive ground state of Gross–Pitaevskii equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01711-9.

Dassi:2022:VEM

[3028] Franco Dassi, Alessio Fumagalli, and Giuseppe Vacca. A virtual element method for the wave equation on curved edges in two dimensions. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01683-w.

Lee:2022:HOU

[3029] Hyun Geun Lee, Jaemin Shin, and June-Yub Lee. A high-order and unconditionally energy stable scheme for the conservative Allen–Cahn equation with a nonlocal Lagrange multiplier. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01735-1.

Fernandez:2022:AHO

[3030] E. Guerrero Fernández, M. J. Castro Díaz, and T. Morales de Luna. An arbitrary high order well-balanced ADER-DG numerical scheme for the multilayer shallow-water model with variable density. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01734-2.

Buchheit:2022:ECL

[3031] Andreas A. Buchheit and Torsten Keßler. On the efficient computation of large scale singular sums with applications to long-range forces in crystal lattices. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01731-5.

Schutz:2022:PTH

[3032] Jochen Schütz, David C. Seal, and Jonas Zeifang. Parallel-in-time highorder multiderivative IMEX solvers. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01733-3.

Fu:2022:ELD

[3033] Lin Fu. An efficient low-dissipation high-order TENO scheme for MHD flows. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01722-6.

Zhou:2022:PEE

[3034] Jianwei Zhou, Huiyuan Li, and Zhimin Zhang. A posteriori error estimates of spectral approximations for second order partial differential equations in spherical geometries. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01696-5.

Fan:2022:TSI

[3035] Linwei Fan, Huiyu Li, and Caiming Zhang. Two-stage image denoising via an enhanced low-rank prior. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01728-0.

Colibazzi:2022:LNE

[3036] Francesco Colibazzi, Damiana Lazzaro, and Andrea Samoré. Learning nonlinear electrical impedance tomography. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01716-4.

Reichel:2022:TAT

[3037] Lothar Reichel and Ugochukwu O. Ugwu. Tensor Arnoldi-Tikhonov and GMRES-type methods for ill-posed problems with a t-product structure. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01719-1.

Buccini:2022:FAD

[3038] Alessandro Buccini. Fast alternating direction multipliers method by generalized Krylov subspaces. Journal of Scientific Computing, 90 (1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01727-1.

Cao:2022:ASV

[3039] Waixiang Cao and Qingsong Zou. Analysis of spectral volume methods for 1D linear scalar hyperbolic equations. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01715-5.

Li:2022:OSM

[3040] Dong Li and Chaoyu Quan. The operator-splitting method for Cahn– Hilliard is stable. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01740-4.

Su:2022:HEE

[3041] Haiyan Su and Guo-Dong Zhang. Highly efficient and energy stable schemes for the 2D/3D diffuse interface model of two-phase magnetohydrodynamics. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01741-3.

Tan:2022:TDA

[3042] Raynold Tan and Andrew Ooi. Two dimensional analysis and optimization of hybrid MDCD-TENO schemes. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01743-1.

Tang:2022:TFP

[3043] Min Tang, Lina Chang, and Yihong Wang. Tailored finite point method for diffusion equations with interfaces on distorted meshes. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01717-3.

Jiang:2022:HOL

[3044] Chaolong Jiang, Jin Cui, and Songhe Song. High-order linearly implicit structure-preserving exponential integrators for the nonlinear Schrödinger equation. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01739-x.

delTeso:2022:FDM

[3045] Félix del Teso and Erik Lindgren. A finite difference method for the variational p-Laplacian. Journal of Scientific Computing, 90(1):??, January 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01745-z.

Yang:2022:MMM

[3046] Zhipeng Yang, Ju Ming, and Xiaoming He. A multigrid multilevel Monte Carlo method for Stokes–Darcy model with random hydraulic conductivity and Beavers–Joseph condition. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01742-2.

Zhang:2022:IMD

[3047] Yu Zhang, Songsong Li, and Shan Du. Image multiplicative denoising using adaptive Euler’s elastica as the regularization. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01721-7.

Zhang:2022:FDF

[3048] Xiaodi Zhang and Xiaorong Wang. A fully divergence-free finite element scheme for stationary inductionless magnetohydrodynamic equations. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01708-4.

Yao:2022:SEM

[3049] Yonghong Yao, Olaniyi S. Iyiola, and Yekini Shehu. Subgradient extragradient method with double inertial steps for variational inequalities. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01751-1.

Feng:2022:EIM

[3050] Xiaobing Feng, Yan Luo, and Zhu Wang. An efficient iterative method for solving parameter-dependent and random convection–diffusion problems. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01737-z.

Avijit:2022:EDT

[3051] D. Avijit and S. Natesan. An efficient DWR-type a posteriori error bound of SDFEM for singularly perturbed convection–diffusion PDEs. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01749-9.

Bozorgmanesh:2022:TDC

[3052] Hassan Bozorgmanesh and Masoud Hajarian. Triangular decomposition of CP factors of a third-order tensor with application to solving nonlinear systems of equations. Journal of Scientific Computing, 90 (2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01758-8.

Herrmann:2022:CDR

[3053] Lukas Herrmann, Joost A. A. Opschoor, and Christoph Schwab. Constructive deep ReLU neural network approximation. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-021-01718-2.

Yang:2022:AHO

[3054] Jiang Yang, Zhaoming Yuan, and Zhi Zhou. Arbitrarily high-order maximum bound preserving schemes with cut-off postprocessing for Allen– Cahn equations. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01746-y.

Niu:2022:ECS

[3055] Cuixia Niu, Heping Ma, and Dong Liang. Energy-conserved splitting multidomain Legendre–Tau spectral method for two dimensional Maxwell’s equations. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-021-01744-0.

Laughton:2022:FBB

[3056] Edward Laughton, Vidhi Zala, and David Moxey. Fast barycentric-based evaluation over spectral/hp elements. Journal of Scientific Computing, 90(2):??, February 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01750-2.

Gassner:2022:SIE

[3057] Gregor J. Gassner, Magnus Svärd, and Florian J. Hindenlang. Stability issues of entropy-stable and/or split-form high-order schemes. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01720-8.

vanGestel:2022:CEC

[3058] R. A. M. van Gestel, M. J. H. Anthonissen, and W. L. IJzerman. Correction to: An Energy Conservative hp-method for Liouville’s Equation of Geometrical Optics. Journal of Scientific Computing, 90(3): ??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-021-01754-y. See [2913].

Xi:2022:NMU

[3059] Yingxia Xi and Xia Ji. A new method using C^{0} IPG for the biharmonic eigenvalue problem. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01762-6.

Loya:2022:HMD

[3060] Allen Alvarez Loya and Daniel Appelö. A Hermite method with a discontinuity sensor for Hamilton–Jacobi equations. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01766-2.

Zhao:2022:SDM

[3061] Lina Zhao, Dohyun Kim, and Eric Chung. Staggered DG method with small edges for Darcy flows in fractured porous media. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01760-8.

Fu:2022:HOD

[3062] Pei Fu, Thomas Frachon, and Sara Zahedi. High order discontinuous cut finite element methods for linear hyperbolic conservation laws with an interface. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01756-w.

Ryzhakov:2022:UAL

[3063] P. B. Ryzhakov, J. Marti, and N. Dialami. A unified arbitrary Lagrangian–Eulerian model for fluid-structure interaction problems involving flows in flexible channels. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-021-01748-w.

Reiss:2022:PTN

[3064] Julius Reiss. Pressure-tight and non-stiff volume penalization for compressible flows. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-021-01747-
x.

Xu:2022:ASA

[3065] Xin Xu and Xiaopeng Luo. Adaptive sparse approximations of scattered data. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01752-0.

Chen:2022:TGD

[3066] Shuangshuang Chen, Qiumei Huang, and Fei Xu. A two-grid decoupled algorithm for a multi-dimensional Darcy–Brinkman fracture model. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-021-01738-y.

Feng:2022:NVE

[3067] Fang Feng, Weimin Han, and Jianguo Huang. A nonconforming virtual element method for a fourth-order hemivariational inequality in Kirchhoff plate problem. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01759-1.

Li:2022:SMS

[3068] Xiaoli Li and Hongxing Rui. Superconvergence of MAC scheme for a coupled free flow-porous media system with heat transport on non-uniform grids. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01763-5.

Han:2022:OPG

[3069] Hao Han and Chengjian Zhang. One-parameter Galerkin finite element methods for neutral reaction-diffusion equations with piecewise continuous arguments. Journal of Scientific Computing, 90(3): ??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01769-z.

Sanchez:2022:EAU

[3070] Nestor Sánchez, Tonatiuh Sánchez-Vizuet, and Manuel Solano. Error analysis of an unfitted HDG method for a class of non-linear elliptic problems. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01767-1.

Wang:2022:OCL

[3071] Bin Wang and Yaolin Jiang. Optimal convergence and long-time conservation of exponential integration for Schrödinger equations in a normal or highly oscillatory regime. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01774-2.

Dolz:2022:RCE

[3072] Jürgen Dölz, Olena Palii, and Matthias Schlottbom. On robustly convergent and efficient iterative methods for anisotropic radiative transfer. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01757-9.

Hou:2022:IPP

[3073] Liangshao Hou, Xun Qian, and Jie Sun. An interior point parameterized central path following algorithm for linearly constrained convex programming. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01765-3.

Chouchoulis:2022:JFE

[3074] Jeremy Chouchoulis, Jochen Schütz, and Jonas Zeifang. Jacobian-free explicit multiderivative Runge–Kutta methods for hyperbolic conservation laws. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-021-01753-z.

Boscheri:2022:CCS

[3075] Walter Boscheri, Maurizio Tavelli, and Lorenzo Pareschi. On the construction of conservative semi-Lagrangian IMEX advection schemes for multiscale time dependent PDEs. Journal of Scientific Computing, 90(3):??, March 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01768-0.

Cao:2022:OSC

[3076] Wanrong Cao, Zhaopeng Hao, and Zhongqiang Zhang. Optimal strong convergence of finite element methods for one-dimensional stochastic elliptic equations with fractional noise. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01779-x.

Li:2022:CPP

[3077] Buyang Li, Weifeng Qiu, and Zongze Yang. A convergent post-processed discontinuous Galerkin method for incompressible flow with variable density. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01775-1.

Wang:2022:LCB

[3078] Cheng Wang, Wanli Wang, and Fuyu Zhao. A local curvature based adaptive particle level set method. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01772-4.

Mossier:2022:ADG

[3079] Pascal Mossier, Andrea Beck, and Claus-Dieter Munz. A p-adaptive discontinuous Galerkin method with hp-shock capturing. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01770-6.

Peng:2022:RBM

[3080] Zhichao Peng, Yanlai Chen, and Fengyan Li. A reduced basis method for radiative transfer equation. Journal of Scientific Computing, 91 (1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01782-2.

Bevilacqua:2022:OIC

[3081] R. Bevilacqua, G. M. Del Corso, and L. Gemignani. Orthogonal iterations on companion-like pencils. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01777-z.

Veilleux:2022:SSD

[3082] Adèle Veilleux, Guillaume Puigt, and Guillaume Daviller. Stable spectral difference approach using Raviart–Thomas elements for 3D computations on tetrahedral grids. Journal of Scientific Computing, 91 (1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01790-2.

Bermudez:2022:NSA

[3083] A. Bermúdez, B. López-Rodríguez, and P. Venegas. Numerical solution of an axisymmetric eddy current model with current and voltage excitations. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01780-4.

Ostermann:2022:FDL

[3084] Alexander Ostermann and Fangyan Yao. A fully discrete low-regularity integrator for the nonlinear Schrödinger equation. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01786-y.

Kivva:2022:ESF

[3085] Sergii Kivva. Entropy stable flux correction for scalar hyperbolic conservation laws. Journal of Scientific Computing, 91(1):??, April 2022. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-022-01792-
0.

Xu:2022:LEE

[3086] Yuan Xu, Di Zhao, and Qiang Zhang. Local error estimates for Runge–Kutta discontinuous Galerkin methods with upwind-biased numerical fluxes for a linear hyperbolic equation in one-dimension with discontinuous initial data. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01793-z.

Diao:2022:PLS

[3087] Xuhao Diao, Jun Hu, and Suna Ma. Preconditioned Legendre spectral Galerkin methods for the non-separable elliptic equation. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-021-01755-x.

Zhang:2022:CET

[3088] Hong Zhang, Jingye Yan, and Songhe Song. Correction to: Explicit Third-Order Unconditionally Structure-Preserving Schemes for Conservative Allen–Cahn Equations. Journal of Scientific Computing, 91(1): ??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01764-4. See [2986].

Luo:2022:OEE

[3089] Hao Luo and Xiaoping Xie. Optimal error estimates of a time-spectral method for fractional diffusion problems with low regularity data. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01791-1.

Liu:2022:IHS

[3090] Xiao Liu, Jianlin Xia, and Xiaofeng Ou. Interconnected hierarchical structures for fast direct elliptic solution. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01761-7.

Gosea:2022:DDM

[3091] Ion Victor Gosea and Serkan Gugercin. Data-driven modeling of linear dynamical systems with quadratic output in the AAA framework. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01771-5.

Manohar:2022:LPE

[3092] Ram Manohar and Rajen Kumar Sinha. Local a posteriori error analysis of finite element method for parabolic boundary control problems. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01788-w.

Jiang:2022:LCR

[3093] Jiahua Jiang, Fatoumata Sanogo, and Carmeliza Navasca. Low-CP-rank tensor completion via practical regularization. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01789-9.

Liu:2022:NSM

[3094] Yang Liu, Mark Sussman, and Kourosh Shoele. A novel supermesh method for computing solutions to the multi-material Stefan problem with complex deforming interfaces and microstructure. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01783-1.

Li:2022:PPE

[3095] Yanjun Li, Hai Bi, and Yidu Yang. The a priori and a posteriori error estimates of DG method for the Steklov eigenvalue problem in inverse scattering. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01787-x.

Kuzmin:2022:BPF

[3096] Dmitri Kuzmin, Manuel Quezada de Luna, and Johanna Grüll. Boundpreserving flux limiting for high-order explicit Runge–Kutta time discretizations of hyperbolic conservation laws. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01784-0.

Jia:2022:SSG

[3097] Lueling Jia, Huiyuan Li, and Zhimin Zhang. Sparse spectral-Galerkin method on an arbitrary tetrahedron using generalized Koornwinder polynomials. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01778-y.

Adak:2022:CAV

[3098] D. Adak, D. Mora, and S. Natarajan. Convergence analysis of virtual element method for nonlinear nonlocal dynamic plate equation. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01794-y.

Lin:2022:RAE

[3099] Matthew M. Lin and Moody T. Chu. Rank-1 approximation for entangled multipartite real systems. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01805-y.

Antoine:2022:GFA

[3100] X. Antoine and E. Lorin. Generalized fractional algebraic linear system solvers. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01785-z.

Law:2022:HOF

[3101] Yann-Meing Law and Jean-Christophe Nave. High-order FDTD schemes for Maxwell’s interface problems with discontinuous coefficients and complex interfaces based on the correction function method. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01797-9.

Li:2022:NDF

[3102] Dongfang Li, Mianfu She, and Xiaoqiang Yan. A novel discrete fractional Grönwall-type inequality and its application in pointwisein-time error estimates. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01803-0.

Kuhn:2022:IEG

[3103] Martin J. Kühn, Carola Kruse, and Ulrich Rüde. Implicitly extrapolated geometric multigrid on disk-like domains for the gyrokinetic Poisson equation from fusion plasma applications. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01802-1.

Qian:2022:MRL

[3104] Elizabeth Qian, Jemima M. Tabeart, and Akil Narayan. Model reduction of linear dynamical systems via balancing for Bayesian inference. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01798-8.

Paskal:2022:EJM

[3105] Nicholas Paskal and Maria Cameron. An efficient jet marcher for computing the quasipotential for 2D SDEs. Journal of Scientific Computing, 91(1):??, April 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01807-w.

Palitta:2022:EMS

[3106] Davide Palitta and Sanda Lefteriu. An efficient, memory-saving approach for the Loewner framework. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01800-3.

Fjordholm:2022:CRM

[3107] Ulrik Skre Fjordholm and Kjetil Olsen Lye. Convergence rates of monotone schemes for conservation laws for data with unbounded total variation. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01806-x.

Ho:2022:NGS

[3108] Myong-Song Ho, Ji-Song Pak, and Ju-Song Kim. Neighbor-gradient single-pass method for solving anisotropic eikonal equation. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01773-3.

Sun:2022:NAF

[3109] Jing Sun, Weihua Deng, and Daxin Nie. Numerical approximations for the fractional Fokker–Planck equation with two-scale diffusion. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01812-z.

Faghih:2022:NLS

[3110] A. Faghih and P. Mokhtary. Non-linear system of multi-order fractional differential equations: Theoretical analysis and a robust fractional Galerkin implementation. Journal of Scientific Computing, 91 (2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01814-x.

Bentbib:2022:ERA

[3111] A. H. Bentbib, M. El Ghomari, and K. Jbilou. An extended-rational Arnoldi method for large matrix exponential evaluations. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01808-9.

Hong:2022:MTS

[3112] Zheng Hong, Zhengyin Ye, and Kun Ye. A modified TENO scheme with improved efficiency. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01809-8.

Natale:2022:MFE

[3113] Andrea Natale and Gabriele Todeschi. A mixed finite element discretization of dynamical optimal transport. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01821-y.

Liu:2022:MSS

[3114] Yiyang Liu, Zaiwen Wen, and Wotao Yin. A multiscale semi-smooth Newton method for optimal transport. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01813-y.

Wang:2022:CDD

[3115] Shuyi Wang, Zixu Zhou, and Dongbin Xiu. Construction of discontinuity detectors using convolutional neural networks. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01804-z.

Marchetti:2022:ERB

[3116] Francesco Marchetti and Emma Perracchione. Efficient reduced basis algorithm (ERBA) for kernel-based approximation. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01818-7.

Khandelwal:2022:PPE

[3117] Rohit Khandelwal and Kamana Porwal. Pointwise a posteriori error analysis of a finite element method for the Signorini problem. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01811-0.

Huang:2022:SAR

[3118] Chaobao Huang and Martin Stynes. A sharp α-robust L^{∞}(H^{1}) error bound for a time-fractional Allen–Cahn problem discretised by the Alikhanov L2 − 1_{σ} scheme and a standard FEM. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01810-1.

Ming:2022:NHM

[3119] Pingbing Ming and Siqi Song. A Nitsche hybrid multiscale method with non-matching grids. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01817-8.

Zhou:2022:EEF

[3120] Kaiye Zhou and Wei Gong. Error estimates for finite element approximation of Dirichlet boundary control for Stokes equations in L^{2}(Γ). Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01831-w.

Pinto:2022:VFS

[3121] Martin Campos Pinto, Katharina Kormann, and Eric Sonnendrücker. Variational framework for structure-preserving electromagnetic particlein-cell methods. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-022-01781-
3.

Kang:2022:ESB

[3122] Yuanyuan Kang and Hong lin Liao. Energy stability of BDF methods up to fifth-order for the molecular beam epitaxial model without slope selection. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01830-x.

Aretaki:2022:EHO

[3123] Aikaterini Aretaki, Efthymios N. Karatzas, and Georgios Katsouleas. Equal higher order analysis of an unfitted discontinuous Galerkin method for Stokes flow systems. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01823-w.

Wang:2022:CAF

[3124] Jiangxing Wang, Kejia Pan, and Xiaofeng Yang. Convergence analysis of the fully discrete hybridizable discontinuous Galerkin method for the Allen–Cahn equation based on the invariant energy quadratization approach. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01822-x.

Gnanasekaran:2022:HOF

[3125] Abeynaya Gnanasekaran and Eric Darve. Hierarchical orthogonal factorization: Sparse least squares problems. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01824-9.

Arbogast:2022:RWR

[3126] Todd Arbogast, Chieh-Sen Huang, and Ming-Hsien Kuo. RBF WENO reconstructions with adaptive order and applications to conservation laws. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01827-6.

Wang:2022:EBS

[3127] Siyang Wang, Daniel Appelö, and Gunilla Kreiss. An energy-based summation-by-parts finite difference method for the wave equation in second order form. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01829-4.

Li:2022:FEA

[3128] Yukun Li, Shuonan Wu, and Yulong Xing. Finite element approximations of a class of nonlinear stochastic wave equations with multiplicative noise. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01816-9.

Jia:2022:NAF

[3129] Jinhong Jia, Hong Wang, and Xiangcheng Zheng. Numerical analysis of a fast finite element method for a hidden-memory variableorder time-fractional diffusion equation. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01820-z.

Aceto:2022:ECT

[3130] Lidia Aceto and Paolo Novati. Exponentially convergent trapezoidal rules to approximate fractional powers of operators. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01837-4.

Zheng:2022:DAO

[3131] Xiangcheng Zheng and Hong Wang. Discretization and analysis of an optimal control of a variable-order time-fractional diffusion equation with pointwise constraints. Journal of Scientific Computing, 91 (2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01795-x.

Zhou:2022:CAS

[3132] Guanyu Zhou. The convergence analysis of semi- and fully-discrete projection-decoupling schemes for the generalized Newtonian models. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01828-5.

Ma:2022:TOA

[3133] Xiaojun Ma, Hongwei Liu, and Xiaoyin Li. Two optimization approaches for solving split variational inclusion problems with applications. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01832-9.

DAmore:2022:SST

[3134] Luisa D’Amore, Emil Constantinescu, and Luisa Carracciuolo. A scalable space-time domain decomposition approach for solving large scale nonlinear regularized inverse ill posed problems in 4D variational data assimilation. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01826-7.

Huang:2022:CRN

[3135] Kevin Huang, Junyu Zhang, and Shuzhong Zhang. Cubic regularized Newton method for the saddle point models: a global and local convergence analysis. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01819-6.

Zeng:2022:MEA

[3136] Jinshan Zeng, Wotao Yin, and Ding-Xuan Zhou. Moreau envelope augmented Lagrangian method for nonconvex optimization with linear constraints. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01815-w.

Yin:2022:CHI

[3137] Jianyuan Yin, Zhen Huang, and Lei Zhang. Constrained high-index saddle dynamics for the solution landscape with equality constraints. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01838-3.

Germain:2022:DTD

[3138] Maximilien Germain, Mathieu Laurière, and Xavier Warin. DeepSets and their derivative networks for solving symmetric PDEs. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01796-w.

Ri:2022:NCU

[3139] Kuk Hwan Ri and Kuk Hyon Sonu. A numerical construction of the universal feedback control in problems of nonlinear controls under disturbance. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01799-7.

Lindblad:2022:MAM

[3140] Daniel Lindblad, Christian Frey, and Niklas Andersson. Minimizing aliasing in multiple frequency harmonic balance computations. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01776-0.

Ji:2022:IRT

[3141] Haifeng Ji. An immersed Raviart–Thomas mixed finite element method for elliptic interface problems on unfitted meshes. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01839-2.

Shi:2022:NPA

[3142] Jia Shi, Ruipeng Li, and Maarten V. de Hoop. A non-perturbative approach to computing seismic normal modes in rotating planets. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01836-5.

Safari:2022:CAL

[3143] Z. Safari, G. B. Loghmani, and M. Ahmadinia. Convergence analysis of a LDG method for time-space tempered fractional diffusion equations with weakly singular solutions. Journal of Scientific Computing, 91(2):??, May 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01835-6.

Ciaramella:2022:SCS

[3144] Gabriele Ciaramella and Tommaso Vanzan. Spectral coarse spaces for the substructured parallel Schwarz method. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01840-9.

Sun:2022:CPT

[3145] Yafei Sun, Shu-Lin Wu, and Yingxiang Xu. Correction to: A Parallelin-Time Implementation of the Numerov Method for Wave Equations. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01847-2. See [2998].

Lukacova-Medvidova:2022:EEG

[3146] Mária Lukácová-Medvid’ová, Bangwei She, and Yuhuan Yuan. Error estimates of the Godunov method for the multidimensional compressible Euler system. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01843-6.

Li:2022:NUB

[3147] Qin Li, Dong Sun, and Xiao Huang. Nonlinear upwind-biased freestream-preserving schemes for compressible Euler equations. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01833-8.

Xu:2022:FVM

[3148] Xiao Xu and Guoxi Ni. A finite volume method for the 3D Lagrangian ideal compressible magnetohydrodynamics. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01851-6.

Le:2022:GDM

[3149] Thuy T. Le and Loc H. Nguyen. The gradient descent method for the convexification to solve boundary value problems of quasi-linear PDEs and a coefficient inverse problem. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01846-3.

Ding:2022:IPD

[3150] Mingcai Ding, Xiaoliang Song, and Bo Yu. An inexact proximal DC algorithm with sieving strategy for rank constrained least squares semidefinite programming. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01845-4.

Brenner:2022:MFE

[3151] Susanne C. Brenner, José C. Garay, and Li-Yeng Sung. Multiscale finite element methods for an elliptic optimal control problem with rough coefficients. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01834-7.

Lin:2022:CID

[3152] Bo Lin, Qianxiao Li, and Weiqing Ren. Computing the invariant distribution of randomly perturbed dynamical systems using deep learning. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01844-5.

Liu:2022:EII

[3153] Ching-Sung Liu. Exact and inexact iterative methods for finding the largest eigenpair of a weakly irreducible nonnegative tensor. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01852-5.

Xiao:2022:SIR

[3154] Yao Xiao, Jan Glaubitz, and Guohui Song. Sequential image recovery from noisy and under-sampled Fourier data. Journal of Scientific Computing, 91(3):??, June 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01850-7.

Chen:2022:ILM

[3155] Xianjin Chen, Zhaoxiang Li, and Jianxin Zhou. An improved local-min-orthogonal method for finding multiple solutions to nonlinear elliptic PDEs. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01842-7.

Erath:2022:MCE

[3156] Christoph Erath, Lorenzo Mascotto, and Alexander Rieder. Mortar coupling of hp-discontinuous Galerkin and boundary element methods for the Helmholtz equation. Journal of Scientific Computing, 92(1): ??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01849-0.

Fatone:2022:DMM

[3157] L. Fatone, D. Funaro, and G. Manzini. A decision-making machine learning approach in Hermite spectral approximations of partial differential equations. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01853-4.

Fu:2022:EFE

[3158] Taibai Fu, Changfa Du, and Yufeng Xu. An effective finite element method with shifted fractional powers bases for fractional boundary value problems. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01854-3.

Balcells-Quintana:2022:CIG

[3159] Oscar Balcells-Quintana, David Codony, and Sonia Fernández-Méndez. C0-IPM with generalised periodicity and application to flexoelectricitybased 2D metamaterials. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01848-1.

Grant:2022:PRK

[3160] Zachary J. Grant. Perturbed Runge–Kutta methods for mixed precision applications. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01801-2.

Ding:2022:MFG

[3161] Lisang Ding, Wuchen Li, and Wotao Yin. A mean field game inverse problem. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01825-8.

Zheng:2022:TCF

[3162] Yu-Bang Zheng, Ting-Zhu Huang, and Qibin Zhao. Tensor completion via fully-connected tensor network decomposition with regularized factors. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01841-8.

Wang:2022:CDG

[3163] Yue Wang, Fuzheng Gao, and Jintao Cui. A conforming discontinuous Galerkin finite element method for linear elasticity interface problems. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01857-0.

Shokeen:2022:PHM

[3164] Ravina Shokeen, Ajit Patel, and Amiya K. Pani. Primal hybrid method for quasilinear parabolic problems. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01858-z.

Keram:2022:AHI

[3165] Aytura Keram and Pengzhan Huang. The arrow-Hurwicz iterative finite element method for the stationary thermally coupled incompressible magnetohydrodynamics flow. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01867-y.

Beltran:2022:ADD

[3166] Víctor Beltrán, Soledad Le Clainche, and José M. Vega. An adaptive data-driven reduced order model based on higher order dynamic mode decomposition. Journal of Scientific Computing, 92(1): ??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01855-2.

Ling:2022:PPS

[3167] Min Ling, Weimin Han, and Shengda Zeng. A pressure projection stabilized mixed finite element method for a Stokes hemivariational inequality. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01871-2.

Zhong:2022:ESG

[3168] Xinghui Zhong and Chi-Wang Shu. Entropy stable Galerkin methods with suitable quadrature rules for hyperbolic systems with random inputs. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01866-z.

Chun:2022:DCP

[3169] Sehun Chun and Taejin Oh. Divergence/Connection preservation scheme in the curvilinear domain with a small geometric approximation error. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01865-0.

Kwak:2022:MMC

[3170] Soobin Kwak, Hyun Geun Lee, and Junseok Kim. Motion by mean curvature with constraints using a modified Allen–Cahn equation. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01862-3.

Chen:2022:SPV

[3171] Xi Chen and Yuwen Li. Superconvergent pseudostress-velocity finite element methods for the Oseen equations. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01856-1.

She:2022:CUS

[3172] Zi-Hang She. A class of unconditioned stable 4-point WSGD schemes and fast iteration methods for space fractional diffusion equations. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01860-5.

Longo:2022:AQM

[3173] Marcello Longo. Adaptive quasi-Monte Carlo finite element methods for parametric elliptic PDEs. Journal of Scientific Computing, 92(1): ??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01859-y.

Wang:2022:UES

[3174] Chenxi Wang, Yichen Guo, and Zhen Zhang. Unconditionally energy stable and bound-preserving schemes for phase-field surfactant model with moving contact lines. Journal of Scientific Computing, 92(1): ??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01863-2.

Yang:2022:AGM

[3175] Yanfang Yang, Shubin Fu, and Eric T. Chung. An adaptive generalized multiscale finite element method based two-grid preconditioner for large scale high-contrast linear elasticity problems. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01869-w.

Xiao:2022:GPG

[3176] Guiyun Xiao and Zheng-Jian Bai. A geometric proximal gradient method for sparse least squares regression with probabilistic simplex constraint. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01873-0.

Xu:2022:DSP

[3177] Xingjian Xu and Minghua Chen. Discovery of subdiffusion problem with noisy data via deep learning. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01879-8.

Zhang:2022:SBE

[3178] Baiju Zhang, Hengguang Li, and Zhimin Zhang. Solving biharmonic eigenvalue problem with Navier boundary condition via Poisson solvers on non-convex domains. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01878-9.

Zhou:2022:AFD

[3179] Qin Zhou and Minfu Feng. Analysis of a full discretization for a Fractional/Normal diffusion equation with rough Dirichlet boundary data. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01875-y.

Gu:2022:EEP

[3180] Xuelong Gu, Chaolong Jiang, and Wenjun Cai. Efficient energypreserving exponential integrators for multi-component Hamiltonian systems. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01874-z.

Fu:2022:API

[3181] Jinxue Fu, Weiming Li, and Yanli Wang. An asymptotic-preserving IMEX method for nonlinear radiative transfer equation. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01870-3.

Botti:2022:HMI

[3182] Lorenzo Botti and Francesco Carlo Massa. HHO methods for the incompressible Navier–Stokes and the incompressible Euler equations. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01864-1.

Li:2022:STO

[3183] Siye Li, Zhensheng Sun, and Ding Yao. A spatio-temporal optimal, hybird compact-WENO scheme with minimized dispersion and criticaladaptive dissipation for solving compressible flows. Journal of Scientific Computing, 92(1):??, July 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01884-x.

Schneider:2022:EGI

[3184] Kleiton A. Schneider, José M. Gallardo, and Cipriano Escalante. Efficient GPU implementation of multidimensional incomplete Riemann solvers for hyperbolic nonconservative systems: Applications to shallow water systems with topography and dry areas. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01880-1.

Chen:2022:PPE

[3185] Wenbin Chen, Jianyu Jing, and Xiaoming Wang. A positivity preserving, energy stable finite difference scheme for the Flory–Huggins– Cahn–Hilliard–Navier–Stokes system. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01872-1.

Wang:2022:NMD

[3186] Tian jun Wang. A new multi-domain spectral method for Korteweg– de Vries equation on the whole line. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01887-8.

Zhong:2022:SAF

[3187] Qiuxiang Zhong, Ryan Wen Liu, and Yuping Duan. Spatially adapted first and second order regularization for image reconstruction: From an image surface perspective. Journal of Scientific Computing, 92 (2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01886-9.

Galaris:2022:NBA

[3188] Evangelos Galaris, Gianluca Fabiani, and Constantinos Siettos. Numerical bifurcation analysis of PDEs from lattice Boltzmann model simulations: a parsimonious machine learning approach. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01883-y.

Wang:2022:IAT

[3189] Sifan Wang, Hanwen Wang, and Paris Perdikaris. Improved architectures and training algorithms for deep operator networks. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01881-0.

Ascione:2022:SBA

[3190] Giacomo Ascione and Salvatore Cuomo. A sojourn-based approach to semi-Markov reinforcement learning. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01876-x.

Futai:2022:MPT

[3191] Kouta Futai, Niklas Kolbe, and Tasuku Suzuki. A mass-preserving twostep Lagrange–Galerkin scheme for convection–diffusion problems. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01885-w.

Mulder:2022:MCM

[3192] Wim A. Mulder. More continuous mass-lumped triangular finite elements. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01890-z.

Rodgers:2022:AIN

[3193] Abram Rodgers, Alec Dektor, and Daniele Venturi. Adaptive integration of nonlinear evolution equations on tensor manifolds. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01868-x.

Liao:2022:CAD

[3194] Wei-Hung Liao and Mei-Heng Yueh. A constructive algorithm for disk conformal parameterizations. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01898-5.

Huang:2022:PST

[3195] Xin Huang, Dongfang Li, Hai-Wei Sun, and Fan Zhang. Preconditioners with symmetrized techniques for space fractional Cahn–Hilliard equations. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01900-0.

Cheng:2022:ISG

[3196] Wanyou Cheng, Xiao Wang, and Xiaojun Chen. An interior stochastic gradient method for a class of non–Lipschitz optimization problems. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01897-6.

Chen:2022:MFE

[3197] Yuxiang Chen and Zhihao Ge. Multiphysics finite element method for quasi-static thermo-poroelasticity. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01877-w.

Zhao:2022:AMD

[3198] Lina Zhao. Analysis of a mixed DG method for stress–velocity formulation of the Stokes equations. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01895-8.

Banks:2022:CDG

[3199] J. W. Banks, B. Brett Buckner, and T. Hagstrom. Continuous/ discontinuous Galerkin difference discretizations of high-order differential operators. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01891-y.

Niu:2022:DDS

[3200] Yi-Shuai Niu and Roland Glowinski. Discrete dynamical system approaches for Boolean polynomial optimization. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01882-z.

Sorokina:2022:IGF

[3201] Tatyana Sorokina and Shangyou Zhang. An interpolated Galerkin finite element method for the Poisson equation. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01903-x.

Wu:2022:IIF

[3202] Tingting Wu and Yuesheng Xu. Inverting incomplete Fourier transforms by a sparse regularization model and applications in seismic wavefield modeling. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01906-8.

Chen:2022:RNM

[3203] Qipin Chen and Wenrui Hao. Randomized Newton’s method for solving differential equations based on the neural network discretization. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01905-9.

Marchildon:2022:UPI

[3204] André L. Marchildon and David W. Zingg. Unisolvency for polynomial interpolation in simplices with symmetrical nodal distributions. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01904-w.

Baier-Reinio:2022:APR

[3205] Aaron Baier-Reinio, Sander Rhebergen, and Garth N. Wells. Analysis of pressure-robust embedded–hybridized discontinuous Galerkin methods for the Stokes problem under minimal regularity. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01889-6.

Liao:2022:MRE

[3206] Hong lin Liao, Bingquan Ji, Lin Wang, and Zhimin Zhang. Meshrobustness of an energy stable BDF2 scheme with variable steps for the Cahn–Hilliard model. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01861-4.

Cai:2022:UAN

[3207] Yongyong Cai and Xuanxuan Zhou. Uniformly accurate nested Picard iterative integrators for the Klein–Gordon equation in the nonrelativistic regime. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01909-5.

Banas:2022:NAS

[3208] L’ubomír Baňas, Herbert Dawid, Tsiry Avisoa Randrianasolo, Johannes Storn, and Xingang Wen. Numerical approximation of a system of Hamilton–Jacobi–Bellman equations arising in innovation dynamics. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01892-x.

Deugoue:2022:NAD

[3209] G. Deugoue, J. K. Djoko, V. S. Konlack, and M. Mbehou. Numerical analysis of a Darcy–Forchheimer model coupled with the heat equation. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01920-w.

Calder:2022:BEP

[3210] Jeff Calder, Sangmin Park, and Dejan Slepcev. Boundary estimation from point clouds: Algorithms, guarantees and applications. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01894-9.

Dong:2022:RIP

[3211] Zhaonan Dong and Emmanuil H. Georgoulis. Robust interior penalty discontinuous Galerkin methods. Journal of Scientific Computing, 92 (2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01916-6.

El-Amrani:2022:BBG

[3212] Mofdi El-Amrani, Abdellah El-Kacimi, Bassou Khouya, and Mohammed Seaid. Bernstein–Bézier Galerkin-characteristics finite element method for convection–diffusion problems. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01888-7.

Yu:2022:NSS

[3213] Xiaoxuan Yu, Yan Xu, and Qiang Du. Numerical simulation of singularity propagation modeled by linear convection equations with spatially heterogeneous nonlocal interactions. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01915-7.

Birken:2022:CPI

[3214] Philipp Birken and Viktor Linders. Conservation properties of iterative methods for implicit discretizations of conservation laws. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01923-7.

Ji:2022:CNH

[3215] Zhe Ji, Tian Liang, and Lin Fu. A class of new high-order finite-volume TENO schemes for hyperbolic conservation laws with unstructured meshes. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01925-5.

Liljegren-Sailer:2022:SBM

[3216] Björn Liljegren-Sailer and Nicole Marheineke. On snapshot-based model reduction under compatibility conditions for a nonlinear flow problem on networks. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01901-z.

Bevilacqua:2022:BPD

[3217] Tommaso Bevilacqua and Simone Scacchi. BDDC preconditioners for divergence free virtual element discretizations of the Stokes equations. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01929-1.

Cheng:2022:RAM

[3218] Fangxiong Cheng, Hui Xu, and Xinlong Feng. Retracted Article: Model order reduction method based on machine learning for parameterized time-dependent partial differential equations. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-022-01899-4. See retraction notice [3333].

Di:2022:SEE

[3219] Yana Di, Yifan Wei, Jiwei Zhang, and Chengchao Zhao. Sharp error estimate of an implicit BDF2 scheme with variable time steps for the phase field crystal model. Journal of Scientific Computing, 92 (2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01919-3.

Ju:2022:SES

[3220] Lili Ju, Xiao Li, and Zhonghua Qiao. Stabilized Exponential–SAV schemes preserving energy dissipation law and maximum bound principle for the Allen–Cahn type equations. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01921-9.

Meng:2022:BNE

[3221] Xiangyun Meng and Martin Stynes. Balanced-norm and energy-norm error analyses for a backward Euler/FEM solving a singularly perturbed parabolic reaction–diffusion problem. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01931-7.

Zhao:2022:SNV

[3222] Jikun Zhao, Tianle Wang, and Bei Zhang. The stabilized nonconforming virtual element method for linear elasticity problem. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01927-3.

Cao:2022:FGB

[3223] Yangyang Cao, Alexander Kurganov, Yongle Liu, and Ruixiao Xin. Flux globalization based well-balanced path-conservative central-upwind schemes for shallow water models. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01912-w.

Zhou:2022:TTS

[3224] Han Zhou and Wenyi Tian. Two time-stepping schemes for sub-diffusion equations with singular source terms. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01914-8.

Badia:2022:CHH

[3225] Santiago Badia, Jérôme Droniou, and Liam Yemm. Conditioning of a hybrid high-order scheme on meshes with small faces. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01913-9.

Chen:2022:CNS

[3226] Qingshan Chen, Lili Ju, and Roger Temam. Conservative numerical schemes with optimal dispersive wave relations: Part II. Numerical evaluations. Journal of Scientific Computing, 92(2):??, August 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01908-6.

Li:2022:LPF

[3227] Qingtao Li and Guangzhi Du. Local and parallel finite element methods based on two-grid discretizations for a transient coupled Navier– Stokes/Darcy model. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01946-0.

Dao:2022:HOR

[3228] Tuan Anh Dao and Murtazo Nazarov. A high-order residual-based viscosity finite element method for the ideal MHD equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01918-4.

Guo:2022:TGS

[3229] Yu ling Guo and Zhong qing Wang. A two-grid spectral deferred correction method for the multi-order fractional differential equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01942-4.

Lu:2022:PEP

[3230] Zhongjie Lu and Yan Xu. A parallel eigensolver for photonic crystals discretized by edge finite elements. Journal of Scientific Computing, 92 (3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01938-0.

Ren:2022:RQQ

[3231] Huan Ren, Ru-Ru Ma, Qiaohua Liu, and Zheng-Jian Bai. Randomized quaternion QLP decomposition for low-rank approximation. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01917-5.

Eriksson:2022:WVS

[3232] Gustav Eriksson and Ken Mattsson. Weak versus strong wall boundary conditions for the incompressible Navier–Stokes equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01941-5.

Sun:2022:ISN

[3233] Hongpeng Sun. An investigation on semismooth Newton based augmented Lagrangian method for image restoration. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01907-7.

Li:2022:NTI

[3234] Ben-Zheng Li, Xi-Le Zhao, Teng-Yu Ji, Xiong-Jun Zhang, and Ting-Zhu Huang. Nonlinear transform induced tensor nuclear norm for tensor completion. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01937-1.

Hegarty:2022:NMS

[3235] A. F. Hegarty and E. O’Riordan. A numerical method for singularly perturbed convection-diffusion problems posed on smooth domains. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01896-7.

Xu:2022:MLD

[3236] Fei Xu, Qiumei Huang, Haishen Dai, and Hongkun Ma. Multilevel local defect-correction method for nonsymmetric eigenvalue problems. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01926-4.

Fu:2022:HOV

[3237] Guosheng Fu. A high-order velocity-based discontinuous Galerkin scheme for the shallow water equations: Local conservation, entropy stability, well-balanced property, and positivity preservation. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01902-y.

Montoya:2022:UAF

[3238] Tristan Montoya and David W. Zingg. A unifying algebraic framework for discontinuous Galerkin and flux reconstruction methods based on the summation-by-parts property. Journal of Scientific Computing, 92 (3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01935-3.

Cuomo:2022:SML

[3239] Salvatore Cuomo, Vincenzo Schiano Di Cola, Fabio Giampaolo, Gianluigi Rozza, Maziar Raissi, and Francesco Piccialli. Scientific machine learning through physics-informed neural networks: Where we are and what’s next. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01939-z.

Duan:2022:FOS

[3240] Yuping Duan, Qiuxiang Zhong, Xue-Cheng Tai, and Roland Glowinski. A fast operator-splitting method for Beltrami color image denoising. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01910-y.

Wang:2022:HIM

[3241] Li Wang and Ming Yan. Hessian informed mirror descent. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01933-5.

Gong:2022:OCB

[3242] Wei Gong, Buyang Li, and Huanhuan Yang. Optimal control in a bounded domain for wave propagating in the whole space: Coupling through boundary integral equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01953-1.

Barnafi:2022:FEM

[3243] N. A. Barnafi, B. Gómez-Vargas, W. J. Lourenço, R. F. Reis, B. M. Rocha, M. Lobosco, R. Ruiz-Baier, and R. Weber dos Santos. Finite element methods for large-strain poroelasticity/chemotaxis models simulating the formation of myocardial oedema. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01944-2.

Che:2022:EAC

[3244] Maolin Che and Yimin Wei. An efficient algorithm for computing the approximate t-URV and its applications. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01956-y.

Yang:2022:SBE

[3245] Minghan Yang, Dong Xu, Zaiwen Wen, Mengyun Chen, and Pengxiang Xu. Sketch-based empirical natural gradient methods for deep learning. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01911-x.

Guo:2022:RMC

[3246] Jun Guo and Minfu Feng. A robust and mass conservative virtual element method for linear three-field poroelasticity. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01960-2.

Desiderio:2022:CBC

[3247] Luca Desiderio, Silvia Falletta, Matteo Ferrari, and Letizia Scuderi. CVEM-BEM coupling with decoupled orders for 2D exterior Poisson problems. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01951-3.

Fu:2022:EPM

[3248] Zhaohui Fu, Tao Tang, and Jiang Yang. Energy plus maximum bound preserving Runge–Kutta methods for the Allen–Cahn equation. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01940-6.

Tan:2022:LMN

[3249] Tan Tan, Wei-Ping Bu, and Ai-Guo Xiao. L1 method on nonuniform meshes for linear time-fractional diffusion equations with constant time delay. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01948-y.

Chen:2022:ECE

[3250] Yanping Chen, Lina Wang, and Lijun Yi. Exponential convergence of hp-discontinuous Galerkin method for nonlinear Caputo fractional differential equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01947-z.

Berrone:2022:VPI

[3251] Stefano Berrone, Claudio Canuto, and Moreno Pintore. Variational physics informed neural networks: the role of quadratures and test functions. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01950-4.

Rong:2022:MVR

[3252] Y. Rong, J. A. Fiordilino, F. Shi, and Y. Cao. A modular Voigt regularization of the Crank–Nicolson finite element method for the Navier– Stokes equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01945-1.

deLuna:2022:MPP

[3253] Manuel Quezada de Luna and David I. Ketcheson. Maximum principle preserving space and time flux limiting for diagonally implicit Runge– Kutta discretizations of scalar convection–diffusion equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01922-8.

Gatica:2022:PBM

[3254] Gabriel N. Gatica, Cristian Inzunza, and Filánder A. Sequeira. A pseudostress-based mixed-primal finite element method for stressassisted diffusion problems in Banach spaces. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-022-01959-9.

Lu:2022:SDF

[3255] Xiaoli Lu, Pengzhan Huang, Xinlong Feng, and Yinnian He. A stabilized difference finite element method for the 3D steady incompressible Navier–Stokes equations. Journal of Scientific Computing, 92(3): ??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01928-2.

Gao:2022:ASP

[3256] Guangwei Gao and Shuonan Wu. Auxiliary space preconditioners for a C^{0} finite element approximation of Hamilton–Jacobi–Bellman equations with Cordes coefficients. Journal of Scientific Computing, 92 (3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01957-x.

Tang:2022:PTO

[3257] Yuchao Tang, Meng Wen, and Tieyong Zeng. Preconditioned threeoperator splitting algorithm with applications to image restoration. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01958-w.

Ying:2022:PRN

[3258] Lexing Ying. Pole recovery from noisy data on imaginary axis. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01963-z.

Xie:2022:AGD

[3259] Xihui Xie. On adaptive grad–div parameter selection. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01964-y.

Liu:2022:OFD

[3260] Yong Liu, Jianfang Lu, Qi Tao, and Yinhua Xia. An oscillation-free discontinuous Galerkin method for shallow water equations. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01893-w.

Wang:2022:RGW

[3261] Bingkun Wang, Lei Liu, Yuchong Li, and Mohammad Khishe. Robust grey wolf optimizer for multimodal optimizations: a cross-dimensional coordination approach. Journal of Scientific Computing, 92(3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01955-z.

Qian:2022:PPW

[3262] Xu Qian, Jian Dong, and Songhe Song. Positivity-preserving and well-balanced adaptive surface reconstruction schemes for shallow water equations with wet–dry fronts. Journal of Scientific Computing, 92 (3):??, September 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01943-3.

Huang:2022:LMR

[3263] Wen Huang and Kyle A. Gallivan. A limited-memory Riemannian symmetric rank-one trust-region method with a restart strategy. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01962-0.

Sun:2022:PGM

[3264] Baochen Sun and Huibin Chang. Proximal gradient methods for general smooth graph total variation model in unsupervised learning. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01954-0.

Ren:2022:HOA

[3265] Yupeng Ren, Yulong Xing, Dean Wang, and Jianxian Qiu. High order asymptotic preserving Hermite WENO fast sweeping method for the steady-state S_{N} transport equations. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01965-x.

Wang:2022:BFS

[3266] Yue Wang, Xiangyun Meng, and Yonghai Li. The Bogner–Fox–Schmit element finite volume methods on the Shishkin mesh for fourth-order singularly perturbed elliptic problems. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01969-7.

Li:2022:DLG

[3267] Jian Li, Jing Yue, Wen Zhang, and Wansuo Duan. The deep learning Galerkin method for the general Stokes equations. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01930-8.

Larios-Cardenas:2022:ECN

[3268] Luis Ángel Larios-Cárdenas and Frédéric Gibou. Error-correcting neural networks for two-dimensional curvature computation in the level-set method. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01952-2.

Liu:2022:ALT

[3269] Tianhang Liu and Yimin Wei. The abstract Laplacian tensor of a hypergraph with applications in clustering. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01973-x.

Al-Maskari:2022:EEA

[3270] Mariam Al-Maskari and Samir Karaa. Error estimates for approximations of time-fractional biharmonic equation with nonsmooth data. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01971-z.

Wu:2022:ECI

[3271] Tingting Wu, Zhihui Mao, Zeyu Li, Yonghua Zeng, and Tieyong Zeng. Efficient color image segmentation via quaternion-based L_{1} / L_{2} regularization. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01970-0.

Lepe:2022:PAM

[3272] Felipe Lepe, Gonzalo Rivera, and Jesus Vellojin. A posteriori analysis for a mixed FEM discretization of the linear elasticity spectral problem. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01972-y.

Petras:2022:MSL

[3273] Argyrios Petras, Leevan Ling, and Steven J. Ruuth. Meshfree semi-Lagrangian methods for solving surface advection PDEs. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01966-w.

Cao:2022:IVE

[3274] Shuhao Cao, Long Chen, Ruchi Guo, and Frank Lin. Immersed virtual element methods for elliptic interface problems in two dimensions. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01949-x.

Tao:2022:AED

[3275] Qi Tao, Liangyue Ji, Jennifer K. Ryan, and Yan Xu. Accuracyenhancement of discontinuous Galerkin methods for PDEs containing high order spatial derivatives. Journal of Scientific Computing, 93 (1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01967-9.

Zhang:2022:SOM

[3276] Xiaotao Zhang, Tiegang Liu, Changsheng Yu, Chengliang Feng, Zhiqiang Zeng, and Kun Wang. A second-order modified ghost fluid method (2nd-MGFM) with discontinuous Galerkin method for 1-D compressible multi-medium problem with cylindrical and spherical symmetry. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01975-9.

Caboussat:2022:LSM

[3277] Alexandre Caboussat, Roland Glowinski, and Dimitrios Gourzoulidis. A least-squares method for the solution of the non-smooth prescribed Jacobian equation. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01968-8.

Zhang:2022:PET

[3278] Jiachuan Zhang, Ran Zhang, and Xiaoshen Wang. A posteriori estimates of Taylor–Hood element for Stokes problem using auxiliary subspace techniques. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01924-6.

Erickson:2022:NSS

[3279] Brittany A. Erickson, Jeremy E. Kozdon, and Tobias Harvey. A nonstiff summation-by-parts finite difference method for the scalar wave equation in second order form: Characteristic boundary conditions and nonlinear interfaces. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01961-1.

Avrachenkov:2022:RLG

[3280] Konstantin Avrachenkov, Patrick Brown, and Nelly Litvak. Red light green light method for solving large Markov chains. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01976-8.

Liang:2022:WGF

[3281] Qigang Liang, Xuejun Xu, and Liuyao Yuan. A weak Galerkin finite element method can compute both upper and lower eigenvalue bounds. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01986-6.

Wu:2022:IDN

[3282] Wei Wu, Xinlong Feng, and Hui Xu. Improved deep neural networks with domain decomposition in solving partial differential equations. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01980-y.

Zhu:2022:TCI

[3283] Hongqiang Zhu, Zhihuan Wang, Haiyun Wang, Qiang Zhang, and Zhen Gao. Troubled-cell indication using K-means clustering with unified parameters. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01987-5.

Sun:2022:EEE

[3284] Yabing Sun and Quan Zhou. Error estimate of exponential time differencing Runge–Kutta scheme for the epitaxial growth model without slope selection. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01977-7.

Kang:2022:EPE

[3285] Shinhoo Kang and Emil M. Constantinescu. Entropy-preserving and entropy-stable relaxation IMEX and multirate time-stepping methods. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01982-w.

Iyiola:2022:IVG

[3286] Olaniyi S. Iyiola and Yekini Shehu. Inertial version of generalized projected reflected gradient method. Journal of Scientific Computing, 93 (1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01989-3.

Zhang:2022:SSM

[3287] Fengshan Zhang, Yongkui Zou, Shimin Chai, Ran Zhang, and Yanzhao Cao. Splitting-up spectral method for nonlinear filtering problems with correlation noises. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01994-6.

Shravani:2022:AEP

[3288] N Shravani, G Murali Mohan Reddy, and A. K. Pani. Anisotropic a Posteriori error analysis for the two-step backward differentiation formula for parabolic integro–differential equation. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01996-4.

Rudel:2022:NFP

[3289] Clément Rudel, Sébastien Pernet, and Jean-Philippe Brazier. Numerical factorization of propagation operator for hyperbolic equations and application to one-way, true amplitude one-way equations and Bremmer series. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01985-7.

Liu:2022:LDC

[3290] Wenjia Liu and Shuo Zhang. A lowest-degree conservative finite element scheme for incompressible Stokes problems on general triangulations. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-01974-w.

Dong:2022:AEI

[3291] Wenbin Dong, Yingjie Liu, and Hansong Tang. Accuracy enhancing interface treatment algorithm: The back and forth error compensation and correction method. Journal of Scientific Computing, 93 (1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01988-4.

Ji:2022:SSO

[3292] Shaolin Ji, Shige Peng, Ying Peng, and Xichuan Zhang. Solving stochastic optimal control problem via stochastic maximum principle with deep learning method. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01979-5.

Shi:2022:CHO

[3293] Jiankang Shi, Minghua Chen, Yubin Yan, and Jianxiong Cao. Correction of high-order L_{k} approximation for subdiffusion. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-01984-8.

Xu:2022:CVE

[3294] Yang Xu, Zhenguo Zhou, and Jingjun Zhao. Conforming virtual element methods for Sobolev equations. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01997-3.

Zhang:2022:SAC

[3295] Hong Zhang, Zhengyu Liu, Emil Constantinescu, and Robert Jacob. Stability analysis of coupled advection–diffusion models with bulk interface condition. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01983-9.

Lyu:2022:SFO

[3296] Pin Lyu and Seakweng Vong. A symmetric fractional-order reduction method for direct nonuniform approximations of semilinear diffusionwave equations. Journal of Scientific Computing, 93(1):??, October 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02000-9.

Kochi:2022:DGO

[3297] S. R. Siva Prasad Kochi and M. Ramakrishna. A discontinuous Galerkin overset scheme using WENO reconstruction and subcells for two-dimensional problems. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01991-9.

Penner:2022:AHO

[3298] David A. Craig Penner and David W. Zingg. Accurate high-order tensor-product generalized summation-by-parts discretizations of hyperbolic conservation laws: General curved domains and functional superconvergence. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-01990-w.

Sun:2022:CIQ

[3299] Zhengjie Sun, Wenwu Gao, and Ran Yang. A convergent iterated quasiinterpolation for periodic domain and its applications to surface PDEs. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01998-2.

Tang:2022:AHO

[3300] Tao Tang, Xu Wu, and Jiang Yang. Arbitrarily high order and fully discrete extrapolated RK-SAV/DG schemes for phase-field gradient flows. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-01995-5.

Li:2022:OCS

[3301] Yaping Li, Weidong Zhao, and Wenju Zhao. Optimal convergence of the scalar auxiliary variable finite element method for the natural convection equations. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-01981-x.

Gao:2022:KIS

[3302] Zixuan Gao, Jiuyang Liang, and Zhenli Xu. A kernel-independent sum-of-exponentials method. Journal of Scientific Computing, 93(2): ??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01999-1.

Huang:2022:THJ

[3303] Jinzhi Huang and Zhongxiao Jia. Two harmonic Jacobi–Davidson methods for computing a partial generalized singular value decomposition of a large matrix pair. Journal of Scientific Computing, 93 (2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01993-7.

Cavalcante:2022:MFA

[3304] T. M. Cavalcante, R. J. M. Lira Filho, A. C. R. Souza, D. K. E. Carvalho, and P. R. M. Lyra. A multipoint flux approximation with a diamond stencil and a non-linear defect correction strategy for the numerical solution of steady state diffusion problems in heterogeneous and anisotropic media satisfying the discrete maximum principle. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-01978-6.

Fu:2022:NAS

[3305] Lin Fu and Tian Liang. A new adaptation strategy for multiresolution method. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02012-5.

Liu:2022:EQA

[3306] Jun Liu and Hongfei Fu. An efficient QSC approximation of variableorder time-fractional mobile–immobile diffusion equations with variably diffusive coefficients. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-02007-2.

Sun:2022:SOQ

[3307] Tao Sun, Zhi Wang, Hai-Wei Sun, and Chengjian Zhang. A sixth-order quasi-compact difference scheme for multidimensional Poisson equations without derivatives of source term. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02003-6.

Zhou:2022:OEE

[3308] Lingling Zhou and Ruihan Guo. Optimal error estimates of the local discontinuous Galerkin method and high-order time discretization scheme for the Swift–Hohenberg equation. Journal of Scientific Computing, 93 (2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02014-3.

He:2022:TCG

[3309] Hongjin He, Chen Ling, and Wenhui Xie. Tensor completion via a generalized transformed tensor t-product decomposition without t-SVD. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02006-3.

Saez-Mischlich:2022:SDR

[3310] G. Sáez-Mischlich, J. Sierra-Ausín, and J. Gressier. The spectral difference Raviart–Thomas method for two and three-dimensional elements and its connection with the flux reconstruction formulation. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02002-7.

Lyu:2022:MDI

[3311] Cheng-Yao Lyu, Xi-Le Zhao, Ben-Zheng Li, Hao Zhang, and Ting-Zhu Huang. Multi-dimensional image recovery via fully-connected tensor network decomposition under the learnable transforms. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02009-0.

Caucao:2022:PEA

[3312] Sergio Caucao, Gabriel N. Gatica, Ricardo Oyarzúa, and Paulo Zúñiga. A posteriori error analysis of a mixed finite element method for the coupled Brinkman–Forchheimer and double-diffusion equations. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02010-7.

Gong:2022:HOA

[3313] Na Gong and Wanshan Li. High-order ADI–FDTD schemes for Maxwell’s equations with material interfaces in two dimensions. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02011-6.

Wang:2022:SNS

[3314] Ruyu Wang, Chao Zhang, Lichun Wang, and Yuanhai Shao. A stochastic Nesterov’s smoothing accelerated method for general nonsmooth constrained stochastic composite convex optimization. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02016-1.

Guo:2022:DPR

[3315] Shimin Guo, Wenjing Yan, Can Li, and Liquan Mei. Dissipationpreserving rational spectral-Galerkin method for strongly damped nonlinear wave system involving mixed fractional Laplacians in unbounded domains. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02008-1.

Koc:2022:VDD

[3316] Birgul Koc, Changhong Mou, Honghu Liu, Zhu Wang, Gianluigi Rozza, and Traian Iliescu. Verifiability of the data-driven variational multiscale reduced order model. Journal of Scientific Computing, 93(2): ??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02019-y.

Wang:2022:PEC

[3317] Wansheng Wang, Mengli Mao, and Yi Huang. A posteriori error control and adaptivity for the IMEX BDF2 method for PIDEs with application to options pricing models. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02013-4.

Busto:2022:NFT

[3318] Saray Busto and Michael Dumbser. A new family of thermodynamically compatible discontinuous Galerkin methods for continuum mechanics and turbulent shallow water flows. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02017-0.

Cicci:2022:DHD

[3319] Ludovica Cicci, Stefania Fresca, and Andrea Manzoni. Deep-HyROMnet: a deep learning-based operator approximation for hyper-reduction of nonlinear parametrized PDEs. Journal of Scientific Computing, 93 (2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02001-8.

Phan:2022:EIS

[3320] Duy Phan and Alexander Ostermann. Exponential integrators for second-order in time partial differential equations. Journal of Scientific Computing, 93(2):??, November 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02018-z.

Bruchhauser:2022:IAM

[3321] Marius Paul Bruchhäuser, Uwe Köcher, and Markus Bause. On the implementation of an adaptive multirate framework for coupled transport and flow. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02026-z.

Abedian:2022:RWF

[3322] Rooholah Abedian and Mehdi Dehghan. A RBF–WENO finite difference scheme for non-linear degenerate parabolic equations. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02022-3.

Duan:2022:NTM

[3323] Shan-Qi Duan, Xue-Feng Duan, and Xi-Le Zhao. A new tensor multirank approximation with total variation regularization for tensor completion. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-02005-4.

Chada:2022:IEM

[3324] N. K. Chada, H. Hoel, A. Jasra, and G. E. Zouraris. Improved efficiency of multilevel Monte Carlo for stochastic PDE through strong pairwise coupling. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02031-2.

Gao:2022:TOF

[3325] Weiguo Gao, Yingzhou Li, and Bichen Lu. Triangularized orthogonalizationfree method for solving extreme eigenvalue problems. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02025-0.

Li:2022:ECT

[3326] Xi Li, Yan Luo, and Minfu Feng. An efficient Chorin–Temam projection proper orthogonal decomposition based reduced-order model for nonstationary Stokes equations. Journal of Scientific Computing, 93 (3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02032-1.

Yang:2022:FEM

[3327] Yidu Yang, Shixi Wang, and Hai Bi. The finite element method for the elastic transmission eigenvalue problem with different elastic tensors. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02030-3.

Li:2022:NFM

[3328] Yu Li, Manting Xie, and Chunguang Xiong. A new family of mixed method for the biharmonic eigenvalue problem based on the first order equations of Hellan–Herrmann–Johnson type. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02024-1.

Prater-Bennette:2022:POL

[3329] Ashley Prater-Bennette, Lixin Shen, and Erin E. Tripp. The proximity operator of the log–sum penalty. Journal of Scientific Computing, 93 (3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02021-4.

Baffet:2022:EEA

[3330] Daniel H. Baffet, Yannik G. Gleichmann, and Marcus J. Grote. Error estimates for adaptive spectral decompositions. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02004-5.

Wu:2022:FCR

[3331] Chao Wu, Yunqing Huang, Nianyu Yi, Huayi Wei, and Jinyun Yuan. Function and curl recovery for the lowest order triangular edge element. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02027-y.

Chang:2022:GRR

[3332] Xiaokai Chang and Junfeng Yang. GRPDA revisited: Relaxed condition and connection to Chambolle–Pock’s primal–dual algorithm. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02033-0.

Cheng:2022:RNM

[3333] Fangxiong Cheng, Xu Hui, and Xinlong Feng. Retraction note to: Model order reduction method based on machine learning for parameterized time-dependent partial differential equations. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-022-02036-x. See [3218].

Zhu:2022:FDM

[3334] Jun Zhu and Jianxian Qiu. A finite difference mapped WENO scheme with unequal-size stencils for hyperbolic conservation laws. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02034-z.

Wu:2022:CAL

[3335] Zhizhang Wu and Zhiwen Zhang. Convergence analysis of the localized orthogonal decomposition method for the semiclassical Schrödinger equations with multiscale potentials. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02038-9.

Luo:2022:FPI

[3336] Songting Luo and Qing Huo Liu. A fixed-point iteration method for high frequency Helmholtz equations. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02039-8.

Yang:2022:UFE

[3337] Fanyi Yang and Xiaoping Xie. An unfitted finite element method by direct extension for elliptic problems on domains with curved boundaries and interfaces. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02035-y.

Leitenmaier:2022:HMM

[3338] Lena Leitenmaier and Olof Runborg. Heterogeneous multiscale methods for the Landau–Lifshitz equation. Journal of Scientific Computing, 93 (3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-01992-8.

Hong:2022:RAG

[3339] Qingguo Hong, Johannes Kraus, Miroslav Kuchta, Maria Lymbery, Kent-André Mardal, and Marie E. Rognes. Robust approximation of generalized Biot–Brinkman problems. Journal of Scientific Computing, 93 (3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02029-w.

Gu:2022:ALS

[3340] Shuting Gu, Hongqiao Wang, and Xiang Zhou. Active learning for saddle point calculation. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02040-1.

Duan:2022:RBP

[3341] Huoyuan Duan and Qiuyu Zhang. Residual-based a posteriori error estimates for the time-dependent Ginzburg–Landau equations of superconductivity. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02041-0.

Lei:2022:SPP

[3342] Zhen Lei, Lei Shi, and Chenyu Zeng. Solving parametric partial differential equations with deep rectified quadratic unit neural networks. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02015-2.

Wang:2022:NMP

[3343] Lai Wang, Will Trojak, Freddie Witherden, and Antony Jameson. Nonlinear p-multigrid preconditioner for implicit time integration of compressible Navier–Stokes equations with p-adaptive flux reconstruction. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02037-w.

Linders:2022:EPS

[3344] Viktor Linders. On an eigenvalue property of summation-by-parts operators. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-02042-z.

Thein:2022:ENS

[3345] Ferdinand Thein, Evgeniy Romenski, and Michael Dumbser. Exact and numerical solutions of the Riemann problem for a conservative model of compressible two-phase flows. Journal of Scientific Computing, 93 (3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02028-x.

Wang:2022:TCE

[3346] Yinghua Wang, Bao-Shan Wang, Leevan Ling, and Wai Sun Don. A time-continuous embedding method for scalar hyperbolic conservation laws on manifolds. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02023-2.

Duan:2022:DFM

[3347] Huoyuan Duan, Can Wang, and Zhijie Du. A div FOSLS method suitable for quadrilateral RT and hexahedral RTN H(div)-elements. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02043-y.

Li:2022:CES

[3348] Meng Li. Cut-off error splitting technique for conservative nonconforming VEM for N-coupled nonlinear Schrödinger–Boussinesq equations. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02050-z.

Zhang:2022:TGV

[3349] Huayan Zhang and Zhichao Peng. Total generalized variation for triangulated surface data. Journal of Scientific Computing, 93(3):??, December 2022. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02047-8.

He:2023:EQT

[3350] Zhuo-Heng He, Xiang-Xiang Wang, and Yun-Fan Zhao. Eigenvalues of quaternion tensors with applications to color video processing. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02058-5.

Gruber:2023:MMC

[3351] Anthony Gruber, Max Gunzburger, Lili Ju, and Zhu Wang. A multifidelity Monte Carlo method for realistic computational budgets. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02051-y.

Huang:2023:CPF

[3352] Jinzhi Huang and Zhongxiao Jia. A cross-product free Jacobi–Davidson type method for computing a partial generalized singular value decomposition of a large matrix pair. Journal of Scientific Computing, 94 (1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02053-w.

Nonino:2023:PBS

[3353] Monica Nonino, Francesco Ballarin, Gianluigi Rozza, and Yvon Maday. Projection based semi-implicit partitioned reduced basis method for fluid-structure interaction problems. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02049-6.

Yan:2023:KBM

[3354] Qile Yan, Shixiao W. Jiang, and John Harlim. Kernel-based methods for solving time-dependent advection–diffusion equations on manifolds. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02045-w.

Zeng:2023:RPC

[3355] Chao Zeng. Rank properties and computational methods for orthogonal tensor decompositions. Journal of Scientific Computing, 94(1): ??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-02054-9.

Huang:2023:MLM

[3356] Juntao Huang, Yingda Cheng, Andrew J. Christlieb, and Luke F. Roberts. Machine learning moment closure models for the radiative transfer equation III: Enforcing hyperbolicity and physical characteristic speeds. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02056-7.

Marino:2023:SNL

[3357] Raffaele Marino and Nicolas Macris. Solving non-linear Kolmogorov equations in large dimensions by using deep learning: a numerical comparison of discretization schemes. Journal of Scientific Computing, 94 (1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02044-x.

Wang:2023:TGF

[3358] Yanli Wang and Jie Xu. Q-tensor gradient flow with quasi-entropy and discretizations preserving physical constraints. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02060-x.

Yin:2023:CEH

[3359] Xu Yin and Waixiang Cao. A class of efficient Hamiltonian conservative spectral methods for Korteweg–de Vries equations. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02061-w.

Cheng:2023:GLA

[3360] Sibo Cheng, Jianhua Chen, Charitos Anastasiou, Panagiota Angeli, Omar K. Matar, Yi-Ke Guo, Christopher C. Pain, and Rossella Arcucci. Generalised latent assimilation in heterogeneous reduced spaces with machine learning surrogate models. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02059-4.

Nie:2023:DRO

[3361] Jiawang Nie, Liu Yang, Suhan Zhong, and Guangming Zhou. Distributionally robust optimization with moment ambiguity sets. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02063-8.

Xiang:2023:OCA

[3362] Yahong Xiang, Can Huang, and Huangxin Chen. Optimal convergence analysis of a fully discrete scheme for the stochastic Stokes–Darcy equations. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-02057-6.

Tominec:2023:RVS

[3363] Igor Tominec and Murtazo Nazarov. Residual viscosity stabilized RBF– FD methods for solving nonlinear conservation laws. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02055-8.

Li:2023:SIM

[3364] Peng Li, Tingting Li, Wai-Sun Don, and Bao-Shan Wang. Scaleinvariant multi-resolution alternative WENO scheme for the Euler equations. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-02065-6.

Liu:2023:BQN

[3365] Ren Liu, Fengmiao Bian, and Xiaoqun Zhang. Binary quantized network training with sharpness-aware minimization. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02064-7.

Jiang:2023:FAA

[3366] Nan Jiang and Huanhuan Yang. Fast and accurate artificial compressibility ensemble algorithms for computing parameterized Stokes– Darcy flow ensembles. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02069-2.

Upperman:2023:FOP

[3367] Johnathon Upperman and Nail K. Yamaleev. First-order positivitypreserving entropy stable scheme for the 3-D compressible Navier– Stokes equations. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02062-9.

Li:2023:SBP

[3368] Ji Li. Solving blind ptychography effectively via linearized alternating direction method of multipliers. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02072-7.

Zhou:2023:DSM

[3369] Han Zhou and Wenjun Ying. A dimension splitting method for time dependent PDEs on irregular domains. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02066-5.

Duan:2023:AFH

[3370] Songyao Duan and Haijun Wu. Adaptive FEM for Helmholtz equation with large wavenumber. Journal of Scientific Computing, 94 (1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02074-5.

Zhong:2023:MRM

[3371] Min Zhong, Quoc Thong Le Gia, and Ian Hugh Sloan. A multiscale RBF method for severely ill-posed problems on spheres. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02046-9.

Franchini:2023:LSB

[3372] Giorgia Franchini, Federica Porta, Valeria Ruggiero, and Ilaria Trombini. A line search based proximal stochastic gradient algorithm with dynamical variance reduction. Journal of Scientific Computing, 94 (1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02084-3.

Chiocchetti:2023:ECF

[3373] Simone Chiocchetti and Michael Dumbser. An exactly curl-free staggered semi-implicit finite volume scheme for a first order hyperbolic model of viscous two-phase flows with surface tension. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02077-2.

Silva:2023:DAP

[3374] Vinicius L. S. Silva, Claire E. Heaney, Yaqi Li, and Christopher C. Pain. Data assimilation predictive GAN (DA-PredGAN) applied to a spatiotemporal compartmental model in epidemiology. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02078-1. See correction [3470].

Lee:2023:LFL

[3375] Sanghyun Lee and Son-Young Yi. Locking-free and locally-conservative enriched Galerkin method for poroelasticity. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-022-02079-0.

Corallo:2023:STD

[3376] Daniele Corallo, Willy Dörfler, and Christian Wieners. Space-time discontinuous Galerkin methods for weak solutions of hyperbolic linear symmetric Friedrichs systems. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02076-3.

Fong:2023:SEN

[3377] Ki Wai Fong and Shingyu Leung. Spherical Essentially Non-oscillatory (SENO) interpolation. Journal of Scientific Computing, 94(1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02080-7.

Kang:2023:EQR

[3378] Hongchao Kang and Hong Wang. An efficient quadrature rule for the oscillatory infinite generalized Bessel transform with a general oscillator and its error analysis. Journal of Scientific Computing, 94 (1):??, January 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02081-6.

Wang:2023:ESI

[3379] Qinghe Wang, Mingyang Pan, Yu-Hau Tseng, and Dongdong He. An energy stable immersed boundary method for deformable membrane problem with non-uniform density and viscosity. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02092-3.

Zhang:2023:SPC

[3380] Mingzhu Zhang and Lijun Yi. Superconvergent postprocessing of the continuous Galerkin time stepping method for nonlinear initial value problems with application to parabolic problems. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02086-1.

Zhang:2023:AHO

[3381] Gengen Zhang, Chaolong Jiang, and Hao Huang. Arbitrarily high-order energy-preserving schemes for the Zakharov–Rubenchik equations. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02075-4.

Yu:2023:ESA

[3382] Yue Yu, Jiansong Zhang, and Rong Qin. The exponential SAV approach for the time-fractional Allen–Cahn and Cahn–Hilliard phase-field models. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-02085-2.

Dang:2023:NAS

[3383] Haikun Dang, Hehu Xie, Gang Zhao, and Chenguang Zhou. A nonnested augmented subspace method for elliptic eigenvalue problems with curved interfaces. Journal of Scientific Computing, 94(2): ??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-022-02089-y.

Xu:2023:NMM

[3384] Wei-Wei Xu and Michael K. Ng. A new matrix maximization model for computing ratios of generalized singular values from high-order GSVD. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02071-8.

Pasini:2023:HMR

[3385] Massimiliano Lupo Pasini and Simona Perotto. Hierarchical model reduction driven by machine learning for parametric advection–diffusion– reaction problems in the presence of noisy data. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02073-6.

Xie:2023:EAM

[3386] Hehu Xie, Manting Xie, Xiaobo Yin, and Gang Zhao. An efficient adaptive mesh redistribution method for nonlinear eigenvalue problems in Bose–Einstein condensates. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02093-2.

Siena:2023:DDR

[3387] Pierfrancesco Siena, Michele Girfoglio, Francesco Ballarin, and Gianluigi Rozza. Data-driven reduced order modelling for patient-specific hemodynamics of coronary artery bypass grafts with physical and geometrical parameters. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02082-5.

Hou:2023:IES

[3388] Dianming Hou and Zhonghua Qiao. An implicit–explicit second-order BDF numerical scheme with variable steps for gradient flows. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02094-1.

Dang:2023:EEE

[3389] Haikun Dang, Yifan Wang, Hehu Xie, and Chenguang Zhou. Enhanced error estimates for augmented subspace method. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02090-5.

Carter:2023:SOU

[3390] John Carter, Daozhi Han, and Nan Jiang. Second order, unconditionally stable, linear ensemble algorithms for the magnetohydrodynamics equations. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-022-02091-4.

Xu:2023:MMN

[3391] Fei Xu, Manting Xie, and Meiling Yue. Multigrid method for nonlinear eigenvalue problems based on Newton iteration. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-022-02070-9.

Watanabe:2023:EAV

[3392] Yoshitaka Watanabe, Takehiko Kinoshita, and Mitsuhiro T. Nakao. Efficient approaches for verifying the existence and bound of inverse of linear operators in Hilbert spaces. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02097-6.

Li:2023:FFM

[3393] Yue Li, Lin Fu, and Nikolaus A. Adams. A family of fast multi-resolution ENO schemes for compressible flows. Journal of Scientific Computing, 94(2):??, February 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02095-0.

Novello:2023:GOS

[3394] Paul Novello, Gaël Poëtte, David Lugato, and Pietro Marco Congedo. Goal-oriented sensitivity analysis of hyperparameters in deep learning. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-022-02083-4.

Gu:2023:FDS

[3395] Yiqi Gu and Jie Shen. A fictitious domain spectral method for solving the Helmholtz equation in exterior domains. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02098-5.

Mao:2023:CEE

[3396] Shipeng Mao and Wendong Xue. Convergence and error estimates of a mixed discontinuous Galerkin-finite element method for the semistationary compressible Stokes system. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02096-7.

Freret:2023:HOC

[3397] L. Freret, C. N. Ngigi, T. B. Nguyen, H. De Sterck, and C. P. T. Groth. High-order CENO finite-volume scheme with anisotropic adaptive mesh refinement: Efficient inexact Newton method for steady three-dimensional flows. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-022-02068-3.

Michel:2023:SAH

[3398] Sixtine Michel, Davide Torlo, Mario Ricchiuto, and Rémi Abgrall. Spectral analysis of high order continuous FEM for hyperbolic PDEs on triangular meshes: Influence of approximation, stabilization, and timestepping. Journal of Scientific Computing, 94(3):??, March 2023. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-022-02087-
0.

Qian:2023:FDA

[3399] Lingzhi Qian, Chunya Wu, Huiping Cai, Xinlong Feng, and Yuanyang Qiao. A fully-decoupled artificial compressible Crank–Nicolson-leapfrog time stepping scheme for the phase field model of two-phase incompressible flows. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-022-02048-
7.

Wang:2023:SRN

[3400] Jiani Wang, Xiao Wang, and Liwei Zhang. Stochastic regularized Newton methods for nonlinear equations. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02099-4.

Leng:2023:PEA

[3401] Haitao Leng. A posteriori error analysis for pressure-robust HDG methods for the stationary incompressible Navier–Stokes equations. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02104-w.

Fang:2023:SSP

[3402] Shuixin Fang, Weidong Zhao, and Tao Zhou. Strong stability preserving multistep schemes for forward backward stochastic differential equations. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02111-x.

Calhoun:2023:CGA

[3403] Donna Calhoun, Erik Chudzik, and Christiane Helzel. The Cartesian grid active flux method with adaptive mesh refinement. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02106-8.

Tomobe:2023:EBO

[3404] Haruka Tomobe, Vikas Sharma, Harusato Kimura, and Hitoshi Morikawa. An energy-based overset finite element method for pseudostatic structural analysis. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02113-9.

Meng:2023:SIP

[3405] Jian Meng, Lourenço Beirão da Veiga, and Lorenzo Mascotto. Stability and interpolation properties for Stokes-like virtual element spaces. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02112-w.

Jin:2023:APN

[3406] Shi Jin, Zheng Ma, and Keke Wu. Asymptotic-preserving neural networks for multiscale time-dependent linear transport equations. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02100-0.

Liu:2023:MNC

[3407] Chang Liu and William Henshaw. Multigrid with nonstandard coarselevel operators and coarsening factors. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02103-x.

Bencomo:2023:DAC

[3408] Mario J. Bencomo and Jesse Chan. Discrete adjoint computations for relaxation Runge–Kutta methods. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02102-y.

Ma:2023:UCP

[3409] Yumin Ma, Xingju Cai, Bo Jiang, and Deren Han. Understanding the convergence of the preconditioned PDHG method: a view of indefinite proximal ADMM. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02105-9.

Han:2023:IBU

[3410] Yihui Han, Xiao-Ping Wang, and Xiaoping Xie. An interface/boundaryunfitted EXtended HDG method for linear elasticity problems. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02107-7.

Li:2023:SPG

[3411] Shengyue Li and Wanrong Cao. On spectral Petrov–Galerkin method for solving optimal control problem governed by fractional diffusion equations with fractional noise. Journal of Scientific Computing, 94 (3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-022-02088-z.

Lai:2023:CSB

[3412] Ming-Jun Lai and Zhaiming Shen. A compressed sensing based least squares approach to semi-supervised local cluster extraction. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-022-02052-x.

Boon:2023:RBM

[3413] Wietse M. Boon and Alessio Fumagalli. A reduced basis method for Darcy flow systems that ensures local mass conservation by using exact discrete complexes. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02119-3.

Garner:2023:LCF

[3414] Casey Garner and Shuzhong Zhang. Linearly-convergent FISTA variant for composite optimization with duality. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02101-z.

Leng:2023:EEE

[3415] Haitao Leng. Error estimates of EDG–HDG methods for the Stokes equations with Dirac measures. Journal of Scientific Computing, 94 (3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02116-6.

Scribano:2023:DFE

[3416] Carmelo Scribano, Giorgia Franchini, Marco Prato, and Marko Bertogna. DCT-Former: Efficient self-attention with discrete cosine transform. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02125-5.

Zheng:2023:SGM

[3417] Chuqi Zheng, Jiayu Qiu, Qin Li, and Xinghui Zhong. Stochastic Galerkin methods for time-dependent radiative transfer equations with uncertain coefficients. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02134-4.

Yang:2023:BMS

[3418] Jinda Yang, Haiming Song, Xinxin Li, and Di Hou. Block mirror stochastic gradient method for stochastic optimization. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02110-y.

Hu:2023:UCC

[3419] Dongdong Hu, Yayun Fu, Wenjun Cai, and Yushun Wang. Unconditional convergence of conservative spectral Galerkin methods for the coupled fractional nonlinear Klein–Gordon–Schrödinger equations. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02108-6.

Lenz:2023:CTF

[3420] Christoph Lenz, Luiz Felipe Aguinsky, Andreas Hössinger, and Josef Weinbub. A complementary topographic feature detection algorithm based on surface curvature for three-dimensional level-set functions. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02133-5.

Dehghan:2023:OEE

[3421] Mehdi Dehghan, Zeinab Gharibi, and Ricardo Ruiz-Baier. Optimal error estimates of coupled and divergence-free virtual element methods for the Poisson–Nernst–Planck/Navier–Stokes equations and applications in electrochemical systems. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02126-4.

Izuchukwu:2023:SCF

[3422] Chinedu Izuchukwu, Simeon Reich, Yekini Shehu, and Adeolu Taiwo. Strong convergence of forward-reflected-backward splitting methods for solving monotone inclusions with applications to image restoration and optimal control. Journal of Scientific Computing, 94(3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02132-6.

Romor:2023:NLM

[3423] Francesco Romor, Giovanni Stabile, and Gianluigi Rozza. Non-linear manifold reduced-order models with convolutional autoencoders and reduced over-collocation method. Journal of Scientific Computing, 94 (3):??, March 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02128-2.

Tang:2023:UES

[3424] Yangyang Tang, Guang an Zou, and Jian Li. Unconditionally energystable finite element scheme for the chemotaxis–fluid system. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02118-4.

Hamadi:2023:DDK

[3425] M. A. Hamadi, K. Jbilou, and A. Ratnani. A data-driven Krylov model order reduction for large-scale dynamical systems. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02122-8.

Zlotnik:2023:CPC

[3426] Alexander Zlotnik and Raimondas Ciegis. On construction and properties of compact 4th order finite-difference schemes for the variable coefficient wave equation. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02127-3.

Liao:2023:THL

[3427] Yulei Liao, Pingbing Ming, and Yun Xu. Taylor–Hood like finite elements for nearly incompressible strain gradient elasticity problems. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02135-3.

Zou:2023:FDD

[3428] Guang an Zou, Zhaohua Li, and Xiaofeng Yang. Fully discrete discontinuous Galerkin numerical scheme with second-order temporal accuracy for the hydrodynamically coupled lipid vesicle model. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02129-1.

Bertoluzza:2023:HME

[3429] Silvia Bertoluzza, Giovanna Guidoboni, Romain Hild, Daniele Prada, Christophe Prud’homme, Riccardo Sacco, Lorenzo Sala, and Marcela Szopos. A HDG method for elliptic problems with integral boundary condition: Theory and applications. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02109-5.

Kassem:2023:SOS

[3430] Mohamed Abd El-Hady Kassem. Second-order symmetric duality for multiple objectives nonlinear programming under generalizations of cone-preinvexity functions. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02114-8.

LeBorne:2023:GRF

[3431] Sabine Le Borne and Willi Leinen. Guidelines for RBF–FD discretization: Numerical experiments on the interplay of a multitude of parameter choices. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02123-7.

Lu:2023:TLS

[3432] Peipei Lu, Andreas Rupp, and Guido Kanschat. Two-level Schwarz methods for hybridizable discontinuous Galerkin methods. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02121-9.

Kheirfam:2023:CPI

[3433] B. Kheirfam. Corrector–predictor interior-point method with new search direction for semidefinite optimization. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02137-1.

Yamaleev:2023:HOP

[3434] Nail K. Yamaleev and Johnathon Upperman. High-order positivitypreserving entropy stable schemes for the 3-D compressible Navier– Stokes equations. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-023-02136-
2.

Cheridito:2023:CCE

[3435] Patrick Cheridito and Balint Gersey. Computation of conditional expectations with guarantees. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02130-8.

Lee:2023:PHM

[3436] Kisun Lee and Xindong Tang. On the polyhedral homotopy method for solving generalized Nash equilibrium problems of polynomials. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02138-0. See correction

[3505].

Ma:2023:FDD

[3437] Huimin Ma and Pengzhan Huang. A fully discrete decoupled finite element method for the thermally coupled incompressible magnetohydrodynamic problem. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02131-7.

Huang:2023:GTM

[3438] Lei Huang, Jiawang Nie, and Ya-Xiang Yuan. Generalized truncated moment problems with unbounded sets. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02139-z.

Qiu:2023:EAF

[3439] Hailong Qiu. Error analysis of fully discrete scheme for the Cahn– Hilliard–Magneto–Hydrodynamics problem. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02147-z.

Chamarthi:2023:IGB

[3440] Amareshwara Sainadh Chamarthi, Natan Hoffmann, Hiroaki Nishikawa, and Steven H. Frankel. Implicit gradients based conservative numerical scheme for compressible flows. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02141-5.

Xiao:2023:FRS

[3441] Tianbai Xiao, Jonas Kusch, Julian Koellermeier, and Martin Frank. A flux reconstruction stochastic Galerkin scheme for hyperbolic conservation laws. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02143-3.

Johnson:2023:PPL

[3442] Erica R. Johnson, James A. Rossmanith, and Christine Vaughan. Positivity-preserving Lax–Wendroff discontinuous Galerkin schemes for quadrature-based moment-closure approximations of kinetic models. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02117-5.

Wang:2023:PFO

[3443] Ruishu Wang, Zhuoran Wang, and Jiangguo Liu. Penalty-free any-order weak Galerkin FEMs for linear elasticity on quadrilateral meshes. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02151-3.

Kuzu:2023:RFB

[3444] Serpil Yalcin Kuzu. Random forest based multiclass classification approach for highly skewed particle data. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02144-2.

Wang:2023:SGA

[3445] Yue Wang and Jiequan Li. Stiffened gas approximation and GRP resolution for compressible fluid flows of real materials. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02140-6.

Ivagnes:2023:TML

[3446] Anna Ivagnes, Nicola Demo, and Gianluigi Rozza. Towards a machine learning pipeline in reduced order modelling for inverse problems: Neural networks for boundary parametrization, dimensionality reduction and solution manifold approximation. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02142-4.

Ma:2023:EEF

[3447] Ying Ma and Teng Zhang. Error estimates of finite difference methods for the biharmonic nonlinear Schrödinger equation. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02124-6.

Yang:2023:RDS

[3448] Bohan Yang and Hongxing Rui. A robust discrete scheme based on staggered grids for poroelastic–elastic coupled problems. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02149-x.

Ma:2023:FCM

[3449] Zheng Ma and Chengming Huang. Fractional collocation method for third-kind Volterra integral equations with nonsmooth solutions. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02155-z.

Wang:2023:PNP

[3450] Yinkun Wang. Parallel numerical Picard iteration methods. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02156-y.

Jiao:2023:JLS

[3451] Yuling Jiao, Dingwei Li, Min Liu, Xiliang Lu, and Yuanyuan Yang. Just least squares: Binary compressive sampling with low generative intrinsic dimension. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02158-w.

Huang:2023:ATS

[3452] Shijie Huang, Xufeng Xiao, and Xinlong Feng. An adaptive time-stepping method for the binary fluid-surfactant phase field model on evolving surfaces. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02150-4.

Wu:2023:LSB

[3453] Keyi Wu, Thomas O’Leary-Roseberry, Peng Chen, and Omar Ghattas. Large-scale Bayesian optimal experimental design with derivativeinformed projected neural network. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02145-1.

Guo:2023:LIO

[3454] Ruchi Guo, Jiahua Jiang, and Yi Li. Learn an index operator by CNN for solving diffusive optical tomography: a deep direct sampling method. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02115-7.

Li:2023:EST

[3455] Lin Li, Li-Lian Wang, and Huiyuan Li. An efficient spectral trust-region deflation method for multiple solutions. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02154-0.

Hu:2023:ASR

[3456] Mengqi Hu, Yifei Lou, Bao Wang, Ming Yan, Xiu Yang, and Qiang Ye. Accelerated sparse recovery via gradient descent with nonlinear conjugate gradient momentum. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02148-y.

Huang:2023:TSP

[3457] Jun Huang and Gang Wu. Truncated and sparse power methods with partially updating for large and sparse higher-order PageRank problems. Journal of Scientific Computing, 95(1):??, April 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02146-0.

Ni:2023:NCP

[3458] Naxian Ni and Suchuan Dong. Numerical computation of partial differential equations by hidden-layer concatenated extreme learning machine. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02162-0.

Dolejsi:2023:AEH

[3459] Vít Dolejsí and Georg May. An anisotropic hp-mesh adaptation method for time-dependent problems based on interpolation error control. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02153-1.

Liu:2023:PPE

[3460] Qianqian Liu, Jianyu Jing, Maoqin Yuan, and Wenbin Chen. A positivity-preserving, energy stable BDF2 scheme with variable steps for the Cahn–Hilliard equation with logarithmic potential. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02163-z.

Ke:2023:QNN

[3461] Yifen Ke, Changfeng Ma, Zhigang Jia, Yajun Xie, and Riwei Liao. Quasi non-negative quaternion matrix factorization with application to color face recognition. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02157-x.

Wang:2023:PQN

[3462] Qinsi Wang and Wei Hong Yang. Proximal quasi-Newton method for composite optimization over the Stiefel manifold. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02165-x.

Sun:2023:GGR

[3463] Zheng Sun and Yulong Xing. On generalized Gauss–Radau projections and optimal error estimates of upwind-biased DG methods for the linear advection equation on special simplex meshes. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02166-w.

Aylwin:2023:RNN

[3464] Rubén Aylwin, Fernando Henríquez, and Christoph Schwab. ReLU neural network Galerkin BEM. Journal of Scientific Computing, 95(2): ??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02120-w.

Liu:2023:SNE

[3465] Chein-Shan Liu and Lin Qiu. Solving nonlinear elliptic inverse source, coefficient and conductivity problems by the methods with bases satisfying the boundary conditions automatically. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02167-9.

Koc:2023:UBD

[3466] Birgul Koc, Tomás Chacón Rebollo, and Samuele Rubino. Uniform bounds with difference quotients for proper orthogonal decomposition reduced order models of the Burgers equation. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02160-2.

Popov:2023:STA

[3467] I. S. Popov. Space-time adaptive ADER–DG finite element method with LST–DG predictor and a posteriori sub-cell WENO finite-volume limiting for simulation of non-stationary compressible multicomponent reactive flows. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02164-y.

Mulder:2023:USN

[3468] W. A. Mulder. Unisolvence of symmetric node patterns for polynomial spaces on the simplex. Journal of Scientific Computing, 95(2): ??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02161-1.

Gong:2023:NSR

[3469] Sining Gong, Johnny Guzmán, and Michael Neilan. A note on the shape regularity of Worsey–Farin splits. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02159-9.

Silva:2023:CDA

[3470] Vinicius L. S. Silva, Claire E. Heaney, Yaqi Li, and Christopher C. Pain. Correction to: Data assimilation predictive GAN (DA-PredGAN) applied to a spatio-temporal compartmental model in epidemiology. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02169-7. See [3374].

Zhang:2023:MWE

[3471] Jiahui Zhang, Yinhua Xia, and Yan Xu. Moving water equilibria preserving discontinuous Galerkin method for the shallow water equations. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02174-w.

Li:2023:FST

[3472] Chuan Li, Yiming Ren, Guangqing Long, Eric Boerman, and Shan Zhao. A fast sine transform accelerated high-order finite difference method for parabolic problems over irregular domains. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02177-7.

Qiu:2023:AAS

[3473] Changxin Qiu, Aaron Bendickson, Joshua Kalyanapu, and Jue Yan. Accuracy and architecture studies of residual neural network method for ordinary differential equations. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02173-x.

Ling:2023:AHS

[3474] Dan Ling and Zhiping Mao. Analysis and Hermite spectral approximation of diffusive–viscous wave equations in unbounded domains arising in geophysics. Journal of Scientific Computing, 95(2): ??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02175-9.

Dong:2023:PSA

[3475] Wandi Dong, Gaohang Yu, Liqun Qi, and Xiaohao Cai. Practical sketching algorithms for low-rank Tucker approximation of large tensors. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02172-y.

Barsukow:2023:ASN

[3476] Wasilij Barsukow. All-speed numerical methods for the Euler equations via a sequential explicit time integration. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02152-2.

Baker:2023:LPO

[3477] Justin Baker, Elena Cherkaev, Akil Narayan, and Bao Wang. Learning proper orthogonal decomposition of complex dynamics using heavy-ball neural ODEs. Journal of Scientific Computing, 95(2):??, May 2023. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-023-02176-
8.

Klein:2023:SDG

[3478] Simon-Christian Klein. Stabilizing discontinuous Galerkin methods using Dafermos’ entropy rate criterion: I — one-dimensional conservation laws. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02170-0.

Lai:2023:TSC

[3479] Ming-Jun Lai and Jinsil Lee. Trivariate spline collocation methods for numerical solution to 3D Monge–Ampère equation. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02183-9.

Li:2023:VTB

[3480] Yiqun Li, Hong Wang, and Xiangcheng Zheng. A viscoelastic Timoshenko beam model: Regularity and numerical approximation. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02187-5.

Quan:2023:GTH

[3481] Chaoyu Quan and Xu Wu. Global-in-time H^{1}-stability of L2-1_{σ} method on general nonuniform meshes for subdiffusion equation. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02184-8.

Ma:2023:AFG

[3482] Yunyun Ma and Jiguang Sun. Analysis of a Fourier–Galerkin method for the transmission eigenvalue problem based on a boundary integral formulation. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02197-3.

He:2023:NMS

[3483] Ying He and Xiaofei Zhao. Numerical methods for some nonlinear Schrödinger equations in soliton management. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02181-x.

Wang:2023:SAC

[3484] Lixiu Wang, Qian Zhang, and Zhimin Zhang. Superconvergence analysis of curlcurl-conforming elements on rectangular meshes. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02182-w.

Ham:2023:SOT

[3485] Seokjun Ham, Soobin Kwak, Chaeyoung Lee, Gyeonggyu Lee, and Junseok Kim. A second-order time-accurate unconditionally stable method for a gradient flow for the Modica–Mortola functional. Journal of Scientific Computing, 95(2):??, May 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02198-2.

Cai:2023:ESG

[3486] Haotao Cai. An efficient spectral-Galerkin method for second kind weakly singular VIEs with highly oscillatory kernels. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02180-y.

Marchildon:2023:NIS

[3487] André L. Marchildon and David W. Zingg. A non-intrusive solution to the ill-conditioning problem of the gradient-enhanced Gaussian covariance matrix for Gaussian processes. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02190-w.

Wang:2023:SPA

[3488] Bin Wang and Yaolin Jiang. Structure-preserving algorithms with uniform error bound and long-time energy conservation for highly oscillatory Hamiltonian systems. Journal of Scientific Computing, 95(3): ??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02178-6.

Hou:2023:MDD

[3489] Jiangyong Hou, Dan Hu, Xuejian Li, and Xiaoming He. Modeling and a domain decomposition method with finite element discretization for coupled dual-porosity flow and Navier–Stokes flow. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02185-7.

Dirckx:2023:CSF

[3490] S. Dirckx, D. Huybrechs, and R. Ongenae. On the computation of the SVD of Fourier submatrices. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02171-z.

Xu:2023:EEC

[3491] Yang Xu, Zhenguo Zhou, and Jingjun Zhao. Error estimates of conforming virtual element methods with a modified symmetric Nitsche’s formula for 2D semilinear parabolic equations. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02188-4.

Lin:2023:DMD

[3492] Yifan Lin, Zhen Gao, Yuanhong Chen, and Xiang Sun. A dynamic mode decomposition based reduced-order model for parameterized timedependent partial differential equations. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02200-x.

Cai:2023:NSF

[3493] Yao-Yuan Cai, Hai-Wei Sun, and Sik-Chung Tam. Numerical study of a fast two-level Strang splitting method for spatial fractional Allen–Cahn equations. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02196-4.

Huang:2023:NAM

[3494] Yuchen Huang, Weifeng Qiu, and Weiwei Sun. New analysis of mixed finite element methods for incompressible magnetohydrodynamics. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02189-3.

Li:2023:EAU

[3495] Yuan Li and Rong An. Error analysis of a unconditionally stable BDF2 finite element scheme for the incompressible flows with variable density. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02205-6.

Li:2023:SDF

[3496] Hang Li and Chunmei Su. A semi-discrete first-order low regularity exponential integrator for the “good” Boussinesq equation without loss of regularity. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02201-w.

LeFloch:2023:CMF

[3497] Philippe G. LeFloch and Jean-Marc Mercier. A class of mesh-free algorithms for some problems arising in finance and machine learning. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02179-5.

Ji:2023:HOF

[3498] Zhe Ji, Tian Liang, and Lin Fu. High-order finite-volume TENO schemes with dual ENO-like stencil selection for unstructured meshes. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02199-1.

DeSterck:2023:ERR

[3499] Hans De Sterck, Chi-Wang Shu, and Rémi Abgrall. Enhancing reproducibility of research papers in SISC, JSC and JCP. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02193-7. See correction [3508].

Hanke:2023:RTB

[3500] Andrea Hanke and Manuel Torrilhon. Representation theory based algorithm to compute Boltzmann’s bilinear collision operator in the irreducible spectral Burnett ansatz efficiently. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02168-8.

Gatica:2023:MPM

[3501] Gabriel N. Gatica, Nicolás Núñez, and Ricardo Ruiz-Baier. Mixedprimal methods for natural convection driven phase change with Navier– Stokes–Brinkman equations. Journal of Scientific Computing, 95(3): ??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02202-9.

Huang:2023:DVD

[3502] Qian Huang, Yihong Chen, and Wen-An Yong. Discrete-velocity-direction models of BGK-type with minimum entropy: I. Basic idea. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02211-8.

Zhao:2023:PPE

[3503] Lina Zhao. A priori and a posteriori error analysis of TDNNS method for linear elasticity problem under minimal regularity. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02213-6.

Drzisga:2023:SMF

[3504] Daniel Drzisga, Tobias Köppl, and Barbara Wohlmuth. A semi matrixfree twogrid preconditioner for the Helmholtz equation with near optimal shifts. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02195-5.

Lee:2023:CPH

[3505] Kisun Lee and Xindong Tang. Correction to: On the polyhedral homotopy method for solving generalized Nash equilibrium problems of polynomials. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02192-8. See

[3436].

Amat:2023:ACH

[3506] Sergio Amat, Zhilin Li, Juan Ruiz-Álvarez, Concepción Solano, and Juan C. Trillo. Adapting cubic Hermite splines to the presence of singularities through correction terms. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02191-9.

Yang:2023:CLM

[3507] Zheng Yang and Fanhai Zeng. A corrected L1 method for a timefractional subdiffusion equation. Journal of Scientific Computing, 95 (3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-023-02204-7.

Sterck:2023:CER

[3508] Hans De Sterck, Chi-Wang Shu, and Rémi Abgrall. Correction to: Enhancing reproducibility of research papers in SISC, JSC and JCP. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02224-3. See [3499].

Wang:2023:DGM

[3509] Fei Wang, Sheheryar Shah, and Bangmin Wu. Discontinuous Galerkin methods for hemivariational inequalities in contact mechanics. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02212-7.

Shi:2023:CAS

[3510] Shuxun Shi and Wenbin Chen. Convergence and asymptotic stability of the BDF schemes for the nonlocal partial differential equations with delay. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02214-5.

Wang:2023:ESG

[3511] Zhongqing Wang, Xian Wen, and Guoqing Yao. An efficient spectral-Galerkin method for elliptic equations in 2D complex geometries. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02207-4.

Du:2023:OFB

[3512] Jie Du, Yong Liu, and Yang Yang. An oscillation-free bound-preserving discontinuous Galerkin method for multi-component chemically reacting flows. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02217-2. See cor-

rection [3769].

Gopalakrishnan:2023:DCV

[3513] J. Gopalakrishnan, L. Kogler, P. L. Lederer, and J. Schöberl. Divergenceconforming velocity and vorticity approximations for incompressible fluids obtained with minimal facet coupling. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02203-8.

Eriksson:2023:NCI

[3514] Gustav Eriksson. Non-conforming interface conditions for the secondorder wave equation. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02218-1.

Basava:2023:GRM

[3515] Seshadri R. Basava and Winnifried Wollner. Gradient robust mixed methods for nearly incompressible elasticity. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02227-0.

Tao:2023:SLL

[3516] Min Tao and Xiao-Ping Zhang. Study on L_{1} over L_{2} minimization for nonnegative signal recovery. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02225-2.

Chen:2023:FGB

[3517] Xi Chen, Alexander Kurganov, and Yongle Liu. Flux globalization based well-balanced central-upwind schemes for hydrodynamic equations with general free energy. Journal of Scientific Computing, 95 (3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-023-02221-6.

Weed:2023:QTE

[3518] Jared Weed, Lingyun Ding, Jingfang Huang, and Min Hyung Cho. Quadrature by two expansions for evaluating Helmholtz layer potentials. Journal of Scientific Computing, 95(3):??, June 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02222-5.

DeSterck:2023:EMR

[3519] H. De Sterck, R. D. Falgout, O. A. Krzysik, and J. B. Schroder. Efficient multigrid reduction-in-time for method-of-lines discretizations of linear advection. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02223-4.

Long:2023:CAF

[3520] Xiaonian Long and Qianqian Ding. Convergence analysis of the fully discrete projection method for inductionless magnetohydrodynamics system based on charge conservation. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02226-1.

Cai:2023:FDF

[3521] Dong-Ling Cai, Jingyan Hu, Yi-Bin Xiao, Ping Zeng, and Guanyu Zhou. A fully-discrete finite element scheme and projection–iteration algorithm for a dynamic contact problem with multi-contact zones and unilateral constraint. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02228-z.

Xiao:2023:SIR

[3522] Yao Xiao and Jan Glaubitz. Sequential image recovery using joint hierarchical Bayesian learning. Journal of Scientific Computing, 96(1): ??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02234-1.

Ouyang:2023:NGA

[3523] Wenqing Ouyang, Jiong Tao, Andre Milzarek, and Bailin Deng. Nonmonotone globalization for Anderson acceleration via adaptive regularization. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02231-4.

Pei:2023:SPP

[3524] Lifang Pei and Yifei Li. A structure-preserving parametric finite element method for area-conserved generalized curvature flow. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02236-z.

Hutridurga:2023:DGM

[3525] Harsha Hutridurga, Krishan Kumar, and Amiya K. Pani. Discontinuous Galerkin methods with generalized numerical fluxes for the Vlasov– viscous Burgers’ system. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02230-5.

Wang:2023:SPCa

[3526] Lan Wang, Linghua Kong, Meng Chen, Pengfei Zhu, and Huacheng Guo. Structure-preserving combined high-order compact schemes for multiple order spatial derivatives differential equations. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02219-0.

Giri:2023:PEA

[3527] Subhajit Giri and Shuvam Sen. Phase error analysis of implicit Runge–Kutta methods: New classes of minimal dissipation low dispersion high order schemes. Journal of Scientific Computing, 96(1): ??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02220-7.

Sun:2023:ENA

[3528] Jing Sun, Daxin Nie, and Weihua Deng. An efficient numerical algorithm for the model describing the competition between super- and subdiffusions driven by fractional Brownian sheet noise. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02240-3.

Sebastiano:2023:HOS

[3529] Boscarino Sebastiano. High-order semi-implicit schemes for evolutionary partial differential equations with higher order derivatives. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02235-0.

Trojak:2023:ERS

[3530] Will Trojak, Rob Watson, and Peter Vincent. An extended range of stable flux reconstruction schemes on quadrilaterals for various polynomial bases. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02216-3.

Fan:2023:NRD

[3531] Wei Fan, Xindi Hu, and Shengfeng Zhu. Numerical reconstruction of a discontinuous diffusive coefficient in variable-order time-fractional subdiffusion. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02237-y.

Bensaid:2023:DNN

[3532] Bilel Bensaid, Gaël Poëtte, and Rodolphe Turpault. Deterministic neural networks optimization from a continuous and energy point of view.

Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02215-4.

Liang:2023:LSM

[3533] Yuxiang Liang and Shun Zhang. Least-squares methods with nonconforming finite elements for general second-order elliptic equations. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02246-x.

Zhang:2023:NEI

[3534] Zhengfang Zhang, Xiangjing Gao, and Xiaoliang Cheng. Numerical estimation of the inverse eigenvalue problem for a weighted Helmholtz equation. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02242-1.

Calvo:2023:STO

[3535] Manuel Calvo, Lin Fu, Juan I. Montijano, and Luis Rández. Singly TASE operators for the numerical solution of stiff differential equations by explicit Runge–Kutta schemes. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02232-3.

Armentano:2023:REB

[3536] María Gabriela Armentano, Ariel L. Lombardi, and Cecilia Penessi. Robust estimates in balanced norms for singularly perturbed reaction diffusion equations using graded meshes. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02245-y.

Hu:2023:CAN

[3537] Shaotao Hu, Yuanheng Wang, and Qiao-Li Dong. Convergence analysis of a new Bregman extragradient method for solving fixed point problems and variational inequality problems in reflexive Banach spaces. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02243-0.

Khandelwal:2023:AQF

[3538] Rohit Khandelwal, Kamana Porwal, and Tanvi Wadhawan. Adaptive quadratic finite element method for the unilateral contact problem. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02206-5.

Wang:2023:CVE

[3539] Haimei Wang, Feng Wang, Jinru Chen, and Haifeng Ji. A conforming virtual element method based on unfitted meshes for the elliptic interface problem. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02229-y.

Wang:2023:MDG

[3540] Shixi Wang, Hai Bi, and Yidu Yang. The mixed discontinuous Galerkin method for transmission eigenvalues for anisotropic medium. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02244-z.

Huynh:2023:FAP

[3541] Phuoc-Toan Huynh, Yanzhao Cao, and Thi-Thao-Phuong Hoang. Fast and accuracy-preserving domain decomposition methods for reduced fracture models with nonconforming time grids. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-023-02251-0.

Hilb:2023:PDF

[3542] Stephan Hilb, Andreas Langer, and Martin Alkämper. A primal–dual finite element method for scalar and vectorial total variation minimization. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02209-2.

Alla:2023:HRB

[3543] Alessandro Alla, Hugo Oliveira, and Gabriele Santin. HJB-RBF based approach for the control of PDEs. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02208-3.

Hazra:2023:MGR

[3544] Arijit Hazra, Dinshaw S. Balsara, Praveen Chandrashekar, and Sudip K. Garain. Multidimensional generalized Riemann problem solver for Maxwell’s equations. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02238-x.

Dong:2023:SOC

[3545] Haixia Dong, Zhongshu Zhao, Shuwang Li, Wenjun Ying, and Jiwei Zhang. Second order convergence of a modified MAC scheme for Stokes interface problems. Journal of Scientific Computing, 96(1):??, July 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02239-w.

Bian:2023:ASH

[3546] Wei Bian and Fan Wu. Accelerated smoothing hard thresholding algorithms for ℓ_{0} regularized nonsmooth convex regression problem. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02249-8.

Rebholz:2023:EAA

[3547] Leo G. Rebholz and Mengying Xiao. The effect of Anderson acceleration on superlinear and sublinear convergence. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02262-x.

Yang:2023:HOC

[3548] Kai Yang. High order conservative schemes for the generalized Benjamin– Ono equation on the unbounded domain. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02255-w.

Kumar:2023:LLG

[3549] Lalit Kumar, Sivaji Ganesh Sista, and Konijeti Sreenadh. A linearized L1-Galerkin FEM for non-smooth solutions of Kirchhoff type quasilinear time-fractional integro–differential equation. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02241-2.

Wang:2023:SEE

[3550] Haijin Wang, Xiaobin Shi, and Qiang Zhang. Stability and error estimates of local discontinuous Galerkin methods with implicit– explicit backward difference formulas up to fifth order for convection– diffusion equation. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02264-9.

An:2023:NSA

[3551] Jing An, Ting Tan, and Zhimin Zhang. A novel spectral approximation and error estimation for transmission eigenvalues in spherical domains. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02261-y.

Zhang:2023:HOD

[3552] Guoyu Zhang, Chengming Huang, Anatoly A. Alikhanov, and Baoli Yin. A high-order discrete energy decay and maximum-principle preserving scheme for time fractional Allen–Cahn equation. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02263-w.

Bai:2023:GLH

[3553] Xiangyu Bai, Jiebao Sun, Jie Shen, Wenjuan Yao, and Zhichang Guo. A Ginzburg–Landau-H^{−1} model and its SAV algorithm for image inpainting. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02252-z.

Li:2023:FGM

[3554] Hongyi Li, Zhen Peng, Chengwei Pan, and Di Zhao. Fast gradient method for low-rank matrix estimation. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02266-7.

Boulbe:2023:TFB

[3555] Cédric Boulbe, Blaise Faugeras, Guillaume Gros, and Francesca Rapetti. Tokamak free-boundary plasma equilibrium computations in presence of non-linear materials. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02265-8.

Li:2023:EFG

[3556] Xiaolin Li. Element-free Galerkin analysis of Stokes problems using the reproducing kernel gradient smoothing integration. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02273-8.

Danis:2023:GFD

[3557] Mustafa Engin Danis and Jue Yan. A generalized framework for direct discontinuous Galerkin methods for nonlinear diffusion equations. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02257-8.

Wang:2023:DSN

[3558] Chengjing Wang and Peipei Tang. A dual semismooth Newton based augmented Lagrangian method for large-scale linearly constrained sparse group square-root lasso problems. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02271-w.

Gjesteland:2023:CCS

[3559] Anita Gjesteland and Magnus Svärd. Convergence of Chandrashekar’s second-derivative finite-volume approximation. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02256-9.

Yang:2023:CIS

[3560] Yijie Yang, Yutong Li, Chunlin Wu, and Yuping Duan. A convergent iterative support shrinking algorithm for non-Lipschitz multi-phase image labeling model. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02268-5.

Franchini:2023:CLS

[3561] Giorgia Franchini, Federica Porta, Valeria Ruggiero, and Ilaria Trombini. Correction to: A line search based proximal stochastic gradient algorithm with dynamical variance reduction. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02267-6.

Yu:2023:SOE

[3562] Fan Yu and Minghua Chen. Second-order error analysis for fractal mobile/immobile Allen–Cahn equation on graded meshes. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02276-5.

Chen:2023:PPE

[3563] Hao Chen and Alfio Borzì. Positivity preserving exponential integrators for differential Riccati equations. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02275-6.

Huynh:2023:DGR

[3564] H. T. Huynh. Discontinuous Galerkin and related methods for ODE. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02233-2.

Takhirov:2023:IAH

[3565] Aziz Takhirov, Aytekin Çıbık, Fatma G. Eroglu, and Songül Kaya. An improved Arrow–Hurwicz method for the steady–state Navier– Stokes equations. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02277-4.

Ciallella:2023:AHO

[3566] Mirco Ciallella, Davide Torlo, and Mario Ricchiuto. Arbitrary high order WENO finite volume scheme with flux globalization for moving equilibria preservation. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02280-9.

Fu:2023:HOT

[3567] Hongfei Fu, Bingyin Zhang, and Xiangcheng Zheng. A high-order twogrid difference method for nonlinear time-fractional biharmonic problems and its unconditional α-robust error estimates. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02282-7.

Chung:2023:MAR

[3568] Eric Chung, Wing Tat Leung, Sai-Mang Pun, and Zecheng Zhang. Multi-agent reinforcement learning aided sampling algorithms for a class of multiscale inverse problems. Journal of Scientific Computing, 96 (2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02279-2.

Kurganov:2023:NLD

[3569] Alexander Kurganov and Ruixiao Xin. New low-dissipation centralupwind schemes. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02281-8.

Xu:2023:DGT

[3570] Xiuxiu Xu and Qiumei Huang. Discontinuous Galerkin time stepping for semilinear parabolic problems with time constant delay. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02278-3.

Jiao:2023:CJL

[3571] Yuling Jiao, Dingwei Li, Min Liu, Xiliang Lu, and Yuanyuan Yang. Correction to: Just least squares: Binary compressive sampling with low generative intrinsic dimension. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02269-4.

Yang:2023:MBP

[3572] Xiu Yang, Li-Lian Wang, Huiyuan Li, and Changtao Sheng. Müntz ball polynomials and Müntz spectral-Galerkin methods for singular eigenvalue problems. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02254-x.

Pechstein:2023:UTN

[3573] Clemens Pechstein. A unified theory of non-overlapping Robin–Schwarz methods: Continuous and discrete, including cross points. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02248-9.

Liang:2023:RHT

[3574] Xiubo Liang, Guoqiang Wang, and Bo Yu. A reduced half thresholding algorithm. Journal of Scientific Computing, 96(2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02250-1.

Chu:2023:OOC

[3575] Tianyang Chu, Jilu Wang, Na Wang, and Zhimin Zhang. Optimalorder convergence of a two–step BDF method for the Navier–Stokes equations with H^{1} initial data. Journal of Scientific Computing, 96 (2):??, August 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02270-x.

Garrappa:2023:CAE

[3576] Roberto Garrappa and Andrea Giusti. A computational approach to exponential-type variable-order fractional differential equations. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02283-6.

Kent:2023:EAS

[3577] Benjamin M. Kent, Catherine E. Powell, David J. Silvester, and Małgorzata J. Zimoń. Efficient adaptive stochastic collocation strategies for advection–diffusion problems with uncertain inputs. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02247-w.

Patane:2023:SLT

[3578] Giuseppe Patané. Spectral Laplace transform of signals on arbitrary domains. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02274-7.

Meng:2023:DGM

[3579] Jian Meng. Discontinuous Galerkin method for the interior transmission eigenvalue problem in inverse scattering theory. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02290-7.

Zhao:2023:ATR

[3580] Shimin Zhao, Tao Yan, Kai Wang, and Yuanguo Zhu. Adaptive trustregion method on Riemannian manifold. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-023-02288-1.

Li:2023:ANA

[3581] Jun Li and Yaolin Jiang. Analysis of a new accelerated waveform relaxation method based on the time-parallel algorithm. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02285-4.

Li:2023:MCF

[3582] Xu Li and Hongxing Rui. A modified convective formulation in Navier– Stokes simulations. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02286-3.

Cheng:2023:STD

[3583] Lu Cheng and Kuan Xu. Solving time-dependent PDEs with the ultraspherical spectral method. Journal of Scientific Computing, 96(3): ??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02287-2.

Xie:2023:OEE

[3584] Ruiyi Xie, Boying Wu, and Wenjie Liu. Optimal error estimates for Chebyshev approximations of functions with endpoint singularities in fractional spaces. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02292-5.

Zhou:2023:SGN

[3585] Siyun Zhou, Xin Liu, and Liwei Xu. Stochastic Gauss–Newton algorithms for online PCA. Journal of Scientific Computing, 96(3): ??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02289-0.

Su:2023:RDF

[3586] Hongmin Su, Jinsheng Cai, Shucheng Pan, and Xiangyu Hu. Reformulated dissipation for the free-stream preserving of the conservative finite difference schemes on curvilinear grids. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02295-2.

Zhang:2023:PEA

[3587] Yongchao Zhang, Liquan Mei, and Gang Wang. A posteriori error analysis of the hybrid high-order method for the Stokes problem. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02291-6.

Chen:2023:USP

[3588] Wenbin Chen, Jianyu Jing, and Hao Wu. A uniquely solvable, positivitypreserving and unconditionally energy stable numerical scheme for the functionalized Cahn–Hilliard equation with logarithmic potential. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02296-1.

Wei:2023:CNR

[3589] Deliang Wei, Fang Li, and Shiyang Weng. Cauchy noise removal via convergent plug-and-play framework with outliers detection. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02303-5.

Wu:2023:CAP

[3590] Zhuohang Wu and Yu xin Ren. The compact and accuracy preserving limiter for high-order finite volume schemes solving compressible flows.

Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02298-z.

He:2023:DUE

[3591] Yuyu He and Hongtao Chen. Decoupled and unconditionally energy stable finite element schemes for electrohydrodynamic model with variable density. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02304-4.

Fornberg:2023:CFD

[3592] Bengt Fornberg and Cécile Piret. Computation of fractional derivatives of analytic functions. Journal of Scientific Computing, 96(3): ??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02293-4.

Xiao:2023:SED

[3593] Yao Xiao, Anne Gelb, and Guohui Song. Sequential edge detection using joint hierarchical Bayesian learning. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02297-0.

Grubisic:2023:DNR

[3594] Luka Grubisić, Ralf Hiptmair, and Diego Renner. Detecting near resonances in acoustic scattering. Journal of Scientific Computing, 96 (3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02284-5.

Xiao:2023:VEM

[3595] Wenqiang Xiao and Min Ling. Virtual element method for a historydependent variational-hemivariational inequality in contact problems. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02310-6.

Ahmat:2023:HHM

[3596] Muyassar Ahmat and Jianxian Qiu. Hybrid HWENO method for nonlinear degenerate parabolic equations. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02301-7.

Zhang:2023:PCP

[3597] Fan Zhang, Jian Cheng, and Tiegang Liu. A physical-constraint-preserving discontinuous Galerkin method for weakly compressible twophase flows. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02306-2.

Ye:2023:MMM

[3598] Qihao Ye and Xiaochuan Tian. Monotone meshfree methods for linear elliptic equations in non-divergence form via nonlocal relaxation. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02294-3.

Gordon:2023:CGA

[3599] Dan Gordon. Conjugate gradients acceleration of coordinate descent for linear systems. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02307-1.

Araya:2023:RPE

[3600] Rodolfo Araya and Franz Chouly. Residual a posteriori error estimation for frictional contact with Nitsche method. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02300-8.

Escalante:2023:MSM

[3601] C. Escalante, E. D. Fernández-Nieto, J. Garres-Díaz, and A. Mangeney. Multilayer shallow model for dry granular flows with a weakly non-hydrostatic pressure. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-023-02299-y.

Singh:2023:MCS

[3602] Abhinav Singh, Alejandra Foggia, Pietro Incardona, and Ivo F. Sbalzarini. A meshfree collocation scheme for surface differential operators on point clouds. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-023-02313-3.

Cao:2023:UAO

[3603] Waixiang Cao. Unified analysis of any order spectral volume methods for diffusion equations. Journal of Scientific Computing, 96(3): ??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02309-z.

Zhang:2023:GNC

[3604] Zhechen Zhang, Sanyang Liu, and Zhiping Lin. A generalized non-convex method for robust tensor completion. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02308-0.

Wang:2023:MTS

[3605] Zhong bao Wang, Zhen yin Lei, Xin Long, and Zhang you Chen. A modified Tseng splitting method with double inertial steps for solving monotone inclusion problems. Journal of Scientific Computing, 96 (3):??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02311-5.

Sun:2023:HBB

[3606] Zhong-Feng Sun, Jin-Chuan Zhou, and Yun-Bin Zhao. Heavy-ball-based optimal thresholding algorithms for sparse linear inverse problems. Journal of Scientific Computing, 96(3):??, September 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02315-1.

Xu:2023:SSO

[3607] Xiaojing Xu, Song Jiang, and Wenjun Sun. Spatial second-order positive and asymptotic preserving unified gas kinetic schemes for radiative transfer equations. Journal of Scientific Computing, 96(3): ??, September 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02305-3.

Chen:2023:NFE

[3608] Wei Chen, Jun Hu, and Min Zhang. Nonconforming finite element methods of order two and order three for the Stokes flow in three dimensions. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02317-z.

Guglielmi:2023:QSS

[3609] Nicola Guglielmi, Anton Savostianov, and Francesco Tudisco. Quantifying the structural stability of simplicial homology. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02314-2.

Wang:2023:PGS

[3610] Yunlong Wang, Chungen Shen, Lei-Hong Zhang, and Wei Hong Yang. Proximal gradient/semismooth Newton methods for projection onto a polyhedron via the duality-gap–active-set strategy. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02302-6.

Biswas:2023:MRR

[3611] Abhijit Biswas and David I. Ketcheson. Multiple-relaxation Runge Kutta methods for conservative dynamical systems. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-023-02312-4.

Lei:2023:FED

[3612] Wenyu Lei, George Turkiyyah, and Omar Knio. Finite element discretizations for variable-order fractional diffusion problems. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02318-y.

Hu:2023:FSG

[3613] Jun Hu, Hao Luo, and Zihang Zhang. A fast solver for generalized optimal transport problems based on dynamical system and algebraic multigrid. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02272-9.

Li:2023:SOS

[3614] Jingwei Li, Rihui Lan, Yongyong Cai, Lili Ju, and Xiaoqiang Wang. Second-order semi–Lagrangian exponential time differencing method with enhanced error estimate for the convective Allen–Cahn equation. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02316-0.

Huang:2023:MWA

[3615] Chieh-Sen Huang, Todd Arbogast, and Chenyu Tian. Multidimensional WENO–AO reconstructions using a simplified smoothness indicator and applications to conservation laws. Journal of Scientific Computing, 97 (1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02319-x.

Thomann:2023:TCD

[3616] Andrea Thomann and Michael Dumbser. Thermodynamically compatible discretization of a compressible two-fluid model with two entropy inequalities. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02321-3.

Diaz:2023:PTP

[3617] P. Díaz, E. Mainar, and B. Rubio. Polynomial total positivity and high relative accuracy through Schur polynomials. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02323-1.

Frasca-Caccia:2023:OPN

[3618] Gianluca Frasca-Caccia and Pranav Singh. Optimal parameters for numerical solvers of PDEs. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02324-0.

Lu:2023:NIT

[3619] Song Lu and Xianmin Xu. Numerical investigations on trace finite element methods for the Laplace–Beltrami eigenvalue problem. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02326-y.

Hillebrand:2023:ALN

[3620] Dorian Hillebrand, Simon-Christian Klein, and Philipp Öffner. Applications of limiters, neural networks and polynomial annihilation in higher-order FD/ FV schemes. Journal of Scientific Computing, 97 (1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02322-2.

Garcke:2023:UFE

[3621] Harald Garcke, Robert Nürnberg, and Quan Zhao. Unfitted finite element methods for axisymmetric two-phase flow. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02325-z.

Cen:2023:CTM

[3622] Dakang Cen, Caixia Ou, and Seakweng Vong. Corrected L-type method for multi-singularity problems arising from delay fractional equations. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02329-9.

Xia:2023:VDV

[3623] Yingzhi Xia, Qifeng Liao, and Jinglai Li. VI-DGP: A variational inference method with deep generative prior for solving high-dimensional inverse problems. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02328-w.

Li:2023:PTB

[3624] Tian-Yi Li, Fang Chen, Hai-Wei Sun, and Tao Sun. Preconditioning technique based on sine transformation for nonlocal Helmholtz equations with fractional Laplacian. Journal of Scientific Computing, 97 (1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02332-0.

Acosta-Soba:2023:UES

[3625] Daniel Acosta-Soba, Francisco Guillén-González, and J. Rafael Rodríguez-Galván. An unconditionally energy stable and positive upwind DG scheme for the Keller–Segel model. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02320-4.

Gao:2023:FEM

[3626] Huadong Gao and Wen Xie. A finite element method for the dynamical Ginzburg–Landau equations under Coulomb gauge. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02327-x.

Fang:2023:FDH

[3627] Yiying Fang and Ying Jiang. A fully discrete high-order fast multiscale Galerkin method for solving boundary integral equations in a domain with corners. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02334-y.

Jia:2023:FSB

[3628] Zhongxiao Jia and Kailiang Zhang. A FEAST SVDsolver based on Chebyshev–Jackson series for computing partial singular triplets of large matrices. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02342-y.

Awanou:2023:SBV

[3629] Gerard Awanou. The second boundary value problem for a discrete Monge–Ampère equation. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02340-0.

Liu:2023:SOA

[3630] Chun Liu, Cheng Wang, Steven M. Wise, Xingye Yue, and Shenggao Zhou. A second order accurate, positivity preserving numerical method for the Poisson–Nernst–Planck system and its convergence analysis. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02345-9.

Hu:2023:WPD

[3631] Jingyan Hu and Guanyu Zhou. The well-posedness and discontinuous Galerkin approximation for the non–Newtonian Stokes–Darcy– Forchheimer coupling system. Journal of Scientific Computing, 97 (1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02344-w.

Agarwal:2023:FAS

[3632] Dhwanit Agarwal, Michael O’Neil, and Manas Rachh. FMM-Accelerated solvers for the Laplace–Beltrami problem on complex surfaces in three dimensions. Journal of Scientific Computing, 97(1):??, October 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02336-w.

Chen:2023:VFF

[3633] Xinan Chen, Anh Phong Tran, Rena Elkin, Helene Benveniste, and Allen R. Tannenbaum. Visualizing fluid flows via regularized optimal mass transport with applications to neuroscience. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02337-9.

Kumar:2023:OPE

[3634] Raman Kumar and Bhupen Deka. Optimal a priori error estimates for elliptic interface problems: Weak Galerkin mixed finite element approximations. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02333-z.

Liu:2023:MHR

[3635] M. Liu, J. Dong, Z. Li, and D. F. Li. A modified high-resolution nonstaggered central scheme with adjustable numerical dissipation. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02349-5.

Wang:2023:SPCb

[3636] Lu Wang and Hui Liang. Superconvergence and postprocessing of collocation methods for fractional differential equations. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02339-7.

Bai:2023:CSI

[3637] Xiao Bai and Maojun Li. A conservative sharp-interface numerical method for two-dimensional compressible two-phase flows. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02338-8.

Liu:2023:MPS

[3638] Shiwei Liu and Li Yuan. A modified a posteriori subcell limiter for high order flux reconstruction scheme for one-dimensional detonation simulation. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02347-7.

Li:2023:PNC

[3639] Zhaoxiang Li, Feng Zhang, and Jianxin Zhou. Partial Newton-correction method for multiple fixed points of semi-linear differential operators by Legendre–Gauss–Lobatto pseudospectral method. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02341-z.

Rodgers:2023:IIN

[3640] Abram Rodgers and Daniele Venturi. Implicit integration of nonlinear evolution equations on tensor manifolds. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-023-02352-w.

Sujanani:2023:ASI

[3641] Arnesh Sujanani and Renato D. C. Monteiro. An adaptive superfast inexact proximal augmented Lagrangian method for smooth nonconvex composite optimization problems. Journal of Scientific Computing, 97 (2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02350-y.

Franco:2023:MIN

[3642] Nicola Rares Franco, Andrea Manzoni, and Paolo Zunino. Mesh-informed neural networks for operator learning in finite element spaces. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02331-1.

Frerichs-Mihov:2023:UDN

[3643] Derk Frerichs-Mihov, Linus Henning, and Volker John. Using deep neural networks for detecting spurious oscillations in discontinuous Galerkin solutions of convection-dominated convection–diffusion equations. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02335-x.

Bracco:2023:DDN

[3644] Cesare Bracco, Francesco Calabrò, and Carlotta Giannelli. Discontinuity detection by null rules for adaptive surface reconstruction. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02348-6.

Diethelm:2023:NAS

[3645] Kai Diethelm and Frank Uhlig. A new approach to shooting methods for terminal value problems of fractional differential equations. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02361-9.

Zhu:2023:UPD

[3646] Zhenyuan Zhu, Fan Chen, Junyu Zhang, and Zaiwen Wen. A unified primal–dual algorithm framework for inequality constrained problems. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02346-8.

Peng:2023:UAA

[3647] Zhichao Peng, Daniel Appelö, and Shuang Liu. Universal AMG accelerated embedded boundary method without small cell stiffness. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02353-9.

Jia:2023:CPQ

[3648] Zhigang Jia, Qianyu Wang, Hong-Kui Pang, and Meixiang Zhao. Computing partial quaternion eigenpairs with quaternion shift. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02355-7.

Chen:2023:ISA

[3649] Xinfu Chen, Zhengyang Lu, Jingtang Ma, and Jinye Shen. An implicit scheme for American put options. Journal of Scientific Computing, 97 (2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02356-6.

Huang:2023:REA

[3650] Chaobao Huang, Na An, Hu Chen, and Xijun Yu. α-robust error analysis of two nonuniform schemes for subdiffusion equations with variable-order derivatives. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-023-02357-5.

Buccini:2023:EIG

[3651] A. Buccini, P. Díaz de Alba, F. Pes, and L. Reichel. An efficient implementation of the Gauss–Newton method via generalized Krylov subspaces. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02360-w.

Ma:2023:ACI

[3652] Fugui Ma, Lijing Zhao, Weihua Deng, and Yejuan Wang. Analyses of the contour integral method for time fractional normal-subdiffusion transport equation. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02359-3.

Ye:2023:CMM

[3653] Qihao Ye and Xiaochuan Tian. Correction to: Monotone meshfree methods for linear elliptic equations in non-divergence form via nonlocal relaxation. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02343-x.

Konig:2023:TLB

[3654] Josie König and Melina A. Freitag. Time-limited balanced truncation for data assimilation problems. Journal of Scientific Computing, 97 (2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02358-4.

Cai:2023:SOC

[3655] Mingchao Cai, Huipeng Gu, Jingzhi Li, and Mo Mu. Some optimally convergent algorithms for decoupling the computation of Biot’s model. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02365-5.

Popa:2023:LRT

[3656] Jonathan Popa, Yifei Lou, and Susan E. Minkoff. Low-rank tensor data reconstruction and denoising via ADMM: Algorithm and convergence analysis. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02364-6.

Mossier:2023:EHA

[3657] Pascal Mossier, Daniel Appel, Andrea D. Beck, and Claus-Dieter Munz. An efficient hp-adaptive strategy for a level-set ghost-fluid method. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02363-7.

Wang:2023:EPE

[3658] Yangshuai Wang and Hao Wang. Efficient a posteriori error control of a concurrent multiscale method with sharp interface for crystalline defects. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02362-8.

Guclu:2023:BFF

[3659] Yaman Güçlü, Said Hadjout, and Martin Campos Pinto. A broken FEEC framework for electromagnetic problems on mapped multipatch domains. Journal of Scientific Computing, 97(2):??, November 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02351-x. See cor-

rection [3681].

Gao:2023:BSR

[3660] Yu Gao, Hongyu Liu, Xianchao Wang, and Kai Zhang. A Bayesian scheme for reconstructing obstacles in acoustic waveguides. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02368-2.

Amigo:2023:VEM

[3661] Danilo Amigo, Felipe Lepe, and Gonzalo Rivera. A virtual element method for the elasticity spectral problem allowing for small edges. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02372-6.

Li:2023:NNI

[3662] Fang Li and Xiao-Guang Lv. A nonconvex nonsmooth image prior based on the hyperbolic tangent function. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02366-4.

Han:2023:AEH

[3663] Yongbin Han and Yanren Hou. Analysis of an embedded-hybridized discontinuous Galerkin method for the time-dependent incompressible Navier–Stokes equations. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02369-1.

Olshanskii:2023:SAV

[3664] Maxim Olshanskii, Yerbol Palzhanov, and Annalisa Quaini. A scalar auxiliary variable unfitted FEM for the surface Cahn–Hilliard equation. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02370-8.

Jia:2023:IUE

[3665] Junqing Jia and Xiaoyun Jiang. Improved uniform error bounds of exponential wave integrator method for long-time dynamics of the space fractional Klein–Gordon equation with weak nonlinearity. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02376-2.

Vo:2023:HOT

[3666] Liet Vo. Higher order time discretization method for the stochastic Stokes equations with multiplicative noise. Journal of Scientific Computing, 97 (3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02375-3.

Cesmelioglu:2023:AEH

[3667] Aycil Cesmelioglu, Jeonghun J. Lee, and Sander Rhebergen. Analysis of an embedded-hybridizable discontinuous Galerkin method for Biot’s consolidation model. Journal of Scientific Computing, 97(3): ??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02373-5.

Carrasco:2023:NMF

[3668] Sergio Carrasco, Sergio Caucao, and Gabriel N. Gatica. New mixed finite element methods for the coupled convective Brinkman–Forchheimer and double-diffusion equations. Journal of Scientific Computing, 97 (3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02371-7.

Komori:2023:SRM

[3669] Yoshio Komori and Kevin Burrage. Split S-ROCK methods for highdimensional stochastic differential equations. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-023-02354-8.

Li:2023:AVE

[3670] Yanwei Li, Qiming Wang, and Zhaojie Zhou. Adaptive virtual element method for optimal control problem governed by Stokes equations. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02377-1.

Huang:2023:SPU

[3671] Qiong-Ao Huang, Wei Jiang, Jerry Zhijian Yang, and Cheng Yuan. A structure-preserving, upwind-SAV scheme for the degenerate Cahn– Hilliard equation with applications to simulating surface diffusion. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02380-6.

Berthon:2023:AVG

[3672] Christophe Berthon, Manuel J. Castro Díaz, Arnaud Duran, Tomás Morales de Luna, and Khaled Saleh. Artificial viscosity to get both robustness and discrete entropy inequalities. Journal of Scientific Computing, 97 (3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02385-1.

Jin:2023:CRP

[3673] Qinian Jin. On convergence rates of proximal alternating direction method of multipliers. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02383-3.

Zhang:2023:FNO

[3674] Zhenwei Zhang, Huibin Chang, and Yuping Duan. Fast non-overlapping domain decomposition methods for continuous multi-phase labeling problem. Journal of Scientific Computing, 97(3):??, December 2023. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-023-02382-
4.

Zeng:2023:ADD

[3675] Li Zeng, Xiaoliang Wan, and Tao Zhou. Adaptive deep density approximation for fractional Fokker–Planck equations. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02379-z.

Benatia:2023:BBH

[3676] Nawfel Benatia, Abdellah El Kacimi, Omar Laghrouche, and Ahmed Ratnani. Bernstein–Bézier H(curl)-conforming finite elements for timeharmonic electromagnetic scattering problems. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02381-5.

Vacek:2023:GLN

[3677] Lukás Vacek and Václav Kucera. Godunov-like numerical fluxes for conservation laws on networks. Journal of Scientific Computing, 97 (3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02386-0.

Li:2023:OEE

[3678] Dongfang Li, Xiaoxi Li, and Hai wei Sun. Optimal error estimates of SAV Crank–Nicolson finite element method for the coupled nonlinear Schrödinger equation. Journal of Scientific Computing, 97(3): ??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02384-2.

Bhoriya:2023:ESD

[3679] Deepak Bhoriya, Biswarup Biswas, Harish Kumar, and Praveen Chandrashekhar. Entropy stable discontinuous Galerkin schemes for two-fluid relativistic plasma flow equations. Journal of Scientific Computing, 97 (3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02387-z.

Arun:2023:APE

[3680] K. R. Arun, Rahuldev Ghorai, and Mainak Kar. An asymptotic preserving and energy stable scheme for the barotropic Euler system in the incompressible limit. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02389-x.

Guclu:2023:CBF

[3681] Yaman Güçlü, Said Hadjout, and Martin Campos Pinto. Correction: A broken FEEC framework for electromagnetic problems on mapped multipatch domains. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02399-9. See [3659].

Fu:2023:MEP

[3682] Yayun Fu, Xuelong Gu, Yushun Wang, and Wenjun Cai. Mass-, and energy preserving schemes with arbitrarily high order for the Klein– Gordon–Schrödinger equations. Journal of Scientific Computing, 97 (3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02388-y.

Ge:2023:STM

[3683] Ang Ge, Jinye Shen, and Seakweng Vong. Space-time methods based on isogeometric analysis for time-fractional Schrödinger equation. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02398-w.

Yadav:2023:ESS

[3684] Anshu Yadav, Deepak Bhoriya, Harish Kumar, and Praveen Chandrashekar. Entropy stable schemes for the shear shallow water model equations. Journal of Scientific Computing, 97(3):??, December 2023. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02374-4.

Liao:2024:FSI

[3685] Qichen Liao, Zihao Wang, Jing Chen, Bo Bai, Shi Jin, and Hao Wu. Fast Sinkhorn II: Collinear triangular matrix and linear time accurate computation of optimal transport. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02403-2.

Xiao:2024:SPG

[3686] Guiyun Xiao, Shuqin Zhang, and Zheng-Jian Bai. Scaled proximal gradient methods for sparse optimization problems. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02393-1.

Dai:2024:GPT

[3687] Ping-Fan Dai, Jianchao Bai, and Jicheng Li. A general preconditioner for tensor complementarity problems. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02391-3.

Liu:2024:SPS

[3688] Chen Liu, Yuan Gao, and Xiangxiong Zhang. Structure preserving schemes for Fokker–Planck equations of irreversible processes. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02378-0.

Jing:2024:MLF

[3689] Yang Jing, Jiaheng Chen, Lei Li, and Jianfeng Lu. A machine learning framework for geodesics under spherical Wasserstein–Fisher–Rao metric and its application for weighted sample generation. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02396-y.

Mallik:2024:HHO

[3690] Gouranga Mallik and Thirupathi Gudi. A hybrid high-order method for a class of strongly nonlinear elliptic boundary value problems. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02390-4.

Camargo:2024:AQP

[3691] Liliana Camargo, Bibiana López-Rodríguez, Mauricio Osorio, and Manuel Solano. An adaptive and quasi-periodic HDG method for Maxwell’s equations in heterogeneous media. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.com/

article/10.1007/s10915-023-02367-3.

Langer:2024:IRE

[3692] S. Langer and R. C. Swanson. Implementation, realization and an effective solver of two-equation turbulence models. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02394-0.

Qu:2024:PIC

[3693] Yunfei Qu, Hongjin He, and Deren Han. A partially inertial customized Douglas–Rachford splitting method for a class of structured optimization problems. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02397-x.

Miao:2024:NCE

[3694] Zhen Miao, Bin Wang, and Yao-Lin Jiang. Numerical conservations of energy, momentum and actions in the full discretisation for nonlinear wave equations. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02405-0.

Xu:2024:NDD

[3695] Fei Xu, Shuangshuang Chen, and Fusheng Luo. A novel domain decomposition method for eigenvalue problems. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02392-2.

Yang:2024:USC

[3696] Jinping Yang, Charles Wing Ho Green, Amiya K. Pani, and Yubin Yan. Unconditionally stable and convergent difference scheme for superdiffusion with extrapolation. Journal of Scientific Computing, 98 (1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02395-z.

Chevyrev:2024:FER

[3697] Ilya Chevyrev, Andris Gerasimovics, and Hendrik Weber. Feature engineering with regularity structures. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02401-4.

Serkh:2024:PFM

[3698] Kirill Serkh and James Bremer. Phase function methods for second order inhomogeneous linear ordinary differential equations. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02402-3.

Askham:2024:RWF

[3699] Travis Askham, Carlos Borges, Jeremy Hoskins, and Manas Rachh. Random walks in frequency and the reconstruction of obstacles with cavities from multi-frequency data. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02406-z.

Zheng:2024:PSA

[3700] Wen-Jie Zheng, Xi-Le Zhao, Yu-Bang Zheng, and Ting-Zhu Huang. Provable stochastic algorithm for large-scale fully-connected tensor network decomposition. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02404-1.

Zhang:2024:HFV

[3701] Xinyuan Zhang and Xiang Wang. The Hermite finite volume method with global conservation law. Journal of Scientific Computing, 98 (1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02407-y. See correction [3767].

Hu:2024:NAC

[3702] Xiuling Hu and Lulu Cheng. Numerical analysis of a convex-splitting BDF2 method with variable time-steps for the Cahn–Hilliard model. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02400-5.

Chen:2024:SNS

[3703] Yuan Chen and Xu Zhang. Solving Navier–Stokes equations with stationary and moving interfaces on unfitted meshes. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02414-z.

Moradi:2024:SSP

[3704] Afsaneh Moradi, Ali Abdi, and Gholamreza Hojjati. Strong stability preserving second derivative general linear methods based on Taylor series conditions for discontinuous Galerkin discretizations. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02415-y.

Tushar:2024:VEM

[3705] Jai Tushar, Ramesh Chandra Sau, and Anil Kumar. Virtual element method for control constrained Dirichlet boundary control problem governed by the diffusion problem. Journal of Scientific Computing, 98 (1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02410-3.

Liu:2024:DCU

[3706] Li Liu, Shengping Liu, Hui Xie, Fansheng Xiong, Tengchao Yu, Mengjuan Xiao, Lufeng Liu, and Heng Yong. Discontinuity computing using physics-informed neural networks. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02412-1.

Ahmadi-Asl:2024:RAT

[3707] Salman Ahmadi-Asl, Anh-Huy Phan, and Andrzej Cichocki. A randomized algorithm for tensor singular value decomposition using an arbitrary number of passes. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02411-2.

Wang:2024:AAS

[3708] Guanjie Wang, Smita Sahu, and Qifeng Liao. An adaptive ANOVA stochastic Galerkin method for partial differential equations with highdimensional random inputs. Journal of Scientific Computing, 98(1): ??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-023-02417-w.

Cances:2024:PEA

[3709] Eric Cancès, Gaspard Kemlin, and Antoine Levitt. A priori error analysis of linear and nonlinear periodic Schrödinger equations with analytic potentials. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02421-0.

Chu:2024:MAI

[3710] Hanyu Chu, Yongzhong Song, Haifeng Ji, and Ying Cai. Multigrid algorithm for immersed finite element discretizations of elliptic interface problems. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02416-x.

Yang:2024:DGM

[3711] Lei Yang, Yuan Liu, Yan Jiang, and Mengping Zhang. Discontinuous Galerkin methods for network patterning phase-field models. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02423-y.

Capatina:2024:RFR

[3712] Daniela Capatina, Aimene Gouasmi, and Cuiyu He. Robust flux reconstruction and a posteriori error analysis for an elliptic problem with discontinuous coefficients. Journal of Scientific Computing, 98 (1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02428-7.

Liao:2024:CDE

[3713] Wei-Hung Liao, Tsung-Ming Huang, Wen-Wei Lin, and Mei-Heng Yueh. Convergence of Dirichlet energy minimization for spherical conformal parameterizations. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02424-x.

Glaubitz:2024:ESG

[3714] Jan Glaubitz, Jan Nordström, and Philipp Öffner. Energy-stable global radial basis function methods on summation-by-parts form. Journal of Scientific Computing, 98(1):??, January 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02427-8.

Jing:2024:FVM

[3715] Feifei Jing, Weimin Han, Takahito Kashiwabara, and Wenjing Yan. On finite volume methods for a Navier–Stokes variational inequality. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02408-x.

Ravelo:2024:EEI

[3716] Fernando V. Ravelo, Pedro S. Peixoto, and Martin Schreiber. An explicit exponential integrator based on Faber polynomials and its application to seismic wave modeling. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02413-0.

Baumgart:2024:SSH

[3717] Alexandra Baumgart, Samuel W. Jones, Philipp V. F. Edelmann, and Joshua C. Dolence. A shock stabilization of the HLLC Riemann solver for the carbuncle instability. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02419-8.

Birke:2024:WBS

[3718] Claudius Birke, Walter Boscheri, and Christian Klingenberg. A wellbalanced semi-implicit IMEX finite volume scheme for ideal magnetohydrodynamics at all Mach numbers. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02422-z.

Chen:2024:RTP

[3719] Yaoyao Chen, Yunqing Huang, Nianyu Yi, and Peimeng Yin. Recovery type a posteriori error estimation of an adaptive finite element method for Cahn–Hilliard equation. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02418-9.

Jia:2024:LLI

[3720] Fan Jia, Tiange Wang, and Tieyong Zeng. Low-light image enhancement via dual reflectance estimation. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02431-y.

Huynh:2024:CAV

[3721] Ngoc Mai Monica Huynh. Convergence analysis for virtual element discretizations of the cardiac bidomain model. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02435-8.

Adak:2024:CCA

[3722] Dibyendu Adak, Gianmarco Manzini, Hashem M. Mourad, JeeYeon N. Plohr, and Lampros Svolos. A C^{1}-conforming arbitrary-order twodimensional virtual element method for the fourth-order phase-field equation. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-023-02409-w.

Kovacs:2024:UPE

[3723] Endre Kovács, János Majár, and Mahmoud Saleh. Unconditionally positive, explicit, fourth order method for the diffusion- and Nagumotype diffusion–reaction equations. Journal of Scientific Computing, 98 (2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02426-9.

Bozorgnia:2024:ILE

[3724] Farid Bozorgnia, Leon Bungert, and Daniel Tenbrinck. The infinity Laplacian eigenvalue problem: Reformulation and a numerical scheme. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02425-w.

Li:2024:NMF

[3725] Longfei Li, Hangjie Ji, and Qi Tang. Numerical methods for fourth-order PDEs on overlapping grids with application to Kirchhoff–Love plates. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-023-02430-z.

Dong:2024:NBI

[3726] Heping Dong and Peijun Li. A novel boundary integral formulation for the biharmonic wave scattering problem. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02429-6.

Xu:2024:HOA

[3727] Linfeng Xu, Shengrong Ding, and Kailiang Wu. High-order accurate entropy stable schemes for relativistic hydrodynamics with general Synge-type equation of state. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02440-x.

Li:2024:PEE

[3728] Lingfeng Li, Xue-Cheng Tai, Jiang Yang, and Quanhui Zhu. A priori error estimate of deep mixed residual method for elliptic PDEs. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02432-x.

Haidar:2024:FBP

[3729] Ali Haidar, Fabien Marche, and François Vilar. Free-boundary problems for wave–structure interactions in shallow-water: DG–ALE description and local subcell correction. Journal of Scientific Computing, 98 (2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02420-1.

Wang:2024:LSV

[3730] Gang Wang and Ying Wang. Least-squares virtual element method for Stokes problems on polygonal meshes. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02436-7.

Zhang:2024:PDM

[3731] Weinan Zhang, Pingping Zeng, Gongqiu Zhang, and Yue Kuen Kwok. Pricing discretely monitored Asian options under regime-switching and stochastic volatility models with jumps. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02438-5.

Ji:2024:DRM

[3732] Xia Ji, Yuling Jiao, Xiliang Lu, Pengcheng Song, and Fengru Wang. Deep Ritz method for elliptical multiple eigenvalue problems. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02443-8.

Xu:2024:EEI

[3733] Guixian Xu, Huihui Wang, and Qingping Zhou. Enhancing electrical impedance tomography reconstruction using learned half-quadratic splitting networks with Anderson acceleration. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02439-4.

Zhou:2024:CNS

[3734] Han Zhou and Wenyi Tian. Crank–Nicolson schemes for sub-diffusion equations with nonsingular and singular source terms in time. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02442-9.

Bir:2024:OEE

[3735] Bikram Bir, Deepjyoti Goswami, and Amiya K. Pani. Optimal error estimates of the penalty finite element method for the unsteady Navier– Stokes equations with nonsmooth initial data. Journal of Scientific Computing, 98(2):??, February 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-023-02445-6.

Barsukow:2024:IAF

[3736] Wasilij Barsukow and Raul Borsche. Implicit active flux methods for linear advection. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02433-w.

Kumar:2024:NSB

[3737] Sarvesh Kumar, David Mora, Ricardo Ruiz-Baier, and Nitesh Verma. Numerical solution of the Biot/elasticity interface problem using virtual element methods. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-023-02444-7.

An:2024:ABT

[3738] Chengtao An and Yangfeng Su. An aggregation-based two-grid method for multilevel block Toeplitz linear systems. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02434-9.

Li:2024:DPN

[3739] Haibo Li. Double precision is not necessary for LSQR for solving discrete linear ill-posed problems. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02447-4.

Seal:2024:DSW

[3740] Aniruddha Seal, Srinivasan Natesan, and Suayip Toprakseven. A dimensional-splitting weak Galerkin finite element method for 2D timefractional diffusion equation. Journal of Scientific Computing, 98(3): ??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-023-02448-3.

Neumayer:2024:TBI

[3741] Sebastian Neumayer and Antonia Topalovic. Template-based image reconstruction facing different topologies. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02446-5.

Benedusi:2024:MEC

[3742] Pietro Benedusi, Paola Ferrari, Marie E. Rognes, and Stefano Serra-Capizzano. Modeling excitable cells with the EMI equations: Spectral analysis and iterative solution strategy. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-023-02449-2.

Yu:2024:PPR

[3743] Boyang Yu, Yonghai Li, and Jiangguo Liu. A positivity-preserving and robust fast solver for time-fractional convection–diffusion problems. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JS-COEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://

link.springer.com/article/10.1007/s10915-024-02454-z.

Pulch:2024:HEU

[3744] Roland Pulch and Olivier Sète. The Helmholtz equation with uncertainties in the wavenumber. Journal of Scientific Computing, 98 (3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02450-3.

Ye:2024:SOT

[3745] Shuyu Ye, Qiang Ma, Qinglin Tang, Junzhi Cui, and Zhihui Li. Secondorder three-scale asymptotic analysis and algorithms for Steklov eigenvalue problems in composite domain with hierarchical cavities. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB.

ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02437-6.

Ersing:2024:ESD

[3746] Patrick Ersing and Andrew R. Winters. An entropy stable discontinuous Galerkin method for the two-layer shallow water equations on curvilinear meshes. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02451-2.

Eto:2024:RNM

[3747] Tokuhiro Eto. A rapid numerical method for the Mullins–Sekerka flow with application to contact angle problems. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02469-6.

Song:2024:SCM

[3748] Chenxiao Song and Reiichiro Kawai. Sampling and change of measure for Monte Carlo integration on simplices. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02461-0.

Yao:2024:RHP

[3749] Jianan Yao, Baoling Xie, and Jun Lai. A robust and high precision algorithm for elastic scattering problems from cornered domains. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02453-0.

Verfurth:2024:HOF

[3750] Barbara Verfürth. Higher-order finite element methods for the nonlinear Helmholtz equation. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02459-8.

Pei:2024:CSA

[3751] Lifang Pei, Yifan Wei, Chaofeng Zhang, and Jiwei Zhang. Convergence and superconvergence analysis of a nonconforming finite element variable-time-step BDF2 implicit scheme for linear reaction–diffusion equations. Journal of Scientific Computing, 98(3):??, March 2024. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-024-02456-
x.

Zhai:2024:NLE

[3752] Mengqing Zhai, Supei Zheng, Chengzhi Zhang, and Mangmang Jian. A new S–M limiter entropy stable scheme based on moving mesh method for ideal MHD and SWMHD equations. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02458-9.

Pang:2024:DBC

[3753] Qiyuan Pang and Haizhao Yang. A distributed block Chebyshev– Davidson algorithm for parallel spectral clustering. Journal of Scientific Computing, 98(3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02455-y.

Hong:2024:CEL

[3754] Xue Hong and Jing-Mei Qiu. A conservative Eulerian–Lagrangian Runge–Kutta discontinuous Galerkin method for linear hyperbolic system with large time stepping. Journal of Scientific Computing, 98 (3):??, March 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02452-1.

Kivva:2024:ESF

[3755] Sergii Kivva. Entropy stable flux correction for hydrostatic reconstruction scheme for shallow water flows. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02457-w.

Zhang:2024:TNN

[3756] Zezhong Zhang, Feng Bao, Lili Ju, and Guannan Zhang. Transferable neural networks for partial differential equations. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02463-y.

Archibald:2024:HLW

[3757] Rick Archibald and Ben Whitney. Haar-like wavelets on hierarchical trees. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02466-9.

Shu:2024:QMI

[3758] Yu-Chen Shu, Bing-Ze Lu, Kui-Yo Chen, and Matthew M. Lin. Quantifying measurement-induced disturbance to distinguish correlations as classical or quantum. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02471-y.

Gerhard:2024:PKS

[3759] Pierre Gerhard, Philippe Helluy, Victor Michel-Dansac, and Bruno Weber. Parallel kinetic schemes for conservation laws, with large time steps. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02468-7.

Cao:2024:ETD

[3760] Weichen Cao, Hengli Yang, and Wenbin Chen. An exponential time differencing Runge–Kutta method ETDRK32 for phase field models. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02474-9.

Chen:2024:MLP

[3761] Yannan Chen, Wen Li, and Jingya Chang. Multi-linear pseudo-PageRank for hypergraph partitioning. Journal of Scientific Computing, 99(1): ??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-024-02460-1.

Zhang:2024:OFS

[3762] Xinyue Zhang, Liang Pan, and Waixiang Cao. An oscillation-free spectral volume method for hyperbolic conservation laws. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02470-z.

Kim:2024:FDT

[3763] Jeongho Kim and Bora Moon. Finite difference time domain methods for the Dirac equation coupled with the Chern–Simons gauge field. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02473-w.

Ning:2024:LRI

[3764] Cui Ning, Xiaomin Kou, and Yaohong Wang. Low-regularity integrator for the Davey–Stewartson II system. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02467-8.

Xu:2024:MRQ

[3765] Liwei Xu, Xuejun Xu, and Shangyou Zhang. A modified rotated-Q_{1} finite element for the Stokes equations on quadrilateral and hexahedral meshes. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02477-6.

DeSterck:2024:AAK

[3766] Hans De Sterck, Yunhui He, and Oliver A. Krzysik. Anderson acceleration as a Krylov method with application to convergence analysis. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02464-x.

Zhang:2024:CHF

[3767] Xinyuan Zhang and Xiang Wang. Correction to: The Hermite finite volume method with global conservation law. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02475-8. See [3701].

Christiansen:2024:CNA

[3768] Thomas Christiansen, Katrin Grunert, Anders Nordli, and Susanne Solem. A convergent numerical algorithm for α-dissipative solutions of the Hunter–Saxton equation. Journal of Scientific Computing, 99 (1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-

7691 (electronic). URL https://link.springer.com/article/10.
1007/s10915-024-02479-4.

Du:2024:COF

[3769] Jie Du, Yong Liu, and Yang Yang. Correction to: An oscillation-free bound-preserving discontinuous Galerkin method for multi-component chemically reacting flows. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02480-x. See [3512].

Chudzik:2024:AFM

[3770] Erik Chudzik, Christiane Helzel, and Mária Lukácová-Medvid’ová. Active flux methods for hyperbolic systems using the method of bicharacteristics. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02462-z.

Luo:2024:RCM

[3771] Xin long Luo, Hang Xiao, and Sen Zhang. The regularization continuation method for optimization problems with nonlinear equality constraints. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02476-7.

Wu:2024:VOF

[3772] Yixuan Wu and Yanzhi Zhang. Variable-order fractional Laplacian and its accurate and efficient computations with meshfree methods. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02472-x.

Hou:2024:DGM

[3773] Shijin Hou and Yinhua Xia. Discontinuous Galerkin method based on the reduced space for the nonlinear convection–diffusion–reaction equation. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02486-5.

Lyu:2024:FSB

[3774] Xing-Long Lyu, Heng Tian, Tiexiang Li, and Wen-Wei Lin. Fast SVD-based linear elastic eigenvalue problem solver for band structures of 3D phononic crystals. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02483-8.

Hofmann:2024:PEE

[3775] S. Hofmann and A. Borzì. Pointwise error estimates of numerical solutions to linear quadratic optimal control problems. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02484-7.

Zappon:2024:ROM

[3776] Elena Zappon, Andrea Manzoni, Paola Gervasio, and Alfio Quarteroni. A reduced order model for domain decompositions with non-conforming interfaces. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02465-w.

daVeiga:2024:NVE

[3777] Lourenço Beirão da Veiga, Yi Liu, Lorenzo Mascotto, and Alessandro Russo. The nonconforming virtual element method with curved edges. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-023-02441-w.

Zhang:2024:ISS

[3778] Yu Zhang, Xiaoliang Song, Bo Yu, and Mingcai Ding. An iDCA with sieving strategy for PDE-constrained optimization problems with L^{1−2}- control cost. Journal of Scientific Computing, 99(1):??, April 2024. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-024-02489-
2.

Thiele:2024:NMO

[3779] J. P. Thiele and T. Wick. Numerical modeling and open-source implementation of variational partition-of-unity localizations of space–time dual-weighted residual estimators for parabolic problems. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02485-6.

Cui:2024:SOE

[3780] Ming Cui, Yiyi Niu, and Zhen Xu. A second-order exponential time differencing multi-step energy stable scheme for Swift–Hohenberg equation with quadratic–cubic nonlinear term. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02490-9.

Tan:2024:NPP

[3781] Yan Tan, Qiang Zhang, and Jun Zhu. A new positivity-preserving technique for high-order schemes to solve extreme problems of Euler equations on structured meshes. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02493-6.

Doehring:2024:MSO

[3782] Daniel Doehring, Gregor J. Gassner, and Manuel Torrilhon. Manystage optimal stabilized Runge–Kutta methods for hyperbolic partial differential equations. Journal of Scientific Computing, 99(1):??, April 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02478-5.

Walker:2024:DST

[3783] Shawn W. Walker. A descent scheme for thick elastic curves with self-contact and container constraints. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02487-4.

Liu:2024:PFI

[3784] Xin Liu, Nachuan Xiao, and Ya xiang Yuan. A penalty–free infeasible approach for a class of nonsmooth optimization problems over the Stiefel manifold. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02495-4.

Babbar:2024:APS

[3785] Arpit Babbar, Sudarshan Kumar Kenettinkara, and Praveen Chandrashekar. Admissibility preserving subcell limiter for Lax–Wendroff flux reconstruction. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-024-02482-

Wang:2024:SLL

[3786] Chao Wang, Ming Yan, and Junjie Yu. Sorted L_{1}/L_{2} minimization for sparse signal recovery. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02497-2.

Lei:2024:WHD

[3787] Wenyu Lei, Stefano Piani, Patricio Farrell, Nella Rotundo, and Luca Heltai. A weighted hybridizable discontinuous Galerkin method for drift–diffusion problems. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02481-w.

Riffaud:2024:LRS

[3788] Sébastien Riffaud, Miguel A. Fernández, and Damiano Lombardi. A lowrank solver for parameter estimation and uncertainty quantification in time-dependent systems of partial differential equations. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02488-3.

Storn:2024:SMR

[3789] Johannes Storn. Solving minimal residual methods in W^{−1,p′} with large exponents p. Journal of Scientific Computing, 99(2):??, May 2024. CO-DEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL

https://link.springer.com/article/10.1007/s10915-024-02494-
5.

Chen:2024:HOS

[3790] Wei Chen, Kailiang Wu, and Tao Xiong. High order structure-preserving finite difference WENO schemes for MHD equations with gravitation in all sonic Mach numbers. Journal of Scientific Computing, 99(2): ??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02492-7.

Yu:2024:NCS

[3791] Cui-E Yu, Xin Liu, and Yang Zhang. A new complex structurepreserving method for QSVD. Journal of Scientific Computing, 99(2): ??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02496-3.

Chen:2024:IMI

[3792] Junqing Chen and Zehao Long. An iterative method for the inverse eddy current problem with total variation regularization. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02501-9.

Lee:2024:UPR

[3793] Seulip Lee and Lin Mu. A uniform and pressure-robust enriched Galerkin method for the Brinkman equations. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02503-7.

Tao:2024:HOM

[3794] Zhanjing Tao, Jinming Zhang, Jun Zhu, and Jianxian Qiu. Highorder multi-resolution central Hermite WENO schemes for hyperbolic conservation laws. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02499-0.

Guidotti:2024:FMC

[3795] Nicolas L. Guidotti, Juan A. Acebrón, and José Monteiro. A fast Monte Carlo algorithm for evaluating matrix functions with application in complex networks. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02500-w.

Cherepanov:2024:DBV

[3796] Vladislav Cherepanov, Jian-Guo Liu, and Zhongmin Qian. On the dynamics of the boundary vorticity for incompressible viscous flows. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02498-1.

Zheng:2024:FHO

[3797] Zi-Yun Zheng and Yuan-Ming Wang. Fast high-order compact finite difference methods based on the averaged L1 formula for a time-fractional mobile–immobile diffusion problem. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02505-5.

Lou:2024:QDF

[3798] Yuzhi Lou and Hongxing Rui. A quadratic discontinuous finite volume element scheme for Stokes problems. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02506-4.

Fan:2024:OTL

[3799] Mengxiao Fan and Jicheng Li. The order-p tensor linear complementarity problem for images deblurring. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02502-8.

Liao:2024:ACE

[3800] Hong lin Liao, Xiaohan Zhu, and Hong Sun. Asymptotically compatible energy and dissipation law of the nonuniform L2-1_{σ} scheme for time fractional Allen–Cahn model. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02515-3.

Zhu:2024:HEN

[3801] Hongyi Zhu and Chuanju Xu. A highly efficient numerical method for the time-fractional diffusion equation on unbounded domains. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02508-2.

Brugnano:2024:SAS

[3802] Luigi Brugnano, Kevin Burrage, Pamela Burrage, and Felice Iavernaro. A spectrally accurate step-by-step method for the numerical solution of fractional differential equations. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02517-1.

Yao:2024:GBI

[3803] Wenhui Yao and Chunxiong Zheng. Generalized boundary integral equation method for boundary value problems of two-D isotropic lattice Laplacian. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02507-3.

Hu:2024:FEG

[3804] Jun Hu, Yizhou Liang, and Ting Lin. Finite element grad grad complexes and elasticity complexes on cuboid meshes. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02512-6.

Luo:2024:UNM

[3805] Yuhua Luo and Ting Wei. Uniqueness and numerical method for determining a spatial source term in a time–fractional diffusion wave equation. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02523-3.

Peng:2024:TSO

[3806] Xiangyi Peng, Wenlin Qiu, Ahmed S. Hendy, and Mahmoud A. Zaky. Temporal second-order fast finite difference/compact difference schemes for time-fractional generalized Burgers’ equations. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.springer.

com/article/10.1007/s10915-024-02514-4.

Bai:2024:NIA

[3807] Jianchao Bai, Linyuan Jia, and Zheng Peng. A new insight on augmented Lagrangian method with applications in machine learning. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02518-0.

Wu:2024:EIG

[3808] Tingting Wu, Yue Min, Chaoyan Huang, Zhi Li, Zhongming Wu, and Tieyong Zeng. An efficient inexact Gauss–Seidel-based algorithm for image restoration with mixed noise. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02510-8.

Zeng:2024:AED

[3809] Ping Zeng and Guanyu Zhou. Analysis of an energy-dissipating finite volume scheme on admissible mesh for the aggregation–diffusion equations. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https:

//link.springer.com/article/10.1007/s10915-024-02522-4.

Hao:2024:NMS

[3810] Wenrui Hao, Jonathan D. Hauenstein, Margaret H. Regan, and Tingting Tang. A numerical method for solving elliptic equations on real closed algebraic curves and surfaces. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print),

1573-7691 (electronic). URL https://link.springer.com/article/
10.1007/s10915-024-02516-2.

Li:2024:ECO

[3811] Xinyi Li, Jing Yuan, Xue-Cheng Tai, and Sanyang Liu. Efficient convex optimization for non-convex non-smooth image restoration. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691 (electronic). URL https://link.

springer.com/article/10.1007/s10915-024-02504-6.

Allendes:2024:FED

[3812] Alejandro Allendes, Gilberto Campaña, and Enrique Otárola. Finite element discretizations of a convective Brinkman–Forchheimer model under singular forcing. Journal of Scientific Computing, 99(2):??, May 2024. CODEN JSCOEB. ISSN 0885-7474 (print), 1573-7691

(electronic). URL https://link.springer.com/article/10.1007/
s10915-024-02513-5.